A Complete Bibliography of the Bulletin in Applied Statistics (BIAS) and Journal of Applied Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)

WWW URL: https://www.math.utah.edu/ beebe/

29 August 2024 Version 2.46

Title word cross-reference

#1 [2505].

(1, 1) [2807]. (5, 2) [1320]. (k, l) [3293]. (n; c_{1}, c_{2}) [776]. (p, q) [1196]. (Q, r) [3542]. (r, T) [816]. 0.005 [1567]. 0.15 [1567]. 12 [2808]. 1 < α ≤ 2 [2579]. 2 [4128]. $24.36 [4599]. 2^{k} [1887, 1532]. 2^{k−r} [1743]. 2^{n} [708]. 2 × 2 [1092, 2827, 1082, 1352, 1865, 3781]. 2 × c [3844]. 3 [3295, 3850]. 4A(c_{1}, c_{2})r [826]. 85 [3093]. [0, ∞)^{d} [4294]. ^{(}R) [2532]. _{2} [1536]. 2.5 [4332]. A [1156, 3128, 1581, 1626]. A_{m,n} [774]. AB − BA [976]. α [2112]. AR(1) [3370]. B [830]. X̄ [4100, 4095]. BINMA(1) [3721]. C [3980, 1780, 1721]. C(α) [3817]. C_{pk} [2181, 941, 1947]. C_{pm} [1529]. C_{pyk} [3804]. ChSP [826]. D [1156, 2813, 1761, 943,

1770, 3249, 4680, 1781, 4674, 2016]. D^{2} ; [1747]. D_{s} [2176]. δ [636, 846, 1062]. ; δ = P (X < Y) [2350]. ϵ [3080, 1730]. F ; [973, 1608, 214, 2366, 1567]. G ; [4356, 2904, 4350]. gh [2488]. H [3128]. I ; [1738]. INAR(1) [3708]. K ; [3705, 1833, 1857, 3206, 3190, 4651, 2465, ; 2001, 2901, 3803, 3323, 2266, 4348, 1853, ; 2196, 3890, 3934, 1963, 3289, 2373]. k ≥ 4 ; [715]. L [3226, 3536, 4093]. L^{1} [3372]. L_{1} ; [2698, 3246]. L_{2} [2698]. l_{2,p} [4457]. λ_{L} [1486]. ; M [2665, 3063, 1490, 2811, 3370, 774]. ; VAR(p) [1416]. n [468, 774]. n_{1} [1963]. n_{2} ; [1963]. np [1536]. ν [4338]. P ; [372, 4074, 1567, 3086, 1510, 1769, 3717, 2293, ; 918, 3954, 4323, 4020, 4008, 2027, 2693, 3940]. ; P (λ_{1} < λ_{2}) [2485]. ϕ [37, 2846]. πps [2606]. ; Q [3647, 2512, 986, 3907]. QQ [1708]. R ; [1426, 3226, 2848, 571]. R^{2}

[1914, 3413, 3711, 1800, 1984]. ρ [4029]. S [3406, 2309, 4100, 4095, 2848]. S^{2} [3134]. t [2665, 4144, 2921, 2079, 4081, 1509, 2211, 1104, 4067, 3154, 3276, 4216, 882, 2535, 2418, 3968, 1044, 1688, 4083, 4342]. T^{2} [1936, 4543, 1076, 2283, 3845]. T_{n} [774]. τ [2518]. × [2307, 4133]. U [3162, 882, 1238, 1239, 1237]. U_{n}^{2} [710]. W_{m,n} [774]. X [1186, 2134, 2291, 1494, 1320, 1426]. X^{−} [1566, 2142]. X^{2} [1752, 214]. Y [1494, 1320]. z [2579].

-ary [468]. -based [4083]. -chart [1721]. -contaminated [1730]. -corrected [636, 846, 1062]. -distance [3372]. -distribution [2665, 2921]. -distributions [2079]. -divergence [2846]. -error [4144, 3276]. -estimates [1490]. -estimation [3370]. -estimator [3063, 3406]. -estimator-based [2811]. -estimators [2665]. -EWMA [2134]. -exponential [3907]. -factor [4674]. -family [4350]. -inflated [3293]. -insensitive [3080]. -Invariant [571]. -means [2373, 3705, 1833, 2465]. -median [3940]. -medoids [3890]. -mode [3647]. -moments [3536, 4093]. -negative [2904]. -norm [4457]. -Optimal [1626, 2176, 1156, 1581, 2813, 1761, 943, 1770, 3249]. -parameter [3295]. -Plot [1708]. -plus [3093]. -run [2808]. -sample [1780, 2901]. -skip [4100, 4095]. -stable [2112]. -statistics [3162]. -support [4338]. -test [4081]. -tests [3968, 2366]. -th [4348]. -trace [4680]. -value [918, 3954, 4008, 372]. -values [3086, 3717, 2293, 2693, 4074]. -Variates [1608].

0 [2817, 4669]. 0-47-1-73578-7 [4669].

1 [2113, 1083, 1180, 2822, 3112, 1502, 3247, 2821, 675, 861, 935]. 1-4822-3810-1 [3572]. 105D [252, 621]. 19 [4538, 4536, 4549, 4687, 4650, 4433, 4544, 4553, 4545, 4543, 4534, 4503, 4565, 4683, 4274, 4654, 4552, 4510, 4631, 4445, 4589, 4535, 4618, 4528, 4540, 4533, 4537, 4547, 4491, 4502, 4541, 4539, 4548, 4551, 4500, 4512]. 1930s [159]. 1976 [92]. 1978 [121]. 19’s [4546].

2 [3112, 1610, 4542, 2818, 4599]. 2000 [1200]. 2009 [3166, 2768]. 2009/10 [3689]. 2010 [2230, 2664]. 2010-2014 [4588]. 2011 [2524]. 2014/15 [3689]. 2018 [3993]. 27680 [2652]. 28S [504]. 2nd [3662]. 2nd-power [3662].

3 [1687, 4599, 1099, 2748, 2651].

3-319-69416-2 [4599]. 3-points-for-a-win [3883]. 3CoSE [4448].

4 [748, 2816, 1687, 683]. 4A

[544, 556, 645, 709].

5 [2815, 2745, 2824, 2823, 2653]. 5-year-old [2304]. 500 [4178]. 5D [3621].

6 [1244, 2820]. 6-min [1913]. 6/49 [1668].

7 [4669]. 7-day [4528].

8th [92].

9 [2819, 2747]. 978 [2820, 2822, 2816, 2815, 2817, 2745, 2824, 2819, 2823, 2821, 2818, 2747, 2653, 2748, 2651, 2652]. 978-0-470-46363-5 [2823]. 978-0-470-53331-4 [2816]. 978-0-470-97429-2 [2818]. 978-1-107-60655-5 [2824]. 978-1-118-21798-6 [2820]. 978-1-1183-3257-3 [2748]. 978-1-119-99320-9 [2819]. 978-1-4398-0624-1 [2822]. 978-1-4398-0910-5 [2815]. 978-1-4398-27680 [2652]. 978-1-4398-4068-9 [2747]. 978-1-4398-4657-5 [2653].

978-1-4398-8414-0 [2817]. 978-1-4398-9254-1 [2821]. 978-1-60805-373-5 [2745]. 978-3-319-69416-0 [4599]. 978-981-4291-32-3 [2651].

A-bomb [1997]. AB/BA [3056]. aberration [3773]. ability [280, 3820]. abolition [611]. Aboriginal [1300]. abortion [2279]. abrupt [2397]. absence [644, 3192]. absolute [3007, 1589, 3053, 1308]. absorption [3766, 1285]. Abstracts [426, 425]. abundance [3487, 2583]. Abundant [1673]. Abundant/Rare [1673]. Acacia [1615]. Academic [3162]. Accelerated [3249, 1422, 4039, 3923, 4354, 2314, 1828, 1864, 2116, 1172, 4651, 4271, 881, 4616, 2137]. acceptable [506, 565, 637, 1041, 567, 718, 646, 991]. Acceptance [3907, 1458, 1121, 1580, 1906, 1965, 1977, 3784, 3274, 4052, 596, 2605, 1287, 1314, 910, 1086, 1828, 3371, 1957, 3830, 3851, 4094, 963]. accident [4409]. accidents [2548, 4582, 4258, 4641]. accommodate [2135]. Accompanying [1403]. according [1517]. account [4387, 2926, 4327]. Accounting [962, 1285, 4378, 3659, 4307, 3262]. Accumulate [103]. accumulated [1048]. Accumulating [480]. Accuracy [1380, 2726, 2761, 3039, 4167, 1500, 2247, 2316, 2740, 4508, 1976, 4679, 4209, 3579, 2030, 3148, 3192]. Accurate [4190]. Accusative [1382]. ACE [1475]. acetate [3833]. achievement [353]. achievements [3425]. acid [2624]. acids [3937, 3268]. Acknowledgement [2359, 1611]. Acknowledgements [2825]. Acquisition [4743, 4411, 3770, 519, 2819]. acquisitions [4415]. across [3471, 3468, 3167, 1137, 2864, 4552, 3807, 2930]. across-regime [3471]. action [669]. activation [3567]. Active [4568, 1736, 2874, 3004, 3163, 2623]. activities [3880]. activity [2995, 2876, 4608, 4559, 4531, 863, 2702]. actual [3158, 862]. actuarial [2048]. acute [1478, 2449]. Ad [1664]. AdaBoost [2127]. adaptation [2655, 1329]. Adapting [317, 1118]. Adaptive [988, 1780, 4565, 2295, 3300, 3568, 4281, 2598, 1390, 2940, 2311, 3074, 4359, 2420, 4388, 4375, 1841, 4487, 4365, 4485, 4001, 3259, 2566, 2314, 2221, 4612, 1823, 1827, 4672, 3754, 672, 928, 995, 2841, 3589, 4294, 4637, 4088, 2696, 2812, 3198, 1881]. added [3450, 1347, 3820]. addition [4109]. Additional [1626, 1483, 1854, 4509, 3405]. Additive [2663, 3998, 3755, 3987, 4515, 2279, 3778, 2151, 2916, 2379, 3967, 3209, 1188, 3335, 1484, 3605]. additivity [599]. address [751, 3201]. addresses [377]. Addressing [4441, 3557, 1805]. Adenomas [1389]. adequacy [1279, 887, 2253]. adequate [3358]. Adhaya [2817]. adjust [1855, 4619]. Adjusted [1699, 359, 1474, 3836, 4509, 4620, 1469, 1914, 4335, 3368, 3527, 2207, 334, 3969, 2551, 2373]. Adjusted-loss-function [1474]. Adjusting [4327, 1297, 1812, 1262, 4252, 48, 4207, 2372]. Adjustment [1409, 1624, 1696, 2091, 1290, 4339, 1145, 3819, 4355, 4277, 2315]. admission [1520, 2165]. admissions [1472, 1735]. adopting [4411]. adoption [4090]. adult [743, 1208, 527, 4559]. adults [1236, 2795, 3602]. Advanced [504, 4747, 2948]. Advances [4726, 4743, 2819, 4599, 2745]. advancing [1232]. advantage [1985, 566]. advantages [1938]. adverse [3660, 1126, 3045, 880]. advice [2613]. Advising [2323]. Aegean [3412]. Aerobics [4396]. aesthetics [4051]. affect [2262]. Affected [196, 2851, 550]. Affecting [155, 104, 3311]. affine [2710]. Afghanistan [3173]. Africa [2844, 235, 3166, 281, 4237]. African [2581, 3383, 4331, 3579]. after [3345, 4458, 1709, 2294, 3169, 3986, 2714, 1990, 1200, 3902]. against [4681, 1085, 1117, 3971, 1293, 1122, 540, 2700]. Age [118, 2258, 734, 2212, 4545, 1243, 3795, 2945, 2604, 1053, 1205, 2390, 1798, 1227, 1924, 884, 1556, 4513, 3580]. age-specific [2212, 4545, 1243, 1798, 1227]. age-specificity [1205]. age-stratified [2945]. age-treatment [3795]. agencie [1797]. agency [3517]. agents [4522]. Aggregate [915, 1060, 3281, 1839]. Aggregated [2964, 4636, 4312, 4014]. aggregation [3169, 2482]. aging [690, 4123]. Agree [520]. agreeable [520]. Agreement [3262, 2025, 1796, 3297, 830, 2662, 1749, 2450, 2340, 4129]. Agresti [2823]. Agricultural [209, 3869]. agriculture [3126, 4050]. Aid [188]. aided [3085]. aids [3050, 1050, 1467, 3229, 3326, 3547]. Air [1700, 2472, 584, 3441, 3420, 2360, 4489, 3961, 2601, 2600, 4219, 3550, 4521]. air-trapping [3441]. Airborne [1398]. Aiyi [4599]. Akaike [900]. Al [1970]. Al-Sultan [1970]. al. [3254, 3255, 3200]. Aladag [2745]. Alan [2823]. alarms [1731]. Alaska [2349]. Alaskan [4316]. Alberink [3254]. Alfred [1711]. algebra [1242]. algebraic [659]. Algorithm [1426, 102, 3213, 3988, 1994, 711, 4143, 3525, 2161, 2562, 2561, 2130, 2622, 2300, 2494, 2688, 3064, 3115, 3407, 2641, 2740, 1296, 4479, 682, 2960, 3546, 3043, 4067, 2826, 3017, 4254, 3100, 2847, 4212, 3812, 4410, 3906, 447, 4227, 4091, 1274, 1067, 2062, 4448, 610, 3565, 3001, 4457, 2413, 1475]. algorithm-based [2826]. algorithms [3950, 3035, 4530, 774, 3048, 931, 3362, 4660, 943, 4474, 318, 1148, 3976, 608]. alias [1561]. aligned [961]. Alignment [1639, 2643, 3036]. Alison [165]. all-zero [1925]. allele [1876]. alleles [1295]. allergy [2233]. allocation [806, 2689, 2731, 4565, 2889, 1230, 4112, 2696, 3377]. allowable [819]. allowing [1290, 1586]. allows [2220]. alloy [2863]. almost [3400, 2138]. along [591, 4522]. alongside [3581]. alpha [4139, 4140, 3307]. Alps [1502, 1610]. alternately [3067]. alternating [4010]. Alternative [2292, 278, 1700, 4610, 1361, 3316, 1677, 1652, 4639, 3081, 3512, 3167, 380, 2713, 3662, 914, 1460, 572, 4428, 4466, 3323, 3373, 3593, 3922, 1506, 345, 930, 1982, 4162, 3099, 3261, 226]. alternatives [1936, 988, 2956, 1355, 1076, 2121, 4208, 4430, 1759, 2700]. Alzheimer [1994, 2268, 4504, 2095]. ambient [287]. amended [2709]. Amendments [156]. America [758]. American [2035, 769, 4555, 4619, 319]. amino [3937]. ammonium [604]. amoebae [658]. among [3942, 2243, 735, 3664, 3327, 2178, 1541, 2390, 1735, 739, 742, 2450, 903, 3669, 2173, 3602]. amongst [3571]. amount [1770]. amplitude [3749, 4054]. amyotrophic [4241]. Anagramatic [151]. analog [4068]. analogs [4220]. analogue [4356]. Analyse [1448, 3287]. analysed [1831]. analyses [3517, 2233, 769, 1791, 4187, 1174, 1157, 4255, 1716, 2448]. Analysing [103, 3069, 1844, 2264, 1281, 1212, 2993, 2153, 2656]. Analysis [2304, 4739, 4710, 2709, 2921, 4296, 2002, 3350, 1643, 4712, 117, 210, 3274, 4655, 4697, 1411, 1209, 2070, 1641, 524, 4694, 2298, 4731, 171, 1711, 2833, 2870, 1664, 296, 4749, 1033, 1875, 3020, 3611, 128, 119, 1835, 2319, 3508, 168, 2066, 2236, 2526, 480, 1925, 3474, 280, 4706, 2843, 1617, 3893, 1326, 2325, 1119, 1047, 1766, 448, 1707, 3253, 1908, 4556, 2191, 669, 141, 146, 196, 1432, 1697, 3620, 1632, 904, 2013, 2065, 4738, 2167, 148, 1422, 2470, 4746, 1438, 3897, 321, 4453, 1420, 4703, 2573, 531, 3383, 3689, 757, 3769, 4226, 1249, 2096, 3950, 3256, 499]. analysis [3242, 3079, 1510, 1330, 1748, 1793, 1769, 4163, 4664, 1456, 62, 91, 897, 3022, 2034, 4648, 1291, 3439, 748, 509, 3336, 4171, 1333, 4257, 3834, 2498, 4191, 23, 2676, 3441, 2384, 2239, 485, 3658, 2616, 2423, 3908, 451, 3199, 2971, 2424, 1953, 343, 1796, 2545, 3551, 847, 4286, 1979, 3328, 3828, 2484, 1531, 1811, 728, 4287, 3564, 3609, 3250, 4575, 3706, 3526, 2649, 1060, 2667, 2941, 2807, 2900, 3217, 3224, 2358, 3717, 3182, 391, 2789, 3841, 793, 2476, 1910, 569, 1201, 1551, 1596, 3211, 2942, 1930, 2487, 3113, 1718, 3892, 2926, 3402, 2851, 2675, 2011, 2284, 3174, 1311, 611]. analysis [4347, 3446, 2815, 494, 4424, 860, 4075, 4509, 2560, 416, 862, 3844, 4532, 1784, 3160, 1877, 2754, 2827, 3818, 2294, 4370, 2538, 1578, 2489, 83, 245, 1900, 2532, 3150, 4132, 2285, 915, 3637, 1370, 3023, 4148, 2627, 2981, 3473, 2458, 1470, 2723, 2690, 332, 616, 3419, 3871, 1312, 2247, 3390, 4241, 378, 2360, 4030, 3986, 377, 2714, 58, 1920, 1893, 4102, 303, 507, 3854, 2457, 21, 248, 559, 604, 2566, 4396, 407, 3747, 4239, 814, 3327, 3457, 849, 694, 1052, 1518, 870, 959, 2844, 1153, 1165, 2226, 2364, 2740, 3280, 851, 3109, 873, 3624, 2010, 2875, 3092, 3695, 3775]. analysis [654, 3981, 2445, 1342, 951, 2644, 2375, 2643, 3975, 2282, 2767, 791, 1054, 1129, 2214, 4517, 4493, 2105, 3392, 3952, 3927, 3012, 3360, 3840, 4327, 4504, 4456, 2087, 2379, 1988, 3151, 3339, 3326, 3145, 670, 3367, 3210, 1895, 1503, 2850, 2962, 2405, 2671, 4221, 370, 453, 415, 570, 1022, 1974, 2455, 3749, 4558, 3425, 1106, 2943, 1762, 2592, 3506, 3175, 1940, 4293, 2727, 3827, 3110, 3424, 4607, 419, 3509, 1971, 3522, 4524, 1046, 1978, 3544, 1202, 2949, 2708, 2174, 3097, 299, 2862, 1528, 875, 2750, 2610, 3119, 1132, 4025, 4271, 1233, 4127, 355, 2136, 3021, 1343, 3618, 2041, 2203, 2496, 3321, 2959, 2260]. analysis [297, 318, 4677, 3363, 4120, 738, 3120, 2751, 3599, 736, 2762, 423, 3850, 3024, 3163, 696, 930, 3870, 488, 3251, 265, 3269, 3396, 1798, 2601, 993, 4482, 1821, 1939, 2947, 2478, 3394, 3136, 2218, 2028, 2929, 2793, 854, 2154, 4217, 3308, 3615, 2306, 2716, 768, 1767, 3291, 3476, 3918, 2038, 2456, 2749, 837, 1005, 3299, 2143, 558, 3614, 907, 2204, 3099, 1943, 356, 1730, 110, 2031, 4330, 2624, 3642, 3791, 3670, 2201, 3761, 863, 3040, 4547, 2232, 2795, 550, 1110, 4491, 1918, 2071, 3296, 517, 3070, 3919, 1839, 519, 543, 903, 3553, 4106, 4213, 2229, 2540, 3811, 282, 4386, 965]. analysis [3714, 1051, 4685, 2569, 2095, 3516, 2932, 3713, 3018, 3940, 4680, 4689, 3013, 4401, 966, 2863, 3580, 3584, 4070, 4406, 2168, 4561, 2118, 2938, 2385, 2415, 4342, 4425, 4500, 2975, 4459, 3432, 3602, 3442, 2884, 4044, 2376, 2502, 2261, 2702, 3052, 740, 121, 4101, 2399, 510, 1880, 2822, 2036, 4669, 2823, 2225]. analysis-based [3446, 4561]. analysis-of-means-type [965]. analytic [1517, 781, 4249]. Analytical [717, 221, 2958, 4233, 682]. analytics [4654, 4563, 4533, 2986]. analyze [2003, 3169, 3247, 1053, 4535, 4258, 4311]. analyzer [3816]. analyzes [4403]. Analyzing [4458, 3574, 1937, 2766, 1863, 3514, 1655, 4028, 1713, 4270, 4278, 3006, 1539, 911, 2785, 4367, 3433, 749, 977, 2654, 3859, 4678, 4211, 4082, 3130, 661, 3293, 3607, 3553, 2475, 3059]. ancestral [1300]. ANCOVA [1648]. and/or [3411, 2403]. Andalusian [3268]. Anderson [1567]. Andrew [2819]. Andrews [182, 4540]. angle [3292, 839]. angles [3097, 688]. Anglo [185]. angular [2466]. animal [2139, 2583]. Animals [153, 738, 741, 739]. animation [2528]. annealing [558]. Announcement [290, 324, 364, 414]. Annual [678, 1230, 281, 376]. annuities [3455]. anomalies [924, 281]. anomaly [2768, 4541, 4390, 2806]. ANOVA [3998, 501, 3113, 1087, 1264, 1418, 2437, 2058, 2164, 195]. ANOVA-like [3998]. Ansari [774]. Ansari-Bradley [774]. answer [561]. Answers [188, 1345, 1100]. ante [1468]. ante-dependence [1468]. antecedent [2095]. Antedependence [951]. antenatal [4633]. anthropology [75]. anti [3719, 863]. anti-inflammatory [863]. anti-tumor [3719]. anticipatory [4277]. antigen [3046]. antiretroviral [3664, 3050]. antiviral [1546]. any [2491, 321]. AOQL [1090]. APARCH [4287]. APD [3662]. aperture [662]. ApoA1 [3537]. ApoB [3537]. apple [2452]. applicable [1217, 1218]. applicants [1869]. Application [4505, 1028, 794, 3942, 1643, 1641, 809, 285, 2990, 3747, 489, 254, 3374, 1608, 2671, 1430, 1691, 943, 3166, 1707, 3412, 527, 1973, 3726, 1188, 1108, 1448, 2863, 2472, 2304, 3440, 4004, 4324, 499, 4072, 4163, 2782, 2034, 4648, 4344, 2585, 4013, 4687, 2676, 3534, 1894, 397, 1975, 3239, 1729, 3786, 4415, 958, 4388, 4433, 4043, 562, 3344, 4118, 3194, 2099, 3234, 3153, 2280, 2290, 2240, 3788, 2000, 1725, 1019, 3645, 2355, 3158, 4282, 1926, 3229, 560, 2809, 2926, 1311, 611, 4483, 1966, 3582, 4434, 3941, 1582, 3160, 4451, 2620, 2976, 4413, 3045, 3488, 4484, 1900, 3518, 3417]. application [3977, 2153, 2656, 1530, 2288, 3226, 4195, 2279, 3581, 1314, 1500, 1469, 3477, 2263, 317, 3295, 4241, 2360, 3591, 301, 2622, 248, 4565, 2268, 962, 3115, 3407, 3366, 4215, 3327, 1835, 2945, 959, 2740, 4511, 4201, 3737, 1252, 3983, 4028, 3838, 635, 4373, 2581, 4683, 2375, 2643, 3405, 3313, 3716, 3770, 4343, 4489, 3536, 2597, 4274, 4281, 4327, 4654, 3657, 3125, 3339, 3326, 4440, 4315, 3376, 3225, 3603, 2511, 2738, 2405, 2005, 2534, 2349, 3984, 1511, 4595, 4173, 3928, 786, 701, 1902, 4277, 1848, 4381, 4096, 4419, 1304, 3776, 3032, 2390, 3497, 1026, 2496, 4128, 4254, 2911, 2783, 4403, 3802, 4216]. application [4631, 4260, 3991, 4416, 4555, 2584, 2908, 3972, 2907, 2947, 4162, 3597, 695, 2793, 2662, 837, 1005, 4018, 3938, 4136, 4093, 389, 4540, 4402, 3796, 3924, 2352, 3093, 1358, 4332, 589, 4463, 3761, 4158, 3171, 3040, 4568, 3003, 3996, 591, 2638, 1197, 4491, 3858, 2071, 444, 2719, 3800, 286, 525, 1328, 2169, 2449, 2569, 2865, 2095, 3353, 3914, 4539, 4070, 4406, 2168, 3550, 2137, 2801, 2832, 2946, 3483, 4342, 4453, 4500, 1933, 2444, 4521, 3262, 3625, 4385, 3410, 2884, 4368, 4246, 1517, 4014]. application-based [3171]. Applications [4472, 4712, 1377, 1488, 1616, 2191, 319, 1420, 3319, 4664, 4139, 3998, 2677, 3485, 3753, 4650, 3671, 548, 676, 4318, 3766, 3980, 1566, 4068, 1807, 3700, 4487, 1598, 2744, 3889, 4356, 4134, 4334, 3470, 4192, 3705, 2048, 3710, 4200, 2422, 1145, 1495, 2688, 1073, 4140, 3862, 3325, 4479, 4428, 4466, 2413, 3323, 3896, 4554, 2317, 2434, 2236, 4352, 3130, 1559, 2520, 2598, 4496, 1940, 4607, 3829, 2509, 3596, 4020, 4414, 3772, 4589, 4417, 359, 3396, 2218, 3812, 4116, 1579, 1055, 4431, 2987, 3233, 3876, 1114, 4050, 2197, 4675, 3230, 543, 1600, 3553, 2728, 4123, 3091, 4147, 4689, 4218, 3974]. applications [3132, 4291, 2625, 1016, 4205]. Applied [1715, 2498, 2463, 1700, 210, 4699, 1644, 2890, 115, 4738, 2258, 2647, 39, 1796, 3888, 4499, 977, 3286, 4657, 4418, 4005, 4478, 4148, 2300, 44, 2604, 2893, 3197, 4679, 3225, 952, 1131, 3966, 4524, 3727, 3271, 95, 2164, 4219, 1058, 699, 2382, 879, 2822, 3021]. Applying [2357, 2250, 1402, 1148, 3714, 4536, 3441, 4471]. appointment [3936]. apportionment [4244]. appraisal [3577, 265]. Approach [1426, 1631, 221, 4699, 1409, 1675, 1704, 1691, 1393, 4377, 1697, 200, 4709, 1452, 3260, 855, 3157, 2051, 4226, 2805, 2245, 3035, 1271, 2574, 2874, 3480, 1333, 3512, 3834, 2896, 573, 668, 2553, 4273, 3847, 3658, 1941, 2706, 2880, 2913, 3694, 4432, 958, 3232, 3199, 4259, 2098, 3468, 3168, 2866, 2649, 717, 1782, 2240, 2244, 2224, 2370, 3532, 3700, 3894, 3329, 2582, 1276, 2070, 1309, 1186, 238, 2166, 3334, 2284, 3303, 2332, 2876, 1517, 2586, 4411, 2827, 2605, 2294, 2828, 4609, 3083, 2069, 2689, 3009, 1946, 4036, 970, 2656, 1159, 3104, 3034, 2139, 1144, 2575, 1512, 3306]. approach [642, 2956, 3358, 1779, 3848, 3945, 4113, 1546, 1950, 797, 2506, 1872, 3537, 2404, 3611, 418, 4178, 2673, 1089, 2497, 2234, 3633, 1786, 2010, 4297, 916, 4572, 4174, 1519, 2266, 2214, 1952, 3084, 2105, 2830, 2431, 2989, 4583, 2185, 2877, 3354, 3809, 864, 2242, 3592, 3139, 2334, 1762, 2592, 2336, 2140, 3588, 597, 1056, 1326, 2774, 3823, 3544, 2914, 875, 4608, 1255, 1352, 3681, 465, 3542, 3032, 4501, 3253, 1506, 4285, 1810, 2428, 1267, 3346, 833, 3725, 4078, 2783, 1194, 2061, 3675, 868, 3850, 2219, 3886, 3159, 4589, 4531, 3287, 1497, 993, 4186, 4528, 4217, 2555, 3031, 692, 4600, 2523]. approach [3416, 1834, 3494, 2201, 3761, 3116, 2517, 3990, 2340, 2939, 777, 3895, 3257, 3462, 1346, 2373, 4053, 113, 2199, 2996, 2449, 3801, 3779, 2095, 887, 4401, 3431, 4250, 2812, 3629, 4557, 2253, 3463, 1540, 3201, 4398, 3729, 4368, 2089]. Approaches [1389, 1237, 4443, 2479, 1372, 1951, 2966, 4584, 2346, 492, 4168, 422, 3843, 3795, 1588, 2944, 720, 4580, 1601, 3434, 4440, 4172, 597, 4464, 2401, 3385, 3046, 3352, 1032, 2209, 1876, 2836, 3296, 4185, 955, 1238]. appropriate [644, 2919, 2154]. Approximate [1980, 3734, 1095, 320, 2605, 3292, 1074, 1133, 3052]. Approximating [679, 1889, 4421]. Approximation [1425, 4478, 2059, 2910, 682, 2443, 4403, 1048, 3353]. Approximations [1613, 4165, 3599, 3974]. AQL [643, 1090]. aquatic [2671]. Aquifer [1619]. Arabidopsis [1336]. Aranda [3161]. arbitrary [3351]. ARCH [2107, 1760, 3926]. Archimedean [1897, 3437, 2680]. architecture [4585, 4492]. Arctic [2926]. ARDL [1159]. Area [1417, 1436, 1633, 1956, 4119, 2280, 2603, 2149, 1321, 4214, 1466, 2635, 1105, 3559, 1983, 4589, 692, 3858, 3902, 4570, 1137]. area-level [4119, 2603, 1983]. areal [3737]. areas [569, 3652, 2907, 742, 3192, 1282, 2303]. ARFIMA [2051]. ARFIMAX [1806]. ARFIMAX-TARCH [1806]. Argentina [3309]. Argentinean [1247]. argument [237]. arid [1247]. ARIMA [1359, 409, 3390, 1428, 911, 4548]. arising [4005, 632]. arithmetic [3420]. arm [4464, 4333]. ARMA [2133, 2742, 2211, 1701, 4358, 1196, 1879, 2975].

ARMA-GARCH [2133, 4358]. Armed [3076, 1487, 3589]. arms [1456]. Arnold [510]. array [1331, 875]. Arrays [1634, 1822, 1830, 2574, 2930, 4057, 888, 2650]. arrival [3760, 678]. arrivals [4042]. art [1232, 3110]. arthritis [2876]. article [2290, 4140]. Artificial [1019, 1141, 3525, 2595, 3911, 4180]. ary [468]. Ascertaining [1398]. ascertainment [632]. Ashford [165]. Asia [3814]. Asian [4512, 4162]. Aspects [1731, 117, 805, 1553, 2880, 24, 781, 3105, 3896, 4274, 453, 1739]. assay [1058]. Assembly [4360]. assess [2235, 3899, 1084, 2740, 3264, 1477, 3714]. Assessing [1869, 927, 998, 3061, 667, 3965, 4345, 3775, 635, 2111, 1003, 1465, 4559, 4014, 3906, 581, 2905, 2340, 399, 1817, 1702, 1316, 2253, 2025, 2602, 1228, 3016, 3561, 1352, 1014, 4380, 1921, 4605, 4129, 2527]. Assessment [152, 165, 4731, 4174, 1957, 704, 1113, 2103, 1795, 4549, 843, 329, 2942, 3450, 3161, 2815, 3050, 1719, 1052, 344, 4232, 4327, 4601, 3966, 1753, 2389, 4395, 3412, 1204]. Assessments [1380, 3970, 3999, 440, 3448]. Asset [2503, 1728, 4093, 1834, 3822]. assets [789]. assignable [1554]. assigned [1976]. Assigning [3939]. assignment [3624, 3185]. assist [1552]. assistance [3659]. assisted [4460, 4571]. associate [960]. associated [2302, 3743, 3296]. association [2025, 1923, 2800, 3452, 2560, 3160, 341, 4266, 960, 3981, 3593, 4437, 3272, 3754, 1120, 4403, 3218, 4394, 1876, 1817, 3516, 2527, 397, 4]. associations [4121, 2236, 4595]. assuming [4190]. Assumption [1637, 1410, 1932, 3122, 3117, 2440, 882, 2884]. assumptions [3661, 971, 3053]. asthma [2795]. astronomical [3766, 419]. Asymmetric [3298, 4274, 1618, 695, 1229, 1702, 3769, 4387, 2632, 3639, 805, 2976, 2925, 2776, 4276, 2576, 4552, 3428, 3276, 1255, 3604, 1734, 4154, 1593, 3483, 3074]. Asymmetrical [1560]. Asymmetrically [1619, 1855, 3212]. asymmetries [3957].

Asymmetry [1614, 1394, 2564, 4092, 1173, 3938, 2378, 2979]. Asymptotic [2430, 542, 4701, 2158, 4003, 2868, 588, 1829, 3266, 1915, 1056, 1326, 1097]. asymptotically [1125]. asymptotics [2501]. Ata [4167]. athletes [823]. Athletic [1660, 2361]. Atlantic [754, 4520, 678, 3482]. atmospheric [604, 2071]. attachment [3655]. attack [3495]. attacks [2521]. attempt [106]. attempters [3091]. attendance [342, 2182]. Attention [1664, 4445, 2638]. attention-deficit [2638]. attention-deficit/hyperactivity [2638]. attitude [58]. attitudes [3173, 2851]. attraction [3666]. attribute [871, 596, 3408, 1041, 3513, 2080, 2837, 1747, 3758]. attributed [4218]. attributes [2100, 360, 1721, 247, 1070, 3612, 2029, 2769, 1726]. attrition [4494]. Atypical [398, 2740]. auctions [2780, 2980]. Audit [183, 4325]. auditing [722]. audits [3127, 1511]. Augmentation [1617]. Augmented [4136, 2370, 4255]. Augmented-limited [4136]. Australia [3689, 3581]. Australian [185, 1300, 1382, 1502, 1610, 566]. authentication [2254]. Author [822, 876, 946, 1166, 770]. authored [4669]. authors [878]. authorship [3174]. auto [2797]. auto-regressive [2797]. Autocorrelated [1679, 1655, 1397, 3263, 1720, 4567, 1945, 1947, 2267, 921]. Autocorrelation [1401, 1261, 4100, 2142, 2861, 1042, 1323, 2436, 4327, 1174, 4095, 3370, 3165, 4374]. autodependogram [3317]. Autologistic [1661, 2304, 3595]. automata [2830]. automated [970, 420]. Automatic [3064, 3687, 1371, 4036]. automobile [301]. automotive [4184, 4371]. autoregression [938, 923]. autoregressions [4012, 4496]. Autoregressive [2391, 1448, 4664, 2447, 4268, 1999, 598, 949, 4438, 4223, 2887, 1000, 2976, 2682, 3676, 3813, 2163, 2737, 815, 1054, 4274, 4654, 3231, 3110, 1991, 2774, 4067, 4644, 1036, 2403, 2510, 2983, 3876, 2119, 2197, 3524, 2692, 4346, 4500]. autoregressive-autoregressive [2774]. auxiliary [2830, 1206, 2315, 1207, 3864]. Average [1545, 144, 1388, 4364, 4523, 1917, 2797, 4423, 1180, 1847, 1141, 1042, 1466, 644, 1041, 815, 3144, 2727, 3752, 4659, 1036, 4528, 4006, 2259, 819, 1319, 861, 935, 4405, 1327, 1368]. average-least-squares [2797]. Averaging [4698, 1792, 3700, 3638, 4592, 2434, 2236, 4583, 3577, 1559, 4604, 2012, 1346, 2037]. avian [1219, 1360]. axes [375]. Axial [634]. Axis [1415, 2376]. Azzalini [1096].

B [4155, 1511, 2232]. B-robust [4155]. B-spline [2232]. B. [1729]. BA [3056]. Bachelor [3509]. Bachelor-to-Master [3509]. Back [1413, 285, 2133]. Back-projection [1413]. backcross [4468]. background [3564]. backwards [1819]. bag [1233]. bagging [2127, 3629]. Baker [1659]. Balan [2651]. balance [2312]. Balanced [4015, 3769, 1168, 4650, 1083, 2367, 1338, 1263, 4492, 4295, 1350, 4662, 1463, 1560]. Band [2697, 1222]. Band-limited [2697]. band-recovery [1222]. banding [1230, 1233]. bandit [3589]. bandit-based [3589]. bands [2333, 3465, 2151, 3496]. bandwidth [4149, 4062, 4294, 2311, 3074]. bandwidths [4485]. Bangladesh [3805, 3950, 3065, 3305, 2738, 1814]. bank [3008, 3786, 4415]. banking [3346]. bankruptcy [2115]. banks [4171, 4162]. Barcelona [3797]. Barents [3482]. Barker [1234]. barley [1740]. barnacle [742]. Barro [3487]. Bartlett [4016]. base [2301, 3974]. Baseball [2001, 952, 4125]. Based [1631, 1418, 1445, 1617, 1618, 1416, 1385, 1702, 3937, 3331, 2042, 829, 2805, 3440, 3703, 1507, 3907, 4046, 3028, 3970, 3035, 4072, 522, 3480, 1030, 2741, 4126, 471, 4047, 3455, 4516, 463, 668, 2102, 2974, 3847, 3658, 3865, 4684, 3199, 4375, 4499, 4550, 4670, 3491, 3767, 2568, 3168, 3999, 1523, 2347, 4582, 1083, 2238, 3728, 4458, 1891, 2110, 513, 4581, 3311, 1972, 1455, 3662, 1248, 3391, 3446, 2332, 4340, 1783, 2901, 3638, 999, 4590, 2481, 4161, 4633, 3011, 1149, 3298, 4435, 1877, 3318, 3987, 2294, 2069, 4329, 3417, 970, 2674, 1958, 4184, 4305, 2991, 3226, 4160, 2068, 4030, 1258, 4102]. based [3854, 733, 564, 3631, 3965, 3472, 4396, 3407, 3152, 1088, 3537, 1992, 3598, 2641, 1835, 2673, 2590, 3925, 3878, 1121, 4122, 851, 4371, 960, 4508, 4240, 3695, 2522, 3644, 1846, 2886, 1160, 1490, 2396, 1331, 2214, 2391, 3221, 2215, 2526, 2469, 3185, 3952, 3933, 2116, 1344, 1213, 2350, 4456, 3341, 3367, 2302, 3748, 4651, 1895, 1355, 1661, 1774, 3095, 2446, 3650, 3855, 2465, 3223, 4351, 3546, 3646, 2580, 2592, 3361, 1315, 1023, 2811, 3435, 4299, 2536, 2860, 1365, 3825, 4096, 3951, 3101, 1541, 4231, 4057, 2788, 3058, 3098, 3343, 3183, 3971, 4061, 2826, 1026, 2838, 4208, 4585, 3959, 1529, 4120, 3675, 4445]. based [423, 4236, 3820, 675, 4008, 1336, 4186, 2381, 2784, 2907, 2510, 2160, 4614, 2063, 4212, 4626, 4217, 3291, 3476, 3918, 2955, 1005, 3192, 4006, 4093, 3403, 3955, 4540, 4141, 3589, 839, 4265, 3791, 3370, 4227, 4480, 2898, 2840, 4158, 3171, 3876, 4568, 3990, 4094, 1580, 2638, 3188, 2164, 4481, 3070, 1274, 2150, 3744, 3230, 969, 1955, 1959, 2743, 2588, 3055, 4106, 4522, 4125, 1375, 1716, 2271, 2918, 1302, 1921, 2915, 3516, 4247, 4349, 4513, 887, 3053, 4107, 4133, 3864, 1204, 4308, 4613, 4561, 1738, 2311, 2354, 3668, 4083, 2571, 2698, 2878, 4407, 4521, 4457, 3463, 4459, 3432, 3693, 3656, 4029]. based-kernel [4240]. baseline [3172, 3807, 4378, 4105]. Basic [2273, 77, 661, 240]. Basis [152, 4719, 3769, 2772, 1101, 4015, 3817, 4336, 4637]. basketball [1272, 3532, 4197, 3984]. Basu [3260, 2687, 2261]. Basu-Dhar [3260]. batched [3897]. batches [3769]. Bath [78]. bathtub [2805, 3328, 3520, 3147, 2515, 2220, 4225, 4213, 1804]. bathtub-shaped [2805, 2515, 2220, 4213, 1804]. Batschelet [2175]. Batteries [1641]. batting [4125]. Baumgartner [1541]. Bay [1517]. Bayes [1869, 687, 3181, 2507, 1664, 1495, 3325, 4437, 1024, 3697, 2992, 169, 726, 2940, 3625]. Bayesian [2091, 2815, 3260, 2304, 3388, 4359, 4226, 2805, 2096, 4224, 1419, 4279, 3805, 3256, 4711, 3242, 3575, 3569, 4652, 3970, 1510, 3399, 3679, 2709, 1809, 2970, 4664, 2337, 1291, 1669, 3917, 1795, 3184, 3834, 2896, 4238, 2307, 1954, 2102, 553, 4273, 1894, 2463, 3908, 2387, 451, 3665, 3694, 4432, 2920, 1775, 1792, 1953, 1796, 3780, 3551, 1758, 598, 1212, 1843, 3320, 3828, 3345, 1811, 2098, 3468, 4199, 3564, 3706, 3069, 3526, 2240, 2244, 2900, 4175, 3329, 3182, 3203, 1910, 3450, 1926, 3892, 2632, 3541, 3334, 2655, 4553, 1542, 2715, 1378, 3314, 4321, 2281, 4356, 4134, 4411, 4223, 2605]. Bayesian [3292, 3630, 2887, 2294, 3913, 3045, 4370, 4609, 4731, 2547, 3488, 2682, 2776, 3170, 3543, 3009, 4048, 4036, 3150, 3042, 4625, 2656, 3678, 3883, 1301, 4012, 3581, 2048, 1512, 3636, 1314, 2458, 790, 3337, 2254, 1779, 2729, 3259, 4266, 4082, 3848, 3591, 2019, 3567, 1546, 1950, 797, 594, 3537, 1597, 2404, 2944, 3750, 1719, 1878, 2059, 4328, 1518, 2552, 2945, 2114, 2679, 2485, 2234, 2672, 1254, 3633, 4479, 4240, 2910, 3263, 3737, 3661, 4269, 3418, 2644, 3652, 2365, 1054, 1129, 4734, 3397, 3434, 2434, 912, 1074, 3561, 3716, 4520, 2454, 3601, 3438, 2185, 2796, 1861, 1988, 3826, 4440, 4588, 1895, 2962]. Bayesian [3585, 3577, 2405, 415, 570, 1974, 3749, 4558, 2190, 3715, 2960, 2005, 2474, 3175, 1940, 2937, 1617, 982, 1079, 1393, 2930, 3704, 4595, 3509, 3522, 3829, 4277, 2272, 3307, 3036, 3422, 3823, 4381, 1978, 2172, 1492, 2788, 2630, 3272, 1528, 1132, 1255, 1352, 1304, 2072, 4244, 1707, 3032, 1859, 3762, 4486, 3545, 1810, 2041, 2496, 3699, 2959, 4254, 2972, 3596, 4464, 4644, 4677, 833, 4078, 3916, 2783, 4512, 2762, 4555, 2219, 1696, 4559, 4531, 3100, 3240, 3287, 3911, 3251, 1432, 2908, 3249, 3396, 1744, 1939, 1919, 4341, 2177, 2704, 4528, 2016, 3615, 4255, 1767, 3291, 3476, 3046, 2600, 1866]. Bayesian [2324, 2749, 4062, 1865, 3219, 2669, 3403, 1730, 110, 3589, 4294, 4600, 4637, 2031, 2636, 1579, 2209, 659, 3116, 3040, 1298, 1422, 2795, 706, 2012, 2340, 3696, 3364, 3296, 4080, 3919, 3963, 2223, 3466, 3372, 1346, 2062, 2411, 2854, 3968, 4106, 4053, 4229, 2980, 3811, 3808, 1622, 298, 1702, 1921, 2449, 4123, 4088, 2856, 2855, 2932, 3041, 2642, 4653, 1452, 2168, 4250, 1178, 2137, 2681, 2946, 2625, 3619, 3074, 3442, 3832, 2155, 3729, 4368, 2502, 2261, 4029, 4527, 954, 2748]. Bayesian-frequentist [2900, 2168]. BBC [264]. BCa [2873]. BCPE [2417]. Be [138, 132, 947, 44, 2491, 3621, 4278]. Bechuanaland [287]. bee [3525]. been [2853]. before [377, 2714, 1200]. Behavior [1749, 4588, 2708, 4419, 1565, 2063, 3264, 1960, 4409, 3013, 2382]. behavioral [1926]. behaviors [3853]. Behaviour [722, 2173, 3410]. Behavioural [2616, 1373, 1211, 1228, 2691]. Behrens [464, 1199, 4444]. being [369]. belief [516, 523]. beliefs [2106]. Bell [2653, 3884]. below [1285]. benchmark [3912, 4327]. benefit [3043]. Benford [1645]. bent [3633]. bent-cable [3633]. Bera [3662, 1628]. Bercai [174]. Berlin [2562]. Bernoulli [1729, 797, 2629, 3368, 4352, 4658]. Bernstein [3329, 2166, 3314]. Bessel [4606]. Best [872, 3999, 4714, 886, 599, 354, 2837, 1911, 1308]. Best-subset [3999]. Beta [1429, 1672, 4502, 3353, 1356, 3270, 2049, 2971, 3888, 2869, 2800, 2216, 3030, 1746, 1158, 4397, 1848, 3587, 2412, 3545, 4341, 2624, 3822, 568, 2763, 4217]. Beta-Binomial [1672, 1356, 2049, 1158, 568]. Beta-negative [4502]. beta-regression [2216]. Better [1392, 3637, 3807, 4492, 1111, 2491].

Betting [1668, 304, 1098]. between

[1330, 1128, 4171, 584, 668, 397, 1008, 1836, 4121, 3301, 1294, 1913, 2335, 2481, 4532, 4370, 3309, 3803, 3437, 4042, 2607, 4396, 2300, 852, 787, 1341, 2126, 3593, 2020, 2256, 4567, 2661, 3095, 1230, 3754, 2660, 4379, 4531, 1412, 3817, 1499, 1229, 3006, 3898, 581, 3416, 4138, 1006, 1712, 1876, 1817, 3492, 688, 4054, 4405, 1204, 3547, 2480, 4129, 3262]. between-respondent [1499]. between-study [2661]. beyond [3929, 526]. Bézier [4516]. BIAS [115, 3205, 4388, 2003, 1398, 4001, 744, 1573, 1624, 3641, 3560, 2338, 1941, 1353, 2920, 1311, 1587, 320, 1755, 1239, 65, 3151, 4031, 1984, 80, 1297, 593, 893, 2836, 2385, 740]. Bias-corrected [4388, 3560, 2338]. bias-variance [2385]. biased [4114, 369, 1832, 4019, 4082, 2314, 868, 3140, 1749, 4525, 2556]. biases [753]. BIB [1182]. biclustering [3048, 3159]. bicycle [3705]. bidimensional [1894]. bidirectional [4445]. bifactor [3832]. Big [4443, 3380, 3808]. bilateral [4474, 3839, 4380]. bile [4568]. Bimodal [3676, 3639, 3062, 4389, 3604, 4018, 4050, 4219, 2694, 4362, 4622]. Bimodality [4442]. bin [4142]. Binary [1644, 4159, 1686, 1696, 2304, 1249, 2091, 1125, 3837, 4238, 3107, 4097, 3491, 1272, 2129, 3224, 2003, 3452, 999, 2787, 1877, 4413, 3112, 1584, 1369, 3831, 3458, 1992, 4663, 3624, 4283, 1130, 4636, 2431, 2958, 3326, 3225, 4558, 3890, 3509, 4337, 3610, 322, 2031, 4235, 2340, 3078, 2720, 4429, 3940, 2760, 2832, 4343]. binding [2792]. Binomial [1605, 1672, 504, 3220, 4098, 1809, 1356, 2585, 3485, 4296, 397, 2049, 4438, 2060, 4347, 4678, 2957, 2007, 641, 3042, 1158, 2241, 3989, 2519, 1743, 3114, 4565, 1588, 1153, 1165, 4507, 4493, 4520, 1925, 1351, 1808, 1046, 1541, 2645, 2412, 2904, 3972, 1499, 3289, 4006, 2075, 2259, 1501, 1110, 1339, 4502, 2588, 2565, 1856, 1813, 568, 3885, 2278, 3625]. binomial-beta [2412]. binomial-mixed [4438]. binormal [2333]. biochemical [3512]. biodiversity [2782, 2444]. bioelectrical [4118]. bioequivalence [3239, 2635, 1352]. Biography [207]. Bioinformatics [1845, 2520]. Biological [1681, 792, 4483, 260, 669, 1058]. biologists [750]. Biomarker [2982, 3152, 3374, 3095, 3787, 2449, 2095]. Biomarkers [1380, 2917, 2302, 3148]. Biomedical [548, 2888, 2197]. biomedicine [4676]. biometric [3528]. Biometrics [207, 353]. biometry [19, 111]. biosignals [2643]. Biostatistical [4724, 2394]. Biostatistics [4723, 2506, 4741, 1861, 1862, 2817, 2651]. biplot [3131, 4127]. Biplots [616, 1470]. bipolar [1578, 3817]. bird [760, 1235, 1217, 767, 1218, 765]. birds [794, 732, 1360, 740]. Birnbaum [4013, 3671, 4131, 3582, 3676, 1442, 3404, 4030, 2308, 2893, 2159, 3536, 3333, 4271, 4556, 2218, 2712]. Birnbaum-Saunders [4131]. Birth [147, 263, 3805, 3045, 409, 2738, 4022]. Birth-date [263]. birth-death [4022]. birthmark [3367]. births [531, 1531]. Bissau [1462]. bitcoin [3827]. Bivariate [2800, 1671, 1776, 2153, 1605, 3425, 2349, 4271, 3972, 1647, 1363, 2687, 3260, 3480, 3994, 2235, 4047, 2363, 1975, 2099, 3745, 2060, 3429, 1486, 4365, 3160, 1602, 4214, 2007, 3678, 4192, 3863, 2529, 2629, 3514, 4335, 3118, 4069, 4035, 1284, 3141, 3341, 4322, 2029, 1974, 1885, 3928, 1848, 4087, 2050, 1013, 4501, 4677, 3148, 1982, 4614, 3218, 4006, 2795, 4675, 4261, 3377, 1593, 2999, 3729, 2261, 4527]. biweight [465]. black [743, 4330]. bladder [4254]. Bleaching [1424]. Blind [2886]. bloc [955]. Block [1626, 1663, 102, 1386, 468, 897, 4097, 1979, 948, 1031, 944, 990, 1085, 1338, 4021, 842, 2379, 3617, 4017, 4391, 318, 4492, 4000, 3499, 2154, 2477, 1463, 2687, 2261]. block-structured [1979]. blocked [4133]. blocking [686, 1849]. blockmodels [4682]. blocks [516, 3350, 1182, 4021, 1263, 603, 834, 926]. blockwise [904]. blood [2866, 2759]. Blooms [1651]. blown [2626]. blue [38, 4401]. BLUP [1633]. BMDP [610]. BMI [4370]. Board [2092, 2328, 2756, 2954, 3155, 2543]. body [3256, 2235, 2387, 4118, 3361, 1803, 2251, 2862, 3580]. Bogota [4332]. bole [788]. bomb [1997]. bond [1962, 3814]. bone [2019, 2493, 4123]. Bones [158]. Bonferroni [1023]. Bonus [1512]. Book [219, 204, 216, 227, 228, 229, 230, 231, 242, 243, 244, 255, 256, 257, 258, 269, 270, 271, 272, 273, 274, 275, 276, 288, 289, 291, 292, 293, 294, 295, 306, 307, 308, 309, 310, 311, 312, 323, 325, 326, 327, 347, 348, 349, 350, 351, 352, 401, 402, 427, 428, 429, 430, 431, 432, 433, 434, 435, 436, 437, 438, 439, 455, 533, 534, 535, 536, 537, 538, 545, 582, 638, 639, 783, 803, 813, 821, 832, 841, 850, 922, 933, 945, 987, 1027, 1038, 1049, 1069]. Book [1080, 1123, 1179, 1256, 1268, 1299, 1332, 1376, 1447, 1464, 1471, 1480, 1489, 1504, 1568, 1577, 1585, 1594, 1603, 1612, 1630, 1640, 1649, 1658, 1667, 1676, 1685, 1695, 1705, 1714, 1724, 1732, 1742, 1751, 612, 1788, 2228, 2395, 2820, 1826, 647, 650, 384, 385, 627, 474, 2090, 2124, 623, 190, 365, 649, 4729, 177, 193, 626, 2822, 2816, 2815, 56, 2394, 2817, 2951, 3572, 622, 2194, 2814, 218, 2227, 473, 477, 629, 388, 192, 191, 205, 166, 179, 203, 361, 363, 386, 475, 478, 2036, 189, 180, 1621, 2226, 2277, 2393, 202, 628, 362, 2066]. Book [2192, 2274, 2746, 2953, 2950, 1787, 215, 1880, 1961, 4669, 2745, 2824, 2819, 2125, 2823, 1987, 2276, 2225, 2949, 2064, 2158, 2273, 2195, 648, 3026, 387, 2191, 1824, 1825, 1883, 1882, 1884, 2821, 2948, 3120, 201, 194, 625, 217, 178, 206, 476, 2193, 2156, 1777, 1881, 2123, 2157, 2818, 2275, 624, 2065, 2747, 4599, 2653, 2952, 1245, 2088, 2037, 366, 2748, 1789, 2651, 2652, 367, 2749, 2814]. Boolean [1576]. boosting [4040, 4592]. Bootstrap [211, 3804, 1158, 4620, 2130, 395, 4312, 3209, 3053, 499, 4681, 4268, 1780, 3420, 927, 1294, 3051, 1287, 1183, 1277, 1490, 1860, 1104, 2184, 1590, 848, 1267, 904, 1032, 3906, 4138, 1741, 2717, 3177, 2588]. Bootstrap-adjusted [4620]. bootstrap-based [2588]. bootstrapped [2436]. Bootstrapping [1604, 4602, 1624, 1068, 4007, 1073, 980, 2044, 1091]. borderline [1925]. borne [2478]. Boston [3130]. both [2881, 2298, 389]. bottle [3481]. bottles [3481]. bottom [3297]. bound [1842]. boundaries [1989, 3792]. boundary [1057]. bounded [2585, 3391, 2876, 4625, 3989, 3983, 4428, 1284, 3587, 4071, 2694]. bounds [1011, 4421, 2673, 3895]. Bovidae [3686]. Bow [547]. bowel [3125]. bowler [3623]. box [4386, 4195, 2492, 2722, 3276, 461, 283, 2666, 4106, 2525, 3779, 2623]. boxplots [3223]. Bradley [3499, 2135, 774]. Brain [251, 3706, 2945, 924, 3765]. branches [3008]. branching [4536]. brand [3586, 32]. brant [743]. Branta [742]. Brazil [4503, 4025, 4311, 3442]. Brazilian [3194, 4287, 3162, 4317, 4368]. breadth [3857]. break [4165, 1835, 3258]. breakdown [3391]. breaking [3489]. Breaks [199, 2835, 845, 2737, 2040, 2072]. breast [2429, 1795, 843, 3295, 4456, 1990, 4406, 4218, 2303]. breed [2889]. breeding [1236, 2139, 745, 766, 1227, 589]. Brewing [1692]. Brexit [4169]. BRFSS [3361]. bridge [1008, 891, 3585, 3948]. brief [671]. bringing [3]. Britain [611, 251]. British [757, 4, 3340, 931, 2445, 756]. broken [1793]. brood [1228]. Brownian [1034]. BROWNIE [746]. BSO [45]. bubbles [4178]. Buckley [3057]. budget [3153]. budget-share [3153]. BUGS [2749]. build [3699]. Building [516, 1152, 2551, 2701]. buildings [3210]. built [3570, 3995]. bulb [1503]. Bulletin [115]. Bundesliga [3775, 3620]. burden [3794]. Burman [1191]. burn [3405]. burn-in [3405]. Burnham [1218]. Burr [3503, 4650, 4318, 3541, 4590, 1269, 3891, 817, 2350, 2770, 3497, 2204, 2572, 1996]. Burr-Hatke [4650]. Burr-type [3541]. burst [4018]. bursts [3846]. Business [1614, 1173, 4691, 1553, 1140, 1138, 2713, 4148, 93, 3106, 3529, 4172, 3401, 2140, 701, 2611, 1055, 2986, 1812]. butterfly [3007, 1188]. butterfly-distributed [3007]. buyer [3069].

c [576, 1408, 1545, 1094, 2609, 1189]. c1 [544, 556, 645, 683, 718, 709]. c2 [544, 556, 645, 683, 718, 709]. cable [1784, 3633]. Caesarean [3468]. Cagdas [2745]. calculate [3401]. Calculating [931, 972, 39, 2247, 1567]. Calculation [188, 1380, 346, 1973, 3730]. calculations [1118, 4084, 3352, 2565]. Calculator [504]. Cali [2496]. calibrated [4589, 3924]. calibration [4252, 4473, 711, 2332, 2964, 1968, 4400, 4172, 4604]. calibration-based [2332]. California [3571, 736]. call [2794, 4360]. calling [2650]. Cam [1966, 1202]. Cambanis [4035]. Cambanis-type [4035]. camera [3850]. cameras [2721]. Can [44, 1130, 2491, 4278]. Canada [1236, 754]. Canadian [151]. Canary [1424]. Cancer [1430, 1674, 2429, 1795, 4043, 843, 1755, 2829, 3477, 3295, 4102, 3366, 4462, 2945, 4281, 4456, 3474, 2511, 2410, 3559, 2412, 2390, 4254, 1990, 993, 3046, 2209, 1924, 2232, 1831, 2047, 966, 4406, 4218, 2303, 2137, 2801, 4368]. cancer-related [1924]. Candidate [3981]. Candidate-gene [3981]. candidates [4492]. candlelight [4303]. canine [2560, 2448]. Canonical [117, 2627, 559, 560, 1791, 1470, 1052, 3109, 670, 4249, 558, 2399]. Capability [2893, 1690, 1842, 2810, 1270, 3804, 3985, 2469, 1283, 4002, 1942, 2838, 941, 1529, 4412, 630, 920, 1921, 858, 921]. capacity [3787, 488]. Cape [737]. capita [3858]. capital [4286, 611, 1728, 1148, 1834, 4170, 2997]. capitalists [4173]. Capture [761, 1203, 1200, 900, 1238, 735, 1232, 767, 1228, 1214, 744, 1239, 1226, 764, 745, 751, 1237, 738, 1206, 747, 1223, 1215, 1231, 1114, 752, 2229, 2540, 1207]. Capture-recapture [761, 1203, 1200, 900, 1238, 735, 1232, 1228, 1214, 744, 1239, 1226, 764, 745, 751, 1237, 738, 1206, 1223, 1215, 1231, 1114, 752, 2229, 2540, 1207]. captured [1236]. Capturing [4316, 2165, 4157, 3681]. Carbon [916, 3355, 3771, 4659]. carcinoma [4187]. card [1373]. CARDIA [3942]. cardinality [2622]. cardiomyopathy [1116]. cardiopulmonary [2441]. cardiovascular [4396, 3247, 2844, 4595, 4453]. care [2744, 4633, 4325, 1900, 4503, 3737, 2264, 3602]. careers [3158]. caregiver [4540]. caries [2304, 2056]. Carl [1216]. Carlo [1274, 2805, 3081, 1643, 211, 3491, 1409, 601, 827, 1763, 1835, 2740, 2221, 2910, 2105, 3657, 2897, 1255, 4272, 4512, 1801, 3385, 3228, 866, 3040, 3808, 1484]. Carlo-based [3491]. Carolina [3720]. carpooling [4486]. carried [579, 3178]. CART [3866, 930]. Carter [4245, 1521]. Cascade [2475]. cascading [3846]. Case [3502, 3641, 1688, 3383, 4538, 2245, 2363, 929, 2262, 1948, 3694, 84, 3199, 3279, 2686, 3069, 793, 1141, 1829, 2335, 917, 4075, 1755, 1784, 341, 4005, 2296, 2538, 732, 2532, 585, 1572, 3042, 4184, 845, 1472, 3487, 71, 724, 2742, 4371, 4542, 2010, 346, 619, 4269, 2236, 673, 4651, 4017, 1574, 1163, 355, 3140, 3159, 2081, 4394, 3614, 907, 1943, 1234, 1227, 3721, 2785, 4502, 3919, 752, 2264, 2482, 113, 2856, 3013]. case-cohort [4075]. Case-Control [3641, 3694, 1755, 585, 1572, 2010, 2236, 673, 4394, 2482]. Case-deletion [1688]. cases [1467, 4534, 4552, 3136]. cash [3498, 3669, 4233]. casualty [2721]. casuarina [3524]. catalog [3773]. Catalogue [856, 1404]. catalyst [1725].

CATANOVA [1697, 3312]. catastrophe [4175]. catch [4316, 1840]. catches [1053]. Categorial [4738]. Categorical [4739, 3609, 2573, 3616, 2988, 3665, 4199, 1930, 2822, 3939, 2153, 4248, 947, 4328, 4092, 3593, 2236, 2958, 2256, 3632, 2796, 3145, 1324, 3654, 3723, 3273, 3684, 975, 2963, 3796, 3963, 2339, 1593, 4129, 2823]. Categories [1395, 1394, 2348, 1549, 2378, 2979, 909, 3018]. categorization [1953]. category [2771]. category-wise [2771]. Cattle [1729]. Cauchy [381, 3414]. Causal [3515, 4396, 3669, 3797, 3941, 4042, 4459]. Causality [2676, 1330, 3029, 3706, 4615, 1258, 2043, 2717, 3003, 1109]. Cause [4298, 3868]. Cause-specific [4298]. caused [4262]. causes [3893, 3249, 2310, 3220, 1554]. cautionary [4207, 1014]. Cauvery [3524]. CD [4157]. CD-vine [4157]. CD4 [2000, 3050, 2711, 3747, 3547, 3132]. CDC [4559, 1108]. CDOs [2227, 4716]. ceiling [1354]. Cell [1641, 3567, 1336, 4649]. Cells [1438, 2715, 2740, 1749]. cellular [2830]. cement [2019, 2785]. Censored [1615, 1698, 1453, 3086, 3260, 4144, 4359, 3769, 3440, 4004, 3336, 2741, 4047, 463, 2113, 4375, 2021, 2347, 3122, 4267, 2659, 482, 3648, 1294, 2632, 3541, 2926, 2935, 2318, 4365, 3873, 1516, 3150, 4305, 1530, 3863, 950, 3014, 4194, 3359, 564, 3965, 1981, 1969, 2672, 3508, 3418, 3392, 1344, 2350, 2916, 3947, 4222, 1868, 3748, 4651, 2178, 3646, 3893, 4361, 3493, 4672, 2892, 3497, 4254, 3959, 2546, 2911, 4078, 4236, 3604, 3100, 4497, 1939, 1004, 994, 3291, 1866, 4382, 2143, 2467, 2663, 3218, 4634, 2799, 1539, 4671, 2570, 3882, 4106, 4105, 2936, 4110, 2261]. censoring [2144, 4426, 4226, 2805, 4469, 1809, 4039, 932, 3865, 3784, 3349, 4581, 3429, 1311, 4354, 4075, 4590, 1269, 962, 2372, 3020, 1828, 2724, 3397, 2454, 4612, 4313, 3143, 4351, 3435, 4672, 2630, 2770, 4677, 3916, 2515, 2924, 4482, 3972, 3454, 4662, 2539, 3642, 808, 883, 1110, 1339, 1302, 1804, 1815]. censorship [3055]. Census [122, 2258].

Center [4396]. centered [3007, 2704]. central [1450, 1263, 4585, 1151]. central-peripheral [4585]. centre [1161]. centred [3245]. Centres [1137]. ceramic [721]. cereal [864]. cerebral [2441]. certain [1456, 1596, 484, 460, 496]. cervical [4043, 2740, 4368]. cessation [2336]. Chaco [1247]. Chain [1409, 885, 4512, 2805, 3138, 3794, 843, 977, 3894, 1186, 3211, 1877, 947, 1763, 3612, 3707, 1255, 252, 423, 544, 556, 645, 675, 718, 709, 826, 3726, 3228, 3040, 3808]. chain-deferred [252]. Chains [1444, 4719, 2228, 3158, 3327, 4402, 3852]. challenges [1147]. Cham [4599]. ChAMP [3914]. championships [3543]. chance [3766, 251]. Chang [3572]. Change [3328, 1679, 978, 2390, 1461, 1647, 3157, 3256, 2968, 2117, 4670, 2240, 4340, 4434, 4645, 4532, 4284, 3488, 3470, 3883, 2287, 4228, 3137, 2759, 88, 4363, 2910, 3263, 3561, 2695, 2397, 3123, 3563, 2496, 1945, 2356, 2783, 1240, 2716, 1767, 2841, 2146, 3882, 3741, 1281, 4117, 1481, 4570]. Change-point [3328, 978, 2390, 1461, 2117, 4340, 4434, 4645, 3883, 3137, 3561, 3123, 3563, 2496, 2716, 1767, 3882, 3741]. change-points [2759, 4363, 2695, 2841]. changepoint [3633, 635, 3119]. changes [2235, 4564, 1290, 268, 1359, 2250, 2017, 915, 2015, 2163, 2959, 4265, 4537, 911, 2168, 4346]. Changing [161, 2711, 1197]. Channel [1517, 3618]. Chao [800]. character [2073, 3392]. Characterisation [4736, 2950]. characteristic [2853, 2126, 3192, 557, 4016]. characteristics [1456, 2262, 3274, 3089, 864, 1106, 443, 1908, 3163, 1515, 1112]. Characterization [621, 662, 489, 3716, 2368, 1565]. characterizations [4318]. characterize [1792, 4588]. Characterizing [1960, 2655]. characters [2528]. Chart [1451, 1450, 1397, 1654, 1405, 1683, 4364, 4098, 4523, 4647, 1936, 2874, 3134, 3357, 4039, 360, 1192, 1536, 1566, 1721, 4423, 2718, 2110, 2142, 1141, 3011, 4543, 2861, 2085, 3408, 2674, 2068, 4630, 2252, 4635, 1089, 2014, 3368, 781, 1765, 3742, 3467, 2291, 3073, 2351, 1569, 902, 1747, 3951, 1506, 4430, 2769, 2283, 1838, 1813, 4054, 4494, 4405, 4117, 1327, 3845, 858]. Charting [1556, 3355]. Charts [1399, 1426, 1684, 3235, 2380, 1320, 4234, 2100, 633, 1841, 1847, 2060, 1102, 1458, 1186, 1776, 320, 1149, 4284, 1000, 641, 1958, 1042, 317, 1088, 1720, 2134, 1548, 2580, 4033, 2645, 2309, 956, 2027, 2006, 4146, 2120, 1474, 1554, 2848]. Chatfield [78]. Chaumont [2824]. Chebyshev [3034]. check [4076]. Checking [3459]. chemical [4344, 3470, 317, 2169]. chemistry [604, 4216]. chemotherapy [1990]. Chen [4599, 4004, 4134, 1234]. Chen-G [4134]. Chengdu [3013]. Chi [1450, 1092, 865, 3891, 1252, 1306, 4208, 526]. Chi-square [1450, 865, 1306]. chi-squared [1092, 3891, 1252, 526]. chi-squared-based [4208]. chickpea [692]. child [3340, 3356]. Childhood [3332, 4611, 4419, 2232, 2047]. children [2304, 2113, 2387, 2862, 2795]. Chilean [2254]. China [1411, 2976, 4370, 4547, 3013]. Chinese [531, 3128, 2620, 2594, 1890, 543, 1955]. chloride [3433]. Choice [188, 691, 1452, 1125, 1272, 4655, 2083, 4021, 1760, 563, 4519, 2266, 2598, 3522, 4096, 517, 1750, 1100, 3013, 2884]. Cholesky [3925]. Cholesky-based [3925]. Choose [4694, 2481, 2605, 2013]. Choosing [798, 1294, 3292, 2154, 1876, 36, 1111, 515, 2759]. Chore [188]. chosen [1966]. chronic [1913, 3339]. Chronological [167]. ChSP [718, 544, 556, 645, 683, 675, 709]. ChSP-1 [675]. ChSP-4 [718, 683]. ChSP-4A [544, 556, 645, 709]. churn [3194]. CIDI [2604]. cigarette [2235, 3250]. circles [609, 375]. Circular [1301, 1663, 4187, 1415, 4744, 4587, 296, 378, 377, 2961, 4201, 2396, 1002, 3588, 2716, 1122, 2839, 3540, 614, 3261, 3026]. circularity [602]. circumference [2251]. cited [1491]. cities [1735, 3149]. citril [1229]. citrus [315, 441]. city [4201, 3013, 2702, 2969, 2246, 4332]. civil [4311]. claim [1602, 1349]. claims [4607]. clara [1208]. Class [1644, 801, 1386, 1482, 3086, 4191, 3997, 2971, 2568, 2361, 3234, 823, 2129, 3122, 2789, 3030, 1182, 2740, 960, 2672, 4506, 3092, 4357, 1895, 505, 1076, 1304, 3635, 3969, 4445, 4416, 4589, 4298, 1990, 3269, 1499, 4018, 322, 3278, 2209, 3076, 4450, 3233, 4661, 2902]. Class-size [801]. class-specific [3234]. classes [19, 4499, 1266, 2434]. Classical [3923, 1691, 1837, 3780, 3517, 4356, 3978, 3390, 1492, 4064, 4644, 3296, 298]. Classification [3085, 4544, 2888, 3537, 4455, 2408, 1430, 3436, 3815, 2170, 3035, 4495, 4687, 4246, 794, 515, 2484, 4487, 656, 1517, 1946, 667, 285, 2864, 1992, 2341, 3487, 2641, 4462, 3688, 4508, 2576, 3092, 4312, 1160, 4610, 2643, 3004, 3686, 2982, 3984, 3890, 3189, 2443, 1733, 4445, 3163, 672, 3866, 2792, 2352, 1045, 4480, 655, 4450, 3372, 3656, 4465, 4130]. classifications [599, 1584, 265]. classified [467]. classifier [3598, 2295, 2266, 3699]. classifiers [3526, 2967, 2599, 3592, 4024, 3946]. classify [1184]. Classifying [1675, 4649]. clays [3558]. Clayton [3341]. clear [3949]. clear-sky [3949]. clearance [719]. Climate [3126, 1748, 4191, 4532, 3103, 2613, 3976]. climatology [4219]. climbing [2716]. clinic [261]. Clinical [4693, 4730, 146, 3107, 2129, 2355, 989, 1115, 2492, 3259, 2457, 1950, 2314, 3644, 2746, 1034, 1519, 3434, 1118, 3326, 2084, 2598, 3677, 4333, 4379, 4099, 1759, 2209, 4522, 525, 1281, 4129]. Clipped [4130]. close [2834]. close-to-reality [2834]. closed [2958, 3898]. closed-form [2958]. clothes [2813]. Cloud [2842]. clouds [3682]. clover [43]. clumpiness [2730]. Cluster [3028, 3573, 119, 481, 3665, 3224, 3574, 1966, 2566, 694, 2184, 4669, 1928, 2762, 1297, 2306, 3087, 4112, 2264, 2569, 1316, 3222, 4680]. Cluster-based [3028]. cluster-randomized [4112]. cluster-specific [3574]. clustered [2304, 4441, 2553, 4530, 3551, 3344, 3301, 3217, 3574, 3798, 4515, 3009, 4329, 3843, 4203, 4322, 2738, 4299, 4337, 2467, 3992, 3078, 3222, 3619]. clustered/longitudinal [3992]. Clustering [2337, 1406, 3279, 3705, 2688, 2413, 2465, 3964, 2074, 4629, 3937, 3950, 3816, 4682, 4472, 3573, 4259, 3647, 656, 2586, 492, 3518, 2622, 1833, 2494, 3064, 3115, 4028, 2396, 4517, 4440, 3367, 4552, 3546, 3890, 3588, 3586, 3384, 2608, 1336, 4212, 4626, 3044, 2792, 3494, 3040, 4467, 2743, 2513, 3902, 4454, 3693]. clusters [811, 851, 2621, 3129, 3724, 4112, 376, 4374, 1738]. clusterwise [3810]. Co [3174, 2966, 524, 3303, 4615, 2198, 3112, 2526, 3145, 3506, 2246]. Co-authorship [3174]. co-dependence [4615]. co-directional [2198]. co-inertia [3145]. co-infection [3112]. co-median [3303]. co-operation [524]. co-operatives [2966]. co-persistence [2526]. CO-tucker [3506]. coach [1985]. coalitions [4360]. coarse [1372]. coast [3016]. coastal [1840, 773]. CoDa [4635, 4270]. Coefficient [3373, 1686, 1385, 1249, 2804, 4652, 3679, 4220, 39, 3491, 3447, 3894, 2882, 4316, 1829, 2736, 2187, 1153, 1165, 2973, 2433, 4232, 4489, 3130, 3461, 3438, 1858, 1046, 4210, 2437, 2132, 3538, 2160, 4337, 3607, 4202, 2539, 4336, 4146, 1274, 3462, 2854, 3560, 4586, 4539, 4613, 3198, 2571, 2698, 2975, 3651, 4447, 3378, 1253]. coefficients [1487, 3297, 2689, 2656, 2422, 486, 723, 959, 1374, 2530, 2934, 3826, 4322, 4623, 3370, 2450, 3822, 2999]. coexpression [3927]. cognition [1990]. Cognitive [3172, 1994, 4643, 2250, 2146]. Coherence [1712]. Coherent [3801, 4307]. Cohort [3340, 118, 1997, 4075, 4103, 4355, 3580]. coin [2314]. cointegrated [3083, 1952, 3084, 1109]. Cointegration [1411, 1384, 1507, 3128, 915, 1093, 1892, 1920, 3258, 1013, 2702]. Collaborative [2802]. collapse [4162]. collapsed [2657, 3018]. collecting [3613, 45]. Collection [154, 814, 235]. collections [864]. collective [1894, 3041]. college [4619, 2855, 2884]. Collinearity [3002, 3271, 4118]. Colombia [2496, 2401]. Colombian [4588]. colon [2801]. colonial [1235]. colony [3525]. Colorado [3487]. colorectal [3366, 966]. colored [2929]. column [1191, 498]. COM [2769]. Combination [1500, 1608, 4209, 1636, 1857, 2710, 1177, 4548]. combinational [4522]. combinations [806, 1149]. Combined [2849, 1086, 3483, 4100, 3773, 2134, 2080, 4355, 597, 875, 4380, 4095, 882, 2006]. Combined-penalized [2849]. Combining [417, 4663, 3154, 893, 1652, 4119, 3717, 4638, 2966, 2161, 2981, 280, 4608, 4585, 3385, 2847, 2209, 4420, 4401]. Cometary [171]. Comment [3083, 3082, 3200, 778, 3255, 3084, 2565]. Comments [165, 510, 1238, 562, 1203, 1225, 79, 1218, 3921, 1202, 80, 1215, 1207]. commerce [2728]. commercial [4162]. Commercials [1664]. Commissions [162]. committed [2866]. commodities [3309, 3437]. Common [1421, 828, 4126, 2787, 4598, 3404, 3325, 3867, 4610, 3763, 1457, 2626, 1205, 766, 3308, 4568, 2980, 1204]. Communicate [4720]. Communication [99, 1139, 712, 970, 3846, 1884]. Communications [1425]. communities [4169, 1184]. Community [3919, 3672, 1517, 4555]. commutative [4000]. commutativity [4000]. commuting [2342]. comorbidities [2844]. compact [1289, 450]. Companion [2323]. Comparative [4695, 3163, 655, 3549, 3950, 1912, 1873, 711, 2011, 2967, 3170, 636, 1374, 1753, 1174, 403, 1499, 3540, 2088, 3288, 4185, 3729]. compare [2008, 2196, 1928]. Compared [197]. Comparing [1443, 1644, 1519, 2773, 1380, 1717, 2183, 2662, 2208, 4541, 2533, 3086, 4594, 3380, 2607, 842, 851, 3624, 1021, 796, 804, 2386, 3209, 1575, 1562, 1194, 2217, 1257]. Comparison [900, 763, 1901, 693, 3420, 4167, 792, 746, 1659, 2432, 2281, 3390, 2133, 2300, 1768, 1634, 3434, 3443, 3428, 3946, 3087, 2315, 1628, 2188, 4150, 2104, 1936, 3079, 1964, 2709, 3081, 3455, 3671, 335, 541, 472, 2452, 492, 2595, 1797, 1170, 2519, 2731, 1460, 1840, 3793, 2864, 2268, 827, 4603, 3062, 787, 4240, 4203, 3952, 2431, 4322, 1774, 3905, 952, 3686, 2399, 4128, 1575, 1733, 4008, 2028, 3499, 1032, 3898, 2583, 1006, 866, 2836, 3296, 3070, 688, 2980, 526, 2115, 1484, 4366, 2303, 2625, 3656]. comparison-based [3070]. Comparisons [1387, 130, 2135, 4681, 3121, 1581, 4339, 2739, 1185, 3082, 1588, 4203, 2495, 2683, 4272, 4120, 1903, 2939, 1375, 282, 1051, 2728, 4107]. compartment [1285, 2962]. compartmental [2285, 1534, 3762, 2625]. compatibility [252]. compensation [2215]. competencies [3494]. Competing [1997, 4415, 3356, 4299, 3086, 4004, 4469, 3834, 4301, 2357, 4424, 4365, 3318, 3873, 3295, 3071, 3631, 3508, 2917, 4297, 4673, 3819, 1012, 3339, 3037, 4313, 3139, 3893, 1013, 3249, 4196, 3791, 4513]. Competition [1663, 3595, 1792, 1463]. competitions [304]. competitive [3362, 3886]. Competitors [1628, 1780]. complementary [2490, 3928, 3886, 4619]. complemented [530]. Complete [1338, 1386, 897, 1031, 944, 990, 842, 3062, 2644, 1293, 1939, 2413]. completion [849]. Complex [1883, 3645, 2602, 4157, 3330, 1361, 1465, 1462, 4419, 1148, 4073, 2907, 1004, 3973, 3932, 2213, 3238, 1146, 3547]. Complexities [4577]. Complexity [3441, 4529, 785]. compliance [4005, 910, 4559, 2402]. Component [4436, 3242, 3079, 3751, 23, 2697, 485, 343, 3517, 1979, 3446, 4424, 2560, 2725, 4036, 1791, 3023, 4215, 1976, 3975, 4580, 2286, 4293, 4524, 355, 4120, 1562, 1770, 4056, 2601, 3308, 3299, 2411, 3801, 3580, 4653, 3584, 4561, 4044]. component-amount [1770]. component-slope [1562]. Components [126, 1677, 3267, 828, 2103, 1364, 1525, 1752, 703, 948, 3530, 2367, 2586, 4329, 3315, 640, 691, 853, 873, 1976, 2396, 3944, 2811, 3110, 1071, 3411, 1993, 603, 834, 926, 605, 4258, 3442]. Composite [1387, 3569, 4329, 3734, 1073, 1976, 597, 1263, 4099, 1151, 2927, 3498, 2698, 2604]. composites [3433, 721, 1565]. composition [60, 1053, 2830]. compositional [485, 865, 3479, 3653, 3647, 4198, 4584, 2851, 3303, 3606, 2827, 2018, 3705, 2407, 3626, 3416, 3393, 1298, 3845, 4044]. Compound [1608, 4488, 4378, 1894, 4408, 2894, 2783, 3996]. compounding [4306, 2330, 4382]. compounds [669]. compression [2686]. Compromise [188, 1613]. Computation [2395, 4722, 1739, 4711, 2605, 3292, 2272, 4185]. Computational [1657, 4740, 1606, 3945, 1719, 2530, 2992, 2816]. Computations [1419, 4175]. computed [3847]. Computer [188, 92, 163, 143, 19, 1242, 4251, 809, 705, 2613, 1055, 3811]. computer-simulation [2613]. computerized [663]. Computing [4715, 4016, 774, 2506, 2609, 3563, 3895]. concave [1938, 3924]. concentrated [1578, 2520]. concentration [1525, 2706, 936, 4177, 4484, 2635, 3707]. concentration-time [2635]. concentrations [2246]. concept [3430, 4084]. concepts [445, 413, 676, 3577]. conceptual [3195]. concerning [807, 507]. concerns [1275]. Concise [4704, 2157]. concomitant [787, 2523]. concomitants [1155]. Concordance [2450, 2238, 2882, 4508, 3820]. concordance-type [3820]. concordant [1130]. Concurrent [2055, 3368, 3776, 4666]. condition [3915, 4240, 4000]. Conditional [3838, 1694, 1400, 1394, 3852, 3195, 1482, 3970, 4152, 1017, 1999, 3957, 2358, 999, 4321, 4151, 1000, 2976, 4594, 2561, 3678, 4192, 3071, 2008, 643, 980, 3922, 3681, 3196, 1916, 4056, 2983, 575, 1097, 2217, 3243, 1494, 2271, 698]. conditionally [2774]. Conditioning [4728, 592]. Conditions [1424, 573, 4581, 2538, 3814, 578, 4022, 3558]. conducting [4411]. Conference [121, 2622]. Confidence [2804, 211, 4499, 3301, 4404, 2333, 2957, 2753, 2414, 3879, 2151, 2207, 1511, 4380, 2707, 4093, 4115, 3800, 1557, 338, 3331, 2430, 2307, 1873, 1820, 2384, 1535, 927, 4596, 2129, 3804, 827, 972, 1878, 3465, 2735, 4507, 1490, 2469, 980, 3151, 1858, 2446, 4210, 3718, 2381, 2784, 3236, 4337, 1865, 2570, 2588, 2565, 3496, 2865, 3475, 4107, 4083]. Configural [1438]. confirmatory [4230]. conflicting [513]. confocal [2096]. Conformal [4665]. conformance [3089, 4115]. conforming [2718, 919]. Confounded [118, 906]. confounder [3807]. Confounding [4339, 4459]. Confusion [118]. conjoint [1717]. connected [2442]. Connecticut [1227]. Connecting [1620]. connections [652]. Connectivity [1712, 3765]. Consequence [173]. conservation [760, 759, 750]. Conservative [4210, 4337, 3167]. considerations [374, 3039, 3899, 1566, 3777, 3719, 3500, 2635, 3624, 2724, 1815]. Considering [4011, 1102, 2190, 1515, 3243]. Consistent [862, 900, 1345]. conspecific [1235]. constant [758, 3923, 4354, 2287]. constant-effort [758]. Constrained [3023, 3949, 4302, 2622, 4283, 926, 4341, 3812, 4619]. constraint [2285, 2114, 1904]. constraints [1193, 3667, 4188, 3329, 2486, 2731, 691, 4437, 1304, 2206, 893]. Construct [4694, 2013]. Constructing [102, 3765, 3773, 4290]. Construction [707, 697, 1270, 3735, 1248, 2597, 1263, 1963, 466, 503, 576, 1070, 2864, 2014, 943, 1543, 3131, 3718, 544, 556, 645, 709, 826, 3918, 575, 4311, 4107]. constructions [3949]. Consultancy [207]. Consultant [143, 2323, 520]. Consultation [97, 524]. consulting [514]. Consumer [1408, 4023, 4053]. consumers [901]. consumption [1956, 2082, 3027, 1913, 4532, 915, 901, 2864, 3597, 3858]. containing [3416]. contaminants [936, 2699]. contaminated [4149, 3783, 554, 3675, 1730, 4649, 4500]. content [39, 4113, 518, 2341, 3457]. Contents [822, 876, 877, 946, 1007, 770]. contest [2920]. Contestability [1704]. Context [168, 1583, 2479, 860, 2126, 424]. continent [758]. continent-wide [758]. Contingency [1395, 1445, 1394, 1482, 2715, 4211, 3844, 341, 3652, 1549, 2179, 2041, 3938, 1943, 3670, 3781, 2979, 909, 286, 2657, 3018]. contingent [2813]. continuation [1053]. continuation-ratio [1053]. Continuity [4596]. Continuous [1408, 1388, 3408, 719, 1708, 1403, 2585, 4238, 2988, 1029, 1059, 1545, 3138, 3685, 1280, 3551, 3794, 1273, 3217, 2097, 4545, 3678, 913, 1086, 1094, 3945, 1261, 4309, 1041, 617, 681, 3981, 3354, 4090, 3727, 3587, 3273, 3489, 2177, 4614, 4136, 869, 4071, 3116, 2786, 861, 4181, 4034, 1077, 2170, 3729]. continuous-time [3138, 3354, 4034]. contour [4421]. contoured [2132]. contours [3599]. contraceptive [903]. contraceptive-use [903]. Contracts [1729, 3663, 1859]. contrast [1156]. Contrasting [2563]. Contrasts [195, 1632]. contribution [88, 2727]. Contributions [928, 1459, 1238, 1309, 1237, 3781]. Control [1451, 1443, 1399, 1426, 360, 1137, 2060, 1149, 1397, 1654, 2580, 1883, 2769, 3641, 4720, 1683, 3810, 4364, 3235, 2380, 1320, 4098, 4523, 4647, 1936, 4681, 4234, 2874, 2100, 3134, 3357, 4039, 3355, 3694, 106, 1731, 1841, 4423, 1847, 1458, 3821, 1776, 2110, 1142, 1141, 1581, 1905, 1310, 989, 1755, 3011, 4543, 1000, 2861, 2085, 3408, 641, 2674, 585, 1572, 1958, 2068, 317, 4630, 967, 3082, 3426, 1145, 2252, 1872, 4635, 1088, 2615, 1089, 842, 2014, 4371, 781, 2010, 1720, 2375, 919, 1765, 2236, 2495, 3742, 3467, 2134, 796, 804, 2080, 2291, 673, 2351, 441, 2683, 1569, 902, 2837, 1747]. control [1365, 3951, 4033, 2645, 1506, 1575, 4326, 1194, 2567, 2309, 956, 2283, 4394, 1838, 2027, 1903, 1579, 2006, 4146, 4569, 1058, 2482, 2120, 4054, 4405, 4117, 1327, 1554, 3845, 858, 2848]. control-chart [1765]. controlled [3446]. Controlling [4007, 1398, 4388, 2147, 2057, 4429, 4055]. conventional [320, 2173]. converge [2474]. convergence [1244, 2474]. converted [1606]. convex [3924, 4158]. convolution [3136]. convolutional [3688, 4585]. Conway [4364, 3574]. Cooch [1202]. Cook [1915]. Coordinate [2445]. Coordinate-free [2445]. coordinates [3424]. coordination [2716]. cope [3807, 1174]. coplot [4411]. coprotheres [737]. Copula [4072, 3957, 4367, 3480, 4287, 1877, 2976, 2758, 3843, 3152, 3341, 3095, 2050, 4501, 4186, 3218, 2265, 4265, 2680, 4453, 4557]. copula-APARCH [4287]. Copula-based [4072, 3480, 1877, 3152, 3341, 3095, 4265]. copulae [1897]. copulas [3663, 1486, 3437, 1799, 3101]. copy [2240, 2783, 3068, 2168]. Corinth [3760]. corn [4090, 3498, 3669]. corner [38]. coronary [1346, 3605]. Corporate [3306, 1962, 2296]. corpus [3888, 4499]. corpus-based [4499]. correct [3942, 1584]. corrected [4388, 4596, 2210, 636, 846, 1062, 3560, 2338]. Correction [3142, 3931, 3943, 4206, 4100, 4288, 4289, 4205, 1624, 3205, 320, 3672, 3490, 2599, 2020, 3343, 2836, 1167]. Corrections [1399, 156]. Correlated [3402, 4343, 4090, 1696, 1647, 1656, 1703, 3331, 2304, 4234, 2091, 4296, 2307, 3564, 2313, 4077, 4638, 2715, 4347, 3045, 4598, 4089, 3112, 1466, 3315, 3458, 1153, 1165, 791, 1534, 1761, 1306, 1885, 1046, 3776, 3754, 4292, 3839, 2219, 685, 2668, 2403, 4604, 2764, 2063, 3610, 907, 356, 3721, 4071, 1856, 4107, 1564, 4129]. correlating [2422]. Correlation [1634, 1385, 499, 3634, 471, 2564, 1250, 39, 1792, 1979, 3471, 3447, 2238, 4301, 4384, 3297, 1913, 2011, 4674, 2627, 1470, 1892, 3739, 1052, 3109, 1284, 919, 1374, 4668, 2211, 2934, 3056, 670, 4322, 4249, 3428, 1772, 4210, 3681, 3411, 857, 1998, 4337, 558, 4138, 4227, 2340, 3078, 2109, 4613, 2331, 2999, 1253]. correlations [3142, 703, 3309, 2776, 3562, 4567, 3196, 4379, 3968, 2271]. Correspondence [1784, 2399, 485, 1930, 2487, 616, 2087, 2671, 4221, 519]. corridor [3797]. Corrigenda [368]. Corrigendum [2039, 2091, 2541, 2685, 2859, 3005, 3324, 3347, 3712, 3736]. corrosion [818]. cosine [4038]. Cost [2302, 91, 3784, 717, 1273, 1309, 2033, 2689, 3543, 2486, 2731, 2808, 3366, 2724, 3221, 1904, 3346, 2150, 1815, 3387, 3900]. cost-effectively-sampled-data [2150]. Cost-efficient [2302]. cost-sensitive [3900]. Costs [1228, 1204, 2564, 2879, 2212, 358, 1515]. could [3621]. counselling [3383]. Count [1605, 4738, 4072, 4441, 4579, 4038, 2718, 3574, 767, 2766, 2822, 4356, 2298, 3009, 4089, 3050, 2529, 4565, 3747, 4240, 3737, 4203, 1601, 3884, 3451, 1808, 4164, 2050, 4555, 3293, 3287, 2668, 1188, 1821, 2381, 4196, 4142, 3992, 3761, 2056, 3409, 4307, 3501, 2812, 2938, 1738, 2480, 4548, 1933, 4571, 4385]. counties [4552]. Counting [3718, 3564, 2000, 4135, 4303]. countries [2968, 2099, 3526, 1370, 346, 619, 4428, 4679, 2474]. country [4245, 2491]. counts [4234, 1103, 3045, 641, 2360, 3591, 3505, 1745, 3265, 2584, 2075, 2259, 4502, 2109, 1164, 2692, 3222, 3353, 3132]. coupled [2647]. couples [3848]. Course [4723, 195, 163, 4745, 3157, 2337, 157, 20, 2821, 1075, 4123, 2651]. Courses [188, 39, 340, 1971, 2551]. Cov [4542]. Covariance [4694, 133, 2443, 1632, 3816, 2110, 3904, 3711, 4280, 3925, 3280, 2234, 2522, 4017, 2843, 3175, 1991, 1119, 1343, 3964, 3047, 2154, 1740, 2013, 727, 3459, 3565, 1051, 2169, 2932, 4561, 1468, 2065]. covariance-based [2522]. covariances [940]. Covariate [1699, 1696, 2091, 3336, 1209, 1896, 2298, 3065, 2312, 2629, 4335, 3398, 2400, 3527, 3601, 2398, 2989, 2511, 3807, 2982, 2937, 1234]. covariate-adjusted [4335, 3527]. covariate-dependent [3065]. Covariates [146, 1632, 1648, 997, 3260, 2881, 2113, 4246, 3786, 3665, 3479, 3653, 4198, 4267, 4638, 680, 2926, 4435, 3630, 4594, 3038, 585, 1572, 2833, 3014, 3500, 1950, 3514, 2372, 3020, 2054, 3819, 4520, 2796, 3484, 4552, 3922, 4277, 2389, 4391, 3754, 3684, 1208, 4073, 4482, 3620, 3796, 3953, 2854, 4448, 1995, 1564, 3730, 2571, 4628, 2261]. Cover [78]. coverage [3138, 2620, 3680, 3496]. covid [4521, 4538, 4536, 4687, 4650, 4433, 4544, 4553, 4546, 4545, 4534, 4503, 4565, 4683, 4274, 4654, 4552, 4510, 4445, 4589, 4535, 4618, 4528, 4540, 4533, 4537, 4547, 4491, 4502, 4541, 4539, 4548, 4551, 4500, 4549, 4543, 4631, 4512]. covid-19 [4521, 4538, 4536, 4687, 4650, 4433, 4544, 4553, 4545, 4534, 4503, 4565, 4683, 4274, 4654, 4552, 4510, 4445, 4589, 4535, 4618, 4528, 4540, 4533, 4537, 4547, 4491, 4502, 4541, 4539, 4548, 4551, 4500, 4549, 4543, 4631, 4512]. COVID-19’s [4546]. COWORDS [3682]. Cox [2034, 4638, 1294, 4478, 3038, 4195, 2492, 3661, 2722, 2634, 3807, 3276, 461, 283, 2666, 1562, 4526, 3953, 4106, 2525, 4319, 3605, 4628]. CR [747]. crab [2136]. crack [2019]. Cramér [1265, 2178]. crashes [3294]. Crazy [224]. cRCT [2264].

Creating [2528, 1462, 1138]. creation [3598]. credibility [4174]. Credit [2194, 2393, 4725, 4691, 1869, 4176, 1797, 1070, 1373, 3900, 3832]. Cricket [185, 3929, 3623, 2726]. crime [2141, 4291, 2702]. criminal [4308]. crises [3906]. crisis [4162, 4540]. Criteria [606, 1454, 1056, 4007, 2249, 2481, 4620, 1314, 930]. criterion [900, 4243, 1083, 4514, 4508, 3952, 2421, 2178]. Critical [825, 3577, 1687, 3385, 528, 588, 1082, 35, 51, 265, 4022, 399, 777]. Criticality [4736, 2950]. critique [1178]. Cronbach [3307]. crop [470, 2026, 2603, 3639, 2739, 2105, 1859, 4186, 3375, 1075, 2477]. cropping [869]. crops [43]. Cross [175, 2422, 1892, 2316, 1386, 1975, 1250, 3128, 2993, 1910, 2033, 1031, 3309, 3543, 2844, 4567, 1293, 1983, 3995, 2206, 1874, 1816, 467, 4133]. Cross-correlating [2422]. cross-correlation [1250]. cross-correlations [3309, 4567]. cross-dependence [2033]. cross-listed [3128]. cross-platform [1910]. cross-product [1975]. Cross-sectional [1892, 2844, 1983, 2206, 1874]. cross-sectionally [1816]. Cross-validating [2316]. cross-validation [3995, 4133]. crossed [4025]. crosses [468, 1581, 944, 990, 1085, 1338, 960, 280, 1248]. crossing [1293]. Crossover [3107, 223, 780, 3056, 3617, 1352, 976, 2867, 2402]. crowd [4303]. crowdsourced [3777]. crude [2742, 4260]. crystalline [417, 3449]. CSP [1408, 1545, 1083, 1094, 681, 963, 861, 935]. CSP-1 [1083, 861, 935]. CSP-C [1408, 1545, 1094]. CSP-M [963]. CSP-V [681]. cube [262]. cubes [405]. Cubic [716, 4431, 590, 1766, 1770]. cultivar [3702]. culture [2223]. cultures [4649]. Cumulative [1930, 1572, 2718, 788, 4593, 1537, 1864, 4232, 2111, 3916, 4105]. Cup [3181, 2664]. cure [2687, 2102, 3786, 4415, 2098, 4331, 4143, 2294, 3518, 3417, 2054, 2796, 2412, 2390, 2202, 2219, 3853, 2584, 3649, 4350, 3882, 1831, 2047, 3220, 3730, 3729, 4368]. cured [4378, 3731]. currency [465, 1024]. Current [81, 4297, 4673, 1689, 3987, 3798, 4515, 4627, 2786, 3713, 4513]. curricula [359]. Curtailed [220, 579]. curvature [4665, 3019]. Curve [4470, 1427, 2853, 413, 2333, 2367, 1001, 1875, 3750, 3374, 1054, 1129, 1074, 2067, 2885, 2483, 835, 1766, 1343, 3148, 345, 674, 3076, 4480, 4491, 3238, 2415, 2681, 1077]. curve* [322]. curves [4505, 1333, 4516, 1837, 2647, 4259, 2452, 805, 4594, 571, 3282, 600, 1466, 2635, 1538, 2126, 4204, 782, 2196, 3890, 487, 595, 3209, 1047, 930, 3192, 4540, 3796, 557, 621]. Cuscore [551]. cusp [4175]. Custom [3740]. Customer [4733, 2245, 2903, 3279, 3194, 2818, 1919, 4463, 4494]. cusum [2874, 2466, 3235, 1192, 4340, 4284, 641, 4630, 4565, 3368, 3467, 3073, 1548, 2351, 2075, 1813, 4494, 1586, 1042, 3426, 1150]. CUSUMs [1280]. cut [2879, 4508, 3221]. cut-off [2879, 3221]. cut-points [4508]. cutaneous [1948]. cutoff [2605]. CV [3696]. CVIF [2210]. cyber [3495, 4452]. cybersecurity [3683]. cyberspace [4251]. Cycle [1614, 2035, 678, 1173, 701, 4608, 995]. cycles [4538, 1553, 1569]. Cyclic [142, 589]. cyclical [1271]. cyclo [3543]. cyclo-cross [3543]. cyclones [678]. Cylindrical [1407, 3553]. cylindrical-type [3553].

D [2091, 4128, 3850]. Dagum [3753, 2318]. Daily [1444, 3430, 1855, 1920, 4201, 3481, 2306, 2075, 4548]. damage [3016, 3449]. damaged [1786, 1565]. dangers [1151]. Danish [1604]. Darling [1567]. Dasatinib [3315]. dasometric [1599]. Data [4739, 4710, 1444, 1050, 1467, 103, 152, 213, 1700, 1631, 1710, 3667, 4697, 3207, 154, 4699, 1675, 1644, 1645, 87, 171, 2532, 1301, 303, 4749, 1605, 128, 260, 4101, 119, 1391, 781, 4519, 4715, 1160, 1436, 1698, 4706, 4743, 1430, 1617, 2325, 2322, 2578, 1689, 2611, 2077, 141, 1453, 1696, 196, 1432, 1697, 4707, 605, 4738, 4547, 225, 4746, 1197, 212, 1688, 4443, 1420, 4702, 1039, 3086, 4640, 3260, 2304, 3213, 4252, 4367, 3440, 4004, 3703, 2292, 4156, 2258, 2733, 499, 253, 3575, 3616, 1330, 1793, 4441, 4163, 3816, 2903, 2337, 1291, 3917, 748, 4013, 3655, 2479, 4549, 3837]. data [4257, 497, 687, 2896, 4047, 760, 3751, 4191, 4687, 1289, 1221, 3441, 2384, 2239, 4246, 485, 929, 1124, 2616, 2145, 2387, 3556, 1752, 3350, 4530, 3595, 4432, 532, 4388, 2237, 3980, 4233, 2243, 4043, 2424, 4375, 1796, 865, 3254, 3027, 3279, 2329, 1212, 4579, 4119, 1217, 3344, 3349, 3828, 4643, 3836, 3767, 2098, 3194, 2686, 4199, 3982, 1372, 2021, 749, 778, 1721, 1060, 1523, 2347, 3301, 4596, 2667, 1595, 1782, 2240, 2313, 2900, 3122, 3294, 3532, 3647, 3728, 3859, 4301, 4267, 4461, 2659, 3057, 4458, 4038, 3613, 2912, 3574, 1102, 1458, 3182, 3203, 391, 3648, 4584, 3949, 2476, 1937]. data [2003, 1294, 4657, 2713, 2186, 2802, 4418, 4049, 1930, 4460, 1718, 3872, 3735, 3049, 2926, 767, 1010, 2284, 1972, 2097, 4347, 3582, 1863, 3606, 2822, 2332, 2964, 3117, 2346, 1241, 2967, 3722, 4365, 4509, 2829, 1517, 4321, 112, 4356, 4532, 1784, 4334, 3160, 2725, 4411, 440, 1877, 492, 671, 4168, 2298, 3873, 3987, 4413, 3818, 2922, 2705, 3913, 3045, 3798, 3875, 4515, 3488, 2538, 3470, 2069, 1578, 2018, 3578, 3009, 4329, 1900, 3417, 4132, 4625, 2153, 2656, 4089, 1530, 2288, 3104, 2279, 3066, 3863, 3705, 4148, 3581, 950, 2627, 3473, 2094, 332, 3014, 4149, 3477, 931, 3871, 4081]. data [2870, 1312, 3843, 3295, 409, 3557, 3390, 3877, 4241, 3380, 3259, 3759, 4592, 1914, 3330, 296, 378, 4082, 2360, 3591, 377, 3930, 2961, 3426, 58, 3114, 1778, 744, 880, 1239, 2528, 2529, 733, 2374, 2537, 21, 2640, 1020, 518, 2990, 3359, 925, 559, 564, 3071, 3631, 4565, 2635, 947, 1369, 2711, 4363, 962, 1251, 3064, 3115, 1538, 1909, 3366, 3831, 3747, 3458, 4593, 1033, 1754, 1992, 2341, 3020, 3750, 4663, 3327, 3457, 4660, 1218, 490, 1518, 3782, 373, 3246, 894, 2226, 2364, 3878, 4455, 250, 3062, 3109, 3633, 4479, 4240, 810, 2576, 2875, 3508, 3529, 4069, 4312, 2522, 3737, 3775, 2917, 3374, 682]. data [1720, 4297, 4673, 3365, 3674, 3102, 4572, 4373, 4203, 951, 3418, 4477, 2644, 2396, 1117, 4283, 4668, 2767, 791, 1334, 1601, 2214, 3096, 3004, 3819, 4408, 4493, 3884, 342, 2236, 2433, 2594, 3770, 4232, 4470, 4489, 4642, 3536, 4679, 2597, 3130, 3461, 3601, 2398, 2916, 3012, 3438, 4104, 4490, 2466, 3281, 2796, 3125, 673, 1988, 3339, 3326, 3484, 2803, 3617, 670, 1002, 1844, 3037, 4627, 3376, 2518, 3860, 4651, 2695, 2612, 3603, 2511, 3654, 3430, 2738, 4024, 2405, 4221, 370, 4558, 2465, 3451, 2550, 1907, 3743, 3351, 4355, 304, 4563, 2474, 2592, 1961, 3361, 3893, 4669, 1808, 4607, 4471, 3588, 4299, 4173]. data [3522, 3723, 1574, 2819, 3342, 4474, 701, 1902, 4277, 4164, 2509, 1848, 4361, 1119, 4210, 4314, 3823, 4606, 3544, 2823, 1237, 2225, 2949, 2050, 1766, 1528, 4419, 3123, 3792, 2610, 1132, 3681, 448, 316, 3411, 4025, 3559, 2892, 3478, 1026, 2512, 1993, 3321, 3964, 2289, 4254, 2260, 297, 318, 3227, 3489, 2202, 4120, 4084, 3725, 3579, 2546, 2911, 4078, 1222, 2183, 613, 1175, 4659, 3120, 3802, 4216, 4414, 2061, 3675, 3772, 4278, 3047, 4236, 3756, 3839, 4380, 4555, 3140, 3604, 4338, 2963, 4589, 4676, 2044, 4417, 3100, 3240, 3287, 3691, 1133, 2668, 1025, 1497, 3726, 4597, 2584]. data [4014, 3237, 1188, 4618, 3269, 1744, 2403, 3972, 1821, 2381, 2784, 4497, 2907, 4604, 4632, 1919, 2947, 2764, 4162, 3268, 2310, 2504, 3228, 2028, 2929, 1947, 4626, 1004, 994, 3973, 4217, 3731, 2716, 1767, 3046, 2662, 2038, 3044, 4382, 4337, 2456, 837, 4018, 2270, 2143, 2467, 2663, 2978, 3218, 3610, 4394, 904, 3607, 1816, 3938, 2780, 3614, 3126, 907, 4093, 1943, 2928, 356, 4141, 2539, 3924, 2352, 4600, 4637, 1045, 4330, 4196, 540, 1150, 3464, 4533, 3642, 3791, 3494, 3076, 4599, 4071, 410, 1097, 3992, 3761, 2167, 2839, 3540, 655, 4091, 3393, 3932, 2146, 1360, 2786, 1298, 2056, 591]. data [1539, 2470, 496, 586, 550, 706, 2340, 2638, 4569, 1114, 1110, 3078, 3895, 1918, 2071, 614, 2164, 4050, 4219, 4639, 4467, 2197, 2150, 2694, 2639, 3963, 1058, 2223, 4671, 2719, 4320, 2264, 773, 1723, 2411, 2570, 2720, 2854, 3409, 3453, 3549, 3553, 3897, 3882, 4106, 4105, 4541, 4439, 699, 4688, 3808, 663, 467, 526, 3714, 3501, 4685, 2169, 2569, 2865, 4123, 4054, 3091, 1995, 3288, 4449, 2932, 2936, 3713, 4605, 4513, 3940, 3222, 4689, 4107, 966, 1349, 3580, 4653, 2441, 3868, 3547, 2303, 4561, 3845, 1593, 2812, 2938, 4570, 858, 921, 1468, 1498, 1738, 2354, 2730, 2832, 3900, 4124, 4372, 4253, 4453, 4500, 4444, 3132, 2975]. data [4291, 4571, 4447, 3463, 4110, 3074, 3602, 3201, 1077, 4385, 4465, 3832, 2170, 2502, 4624, 2261, 3052, 740, 4527, 1238, 2091, 2064, 1987, 2090, 2951]. Data-analytic [781, 1517]. Data-based [1160]. Data-driven [3207, 4519, 4547, 2387, 3228, 2411]. data-mining [3126]. Database [2532, 4368]. databases [3824]. dataset [3598, 3807]. datasets [4472, 2346, 3978, 2672, 3140]. date [263]. dates [4187]. dating [2547, 4265, 4452]. Davidson [2709]. Day [3929, 83, 1727, 4528, 2726]. de-duplication [4600]. DEA [2384, 2005]. deal [2468]. Dealing [3380, 4419, 3140, 4597]. Death [1574, 2548, 2113, 2553, 301, 4510, 4022, 4548]. deaths [4534, 1735, 4528, 3054]. debate [4492]. debiased [4416]. deblurring [665]. December [92]. decentralization [3360]. Decision [999, 4731, 162, 1601, 372, 515, 1731, 2815, 4325, 1042, 3557, 3513, 3206, 2837, 786, 4626, 4568, 4390, 4521]. Decision-Making [162, 3206, 2837, 4568].

Decisions

[4691, 3232, 1235, 3164, 2114, 2611, 726]. decks [891]. decline [756, 2146]. Decompose [1395]. Decomposing [2727, 2562]. Decomposition [1691, 151, 1367, 4249, 4659, 2905, 2657, 4291, 4571]. deconvolution [3203, 2349]. DECORATE [2385]. decreasing [3520, 4184, 4225]. dedifferentiation [843]. deep [3688, 4445, 4452]. Default [1843, 4023, 4345]. defaults [2637]. Defective [297, 3731, 2102, 1192, 1536, 3836, 1310, 462]. defects [357, 3449]. deferred [506, 637, 252, 4560, 412, 503]. deficient [4249]. deficit/hyperactivity [2638]. deficits [1258]. defined [3983, 4374]. Defining [1638, 1070]. definition [2627, 1831]. deflated [2914]. Degradation [2538, 3903, 4052, 642, 2045, 2116, 4341, 4616, 3052]. Degree [3157, 135, 196, 1362, 15, 550, 706]. Degrees [1199]. Delaunay [3036]. delay [236, 2377, 3091]. delayed [1227]. Deletion [1852, 1347, 3105, 1688]. Delicious [2452]. Delineating [765]. delineation [3035]. delivery [3468]. delta [3524, 2865]. delta-lognormal [2865]. Demand [1411, 154, 3420, 2721, 3225, 283, 2555, 3602]. dementia [3313]. demerit [3742]. democracy [4501]. demographic [3655, 1238, 3586, 1237]. demography [756, 2987]. demon [917]. demonstration [1781, 2016]. dengue [3910, 4025, 2478, 3136]. Densities [1650, 2972]. Density [1453, 470, 1221, 4059, 4684, 3104, 4149, 418, 2427, 2493, 3245, 3546, 2349, 3384, 3343, 2055, 1814, 4212, 4626, 4062, 2514, 2638, 2417, 2311, 4399, 3074]. density-based [3343, 2638]. dental [2304]. department [3689]. departmental [4270]. Departure [1886, 2348, 909, 3018]. Departures [3158]. Dependence [320, 3437, 1653, 1221, 4246, 3199, 2603, 3957, 2035, 3113, 2033, 4615, 1486, 1571, 4157, 2298, 799, 1362, 3152, 2629, 3514, 2933, 3000, 1054, 1129, 2256, 4567, 3095, 1173, 701, 1366, 2203, 3934, 1515, 1943, 4265, 3053, 1468, 4193]. dependencies [2727]. dependency [4282, 739, 467]. Dependent [1608, 1439, 2988, 3349, 4438, 4581, 3429, 596, 4340, 4075, 4365, 3408, 637, 3065, 3945, 4593, 4507, 4673, 3838, 1064, 3339, 4627, 1957, 1772, 4471, 3951, 461, 4391, 1908, 4482, 1816, 4135, 3796, 503, 2470, 2340, 4429, 2533, 4628]. dependently [4105]. deposit [4171]. depression [3714, 2832]. depressive [2604]. Deprivation [1895]. Depth [3887, 3356, 3594, 3857, 3592, 2354]. Derivation [792]. Derivatives [4725, 2393]. Derived [213, 285, 2917, 2982, 893, 1077]. deriving [600]. descent [3812]. describe [3441, 4341]. Describing [1664, 4282]. Descriptive [168, 3874, 3900]. descriptor [2117]. deserving [22]. Design [2032, 4695, 1426, 3865, 4696, 1083, 2754, 2131, 116, 195, 1613, 1094, 1264, 2457, 3624, 4743, 1670, 1412, 854, 161, 153, 148, 2088, 905, 1099, 1320, 4344, 1583, 3671, 3107, 4530, 1235, 934, 3039, 1841, 1566, 1273, 3252, 1102, 1725, 1186, 20, 1896, 3244, 641, 585, 1572, 906, 3034, 772, 1144, 3259, 2909, 2004, 1818, 2314, 721, 842, 340, 1563, 2297, 3644, 521, 3434, 1161, 2116, 3617, 3371, 952, 2894, 2819, 2860, 2645, 795, 4183, 1352, 881, 1908, 3839, 675, 3851, 3680, 3240, 1881, 4186, 1098, 2283, 547, 1532, 4616, 2078, 777, 444, 2213, 1600, 4374, 394, 4088, 2863, 2461]. design [1554, 1481, 1257, 1248, 2088]. design-based [2860]. design-some [394]. Designed [1713, 1117, 1569, 1148]. Designing [3433, 1917, 1587, 3106, 1543, 3951, 683, 984, 2867, 969, 4263, 412]. Designs [1443, 223, 3613, 1404, 140, 1441, 1626, 1663, 102, 1609, 142, 1386, 1412, 4109, 468, 856, 2732, 1736, 1822, 2574, 1356, 1168, 3903, 1289, 631, 2176, 1191, 404, 577, 546, 1193, 948, 1156, 4487, 2471, 1887, 2439, 1888, 3094, 1479, 1031, 1896, 3298, 3773, 2813, 1357, 2925, 4001, 1085, 1182, 1338, 3330, 3793, 4276, 2808, 964, 590, 498, 708, 720, 1768, 725, 853, 3379, 1021, 4103, 1534, 1761, 2302, 2598, 943, 1928, 4057, 940, 1071, 1904, 580, 1263, 3489, 3146, 4492, 603, 834, 926, 1770, 4597, 4295, 857, 2154, 1108, 1350, 4145, 3589, 540, 1151, 2209]. designs [411, 1463, 3188, 3133, 1849, 4522, 529, 3779, 2606, 2696, 3864, 714, 715, 2104, 1560]. desirability [2370, 3207]. Detect [1645, 182, 3369, 3303, 2085, 3738, 2466, 1569, 3588, 527, 4623, 869]. Detecting [1736, 1782, 2678, 2287, 3567, 1318, 3129, 4221, 334, 3148, 1647, 1879, 2646, 2120, 1100, 3868, 1994, 1775, 4259, 1536, 1901, 4284, 3803, 530, 2759, 2990, 3191, 2291, 3073, 4518, 795, 2783, 2259, 2482, 1813]. Detection [4564, 1710, 3880, 2015, 1434, 406, 3539, 2695, 3608, 141, 779, 4338, 1620, 1412, 2841, 3741, 858, 3703, 3504, 1553, 4665, 3573, 2117, 2257, 631, 2521, 3709, 4432, 2835, 4529, 1305, 2568, 4159, 1966, 3080, 4434, 4645, 3490, 4182, 4451, 2162, 3470, 606, 4228, 495, 4565, 1285, 1852, 4045, 3568, 2886, 1325, 4470, 3474, 1590, 1492, 3123, 3119, 3724, 2345, 376, 4158, 3393, 3932, 1876, 2197, 4420, 4541, 699, 2169, 2449, 3068, 4218, 4390, 4570, 2806, 4291, 4571]. determinant [2169, 4561]. Determinants [4108, 3168, 3304, 2780]. Determination [871, 1017, 3038, 578, 1041, 4560, 4652, 573, 633, 3491, 937, 1743, 3644, 2530, 2469, 3617, 1978, 2172, 2437, 1579, 2062]. determine [2853, 4177]. determined [3537]. Determining [4171, 1423, 3311, 3204, 717, 3912, 3419, 3259, 3163, 1515, 3531]. Deterministic [4005, 413]. detrending [1551]. Developed [1430]. Developing [186, 2531, 1905, 4102, 3537, 346, 619]. Development [3555, 816, 160, 4381, 3022, 4536, 4318, 3517, 1854, 2686, 378, 1556]. developmental [1538, 1764]. Developments [1436, 148, 58, 4026]. deviance [2667]. deviation [1535, 1542, 2374, 2291, 4585, 3053]. deviations [2466]. device [2521]. devices [3344]. DEWMA [4098, 4033]. Dhar [3260]. diabetes [3514, 2251, 3699]. diabetic [3247, 2405, 3296]. diagnosed [2604]. diagnosis [522, 1795, 3172, 2888, 4280, 2604, 2390, 4445, 2377, 4158, 3055, 1889]. Diagnostic [3666, 1662, 2216, 2604, 1380, 1686, 1620, 2761, 2873, 562, 1305, 2101, 2161, 3523, 1867, 2247, 3283, 2673, 2740, 4508, 1976, 2431, 4654, 1861, 4209, 1853, 3787, 4271, 2183, 2030, 3148, 3271, 4614, 3192, 2217, 4491, 2836, 2918, 2801, 3262, 1257]. diagnostic-robust [1867]. Diagnostics [1796, 1442, 3197, 1449, 3333, 3849, 3534, 3957, 3187, 2547, 3962, 1286, 1728, 3042, 330, 3986, 3105, 2159, 3130, 3363, 3265, 3386, 2132, 2403, 2371, 2856, 3220, 3693]. diagonal [4017]. Dialled [1386, 468]. diallel [1581, 1031, 944, 990, 1085, 1338, 960, 280, 1293]. diameter [817]. dichotomised [4676]. dichotomized [3434]. Dichotomous [1414, 3996, 1452, 2903, 4491, 3492]. Dickey [1551]. diesel [1956]. diet [809, 3416]. dietary [4252, 4282]. differ [1992, 1012, 3807]. Difference [3472, 3284, 1394, 2430, 393, 2171, 3301, 4596, 3532, 2607, 15, 2735, 2126, 1864, 2151, 4464, 3817, 2478, 3610, 1865, 581, 2588, 2480]. Difference-based [3472]. Difference-in-difference [3284]. difference-sign [2171]. differences [3942, 4259, 3250, 1799, 4230, 1589, 1465, 3839, 4107]. Different [1652, 2572, 3079, 2479, 763, 2667, 2602, 3804, 4545, 3690, 2519, 3795, 2008, 3661, 661, 4017, 1853, 2580, 3428, 4607, 4474, 3910, 3544, 3802, 3853, 3533, 3312, 2028, 3093, 4547, 4219, 1586, 4150, 2385]. Differential [3771, 3919, 2568, 4483, 1893, 2990, 1950, 2888, 2944, 2857, 4680]. differentially [4293]. differentials [3167]. differing [1229]. difficult [761]. Difficulty [135]. diffraction [4221]. diffusion [3566, 3945, 2615, 4585, 2927, 3216]. Digit [1645, 2768]. Digital [451, 453, 364, 3850]. digits [852, 2295]. Dihedral [3097]. dilated [1116]. Dimension [4403, 3703, 4267, 1527, 3380, 3567, 2703, 3733]. Dimension-wise [4403]. Dimensional [208, 2977, 2096, 4495, 4246, 3647, 4301, 4267, 4461, 4460, 2432, 4435, 4413, 3759, 3103, 2537, 3750, 3824, 3925, 3778, 951, 1794, 3096, 4489, 4024, 782, 3905, 1445, 4210, 2443, 3754, 3227, 4403, 4338, 4676, 4604, 2504, 4626, 3044, 3607, 3444, 2928, 4410, 3171, 2786, 4569, 3208, 3453, 4420, 4439, 371, 4319, 4586, 4133, 4561, 3198, 4124, 4444]. dimensionality [2686, 4485]. Ding [4599]. Ding-Geng [4599]. Dinosaur [2444]. dioxide [3771, 916]. dipper [1220]. Direct [918, 304, 305, 1238, 3471, 1237, 2905]. Directed [1712]. Direction [1407, 4010]. Directional [4280, 370, 1022, 497, 687, 4625, 2198, 1500, 373, 1002, 1528, 3579, 1767]. directions [759, 4566, 714, 715]. Directive [3830]. Dirichlet [1953, 1312, 3626, 2972, 2513, 2475, 4360]. disability [2258, 3448]. Disaggregate [2280, 2603]. Disaggregate-level [2280, 2603]. disaggregation [3027, 2077, 2206, 2919, 2152]. disaster [3206]. discharge [3986]. disciplinary [3173]. disciplines [10]. disclosure [3529, 2567]. discontinuities [2790, 3552]. discontinuity [2925]. discordance [3803, 4508]. discover [4102]. Discovering [4121, 3842]. Discovery [1433, 1434, 1440, 1620, 1622, 1135, 330, 2362, 4569]. Discrepancy [2335, 3298, 4008]. Discrete [4712, 1437, 2066, 3729, 2761, 3242, 4306, 1028, 4650, 3551, 1531, 4068, 3077, 2744, 2901, 3889, 4356, 4334, 2019, 594, 3020, 1989, 4090, 3147, 4087, 1187, 1133, 2792, 1015, 1016, 710, 2569, 2170]. discrete-time [1531]. discretization [3900]. Discretized [1400, 4483]. Discriminant [2233, 71, 3503, 2484, 3609, 4132, 2981, 3854, 299, 610].

Discriminating [397, 2660]. Discrimination [584, 690, 3400, 62, 3899, 2254, 1559, 3273, 653, 2955, 4202]. discriminative [4454]. discriminatory [4345]. Discussion [1238, 1225, 1218, 3644, 3921, 1202, 1215, 1207, 4479]. disease [1994, 4530, 4550, 3794, 977, 3207, 1322, 2409, 2876, 4396, 2268, 4508, 2643, 3004, 2105, 2597, 4504, 3632, 3339, 4397, 4440, 3376, 441, 2591, 3762, 3946, 3866, 2478, 4099, 2388, 3614, 2075, 4158, 1876, 2232, 2517, 2795, 1346, 2229, 2540, 2513, 2918, 2855, 2095, 1889, 3605]. disease-specific [2918]. diseases [3634, 1478, 300, 4595]. Disentangling [1220]. Dismissals [185]. disorder [3286, 3817, 4540, 2638]. disparities [4629, 4070]. disparity [4406]. dispersal [1235, 1360, 740]. dispersed [4334, 1808, 2784]. Dispersion [1449, 1670, 170, 1134, 1583, 3888, 2800, 4423, 3030, 1517, 578, 2279, 3960, 964, 873, 1928, 1821, 2381, 3087, 4003, 3409, 2046, 2763, 3558]. Display [4720]. displays [495]. Disputes [1653, 962]. disruption [2758]. dissemination [2346]. Distance [3703, 2592, 3230, 3937, 1526, 2447, 2471, 1972, 3490, 2922, 1260, 2537, 2295, 3878, 2497, 1341, 1002, 3905, 1915, 3416, 4227, 3372, 4407]. Distance-based [3703, 2592, 3230, 1972]. distances [1163, 1360]. distinct [1571, 982, 1079]. distinctness [1163]. distress [2974, 4540]. distributed [91, 2779, 3007, 3402, 3212, 4514, 2883, 4391, 4556, 3589, 4353, 3876, 1110, 3640, 4449, 1112]. Distribution [1039, 4647, 1637, 1657, 1693, 126, 3980, 1671, 3114, 1421, 1377, 3392, 1608, 4735, 1023, 1611, 3346, 4463, 3990, 225, 688, 1425, 2687, 3260, 2665, 4144, 4364, 1198, 3319, 4359, 3440, 4004, 3869, 1113, 4324, 3907, 4652, 3395, 4306, 4139, 2677, 3485, 1906, 2032, 1977, 3357, 2921, 2741, 4505, 1354, 3332, 4205, 374, 3834, 3510, 4047, 1028, 2820, 3753, 463, 3671, 932, 553, 836, 1894, 2049, 3935, 4576, 3888, 3328, 4155, 2171, 2568, 4068, 1009, 3728, 3745, 2659, 4038, 3574, 380, 3648, 1937, 1949, 1829, 472, 2632, 3541, 3429, 3174, 320, 2744, 2935, 3923, 4354, 1247, 2318, 3520]. distribution [4365, 4590, 3889, 4334, 2725, 3818, 3511, 4484, 1578, 1900, 1286, 3417, 2549, 1269, 1530, 1508, 978, 2288, 4192, 831, 3419, 3477, 2870, 3295, 2015, 4030, 3426, 1871, 2775, 2990, 3965, 1184, 2300, 4140, 852, 1597, 973, 4215, 4239, 1763, 1981, 1969, 894, 2844, 1555, 2014, 3062, 1296, 3280, 3118, 381, 1254, 846, 4479, 1828, 700, 801, 3983, 4466, 4035, 1284, 3879, 3323, 4686, 2724, 3539, 382, 456, 3221, 3397, 2308, 3467, 3536, 1672, 817, 2350, 3947, 2185, 2490, 2885, 1868, 2330, 3354, 4612, 3748, 3414, 1858, 4313, 3592, 4024, 3371, 2446, 315, 1915, 2178, 3743, 3922, 1957, 1465]. distribution [1351, 3435, 3928, 2860, 3948, 1505, 1848, 3920, 3727, 4317, 4488, 3493, 4672, 3147, 4057, 4002, 2630, 3098, 3343, 3971, 1528, 3542, 2892, 1859, 2770, 3497, 4285, 3964, 4383, 2579, 3959, 4677, 4560, 1096, 2175, 3916, 1567, 4216, 3772, 3886, 3851, 2515, 2924, 4298, 4417, 3533, 3806, 4412, 4497, 1939, 1004, 994, 3291, 3476, 3289, 4018, 4225, 263, 322, 2669, 4141, 2624, 2248, 3464, 4431, 2799, 3996, 4094, 1580, 883, 1339, 2572, 2164, 4481, 2535, 2694, 3744, 4362, 985, 2712, 4213, 371, 4016, 1302, 1804, 2865, 4054, 4513, 568, 3353, 3864, 4150, 4027, 2938, 2681, 4342, 4624, 2261]. distribution [740, 3921]. distribution-based [2568, 2164]. Distribution-free [4647, 3114, 1023, 4463, 3990, 3467, 4054]. Distributional [1446, 1690, 3153, 3434, 4467]. Distributions [828, 136, 214, 1687, 1432, 1400, 1688, 997, 3400, 3503, 1842, 3242, 3460, 1964, 3994, 4152, 1965, 3655, 584, 1155, 2670, 4191, 3671, 668, 3121, 3997, 397, 1975, 2329, 620, 3153, 4475, 2358, 3717, 3182, 560, 2901, 3639, 1214, 4321, 4151, 4356, 4134, 2957, 1516, 3962, 2776, 1728, 2079, 3150, 4305, 4200, 4194, 2247, 3358, 2494, 3152, 4603, 594, 1033, 2944, 3862, 3325, 3867, 816, 4578, 2672, 4201, 1252, 2365, 1344, 2557, 2773, 2386, 4031, 3763, 896, 1974, 2520, 4389, 3646, 2633, 953, 1771, 334, 1366, 1313, 3422, 4087, 3276, 3635, 1335, 3790, 975, 3675, 423, 2660, 2132]. distributions [1187, 3269, 2704, 695, 1215, 1547, 1532, 4382, 1749, 1533, 4093, 3622, 4290, 859, 4091, 586, 3257, 2074, 1802, 4675, 4661, 1363, 1521, 710, 2118, 2999, 3974, 1308, 3442, 2155]. District [2269, 931, 86]. District-level [2269]. districts [4511]. disturbances [539, 3231]. disturbing [1960]. diurnal [287]. divergence [4059, 1665, 3104, 1445, 3183, 2846, 909]. Divergences [3304]. diverging [2849]. Diverse [3598, 863]. diversification [3456]. diversified [2997]. Diversity [1592, 1836, 574, 864, 2242]. division [1966]. divorce [4277]. DNA [2240, 2112, 2301, 2489, 2723, 851, 2783]. Do [3908, 1855, 3340, 1619, 1012, 2215, 2680, 1, 2001, 2474, 3019]. document [4450]. documents [4682, 3857]. Does [3456, 127, 3659, 1911]. doing [1225]. domain [601, 1284, 3316, 3669]. domains [3666, 417, 4597]. domesticus} [768]. dominance [4177, 4399]. dominated [500]. dominoes [543]. Done [134]. donors [2866]. Doppler [2536]. Dose [1665, 220, 3912, 4001, 1001, 1277, 1369, 1538, 1074, 3677, 4183, 1575, 1032, 322, 1097]. dose-binary [322]. dose-finding [4001]. Dose-related [1665]. dose-response [220, 1001, 1538, 1074]. dot [2104]. Double [1451, 2100, 2067, 1816, 1627, 4094, 2223, 1680, 1683, 2380, 4523, 4664, 697, 1965, 2032, 1017, 4423, 1776, 2976, 3034, 1246, 4603, 4468, 2465, 3946, 575, 621, 2006, 905]. double-ranked [2380]. Double-sampling [2100, 697, 3946, 905]. double-threshold [4603]. Doubly [4391, 2113, 4089, 950, 1344, 3959, 2270, 2143, 2467, 2663, 2978, 4628]. doubly-inflated [4089]. doves [757, 1233]. down [3520, 1074, 3489]. DP [1959]. DP-based [1959]. Dr [78]. Dr. [138]. dragon [3806]. dragonfly [4527]. dramatic [1870]. drift [1973]. Drive [1729]. driven [2387, 3207, 4519, 2611, 3228, 4547, 2411]. driver [4486]. drivers [4532, 3561]. driving [3294]. drop [2262, 2866]. drop-out [2262, 2866]. Dropout [4377, 2186, 2711, 3178, 3093, 1274]. dropouts [4348, 3047, 4253]. drought [4168, 1848]. drug [880, 2899, 4605]. Drugs [1647]. Dry [1641]. dual [2573, 1040, 2244, 2121]. dual-record [1040]. duct [4568]. due [3634, 1472, 2175, 740]. dueling [2278]. Dummy [2771, 3255, 3200]. Duncan [253]. duplication [4600]. Duration [1653, 2969, 2076, 1999, 2035, 1926, 1311, 3356, 500, 962, 2400, 1173, 4173, 701, 4018, 903]. during [1748, 2420, 2714, 4512, 4162, 1924, 2791, 339]. DX [4406]. dyadic [2475]. Dynamic [2974, 4304, 1170, 3975, 2767, 4343, 4277, 4717, 4073, 2275, 2997, 4360, 3202, 4472, 3595, 2842, 3456, 4303, 4404, 2675, 2976, 2776, 2729, 4630, 1546, 1950, 1323, 2400, 3624, 762, 3467, 4489, 2398, 3438, 2877, 4595, 3422, 4096, 2708, 3196, 3762, 3596, 3449, 1371, 4629, 4053, 2915, 4453]. Dynamics [1664, 2791, 4316, 4251, 4369, 1875, 2020, 3376, 2716, 2388, 4308, 3132]. dyslipidemia [3918]. dysplasia [2560, 2448]. dysregulation [2638].

E-Bayesian [3916]. e-book [3572].

e-commerce [2728]. e-mail [3846]. each [1976, 2451]. ear [4611, 261]. Early [2075, 4545, 4573, 2584, 4611]. early-stage [4573]. earnings [3167, 2884]. earthquake [3760, 2224, 3536, 2349, 3571]. Earthquakes [2334, 3256, 4384, 3354]. east [3016, 1300, 111, 235, 281]. eastern [1209, 3524]. easy [1783, 1068]. easy-to-use [1783]. EBLUP [1633]. Ebola [4237]. ebook [4599]. eccentrics [483]. echo [3570]. echo-state [3570]. echocardiography [3942]. eclipsing [3890]. ECM [4067]. ecological [3842, 4232, 1844, 2399, 2389]. ecology [761, 1203, 1201, 2598, 3966, 751, 1202, 3024]. Econometric [2577, 4173, 2676, 3389]. Econometrics [4721, 2327, 3534, 2277, 3197]. Economic [1320, 4727, 1273, 1102, 1186, 3244, 2500, 3371, 2322, 173, 3655, 2107, 2697, 3214, 3908, 3517, 3797, 1310, 989, 1314, 2563, 1555, 3106, 916, 4679, 462, 2474, 4270, 2645, 4560, 3675, 3772, 2283, 3906, 2248, 1812, 3257, 2702, 2653]. Economic-statistical [1102, 3244, 2283]. Economical [1426, 286]. Economically [1329]. Economics [1377, 1019, 4148, 2577]. economies [3554]. Economy [1642]. ecosystem [3555]. Edema [2405]. EDF [3971]. EDF-based [3971]. edge [4128]. edit [3529]. edit-imputation [3529]. edited [2818, 4599, 2653]. edition [1899, 2816, 2817, 1988, 2824, 2617, 2823, 3120, 2089, 2814, 2191, 2952, 2192]. editor [1715, 114, 105, 241, 454, 176, 259, 94, 124, 131, 68, 41, 139, 54, 30]. Editorial [14, 16, 232, 313, 369, 730, 879, 2092, 2328, 2756, 2954, 3155, 1790, 2093, 2344, 2544, 3302, 3348, 2757, 4148, 784, 4101, 4563, 4032, 3993, 4533, 4443, 2543]. Education [3156, 1531, 3167, 2296, 3170, 1931, 3581, 2406, 4028, 4277]. education/employment [1931]. educational [3164, 42, 3172, 4428, 4466, 3166]. Edwards [1619]. EEC [197]. EEG [3765]. Effect [1625, 2142, 3664, 686, 1707, 2666, 4353, 706, 1081, 3157, 4156, 1722, 2709, 4100, 1997, 1354, 1985, 1560, 1235, 1476, 2603, 3574, 1891, 611, 3314, 2753, 2861, 4400, 962, 2931, 1254, 3644, 1720, 2111, 3601, 1800, 4518, 1172, 2043, 4221, 2351, 902, 4277, 2390, 2904, 358, 2058, 2947, 1196, 4095, 3093, 1741, 3992, 484, 460, 496, 4481, 3165, 2199, 2728, 3902, 4070, 4406, 2946, 4621, 3558, 4190]. effecting [3050]. effective [712, 1001, 1277, 4445, 1032, 2791]. effectively [2150]. Effectiveness [2894, 1985, 3344, 3437, 1853, 1055, 3673]. Effects [1443, 1399, 1236, 140, 1660, 118, 1699, 1663, 919, 392, 1379, 1670, 225, 1986, 757, 1736, 2874, 1128, 1583, 3215, 4296, 3510, 3214, 2262, 2545, 1217, 2900, 3224, 2342, 1209, 1596, 2487, 3872, 3402, 1496, 2887, 3875, 4598, 1219, 2561, 3883, 578, 2139, 4498, 788, 3419, 1087, 3061, 2956, 3793, 1778, 604, 3413, 4393, 1218, 3246, 2364, 2400, 4063, 725, 3102, 1129, 2214, 4646, 3601, 3090, 2989, 4468, 3657, 3376, 3654, 2405, 1885, 3807, 4299, 2410, 4090, 4381, 3752, 835, 4011, 3849, 4015, 4025, 1233, 3196, 3382, 3545, 1735, 3790, 1208, 3140, 2963, 2923, 1739, 1223, 739, 1499, 2467, 4540]. effects [1234, 2402, 884, 2056, 1463, 911, 4502, 2109, 2213, 3856, 1959, 3459, 3732, 529, 3779, 966, 3059, 4111, 4250, 2104, 2884, 4527]. Efficacy [1699, 1546, 4333, 3714]. Efficiencies [808, 1928]. Efficiency [1168, 1642, 1668, 829, 1769, 3894, 3391, 2281, 1900, 970, 1840, 713, 1544, 684, 1573, 3360, 2005, 4317, 2831, 2265]. Efficient [2471, 1783, 2731, 4248, 4743, 4183, 4526, 4374, 4307, 1125, 3908, 4530, 3051, 4660, 4076, 4572, 2302, 2819, 3343, 4416, 886, 2919, 3394, 3278, 3576, 2606, 4392, 4407]. effort [1137, 758, 2583, 1204]. efforts [4170]. Egrioglu [2745]. eiders [1204]. eigenstructure [1979]. Elastic [4425, 3399, 4119, 4638, 3895, 3208, 4147]. elastic-net [3208, 4147]. elasticity [644]. elderly [1735]. election [262, 931, 24, 2762, 4014, 2768, 36, 824]. elections [246, 4270]. electric [1784, 1106]. Electricity [2082, 2797, 4532, 901, 4317, 2555]. electromyographic [2646]. electromyography [3568]. electron [417]. electronic [3824, 2893, 3936, 3401, 3687]. element [532]. elementary [2460, 110, 591]. elements [1011, 1079, 447]. elevation [2775]. elicitation [3698, 2549]. elicited [3970]. Eliciting [2895]. Eliminating [1941, 140]. elimination [1819]. Elite [3382]. ellipses [375]. ellipsoids [3331]. Elliptical [2181, 4151, 1728, 3092, 3197, 3290, 4210, 3251, 2403, 3979]. elliptically [2883, 2132]. elm [814]. Elteren [2486]. EM-test [3909]. emergency [3689, 2360, 342, 1462]. emergent [4053]. emerging [1797, 3554, 1890]. emigration [745, 1230, 741]. emission [2404, 418, 3771, 635, 663, 2625]. emissions [916]. emotion [1556]. emphasizing [2415]. Empirical [1399, 1912, 687, 2568, 2187, 4103, 1672, 1633, 2397, 3098, 4323, 3697, 2992, 2270, 726, 3516, 3864, 3378, 3625, 3181, 1873, 335, 3935, 4684, 562, 2807, 3128, 2965, 4143, 3174, 2507, 3318, 507, 601, 3360, 3855, 2633, 2399, 2914, 3343, 1024, 4208, 4659, 345, 2201, 2638, 3055, 2385, 3668, 2571]. Empirical-distribution-function [1672]. empirically [285]. employability [3169]. employed [1141]. Employee [3432, 3252, 2834]. employer [2834]. employer-employee [2834]. employing [3392]. Employment [1931]. emptying [2946]. Emulation [3103]. en-bloc [955]. enamel [1754]. enclosures [3500]. encoding [3970]. Encyclopedia [4704, 2157]. endemic [4397, 2229, 2540]. endemic-epidemic [4397]. endogeneity [2406, 4090, 4242]. endpoint [3912, 1307]. Endpoints [1391, 4573, 3589]. Energy [3526, 4257, 635, 2308, 3481, 3304, 3800, 1481]. Engel [644, 4204]. Engineering [209, 33, 818, 7]. Engineers [2320, 2078]. England [2141, 1521]. Engle [4721, 2277]. English [1517, 408, 677]. enhance [512]. enhanced [4566, 3914]. Enhancement [1642]. Enhancements [748]. Enriched [1470]. enrichment [4088]. Ensemble [4400, 3598, 3688, 4659, 2847, 2385]. ensembles [3447]. ensuring [2622]. ENTERPRISE [2325, 2637, 1891, 2115]. entrance [2294, 3162]. entrepreneurship [4173]. entropies [2782]. Entropy [3857, 234, 32, 2881, 1891, 2993, 2689, 2300, 3971, 2077, 2559, 3338, 4145, 333, 3364, 2444]. entropy-based [3971]. entry [1997, 304]. envelope [2253]. envelopment [1900, 2532, 2627]. Environment [1419, 828, 749, 3300]. Environmental [1605, 787, 1452, 2284, 4509, 3488, 791, 4471, 1208, 3240, 3607, 1150, 888, 3553, 4170, 785]. environmentally [2813]. Epidemic [4645, 1478, 1467, 4545, 4397, 4539]. epidemics [2259]. epidemiological [3191, 791, 1114]. Epidemiology [90, 1948, 2387, 907]. epigraphs [3223]. epilepsy [2584]. episode [2604]. Epoch [3740]. Epsilon [1421, 3152, 2535]. epsilon-skew- [2535]. Epsilon-Skew-Normal [1421, 3152]. EQ [3621]. EQ-5D [3621]. Equal [1701, 4142]. equal-bin-count [4142]. Equal-bin-width [4142]. Equality [1431, 2965, 1095, 3617, 1853, 2178, 4437, 844, 1066, 1501, 689, 2940]. Equation [1631, 1891, 2149, 4483, 279, 3306, 1893, 1950, 3771, 2375, 3204, 1860, 971, 3597, 2451, 2642, 1540]. Equations [3919, 2003, 1322, 2212, 1323, 3739, 2944, 3389, 2478, 2857]. equidispersion [3708]. equilibrium [1925, 1973]. equine [2352]. equipartition [3017]. Equivalence [1381, 4380, 3352, 2062]. equivalent [252, 446]. equivalents [1860]. era [3995]. Ergodic [2615]. Erlang [2147, 1687]. Erlang-3 [1687]. Erlang-4 [1687]. Erlang-failure [2147]. Erol [2745]. Erratum [2542, 2619, 2755, 570]. Error [1405, 2455, 1686, 1706, 4144, 4252, 2881, 1722, 4665, 3480, 1364, 3942, 1811, 4267, 2142, 4007, 4282, 2409, 2481, 679, 2776, 2532, 3358, 1893, 2293, 2944, 3398, 3280, 2604, 1342, 1160, 4683, 1509, 2020, 1935, 2421, 3316, 1984, 2351, 3276, 875, 4323, 4292, 4559, 4531, 2058, 3178, 3264, 984, 1532, 1065, 4062, 356, 727, 722, 861, 549, 1716, 1995, 3560, 3475, 2681, 4500, 704, 2571, 3651]. error-contaminated [4500]. error-prone [2571]. Errors [2869, 1646, 1647, 1510, 1769, 91, 1291, 2779, 3212, 2887, 2682, 3009, 1286, 3581, 1116, 507, 4085, 2883, 1938, 1063, 3263, 3102, 2211, 2398, 2989, 1047, 4391, 392, 3776, 1529, 1947, 4095, 3436, 389, 692, 3370, 4146, 3659, 2213, 2854, 3757, 1552, 3475, 3845, 4100]. errors-in-covariates [2854]. Errors-in-Variables [1646, 2869, 1291, 1286, 1063, 692, 1552, 3475]. eruption [4335]. Essays [4721, 2277]. essential [3416]. Establishing [1424]. establishment [2346, 3529]. estate [786]. Estimate [1414, 1244, 1354, 2447, 3320, 1241, 3690, 745, 3469, 4608, 465, 1024, 1175, 375, 2583, 746]. estimated [3503, 2995, 744, 4058, 2207, 4014, 3054]. Estimates [136, 1037, 224, 197, 1820, 872, 763, 3214, 2280, 2335, 1084, 758, 2130, 3521, 873, 1490, 1860, 1773, 1511, 3563, 886, 2377, 2908, 1739, 581, 3416, 2514, 1812, 2303, 3625]. Estimating [2384, 2026, 532, 1631, 357, 4043, 3760, 3423, 3234, 1951, 767, 1571, 2828, 732, 1243, 279, 1287, 1116, 2374, 764, 684, 723, 4668, 743, 2493, 4204, 3657, 4471, 874, 4317, 2559, 2061, 2401, 2515, 1427, 4379, 3067, 1240, 742, 869, 3076, 859, 1319, 1360, 2693, 1494, 3673, 3550, 921, 1873, 979, 107, 936, 802, 2003, 1322, 2212, 2073, 4545, 2787, 2721, 3543, 1900, 2414, 1159, 2519, 1757, 644, 1546, 590, 962, 2494, 3739, 2234, 2735, 682, 2009, 65, 2434, 1074, 2469, 1860, 2877, 266, 2190, 1056, 1230, 3752, 4231, 2914, 80, 2390, 3725, 747, 2030, 2265, 3758, 1834]. estimating [3822, 3462, 1996, 3492, 4591, 1564, 4034, 1540, 1077, 1253]. Estimation [3440, 1625, 1399, 3998, 2741, 4126, 1693, 2113, 1643, 211, 1671, 1962, 3835, 482, 3429, 1311, 3755, 4590, 2925, 3523, 2689, 1219, 4305, 1362, 2723, 4113, 301, 1472, 4215, 894, 1609, 853, 815, 4408, 2350, 4583, 4651, 1661, 974, 1457, 3740, 1633, 3435, 1689, 4672, 835, 3497, 424, 685, 2924, 1453, 1620, 493, 1927, 4358, 818, 1632, 3218, 4662, 446, 692, 2539, 3464, 727, 884, 1653, 863, 1599, 586, 752, 1629, 3882, 4105, 1804, 3288, 966, 4181, 2761, 2805, 4324, 4224, 2881, 3193, 2804, 3679, 687, 4243, 3751, 3753, 4687, 4650, 463, 4376, 2697, 1238, 4564, 1560, 3847, 2706]. estimation [3138, 4684, 4576, 1758, 598, 4119, 3471, 2686, 3420, 778, 1266, 1566, 4068, 1967, 2280, 2603, 1523, 3089, 3252, 3057, 2083, 2718, 2582, 1276, 2269, 4316, 1141, 1913, 4460, 1321, 2790, 2632, 2715, 3277, 2249, 3804, 3923, 2507, 1486, 2318, 2829, 3889, 4356, 4134, 4214, 805, 3630, 2561, 729, 3042, 1269, 3104, 4012, 2279, 2492, 1934, 950, 1469, 3477, 4514, 3137, 1466, 809, 2406, 4592, 1914, 1001, 3986, 4248, 4400, 396, 2593, 3854, 1239, 3507, 964, 3359, 564, 3071, 1277, 2300, 502, 918, 594, 2404, 2372, 3325, 1922, 4393, 1763, 3135, 574, 418, 4076, 486, 2740, 3925, 3878, 3062, 2752]. estimation [381, 686, 2973, 3661, 4507, 1252, 4035, 4572, 4373, 1160, 1341, 2581, 3405, 3284, 1334, 2594, 3527, 4343, 3536, 1213, 2899, 3151, 4566, 2084, 3056, 2634, 4322, 4627, 2184, 1131, 2334, 4351, 3546, 4172, 3502, 992, 1772, 2811, 3307, 2626, 3825, 3493, 4067, 3967, 1237, 1255, 1071, 316, 4271, 305, 4323, 1904, 2770, 1026, 3959, 4020, 1983, 3363, 4666, 4292, 613, 3047, 3338, 345, 4589, 928, 4298, 1036, 1461, 2923, 1133, 3533, 1223, 1821, 2846, 3271, 2473, 4162, 3454, 4212, 4626, 995, 739, 2992, 1734, 1998, 1866, 2189, 3044, 4062, 4242, 3812, 1032, 587, 1075, 1282, 2477, 2669, 3403, 2402, 4336]. estimation [1358, 4526, 411, 3370, 4235, 2840, 3243, 3953, 1114, 3364, 4153, 3858, 2572, 4320, 3576, 1303, 2570, 2712, 2927, 549, 4307, 4125, 3565, 2918, 2865, 4261, 2856, 3261, 1484, 2122, 3053, 3902, 4133, 2188, 4150, 2531, 2311, 4253, 4407, 4447, 4110, 3074, 3442, 3832, 4190, 3456]. estimations [2849, 3397, 3090, 3916, 3268, 3560]. Estimator [1396, 1377, 1646, 1633, 1416, 1627, 4144, 3388, 2845, 4114, 1030, 2420, 728, 615, 3391, 2794, 1496, 3051, 2481, 3063, 3873, 4160, 4021, 4179, 1183, 1073, 4310, 3406, 2590, 3225, 3650, 982, 1079, 2811, 1991, 3478, 4086, 4162, 2189, 4294, 893, 3275, 2169, 1308, 3991, 1030, 1375]. Estimators [1040, 1190, 2665, 829, 2082, 2363, 4388, 693, 4155, 380, 4161, 2678, 2128, 4149, 4248, 2537, 3246, 2221, 4122, 2596, 3838, 1348, 1864, 2205, 3141, 2421, 2558, 3443, 4031, 1324, 3316, 953, 992, 2860, 1195, 2831, 4667, 4416, 672, 3697, 3278, 4141, 96, 808, 1303, 2138, 568, 3864, 3377]. Estonia [766]. ethanol [4219]. ethics [70]. Ethiopia [3332]. ethnic [4070]. ethnicity [3580]. ethylene [3833]. etiologic [3191]. Euclidean [2922, 3216]. EURING [1200]. Europe [3199, 3892, 2346, 2078]. European [2994, 4121, 3167, 4679, 1220, 2438, 3830, 18, 2306, 226]. Eurovision [2920]. EV [4447]. evaluate [1322, 970, 1860, 1055, 3494]. Evaluating [1278, 3797, 3294, 1650, 3318, 3723, 1723, 2798, 3820]. Evaluation [1217, 3212, 1378, 2086, 3544, 3599, 930, 1282, 2362, 1358, 3932, 3176, 2164, 1337, 4522, 2801, 468, 2845, 2733, 2903, 2684, 707, 2305, 1353, 3254, 3517, 3344, 3526, 802, 442, 380, 1980, 240, 1357, 3170, 1279, 4103, 3513, 2897, 864, 1590, 3752, 3478, 2551, 4568, 2638, 905, 1099, 4688, 1177, 2884, 1218]. Evaluations [2057, 1368, 4189]. even [3767]. event [2896, 3294, 2912, 3334, 4321, 3690, 3507, 4335, 4069, 3819, 3434, 3770, 3601, 2080, 3125, 1902, 4084, 2546, 2911, 779, 1307, 2584, 3720, 2310, 4409, 3856, 3868, 3132]. event-varying [3601]. events [3485, 1040, 2553, 3556, 4318, 360, 3660, 2637, 2879, 1926, 2851, 998, 3507, 2875, 4069, 3339, 3095, 551, 3067, 3733, 4054, 4405]. every [591]. everything [1210]. Evidence [2920, 1668, 2418, 2733, 2968, 2107, 3214, 3254, 3797, 2361, 4458, 2033, 3930, 601, 3577, 3346, 3579, 2141, 2780, 4170, 2416, 3340, 1619]. evidence-based [4458]. EViews [4710, 2226]. Evolution [3309, 1660, 4472, 1372, 3300, 4542, 753, 4528]. evolutionary [3035, 761, 1203, 1201, 1835, 750, 751, 1202]. EWMA [2259, 3235, 4647, 4234, 4284, 1145, 2252, 2134, 2291, 1405, 3951, 4033, 2645, 1506, 4430, 2120, 4054]. EWMA- [2291]. Exact [1923, 4550, 949, 2659, 4211, 700, 980, 1344, 3236, 1482, 528, 1009, 999, 3844, 1588, 943, 1541, 961, 1097, 286, 4016, 1111, 887]. examined [1837, 934]. Examining [2020, 3165, 2474, 1234]. Example [182, 168, 2423, 2406, 1184, 2483, 2982, 3401, 465, 3478, 2751, 2058, 1874, 110, 481]. Examples [1657, 4699, 102, 4737, 444, 19, 671, 4323, 2652]. exams [1100]. exceedance [4193, 4624]. exceedances [2969, 4332]. exceeding [3707]. excel [2458]. excess [4252, 3972, 4116, 2056, 1164]. excessive [4678, 3880, 1100]. exchange [3214, 1476, 1159, 1920, 2581, 465, 1024, 1293, 1229, 3006, 1955, 1328, 1252]. exchangeability [2440, 4578]. exchangeable [4391]. excise [3250]. Exclusivity [1929]. exercise [112, 4396, 1098]. Exercises [4728, 87, 4724, 2891, 2394, 2824]. exhaustive [2420]. exhibiting [3199]. Exist [127]. existence [1785]. exoplanet [3745]. exoplanets [3040]. exoskeleton [3446]. Expanded [3139]. Expansions [2501, 560]. Expect [4723, 2651]. Expectancy [1674, 3305]. Expectation [1400, 2740, 3043, 4254, 3219, 4410]. expectation-maximization [2740, 4254, 3219]. expectations [2789, 940, 3974]. Expected [1666, 1618, 883, 1702, 1478, 1092, 792, 999, 4145, 4154]. expectile [4348, 4147]. Expendable [126]. expenditure [4257, 4214, 3858]. Expenditures [200, 1425]. expenses [339]. experiences [3330, 340]. Experiment [134, 3433, 1127, 3746, 403, 1562, 854, 1234, 883, 3230, 2229, 2540]. Experimental [1193, 116, 1613, 153, 3133, 3656, 1113, 2240, 3252, 1725, 4581, 1888, 3094, 578, 1144, 721, 2598, 4183, 1233, 580, 411, 394, 1204, 2461, 1481, 2938]. Experimentation [1887, 2439, 1479, 546, 3379, 405]. Experiments [4695, 133, 195, 1713, 665, 142, 1670, 148, 2292, 2845, 897, 2874, 4344, 1487, 1134, 573, 1289, 1191, 1775, 934, 3564, 1156, 220, 3515, 2471, 1232, 1596, 1930, 1944, 2222, 1587, 999, 2754, 2813, 3912, 578, 772, 1743, 713, 1033, 1117, 2930, 1326, 1908, 3146, 1148, 961, 993, 3499, 984, 1231, 1911, 589, 2078, 866, 1539, 3811, 282, 321, 2623, 1481, 4109, 2088]. Expert [3698, 150, 511, 3970, 522, 513, 523, 247, 521, 514, 525, 2895]. experts [3643, 517]. explain [3586]. explained [3305]. explaining [3149]. explanations [1966]. explanatory [2101, 2407, 2451, 2449]. explicit [4582, 3559, 4667, 2712]. EXPLORA [518]. exploration [3356]. Explorative [1697, 2601]. Exploratory [2423, 128, 4120, 3393, 1793, 1596, 670, 3321, 2071, 3250]. explore [1499, 4070, 4406]. Exploring [2603, 671, 2489, 1840, 1106, 1244]. explosive [2521]. Exponential [1709, 1422, 225, 997, 2472, 2977, 3440, 3907, 4139, 3994, 3485, 1977, 4505, 4126, 463, 932, 3328, 4670, 3423, 4167, 2659, 2632, 4143, 2935, 4365, 3519, 1516, 4140, 3325, 3867, 816, 2054, 1589, 2014, 4201, 2419, 1344, 2185, 3125, 2683, 3435, 3959, 3916, 3802, 3533, 3249, 4412, 4255, 3992, 2167, 2799, 4436, 267, 1111, 898, 4459, 2261, 917]. exponential-Poisson [4255]. exponentiality [3855, 4272, 4236, 4247]. exponentially [2259, 4405, 1112, 1327, 1368].

Exponentiated [4134, 3519, 4225, 4359, 1169, 3923, 3889, 3397, 3435, 3548, 3802, 2924, 2904, 4412, 1533, 3992, 4261]. exponentiated- [2904]. Exponentiated-exponential [3519, 3992]. exponentiated-Weibull [1169]. exponents [591]. export [2717, 3003]. export-led [2717, 3003]. exposure [1476, 560, 4509, 4598, 1537, 4518]. exposures [4207, 2523, 3492, 4250]. expression [4432, 2568, 3705, 2990, 2897, 3228, 4123]. Extended [3708, 1348, 2661, 2012, 3211, 1530, 2279, 4200, 3951, 3886, 2904, 3136, 1838, 4402, 2572, 2109, 2936]. Extended-REML [1348]. Extending [2573, 2135, 2941]. Extension [1992, 4127, 1811, 778, 1312, 2247, 3422, 2175, 2570]. Extensions [253, 4208, 4383]. extensive [2638]. external [3007]. extra [901, 1808, 4190]. extra-model [901]. extra-Poisson [4190]. Extraction [3482, 2296, 955]. extracts [863]. extrapolation [1878]. extrasolar [3115, 3407]. Extremal [4021, 2155]. Extreme [775, 2306, 1438, 2996, 1820, 4415, 2637, 793, 1037, 3174, 4413, 2758, 3536, 4524, 1195, 1366, 3422, 4231, 2972, 1157, 4020, 3495, 554, 3806, 3394, 2793, 1955, 4452, 4150, 3088]. extreme-value [2758]. extremes [2828, 3488, 3720, 818]. extropy [4224, 4247]. Eyam [187]. eye [4648].

F [4721, 4185, 3830]. F-test [4185]. F1 [3830]. fabrication [1818]. fabrics [2408]. face [820]. faced [2259]. facet [3176]. facial [702, 2528, 1559]. facilitate [3555]. FACT [3230]. Factor [3188, 2376, 2573, 3816, 1193, 4199, 4077, 4304, 2713, 2655, 3490, 4674, 4598, 3002, 772, 563, 4030, 3611, 4240, 3012, 1106, 1971, 4437, 3901, 3271, 2792, 2071, 2418, 2863]. factor-binding [2792]. Factorial [1666, 1404, 142, 405, 1412, 857, 856, 1887, 2439, 1479, 578, 772, 3793, 2808, 964, 1033, 340, 1773, 4015, 961, 1532, 282, 321, 529, 4123, 3779, 2104]. factorials [2654, 2154]. factoring [2376]. factorization [4189, 3041]. Factors [104, 1412, 1134, 4197, 4544, 4611, 1891, 1126, 3311, 578, 3050, 1950, 3204, 2960, 3166, 4270, 1208, 3820, 3019, 3296, 444, 1346, 2623, 4551]. factual [113]. Faddy [2665, 4144]. Failure [3405, 3769, 3440, 4046, 4039, 2113, 4388, 1913, 3863, 2045, 564, 2147, 724, 1969, 894, 3508, 4297, 4673, 2644, 2350, 2897, 3037, 4651, 3435, 3948, 3147, 3478, 3249, 994, 1866, 2539, 3852, 3800, 4105, 3220, 2137, 4110]. failure-censored [564, 1969, 2350]. failures [3786, 4415, 1781, 2016]. fair [677]. Fairness [1885, 2726]. Faithful [2694]. Falkner [1227]. fallible [3946]. falls [4458]. False [1325, 1620, 1622, 1731, 2362, 4569, 3810]. familial [1116, 2482]. families [4200, 4201, 4290]. Family [3686, 212, 4675, 1420, 2977, 1937, 4657, 3662, 4356, 4134, 3152, 973, 1488, 3862, 4322, 4389, 4351, 4087, 4009, 3756, 2567, 3886, 2904, 1122, 4382, 1533, 4290, 4350, 2482, 2801]. family-type [4322]. Farlie [1155]. Farlie-Gumbel-Morgenstern [1155]. farm [2689, 1514, 226]. Fashions [95]. Fast [4498, 1847, 2740, 1786, 1548, 4326]. fat [1722, 1803, 1335]. fat-tailed [1722, 1335]. fatigue [4013]. fatty [3268, 2624]. fault [4451, 2886]. fauna [792]. Fay [4373, 3858]. FDR [4388]. FDR-controlling [4388]. FE [3284]. Feasibility [4137]. Feature [4227, 3759, 3227, 3680, 4457]. featured [2291]. Features [2257, 1847, 4264, 702, 3860, 1853, 4649, 4319, 4586]. feed [957]. feed-forward [957]. feedback [1290, 3915, 1145]. Feedforward [1290]. feeding [482]. feedstock [1290]. fees [1586]. Fellegi [4349]. Female [2361, 4418, 500, 743]. Fenton [2815]. feral [3970]. Fermentation [1659]. Ferrara [1931]. fertility [2676, 2691, 1902, 833, 1798]. fertilizer [1278, 726]. fever [4025]. FFR [3262]. fibre [1582, 721, 1048]. fibre-optic [1048]. Fibres [17, 181]. fibrosis [2912]. fidelity [1236, 743, 3811]. fiducial [237, 4113, 1597, 4476]. Field [1398, 133, 125, 2484, 3535, 1587, 416, 4593, 3620]. Fields [1681, 3877, 3577, 3958, 424]. FIFA [2664, 3241]. Fifteen [369]. Fill [773]. Fill-in [773]. filter [4495, 2991, 521, 4227]. filter-based [2991]. filtering [2974, 3205, 2802, 3978, 4222, 2462]. filters [2886]. filters-based-fault [2886]. filtration [1718, 2982]. FIMIX [2028]. FIMIX-PLS [2028]. final [981]. Finance [4745, 4148, 2460, 2821]. Financial [2705, 4746, 4171, 2974, 2658, 3595, 1731, 3199, 1951, 2083, 3011, 3488, 2069, 2422, 3358, 2015, 2374, 2133, 2114, 3360, 4274, 2949, 2708, 4162]. finch [1229]. find [2400, 22]. Finding [991, 4001, 3677, 2623, 4669]. findings [2439]. Finite [1414, 1183, 1627, 4234, 1354, 836, 1953, 3234, 609, 1551, 4211, 4019, 3813, 1785, 2133, 3313, 3561, 2982, 2937, 1079, 3675, 4555, 3278, 2523, 3233, 2188, 4117, 3864, 4027, 3625, 4465]. Finite-sample [1183, 1551]. Finnish [1330, 2862]. firm [3439, 2265, 4628]. firms [3232, 2215]. First [3215, 4723, 1645, 163, 1624, 3769, 3440, 1560, 3345, 2866, 4438, 3094, 2561, 852, 2604, 2350, 3435, 2768, 1234, 2539, 2208, 3091, 2651]. first-digit [2768]. first-failure [3435]. first-failure-censored [3769, 3440]. first-failure-censoring [2539]. First-order [3215, 1624, 4438, 3094, 2561]. first-time [2866, 3091]. first-year [3345, 1234]. Fiscal [3360, 488]. Fish [158, 4240, 1053]. Fisher [1964, 528, 374, 380, 464, 999, 679, 116, 207, 2446, 896, 1199, 4444]. Fisher-exact [999]. Fisherian [125]. fishery [1559]. Fit [1387, 1698, 1439, 4224, 372, 2049, 4233, 1843, 2021, 4475, 4515, 1185, 1791, 1158, 3891, 3978, 2316, 4686, 3536, 3285, 1800, 3414, 2633, 3966, 4009, 3098, 3343, 3183, 3971, 1014, 4236, 4574, 4658, 3019, 1337, 1723, 4275, 2354]. fitness [1213]. Fits [1644, 2847, 287]. fitted [490, 4091, 2694]. Fitting [2235, 3728, 1832, 1794, 382, 456, 3675, 1004, 389, 3257, 2936, 2881, 3504, 2288, 640, 682, 4470, 3926, 1035, 971, 3853, 1764, 267, 321]. five [4511, 603, 834]. fixed [2113, 3872, 3875, 1087, 3413, 4393, 3246, 486, 3434, 1105, 4015, 956, 2923, 4112]. Fixing [772]. Fledging [1210]. Flexibility [1421]. Flexible [2429, 3169, 1427, 3396, 1998, 3789, 2098, 657, 2212, 4166, 3826, 2405, 2598, 4573, 3493, 3111, 4631, 3675, 3604, 2220, 653, 3269, 4290, 3576, 3914, 3832]. flight [2455]. flood [3690, 3464]. Florida [1241]. Flow [1619, 2759, 1762, 4486]. flows [2342, 4641]. fluctuation [3258]. fluctuations [3766]. Fluid [1762]. flux [2308]. Flyway [754]. fMRI [1846, 3596, 3924, 3810]. foam [1794]. focal [2626]. focus [2165, 3720]. focused [1550]. foetens [863]. fold [1768]. folding [2368]. foldover [3773]. follow [1191, 2190]. follow-up [1191, 2190]. followed [2397]. following [4259, 1314, 3426, 2372, 4497]. food [3069, 2864, 2893, 3597]. foodservice [2005]. football [1954, 397, 4575, 3735, 3883, 3727, 3382, 4619, 319, 566, 2182]. footwear [4128, 1727]. forage [43]. Force [2153, 1689, 2656, 500]. Forces [3076]. Forecast [1478, 1650, 1935, 3420, 4167, 949, 3212, 597, 4395, 893, 3498]. Forecasting [4726, 1700, 2713, 1905, 1010, 4168, 1454, 901, 2714, 3930, 1606, 1514, 1617, 597, 1753, 4659, 1652, 200, 4548, 3179, 4650, 2647, 413, 4427, 354, 923, 3637, 2991, 3226, 4012, 1500, 24, 2133, 1472, 4580, 2722, 4607, 2260, 4073, 2919, 3375, 2555, 2152, 4026, 36, 1521, 3553, 4307, 1328, 3801, 2115, 4180, 4521, 3432, 2500, 2745]. Forecasts [1608, 3908, 1500, 3594, 3579, 3019, 1177]. foreclosure [3368]. foreign [2387]. forensic [2733, 3254, 4128]. forest [3660, 817, 4231, 4562]. forests [2338]. form [2585, 3834, 4166, 3038, 2958, 583, 3436]. formal [3158]. formalism [552]. formats [1342]. formed [721]. forms [278]. formulae [940]. formulas [1783]. Formulation [1700, 1127]. Forward [4562, 957, 3573, 662, 3570, 3178]. Four [2049, 1441, 3395, 631, 2099, 1137, 3773, 981, 3147, 4285, 3227]. Four-level [1441, 3773]. four-parameter [3395, 3147, 4285]. Four-Staged [3227]. Fourier [2066, 4712, 4166, 4165, 1639, 1336, 3070, 282, 287]. Fourier-transform [1336]. fourth [1893]. fourth-order [1893]. fox [3500]. fractal [2420]. fraction [2687, 1192, 1536, 3194, 1180, 2083, 2060, 919, 2796, 4378, 3731, 2027, 4350, 861, 935, 3882, 1813]. Fractional [2099, 1404, 1034, 342, 1412, 856, 3270, 404, 577, 2998, 2620, 1370, 4165, 772, 2808, 964, 972, 721, 3618, 854, 984, 2806]. Fractionally [3511, 2742, 2797, 1093, 1036]. Fractions [167]. fracture [2641, 3497]. Fragile [1733]. fragility [3785]. Frailty [1799, 2319, 2738, 2357, 1832, 2139, 3863, 3843, 2522, 2219, 4378, 2310, 3731, 1898]. Frame [161]. framework [4648, 4245, 1138, 4037, 4509, 4081, 3927, 2490, 2474, 3944, 2798, 2077, 2704, 4623, 3219, 4229]. France [283]. Francisco [884]. fraud [4529]. Fraudulent [1645]. Fréchet [4205, 3371, 3319]. Frederic [2652]. free [4647, 2912, 3114, 2808, 1573, 2445, 3467, 1023, 4463, 3990, 2728, 4054, 3486, 4124, 4621]. freedom [1199]. freely [4552]. freezing [768]. French [4270, 1668, 4014]. frequencies [1602, 2589, 1012, 3464, 1295]. Frequency [1438, 584, 3888, 4499, 4118, 1116, 3061, 2422, 601, 4520, 2080, 3354, 3617, 1497, 3006, 3669, 4349, 1349]. frequency-based [4349]. Frequentist [4464, 3046, 2900, 2517, 2168]. friendly [2813]. fronted [1236]. frontier [1456, 3837, 3526, 2649, 684, 3346, 1091]. frontiers [3308, 4390, 4101]. FS [3825]. FS-LS-SVR [3825]. FT [1336]. FT-IR [1336]. Full [4664, 2155, 2418, 4624]. full-factor [2418]. Fuller [1551]. Fully [4482, 2369]. Function [1423, 1425, 775, 2761, 3503, 91, 2429, 3215, 3445, 4656, 4043, 1018, 2370, 3207, 3286, 2279, 601, 1073, 3398, 2008, 2931, 801, 3838, 3221, 2830, 3185, 1672, 3341, 896, 2860, 1056, 305, 3959, 886, 3338, 2220, 1939, 3454, 322, 3060, 1513, 3335, 969, 3882, 4105, 4016, 1474, 4591, 4154, 3864, 4372, 2376, 4041]. Functional [1372, 1675, 4532, 3457, 2934, 3223, 1035, 3776, 2601, 2555, 4467, 1706, 3580, 4653, 2337, 2082, 4257, 548, 2145, 2447, 3532, 3277, 2964, 4424, 3672, 3755, 1517, 679, 3578, 3038, 4132, 3473, 3871, 3771, 3109, 4312, 2427, 4580, 2767, 4209, 3787, 583, 4293, 3588, 4666, 3248, 493, 4062, 3299, 2352, 4637, 4330, 3895, 3714, 3801, 4689, 3656]. functionals [4149]. functioning [690]. Functions [1677, 1453, 3769, 2772, 2632, 523, 240, 915, 600, 4165, 831, 2287, 2406, 644, 4593, 71, 2151, 3245, 1131, 2178, 3829, 1366, 3384, 283, 2559, 458, 3067, 1734, 4225, 4336, 4637, 3464, 3462, 2940, 3261, 3887]. fund [2070, 2418]. Fundamentals [2320]. funding [4173]. funds [2826]. fungicide [220]. furnace [1905]. Furnaces [1651]. Further [1596, 834, 1203]. fused [4010, 4539, 4111]. fusion [4504, 2908, 2792]. future [2789, 2212, 759, 2607, 762, 1717, 1204]. futures [3669]. Fuzzy [3985, 3384, 2658, 3158, 3311, 3226, 3064, 2722, 3513, 3840, 2837, 2838, 3494, 4454]. FWER [2057]. FWER-controlling [2057]. FX [4512, 2703].

G [4134, 2818, 4671]. G7 [2474]. gain [2943, 3673]. gait [3286, 2646]. gait-related [2646]. game [3746, 543]. games [2664, 3532, 4197, 4601, 3421]. GAMLSS [3535, 3853, 2417]. Gamma [2664, 1320, 3671, 668, 1975, 3717, 2335, 240, 3525, 3956, 1848, 4488, 3111, 2972, 3675, 4378, 3731, 4018, 4353, 4294, 1726, 1564, 4633, 2938]. Gamma-count [2938]. gamma-distributed [4353]. gamma-Poisson [1726]. gamma-ray [4018]. Gang [4599]. gap [3057, 3305, 3952, 4232, 3095, 1570]. gaps [3164, 2875, 2601]. GARCH [1879, 2807, 3511, 2776, 2017, 2133, 4358, 2418, 2975, 4407]. GARTFIMA [4684]. gas [3027, 1140, 2758, 303, 2404, 3825]. Gasoline [1411, 1855]. Gastric [1430, 4187, 2946]. gate [3952]. Gauge [902, 3950]. Gaussian [4130, 3512, 1693, 2102, 2026, 1305, 4052, 3745, 3203, 3535, 4302, 4478, 2299, 2247, 4503, 3407, 1597, 4239, 4610, 2317, 1449, 3958, 2843, 3175, 3343, 4383, 2061, 3772, 3240, 2371, 4481, 3897, 2806, 3974]. Gaussian-multinomial [1693]. Gaussians [3799]. GEE [4387, 2480]. GEEs [3458]. geese [1236, 1210, 754, 741, 1234]. gelatin [1818]. Gender [823, 3305, 3580]. gene [3213, 3694, 4432, 2568, 1718, 3705, 3981, 3927, 4456, 2897, 3228, 3607, 2792, 1876, 1295, 4123, 2168]. gene-based [4456]. gene-environmental [3607]. gene-gene [3694]. General [4732, 1436, 2433, 2421, 2916, 1616, 3166, 940, 4071, 4360, 3331, 3816, 2032, 3997, 1359, 3872, 4645, 4674, 1287, 3366, 2953, 1054, 3748, 3505, 2392, 2202, 4236, 4589, 4000, 1091, 3990, 4661, 2902, 4055, 2170]. Generalised [91, 2472, 1899, 3658, 2492, 2529, 3231]. generalization [2025, 1561, 941, 3499, 4242, 2402, 1189]. generalizations [4223]. Generalized [2988, 1029, 3534, 1948, 1476, 3821, 2993, 1949, 1247, 1665, 1605, 3862, 3398, 3633, 1389, 142, 1611, 1689, 1987, 1071, 4476, 3335, 3019, 4060, 1189, 3501, 2865, 4346, 4702, 3440, 2881, 2810, 1964, 806, 3485, 1965, 1977, 2741, 4126, 1155, 4516, 1270, 548, 2880, 4415, 2995, 3491, 2637, 2238, 3224, 3728, 1891, 2003, 1322, 464, 2212, 3113, 3892, 2632, 2926, 4143, 3174, 2935, 3804, 3520, 3755, 4151, 2464, 3600, 2976, 3676, 2549, 2656, 600, 2279, 1094, 3050, 4066, 2870, 1867, 1871, 3734, 2187, 3283, 3739, 2629, 3514, 3413, 3247, 2114, 2679, 2054, 1860, 3285, 4104, 4313, 2242, 3333, 4351]. generalized [1771, 1351, 1505, 3422, 2631, 2392, 3209, 3276, 3849, 3253, 1026, 2512, 618, 3684, 2559, 3916, 3338, 3386, 2660, 3870, 3251, 1188, 1004, 592, 2983, 3955, 4402, 2551, 4290, 4431, 2799, 2786, 1580, 2939, 4481, 2535, 1802, 2223, 4350, 4671, 1067, 2854, 2781, 3208, 1199, 2449, 3583, 1316, 2046, 2856, 4401, 1349, 4185, 3431, 2700, 2681, 4457, 2527, 3832]. generalizes [674]. Generalizing [2355]. Generally [1847, 4523, 4423]. Generated [188, 3862, 980, 4389, 3111]. Generating [2097, 1642, 2907, 19, 3949]. Generation [1444, 1124, 1753, 1942, 2838, 3230, 1921, 3353]. generator [450, 264]. genes [4102, 4456]. genesis [1382]. Genetic [1426, 148, 1420, 4246, 2130, 2236, 4422, 4468, 2850, 943, 2826, 1148, 1973, 4394, 608, 3516, 3001, 2527, 2911]. Geng [4599]. genome [2560]. genome-wide [2560]. Genomic [3680, 2355, 3315, 4663, 4063, 2719]. genotype [1876, 2650]. genotypes [3756]. Genotypic [888]. genres [1008]. genus [2136]. Geo [4130]. Geo-Classification [4130]. geodesic [2489, 3097, 2716]. Geof [2816]. Geographic [2844, 168, 743, 2907, 4053]. geographical [4259, 2933, 2303]. geographically [4374]. geographically-defined [4374]. geography [3008]. Geometric [4648, 2809, 3105, 1389, 1978, 3260, 1192, 2866, 4038, 3429, 3519, 3514, 1866, 3992, 3895, 1329, 4527]. geometrically [4300]. Geometry [3937]. Geometry-based [3937]. Geostatistical [3614, 2094]. Geostatistics [2617, 2921]. German [3908, 2346, 3775, 3620].

Germany [3179, 4373, 283]. germplasm [864]. Getting [2986]. GEV [4093]. geyser [4335, 2694]. Ghana [3383]. Gibbs [2383, 609, 2829, 1597, 4239, 2910, 2030, 423, 1959]. Gilles [2651]. Gillian [510]. gillnet [1840]. Gini [2563, 3246, 3338]. Gittins [3589]. given [2733, 3254, 506, 565, 637, 4630, 4345, 567, 643, 646, 1494]. Givens [2816]. glandular [2740]. Glass [4]. gLasso [2443]. Glaucoma [1455]. GLCM [3716]. glim [382, 456]. GLM [4245, 4581, 1395, 4204, 3856]. GLM-based [4581]. GLM-type [3856]. GLMM [4217, 2480]. GLMs [1325]. Global [1638, 1820, 2140, 2648]. glomerular [2982]. GM [3391]. GME [2966]. GMM [3549, 1995]. Go [151]. goal [397, 3034]. Goiás [4025]. gold [2162, 597]. golden [3192]. golf [3181, 1288]. Gompertz [1198, 1937, 3295, 4612, 4662, 4431]. Gompertz-type [1198]. good [1887]. Goodman [3593]. Goodness [4515, 1387, 3978, 4686, 1698, 3285, 1800, 4009, 1439, 4236, 4574, 4224, 372, 1843, 2021, 4475, 1185, 1158, 3891, 3536, 3414, 2633, 3966, 3098, 3343, 3183, 4658, 4275, 2354]. Goodness-of-Fit [1387, 1698, 1439, 4515, 3978, 4686, 3285, 1800, 4009, 4236, 372, 1843, 2021, 4475, 1185, 1158, 3891, 2633, 3966, 3098, 3343, 4658, 4275, 2354]. goods [3746, 1148]. Google [4618]. goose [742, 752]. Gopal [1715, 2230, 2524]. Government [162, 1330, 47]. Governments [200]. GP [4093]. GQL [3549]. grade [937]. grades [3345]. Gradient [2413, 3058]. gradients [791]. Graduate [1707]. graduates [3168, 1931]. graduation [3162]. grain [2939]. Grand [1251]. Granger [3029, 3706, 4615, 2043, 1109]. Granger-causality [1109, 3029, 4615]. grants [3043]. graph [4097, 4670, 4490]. Graphical [1526, 541, 3645, 21, 1279, 1021, 3796, 4720, 1552, 887, 2845, 516, 522, 3512, 3941, 530, 495, 3681, 2289, 3124, 844, 3924]. graphically [671]. Graphics [2456, 2424]. graphs [1008, 1011, 3880, 873, 1928, 3087, 3765, 4446]. gravity [959]. great [763, 611]. greater [931]. Greece [3760, 3412]. Greek [4384]. greenhouse [2404]. Griffon [737]. Grizzle [2402]. Gröbner [1101]. grocery [1181]. ground [3776]. groundfish [4316]. grounds [3899]. Group [4530, 1404, 3788, 2441, 856, 1128, 1906, 1558, 3671, 3456, 494, 3108, 4434, 2085, 2731, 3255, 2440, 1969, 894, 951, 2724, 3513, 2597, 796, 804, 2207, 670, 4322, 2704, 3200, 1866, 2955, 4619, 4094, 3765, 3565, 1815, 4250, 3486, 2878]. group-based [2955]. group-censoring [1815]. group-specific [4322]. group-structured [670]. Grouped [1710, 2594, 1752, 644, 1993, 4091, 3288]. Grouping [199, 620, 1057, 888]. Groups [1648, 2243, 3803, 2864, 1341, 2207, 280, 4669, 2199, 2533]. groves [315]. growing [1595]. Growth [129, 2415, 3022, 2107, 3214, 3908, 413, 676, 3229, 2452, 4483, 1241, 2367, 805, 2131, 923, 2130, 790, 3566, 1875, 3750, 3374, 916, 1054, 1129, 3090, 2885, 2483, 2982, 2474, 487, 835, 1047, 1766, 1343, 1805, 2248, 2717, 3003, 3257, 3238, 4229, 3673, 2997, 2681]. GTL [1802]. GUI [2458]. guidance [4522]. Guide [4714, 2193, 4691, 2610, 608, 2325]. Guided [4728]. guidelines [4559, 284]. Guinea [1462]. Gulf [3760]. gull [736]. gulls [766]. Gumbel [1155, 4068, 1949, 3539, 3971]. Guttman [4221]. GWMA [1838]. Gyps [737]. gypsy [669]. gyration [1456].

H [2822, 2816, 2653]. Habitat [1681, 3970, 794, 1229]. habitation [1983]. habituation [3059]. Hacking [237]. haemodynamic [1846]. Haghighi [4354]. Hakan [2745]. half [3397, 3291, 4261]. half-logistic [3291, 4261]. Hall [1418, 2343]. halves [942]. handball [1886]. Handbook [4693, 4730, 4749, 2036, 2500, 2393, 2503, 4725, 2462, 3024, 4703, 2746, 3120]. handle [4252]. Handling [4461, 4329, 1456, 3982, 2951, 3330, 3145, 3140]. handwritten [2295]. Hansen [2794]. harmonic [1851]. harvests [957]. Harville [304]. Hatke [4650]. having [1906, 339]. Hawkes [4237]. Hazard [1707, 1453, 2429, 2113, 4043, 1531, 2355, 3520, 3038, 2287, 3661, 1864, 4327, 3563, 2202, 3412, 3916, 4378, 2220, 4298, 3067, 4225, 3649, 3882, 4105]. Hazards [152, 165, 2034, 3786, 4415, 3122, 1294, 2355, 2942, 3987, 4515, 4256, 4241, 4230, 1864, 2634, 4651, 3807, 4355, 4299, 1283, 2663, 2208, 4319, 3730]. HBS [4204]. HDI [3442]. Health [152, 165, 90, 925, 1436, 4377, 2907, 173, 63, 2970, 3439, 1354, 2913, 3445, 3660, 1126, 2149, 4282, 3340, 4325, 2754, 2620, 4609, 3591, 4503, 3621, 3936, 1885, 2483, 3043, 4314, 4419, 551, 4618, 3973, 2642, 1933, 3602, 4385, 3330]. health-related [3445]. healthcare [2564, 3859, 3311, 2721, 3824, 3838, 2760, 2628]. heaping [1187, 3201]. heart [1913, 3514, 2897, 1346, 3605]. heartbeat [2420]. heat [4648]. heavily [4108]. Heavy [4065, 3983, 4383, 1564, 3074]. heavy-tail [4383]. heavy-tailed [3983, 1564, 3074]. hedge [2418]. hedging [3437, 2526]. hedonic [3234, 3462]. height [925, 1803]. Hellinger [3878, 4407]. help [4221]. hemisphere [1362]. heritability [2012]. Herriot [4373, 3858]. Heterogeneity [2545, 4377, 2556, 1128, 4575, 3841, 4546, 3913, 1852, 3191, 4357, 3360, 1104, 4558, 1591, 1222, 975, 741, 2028, 4623, 3436, 2271, 4621]. Heterogeneous [1272, 2725, 1081, 4441, 4126, 3510, 3441, 735, 2731, 3782, 4646, 3326, 3763, 2063, 4112, 2639, 2946, 3602]. heterogeneous-variance [2731]. Heteroscedastic [4484, 3548, 3679, 3211, 1000, 3962, 2910, 2774, 3783, 2403, 4311, 2199, 3496]. Heteroscedasticity [3130, 4252, 3072, 2564, 4387, 1292, 3007, 3369, 3082, 539, 2495, 2211, 2885, 2683, 4056, 3757, 698]. heterosis [4432]. Heteroskedastic [3653, 1980, 2055]. heteroskedasticity [2976, 3875, 2148, 2436, 2983]. Heuristic [682, 1951, 3598]. HGLM [1564]. Hidalgo [836]. hidden [2675, 2775, 3842, 4269, 2317, 4608, 3412, 3622, 3822, 3202]. Hierarchic [829]. Hierarchical [1666, 4188, 1910, 4553, 1378, 1664, 2937, 2795, 4279, 3805, 1954, 2463, 2920, 1775, 2244, 2358, 3452, 1926, 3170, 3543, 4048, 3562, 2404, 694, 3715, 2930, 3704, 3522, 3411, 4486, 4677, 4589, 2044, 3691, 3697, 4528, 2340, 3919, 3549, 4229, 3808, 2760, 3041, 2946, 2878]. Hierarchies [4169]. High [1684, 4275, 2977, 3703, 4495, 4246, 417, 4118, 3797, 3647, 4301, 4461, 2718, 4460, 3391, 4435, 4413, 3690, 3061, 1867, 2422, 3759, 3103, 2537, 2147, 2734, 3537, 2341, 3750, 3457, 3824, 3778, 951, 1064, 1765, 3096, 4024, 3905, 4210, 3189, 2443, 3754, 3227, 4403, 4338, 4676, 1497, 4604, 2504, 3044, 3607, 2928, 4410, 2786, 4569, 3208, 3453, 4420, 371, 4319, 4586, 2650, 4133, 4561, 4570, 3198, 4444, 526]. high-content [2341, 3457]. high-dimensional [2977, 4246, 3647, 4461, 4460, 4435, 4413, 3759, 3103, 2537, 3750, 3824, 951, 3096, 4024, 3905, 4210, 2443, 3754, 3227, 4403, 4338, 4676, 4604, 2504, 3044, 3607, 2928, 4410, 2786, 4569, 3208, 3453, 4420, 4319, 4586, 4133, 4444]. high-frequency [3061, 2422]. high-quality [2718, 1064, 1765]. high-throughput [3457, 2650, 4570]. Higher [223, 4058, 3077, 2296, 3170, 3581, 4028]. higher-order [3077]. highest [2674]. highly [1937, 4598, 3409]. highs [1920]. highway [3625]. hiker [608]. Hip [2641, 2560, 2019, 2251, 2448]. hippocampus [924]. histogram [4455, 4142]. histogram-valued [4455]. historic [3210, 1727]. historical [1872]. historie [1225]. histories [1224, 747].

History [160, 977, 1228, 812, 753, 1902]. Hitch [608]. Hitch-hiker [608]. Hits [1177]. Hits-and-misses [1177]. hitting [2208]. HIV [3383, 1050, 1467, 2113, 3794, 1413, 2000, 3941, 3112, 3664, 2729, 1546, 2398, 2493, 3632, 2877, 3376, 884, 3132]. HIV-1 [3112]. HIV-1-infected [2113]. HIV-infected [2000, 3664]. HIV/AIDS [1467]. HLA [410]. HMM [3354]. hoc} [1716]. Hoeting [2816]. Hofmann [1363]. Holan [2653]. Hold’em [2652, 4737]. holding [1515]. Holgersson [3255, 3200]. Holt [1700]. home [4243, 1985, 1900, 566, 593]. homicides [611, 2636]. Homogeneity [2243, 3578, 4474, 3844, 2348, 3909, 3783, 3272, 3839, 4476, 1461, 1592, 1916, 3385, 2231, 3610, 3370, 3078, 526, 965]. homogeneous [3950, 3158, 3104, 2775, 2268, 781, 379, 457, 1708, 2937, 3726, 2259, 4332, 1564]. homogeneously [4364]. homoscedastic [4273]. homoscedasticity [1760]. homosexual [884]. Hong [3906]. Honor [4721, 2277]. horse [895]. horse-race [895]. horticultural [1499]. horticulture [1908]. Hosmer [4658]. Hosmer-Lemeshow [4658]. Hospital [1709, 1520, 1472, 2005, 1735, 2795, 4034]. hospitalizations [3634]. hospitals [3468]. host [2343]. Hot [4540, 4571, 1310, 4291]. Hot-spots [4571, 4291]. hotel [4116]. Hotelling [1936, 4543, 1076, 3845]. hotspot [4565]. hours [4201]. house [768]. Household [3153, 3166, 2620, 4623, 3858]. households [303, 2872]. housing [3130, 3462, 3483, 4108]. HP [504]. HP-28S [504]. HSV [3112]. HSV-2 [3112]. Huberian [3286]. huge [3140]. Human [210, 1644, 1704, 4381, 3022, 2686, 4037, 3777, 3306, 4689, 4360, 3432]. humidity [3109]. Hungarian [755]. hunger [3126]. hunting [757, 731]. Huntington [4099]. hurdle [4441]. Hurst [1951]. Hurwitz [2794]. Huygen [1502]. Hybrid [4643, 2729, 939, 2784, 4039, 2658, 3865, 3780, 2900, 4365, 4590, 2622, 816, 4612, 4313, 4351, 2554, 3497, 3100, 2168, 2582]. Hybrid-based [2784]. hybridization [3337]. hydraulics [3219]. hydroelectric [2404, 3304]. hydrologic [3485]. hydrological [4318]. hydrology [1940]. hypberbola [674]. hyperactivity [2638]. hyperbolic [680, 685]. hypercube [4057, 857, 4145]. hypercubes [1331]. hyperspectral [3656]. hypertension [4119]. hypographs [3223]. Hypotheses [1387, 1699, 1487, 445, 4388, 3039, 1009, 1185, 797, 2552, 747, 1194, 2693, 3149]. Hypothesis [2853, 1384, 163, 2550, 172, 1632, 3641, 2444, 3030, 238, 2057, 2931, 889, 2474, 939, 49, 3209, 3119, 2717, 3003, 2902, 298, 338, 2556, 4029]. Hypothesis-Testing [172, 49].

I.C.I. [17]. I.D.A. [336]. I.O.S [121]. IBM [3120, 4749]. IBNR [3211]. ICA [2908]. ice [2716]. ICT [3320]. idea [2368]. Ideas [517, 394, 2502]. identical [557]. identically [3400]. identifiability [2947, 3459]. Identification [1134, 906, 3315, 925, 2734, 2621, 2917, 4456, 3827, 848, 2053, 3189, 4439, 1738, 4551, 2448, 3028, 1830, 4530, 2582, 2809, 3446, 4137, 1867, 3692, 2673, 1737, 3374, 820, 2399, 2852, 3969, 3691, 3806, 4540, 2792, 911, 2996, 2122, 4561, 4055]. identifications [3704]. identified [2853]. identify [3950, 4511, 3258, 1733, 3031]. Identifying [4538, 2701, 4197, 3941, 1524, 3562, 694, 1945, 2356, 1670, 4080, 1475, 2173, 335, 3666, 1782, 3537, 4360]. idiopathic [2912]. if [2853]. iFR [3262]. ignorability [2511]. ignorable [4047, 3828, 947, 4377, 4071, 2832]. ignored [4278]. II [4359, 2805, 4004, 4647, 463, 2021, 2129, 4581, 4354, 4365, 4305, 3137, 3259, 1872, 3418, 3397, 3770, 1698, 1344, 4612, 3748, 4672, 2770, 3916, 4236, 3850, 4333, 2515, 2924, 3178, 3291, 4662, 1730, 2799, 1539, 1110, 1339, 1302]. iid [1829]. III [3503, 4590, 3497]. Illustrated [1389]. Illustration [1644, 208, 1330, 2914]. illustrative [3401]. Image [451, 4712, 2772, 657, 1582, 453, 2096, 364, 421, 3847, 4302, 860, 416, 890, 420, 3688, 654, 665, 4281, 415, 570, 1559, 419, 4585, 3702, 655, 4570, 3656]. image-based [4585]. image-warping [890]. imagery [662]. Images [1455, 1639, 417, 452, 1582, 418, 1786, 651, 4128, 2959, 423, 3850, 605, 3216]. imaging [3441, 3457, 2962, 4585]. Imbalance [1666]. imbalanced [3807, 3900, 4372]. immigrant [3164]. immigrant-native [3164]. immigration [3892]. immune [3664]. immunes [1169]. immunity [2523]. immunologic [2711]. immunotherapies [4102]. Impact [1353, 3877, 1814, 4618, 1674, 2684, 1731, 1531, 2290, 1126, 4546, 1466, 2933, 2803, 4090, 1473, 1335, 2058, 1565, 3906, 4641, 3602, 4628]. impact-damaged [1565]. impacting [4551]. impacts [3797, 3456, 4270, 3238]. impedance [4118]. imperfect [872, 357]. Imperfectly [891]. implementation [1141, 2030, 1881, 4275, 2606]. implications [2076, 2237, 1979, 3153, 3639, 2278]. Importance [90, 4118, 75, 3033, 2251, 3620, 3659, 2127, 3629]. important [359, 3531]. Imports [197]. imposed [3444]. imposing [893]. imprecise [2027]. imprecision [902]. impressions [4128]. improper [3834, 2286]. improve [1905, 4663, 3241, 4204, 4209, 4057, 3019]. Improved [2380, 936, 4348, 2558, 1418, 4351, 2848, 1288, 3226, 1934, 3926, 4090, 4672, 1810, 3959, 1303, 3553]. Improvement [4079, 771, 1137, 3094, 3050, 1757, 721, 2009, 1908, 2309, 3385, 3013]. Improvements [1660, 1670, 3999]. improves [3336]. Improving [1250, 1139, 2722, 3787, 4051, 3948, 3248, 3268, 4658, 4372, 1308, 2521]. improvised [2521]. impulse [2099]. impulses [4286]. imputation [3336, 3647, 4458, 1972, 3014, 1914, 2488, 2593, 3529, 3365, 1361, 4517, 2989, 3206, 2788, 3040, 3963, 1468]. Imputed [4642, 3117, 4373]. in-control [1320]. in-depth [3356]. in-sample [3019]. INAR [4650, 4300, 4135, 2120, 2369]. incentive [2215]. incentives [2141]. incidence [4582, 3514, 4510, 2377, 3614, 4526, 2303, 2730]. including [2222, 3335]. inclusion [1105, 2937]. Income [4501, 212, 2361, 4370, 3338, 3806, 3201]. incompatible [1342]. Incomplete [4517, 458, 3213, 2003, 1225, 1031, 2951, 2739, 1338, 1914, 3747, 3458, 2875, 1224, 3012, 3617, 280, 316, 2041, 3265, 2801, 4385]. Inconsistency [155, 84, 174]. Incorporating [2761, 4655, 3490, 1950, 3352]. Increase [1641, 1233, 779, 3659]. increased [3129]. increases [1976, 2120]. increasing [3235, 3520, 3415, 4225]. indebtedness [2280]. independence [449, 3665, 2171, 3722, 1886, 3778, 3652, 2906, 3681, 3670, 491]. Independent [2977, 2071, 3267, 3213, 4469, 2957, 2365, 2773, 1773, 2196, 2178, 4051, 3075, 3466, 3531, 321]. Index [822, 876, 878, 946, 1455, 642, 1430, 410, 4640, 3769, 2245, 3399, 81, 2235, 1354, 1270, 1923, 3445, 3865, 3663, 2686, 3555, 2083, 3835, 792, 1276, 3804, 3011, 2736, 4021, 1469, 2864, 2640, 3965, 4085, 2973, 1976, 892, 3461, 3961, 3281, 2871, 2511, 3361, 2251, 2554, 2862, 941, 4020, 3338, 1814, 3820, 3934, 4412, 3589, 2201, 1921, 2642, 3580, 3486, 3651, 4399, 2491, 770, 3309, 4381]. indexed [780, 1041, 4552, 718, 2867, 1090, 819]. indexes [2585, 4404, 690, 4003, 920, 921]. India [2280, 845, 915, 1159, 3929, 160, 3524]. indicator [552]. Indicators [2532, 4121, 3170, 4496, 2914, 2551]. Indices [1446, 1690, 1842, 758, 4177, 3985, 2563, 2263, 346, 619, 2893, 2469, 1803, 4002, 1942, 2838, 630, 1604]. Indirect [3185, 3419, 2905]. Individual [2872, 2262, 2647, 2076, 1595, 1209, 1225, 2864, 733, 2252, 723, 3842, 1224, 3846, 1213, 3657, 2666, 1208, 376, 3494, 3498]. individual- [1209]. individual-based [733]. Individual-level [2872, 3842, 3657]. individualized [4646]. individually [3090]. individuals [3537, 762, 2190, 1924]. Indonesia [3163, 4186, 3858]. Indonesian [3972]. Indrayan [2817]. induced [4001, 2279, 2490, 3731, 4022]. inducing [4516]. Industrial [152, 1638, 165, 27, 2713, 1930, 1469, 1145, 346, 619, 3360, 1147, 392, 59, 1150]. industries [235, 2893]. Industry [4, 4344, 4184, 1144, 2435, 4371, 3401, 1908, 1148, 2265, 4116]. inefficiency [3346]. Inequality [1417, 4195, 2563, 2438, 4154]. inertia [1456, 3145]. Infant [2998, 1907, 1574, 1126, 2690, 2401, 3059]. infarction [3537]. infected [2113, 2000, 3664, 4542]. Infection [1689, 4538, 3112, 2493, 3762]. infections [3383, 4611, 2212, 2523]. infectious [2409, 4397, 2591, 3762, 2478, 2388, 2855]. Inference [4004, 3460, 1669, 155, 1631, 3349, 4483, 4354, 3962, 2241, 469, 3404, 2883, 3867, 3947, 4313, 4735, 4722, 4497, 1947, 4006, 2523, 1656, 1712, 2799, 3409, 1302, 4406, 4359, 1809, 900, 4152, 4013, 3427, 2102, 3997, 3780, 3039, 4550, 1781, 2240, 1980, 3541, 2926, 1542, 2507, 4211, 4365, 4019, 4633, 4176, 807, 2925, 4515, 4498, 3636, 2729, 4266, 3114, 3734, 2985, 1597, 1878, 2497, 2485, 4240, 2910, 3102, 3652, 3819, 2454, 3438, 3341, 3037, 4612, 2518, 3748, 3646, 1771, 2930, 4361, 4314, 2630, 1304, 4064, 4254, 1157, 1096, 2908, 4412, 2016, 3955, 1550, 4196, 2426, 2898, 706, 2150, 4362, 4661, 3466, 3177]. inference [3583, 4653, 3885, 4459, 4399, 2820, 2395]. Inferences [4426, 3556, 2607, 1857, 3763, 1513, 1385, 1673, 2999, 2868, 2313, 2238, 2659, 3515, 3266, 4085, 1254, 3826, 3810, 2571, 3378]. Inferential [733, 1446, 1690, 4613, 3531]. inferiority [1783, 4464, 3192, 2062]. Inferring [3213]. infinite [2337, 4674, 1870]. Inflammation [2911]. inflammatory [863]. Inflated [3448, 4098, 4072, 3270, 3888, 4331, 3859, 3574, 3788, 2766, 4347, 3600, 4207, 3009, 3042, 4089, 3989, 2360, 3114, 2596, 4069, 3323, 3896, 1601, 3884, 4300, 4352, 4357, 3451, 3505, 1808, 3489, 3293, 2473, 2624, 3992, 3932, 1058, 1164, 2856, 2122, 2700, 3885, 2812, 4368]. Inflation [2107, 4441, 3908, 107, 3002, 50, 3579, 3271, 2046, 1933]. Influence [3331, 1289, 1211, 329, 1286, 1442, 1728, 1042, 2159, 1131, 2710, 1175, 1694, 2132, 3870, 2403, 2218, 2371, 1688, 3220, 2921, 328, 958, 2290, 3294, 3340, 3391, 3187, 2547, 1260, 3042, 2372, 607, 1052, 810, 889, 1509, 1503, 3469, 1343, 3363, 3265, 3386, 2983, 3761, 2839, 3165, 699, 1316, 2856, 2118, 704]. influences [328, 4096]. influencing [1134]. Influential [330, 332, 355, 331, 1706, 335, 4159, 2678, 3061, 495, 406, 1524, 694, 1063, 2053, 2852, 3189, 4080]. Influenza [1472, 4422, 2872]. Information [1625, 63, 1454, 128, 3695, 1698, 3988, 1964, 900, 1125, 4458, 4007, 2249, 2901, 4509, 679, 4620, 901, 234, 3405, 1130, 743, 3650, 2943, 2937, 392, 1884, 675, 4577, 3476, 2523, 4526, 722, 4401]. information-based [2901]. informative [2186, 3818, 2372, 3621, 3596, 4555, 4298, 3178, 3642, 3713]. informatively [4651]. infra [662]. infra-red [662]. infrared [2964, 1336, 4604]. inheritance [2527]. inhomogeneity [1549]. inhomogeneous [4638, 2927]. initial [1847, 1548, 2001, 4326, 4022, 1876]. initiation [2019]. injecting [3001]. injuries [2387]. Injury [1709, 2449]. ink [2589]. innovation [4344, 4135, 4628]. Innovations [4714, 4438, 3813, 4300, 4067, 3721]. inpatients [4034]. input [601, 3102, 3219]. insect [266, 3966]. insensitive [3080, 2331, 4454]. insertion [609]. insight [3566]. Insights [2919]. inspected [1180, 861, 935]. Inspection [145, 156, 1388, 1558, 553, 84, 357, 1192, 1841, 1273, 1310, 885, 719, 3612, 252, 392, 3615, 576, 905, 963, 861, 1329]. Inspections [1683]. Inspired [207, 4040]. instantaneous [4615]. institutes [3908]. Instrumental [4666, 2571]. Instruments [3830, 3205, 4592, 4604, 4229]. Insurance [4732, 4702, 4163, 2970, 3663, 3788, 3639, 2620, 1512, 2953, 4607, 1987, 1859, 3675, 3772, 4186, 2265, 2987, 3996, 1349, 2628, 3602]. insurances [301]. INT [561]. intake [4252, 4282, 3481]. Integer [2737, 2692, 4438, 4645, 4223, 4006, 4346]. Integer-valued [2692, 4438, 4645, 4223, 4006, 4346]. integral [3855, 896]. Integrated [4325, 1697, 1005, 2797, 4049, 3511, 4478, 1093, 2742, 3958, 1036, 1109]. integrated-cointegrated [1109]. integrating [4526, 4450, 4541]. integration [2099, 2998, 1370]. Integrative [4709, 2089, 4403, 3811]. intelligence [2534]. intelligent [394]. intensities [4488]. intensity [3555, 886]. intensive [4232]. Inter [849, 1792, 2716, 2340]. inter- [2340]. inter-limb [2716]. inter-tree [1792]. Inter-university [849]. interacting [2708]. interaction [4495, 542, 2186, 2487, 3795, 4309, 3657, 505, 4556, 2058, 3492]. Interactions [1412, 3694, 245, 2198, 578, 906, 4605]. interactive [423]. intercept [1291, 4665, 1811, 1509]. intercepts [3247]. intercropping [356]. interdependence [2676]. Interdependency [3571]. interdependent [1126]. interdigital [569]. Interest [1617, 2968, 1500, 3930, 1753, 2791]. Interest-rate [1617]. Interestingness [1440]. interface [519]. Interlinking [10]. intermediaries [1717]. Intermediate [4706]. intermittency [4097]. internal [2416, 234, 2255]. International [1614, 2604, 1037, 3309, 1173]. Internationals [3929]. Internet [2203]. interpersonal [3714]. Interpoint [3905]. Interpolation [1639, 1039, 343, 1778, 2785, 3216]. interpolator [3576]. Interpretable [1527, 4566, 3662, 2586]. Interpretation [1583, 830, 2437, 393, 703, 245, 3103, 518, 3746]. interpretatively [2008]. Interpreting [404, 224, 2322]. interregional [3250]. interrelation [2020]. interrupted [2726]. intersection [2198]. intersection-union [2198]. intersections [4409]. Interstate [1653]. Intertemporal [2628]. Interval [2363, 211, 1967, 3089, 2128, 1001, 574, 1615, 1868, 1654, 953, 4078, 1032, 4426, 2804, 2307, 1820, 1535, 4375, 2329, 3784, 3349, 778, 4596, 1294, 1310, 1966, 4075, 3244, 1042, 4514, 4113, 3965, 1277, 4635, 894, 2576, 3508, 3983, 3879, 1160, 719, 3392, 2454, 2207, 3947, 4651, 2738, 3893, 3727, 3101, 3321, 3964, 613, 1194, 2381, 994, 1865, 1358, 2006, 4671, 3800, 3882, 4685, 2865, 3475, 4107, 4425, 4110]. interval- [1294]. Interval-censored [4078, 3947, 4651, 3893, 994, 4671, 3882, 4110]. interval-screened [2576]. interval-valued [3101, 3321, 3964, 4685, 4425]. Intervals [103, 2430, 1873, 2384, 4268, 4499, 3420, 3301, 2129, 2333, 3804, 2753, 2414, 2492, 827, 972, 2735, 1490, 980, 1344, 2134, 3151, 1858, 2446, 487, 4210, 3718, 4380, 2707, 956, 2784, 3236, 4337, 1533, 4093, 2588, 2565, 1557, 338, 1474, 4083, 4621]. Intervention [3383, 611, 2428, 1371, 837]. interventions [1782, 1005, 4547]. Interview [2604, 2907, 1364]. interview-reinterview [1364]. interzone [2342]. intra [2340]. intra-agreement [2340]. intraclass [2715, 3562, 2934, 3056, 4337]. intracluster [4668]. intraday [2647]. Intriguing [3655]. Intrinsic [4516, 1548]. Introduction [2321, 4705, 4734, 2392, 2750, 4737, 4746, 1899, 1863, 4669, 2460, 2949, 2749, 2502, 2156, 2652]. Introductory [4713, 4692, 4700, 2192, 2276, 340]. intuitionistic [3513]. intuitive [3115, 2960]. invalid [3073]. INVALSI [3140]. Invariant [571, 599, 2285, 1909, 1045, 4444]. invariants [453]. invasion [3702]. invasive [4368]. inventory [3542, 3387, 3076]. Inverse [4324, 579, 1449, 4382, 3219, 2102, 4052, 1967, 3745, 3648, 3891, 2299, 4503, 3407, 1597, 4215, 4239, 4610, 3748, 1306, 2843, 3175, 3928, 4087, 3058, 2630, 3343, 3032, 3772, 685, 587, 2669, 4094, 4481, 1564]. inversion [422, 4016]. Inverted [1608, 3440, 2935, 3862, 3397, 3916, 2924]. inverting [447]. Investigating [757, 3445, 3595, 3168, 4635, 3720, 3238, 4053, 3719, 2523, 2146, 893]. Investigation [1956, 4384, 3795, 721, 390, 2035, 2933, 22, 1109, 4330]. Investigations [1239]. investment [3320, 2114, 2997]. investments [3797, 3304]. involving [482, 506, 640, 412, 4261]. IR [1336]. Iran [3614]. Iranian [2768, 925]. Ireland [2342]. Irregular [1441, 894, 2510]. irregularly [4653]. Irrelevance [237]. Irreverent [169]. Irreversible [1614]. IRT [2305]. ISBN [2820, 2822, 2816, 2815, 2817, 3572, 4669, 2745, 2824, 2819, 2823, 2821, 2818, 2747, 4599, 2653, 2748, 2651, 2652]. ISBN-13 [2820, 2822, 2816, 2815, 2817, 2745, 2824, 2819, 2823, 2821, 2818, 2747, 2653, 2651, 2652]. Ising [4410]. Island [3487, 1227]. Isolated [1680]. isolation [2886]. Isotonic [4320]. isotope [929]. isotropic [802, 3436]. isotropy [1122]. Israel [315]. Issue [364, 836, 4148, 4101, 3644, 4563, 3140, 3993, 4533, 4443]. Issues [1670, 1235, 2131, 3842, 2251]. Italian [2262, 3279, 3168, 2966, 3645, 2294, 2279, 1735, 3159]. Italy [3164, 2416, 3169, 1885, 1902, 283, 1570]. item [3698, 2903, 2479, 2534, 2704, 2551]. item-response [2479]. items [4052, 1310, 2027, 3176]. iteration [600]. iterative [728, 2562, 502, 2365, 411, 1540]. IV [4318, 4005, 1108].

J [4669, 2819, 2818, 1313]. J-shaped [1313]. J. [165]. JaB [2873]. jack [610]. jack-knife [610]. Jackknife [3935, 2965, 2347]. jackknife-based [2347]. Jackknifed [3275]. Jacknife [73]. JAGS [4328]. James [3057]. James-type [3057]. Jan [2747]. Jana [2747]. Japan [1674, 4502]. Japanese [1924, 2201, 591]. Jarecová [2747]. Jarque [3662, 1628]. jemmies [67]. Jennifer [2816]. Jin [4599]. job [1353, 3168, 20, 2400]. John [2820, 4669]. Johnson [1104, 1044]. joinpoint [2945, 4519]. Joint [2896, 1426, 3138, 4321, 2294, 4609, 3112, 3831, 3458, 3118, 4069, 3481, 3632, 3339, 3326, 3095, 1473, 4377, 3684, 2546, 2911, 4216, 2584, 3116, 3856, 2570, 2918, 4557, 4226, 4647, 4238, 2553, 2800, 949, 2224, 2912, 3334, 2876, 3818, 3045, 3009, 3507, 4363, 2516, 2917, 4572, 669, 3046, 4662, 4480, 3733, 2525, 2932, 2832]. Jointly [4424, 3484, 4284]. jointness [3577]. Jonckheere [988]. Jonckheere-type [988]. Jones [2665, 4144]. Jordan [2245]. Journal [1715, 2290]. journals [236]. journey [3392]. judgements [2895]. judging [2104]. jump [2790, 4470, 1255, 3421]. Jumps [3929, 3822]. justice [4308]. justification [3977]. justified [48].

K-Bessel [4606]. Kalman [2974, 2991, 4222]. Kanji [2230, 2524, 105, 1715]. Kappa [1686, 4285, 830]. Kappa- [830]. Kaufman [4669]. Keep [2923]. Kendall [4495, 3071, 2518]. Kenya [254, 443]. Kepler [3745]. Kernel [4683, 3681, 3873, 4149, 4240, 4281, 4490, 3546, 1026, 4212, 4626, 4141, 2792, 4294, 2514, 3897, 2311, 3074]. kernels [2925, 2566]. key [4197]. keys [4492]. kidney [3184, 3339, 3478]. kind [4114, 46, 66, 89, 76, 100]. kinds [2205]. kinetic [2962]. kinetics [4341]. King [1711]. Kingdom [3930, 3621]. kings [3806]. kite [1241]. kittens [46]. kittiwake [1211]. knee [3446]. knife [610]. Knockoff [4569]. knockout [908]. know [1]. knowing [3078]. Knowledge [513, 2296, 519, 3698, 4411, 512, 1005]. Knowledge-based [513, 1005]. Known [145, 156, 1415, 1965, 3539, 3475, 2527]. Knox [1406]. Kohonen [3362]. Kolmogorov [1265, 1675, 636, 846, 1062, 2126, 1687, 2840]. Kong [3906]. Korea [923, 3906]. Kozak [2176]. kppm [4638]. kppmenet [4638]. Kriging [4057, 471, 3933]. Kronecker [1119]. Kruskal [3593]. Kuiper [372]. Kullback [3317, 1698, 675, 4008]. Kumar [2819]. Kumaraswamy [3862, 4313, 3922]. kurtosis [3655, 3662, 1477, 2417].

L [2821, 2200, 2570]. L-moment [2570]. L-moments [2200]. label [1045]. label-invariant [1045]. labels [4506]. labor [4331, 4418, 2033]. Laboratory [153]. labour [3035, 2294, 2656, 500, 88, 2577, 2153]. lack [4339, 1337, 1723]. lactation [1837]. LAD [2107, 4045, 3812]. LAD-LASSO [4045, 3812]. ladder [3211]. Lag [1991, 1760, 3876, 3003]. lag-length [1760]. lagged [4090, 461]. Laguerre [1794]. Lake [86]. lambda [3593]. lameness [2352]. lamination [3833]. Lamothe [2651]. land [3869, 1372, 3949, 446]. landbirds [759]. landmark [1045]. Langevin [371]. Language [1382, 408, 1300, 677]. Lanka [1840]. LANZA [3938]. Laplace [490, 4478, 1636, 2164]. Laplace-normal [490]. Large [2507, 2446, 2600, 1680, 1678, 1819, 2033, 2967, 2296, 3978, 1277, 827, 2364, 3925, 1938, 4201, 2215, 3603, 753, 4496, 2368, 3635, 1032, 4658, 4439, 1813, 2569, 4453]. large-dimensional [3925, 4439]. large-sample [1819, 1277, 1032]. Large-scale [2507, 2600, 1938, 2569]. larger [1111]. larval [482]. laser [662]. Lasso [3527, 3812, 2597, 4420, 3565, 4111, 2812, 4453, 4045, 4010, 4179, 2830, 2841, 2878]. lasso-regularized [4453]. Lasso-type [3527]. last [3178]. lasting [2523]. late [1997, 2832]. late-life [2832]. Latent [4549, 2789, 1644, 1990, 3828, 979, 2361, 2582, 4543, 3169, 4598, 2019, 2740, 3313, 4357, 2490, 3860, 1895, 2483, 3422, 3123, 1304, 3293, 2177, 2947, 2704, 1499, 2955, 4071, 3963, 3238, 4229, 4494, 3220, 4250, 1593, 3202]. lateral [4241]. latin [1331, 1176, 4057, 405, 4145]. latitude [4442]. Lattice [547, 2382]. Lattice-oriented [2382]. lattices [4051, 716]. Law [1616, 1422, 3869, 262, 852, 2708, 886, 1645, 1610, 129, 164]. laws [2112]. layer [4445]. Layered [3558]. layout [541]. Layouts [1666, 505]. LDA [4450]. lead [929, 459, 572, 700]. leadership [3915]. leading [1335]. leaf [4570]. league [2182, 952]. Leagues [4121]. lean [4118]. leapfrog [1081]. learner [4462, 3824, 4459]. Learning [3407, 4632, 4344, 3694, 4643, 4461, 4584, 4451, 2288, 3362, 4178, 3004, 4646, 4504, 3686, 1805, 3554, 2636, 4568, 4452, 4586, 3041, 4443]. Learning-based [3407, 4568]. Least [1646, 1002, 1448, 1507, 4387, 2995, 979, 2797, 4198, 2249, 4509, 2481, 3292, 2018, 1590, 1162, 305, 3131, 4242, 3607, 3444, 411, 321, 3053, 3584, 1498]. least-square [4387]. least-squares [2995, 979, 2481, 3131, 3607]. leave [3356]. leaves [814, 692]. led [2717, 3003]. Lee [2748, 4245, 1521]. left [1997, 3122, 3648, 3650, 4677, 3218, 4106]. left-censored [3648]. left-censoring [4677]. left-truncated [3122, 3650, 3218, 4106]. legal [962]. legislative [4588]. Lehmer [4020]. Leibler [3317, 1698, 675, 4008]. Leiden [3093]. leishmaniasis [1948].

Lemeshow [4658]. Length

[1693, 2171, 1595, 1141, 1905, 1311, 1042, 1760, 4082, 4503, 1720, 3707]. length-bias [1311]. length-biased [4082]. lengths [4083]. Lenth [3779]. Leonard [4669]. Leone [1508, 4497, 1802]. Lepage [1265]. Leptograpsus} [2136]. leptokurtically [3212]. lesion [3716]. lesser [752]. Letter [454, 259, 124, 131, 114, 139]. Letters [105, 241, 176, 94, 68, 41, 54, 30]. leucopsis} [742]. Level [1619, 856, 2733, 3460, 1830, 1029, 1059, 3254, 4119, 2280, 2603, 717, 1404, 2269, 4339, 3773, 4214, 2689, 772, 1441, 506, 565, 637, 913, 964, 1041, 617, 3842, 4201, 1634, 567, 3443, 3657, 2803, 2872, 3807, 4263, 2040, 3752, 1983, 718, 2919, 984, 2477, 3087, 484, 460, 496, 911, 1879, 3188, 4080, 963, 282, 2104]. levels [793, 1201, 1369, 1285, 1202, 1575, 4616, 646, 321]. Leverage [2174, 1867, 2734, 3189, 4561]. leveraged [2717]. Lévy [3935, 4686]. lexical [3888]. Lexigenesis [1300]. Li [4599]. Library [149]. LIC [4514]. Life [1637, 78, 126, 4694, 2491, 3249, 1674, 3907, 4139, 2585, 1906, 2032, 3455, 3923, 4354, 1228, 469, 2457, 3631, 4140, 1121, 3305, 2899, 480, 2330, 3371, 3577, 3143, 508, 753, 881, 3916, 3886, 4618, 2013, 2987, 4094, 1580, 808, 1815, 2832]. life-history [1228]. life-line [508]. Lifetime [1422, 2687, 3769, 2805, 3395, 1809, 3865, 3836, 3094, 2299, 3965, 3020, 816, 2672, 2522, 1172, 2490, 2796, 1868, 1907, 4271, 3111, 2202, 1175, 3269, 1110, 1804, 3729]. lifetime-performance [3865]. lifetimes [828, 3650]. Liga [4304]. light [3890, 3032]. like [3998, 4610]. Likelihood [2971, 2214, 1661, 3646, 1379, 1677, 2473, 4394, 3876, 2197, 4144, 2051, 3193, 2733, 3081, 463, 2102, 3935, 4233, 1953, 3254, 4579, 728, 927, 3999, 778, 4655, 615, 2313, 2965, 464, 1913, 2301, 1037, 2318, 2849, 805, 3318, 4620, 1269, 79, 3956, 3738, 4066, 3843, 3507, 3734, 2187, 1938, 2752, 381, 2581, 665, 3397, 2431, 2398, 3073, 4104, 1774, 3650, 3855, 4351, 1056, 1326, 1107, 2397, 3098, 3343, 316, 4323, 4285, 4208, 1157, 613, 1175, 4589, 4008, 1133, 2381, 2270, 2209, 2898, 3953, 2638, 808, 3257, 2150, 3055, 549, 1199, 3516, 4591, 3668, 4372, 2571, 3378, 3463, 4110, 2376]. Likelihood-based [2214, 3646, 3876, 2381, 2898, 3516, 3463]. likelihoods [1281]. limb [2716]. Limit [1423, 1018, 1042, 4630, 3474, 2197, 4420, 1111, 2449]. limitation [3529]. limitations [4295]. limited [2988, 2697, 4066, 4136]. limiting [506, 565, 637, 567, 718, 646]. Limits [1403, 360, 1285, 3879, 2469, 3467, 1511, 1233, 838, 2283, 4115]. Lindley [3921, 4359, 4156, 2677, 4687, 4155, 3648, 4678, 3923, 3889, 3710, 3743, 3920, 4417, 4382, 3955, 3501]. Lindley-binomial [4678]. line [2296, 508, 3421]. Linear [1669, 2779, 147, 1423, 4485, 1665, 4132, 3066, 2459, 4194, 2407, 1605, 1389, 1608, 670, 1636, 1689, 4717, 1632, 1400, 3732, 3496, 4702, 3331, 470, 2881, 4279, 3028, 3072, 1510, 3679, 2082, 3917, 806, 2988, 335, 2676, 548, 872, 1899, 3658, 1948, 2880, 2873, 277, 314, 337, 4233, 3491, 4587, 3609, 1273, 3224, 3479, 3653, 4611, 398, 391, 442, 482, 680, 2926, 4159, 3277, 1247, 2367, 3755, 1149, 2464, 3244, 3063, 4214, 2887, 571, 2682, 3523, 3470, 4484, 2018, 1260, 729, 3985, 3150, 2549, 2285, 600, 4165, 4160, 2453, 642, 2599, 788, 4066, 1867, 3137, 2045]. linear [1840, 644, 396, 2808, 3734, 2187, 3283, 3366, 1524, 1852, 2004, 2883, 3413, 3711, 3247, 640, 4393, 3135, 3406, 3398, 3771, 2679, 2221, 3778, 2234, 889, 1737, 682, 4079, 4683, 2365, 1519, 4408, 2205, 2433, 3527, 1860, 1857, 3130, 3285, 2421, 2558, 2916, 673, 1095, 4627, 2024, 1823, 2242, 4074, 3646, 3231, 992, 2727, 3588, 3948, 4361, 3967, 1987, 2392, 2397, 2174, 1163, 2852, 3189, 2750, 3849, 461, 4391, 2892, 4323, 2831, 3478, 1473, 618, 3684, 1562, 4292, 3386, 672, 1990, 493, 1739, 3396, 1371, 1004, 592, 1734, 1065, 2275, 4062, 4616, 558, 3352, 389, 587, 4353, 2786, 888, 399, 2371, 4060, 2639]. linear [2223, 1303, 1337, 1723, 2781, 3208, 3459, 549, 2271, 3501, 2449, 1316, 1995, 2138, 3713, 4689, 4133, 4392, 1349, 3979, 4180, 704, 4447, 1308, 1540]. linear-circular [4587]. linearity [1544, 3186, 1097]. linearization [2561, 2213]. linearized [4310, 4010]. linearizing [1043]. linearly [2302]. Lines [561, 3766, 2626]. LINEX [4421]. Linguistic [1610, 1863, 1502]. linguistics [3888]. link [3215, 3398, 3335, 3486, 1202]. link-free [3486]. linkage [1262, 4349, 2448]. linked [4448]. links [1727]. Liquidity [1962]. Lisbon [1850]. List [3701, 800]. listed [3128]. literacy [3305]. Literary [1711, 1008, 1461, 1592]. literature [1008]. Lithuania [2180]. Liu [2481, 3991, 4599, 1030, 2678, 2590, 3225, 4162]. Liu-type [2481, 4162]. live [4049]. liver [3986, 3796]. livestock [1729]. Living [198, 3941]. LM [1340]. LND [2720]. load [2647, 3547]. loading [881]. Loadings [703]. loads [4231]. loan [2637]. loblolly [1792]. Local [3035, 4485, 702, 1052, 810, 889, 1509, 1366, 3469, 1343, 1026, 1694, 741, 3265, 3386, 2927, 2303, 2118, 2921, 958, 1551, 2048, 931, 3695, 2511, 2289, 2983, 47, 3761, 1316, 4591, 1738]. local-to-unity [1551]. localized [3643]. locally [599, 3570, 4485]. located [602]. locating [2783]. Location [3966, 1390, 3235, 4647, 3816, 4126, 463, 3232, 3423, 3051, 964, 3325, 4122, 2683, 2580, 3505, 4009, 3363, 4216, 672, 1982, 96, 581, 4463, 4071, 4401]. Location-scale [3966]. loci [2448]. locus [4468]. Log [4331, 564, 1555, 4479, 673, 1964, 1965, 3753, 4043, 1156, 1009, 391, 4418, 2335, 3402, 2715, 4478, 1442, 3038, 642, 1087, 1121, 4466, 2557, 2773, 1868, 3763, 3881, 2892, 3548, 3363, 3148, 2204, 2898, 4094, 4481, 4350, 4671, 2999, 4372]. log-Birnbaum [1442]. log-Burr [2204]. log-Dagum [3753]. log-exponentiated [3548]. log-Gaussian [4478]. log-likelihood [4372]. Log-linear [673, 391, 642, 2892]. log-location [3363]. log-log [4466]. Log-logistic [564, 1555, 1965, 4043, 1121, 1868, 4094, 4481]. log-logistic- [4350]. log-logistic-G [4671]. Log-mean [4479]. Log-Normal [4331, 1964, 1009, 2335, 1087, 2557, 2773, 3763, 3148, 2898, 2999]. log-normally [3402]. log-symmetric [4418]. Logarithmic [4507, 3323]. logging [260]. logic [2897]. logical [408, 677]. logicFS [2897]. Logistic [1520, 1060, 1636, 2845, 4114, 1356, 1965, 3336, 2913, 3556, 4043, 2842, 1294, 4023, 3892, 4483, 3630, 585, 1584, 1001, 3986, 564, 2734, 4593, 4660, 1121, 1555, 4063, 3465, 682, 4283, 2282, 3397, 3186, 2597, 2067, 1868, 1915, 3166, 4009, 3098, 2053, 3382, 1014, 3756, 3820, 1916, 1798, 2846, 3394, 4162, 1499, 3291, 3918, 4094, 3296, 4481, 4350, 4671, 3549, 4429, 4261, 2760, 2173, 3900, 4372]. Logit [4343, 615, 3700, 442, 1496, 2595, 1572, 2656, 3983, 2111, 2958, 1304, 339]. logits [1053]. LogLindley [4296]. LogLindley-Binomial [4296]. lognormal [1975, 3359, 3403, 2865]. Logratio [2827, 3245]. logratios [4584]. Loïc [2824]. Lomax [4163, 3965]. Long [1643, 3146, 1850, 2235, 3199, 4043, 1901, 1782, 1219, 4241, 3594, 1033, 3280, 2490, 2601, 2141, 2523, 3003, 903]. long-lasting [2523]. Long-range [2203, 3199]. long-run [2141, 3003]. long-tailed [1033, 3280]. Long-term [3146, 1850, 2235, 4043, 1219, 4241, 3594, 2490, 903]. longevity [4174, 2850, 508]. Longitudinal [2479, 2900, 4611, 2000, 2876, 3818, 4396, 3327, 3424, 4377, 1432, 1647, 1918, 4640, 4367, 499, 3575, 3917, 3184, 4296, 2896, 3510, 4238, 3767, 3982, 2941, 3207, 2912, 2476, 2003, 2186, 3334, 3117, 4321, 1877, 2656, 3678, 2833, 2690, 1312, 3114, 1778, 3507, 947, 4363, 3366, 3831, 2516, 3747, 3750, 3878, 2917, 3374, 3365, 3674, 4572, 2214, 4232, 3461, 3339, 3326, 3484, 3376, 2518, 3860, 3654, 3095, 2405, 2592, 3272, 3209, 2512, 3964, 2546, 2911, 3047, 3265, 2963, 3293, 3287, 2473, 3046, 1960, 1874, 3093, 3642, 3992, 3761, 3116, 2197, 3963, 3856, 2854, 4429, 1995, 2932, 2022, 2278, 2480, 2832, 3202, 4253]. longitudinal [4453, 4557, 3132, 2975, 4447, 4385, 4465]. longitudinally [1307]. Longtailed [1208]. Look [184, 1284, 3772, 777, 2480]. looking [662]. Lorenz [345]. Loss [1423, 713, 1088, 3829, 1446, 1618, 1702, 3769, 4529, 1018, 1102, 3182, 2632, 4421, 4076, 4345, 4607, 1734, 1048, 4311, 1839, 4525, 969, 1474, 4154, 4680]. Loss-based [1088]. losses [2970, 3996]. lost [2190, 3723]. Lot [1388, 320, 1086, 1041, 3612, 4263, 392, 905, 963, 1099]. Lot-size [1388]. Lots [1409, 1680]. Lottery [1098]. Lotto [1668]. low [3703, 3805, 285, 635, 4010, 3927, 4489, 2493, 4403, 2928]. low-dimensional [4489]. low-energy [635]. low-lumbar [2493]. low-rank [4010, 3927, 4403]. low-sample [2928]. Lower [2548, 1486, 4421, 2469, 1111, 898, 899]. lows [1920]. LQL [643]. LS [3825, 3825]. LS-SVR [3825]. LSTM [4445]. LTS [2590]. lumbar [2493]. Lund [399]. lung [2829, 2945, 344, 3004, 2449].

M [2822, 2815, 52, 510, 2748, 1030, 3520, 2189, 963, 1375]. M-estimator [1030, 1375, 2189]. M-shape [3520]. Machine [4344, 199, 4443, 4643, 2797, 4461, 3381, 4451, 4264, 4178, 4281, 3686, 2792]. machines [2133, 4455, 2282, 3004, 1106]. MacMahon [262]. Macro [1454]. macrobenthic [1517]. macular [2405]. Made [181, 2458]. Madrid [3797]. magnitude [3256, 4046, 931, 3354]. magnitude-frequency [3354]. magnitudes [3571]. Mahalanobis [1526, 3490, 2537, 1747]. mail [3846]. main [4384, 2487, 578, 529]. Maintaining [2440]. maintenance [440]. maize [609]. major [867, 2832, 952]. make [2667]. Making [162, 3513, 3206, 2837, 2496, 4568, 3013, 1140]. malaria [2094, 2190]. Malawi [4511].

Malaysia [3910, 2478, 2717]. malicious [2343]. malignant [3716]. Malling [111]. Mallows [3499]. Malmquist [1604]. malnutrition [3327]. malpractice [962]. Malus [1512]. malware [4251]. mammalian [1503]. Man [181]. Man-Made [181]. Management [1644, 4733, 4716, 4691, 760, 2424, 1139, 1891, 2070, 759, 2069, 789, 754, 1514, 786, 3432, 2503, 2194, 2227, 2818]. Manchester [11]. Mangium} [1615]. manifolds [2453]. Manipulation [2090, 4707, 1228]. Mann [2486, 968, 2941]. Mantel [1174, 3099]. manual [1841]. manufacturing [3232, 1841, 1314, 88, 1146]. manufacturing/inspection [1841]. Many [2452, 1454, 3515, 4303, 4592, 2552, 4290, 3176, 4107]. many-facet [3176]. Many-to-one [2452, 4107]. MAP [4269]. mapping [2560, 2094, 2349, 4130, 2478, 2232, 2517, 2795, 2303]. maps [4648, 758]. MAR [3613]. Marc [2824]. margin [2967]. Marginal [3452, 3451, 1549, 2979, 2109, 2701, 380, 3734, 2348, 1749, 3222]. marginalised [3215]. marginality [1035]. Marginalized [3600, 3859, 2357, 2516, 2214]. Marginally [1761]. marginals [2529, 1848, 4087, 1856]. margins [3341]. Marine [1693, 3709, 1517, 642]. marital [2832]. Mark [3572, 736, 748, 1221, 1240]. mark-recapture [748, 1221, 1240]. Mark-resighting [736]. Marked [1703, 759, 762, 2937, 3495]. Marked-point [1703]. marker [3474, 1307]. Markers [2911, 4609, 3507, 3315, 3247, 4209, 3820, 4614, 4420, 1257]. Market [3618, 996, 4169, 3279, 4287, 1962, 2035, 64, 4177, 2976, 2294, 3814, 3929, 2263, 3061, 1544, 3775, 3603, 3626, 1304, 249, 4093, 1890, 2382]. market-opening [4093]. marketing [2076, 4188, 2296]. Markets [1668, 3035, 3595, 4286, 1037, 2675, 1797, 3437, 2020, 1890, 3498, 1586]. Marking [677]. Markov [2228, 4512, 2805, 1444, 4072, 1681, 2239, 3138, 3029, 4097, 3828, 3794, 843, 977, 3894, 1409, 1186, 3158, 3735, 2675, 3535, 1877, 4005, 3408, 3065, 1183, 2775, 947, 2268, 4593, 3327, 1763, 4269, 719, 1064, 2317, 2020, 3707, 3958, 415, 570, 2334, 2336, 1772, 4608, 1255, 3412, 423, 3726, 3228, 2992, 4402, 3622, 4719, 3040, 3963, 3852, 3800, 1015, 1016, 1164, 3808, 1623, 3202, 4034]. Markov-dependent [3408, 1772]. Markov-switching [1183, 2020]. Markovian [3760, 3077, 1973]. marriage [377]. married [1311, 3848]. MARS [3512, 1889]. Marshall [4305, 1530, 3863, 3477, 2936]. martins [755]. masked [4665, 3349]. mass [2235, 4118, 2386, 3361, 2251, 2862, 4311, 3580]. massive [2802, 4660, 3782, 3421, 4449]. Master [3509]. Matatu [254]. match [1886, 981]. matche [4121]. matched [707, 2766, 585, 1572, 1367, 1130, 673, 4112, 2482]. matches [2726, 1911]. Matching [1435, 390, 3214, 4049, 912, 3033, 4204, 2929]. material [717, 3958]. materials [1481]. maternal [3340]. maternity [3356]. Mateu [2818]. math [3172]. Mathematical [4719, 39, 2460]. mathematicians [19]. Mathematics [31, 10, 65, 3425, 80]. MATLAB [749]. Matrices [151, 1979, 2922, 1367, 4280, 3925, 2154, 447, 1303, 2475]. matrix [1125, 1792, 2110, 1011, 4189, 3472, 3904, 2688, 1561, 4010, 1991, 1119, 4403, 1240, 2992, 727, 444, 3041, 4117, 3483]. maturity [1227]. Max [1575, 3666]. Max-min [1575]. maximal [1913]. maximin [2471]. maximization [2740, 3043, 4254, 3219, 4410]. Maximizing [1584, 1420, 471]. Maximum [3193, 4376, 778, 1913, 2318, 1269, 2752, 381, 1661, 1390, 316, 613, 3338, 4145, 333, 549, 4110, 4144, 2881, 3081, 463, 1820, 4233, 728, 615, 1891, 805, 3507, 2581, 665, 3397, 3592, 3650, 4351, 2077, 2559, 1175, 1133, 819, 3953, 808, 3364, 3257, 4591, 2376]. Maximum-likelihood [2752, 4233].

Maxwell [4364, 3574, 4324]. may [947]. MCAR [3613]. MCB [3721]. McCool [2820]. McDonald [3279]. McElroy [2653]. MCMC [1510, 1582, 3945, 4479, 1432, 2405]. MDA [358]. MDS [412, 506]. MDS-1 [412, 506]. Mean [1451, 1666, 577, 1423, 1450, 875, 1624, 1627, 77, 2380, 1873, 3270, 404, 3980, 598, 3420, 2800, 4423, 936, 3456, 1542, 2073, 2481, 4284, 2298, 2085, 2279, 1934, 3014, 3404, 1757, 4030, 2374, 3327, 4479, 1737, 2009, 4610, 4642, 1857, 2773, 3129, 1935, 2291, 1095, 3316, 3763, 3143, 1131, 953, 2397, 940, 2579, 4020, 1983, 1945, 2356, 3916, 3047, 1973, 3454, 2716, 2189, 3044, 3060, 3800, 549, 2120, 2525, 2865, 2932, 3902, 2188]. mean-of-order- [4020]. mean-shift [1737]. mean-squared [3316]. mean-standard [2374]. mean-variance [4030]. Means [213, 1418, 1431, 4499, 541, 3113, 4214, 3705, 1779, 1840, 644, 1833, 3611, 1279, 2735, 2317, 1334, 2557, 3316, 266, 2465, 3253, 2784, 1759, 2373, 965]. Measure [4728, 2348, 3018, 3751, 588, 3355, 3207, 4544, 1888, 1260, 574, 4508, 3305, 1853, 2943, 1023, 4608, 2437, 580, 1120, 1175, 3934, 3938, 4353, 3416, 2450, 909, 467, 3629, 1570]. measured [1307, 4229, 1995]. Measurement [1704, 2029, 1405, 2351, 1706, 4252, 4100, 4665, 3942, 1811, 4267, 1138, 2142, 4282, 1897, 2409, 679, 3009, 2532, 3581, 507, 4085, 3398, 1573, 2604, 1342, 4683, 1509, 2398, 2989, 1984, 1047, 2798, 1529, 4559, 4531, 3264, 3597, 1947, 1065, 4095, 4146, 2449, 3560, 3387, 3845, 704, 3651]. Measurements [1383, 1249, 1278, 2896, 3510, 3942, 1115, 4428, 4466, 635, 1519, 1118, 4222, 3849, 2838, 1350, 2836, 3409, 4557]. Measures [341, 1791, 130, 3152, 1187, 2955, 1688, 2331, 3331, 3616, 2103, 3685, 4615, 2753, 1508, 35, 3000, 1800, 4437, 830, 1119, 2710, 833, 18, 4008, 1739, 993, 1749, 2928, 4634, 4265, 1006, 1599, 2213, 4429, 3861, 4641, 4617, 2730, 4398]. Measuring [2684, 1221, 631, 3857, 3593, 3984, 3830, 1618, 173, 3492, 2199, 372, 2789, 4037, 2662]. mechanical [3322, 1827, 1565]. Mechanics [934]. mechanism [4581, 4137, 3914]. mechanisms [2679, 3806]. media [560]. Median [1089, 1415, 2978, 2451, 3420, 3301, 2809, 3303, 3051, 4161, 2068, 1001, 604, 918, 1195, 2785, 3940, 2531]. medians [4667, 4556, 3887]. mediating [3941]. mediation [4396, 3443, 4080]. Medical [127, 4741, 1377, 2673, 522, 2548, 23, 3039, 3344, 2744, 3582, 523, 3381, 3722, 3127, 1357, 2890, 962, 3366, 1938, 4479, 2431, 2962, 3761, 339, 2817]. Medicare [1511]. medicine [2594, 1861, 2960, 2836]. Mediterranean [768]. medium [2637]. medoids [3890]. meet [317]. megalithic [85]. melanoma [3417]. membership [3185]. memberships [4250]. Membranes [658]. Memory [1643, 1901, 1782]. men [1272, 2493]. menarcheal [4513]. Mental [4377, 4609, 4618]. Menten [971]. merge [2069]. merger [3770]. mergers [2595]. Merging [1625]. merits [3318]. MERVAL [3309]. meta [2545, 1060, 3717, 1910, 2942, 3151, 4558, 3021, 4491, 3553, 2095, 4401]. meta-analysis [2545, 1060, 3717, 1910, 2942, 3151, 4558, 3021, 4491, 3553, 2095, 4401]. metabolic [4595, 4531]. metagenomic [3672]. metagenomics [2434]. metallic [4320]. metallographical [2793]. metapopulation [1229]. Method [1657, 1693, 1638, 1398, 1613, 1389, 1427, 1108, 2304, 2732, 1869, 4516, 2117, 4387, 2873, 515, 1305, 2484, 979, 354, 1009, 2654, 2965, 802, 3717, 2993, 3322, 2269, 2249, 4182, 4411, 3912, 4040, 4598, 530, 954, 3226, 600, 4160, 2759, 4082, 2147, 3472, 2300, 1818, 640, 1878, 4045, 2059, 1938, 3568, 1279, 4519, 2365, 665, 1074, 4010, 1860, 2557, 4504, 1104, 3206, 1306, 3506, 3401, 2788, 1026, 3482, 3754, 4292, 3991, 4416, 844, 3178, 4162, 2847, 1282, 4540, 4158, 3822, 517, 3372, 3055, 3453, 4275, 699, 3531, 321, 3779, 2115, 3261, 4591, 2650, 4392, 1586, 3068, 2623, 4561, 2531, 3486, 2385, 2806, 3052]. Method [2562, 1178]. methodological [1854]. methodologies [4376, 851]. Methodology [1434, 4742, 1652, 2574, 4246, 1217, 4184, 2459, 1218, 3633, 4210, 4127, 1091, 525, 1015, 2747]. Methods [1643, 211, 4748, 2787, 1639, 4733, 2323, 4706, 4690, 4718, 3264, 212, 4747, 4129, 2380, 1478, 3035, 3142, 3081, 1553, 2034, 364, 2257, 763, 2463, 343, 3780, 1212, 4579, 4487, 652, 3804, 3923, 2507, 4339, 4176, 2951, 1287, 2198, 890, 4148, 2519, 3521, 1087, 3380, 3793, 2593, 1277, 2312, 2057, 1763, 2641, 4328, 3688, 3824, 3528, 3062, 2581, 1161, 2398, 2577, 4204, 1988, 1095, 1774, 415, 570, 1590, 3351, 4607, 2837, 3544, 4270, 3752, 1462, 4395, 881, 4644, 2567, 3163, 4476, 3533, 4295, 2818, 1240, 3312, 2992, 2662, 1032, 2477, 1368, 319, 1358, 2425, 2987, 3932, 2952, 2572, 2482]. methods [4605, 1484, 2760, 2188, 4150, 4366, 2303, 3619, 4385, 3656, 3572, 2195, 2948]. methylated [4293]. metric [3978, 1361, 4062, 3584]. metrics [2602, 4125]. metropolitan [3013]. MEWMA [4433, 4635]. MEWMA-CoDa [4635]. Mexico [2969, 836, 1469, 209, 4332]. Meyer [3779, 2623]. mice [3719]. Michaëlis [971]. Michaëlis-Menten [971]. micro [4049, 666]. micro-texture [666]. microaggregation [3017]. microarray [2337, 1910, 1718, 2465, 2362, 2164]. microbial [3672]. microcomputer [264]. Microcomputers [222]. microdata [2907, 4055]. microelectronics [934]. microenvironments [1261]. microorganisms [4266]. microRNA [2990]. microsatellite [2243]. microscopy [2096, 417]. microstructure [4320]. microtubule [1875]. middle [3020]. middle-censoring [3020]. Midlands [1472]. might [3814]. migrants [2416]. Migration [4259, 3065]. migratory [759, 2591]. MIL [252, 621]. MIL-STD-105D [252, 621].

Milevsky [4402]. Militarized [1653]. military [1476, 1836, 3076]. milk [3481]. Mill [1651, 1310]. min [1913, 1575]. mineral [2493]. ‘Mini [92]. Minimal [1350, 3887]. Minimax [2813, 590, 1609, 2343]. minimise [2521]. Minimizing [4154, 1917, 1589]. Minimum [1180, 1388, 2808, 1252, 718, 1680, 861, 935, 4561, 1092, 2447, 3773, 506, 3878, 1938, 2497, 3017, 2846, 412, 839, 2169]. Mining [3846, 2509, 2578, 2610, 3725, 3126]. Minitab [1723]. minority [3076]. Minute [480]. Minute-Accumulating [480]. mis [2101]. mis-specification [2101]. Misclassification [1696, 2091, 2879, 3108, 3581, 1116, 4078, 358, 991, 3659, 3619]. misclassified [3630, 3037, 3823, 1744, 4410]. Mises [374, 1265, 472, 978, 3426, 2494, 382, 456, 1974, 2178, 2175]. misinterpretation [2293]. mismeasured [3484]. misses [1177]. Missing [2237, 3206, 1176, 3351, 4640, 4367, 4047, 1289, 3350, 3828, 4461, 4458, 3735, 2926, 4329, 1085, 3581, 3014, 3557, 3877, 3330, 3500, 502, 2679, 4297, 4673, 3365, 3674, 3975, 2396, 1117, 4408, 4470, 2398, 2989, 2084, 3603, 2511, 2550, 3893, 3684, 2183, 3756, 2963, 4597, 2177, 540, 411, 4071, 2146, 773, 1262, 1468, 3463]. missing-at-random [4640]. missingness [3859, 2833, 2711, 4232, 3484, 3654, 3351, 3914, 2832]. Misspecification [1662, 4429, 2237, 1275, 4262]. mist [758]. mist-netting [758]. mitochondria [658]. MIVQUE [3443]. mix [4188]. Mixed [1669, 2909, 1689, 2389, 4025, 1432, 3093, 3157, 3235, 3085, 4156, 1510, 3816, 4100, 3917, 4039, 4296, 2896, 2262, 2971, 3224, 4438, 2912, 2926, 3314, 4214, 492, 2887, 3690, 2561, 4620, 3678, 4498, 788, 3734, 2300, 3413, 3711, 3247, 4393, 3644, 4079, 4683, 2365, 1519, 4408, 1860, 3285, 1800, 2989, 3632, 3376, 3654, 4602, 2850, 2242, 3966, 2410, 2174, 3272, 4011, 3849, 3382, 3478, 3545, 3790, 2061, 3386, 3293, 1739, 2177, 2947, 1499, 1998, 1866, 2467, 4095, 3464, 3370, 4071, 3116, 2799, 1959, 3459, 3195, 282, 1316, 1349, 3979, 4111, 2946, 3202, 3629, 4110, 4465, 2170, 4527, 4179]. mixed- [4100, 4095]. mixed-data [2170]. Mixed-effect [3093, 3157, 4156, 3314]. Mixed-effects [4025, 2262, 3224, 2561, 4498, 788, 2989, 3376, 2410, 3545, 3790, 1499, 2467, 4527]. mixed-mode [492]. Mixed-Poisson [2389]. mixed-POT [3690]. mixed-Weibull [2300]. mixing [3433, 4650, 2052, 4182]. Mixture [2286, 603, 926, 1623, 1425, 997, 2292, 2845, 3503, 2810, 3433, 2337, 1354, 3751, 3658, 2484, 3899, 948, 843, 3229, 1944, 2222, 4509, 1832, 3298, 2887, 2162, 2705, 2682, 2723, 2690, 1127, 2247, 4030, 2133, 2566, 3407, 4603, 3643, 250, 489, 3062, 2672, 1522, 2621, 3313, 3561, 3909, 4518, 2493, 4440, 2982, 4361, 3783, 2630, 2055, 2136, 3964, 2972, 3790, 1562, 834, 1770, 2219, 3851, 4476, 3697, 1939, 1004, 1122, 2523, 4091, 3233, 4115, 3789, 4258, 3963, 903, 4436, 1831, 2047, 2513, 3799, 4027, 4360, 3730, 3625, 3442, 4465]. mixture-based [3658]. Mixtures [3816, 2520, 2774, 3242, 4273, 836, 1953, 3234, 1156, 3182, 2471, 4633, 3962, 3150, 4194, 3813, 2494, 490, 2234, 4552, 3646, 2937, 3361, 3273, 1222, 3675, 3815, 3059, 2118]. ML [3443, 1730]. ML-II [1730]. MLEs [1981, 3744]. Mobile [2721, 1425]. mobility [3982, 4037, 2445, 3421]. MoCap [2528]. Mode [1807, 60, 3647, 3590, 492, 4659, 2527, 4182]. Mode-Mixing [4182]. Model [2042, 2984, 1830, 3847, 1682, 147, 1662, 4698, 1709, 4460, 2249, 1378, 3638, 186, 4592, 182, 3500, 1410, 198, 3643, 1646, 1615, 2396, 1395, 1672, 1213, 1661, 300, 1616, 1617, 3342, 299, 1707, 1801, 187, 4604, 1647, 4480, 1422, 1394, 1706, 2037, 3640, 4125, 1623, 1425, 1452, 4124, 4447, 2135, 3331, 4640, 4144, 3157, 2304, 4367, 2805, 3085, 3988, 4156, 470, 3805, 2733, 4114, 1510, 1793, 3679, 2709, 4163, 1809, 3816, 522, 900, 2034, 2337, 4473, 1128, 3917, 2779, 3512, 4296, 4245, 2025, 3510, 4238, 2107, 1954, 4650, 4376, 548, 2969, 2239, 1894, 2664, 4131, 3270, 2262, 2658]. model [2880, 2180, 3786, 4415, 2176, 3595, 2995, 4318, 2237, 4233, 501, 1792, 4433, 1999, 3254, 4529, 3784, 4119, 1843, 2637, 3423, 1169, 1811, 2098, 3194, 3168, 3899, 3999, 778, 2361, 3234, 4331, 3706, 2797, 542, 3794, 555, 948, 1156, 2603, 1962, 3048, 977, 1595, 1782, 1970, 2240, 2807, 2900, 3456, 3532, 3653, 3700, 3859, 4175, 4438, 2101, 3329, 2993, 2149, 2342, 4023, 1141, 4007, 4418, 4282, 1926, 3229, 3735, 4143, 2675, 989, 4347, 4678, 3277, 3582, 3606, 2964, 4339, 2357, 4424, 3755, 862, 679, 3011, 4543, 4137, 3160, 1602, 1877, 3915, 4214, 4157, 4223, 2007, 3519, 2887, 3987, 4674, 2620]. model [2976, 2162, 2294, 2736, 3511, 4515, 4598, 4005, 3488, 4484, 1968, 3690, 3676, 3170, 3543, 2017, 1286, 1728, 2079, 3518, 3417, 1260, 729, 3038, 585, 3042, 606, 1158, 4620, 4625, 2153, 2656, 3678, 2130, 4369, 1934, 4160, 3581, 3989, 3710, 642, 3891, 4256, 790, 788, 901, 3065, 3738, 2263, 2459, 2299, 2691, 3566, 2045, 3813, 1785, 4241, 2759, 967, 3591, 4030, 3986, 4503, 880, 1239, 1183, 3016, 2529, 733, 2775, 3103, 2640, 3359, 2566, 3567, 947, 797, 3407, 1285, 3366, 2516, 4593, 2004, 2629, 1992, 3413, 3711, 3750, 4215, 640, 1878, 1518, 1768, 4076, 2945, 3771, 486, 959, 2740]. model [250, 2054, 1121, 489, 1938, 2400, 1318, 4542, 2596, 889, 1063, 1737, 2621, 2910, 3263, 3695, 4069, 4519, 344, 3661, 2604, 2917, 3644, 1537, 4428, 4373, 3896, 2375, 1509, 1064, 1519, 3313, 1054, 1129, 1864, 2151, 2391, 3764, 2434, 2020, 2205, 3186, 3392, 3561, 3527, 3770, 4010, 3952, 3933, 4520, 1860, 3601, 4357, 2421, 2558, 2916, 3090, 3443, 3360, 3840, 4583, 3125, 2877, 4397, 4315, 2634, 4627, 4588, 3376, 2871, 2029, 3809, 3654, 4313, 2242, 2962, 3577, 441, 2465, 3139, 2190, 4051, 1885, 3715, 1275, 1907, 3743, 2483, 304, 3231, 2843, 3175, 2336, 3505, 2930, 2811, 3704, 3110, 4607, 4601, 4471, 3588, 4299, 3928]. model [1971, 4524, 1902, 2536, 4164, 4361, 3910, 3469, 2626, 3825, 4606, 4096, 2631, 4510, 3101, 4231, 2050, 2397, 2708, 1163, 835, 4419, 3123, 3587, 1255, 1304, 4244, 1024, 3542, 4015, 4271, 2412, 2892, 3548, 3776, 2608, 2826, 3382, 3762, 4064, 2437, 3563, 1343, 4237, 4254, 4556, 2972, 3596, 4644, 2202, 3363, 1562, 3124, 737, 4292, 738, 613, 3802, 4414, 4631, 2052, 4130, 3820, 4378, 1805, 3604, 3853, 685, 2919, 2963, 4535, 3293, 2220, 930, 4559, 1973, 3870, 1336, 1739, 1744, 4186, 3375, 3972, 2510, 2177, 2947, 1982, 3136, 2310, 2704, 3421, 3597, 3499, 1152, 1215, 4217, 3615, 4262]. model [3046, 4382, 1065, 4062, 2663, 2978, 4394, 2265, 1196, 3649, 4135, 3219, 3682, 4410, 389, 3758, 1234, 2539, 2402, 4336, 1358, 1834, 4526, 4332, 2703, 2378, 2979, 727, 3992, 2898, 3243, 3953, 3060, 2786, 4568, 3176, 591, 550, 2012, 3275, 3019, 4258, 339, 2109, 4060, 4050, 4219, 4080, 2223, 3524, 4622, 2719, 3075, 1898, 4629, 2418, 903, 1346, 1521, 3733, 4106, 4429, 549, 4688, 3195, 1494, 2271, 467, 1051, 3501, 4319, 4261, 2856, 2915, 2936, 3713, 4349, 887, 3475, 4689, 2122, 2475, 3914, 4539, 1349, 3979, 4181, 3220, 4111, 4360, 2168, 3605, 3885, 1593, 3486, 3730, 1738, 2137, 2832, 2946, 4034, 4548]. model [2698, 2763, 3651, 2625, 1540, 3693, 4110, 4399, 4465, 3832, 3410, 4368, 4527, 1932, 1891, 3341]. Model-assisted [4460]. model-averaging [4583]. Model-based [3847, 2396, 1213, 4480, 4125, 3168, 1992, 3644, 2465, 3820, 4217, 2271, 1738, 3693]. Model-free [4124]. modeled [3224, 3137, 2197]. Modeling [1681, 2395, 2564, 1837, 4727, 1169, 3211, 4546, 3639, 4633, 4369, 1538, 4335, 4715, 2308, 2796, 3958, 4735, 3043, 3792, 2064, 3635, 4377, 3754, 3489, 3495, 3683, 4722, 2377, 1874, 4719, 2182, 4229, 2047, 4452, 3222, 2022, 4250, 2278, 4527, 2292, 3480, 3485, 2896, 4238, 2820, 2420, 2880, 3760, 2842, 4579, 3320, 4643, 3836, 3345, 2568, 4287, 2800, 4582, 2224, 4611, 2358, 4038, 2000, 3949, 2476, 2070, 3049, 3334, 4553, 1806, 2876, 2725, 4411, 4251, 2705, 3045, 2682, 3112, 2139, 3306, 2094, 3557, 4241, 2019, 2711, 3562, 1754, 4466, 4572, 4203, 3204, 1171, 4103, 3716, 3144, 743, 2067, 4654]. modeling [3481, 3354, 3484, 3860, 3430, 2738, 2872, 3922, 2982, 4355, 2937, 1808, 2617, 3422, 2610, 4608, 3559, 3111, 2428, 3684, 2546, 2911, 4216, 2246, 4528, 2028, 3973, 2388, 3649, 2624, 3116, 2517, 4502, 4639, 2535, 4053, 4123, 2855, 3799, 2642, 3088, 4308, 2415, 2480, 4398]. modelings [3628]. Modelling [996, 836, 3232, 2329, 4670, 2637, 2321, 734, 2966, 3841, 1417, 1651, 2589, 1429, 789, 831, 1512, 4732, 500, 2758, 3487, 1061, 4300, 817, 2442, 1574, 4716, 1222, 666, 1547, 1740, 1227, 3791, 2417, 1521, 663, 1328, 1349, 1564, 1485, 4041, 3634, 2970, 2684, 2429, 4536, 2553, 1526, 2002, 2647, 2920, 3199, 2701, 4375, 562, 2244, 1126, 1225, 4424, 4321, 1103, 3818, 4609, 923, 3883, 2991, 409, 2227, 1840, 601, 3255, 3831, 1846, 2953, 4580, 2317, 1224, 342, 673, 2803, 3746, 2612, 3095, 3926, 2005, 2353, 4595, 3727, 4057, 3272, 4183, 281, 4025, 4395, 1206, 2401, 1814]. modelling [3911, 2584, 1990, 994, 3200, 2600, 4116, 1550, 4450, 1298, 2232, 777, 1048, 3238, 1295, 1207, 1552, 785, 3547, 4557]. Models [1419, 1669, 1364, 82, 1659, 1644, 150, 4694, 1665, 1442, 1605, 2319, 1389, 1449, 1418, 1689, 595, 1439, 4717, 4736, 1432, 4000, 1448, 4022, 1652, 1629, 3091, 4702, 775, 997, 2472, 1520, 3173, 1986, 4279, 2258, 2984, 895, 2810, 3028, 3072, 516, 4072, 4441, 4664, 522, 2903, 3998, 4665, 1125, 3917, 3184, 806, 1356, 4013, 3215, 2235, 1354, 2305, 2479, 4549, 3837, 3751, 2988, 335, 2102, 2676, 4273, 3534, 1899, 3658, 1837, 1353, 1948, 2913, 3556, 3942, 2873, 4268, 2076, 3029, 2920, 3766, 2971, 1775, 3027, 1758, 735, 1359, 3828, 4670, 3491, 3836, 4587, 1854, 4199, 1292, 2869]. models [734, 1272, 3234, 2800, 949, 220, 843, 3294, 3224, 3700, 4188, 4077, 2216, 2912, 398, 3788, 391, 442, 2582, 482, 2149, 1910, 657, 4316, 1294, 1322, 4657, 3645, 3452, 2713, 2301, 3872, 1905, 2926, 2655, 3535, 2409, 3212, 1225, 2073, 1247, 3314, 4340, 3941, 1228, 2367, 416, 3525, 1214, 1832, 914, 2281, 3127, 4325, 2464, 2887, 3318, 3913, 3875, 4207, 2682, 3523, 3470, 3962, 2595, 2131, 2776, 2018, 4478, 1219, 2561, 3009, 4048, 4534, 1572, 3150, 2549, 2285, 2279, 3863, 1170, 1143, 1799, 4498, 2453, 2690, 3172, 4179, 1087, 3337, 4066, 4228, 4194, 3843, 1760, 3390, 3594, 2729, 4592, 1778]. models [3507, 2148, 2771, 3734, 1546, 1950, 2187, 3283, 4085, 4363, 2268, 1909, 2883, 3514, 2404, 2944, 3247, 1875, 3325, 4421, 4393, 1719, 3135, 3406, 3643, 2163, 1226, 3398, 1835, 3782, 2008, 2679, 723, 2221, 4511, 1296, 2737, 3778, 4345, 3465, 1576, 1522, 2973, 3842, 4269, 3674, 951, 2581, 4079, 4683, 2950, 2365, 2317, 2767, 1224, 1021, 2214, 4408, 3197, 2159, 3884, 2433, 4232, 3144, 4489, 2116, 4679, 2111, 2211, 3130, 3285, 3461, 1800, 2398, 2989, 3012, 3438, 4274, 4490, 2493, 2286, 1534, 3632, 3657, 2885, 2934, 4440, 4566, 4651, 3225, 1895, 2511, 2850, 4552, 3290, 3926, 3333, 415, 570, 1514, 1773, 2455, 952]. models [1915, 3451, 3425, 2534, 3646, 2943, 3807, 3428, 658, 1549, 419, 4173, 3522, 3626, 3966, 2410, 1753, 2774, 3901, 1283, 4067, 3967, 751, 1987, 2392, 2914, 2174, 3209, 3276, 4011, 690, 1304, 3681, 3849, 4391, 3111, 2055, 2390, 4323, 2831, 3478, 4486, 3545, 2496, 618, 1983, 3790, 3969, 3684, 3412, 4666, 1014, 1733, 731, 1175, 4512, 1770, 2061, 4073, 423, 4555, 2219, 3820, 3386, 345, 527, 1036, 2923, 4008, 3248, 3538, 1133, 2668, 4056, 3251, 3697, 2081, 1188, 2908, 1223, 1798, 2403, 4341, 1371, 2473, 2218, 3597, 4358, 1004, 4255, 3731, 592, 1998, 4574, 1866, 2955, 2275]. models [4623, 4616, 2143, 2467, 3218, 4409, 2983, 3352, 907, 2451, 2204, 4006, 4136, 1091, 587, 4402, 4202, 3924, 3093, 3622, 2013, 2583, 2523, 1358, 2208, 2551, 3370, 4658, 4071, 4235, 2146, 3876, 1373, 2056, 3996, 481, 2119, 911, 4115, 1114, 399, 3335, 3789, 3858, 2371, 3296, 2164, 4562, 1274, 752, 3963, 1058, 1839, 3466, 3165, 3856, 3800, 1015, 1016, 1164, 1337, 1959, 2854, 2927, 2781, 3055, 3208, 3459, 3549, 3732, 3897, 4436, 3808, 2692, 1831, 2047, 2513, 3673, 2449, 2525, 1316, 3560, 2046, 4586, 2932, 3498, 2657, 4653, 4027, 2173, 2118, 2415, 2681, 2946, 3202, 3198, 4253, 4453, 4548, 4500, 704, 2253, 2571, 3132]. models [2975, 4407, 4447, 2625, 3378, 1570, 1077, 1257, 3625, 2170, 3729, 4190, 4193, 4624, 1879]. moderate [4531]. moderating [1476]. moderation [4406]. Modern [3025, 3674, 2195, 1727, 4690, 4718, 4461, 3528, 4563]. Modern-day [1727]. modes [2045, 3289]. Modification [1036, 3598, 1591, 840, 2127, 2623]. modifications [1251]. Modified [4144, 1249, 1406, 1059, 1408, 1684, 3956, 4076, 567, 2557, 3341, 3612, 4315, 1416, 2169, 1081, 2968, 2741, 1837, 4318, 1265, 1970, 3391, 2678, 2085, 3891, 1763, 1981, 2205, 2184, 1541, 3881, 881, 4298, 4417, 2283, 3938, 3758, 466, 646, 3275, 3744, 1375, 3052]. Modifying [1588]. Modular [2743]. Modular-transform [2743]. module [4218]. modules [1953, 3830, 3833]. molecular [2655, 3449]. molecules [2489]. mollissima} [1204]. Moment [1631, 568, 1456, 2238, 4151, 1284, 3855, 453, 2570]. Moment-Based [1631, 2238, 3855]. momentary [4232, 2389]. Moments [1693, 1313, 954, 2200, 2300, 4058, 3536, 3921, 3920, 1026, 4093, 4525, 1178]. momentum [4197]. monetary [3814, 2791]. money [283]. monitor [865, 4543]. monitored [3973]. Monitoring [4098, 4046, 4469, 1271, 4433, 1359, 4423, 1450, 3137, 1655, 3368, 1569, 1402, 2063, 2027, 4146, 1856, 2380, 4523, 4647, 4234, 2023, 3480, 4039, 4472, 2521, 3355, 1192, 4670, 2110, 4413, 3960, 967, 1261, 4565, 4280, 3742, 4300, 4636, 2080, 4104, 2024, 2671, 2769, 4353, 1150, 4463, 3990, 1146, 4054, 4494, 1112, 1327, 3550, 2848]. monocropping [1740]. Monotone [1764, 4640, 3365]. monotonic [3369, 2166, 1067]. monotonicity [3073]. monotony [3445]. monoxide [3355, 4659]. Monte [827, 4512, 2805, 3081, 1643, 211, 3491, 1409, 601, 1763, 1835, 2740, 2221, 2910, 2105, 3657, 2897, 1255, 4272, 1801, 3385, 3228, 866, 3040, 1274, 3808, 1484]. Monte-Carlo [1801]. Monthly [1606, 1025, 4026]. Monty [2343]. mood [2638]. Mop [151]. Moran [1738]. morbidity [1478]. Morgenstern [1155]. Morlaix [1517]. morphology [664]. Mortality [1417, 3332, 2998, 4545, 2690, 2945, 4511, 4580, 1907, 731, 2401, 3444, 4402, 1924, 3054, 1521, 3801, 4689, 4551]. most [1491, 3531]. most-cited [1491]. Motacilla [1208]. moth [669]. motility [3125]. Motion [101, 1034, 665]. motivated [2341, 2897]. mourning [1233]. move [2400]. movement [3504, 744, 1226, 742]. movements [3929, 764]. mover [843]. mover-stayer [843]. Moving [1368, 4364, 4523, 2797, 4423, 1847, 815, 3144, 2727, 4659, 1036, 3394, 4528, 4006, 2259, 4405, 1327]. Mozambique [3632]. Mr. [105]. MRI [4462]. Mrs. [52]. MRSI [1374]. MRSI-observed [1374]. MSE [1303]. MTAR [1518]. mTOR [4632]. much [4049]. Müller [2818]. MULT [746]. Multi [1426, 136, 1709, 4462, 3041, 1623, 3173, 4245, 4118, 2568, 3010, 4544, 3049, 4339, 1228, 4325, 4284, 4005, 246, 2640, 2348, 3325, 4215, 1226, 951, 743, 3443, 2379, 4017, 304, 3752, 751, 2041, 1547, 2662, 3589, 4080, 920, 4053, 3811, 2918]. multi-armed [3589]. multi-block [2379]. multi-charts [4284]. multi-class [2568]. multi-component [4215]. multi-country [4245]. multi-disciplinary [3173]. multi-disease [2918]. multi-entry [304]. multi-fidelity [3811]. multi-frequency [4118]. multi-group [951]. multi-index [2640]. multi-IV [4005]. multi-level [4339, 3443, 3752, 4080]. Multi-Nomial [136]. Multi-Objective [1426]. multi-omic [4544]. multi-parametric [4462]. multi-party [246]. multi-rater [2662]. Multi-relational [3041]. Multi-resolution [4462]. multi-response [920]. multi-sample [4017]. multi-scale [4053]. multi-stage [3010, 4325, 1547]. Multi-state [1709, 1623, 3049, 1228, 3325, 1226, 743, 751]. multi-way [2348, 2041]. multibiometric [2908]. multiblock [3609]. multicenter [3217]. MulticlusterKDE [4212]. Multicolinearity [3757]. Multicollinearity [2114, 3556, 3310, 2468, 3283, 3472, 4310, 3881, 3031, 1498]. multicomponent [4497, 4213]. Multicriteria [4390]. Multidimensional [2903, 1142, 3504, 2994, 4594, 3978, 1251, 109, 2927]. multifractal [3309]. multigraph [4574]. Multilevel [2305, 2803, 1808, 4703, 3173, 4441, 2235, 1353, 2237, 1531, 2406, 2036, 2516, 3657, 3290, 3425, 3110, 3801, 2173, 3885]. Multimodal [4201, 4302, 2288, 4488]. multimodality [4116]. Multinomial [1413, 1693, 2670, 2913, 2802, 2595, 606, 1082, 1312, 3986, 3166, 334, 4419, 1461, 1744, 467, 3222]. Multinomials [1387]. Multiparametric [1389]. multiplayer [3421]. Multiple [253, 4681, 3336, 188, 1710, 3479, 3077, 2326, 1437, 3014, 2488, 601, 2004, 1391, 3365, 2419, 1713, 2495, 3461, 2683, 1393, 252, 141, 837, 1903, 200, 3963, 1622, 1375, 1648, 1468, 2573, 4426, 3121, 4564, 3089, 4458, 2186, 2487, 2809, 3108, 4434, 4424, 3127, 4598, 3488, 530, 4132, 600, 2287, 506, 637, 1867, 2045, 1914, 3082, 3507, 797, 2057, 1524, 2734, 395, 3624, 889, 3263, 344, 4203, 1374, 4517, 3513, 2431, 2989, 2087, 3258, 4209, 2695, 1569, 3588, 2837, 4573, 4210, 3951, 2053, 2852, 3189, 3123, 2072, 4064, 1473, 1575, 4560, 1194, 2030, 2584, 4186, 2310, 2704, 1734, 1005]. multiple [2841, 2152, 3682, 4634, 412, 503, 3540, 1723, 1750, 4541, 4229, 688, 1100, 1716, 1051, 1995, 4449, 2475, 4366, 1554, 4070, 2331, 4250, 898, 899, 2801]. multiple-attribute [2837]. multiple-choice [1750]. Multiple-index [3461]. multiple-interventions [1005]. multiplicative [542, 4180]. multipliers [4010]. multiply [463, 3117, 4400, 4373]. multipoint [2448]. Multipopulation [4580]. Multiresolution [2806, 3482]. Multiscale [3706, 3502]. multisite [2601]. multisite-multivariate [2601]. multispectral [3702]. multistage [3379, 4568]. ’Multistate [1225, 1224]. multitesting [3954]. multitudinal [3999]. Multivariate [3173, 1710, 1641, 1650, 494, 2464, 2674, 3473, 3337, 2254, 3960, 1258, 4749, 1401, 4363, 1495, 3247, 1655, 814, 1397, 2850, 3522, 1747, 1402, 2260, 1690, 2206, 2668, 976, 4295, 2345, 2793, 1416, 3858, 2639, 1688, 1198, 4367, 1113, 3028, 4665, 3357, 374, 2896, 2117, 929, 2423, 4268, 1811, 3194, 4199, 3153, 2313, 3894, 2912, 1019, 2110, 1142, 2355, 2942, 2284, 2097, 4347, 671, 2776, 2086, 1260, 3034, 4369, 469, 4228, 2488, 406, 1078, 3152, 3458, 2883, 1719, 4578, 4511, 4122, 1296, 810, 3105, 2522, 2408, 3374, 2396, 3313, 4580, 3764, 2469, 3742, 3467, 1857, 2211, 2773, 2386, 2379, 3376, 2695]. multivariate [864, 3654, 3333, 2520, 2633, 2410, 1827, 4164, 1119, 4067, 4510, 2645, 2788, 3123, 4244, 4391, 4127, 3635, 1506, 3618, 1993, 4383, 3754, 3683, 2911, 3120, 3240, 1133, 1739, 2601, 2769, 4212, 4626, 1547, 1960, 4616, 2928, 4634, 2267, 4294, 605, 1055, 630, 4071, 2418, 2781, 3565, 4452, 1327, 4070, 4617, 2354, 4342, 4453, 4557, 2778, 3463, 3656]. Multivariate-based [1258]. multiway [515]. Municipal [4381]. muscle [3446, 1582]. muscles [3568]. museums [3645]. music [2830]. mutants [1336]. mutation [4422]. mutidimensional [3887]. Mutual [3988, 1929, 2070, 2901]. MX [4304]. My [46, 66, 89, 76, 100]. myocardial [3537]. myoelectric [3446, 3070]. myoelectric-controlled [3446]. Myopia [123].

N [2815, 3520]. N-shape [3520]. n1 [576]. n2 [576]. Nadarajah [4354]. Nadu [3524]. Nairobi [254]. Naive [62, 1869, 3699, 3377]. nanoparticle [3390]. nanoparticles [1818]. natal [740]. National [925, 129, 164, 2245, 3543, 2844, 1885, 2532, 2907, 1098]. Nations [4360]. native [3164]. natural [3027, 807, 303, 1544, 4327, 731]. Naturalistic [3294]. nature [2613, 2554]. NB [3721]. NBA [1072]. NBER [2575]. NCAA [1272]. NDNQI [2532]. near [2964, 4642, 4419, 4604, 4409, 887]. near-accident [4409]. near-infrared [2964, 4604]. near-sited [4642]. near-to-separation [4419]. nearest [2097, 2266, 3206, 3190]. Nearly [1560, 961]. necessary [247]. necessitating [3562]. neckbands [1236]. need [3634, 3613, 65, 80]. needed [1916]. needs [317]. Negative [1605, 3411, 3220, 397, 4286, 2957, 2007, 3042, 4189, 4520, 4456, 1808, 2412, 2904, 3972, 3289, 2075, 2259, 4502, 3885, 3625]. negatively [1937, 2416, 4607]. neglected [1309]. Negligible [4309]. neighbor [3206, 3190]. neighbors [2097, 2266]. Neighbouring [1663]. Neil [2815]. Nelson [3019]. nematode [4369]. neoplasia [2740]. Neosporosis} [2162]. neotropical [1240]. nest [743]. Nested [4530, 584, 3700, 4478, 1768, 725, 853, 940, 1071, 4025, 2356, 984, 4053]. net [3399, 4119, 4638, 3208, 4147]. Netherlands [283]. netting [758]. Network [4102, 957, 1520, 2387, 3595, 4670, 3526, 4582, 652, 1141, 4411, 4168, 2250, 2595, 970, 3927, 4490, 2960, 4595, 4096, 4585, 4445, 3240, 1045, 3165, 4448, 4680, 4308]. Network-based [4102, 4096, 4308]. Networks [1659, 4731, 3213, 4279, 3970, 3512, 3069, 4582, 1019, 3645, 3450, 3174, 3570, 4543, 4451, 3337, 2133, 3688, 2788, 4585, 3724, 1884, 653, 3911, 2499, 1919, 2636, 1048, 4180, 4218, 1883, 2815]. Neural [1659, 653, 1712, 957, 1520, 652, 1019, 1141, 4168, 2250, 2595, 2133, 3688, 4585, 3911, 1045, 4180]. neurochemicals [1374]. neurocognitive [4540]. neurodegenerative [3286]. neuron [4180]. neuroscience [4643, 2290, 2612]. neutral [684]. News [426, 13, 2851]. newsletter [55]. newsprint [1061]. next [1201, 1202, 3353]. Neyman [1650, 3641, 4350]. NF [2377]. NFL [4601]. Ngau [543]. NGS [3672]. NHPP [1629]. nicotine [1860]. Nigeria [4633]. Nigerian [465, 200]. nitrogen [527]. NMR [3268]. No [29, 53, 1007, 599, 3936, 561]. no-shows [3936]. Noise [1393, 1412, 4454, 2414, 601, 815, 2894, 605, 3799, 2806]. Noise-insensitive [4454]. noises [4077]. noisy [417, 3681]. Nomial [136]. Nominal [1394, 1923, 2153, 2656, 3593, 2214, 2256, 1549, 1120, 4202, 4071, 2378, 909]. Nomogram [3918, 3699]. Nomograms [98]. Non [1444, 1399, 2686, 1413, 1638, 1450, 3080, 908, 950, 2094, 870, 1341, 1708, 1418, 4377, 2219, 2327, 558, 2006, 170, 1629, 4359, 470, 3242, 1236, 1478, 1271, 19, 445, 553, 1975, 1901, 1758, 3320, 3828, 3767, 3369, 3217, 802, 1102, 1458, 3158, 1829, 4049, 2487, 4302, 3314, 1783, 4134, 2922, 3104, 600, 4165, 788, 4189, 3137, 4630, 644, 285, 2775, 947, 2268, 1323, 2440, 1544, 4230, 849, 539, 684, 842, 1938, 781, 4479, 892, 2365, 919, 2317, 2767, 379, 457, 3284, 2105, 2134, 4327, 1104, 670, 1355, 2511, 2242, 2661, 300, 2937, 992, 2727, 701, 4033, 1255, 3411, 4237, 4464]. non [1745, 318, 2061, 1814, 685, 3163, 4298, 3240, 3726, 4056, 493, 993, 2510, 4409, 3192, 3436, 1091, 322, 389, 587, 4353, 2539, 2402, 2259, 4332, 411, 3721, 4071, 2201, 866, 863, 1599, 3, 888, 496, 586, 2785, 339, 2062, 3055, 1044, 1831, 2047, 965, 2728, 4513, 2606, 4180, 3584, 1564, 2832, 3974, 704, 3216, 1540]. non-active [3163]. non-Bayesian [4359, 4134, 4479]. non-breeding [1236]. Non-central [1450]. non-completion [849]. non-compliance [2402]. non-concave [1938]. non-conforming [919]. non-continuous [3217]. non-epidemic [1478]. non-Euclidean [2922, 3216]. non-Gaussian [2317, 2061, 3240, 3974]. Non-homogeneous [1708, 3158, 3104, 2775, 2268, 781, 379, 457, 2937, 3726, 2259, 4332, 1564]. non-ignorability [2511]. Non-ignorable [4377, 3828, 947, 4071, 2832]. non-iid [1829]. non-inferiority [1783, 4464, 3192, 2062]. non-informative [4298]. non-isotropic [3436]. non-iterative [2365, 411, 1540]. Non-linear [558, 470, 600, 4165, 788, 3137, 644, 670, 992, 2727, 493, 389, 587, 888, 4180, 704]. non-linearity [1544]. non-mathematicians [19]. non-metric [3584]. Non-mixture [2219, 1831, 2047]. non-monotonic [3369]. non-negative [4189, 3411]. non-neutral [684]. non-Normal [2728, 553, 1102, 1458, 4033, 1814, 685, 322, 2539, 496, 586]. Non-normality [1418, 2006, 539, 2134, 1104, 4353, 1044]. non-numerate [3]. non-orthogonal [318]. Non-Parametric [1399, 2327, 170, 2686, 1413, 1638, 908, 950, 1341, 1629, 3242, 1271, 1901, 1758, 3320, 3767, 802, 4049, 3314, 842, 892, 2767, 1355, 701, 4237, 4056, 993, 1091, 2201, 866, 863, 1599, 3055, 965]. Non-perennial [1444]. non-persistent [300]. non-proportional [4230]. non-random [4630, 2511]. non-randomized [1975, 2440]. non-recall [4513]. non-rejective [2606]. non-response [947, 3284]. non-rigid [4302]. non-selected [1745]. non-signalized [4409]. Non-sparse [3080]. non-spatial [4327]. non-specific [285]. non-stationarity [2785].

Non-stationary [2094, 1323, 2242, 1255, 2510, 3721]. non-statisticians [445]. non-symmetric [2487]. Non-traditional [870]. non-uniform [2105]. non-zero [2661, 339]. nonbinary [3730]. noncoding [2112]. noncompliance [2084]. Nonconformance [3961]. nonconforming [2060, 2027, 1813]. nonconformity [4184]. nonhomogeneous [4524]. nonignorable [2679, 4348, 2041]. noninferiority [4084]. Nonlinear [2258, 3813, 3782, 1846, 2462, 3790, 1475, 3503, 2117, 4273, 3667, 4282, 2452, 3962, 2561, 2130, 4498, 1840, 2316, 2317, 2159, 2989, 3326, 3376, 2777, 1753, 3097, 3545, 3911, 2081, 2403, 4341, 1998, 3126, 3370, 4153, 4502, 3075, 3466, 1959, 3560, 4181, 2946, 3378]. Nonlinearities [1614]. Nonlinearity [3144, 3556]. nonnegative [3074]. Nonnested [3030]. nonnormal [3667]. Nonparametric [3427, 3903, 1486, 3987, 3415, 2593, 3071, 1922, 4742, 2973, 2140, 61, 2136, 2428, 4260, 2044, 3454, 2231, 4196, 2426, 1653, 2150, 1839, 3388, 968, 3709, 2790, 3552, 4615, 3873, 4625, 2068, 3793, 972, 2735, 3838, 3186, 4470, 3603, 2196, 1275, 1762, 3618, 3850, 4295, 4482, 2160, 2154, 4141, 2425, 2747, 3560, 2940, 3261, 3041]. Nonparametric-likelihood [2150]. Nonparametrics [2324]. nonregular [3188]. nonresponse [2041]. norm [4457]. Normal [2329, 4331, 1421, 1611, 2728, 1113, 4652, 1964, 4665, 1669, 553, 2363, 4273, 836, 3239, 620, 1843, 1009, 3745, 1102, 1458, 482, 2335, 1542, 3913, 2682, 3523, 3962, 4484, 1968, 3150, 565, 1087, 4194, 3813, 1779, 4030, 2990, 3407, 3152, 973, 1488, 490, 4122, 810, 2621, 2576, 2365, 1857, 2557, 2773, 3012, 2493, 3947, 2885, 4031, 1858, 3763, 2446, 896, 3646, 3505, 1971, 1543, 1505, 4361, 4033, 3783, 4002, 1528, 3790, 554, 3148, 1814, 685, 2963, 3385, 2704, 1963, 322, 2539, 576, 2898, 4091, 484, 460, 496, 586, 4115, 776, 3583, 4027, 2118, 899, 2311, 2999, 4444]. normal [2104]. normal-inverse [3407]. normal/independent [3075, 3466]. normalisation [3814]. Normality [2200, 1410, 1635, 1628, 929, 1752, 501, 2021, 1932, 3662, 469, 539, 2134, 1104, 1418, 1315, 1567, 840, 2058, 4353, 2006, 4003, 1044, 986, 2778]. Normalization [2014, 3666, 691, 3995]. Normalized [1044, 621]. normalizing [1112]. normally [91, 3402, 3589]. normative [249]. Norms [1659]. North} [3767, 1209, 3524, 758, 769, 1604, 4520, 3412, 4026]. north-eastern [1209]. northeast [3088]. northern [1362, 2342]. nose [261, 3687]. not-at-home [593]. Not-so-linear [277, 314, 337]. Note [103, 213, 1691, 174, 170, 1425, 1452, 1128, 19, 555, 1970, 1103, 4207, 812, 507, 1161, 3746, 2634, 1351, 1315, 1562, 1014, 1567, 1515, 587, 2477, 630, 286, 2650]. Notes [13, 2084, 3056, 2764, 2906]. nothing [34]. Notice [4288, 4289]. notion [2097]. novel [4139, 2797, 4434, 3591, 2622, 4140, 4511, 4456, 2029, 3681, 4585, 3227, 3820, 3721, 2638, 4153, 4689, 3629, 4372, 4548]. novice [3561]. NP [1683]. NPCA [2966]. nucleotide [2569, 2650]. Nuisance [1420]. null [1291, 4665, 4388, 1811, 2568, 1185, 1509, 3952, 2693]. nullity [1128]. Number [1673, 1917, 357, 693, 1266, 1595, 2240, 2849, 4546, 1571, 990, 264, 910, 1369, 4421, 1976, 2434, 1161, 3210, 982, 1079, 2811, 4015, 2783, 4528, 4112, 3054, 963, 3068, 2168, 3442]. Numbers [1606, 1678, 472]. numerate [3]. Numerical [284, 1980, 4323, 4016]. nursery [3618]. Nursing [2532, 1900, 3802]. nutrition [4417, 3659]. nutritional [4657].

O.R. [33]. obesity [3659, 4629, 4070]. Obituary [479]. object [912, 659]. Objective [1426, 3564, 3476, 2995, 4608, 4061, 3396, 3291]. objects [602]. Observable [1143]. Observation [103, 607, 2875, 3339, 3178, 3642, 4261]. Observational [152, 3556, 2766, 2364, 4477, 3492, 4605, 1204, 3619]. Observations [182, 1254, 1634, 1386, 1656, 1706, 4665, 335, 3667, 398, 482, 1926, 2678, 1115, 330, 2607, 406, 248, 2252, 1524, 694, 3368, 1063, 3090, 3874, 1534, 1761, 2024, 2053, 2852, 3189, 4015, 355, 1473, 554, 2666, 331, 4080, 4439]. Observed [4575, 792, 1542, 4329, 1374, 815, 3376, 4313, 2389, 3449, 2693, 2369]. observer [740]. observing [1781, 2016]. obstetric [4331]. obtained [763, 4199, 2635, 1539]. OC [621]. Occam [1202, 1201, 785]. occupancy [4012]. Occupational [90, 1933]. occurrence [2969, 3760, 2294, 3650, 3852]. occurrences [2334]. occurring [4468, 3067]. October [3720]. odd [4356, 4134, 2892, 4481, 4350, 4671]. odds [2715, 3117, 2348, 2084, 3809, 4366, 4110, 2527, 3410]. odor [3687]. odour [1503]. OECD [2968, 3526, 1370, 2491, 4428]. off [2879, 3221]. offline [2835]. offs [952, 752]. offset [4207]. offspring [3981]. often [2001, 1911]. oil [3203, 3437, 2758, 4042, 2742, 4260, 3006]. oils [3268]. Old [184, 2304, 2346, 2694]. Old-Faithful [2694]. older [2493]. olfactory [1503]. olive [3268]. Olkin [4305, 1530, 3863, 3477, 2936]. OLS [3284]. Olympic [2851, 4061]. Olympics [2851]. omega [4576]. omic [4544]. omitted [1941, 3205]. Omnibus [1718, 1635, 986, 1315]. On- [2835]. on-field [3620]. on-line [2296, 3421]. Oncology [4693, 4730, 4246, 2746]. on‘connector [26]. oncotype [4406]. One [2665, 1873, 1392, 1423, 1450, 1387, 3266, 842, 1403, 1656, 4306, 1128, 4152, 3935, 1191, 3780, 1193, 541, 4423, 2718, 2452, 4334, 944, 990, 1085, 931, 1087, 918, 1285, 4686, 1548, 2962, 1853, 3743, 1293, 2309, 1982, 3499, 4626, 2726, 1759, 1903, 1579, 3171, 4450, 4146, 550, 963, 2588, 1044, 1375, 4405, 4107, 3463, 3929]. one-class [4450]. one-compartment [1285, 2962]. one-day [2726]. one-dimensional [4626, 3171]. one-level [963]. one-parameter [4306, 4334]. One-sample [1656, 918, 1982, 1044]. One-sided [1423, 3266, 842, 1403, 3935, 2718, 4686, 1548, 2309, 1759, 4146, 4405, 3463]. One-step [2665, 1375]. One-Way [1387, 1128, 541, 1087, 550]. ones [3707]. Online [4413, 4228, 4283, 2780, 3068, 4570]. only [3613, 1542]. onset [2604]. onto [2589]. Open [3163, 748, 735, 3176, 3133, 157]. opening [4093]. operating [828, 2853, 2538, 2126, 3192, 557]. operation [524, 4006]. Operational [268, 1986]. operations [254, 2005]. operative [966]. operatives [2966]. opinion [246, 24, 459, 36]. opinions [513]. Opportunities [1147, 2140]. opportunity [1755]. Opposite [1389]. optic [1048]. optics [422]. Optimal [806, 1356, 3671, 4155, 2879, 948, 2083, 64, 1082, 4256, 4149, 2486, 4081, 1760, 1626, 1663, 102, 2576, 2724, 1331, 2116, 1534, 3830, 4263, 1230, 881, 3851, 3240, 4616, 4112, 2259, 3833, 1600, 2461, 1420, 1278, 3903, 4059, 2176, 1156, 3322, 1581, 3773, 2813, 302, 641, 3104, 4514, 4630, 1088, 2552, 4646, 1161, 1761, 943, 1991, 4672, 1770, 3249, 1982, 1122, 4099, 1515, 4634, 3003, 726, 1329, 1112, 4117, 2696, 2997, 3377]. Optimising [261]. Optimization [1426, 1725, 1944, 2222, 1127, 1818, 1393, 1554, 3322, 4159, 1264, 3680, 4162, 1730, 4158, 4181]. optimize [3433, 1148]. Optimized [774, 4284, 4451]. Optimizing [2375]. Optimum [3784, 1423, 1896, 3034, 1424, 2860, 1904, 1057, 573, 717, 578]. option [2020, 3974]. orbits [4442]. Ordaz [3161]. Order [223, 1665, 4064, 3267, 3213, 2709, 2994, 3215, 2741, 4126, 1155, 668, 1560, 3728, 4438, 2582, 3077, 1887, 2439, 1888, 3094, 1084, 1479, 408, 807, 1149, 2561, 1893, 4276, 590, 972, 1285, 4058, 2163, 1609, 3118, 3323, 4610, 3373, 3840, 2108, 1991, 4437, 2914, 2770, 580, 1624, 4020, 405, 2052, 423, 488, 3289, 1759, 677, 2898, 3990, 4320, 2331, 4109]. order-insensitive [2331]. order-of-addition [4109]. Order-restricted [1665]. Ordered [167, 1395, 3995, 988, 3700, 1930, 1493, 4264, 3939, 2348, 4092, 2236, 2405, 1591, 4208, 975, 2923, 4623, 1759, 2979, 3018, 1593]. ordered-heterogeneity [1591]. ordering [872, 3171]. orderings [1508]. orders [2179]. Ordinal [1322, 1697, 3308, 2339, 3575, 4296, 2025, 2239, 1923, 391, 1972, 3117, 2153, 3678, 2833, 2640, 4593, 4508, 3365, 4668, 3185, 3285, 2087, 2256, 3145, 2671, 4221, 3823, 3382, 1908, 1120, 2963, 3293, 2177, 3312, 2662, 3938, 3649, 1943, 4202, 4071, 3116, 3789, 3296, 3861, 3410, 4343]. ordinal-nominal [2256]. ordinal-scale [2087]. ordinary [3429, 1893, 1162, 1498]. ore [937]. organic [3069]. organised [4736, 2950]. organizational [1836]. organs [2918]. orientation [3899, 377]. oriented [238, 1334, 2382]. Origin [1382, 561, 2014, 595, 1267]. origin-shifted [2014]. Origins [1502, 1610]. orthant [4625]. Orthogonal [1560, 1634, 1849, 282, 1822, 1830, 2574, 2874, 948, 3570, 418, 1331, 2087, 2256, 4057, 1263, 318, 603, 834, 926, 4000, 3312, 321]. orthogonal-array-based [1331]. orthogonal-saturated [2874]. orthopaedic [2019]. Oscar [1008]. osteoporosis [4248, 2594]. other [43, 1588, 3577, 3304, 1151, 2213, 2303]. out-of-sample [3019]. outbreak [2591, 2345, 4537]. Outcome [1696, 2103, 2091, 3491, 2129, 2876, 3381, 4329, 2372, 3326, 3037, 4099, 4634, 1262, 3492]. Outcomes [1709, 4377, 895, 4238, 2913, 3551, 1272, 4655, 3224, 1126, 2149, 4282, 2766, 4424, 999, 2787, 3045, 2656, 3112, 2833, 2516, 3458, 1153, 1165, 3624, 1925, 3632, 2084, 3376, 3654, 2483, 3505, 4090, 1046, 3587, 3754, 4676, 4379, 2760, 3202]. outgoing [1545, 1041, 819]. Outlier [3490, 4045, 141, 1620, 2122, 3703, 3028, 3072, 2257, 1305, 937, 3080, 3738, 4511, 1737, 2419, 1590, 848, 3969, 2031, 4158, 3393, 3932, 2646, 699, 2169]. outlier-resistant [3072]. Outliers [1710, 1063, 2051, 3079, 497, 335, 555, 2809, 3303, 530, 606, 3738, 3782, 2419, 3539, 583, 3588, 334, 4002, 3189, 4292, 3675, 4338, 3806, 3031, 3464, 399, 614, 1996, 898, 899, 4253]. Outlying [182, 925, 3608, 4439]. outpatient [261]. Output [130, 1912, 3837, 2099, 527, 1570]. outputs [3103, 1358]. Ovarian [1674]. over- [2589, 3411, 1821]. over-dispersed [4334, 1808, 2784]. over-dispersion [3409]. over/underdispersion [4116]. overall [2231]. overburden [1956]. overdifference [2742]. overdispersed [4296, 4203, 3896, 4493, 3265, 3287, 2109]. overdispersion [2472, 4441, 3708, 2207, 2401, 3386]. overlapping [2160]. overly [2215]. overnight [4026]. Overpayment [3127, 1511]. overview [370, 651, 4383, 731, 1206, 1207]. owls [1243]. ownerships [3252]. Oxford [2393, 2500, 2503, 4725, 2462]. oxygen [1913]. ozone [2472, 2969, 3726].

p [633, 320, 2887, 1189]. p-charts [633, 320]. Pacific [3814]. paediatric [3184, 4329]. page [12, 770, 822, 946]. Paik [2652]. pain [285]. paint [1127]. painting [3180]. pair [3539, 4112, 3550]. Paired [1391, 952, 2135, 2709, 4596, 1170, 4081, 3096, 2550, 3506, 3723, 3489, 4278, 2764, 3610, 3192, 2638, 2918, 4261, 4107, 1257]. pairing [3723]. pairs [4544, 585, 1572, 1130]. Pairwise [4096, 3121, 4584, 4566, 423, 4008, 1051]. Pakistan [3006]. paleoecological [2145]. palm [2742]. palmar [569]. pandemic [4536, 4433, 4274]. panel [2968, 2384, 1595, 1782, 3859, 3872, 3913, 3875, 1892, 4592, 1518, 3246, 3284, 3438, 3430, 2474, 1026, 4260, 1816, 2780, 4196, 2470, 3602]. panel-data [3430]. panels [2033]. Paper [165, 836, 2589, 1314, 1569]. papers [236, 1491]. parabolic [1927]. paradigm [2709]. paradigms [2278]. Paradox [1440]. Paradoxes [2985]. Parafac [3606]. Parallel [3808, 268, 2622, 3627, 3001]. Parameter [1643, 1671, 1566, 3252, 2561, 2017, 1377, 4035, 3493, 2770, 249, 1420, 2805, 3235, 3395, 4469, 4306, 1822, 2574, 4126, 3510, 4238, 584, 932, 1190, 4059, 3138, 4576, 3780, 1359, 3471, 927, 1242, 4068, 1009, 2659, 1951, 2718, 3889, 4334, 3063, 1516, 279, 2130, 3104, 4160, 3477, 3295, 2759, 2300, 3325, 3867, 1763, 3062, 4122, 846, 3879, 3539, 2308, 4636, 3090, 2877, 2446, 2894, 3743, 3829, 3147, 3272, 1255, 1304, 4285, 3959, 4086, 956, 1133, 1739, 1821, 2381, 3291, 3476, 1903, 2840, 4153, 2712, 2927, 1804, 2122, 1308]. parameterizations [1562]. parameterized [4503]. Parameters [155, 182, 4647, 1965, 463, 3997, 2420, 4433, 620, 4155, 2021, 2800, 3322, 1186, 1141, 2849, 1269, 2723, 644, 2494, 1818, 4215, 1981, 894, 764, 2485, 1252, 4610, 3033, 1457, 4051, 2534, 2683, 4471, 2608, 4292, 4216, 2515, 3248, 3934, 4341, 4358, 2992, 1866, 3218, 3219, 2669, 1579, 1834, 586, 2572, 3744, 1996, 3459, 3409, 1554, 1540]. Parametric [1399, 1266, 1613, 2735, 1615, 3819, 4412, 2327, 907, 170, 957, 3242, 4681, 1271, 2429, 1837, 4432, 1901, 1758, 3320, 3767, 3468, 2686, 3234, 1962, 1413, 802, 1638, 4049, 3314, 1832, 908, 4609, 1362, 2492, 950, 2247, 2759, 1277, 3472, 4462, 842, 4297, 3674, 892, 1341, 2767, 980, 1355, 2511, 2334, 2353, 701, 2774, 3967, 1707, 2496, 4237, 2972, 2220, 4056, 3251, 993, 1091, 4138, 2201, 866, 863, 1599, 3953, 3822, 1513, 1600, 1629, 3055, 3897, 965, 2137, 1257]. Parametrizing [3745]. parasuicide [376]. parent [3981]. parent-offspring [3981]. parental [2139, 280]. Pareto [3769, 1637, 1809, 2677, 1671, 3728, 2659, 1377, 2724, 1302, 1564]. Pareto-inverse [1564]. Parkinson [4158, 2513]. parrot [1240]. parse [4308]. parsimony [1325]. Part [1511, 544, 645, 709, 4252, 277, 314, 337, 2620, 2530, 3186, 2483, 4559, 3560, 1502, 1610, 556, 1015, 1016]. Partial [3310, 4198, 960, 2282, 1386, 1686, 488, 1712, 3919, 2857, 3733, 2994, 4656, 979, 4611, 4509, 2682, 2010, 2530, 4361, 3967, 3681, 3131, 3476, 4062, 3192, 1759, 4060, 722, 3968, 3584, 4447, 3832, 1248]. Partial-order [488]. Partially [3470, 4619, 2926, 4354, 3755, 2736, 3523, 4484, 4329, 4081, 2187, 3135, 3406, 2221, 3778, 2644, 4683, 3096, 2433, 3130, 2916, 3376, 4313, 2389, 4278, 2764, 3560, 4181, 3378]. participated [4303]. participation [2361, 3581]. particle [4159, 4162]. Particular [162, 872]. partition [2094, 2796, 3123, 3693]. Partitioning [3228, 3847, 4256, 2641, 2331]. partitions [515, 1976]. Partly [2453]. party [246, 129, 164]. passage [3327]. Passenger [1700, 3420, 4486]. Passer [768]. passerines [769, 764]. past [2212, 4090, 2523]. patch [631]. patches [795]. patent [3857]. Path [3419, 2028]. Pathogens [1398]. pathological [4063]. pathology [2366]. paths [3785]. pathway [4632]. pathways [1986, 3941]. Patient [2942, 3677, 1986, 4655, 3207, 4005, 4102, 3807, 1943]. patient-level [3807]. Patient-specific [2942, 3677]. patients [4544, 2912, 2000, 1913, 2876, 3664, 3050, 4503, 3247, 4187, 1374, 2390, 4631, 1990, 966, 2441]. Pattern [1433, 1434, 1435, 1440, 1181, 2602, 3065, 654, 1762, 2121, 3024, 2693, 2376]. Patterns [3167, 1679, 708, 1072, 4037, 4633, 3365, 3004, 3846, 3827, 1902, 833, 4073, 527, 1798, 376, 2646]. Pauling [279]. Paulo [2702]. pavement [1810]. pay [2813]. payload [4057]. PBIB [1108]. PCA [1184]. PCR [3544]. PDO [3268]. peak [3568, 4201]. peaks [3237, 2996]. peaks-over-threshold [3237, 2996]. peanut [2233]. Pearl [129, 164]. Pearson [1657, 588, 4475, 4613]. Pedagogical [1966]. pedestrian [4409]. pediatric [4453]. pelagic [1184]. penalization [4516, 2594]. Penalized [3504, 4579, 4646, 1559, 4285, 1677, 1157, 3444, 4119, 3048, 936, 2849, 2944, 4045, 1938, 3587, 4086, 2846, 4219, 4147, 4372]. Penalty [1677, 3881, 3640]. Peninsular [3910]. people [4303, 3941, 2789]. per-allele [1876]. per-field [2484]. per-genotype [1876]. perceived [4549, 4136]. percentage [1074]. percentile [918, 801, 3673]. percentiles [1457]. perception [3899, 3306, 3632, 1727]. perceptions [1836, 3069, 3382]. perceptual [3777]. percolation [2382]. perennial [1444]. Performance [3164, 1000, 1454, 1774, 3752, 1618, 1025, 4614, 4719, 1702, 3845, 3157, 3769, 4171, 1819, 2257, 3865, 4259, 1836, 4121, 4575, 3526, 1721, 823, 1288, 2718, 1891, 1944, 2222, 3735, 3174, 3161, 4166, 3929, 1127, 1867, 2864, 3965, 4635, 3688, 1720, 3162, 3623, 2727, 902, 3587, 2826, 1810, 4430, 1148, 3248, 4619, 888, 3165, 1921, 3498, 1586, 1072, 2702]. performances [4304, 3984, 3159, 4295]. performing [3099]. Period [175, 118, 2447, 3745, 3580]. period-radius [3745]. periodic [1909, 3846, 3261]. peripheral [4585]. Permutation [491, 1872, 2440, 2266, 1993]. persistence [2968, 2526, 2826, 1036, 2628]. persistent [300, 462, 1960]. person [3259]. person-time [3259]. personal [20]. Personality [3076, 3586]. Personalized [4477, 4040, 2960, 4634]. personnel [1084]. perspective [760, 4169, 4097, 759, 4321, 762, 2391, 3513, 1173, 4560, 1197, 4153, 2223, 2127]. perspectives [4533]. perturbation [563, 3472]. Peter [4669, 2748]. Petersen [2819]. petroleum [2435, 2140]. Ph.D. [3169]. pharmaceutical [3439, 2326]. pharmacodynamic [1600]. pharmacokinetic [1600]. pharmacokinetics [1285]. Phase [4647, 2129, 3137, 2147, 1872, 2314, 682, 1161, 1324, 2930, 3677, 4333, 1834, 4522, 1481, 2068, 3259, 1720, 2209, 4463]. Phase-I [4463]. PhD [1931]. phenological [248]. Phenomena [2320, 676, 2626]. phenomenon [2888, 3846]. phenotype [3981]. phenotypes [2560]. phenotypic [4663]. Phi [1665, 1445, 3183]. Phi-divergence [1665, 1445, 3183]. Phillips [305]. philosophical [2613]. Phono [1382]. Phono-genesis [1382]. phosphorus [2982]. phosphorylation [4632]. photographs [820]. photon [663]. photovoltaic [3725]. Phylogenetic [2373, 4266]. physical [2995, 75, 4559, 4531]. Physically [3949]. physico [2658]. physico-financial [2658]. physics [1884]. Physiological [210]. Piccarreta [1923]. Picek [2747]. pickle [957]. Picture [101]. Picturing [2274, 4720]. piece [4273, 695]. Piecewise [1709, 4005, 680, 2549, 2287, 4363, 3090, 1990]. Piecewise-Deterministic [4005]. piecewise-linear [2549]. pig [3970, 3410]. pigeons [757]. Pine [1383, 1792, 790]. pinnule [692]. pipeline [4493]. Piracicaba [4503]. pit [2785]. Pitfalls [152, 4207]. Pituitary [1389]. Pivotal [3748, 3431]. pixel [1639]. pixels [4128]. placement [3345]. placements [3737]. Plackett [1191]. Plackett-Burman [1191]. Plague [187]. Plaid [1357]. Plan [1408, 1388, 1906, 1965, 2032, 1545, 1917, 1280, 4052, 1083, 1180, 1273, 990, 885, 913, 1086, 1094, 1969, 1041, 1555, 2375, 462, 4263, 4560, 675, 2867, 503, 819, 4094, 905, 963, 1726, 861, 935, 969]. planar [4421, 3599]. planets [3115, 3407]. Planning [1472, 1815, 1900, 3282, 2486, 1563, 2297, 33]. Plans [1680, 4226, 3907, 1977, 1029, 1059, 3671, 932, 3865, 3784, 596, 3773, 1516, 1287, 944, 1338, 1314, 506, 637, 816, 617, 681, 1828, 643, 1868, 3612, 3371, 1957, 252, 392, 1293, 544, 556, 645, 683, 718, 709, 826, 3851, 400, 3249, 412, 575, 557, 621, 839, 2075, 2259, 1339, 905, 963, 1099]. Plant [1398, 1642, 470, 1278, 1775, 2105, 2242, 3618, 1336, 589, 4570]. Plantation [1615]. plantations [790]. plasma [1285]. plastic [3449]. plateaus [1994]. platform [1910]. Plato [233]. play [952]. play-offs [952]. player [4047, 4121, 4575, 3382]. Players [1668, 3181, 3775, 2001]. playing [3382, 1098]. plethora [4492]. Plot [1483, 1708, 1620, 1134, 546, 2101, 1596, 1896, 3487, 3379, 3874, 2512, 3146, 854, 1849, 4386]. Plots [182, 2181, 3666, 2216, 4603, 1347, 2765, 1105, 2104]. plotting [2765]. PLS [4575, 2379, 2914, 2028]. PLS-PM [4575]. plug [2562]. plug-in [2562]. plus [3927, 3093]. Pluto [1267]. PM [4575, 2914, 4332]. PM2.5 [4244, 4201]. pneumococcal [2212]. Point [4687, 1820, 4534, 2657, 855, 780, 2117, 3232, 4043, 3328, 3423, 609, 2240, 4638, 4340, 4434, 4645, 3488, 3470, 302, 3883, 978, 3137, 3221, 1074, 3561, 4601, 2626, 2389, 2397, 3123, 2390, 1904, 3563, 2496, 3495, 3024, 1461, 2716, 1767, 2867, 2146, 1048, 3882, 1703, 3741, 4117]. point-pattern [3024]. Point-symmetry [2657]. pointing [3655]. points [3256, 1553, 1966, 3883, 2575, 2287, 1867, 2759, 4363, 2734, 4421, 4508, 3258, 2695, 874, 3727, 3189, 2783, 2841, 1151, 859, 3940, 4561]. Poisson [2677, 4364, 3634, 4152, 3357, 548, 1190, 1894, 2664, 397, 3980, 4233, 3708, 2866, 1413, 4438, 3574, 3788, 1276, 2000, 450, 3722, 807, 2725, 1103, 2250, 3600, 2705, 4089, 4192, 4369, 3050, 2870, 3259, 2529, 4310, 1296, 2485, 781, 2596, 379, 457, 4493, 3742, 4300, 4357, 2185, 3125, 2330, 4315, 1771, 2937, 4524, 4164, 504, 4488, 2389, 4011, 1132, 2496, 4254, 2783, 3991, 4378, 3886, 3911, 3533, 2769, 2784, 4255, 1579, 4332, 3275, 4258, 1726, 1015, 1016, 1164, 2120, 3501, 2046, 2856, 2122, 2700, 4346, 4548, 4571, 4190]. Poisson-exponential [2185, 3125, 3533].

Poisson-Lindley [3501]. polarization [2437]. police [2702]. policies [64, 3627]. Policy [173, 3899, 3500, 2613, 2496, 1098, 2600, 2791, 1329, 3013]. policy-making [2496, 3013]. polish [2785]. polishing [604]. political [483]. politics [50]. poll [2, 824]. pollen [4484]. pollination [3133]. polling [24, 1745]. polls [246, 459, 36]. pollutant [584]. pollution [2472, 3709, 793, 2360, 791, 2601, 2600]. Pólya [4633]. Pólya-Gamma [4633]. polycross [3133]. polygon [3599]. polygonal [452]. polygons [1582]. polymer [3558]. polymorphism [2569, 2650]. Polynomial [1613, 287, 3329, 938, 874, 321]. polynomials [2166, 3314, 2087, 2256, 3312, 447, 321]. polyphenylene [3322]. Polystats [9]. Polytomous [585]. polyvinyl [3433]. pond [4240]. Pontypool [17]. Pooled [1631, 1380, 3456, 3807]. poor [4005]. POPAN [748, 1244]. POPAN-4 [748]. POPAN-6 [1244]. Population [1666, 1414, 754, 4377, 1627, 1114, 1081, 2804, 763, 735, 1266, 4331, 1542, 759, 1241, 4019, 2131, 2414, 4369, 1370, 1757, 3015, 4542, 2009, 762, 953, 4172, 3361, 1352, 1735, 756, 741, 736, 1240, 2152, 1759, 3278, 884, 2834, 4229, 2855, 2188, 3864, 3550, 2946, 4527]. Populations [1693, 1708, 748, 4126, 760, 1873, 4687, 3239, 4211, 3404, 1992, 2735, 3092, 4507, 4610, 266, 1465, 1079, 4510, 765, 3946, 4555, 1927, 1229, 4662, 1903, 2197, 3149, 2188]. porcine [2597]. port [983, 1167]. portable [450]. portfolio [3455, 3904, 4058, 3812, 3431]. portion [4278]. Portmanteau [698, 1171, 4602]. Portugal [1050, 1467, 4026]. posed [1151]. position [3241]. Positional [140]. positions [2765]. Positive [1413, 4625, 4389, 3810, 4398]. positively [2416]. positives [2853]. positivity [4339]. positron [418, 2625]. possibility [2674, 2177]. Possible [1634, 492, 4674]. possibly [4077, 1255]. Post [4055, 1972, 2584, 966, 1716]. Post-hoc} [1716]. post-imputation [1972]. post-operative [966]. post-randomisation [2584]. Post-randomization [4055]. post-test [1291]. Posterior [2733, 3254, 2548, 3854, 1957, 722]. postfledging [1227]. postponing [3543]. POT [3690]. Potential [155, 174, 1239, 353, 2673, 2511, 3725]. potentially [495, 2742]. potentials [1867]. Poverty [1704, 2438, 2994, 2269, 4404, 4130, 2213]. Power [3217, 3224, 2301, 1650, 973, 1380, 1642, 2708, 838, 3670, 866, 1628, 1422, 909, 997, 4359, 3869, 4324, 4139, 1125, 4505, 3753, 4059, 3666, 1265, 4155, 3648, 1321, 1010, 3662, 2560, 999, 3104, 4195, 1043, 1871, 733, 4140, 1488, 4345, 1054, 1129, 2722, 4348, 283, 4667, 3802, 886, 1939, 2473, 4382, 1196, 3099, 2362, 484, 496, 2638, 985, 1723, 4436, 3757, 2527, 4041]. Power-divergence-type [909]. Power-law [2708, 3869]. power-transformed [1010]. power-weighted [4667]. powerdivergence [1185]. Powerful [2196, 2654, 3662, 3844, 4602, 4464, 2778]. powers [3741]. pp [4599, 3558]. PPP [1384]. Practical [117, 2476, 140, 3719, 19, 1276, 1863, 317, 2413, 3401, 2749, 4446]. Practice [4732, 3781, 2953]. Practices [4714]. practitioner [759, 2610]. practitioners [1357]. Pradel [1239]. Pradesh [2280]. pragmatic [1144]. Pranab [2747]. pre [2042, 2584]. pre-randomisation [2584]. pre-smoothing [2042]. preceding [4384]. precipitation [3720]. Precision [196, 4376, 1162, 2061, 1066, 550, 706, 4275]. predict [2250, 788, 4421, 4178, 2251, 2412, 1152, 3918, 4402, 824]. Predictability [1701, 2418]. predicted [2534]. Predicting [957, 2106, 2664, 2995, 3660, 1126, 1920, 3774, 1803, 1163, 2591, 4676, 3702, 4099, 1520, 1372, 1272, 2866, 3211, 2595, 3986, 4102, 2775, 4520, 3495, 2587, 1586]. Prediction [471, 2912, 724, 4063, 1607, 3936, 4490, 1642, 1827, 487, 1383, 618, 1065, 4688, 3440, 2082, 1954, 2974, 4268, 4118, 2797, 1493, 938, 2507, 2964, 3381, 4590, 4048, 302, 3337, 3877, 4663, 2641, 3824, 1938, 2742, 3879, 3418, 2391, 4103, 3933, 1344, 2421, 3281, 3210, 1773, 2455, 3502, 3687, 875, 2770, 4486, 1810, 3449, 3599, 2246, 3100, 3240, 3248, 3126, 3403, 2636, 549, 1557, 1716]. Predictions [566, 668, 4442, 3212, 4683, 1054]. Predictive [4226, 4679, 1283, 3587, 829, 3427, 4546, 2316, 1254, 4566, 2727, 3820, 3195]. Predictor [3033, 2977, 2331]. predictors [806, 2876, 4309, 4328, 4079, 2958, 1874, 2173, 2339]. Preference [1638, 794, 3252, 1727, 1499, 3171]. Preferences [1681, 3279, 4028, 3586]. preferential [3655]. pregnancy [4609]. Prehistoric [158]. Preliminary [992, 8]. premium [2416, 1512]. Preprocessing [3245, 3976]. Prescribing [1647]. prescription [4390]. Presence [1634, 1393, 1686, 997, 2687, 3260, 3079, 4469, 1354, 2564, 3556, 3865, 4043, 4118, 1169, 555, 3648, 3735, 4075, 3818, 2532, 2241, 1116, 3014, 3507, 4310, 3020, 2436, 2796, 2885, 3339, 1984, 2550, 1569, 4002, 2608, 1529, 2202, 3756, 4482, 1947, 1749, 3031, 3796, 4600, 2146, 4146, 1295, 3845, 3729, 2261]. Presentation [87, 332, 284]. presentations [3689]. presidential [4492, 4014, 2768]. PRESS [3739]. pressure [1472, 3984, 1874]. prestige [2290]. pretest [1291]. pretest/post [1291]. pretest/post-test [1291]. prevalence [1758, 4119, 4550, 3049, 2493, 3946, 2377, 4629]. prevention [4458]. price [1912, 2864, 3840, 597, 1983, 3006, 2780, 2201, 3462, 2980, 3498, 4108]. prices [1855, 3234, 2797, 3309, 4042, 8, 2742, 3827, 3110, 3342, 4093, 3669, 3483]. pricing [1728, 1834, 3974]. primaries [4492]. primary [609, 2264]. prime [4015]. primer [2508]. Principal [23, 485, 343, 3446, 2560, 1677, 3940, 3079, 1525, 3517, 703, 1979, 4424, 2586, 2489, 4036, 1791, 3023, 3315, 691, 3975, 4580, 3944, 4293, 3097, 355, 4120, 375, 2601, 3308, 3299, 859, 2411, 3801, 3580, 4653, 3584, 4561, 4044, 2376]. Principle [237, 1502]. Principles [4748, 4732, 1512, 2953, 3572]. printed [234]. ’Prior [1215, 1113, 1214, 2549, 3460, 553, 1894, 1796, 1476, 2686, 1312, 2672, 1254, 1105, 3650, 3596, 4378, 1730, 722]. prior-data [1796]. Priority [1611]. priors [2887, 3621, 1925, 419, 3522, 4298, 3697, 358]. prison [448]. Privacy [3685]. private [2980]. Prize [2230, 2524]. Probabilistic [3837, 113, 2913, 3650, 4008, 3682]. probabilists [2857]. Probabilities [2664, 554, 2548, 735, 4007, 3912, 1219, 1797, 744, 3854, 2293, 1105, 2190, 3401, 504, 3726, 1227, 1345, 1048]. Probability [895, 11, 2228, 1396, 3766, 1414, 4728, 120, 4701, 150, 2320, 4737, 4719, 4708, 3970, 631, 3138, 1923, 2828, 729, 4113, 3405, 3961, 2765, 1105, 3855, 3384, 2579, 4631, 2992, 2248, 800, 2217, 3257, 2074, 339, 3496, 2652, 2104, 2158, 2824, 1882]. Probit [4394, 3215, 3700, 2148]. Problem [1633, 1390, 1403, 161, 1656, 170, 3239, 1780, 937, 464, 238, 917, 4485, 4182, 3557, 3514, 632, 3909, 2889, 3954, 3032, 4476, 1057, 2343, 592, 1767, 3372, 3553, 688, 1199, 3583, 3940, 3201]. Problems [509, 107, 1301, 644, 407, 1421, 1096, 153, 3460, 4495, 4687, 2706, 1242, 2347, 3340, 2326, 422, 978, 469, 2406, 75, 3528, 3627, 235, 346, 22, 820, 493, 1734, 1151, 45, 3540, 3177, 4444, 4465, 510]. Procedural [155, 174]. Procedure [963, 1702, 3400, 2881, 372, 3181, 2670, 1558, 3709, 4375, 3785, 908, 1785, 3904, 2341, 395, 3778, 4506, 2466, 2184, 266, 3273, 1859, 1575, 1336, 1133, 3099, 3898, 3924, 1501, 2217, 4569, 1726, 3208, 4420, 1921, 2915, 4349].

Procedures

[2852, 1379, 141, 544, 556, 645, 709, 826, 153, 575, 776, 3079, 106, 1523, 3390, 4248, 2552, 842, 2576, 919, 1334, 796, 804, 2207, 1771, 4672, 2077, 1801, 3540, 3288, 4613, 1498]. Process [1423, 1409, 1450, 1659, 1692, 4005, 4002, 164, 1446, 1618, 1702, 1112, 2380, 1842, 4523, 4647, 2810, 4046, 2023, 4650, 2420, 1729, 4684, 3232, 3760, 865, 598, 4286, 4052, 609, 1137, 2866, 4423, 3532, 2060, 4638, 3821, 1142, 2222, 2165, 4302, 3804, 3470, 4478, 2085, 4534, 1900, 3985, 3960, 967, 3945, 301, 4335, 721, 3180, 781, 4554, 2375, 919, 1064, 379, 457, 815, 2469, 4300, 2116, 1935, 2291, 4204, 1514, 462, 1957, 2389, 1942, 2838, 2972, 941, 1529, 2356, 737, 2783, 4326, 886, 4378, 1973, 4412, 3421, 2189, 4022, 2197, 1146, 2513, 1111, 1921, 2169, 2475, 1327, 3868, 4360, 858, 921]. process [2848]. Processes [1684, 4728, 1679, 1655, 1397, 1701, 1428, 1703, 2051, 4098, 1271, 4536, 3655, 471, 1545, 2447, 2835, 343, 3708, 802, 2718, 3203, 2110, 1000, 3408, 2147, 2615, 719, 1765, 2893, 3742, 2442, 1870, 2937, 1772, 3422, 4033, 2798, 3495, 1945, 1547, 2267, 4353, 1150, 920, 1856, 2120, 2369, 4307, 4346]. Processing [451, 4712, 453, 364, 3322, 112, 317, 2055, 3558]. Procrustes [510, 509, 407, 1940]. produce [2907]. producers [3618]. Product [1423, 3123, 829, 4344, 1975, 546, 1944, 2222, 1127, 1284, 1119, 3542, 1627]. product-based [829]. Production [1388, 122, 1545, 2713, 2281, 1469, 2406, 1041, 346, 619, 3838, 1569, 1772, 4260, 2477, 3800, 2863]. productions [3777]. Productivity [758, 104, 130, 771, 88]. Products [1638, 3769, 1148, 1499]. professional [3532, 2182]. Professionalized [138, 132]. Professor [1715]. professors [4189]. proficiency [4473]. Profile [897, 3843, 2381, 3985, 4063, 4104]. Profiles [3691, 4246, 2876, 2432, 3244, 3137, 4636, 2024, 4353]. profiling [3672, 2760]. profitability [3456]. Prognostic [1430]. prograde [4442]. program [240, 3659]. programme [758]. programmes [1353]. Programming [4697, 3034]. Programs [1707, 746]. Progress [183]. progression [843, 3207, 4544, 2268, 4621]. progressive [4359, 4226, 2805, 4469, 1809, 932, 3784, 3349, 4581, 4354, 4365, 4590, 3965, 2724, 3397, 2454, 3947, 4612, 4313, 4351, 3435, 4236, 2515, 2924, 4662, 2539, 1924, 883, 1110, 1339, 1302, 1804, 1815]. Progressively [1516, 3769, 3440, 4004, 2935, 4305, 1981, 2350, 3748, 4672, 4497, 3291, 2799]. Projected [1675, 938, 1528]. projecting [1143]. projection [1413, 4210, 3188]. projection-based [3188]. projections [1050, 1467, 421]. Promislov [4402]. promote [3456]. promotion [837]. Promotions [1181]. prone [2701, 2571]. proof [572]. propagation [4251, 2133]. Propensity [3824, 3214, 1476, 2766, 3752, 3619]. propertie [3322]. Properties [1721, 2718, 1537, 1446, 1690, 4324, 1722, 1244, 4650, 278, 4318, 4068, 615, 1887, 579, 2744, 1496, 3889, 4356, 4134, 4334, 3915, 3977, 4192, 3710, 1183, 733, 2413, 834, 4431, 866, 4362, 4675, 4661, 4525, 2855]. property [2368, 1827]. prophets [2106]. Proportion [1414, 4640, 4156, 3193, 2868, 4388, 1967, 2241, 2519, 3015, 3266, 827, 1588, 3743, 3758, 3076, 2693, 1058, 2223, 2588, 4107, 3550, 4465]. Proportional [3809, 1707, 2034, 3786, 4415, 3122, 1294, 2355, 2942, 4347, 3117, 4256, 4241, 4230, 2634, 4355, 1283, 4084, 3649, 2208, 2854, 4319, 3730, 4110, 3410]. proportional-hazards [4415, 4241]. Proportions [1381, 1429, 3433, 2430, 4596, 3089, 4678, 2957, 2753, 1584, 745, 4507, 3764, 2207, 4397, 3608, 1541, 4414, 3610, 1865, 2624, 1501, 4115, 3078, 4622, 2588, 3222, 2895]. Proposal [155, 4049, 2765, 3145, 174]. proposed [2269]. prospective [1566]. prospects [3482]. prostate [4462, 4281, 2410, 2412, 2390, 3046, 2137]. prostate-specific [3046]. prostatectomy [4639]. Protection [1408, 3685, 4170]. protein [3504, 2930, 3704, 4632, 2514]. proteins [3036]. protests [4303]. Proto [1382]. Proto-Australian [1382]. protocol [888]. prototypes [421]. provided [901]. providers [2760]. Proximal [3381]. proximities [4231]. proximity [2688, 3305]. proximity-relation [2688]. pseudo [3994, 3491, 264, 1107, 2470]. pseudo-coefficient [3491]. pseudo-exponential [3994]. pseudo-likelihood [1107]. pseudo-panel [2470]. pseudo-random [264]. pseudolikelihood [2803]. psychiatric [2604, 1197]. psychotherapy [3714]. Public [154, 3356, 4503, 58, 3746, 4618, 2907, 3013]. public-use [2907]. publication [236, 3151]. publish [3908]. Pulak [1970]. pulmonary [2912]. pulp [1314]. pulse [1874]. punishment [611]. pupil [3425]. pupils [591]. purchases [983, 4053, 1167]. pure [301, 589]. purpose [2864]. Purposes [1430]. putative [791]. Puzzle [12].

Q [111, 1315]. Qatar [4433]. QSN [55]. QSVSS [1090]. QTLs [1775]. Quad [2368]. quadrat [1571, 982]. Quadratic [3857, 948, 1770, 2444]. quadrature [2026]. qualitative [2186]. Quality [1451, 1423, 2532, 1397, 2380, 2585, 1545, 1985, 106, 771, 979, 717, 1018, 3089, 2718, 1102, 2966, 792, 989, 320, 3170, 506, 565, 637, 2457, 1041, 4371, 567, 1064, 1765, 3742, 4489, 3961, 1365, 2798, 1908, 718, 2908, 1229, 1515, 1579, 646, 819, 2834, 4219, 3550, 4521]. quality-of-life [2585]. Quantal [1854, 2292]. Quantification [4262, 2352, 3219]. Quantifying [3970, 4230, 1984, 4128, 2106]. Quantile [3872, 3631, 873, 2871, 3860, 1928, 3198, 2383, 2984, 3460, 3575, 3569, 3399, 4656, 3057, 3835, 4657, 4418, 3161, 3356, 3873, 3678, 4195, 831, 2316, 4400, 2640, 2312, 1073, 4393, 4076, 1089, 3737, 4428, 4466, 3819, 3809, 3190, 3509, 4381, 2631, 4644, 3995, 3454, 3087, 1741, 893, 3060, 3233, 3876, 3696, 3364, 2927, 2915, 2311, 2878]. Quantile-adaptive [3198]. quantile-treatment-effect [1741]. quantiles [1523, 1598, 3690, 972, 3922, 4093, 893, 3462, 2533]. Quantitative [344, 2503, 148, 1139, 1493, 2073, 4468, 864, 1187, 2624, 2587, 2482, 3516, 2448]. Quantity [144, 1092, 3431]. quantization [656]. quantum [3636]. quarantine [3615]. quartiles [3278]. Quasi [3989, 2552, 1107, 1448, 2378, 1807, 1232, 4620, 4066, 2581, 3162, 1056, 1326, 1231, 1281]. Quasi-asymmetry [2378]. Quasi-binomial [3989]. quasi-experiments [1232, 1231]. Quasi-Least [1448]. Quasi-likelihood [1107, 4620, 4066, 1056, 1326]. quasi-likelihoods [1281]. quasi-maximum-likelihood [2581]. Quasi-optimal [2552]. quasi-range [1807]. Queen [2902]. query [2059]. question [4065]. Questionnaire [3613]. questionnaires [4199, 3264]. questions [751, 1750, 113]. queueing [2969]. Queues [4719, 2228, 1564]. queuing [261, 970]. queuing-theory-based [970]. Quick [596, 1745, 1090, 2548, 707, 697, 910, 567, 466, 581, 646]. Quiz [29, 53, 40]. quotations [25].

R [2091, 165, 2814, 2192, 2653, 4019, 4711, 2618, 2616, 2423, 2424, 4697, 4699, 4713, 4729, 4692, 4700, 1493, 1863, 2464, 2508, 2272, 2578, 2609, 3542, 3253, 2191, 4717, 4744, 2456, 4707, 4746, 3696, 4708, 2461, 544, 556, 645, 709, 2276, 2275, 2952, 2090, 1961, 2949, 3021, 3026, 1882, 1881]. r-Part [556]. r-size-biased [4019]. R-squared [1493]. R. [207]. R/super [3696]. R/super-quantile [3696]. race [895, 304]. race-track [304]. racial [4070, 4406]. racing [3081]. radar [662, 3576]. radial [2772, 4336]. radiators [3180]. radical [4639]. radio [1230, 738]. radio-tagged [738]. radio-tagging [1230]. radiocarbon [2547]. radius [1456, 3745, 3129]. radon [2701]. radon-prone [2701]. ragweed [4484]. rail [3797]. rain [1131, 2726]. raindrops [46]. rainfall [3950, 2828, 2288, 3715, 4524, 281, 3776, 755, 3576, 3088]. rainier [3767]. rainy [2106]. raise [3456]. Raluca [2651]. Rand [2581]. RandGA [3001]. Random [1986, 1419, 1681, 167, 4728, 416, 1636, 2320, 1884, 1223, 3992, 2146, 2056, 3059, 4640, 4046, 1128, 3215, 3573, 4296, 3510, 2545, 3784, 1217, 3660, 4670, 2603, 2900, 3574, 2342, 1829, 450, 472, 4581, 3402, 3535, 1219, 2656, 1269, 264, 1757, 2956, 1785, 3015, 3877, 2015, 4630, 1778, 4363, 4593, 3247, 3611, 3300, 4393, 1218, 1768, 486, 787, 4063, 725, 1576, 2009, 1021, 1129, 2214, 2111, 3438, 3443, 1548, 2511, 2405, 3958, 1885, 2894, 3807, 462, 4299, 4210, 4381, 4231, 835, 4011, 4015, 3111, 1529, 424, 685, 2963, 3538, 2947, 1947, 4574, 4623, 3628, 4022, 883, 4502, 2109, 4481]. random [4562, 1274, 3856, 3055, 3732, 1302, 3222, 2188, 4446, 2338, 4398, 4190]. random-coefficient-based [1274]. random-effects [3510, 2545, 2342, 2109]. randomisation [2584]. randomization [851, 4088, 4055]. Randomized [1414, 3193, 897, 1975, 3350, 3685, 3224, 1934, 3521, 2440, 842, 3624, 2084, 1928, 1297, 3236, 3087, 3758, 4112, 2264]. Randomly [1453, 3336, 2632, 3541, 1966, 2936]. randomness [3001]. Range [253, 4234, 4243, 3199, 1807, 1542, 4284, 3073, 4567, 2203, 1477]. ranges [2110]. Rank [3277, 1391, 1150, 3142, 3767, 2941, 4301, 2882, 3297, 4316, 3638, 3118, 4010, 3927, 2386, 4567, 4024, 4249, 505, 2108, 2140, 3154, 3967, 4403, 961, 2929, 1055, 4431, 4029]. rank-based [3767, 4029]. rank-deficient [4249]. rank-sum [3154]. Rank-Sum-Type [1391]. Ranked [1451, 1671, 2380, 2804, 872, 1190, 3350, 4161, 2909, 4248, 2491, 787, 3141, 1355, 1195, 1365, 1982, 3394, 3377, 3550, 1308, 4035]. ranked-set [872, 1355, 1982]. Ranking [1638, 1750, 3081, 3181, 1985, 3279, 2070, 1779, 2505, 3627, 3241, 2433, 3499, 3308, 3171, 3629]. rankings [426, 3297, 3803, 3586, 4061, 3934]. rankit [1089]. ranks [2450, 698]. Rao [3662, 1260, 3857, 1343, 2444]. Rapid [4570, 4291]. Rare [1673, 3556, 2637, 2879, 998, 3434, 2366, 551, 779]. RAROC [2826]. Rasch [2684, 3176]. Rate [1707, 1622, 3805, 2968, 3214, 2098, 107, 4143, 758, 3520, 1241, 3518, 3417, 1500, 3930, 2404, 2054, 3661, 1131, 2982, 1617, 3948, 3147, 3058, 465, 1024, 2202, 3916, 779, 3853, 2220, 3006, 2362, 1579, 4526, 1924, 991, 4569, 2791, 722, 1955, 3220, 2415, 4291, 3729]. ratemaking [3788, 1859]. ratemeter [108]. rater [2025, 1754, 2662, 1066]. Rates [1619, 1429, 1620, 1278, 3480, 2102, 3760, 3468, 2998, 734, 482, 4545, 807, 732, 1243, 1159, 1920, 1226, 849, 3368, 1160, 2581, 3764, 753, 1753, 1230, 4510, 3495, 4512, 4414, 766, 2584, 3178, 1240, 768, 4402, 2703, 755, 2791, 4622, 3852, 1328, 3810]. rating [3663, 1797, 2505, 2662, 867]. ratings [1317, 3777, 1797]. Ratio [136, 1627, 829, 2968, 2307, 929, 1975, 927, 4655, 464, 2355, 2301, 2249, 1783, 4545, 4166, 3318, 3038, 2414, 79, 3956, 3738, 3537, 2436, 4507, 1053, 3073, 4104, 2084, 1774, 3855, 2860, 2251, 3727, 2397, 3058, 3098, 3343, 2707, 4008, 2784, 1240, 2270, 4394, 2217, 3054, 2638, 1199, 4591, 3353, 4366, 3377, 2303, 4399, 2527]. ratio-based [829, 1783, 1774, 3098]. Ratio-Estimates [136]. Ratio-product [1627]. Rationality [1668]. ratios [2733, 3254, 4499, 2715, 4214, 2374, 4474, 446, 2570]. rats [527]. Rattle [2578]. raw [717]. ray [3766, 4221, 4018]. Rayleigh [1964, 4047, 3710, 3891, 3879, 3748, 2660, 2924, 3476, 1580]. Raynaud [2888]. razor [785]. Re [1233, 3028, 1837, 1250, 2165, 4503]. re-admission [2165]. Re-analysis [1233]. re-examined [1837]. re-parameterized [4503]. re-sampling [1250]. re-weighted [3028]. react [3814, 3807]. reaction [880]. reactions [3061]. reactor [1152]. read [3353]. reader [1466, 2918]. reader-based [2918]. readers [3942]. Reading [207, 3425]. readings [1466]. readmission [4034]. Readouts [3570]. reads [3672, 3068]. reaggregation [3027]. Real [3355, 1232, 1660, 1231, 4664, 2034, 4139, 2968, 2099, 3722, 3390, 1261, 4140, 233, 3544, 2260, 2662, 113]. real-life [4139, 4140]. Real-time [3355, 3390, 1261, 3544, 2260]. reality [2834]. realized [1806, 3212, 3276]. really [3659]. realms [3929]. reasoning [1005]. recall [3791, 4513]. recall-based [3791]. recapture [900, 748, 1238, 1221, 763, 735, 749, 761, 1203, 1232, 1225, 1228, 1214, 744, 1239, 733, 1226, 764, 745, 254, 1224, 751, 1237, 738, 1206, 1223, 1240, 1215, 1231, 1200, 1114, 752, 2229, 2540, 1207]. recapture-sampling [254]. recaptures [1243]. receiver [2853, 2126, 3192]. receiving [3664]. receptor [4244]. recession [2714]. recessions [3774]. recipients [3184]. recoding [2585]. recognition [2295, 654, 3367, 2121, 659, 1556]. recombination [4468]. Reconceptualizing [4008]. reconciliation [3667]. Reconsidering [1340]. Reconstructing [2777, 418, 801]. Reconstruction [452, 421, 4302]. record [1040, 3476, 3403, 3955, 3744, 1262, 4247, 4349]. recorded [1850, 706]. Records [1660, 3936, 2353]. recover [1130]. recoveries [757, 732, 769, 1243, 765]. recovering [1786]. recovery [734, 749, 737, 1222, 4631, 2376]. Recreation [186, 547]. recruitment [1235, 1211]. rectangle [3092]. rectangle-screened [3092]. rectangular [674]. rectification [462]. recurrent [2553, 3294, 3057, 3637, 3507, 4335, 2875, 4069, 3601, 3125, 3095, 3067, 2310, 3733]. recursion [4275]. Recursive [1624, 955, 3692, 2641]. recycling [1197]. red [38, 662, 2254, 2452, 2902]. Redefining [3361]. reduce [4100, 2861, 4095, 3714]. reduced [4396]. Reduction [1765, 1624, 2913, 3456, 4267, 1527, 3380, 3305, 1935, 4031, 4496, 605, 3733]. reductions [2721]. redundancy [1242]. reentry [4442]. referees [2359]. Reference [144, 162, 872, 2739, 2492, 814, 1462, 905, 963, 2022, 2311, 4444]. refined [2877]. refining [2435]. Reflective [1415, 2914, 3792]. REFLEX [3494]. reforms [962]. regime [1941, 3471, 2263, 3945, 4537, 1586]. regime-switching [1941, 2263]. regimens [3624]. regimes [3214, 3167, 4646, 2315]. Region [4026, 3383, 2181, 528, 3912, 2828, 2674, 590, 1762, 2570, 4330]. Regional [4244, 1758, 4119, 4373, 281, 3237, 3669]. regions [3517, 927, 949, 2112, 4545, 2957, 1082, 2438, 4293, 4629]. registration [4302, 4480]. registry [2511, 2801]. Regression [4375, 3647, 1662, 4075, 4485, 2922, 1429, 1442, 2360, 1421, 4421, 1153, 1165, 1576, 2875, 1607, 1449, 1616, 1195, 1046, 3101, 2050, 2195, 197, 2751, 3991, 4690, 4718, 994, 1448, 200, 1475, 3713, 1623, 4193, 4624, 3331, 775, 4144, 2042, 3388, 4252, 1259, 1249, 3085, 2881, 2383, 2984, 3460, 3575, 4114, 3569, 3028, 3970, 1769, 3399, 3679, 4163, 1456, 91, 2082, 1291, 1356, 1030, 3336, 2779, 3837, 335, 2102, 4273, 3270, 2145, 4415, 2913, 3556, 2873, 4656, 2995, 958, 277, 314, 337, 1775, 2637, 3836, 4587, 1292, 2869, 4331, 2800, 3310, 1060, 1595, 2941, 3479, 3653, 3590, 4198]. regression [3057, 2216, 3329, 3788, 3835, 1276, 680, 4657, 3211, 4023, 2355, 2942, 2166, 2790, 3552, 3872, 3892, 1493, 2335, 3402, 3161, 4159, 4302, 3080, 3391, 3314, 4434, 3356, 3570, 3755, 4509, 862, 3525, 3011, 3187, 3063, 1746, 3292, 2250, 3630, 2007, 3519, 3600, 2705, 571, 2925, 3798, 4515, 4598, 3962, 812, 4036, 302, 729, 1958, 3150, 3042, 3002, 4089, 3678, 600, 1101, 4160, 3989, 3050, 4179, 3871, 2956, 4194, 1867, 495, 1584, 563, 3390, 3380, 3591, 3986, 4503, 396, 507, 2529, 3255, 2640, 406, 3359, 2407, 3631, 3472, 2268, 2312, 1524, 1852, 2734, 2004, 2883, 3135, 3406, 4660, 3782, 4076]. regression [2945, 3771, 2114, 486, 539, 723, 2590, 2221, 4063, 3465, 4479, 2596, 4240, 889, 1737, 3105, 3263, 3768, 4519, 3737, 344, 4428, 4466, 1509, 4283, 3197, 2159, 3764, 3884, 2205, 3185, 3186, 4010, 3933, 4348, 4352, 4520, 4470, 4489, 4642, 2211, 3130, 3601, 4357, 3141, 2067, 2958, 3033, 2421, 2558, 3874, 3826, 1095, 4315, 1002, 2871, 3190, 3860, 3225, 3603, 2024, 1823, 3585, 4074, 3143, 3333, 1590, 1915, 3451, 3646, 2943, 3743, 2777, 992, 1808, 3509, 3166, 3626, 874, 1902, 1827, 4164, 4361, 3469, 2626, 4381, 4606, 2631, 2053, 2852, 3189, 2750, 3587, 461, 2412, 2892, 3548, 3776, 3382, 4127, 4323]. regression [2710, 2831, 3563, 3618, 1473, 2496, 3969, 3684, 3363, 4666, 1014, 4292, 3802, 4216, 4414, 4631, 634, 868, 3756, 3820, 3604, 4338, 4298, 1916, 3385, 653, 3870, 3538, 1133, 3911, 359, 3396, 993, 3972, 2846, 857, 2473, 2218, 3200, 1499, 1734, 3918, 4382, 2978, 2265, 3607, 3649, 3126, 2451, 2204, 4136, 4202, 2031, 2208, 727, 4235, 2898, 893, 863, 1599, 2056, 399, 3696, 3364, 3275, 4153, 3296, 4050, 4311, 4219, 4481, 2639, 4671, 3075, 3466, 3177, 4320, 1189, 1303, 1015, 1016, 1164, 1337, 1723, 1996, 2411, 2927, 3055, 3409, 3549, 3732, 4448, 1494, 1716, 3496, 3531, 3757, 2996, 1552]. regression [1995, 2046, 2856, 2915, 2138, 3799, 4147, 2760, 4133, 4392, 2863, 4027, 3220, 3059, 4641, 2339, 2700, 1593, 2531, 2118, 1498, 2338, 3900, 4372, 4342, 2698, 2763, 4457, 3625, 4465, 4628]. regression-kriging [3933]. regressions [4387, 1980, 3030, 3083, 4195, 3956, 2316, 3255, 3246, 1952, 3084, 3200, 3233]. regressive [2797]. regressors [3331, 1851]. regret [1083]. regret-balanced [1083]. regular [4051, 1230]. Regularisation [3216]. Regularity [2076]. regularization [4638, 3788, 2300, 4490, 4111]. Regularized [4235, 4448, 3570, 4561, 4453]. regulation [2897]. regulatory [3213, 527]. reinforcement [1168]. reinspection [4263]. reinterview [1364]. Reivew [436]. Reject [4176]. rejection [1781, 2016]. rejective [2606]. related [4687, 588, 3445, 2290, 3048, 2753, 1665, 3023, 2360, 1296, 4654, 1514, 2930, 4255, 2636, 1924, 2646]. Relating [153, 470]. relation [1330, 3655, 2688]. relational [4600, 3041]. relations [4383]. Relationship [4733, 4171, 1836, 4532, 4370, 3038, 1931, 4042, 852, 3588, 4531, 2818, 3006, 755, 1890, 3547]. Relationships [1475, 2399, 493, 689, 1204]. Relative [3136, 1702, 2787, 3318, 377, 3688, 3033, 4474, 2251, 2831, 4003]. Relativity [120]. relaxation [4399]. Release [3623, 733]. release-recapture [733]. releasing [4055]. relevance [2282, 2793]. Reliability [4047, 2820, 932, 440, 1828, 619, 3952, 4735, 1339, 1629, 4213, 2363, 4427, 3784, 3274, 4052, 4334, 4305, 1508, 2147, 3325, 3867, 720, 1969, 1555, 1805, 400, 4497, 3060, 4261]. reliable [4049, 4366]. relocation [2416]. Remark [1635]. Remarks [954, 26, 1178, 22, 3597]. remedial [4485, 4353]. REML [1348, 3443]. remote [1372, 1778]. remote-sensing [1778]. remotely [3949]. removal [3784, 4581, 2583]. removals [1809, 883, 1110, 1339, 1302]. removing [2646]. Ren [2091]. renal [3049, 2372, 3139]. renewable [3304]. renewal [4046]. repair [1784, 440]. repairing [4541]. Reparameterization [2081, 775]. reparameterized [4512].

Reparametrization [470, 1172].

repeatability [2061]. repeated [1249, 3193, 1926, 3049, 1182, 3521, 635, 3285, 1118, 1800, 896, 1119, 3849, 4667, 1739, 993, 1350, 2928, 2109, 3409, 4429, 1349, 4617, 4398]. repeated-prevalence [3049]. repetitious [1539]. Repetitive [1558, 3134, 3357, 3671, 1969, 4033]. replace [4420]. Replacement [1396, 3647, 3049, 4019, 2019]. replicate [1115]. Replicated [1626, 1448, 578, 1176, 4658]. replicates [887]. replications [599]. Reply [138, 52, 28, 1166]. Report [11, 121, 93, 2995]. reported [4608, 3201]. Reporting [1414, 734, 737, 2377, 3091, 1167]. Reports [162]. Representation [1398, 1587, 3796, 4637]. representing [3785]. reproducibility [2061]. reproduction [1228, 1204]. reproductions [2602]. reproductive [1235, 1211, 4546, 2597, 1204, 4041]. required [1351]. requirements [1066]. resample [4420]. resample-replace [4420]. resampled [3904]. resampling [1478, 2782, 1923, 1277, 2312, 2610, 995, 1032, 2952]. rescheduling [1797]. Research [4, 268, 207, 181, 2323, 426, 23, 2616, 3908, 2913, 97, 42, 3162, 4437, 1801, 2056, 1197]. reserve [3211]. reserving [3182]. reservoirs [2404]. RESET [1722]. resident [741]. residential [8, 4108]. Residual [1443, 1679, 4163, 501, 1535, 1958, 3631, 3143, 3363, 3916, 4445, 3454, 3060, 858]. Residualization [3977]. Residuals [1347, 373, 2204, 3007, 1746, 1852, 708, 2777, 334, 3969, 698]. resighting [1241, 736]. resilience [4087]. resistance [4396]. resistant [3072]. resolution [855, 417, 1372, 513, 4462]. Resolvable [498, 142]. Resolve [155, 174]. Resource [1644, 3432]. resources [4565]. respect [864, 1853, 1903]. respiratory [3634, 2597].

Respond [1619]. respondent [1499]. Response [1414, 1216, 1613, 1441, 1609, 3084, 1393, 1154, 1565, 2292, 3193, 3698, 2903, 2548, 1364, 2479, 3107, 3270, 3685, 3320, 3250, 220, 2244, 3479, 3653, 4198, 1847, 391, 1493, 1888, 2073, 2794, 4321, 3630, 2620, 1934, 1743, 3521, 1001, 947, 1369, 2711, 590, 1538, 723, 686, 4428, 1846, 1342, 3284, 1074, 1161, 2934, 1548, 2534, 3823, 1035, 4419, 4183, 580, 1908, 4556, 4326, 930, 674, 2923, 2704, 3236, 4623, 1874, 322, 3758, 2551, 1097, 1197, 920, 3856, 1849, 2720, 2696]. response-adaptive [2696]. response-surface [1035]. Responses [1679, 1713, 4367, 4296, 2099, 3217, 3452, 2926, 1228, 3939, 2153, 3678, 3285, 3272, 2963, 3293, 1990, 2177, 3649, 4353, 3093, 4071, 3761, 3116, 3789, 1750, 2278, 2480, 3410, 4628]. Responsibility [3306]. responsible [3597]. responsiveness [1503]. resting [764]. restoration [3664]. Restricted [1469, 4172, 4114, 3314, 1665, 3014, 4179, 4122, 3465, 2621, 4610, 4683, 2205, 1761, 3225, 3143, 3829, 2831, 4064, 2152, 2138]. restrictions [807, 2627, 3197, 4320]. result [2673]. resulting [2523]. Results [1793, 1648, 1954, 4121, 4048, 262, 3621, 2762, 1801, 3757]. resumptions [301]. resuscitation [2441]. retail [1855, 4679]. retest [3262]. retinopathy [3296]. retirement [1715, 4385]. RETRACTED [4140]. Retraction [4139]. retrospective [4277]. Return [753, 4171, 1731, 1226, 731, 3822]. returns [3427, 1037, 3358, 2133, 870, 2727, 4395, 2680]. revealed [2716]. Revealing [2212]. revenue [1330, 4186]. reverse [2930]. Reversibility [2499]. reversible [1255]. Review [242, 243, 244, 428, 429, 430, 431, 432, 433, 434, 435, 437, 438, 439, 455, 545, 582, 612, 1788, 2228, 2395, 2820, 1826, 223, 2090, 2124, 190, 177, 193, 2822, 2816, 2815, 860, 2394, 2817, 2951, 3572, 2194, 2814, 2227, 192, 191, 166, 179, 475, 2036, 189, 180, 1621, 4328, 2226, 2277, 2393, 2066, 2192, 2274, 2746, 2953, 2950, 1787, 1880, 1961, 383, 4669, 2745, 2824, 2819, 2125, 2823, 1987, 2276, 2225, 2949, 2064, 2158, 2273, 2195, 3026, 2191, 1824, 1825, 1883, 1882, 1884, 2821, 2948, 3120, 141, 194, 178, 2193, 2156, 1777, 1881, 2123, 2157, 2818]. Review [2275, 2065, 2747, 4599, 2653, 2952, 2037, 2748, 1789, 2651, 2652, 1854, 671, 890, 4272, 109, 1245]. Reviewers [3701]. Reviews [219, 204, 216, 228, 229, 230, 231, 255, 256, 257, 258, 269, 270, 271, 272, 273, 274, 275, 276, 288, 289, 291, 292, 293, 294, 295, 306, 307, 308, 309, 310, 311, 312, 323, 325, 326, 327, 347, 348, 349, 350, 351, 352, 401, 402, 427, 533, 534, 535, 536, 537, 538, 638, 639, 783, 803, 813, 821, 832, 841, 850, 922, 933, 945, 987, 1027, 1038, 1049, 1069, 1080, 1123, 1179, 1256, 1268, 1299, 1332, 1376, 1447, 1480, 1489, 1612, 1630, 1640, 1649, 1658, 1667, 1676, 1685]. Reviews [1695, 1705, 1714, 1724, 1732, 1742, 1751, 647, 650, 384, 385, 627, 474, 623, 365, 649, 626, 622, 218, 473, 477, 629, 388, 205, 203, 361, 363, 386, 478, 202, 628, 362, 215, 648, 387, 201, 625, 217, 206, 476, 624, 2088, 366, 367, 1464, 1471, 1504, 1568, 1577, 1585, 1594, 1603, 56]. revised [112]. Revising [1430]. revisions [2260]. Revisit [4257]. revisitation [2717]. Revisited [1381, 1691, 836, 577, 2333, 4430, 785]. Revisiting [4493, 3002]. Reviwes [227]. reward [4046]. rheumatoid [2876]. rheumatology [2103]. rhythm [3286]. rice [2939]. Richards [2130]. richness [767, 2184, 4086]. Ridge [729, 2845, 958, 4375, 3310, 3187, 3063, 3002, 4160, 2956, 563, 4310, 2114, 2590, 2596, 4683, 2558, 4315, 3469, 3478, 3271, 1189, 3757, 2863, 2339, 2531, 1498]. ridge-type [4315]. ridit [248]. Riemannian [2453]. right [2144, 2741, 4047, 2347, 3122, 3429, 1311, 20, 3873, 3863, 2372, 2916, 3143, 2178, 3972, 3454, 3218, 4634, 2570, 3055, 4106, 4105, 4108]. right-censored [2347, 3122, 3873, 3863, 2916, 3218, 4634, 4106, 4105]. right-censoring [1311]. right-tailed [4108]. rights [4360]. rigid [4302]. ring [749, 765, 737, 731]. ring-recovery [749, 737]. ring-return [731]. ringed [732, 1360]. ringing [757, 760, 1212, 1217, 1218]. rise [1924]. Risk [3751, 4287, 4731, 2194, 4732, 4716, 1686, 4691, 1346, 4549, 2307, 3455, 1894, 3355, 3660, 2224, 3294, 3456, 4611, 3835, 1126, 4023, 2942, 1897, 998, 4424, 2787, 2069, 2094, 2227, 3016, 4396, 3537, 3247, 1922, 4122, 4371, 3368, 3508, 4174, 2151, 4103, 3561, 2597, 4327, 2493, 3632, 2490, 3139, 2960, 4355, 4607, 4474, 2536, 4277, 3825, 2591, 4464, 3839, 3820, 4186, 3136, 4262, 3046, 4136, 3243, 1955, 4513, 1889, 4034, 4055, 2815, 2953]. risk-adjusted [3368]. risk-based [4371]. risks [3086, 4004, 4469, 1997, 3834, 4415, 4301, 2357, 3356, 4365, 3873, 506, 2758, 3295, 3071, 3631, 2917, 4297, 4673, 3819, 3037, 4313, 3893, 4299, 3928, 3683, 718, 3886, 412, 4196, 3791, 4452]. River [547, 792]. RL [1536]. RNA [4432]. RNA-seq [4432]. road [2548, 4582, 4258]. Robert [2277, 4721]. Roberts [781]. Robust [1259, 4220, 1030, 3121, 2873, 4268, 4656, 1535, 847, 1710, 1841, 2313, 3182, 2926, 3094, 3108, 4019, 3063, 1085, 3104, 3759, 396, 2566, 3135, 4660, 4742, 2590, 3878, 4654, 3151, 1418, 4552, 4314, 4067, 3967, 4667, 1670, 141, 3047, 3756, 696, 1431, 2648, 1620, 1532, 1632, 3044, 96, 2425, 2535, 1996, 2482, 2712, 1648, 3531, 2728, 4586, 3799, 3547, 3885, 4617, 4253, 2698, 4407, 4044, 1936, 1822, 2574, 1583, 1270, 3658, 546, 3517, 4155, 1523, 2101, 1756, 1888, 3606, 3051, 4411, 4036, 4160, 4498, 1867, 3358, 1892, 4400, 1818, 4045, 3775, 1334, 4646, 2894, 992, 2811, 2906, 1076, 1591, 1908, 4292, 2189]. robust [2747, 4235, 3932, 1513, 777, 3453, 2199, 3496, 4494, 3516, 2122, 4392, 4342].

Robustness

[1265, 979, 944, 990, 1277, 2252, 1117, 2134, 1386, 2108, 1293, 672, 540, 882, 1656, 689, 1894, 2048, 1512, 1093, 1369, 3443, 583, 1550]. ROC [4505, 2333, 4594, 2247, 1466, 4603, 2196, 3148, 1427, 3076, 4491, 1077]. rock [2136]. ROK [3076]. Role [771, 3356, 2613, 3954, 59, 4170, 2628]. roll [4360]. Rolling [1651, 1310, 2717, 3003]. rooftop [3725]. room [2360, 342]. Root [1416, 1428, 1309, 1756, 4166, 825, 1460, 1870, 2040, 2072, 1025, 1816, 1340, 3053]. roots [609, 845, 4165, 1246, 1318, 1335, 1497]. RORAC [3455]. Roseate [1227]. roses [46]. Ross [1234]. rossii} [1234]. rotatability [1888, 580, 714, 715]. rotatable [3094, 4276]. rotating [1106, 3944]. Rotation [691]. rotations [1075]. Rothamsted [159]. rounded [3647, 2167, 586]. Rounding [221, 1379, 225, 942, 1972, 2058, 484, 460, 496]. Rousseeuw [4669]. routine [1126, 4329]. row [498]. row-column [498]. Royal [162]. RR} [3874]. RR}-plot [3874]. RSM [2966]. rule [3311, 3883, 2727]. rule-based [3311]. rules [2023, 3894, 999, 667, 3537, 4610, 2726, 566, 4146]. ruminal [4341]. run [1545, 2420, 3894, 3821, 1141, 1905, 1042, 2808, 1720, 2141, 2267, 3003, 4146, 935, 3498]. Runs [2023, 1830, 2171, 1479, 2085, 1161, 3707]. rural [4370]. Ryder [3181].

S [2821, 2653, 334, 4278, 2646, 902]. SP [4178]. s-multinomial [334]. saddle [1151, 1048]. saddle-point [1048]. Saddlepoint [3086, 3974]. safely [4278]. safety [2181, 3344, 2721, 3625]. saga [85]. Sahel [755]. sales [303, 837]. salmon [3482]. same [763, 1107, 2451]. Sample [697, 633, 154, 3282, 1387, 2519, 1743, 733, 1389, 2469, 1380, 3617, 1661, 1390, 4084, 1431, 1403, 1066, 1627, 1759, 212, 3730, 1482, 4359, 3703, 4652, 1819, 1873, 1917, 4131, 3980, 1265, 1780, 2347, 615, 2083, 1551, 2110, 579, 1542, 1496, 1783, 2901, 3490, 1185, 2486, 2731, 4081, 3015, 1183, 3965, 1277, 797, 827, 918, 2752, 3118, 1563, 2297, 572, 700, 3644, 1720, 2419, 3539, 3096, 1118, 3909, 1104, 2024, 2446, 4017, 3749, 953, 2368, 1351, 1162, 1978, 2172, 1230, 4011, 3497, 3959, 375, 1459, 4476, 2907, 1939, 1982, 2764, 2283, 1866, 3352, 1032, 2928, 2362, 1579, 1656, 2638, 3019, 36]. sample [2062, 1044, 2565, 1648, 2728, 1474, 4449, 4446, 898, 2385, 4444, 2376, 866]. Sample-based [1661]. sample-size [1579]. Sampled [225, 2150]. sampler [2829, 1597, 4239, 2030]. samples [463, 3556, 2318, 3051, 2538, 4305, 1087, 1785, 4630, 4248, 1981, 745, 1344, 2350, 4583, 3748, 2178, 2550, 1275, 3351, 1511, 848, 1195, 3493, 4295, 3291, 2799, 2470, 4617, 899, 1308]. Sampling [1451, 780, 1408, 1396, 167, 145, 156, 1671, 1388, 472, 579, 2614, 2864, 1654, 161, 1627, 1680, 3907, 2380, 2804, 2383, 4100, 2100, 697, 1906, 1965, 2032, 1977, 3134, 3357, 871, 1017, 1029, 1059, 1545, 1558, 3671, 932, 553, 872, 1190, 631, 1250, 3665, 3865, 1192, 1280, 3784, 3274, 4052, 1967, 1273, 1776, 1084, 596, 1496, 1571, 4161, 1832, 4019, 4325, 3244, 3025, 3818, 1516, 2861, 1287, 1042, 3034, 1314, 506, 637, 885, 913, 910, 1086, 1094, 1757, 3390, 809, 4630, 2909, 2635, 4635, 816, 1969, 1041, 617, 681, 1121, 2931, 787, 3106, 1828]. sampling [2910, 254, 3661, 2009, 643, 2365, 4103, 2134, 3141, 1868, 3612, 1355, 1105, 1324, 2184, 3371, 2580, 1957, 982, 1079, 3830, 4263, 1543, 1365, 3951, 3058, 4560, 3946, 544, 556, 645, 675, 718, 709, 826, 956, 3851, 400, 4597, 1744, 3394, 1927, 1963, 333, 4095, 331, 412, 503, 576, 575, 557, 621, 839, 1090, 800, 819, 2006, 4094, 1580, 1339, 2213, 3238, 776, 905, 963, 1099, 1726, 4525, 861, 969, 1959, 1474, 2606, 2188, 3864, 2127, 3377, 3550, 4035]. San [884]. sand [2626, 755]. sandwich [1565]. São [2702]. SAR [4488]. sardines [1840]. SARIMAX [2647]. Sarmanov [2529]. SARS [4542]. SARS-Cov-2-infected [4542]. SAS [2225, 1881, 2890, 4706, 2325, 2509, 3544]. satellite [3949, 2959, 4659]. satisfaction [2245, 2903, 3168, 3311, 1971, 1919, 1943]. SATORDs [4276]. saturated [2874, 404, 577]. Saunders [4013, 3671, 4131, 3582, 3676, 1442, 3404, 4030, 2308, 2893, 2159, 3536, 3333, 4271, 4556, 2218, 2712]. SBMH [3721]. SCAD [3640]. Scalable [4218]. scalar [4656, 3732, 491]. scalar-on-function [4656]. scalar-on-surface [3732]. Scale [661, 4647, 463, 4273, 421, 1780, 3182, 2507, 3051, 2682, 3962, 3150, 4194, 3813, 3867, 1938, 235, 4610, 3539, 2087, 2598, 4351, 3646, 3966, 4361, 3321, 2959, 3482, 3959, 3790, 4216, 956, 2662, 2600, 1533, 1903, 4463, 4053, 2728, 2569, 4401, 2118, 4444, 1308]. scale-free [2728]. scale-invariant [4444]. scale-like [4610]. scale-mixture [4361]. Scale-space [661, 3321]. Scaled [213]. scales [2025, 2933, 661, 4136]. scaling [1142, 3978, 4281, 2708, 109]. Scan [1437, 3077, 3129]. Scandinavian [2099]. Scatter [1483]. scattering [3032]. scenario [3970]. scenario-based [3970]. scenarios [4546, 3127]. schedules [2691]. Scheduling [2622, 1148]. Scheffé [1522, 1562]. scheme [4426, 4100, 760, 871, 1017, 3865, 1905, 3429, 4590, 4451, 3695, 2724, 2080, 4351, 1543, 3227, 4677, 3946, 995, 1963, 2867, 4662, 4095, 677, 576, 800, 4463, 579]. schemes [2023, 553, 3894, 2864, 960, 4636, 2580, 4672, 4061, 252, 1927, 3990, 776]. Schindler [1541]. schizophrenia [3817]. Schoenberg [2652]. School [11, 526, 2, 3673]. schools [72, 591]. Schwarz [1216]. science [712, 2505, 2613, 2071, 4599]. Sciences [4694, 2754, 260, 3577, 2483, 4314, 2013]. Scientific [4748, 1645, 3572, 4715, 504, 3174, 712, 2985, 236, 3577]. scientists [38, 2502]. scleroderma [3004]. sclerosis [4241, 1374].

Sclove [2821]. Score [2700, 3214, 278, 4596, 3532, 3662, 4435, 2250, 3676, 3824, 2917, 2236, 2111, 3752, 2908, 3938, 3673, 1889, 3619]. scorecard [1070]. scores [2585, 2233, 2876, 3939, 1900, 1184, 2839]. scoring [397, 4176, 1070, 2597, 3312]. Scots [1383]. Scott [2653]. Scottish [158, 129, 786, 164]. scrambled [2073, 2794]. scraping [4450]. screened [2576, 3092]. Screening [1755, 2410, 1678, 2977, 2874, 1869, 4495, 1795, 4301, 1126, 3722, 4040, 1115, 3759, 2341, 3778, 2906, 3188, 3453, 2623, 3198, 4124, 4129]. Sea [3412, 4330, 3066, 2775, 3482, 1604, 3482]. seabird [1184]. seal [809]. sealant [3518]. Search [1031, 1989, 1438, 522, 657, 2294, 4160, 2960]. searches [3573]. Searching [2699]. season [4090]. Seasonal [3689, 1682, 833, 531, 4664, 2880, 2562, 3690, 825, 1318, 4422, 3749, 4524, 4073, 1025, 1497, 3726, 2259, 2905]. Seasonality [3634, 4727]. seasonally [48, 1574]. seasons [2106, 3690]. Second [1609, 2191, 1560, 2582, 1888, 2816, 2794, 2814, 590, 2192, 3373, 1161, 3840, 2824, 2617, 1753, 2914, 580, 3120, 405, 423, 2952, 3019, 2980, 2089]. second-generation [1753]. Second-order [1609, 2582, 1888, 590, 3373, 3840, 2914, 405, 423]. second-price [2980]. secondary [2721, 745]. section [97]. sectional [1892, 2844, 1983, 2206, 1874]. sectionally [1816]. sedentary [3264]. seed [4090]. seeking [4419]. Seemingly [3389, 3768, 3469, 727, 4153]. SEER [2511]. Segmental [2711]. Segmentation [2961, 2643, 3279, 1582, 3692, 4269, 1304, 1499, 249]. Segmented [3092, 4388, 3066, 2263, 1902, 4548]. segments [996]. Seiners [1604]. SEIRD [4433]. seismic [2967, 3412]. seizures [2584]. select [4246, 3358]. selected [2416, 1745, 3124]. Selecting [668, 1043, 3779, 4529, 1345]. Selection [1657, 735, 4698, 506, 565, 637, 1369, 617, 681, 2672, 643, 2837, 2536, 3538, 4623, 4634, 839, 646, 819, 2042, 3988, 2383, 2984, 3679, 900, 2034, 2732, 3917, 697, 4243, 2670, 4131, 1353, 3786, 1792, 3471, 4587, 4118, 3999, 3010, 3447, 3590, 652, 4458, 3329, 2149, 4007, 2166, 1084, 2249, 4434, 862, 3525, 2887, 2294, 2736, 302, 4620, 3104, 4369, 4514, 1779, 3500, 797, 3739, 3711, 3643, 4045, 4328, 4076, 3627, 1938, 1737, 4519, 3674, 4477, 2427, 567, 2391, 4408, 4440, 2871, 1991, 3342, 3901, 4381, 3967, 3881, 299, 1255, 1304, 3596, 4086, 3227, 4644, 4403, 1801, 544]. selection [556, 645, 709, 826, 345, 3680, 930, 3237, 2846, 1963, 2867, 4062, 3812, 4410, 1759, 4619, 466, 503, 576, 575, 4637, 2587, 4227, 2786, 3003, 3364, 3335, 3188, 4562, 776, 963, 1726, 2037, 2411, 2373, 2781, 3208, 3640, 4436, 3195, 2525, 4319, 2915, 3914, 4392, 3979, 4027, 2441, 3431, 2812, 3486, 3001, 3629, 2698, 2878, 2763, 4447, 4457, 3074, 3227]. selections [3618]. selective [1371]. selector [4485, 4294]. Self [3471, 2494, 4736, 2995, 4608, 4395, 2289, 3649, 3201, 2950]. self-assessment [4395]. self-modeling [3649]. Self-organised [4736, 2950]. self-report [2995]. self-reported [4608, 3201]. Self-selection [3471]. self-similarity [2289]. Self-updating [2494]. SEM [2966]. semantic [512]. Semi [502, 2353, 1707, 4432, 4233, 1901, 3468, 1962, 802, 3735, 4609, 1362, 2759, 3071, 3631, 3472, 3674, 2334, 2121, 2774, 3967, 3792, 4065, 2496, 2972, 3412, 3251, 4062, 907, 3953, 3822, 3852, 3800, 3897]. semi-analytical [4233]. semi-competing [3071, 3631]. semi-heavy [4065]. Semi-iterative [502]. semi-Markov [3735, 2334, 3412, 3852, 3800]. semi-metric [4062]. Semi-parametric [2353, 1707, 4432, 1901, 3468, 1962, 4609, 1362, 2759, 3472, 3674, 2774, 3967, 2496, 2972, 3251, 907, 3953, 3822, 3897]. semi-reflective [3792]. semi-sequential [2121]. semi-variograms [802]. semiconductor [3405]. semicontinuous [3767, 2483]. Seminar [92, 93]. Semiparametric [4611, 1276, 3798, 1301, 1909, 2010, 4352, 3143, 4361, 1821, 2143, 2467, 3642, 2946, 3132, 1257, 4385, 2262, 3694, 3859, 3334, 4370, 3843, 3813, 4082, 2679, 3186, 3770, 4456, 4627, 1275, 3995, 3604, 3853, 4219, 3815, 3713, 3914]. Semivarying [3607]. Sen [2747]. Senate [4588]. senescent [2690]. sensed [3949]. sensing [1372, 1778]. sensitive [3193, 3297, 2073, 2332, 2085, 3900]. sensitivities [2918]. Sensitivity [553, 2285, 1171, 3934, 2947, 3761, 4500, 4459, 2884, 3235, 2706, 3445, 4550, 1551, 1756, 3315, 3747, 2511, 2836, 2720, 2022, 2331]. sensor [4451]. Sensory [2379, 690]. sentiment [2851]. Seoul [4244]. Separability [2928, 3964]. separate [3255, 3200]. Separation [1226, 731, 2886, 4419, 605]. sepsis [4040]. September [1007]. seq [4432]. sequence [2015, 4630, 635, 3506, 1461]. Sequenced [146]. sequences [3077, 2112, 4340, 851, 1817, 4117]. sequencing [3353, 3068]. Sequential [968, 1409, 2599, 2981, 4276, 3119, 2720, 3400, 2129, 1725, 772, 2575, 1785, 3624, 796, 804, 266, 1989, 1351, 2121, 1497, 1734, 800, 1501, 2217, 4568, 2448]. Sequential-type [968]. serial [3634, 2171, 4674, 2635, 1772, 1998]. serially [4438, 2063, 4135]. Series [4710, 4726, 103, 4727, 2321, 121, 4713, 4721, 1454, 1880, 1402, 1439, 2191, 4722, 1448, 3383, 3988, 4072, 62, 3480, 1487, 2498, 2395, 2117, 2697, 2239, 4268, 3199, 1359, 1305, 3349, 3423, 609, 555, 1951, 2851, 1010, 938, 4645, 4168, 2562, 2739, 825, 3226, 1093, 3880, 1043, 1469, 3337, 2422, 2360, 1544, 418, 2226, 2277, 1318, 2910, 521, 3323, 1171, 2722, 4679, 3840, 4397, 4567, 2612, 4602, 4552, 2745, 1462, 2276, 2708, 3792, 1132, 1255, 4395, 2428, 1983, 4073, 2919, 527, 3249, 2601, 2510, 2473, 1850, 3976, 2306, 4382, 2555, 1005, 2152, 4138, 481, 4649, 3765, 1015, 1016, 1164, 3553, 699]. series [4180, 4425, 4548, 4500, 1933, 3432, 1570, 2653]. Serinus [1209, 1209]. seroconversion [884]. serological [2523, 410]. server [3480]. Service [188, 63, 1476, 3311, 1885, 4486, 3897, 3013]. services [2063, 339]. SERVQUAL [2993]. sessions [2622]. Set [1451, 1671, 1409, 4035, 4402, 2380, 2804, 499, 872, 1190, 1896, 4161, 3417, 2909, 4248, 787, 3141, 1355, 981, 4355, 4471, 1195, 1365, 1739, 1982, 3394, 3044, 3377, 3550, 2895, 1308]. Set-up [1409, 1739]. set-ups [1896]. sets [3213, 4404, 3226, 1914, 1576, 2368, 3227, 3691, 3464, 2450]. Setting [1693, 2245, 2440, 2518, 2022]. Settings [1413, 3173, 4460]. settlement [914]. several [4126, 4687, 3404, 2557, 796, 804, 3763, 1575, 844, 3268, 1903, 1501, 2450]. severe [2638]. severities [1602]. severity [4544, 2233, 1322, 3817]. sex [2636]. sex-related [2636]. sexual [3899]. SFILTEX [521]. shade [2842]. shadow [1201, 1202]. Shape [1455, 4302, 1377, 1407, 4469, 1965, 1333, 3328, 1009, 569, 657, 3314, 3520, 2489, 3871, 1909, 3661, 896, 1003, 1022, 4351, 1465, 658, 3424, 3599, 3850]. Shape-based [1455]. Shape-constrained [4302]. shape-invariant [1909]. shape-restricted [3314]. shape-scale [4351]. shaped [3400, 2805, 2515, 2220, 4213, 1804, 1538, 1313]. shapes [4421, 1465]. Shapiro [840]. share [3153, 3626, 3338]. Shared [2310, 3510, 4238, 2522, 4051]. shared-parameters [4051]. shares [996, 3128, 799]. sharing [3705]. Sharpe [2079, 4591]. shear [490, 250]. sheets [3833]. Sheffield [11, 92, 8]. Shelf [2899]. Shelf-life [2899]. Shewart [956]. Shewhart [1684, 2068]. Shewhart-type [2068]. shift [3738, 684, 1737, 2308, 462, 2840, 1485]. shifted [3423, 2014]. shifting [3392]. shifts [1536, 2085, 2291, 3073, 1870, 4463, 911, 1879, 1813, 1586]. shipping [3502]. Shiryayev [781]. Shiryayev-Roberts [781]. shocks [4384]. shooting [3984]. shopping [2076, 1181]. shops [440]. Short [3179, 4705, 2156, 3003, 3498, 1545, 3121, 3672, 4012, 3594, 4567, 4073, 1990, 2555, 2267, 3070, 935, 3068]. Short- [3003, 3594]. short-range [4567]. Short-run [3498, 2267, 935]. short-tailed [3121]. Short-term [3179, 4012, 4073, 1990, 2555]. short-time [3070]. shortage [3387]. Should [138, 132, 6]. shoulder [3568]. shovels [1784]. showing [2692]. shows [3936]. Shrinkage [4161, 3359, 4393, 2128, 3380, 1495, 4416, 1223]. shrinkaged [3904]. shuffled [891]. Sichel [2052]. Side [1098, 2085]. sided [3935, 1423, 2718, 3266, 1871, 842, 4686, 1548, 3718, 2309, 1403, 1759, 3822, 4146, 1802, 2694, 4405, 3463]. Sideritis [863]. SIDS [1574]. Siegel [3019, 170]. Sigma [1136, 4184, 3204]. sigmoid [220]. Sigmoids [137]. sign [1081, 2171, 4188, 579, 842, 4430, 1982, 3583]. sign-test [1081]. sign-type [842]. Signal [4712, 3706, 2414, 3568, 1765, 2811, 605, 955]. signal-to-noise [2414]. Signaling [2575, 4632]. signalized [4409]. signalling [880]. Signals [210, 2420, 2658, 2888, 3070]. signature [2624, 4088]. signatures [3868]. signed [4316, 3967, 4029, 1174]. signed-rank [4316, 3967]. Significance [4567, 4074, 110, 1678, 3282, 264, 459, 572, 1306, 2579, 263, 581, 460, 496, 2155, 2104]. significant [4458, 3779]. significantly [1012]. SILC [4204]. SILC-HBS [4204]. silico [3544]. SIMEX [2349]. similarities [485, 1799]. similarity [3239, 4128, 2289, 1100]. Simon [3417]. SIMPCA [3944]. Simple [1396, 2238, 3530, 1011, 163, 1377, 2087, 3826, 1620, 267, 529, 2810, 3616, 3655, 471, 1305, 1009, 2468, 391, 3244, 3985, 1185, 2458, 1757, 640, 787, 1095, 2024, 2368, 1343, 318, 2923, 1109, 4353, 1006, 1494, 689, 1111, 4349, 2188]. simplex [3442]. Simplicity [785, 1309]. simplified [573]. SIMULATE [1244]. Simulated [3507, 400, 558, 2253]. Simulating [37, 2613, 1013, 1703]. Simulation [163, 172, 1431, 4719, 2834, 2804, 3138, 343, 615, 380, 238, 2332, 3511, 4329, 2414, 3419, 1020, 827, 2221, 2613, 3241, 4343, 3657, 4315, 2554, 4020, 1801, 696, 2662, 2717, 2839, 3540, 4522, 1498, 2228]. Simulations [77, 264, 601, 3449, 3219, 4547]. Simultaneous [964, 539, 3465, 1737, 3529, 3037, 2121, 4116, 1051, 4107, 2564, 2828, 3506, 1501, 3496, 1891]. Simultaneously [1647, 4017, 3222]. sine [2520, 685]. single [4640, 3399, 871, 3835, 1276, 2736, 3408, 913, 4085, 4603, 607, 2973, 2291, 2871, 4333, 2584, 839, 1090, 963, 1726, 3882, 663, 2569, 2650, 4180, 3864, 3486, 3651]. single- [4603]. single-arm [4333]. single-featured [2291]. single-index [2736, 4085, 2973, 2871, 3486, 3651]. single-level [913]. single-nucleotide [2650]. single-photon [663]. Singly [1626]. Singular [1367, 3637, 2714, 3210, 2260, 4659, 727]. singularity [4005]. sinh [3075, 3466, 1937]. sinh-normal [3075, 3466]. sinh-normal/independent [3075, 3466]. sinusoidal [2811]. SIR [4539]. site [743, 4391, 2792]. site-dependent [4391]. sited [4642]. sites [792]. situation [1462]. situations [1760]. Six [1136, 4184, 3204]. Size [1625, 569, 1380, 1403, 1680, 3703, 3869, 4652, 1125, 697, 811, 633, 2545, 1388, 4303, 1210, 1542, 1783, 1228, 999, 4019, 2753, 3282, 4021, 2519, 1743, 2486, 2731, 733, 797, 1041, 3644, 801, 1720, 2266, 2469, 1118, 2080, 2207, 2386, 3617, 3749, 953, 1351, 4607, 1162, 1978, 2172, 403, 4084, 3352, 1066, 2928, 1759, 2362, 1579, 1114, 3149, 2062, 2565, 2199, 3730, 2376]. size-of-loss [4607]. SiZer [1439]. sizes [532, 1230, 4011, 2283, 3670, 2728, 1474, 3222, 2385]. Skagerrak [1604]. Skew [1669, 2329, 3913, 1421, 3012, 2885, 3654, 3901, 4144, 4505, 4191, 1843, 1811, 4151, 3523, 3962, 4484, 1968, 3678, 4194, 3813, 4030, 2990, 3152, 2396, 3484, 4031, 2446, 3646, 3505, 1971, 3790, 4216, 2963, 3100, 2218, 2704, 4091, 2535, 4027, 2118]. skew-Birnbaum [2218]. skew-elliptical [4151]. Skew-mixed [3654]. Skew-Normal [2329, 1669, 3913, 3012, 2885, 1843, 3523, 4484, 1968, 2990, 4031, 2446, 3646, 1971, 3790, 2963, 2704, 4091, 4027, 2118]. skew-t [4031]. Skewed [1842, 2665, 1873, 1937, 1185, 4363, 2735, 953, 4607, 4314, 1859, 3116]. skewed-longitudinal [4363]. Skewness [588, 4252, 3655, 1272, 3662, 3009, 2711, 2446, 3926, 4463, 1006, 2417]. skewnormal [1096]. skip [4100, 4263, 4095, 905, 963, 1099]. skip-lot [4263, 905, 963, 1099]. Skipped [1385]. SkSP [1099]. SkSP-3 [1099]. sky [3949]. slab [4369, 4319]. Slack [1522]. Slack-variable [1522]. slash [790, 3983, 4417]. sleep [635, 4608, 1197]. sleep-wake [4608]. sleeping [4257]. slice [3661, 4124]. Slope [2372, 1609, 714, 715, 1888, 590, 1095, 2040, 580, 1562, 1319, 3019]. slope-rotatability [1888, 580]. slopes [689]. Slovenian [2303]. Small [615, 1417, 579, 4214, 1185, 1387, 1436, 1633, 1983, 1648, 3902, 1482, 421, 3556, 417, 4119, 2637, 774, 1967, 2280, 652, 1321, 1496, 3051, 999, 3490, 3015, 4630, 235, 3652, 3749, 1275, 3351, 953, 3559, 375, 4589, 4295, 2907, 1282, 3670, 3858, 2303, 4617, 866]. small-area [2280]. Small-Sample [1387, 1185, 1648, 1482, 866]. small-scale [235]. Smirnov [1265, 1675, 636, 846, 1062, 2126, 1687, 2840]. smoking [2235, 3250, 2336, 2389]. Smooth [680, 1650, 2279, 2588, 4165, 1246, 3856, 4571]. Smoothed [3109, 2840]. smoothers [4219]. Smoothing [1275, 4641, 2042, 3805, 4516, 4167, 2759, 1020, 2944, 2010, 3245, 3827, 1766, 3444, 1067, 3261]. smoothness [3444]. snail [1241]. Snell [1610]. snow [1210, 3594, 741, 752]. SNP [2587]. Soccer [4121, 1911, 4047, 3775]. Social [3306, 119, 2322, 173, 38, 4670, 2966, 2445, 4437, 1801, 4618, 593, 3165, 4170]. socially [4510, 3597]. society [72]. socio [3517, 3586, 4270]. socio-demographic [3586]. socio-economic [3517, 4270]. sociology [3577]. Software [1953, 4697, 383, 2038, 1629, 512, 750, 3367, 520, 3544, 1805, 3060]. soil [559, 604, 3430, 4216]. soils [86]. solar [3833]. solid [3315]. solution [4233, 2018, 604, 3528, 2958, 2343, 558, 1111]. Solutions [4724, 4472, 2394, 2891, 1552]. solved [4249]. solvers [1893]. Solving [1242, 2406, 4357]. Somateria [1204]. Some [1443, 1553, 1822, 152, 155, 562, 117, 391, 165, 239, 4161, 4177, 207, 79, 35, 24, 58, 102, 119, 340, 2596, 346, 2317, 1864, 796, 804, 22, 2906, 701, 80, 197, 318, 141, 51, 3597, 557, 3781, 45, 148, 4525, 1648, 2354, 335, 2257, 485, 1217, 1523, 1970, 2335, 807, 1508, 1370, 931, 1277, 1218, 607, 3896, 1334, 1774, 1514, 1753, 1313, 448, 392, 1335, 2401, 1032, 1368, 540, 2836, 543, 3861, 3496, 394, 4117]. Something [34]. song [2920]. Sonority [1675]. Sons [4669]. Sound [4691]. source [2886, 4244]. Sources [975, 855, 3766, 3564, 2186, 1220, 3304, 4577, 4541]. South [3166, 3412, 3720, 1300, 923, 2844, 2581, 3579, 3906]. south-east [1300]. Southeast [3199]. Southern [3571]. Sovereign [1797, 3814]. sow [1514]. soybean [3639]. Space [1406, 2011, 1402, 2977, 811, 3706, 2284, 4543, 2489, 1143, 2317, 2767, 3933, 661, 3231, 3910, 3321, 2959, 3482, 2905, 4688, 3795]. space-time [3933, 4688]. spacecraft [4442]. spaced [4653]. spaces [3829]. Spacing [1615, 1289, 321]. spacings [4236]. Spain [3919, 1209]. Spanish [3008, 3346, 2182]. spanning [3017, 3887]. sparrows [768].

Sparse [1644, 2956, 3854, 4489, 3924, 2569, 2878, 3595, 4475, 3949, 3080, 4264, 3315, 4063, 4312, 3927, 4627, 3754, 4403, 4586, 3799, 4653, 4561, 3486, 4571]. Sparsely [3344]. sparseness [4678]. sparsifying [3944]. sparsity [4516]. Spatial [3805, 1681, 1682, 2180, 1792, 2409, 3716, 3196, 3449, 3622, 2517, 3524, 1295, 4685, 3088, 3439, 3332, 449, 3441, 631, 1948, 3232, 2603, 2224, 802, 2342, 3872, 3535, 1247, 1517, 4633, 3913, 2828, 4478, 2019, 3734, 2933, 2400, 3695, 3737, 379, 457, 4642, 3360, 4327, 4588, 2242, 419, 4299, 2617, 3910, 552, 4555, 3024, 3100, 3240, 857, 2600, 1740, 605, 1298, 2371, 1817, 3576, 549, 4374, 2137, 3483]. Spatial-temporal [1682, 2180, 3524, 3695, 3910]. Spatially [3559, 2304, 602, 4593, 3368, 1374, 791, 4051, 3776]. Spatio [2145, 2521, 4582, 2105, 4743, 2246, 3375, 2388, 2232, 3919, 4308, 4521, 4553, 3016, 4566, 3715, 2399, 4535, 3136, 4502, 4467, 2819]. Spatio-Temporal [4743, 2145, 2521, 4582, 2105, 2246, 3375, 2388, 2232, 3919, 4308, 4521, 4553, 3016, 4566, 3715, 2399, 4535, 3136, 4502, 4467, 2819]. Spatiotemporal [4524, 3949, 4316, 4510, 4053]. SPC [1956, 1137, 1140, 1138, 1692, 1402, 2267, 1485]. Spearman [4029]. Special [364, 3780, 1141, 4148, 4101, 4563, 3993, 4533, 3868, 4443, 120]. special-cause [3868]. Species [1681, 1673, 693, 767, 1247, 1571, 1517, 1184, 3487, 2434, 2184, 982, 1079, 4086, 1227, 3622]. specific [3234, 3574, 1209, 2942, 2212, 4545, 1243, 285, 744, 1699, 1976, 4322, 3807, 3677, 4298, 1798, 2847, 3046, 1227, 4229, 2918]. Specification [1018, 1423, 555, 2101, 596, 4339, 4192, 2933, 1111]. specifications [1842]. specificities [2918]. specificity [4550, 1205, 2836, 2022]. specified [302, 4113, 2944, 1501, 3496]. specifying [2913]. speckled [4606]. spectra [3766].

Spectral

[210, 4684, 1782, 4249, 1569, 3268, 3656]. spectral-based [3656]. spectroscopy [2964, 1336]. spectrum [3637, 2714, 1835, 3210, 2260]. Speech [1675, 3777]. speed [3179, 2880, 2002, 2180, 3797]. speeds [443]. spells [500]. spending [1330, 2628]. sphere [4625, 375]. spheres [371]. spherical [4516, 3115, 3424]. Spike [4319, 4369, 3981]. Spike-and-slab [4319]. spillover [2043]. spinal [1333, 1003]. spine [2493]. splice [1048]. spline [2082, 3137, 2010, 344, 2232]. spline-modeled [3137]. splines [680, 2379, 3245, 3827, 1827, 1766]. Split [546, 3613, 3379, 1134, 1596, 1896, 854, 1849]. Split-plot [546, 3379, 1134, 1596, 1896, 854, 1849]. spoilt [713]. sport [3984]. sporting [1456]. sports [3081, 1317, 867]. Spot [4492, 2797, 597, 4540]. Spotify [4217]. spots [4291, 4571]. spouses [377]. spraying [2835]. spread [3774, 2872, 300, 441, 4601, 4237, 4535, 4547, 3524]. spreads [1962]. spreadsheet [1141]. Springer [4599]. SPSS [2089, 4749, 4709, 1498, 3120]. Spuriosity [497]. spurious [1595, 3083, 1952, 3084, 3177]. Square [1248, 1395, 1394, 1482, 4387, 865, 1450, 2481, 3490, 1306, 1549, 3938, 2378, 2979, 909, 4675, 549, 2657, 3018]. squared [1092, 471, 1493, 3891, 1252, 1935, 3316, 875, 4208, 51, 526, 698]. Squares [1666, 1646, 1448, 1507, 2995, 404, 577, 979, 2797, 4198, 2249, 4509, 2481, 2018, 2421, 1176, 1590, 1162, 940, 305, 3131, 4242, 3607, 3444, 411, 321, 3584, 1498]. Sri [1840]. stabilisation [3502]. Stability [1609, 3914, 4037, 2294, 2899, 249]. stabilized [4435]. Stabilizing [3007, 2719]. Stable [1616, 2112, 3511, 1760, 2015, 1252, 2579, 4358, 1122]. stage [1507, 4376, 3665, 3010, 2358, 4325, 4598, 3259, 1062, 2598, 4573, 2356, 1547, 4242, 3758, 4353, 2047].

Staged [3227]. stages [3808]. staggered [725, 853, 940, 1071]. staging [742]. stamp [836]. Standard [2889, 4576, 1535, 541, 1542, 2162, 2374, 607, 2291, 3385, 3192, 1501, 2022]. Standardization [1659]. Standardized [3054, 2355, 4545, 764, 4283, 3826, 2303]. standards [84, 3172, 3361]. standing [788]. stands [817]. start [3573, 1781, 4173, 2016, 1831]. start-up [1781, 4173, 2016]. started [2986]. Starting [728]. State [1402, 3973, 4433, 3250, 3706, 3794, 2280, 1709, 1232, 3785, 3049, 1228, 3570, 2829, 2285, 1143, 3636, 506, 637, 3945, 4635, 3325, 1226, 2317, 2767, 743, 462, 3951, 751, 4025, 4560, 3720, 412, 503, 2905, 1623]. state-dependent [3945]. state-space [3706, 1143, 2317]. Statement [4139, 2229]. states [3327, 3621, 3762, 488, 4415, 4553, 303, 3621, 4629, 4539]. states* [4410]. STATICO [2399]. statics [1912]. stationarity [2033, 1886, 4260, 2648, 2785]. Stationary [1701, 471, 1305, 938, 2094, 1323, 4554, 2242, 1255, 2510, 3721, 858]. stations [3950, 970, 1745]. Statistic [1418, 1687, 1081, 3980, 4233, 865, 3077, 2481, 3739, 3413, 3711, 4511, 1171, 3258, 1315, 1747, 3951, 1541, 3058, 4056, 1189, 3195, 4016, 4613, 986]. Statistical [67, 3022, 364, 3480, 4152, 152, 4536, 4682, 3997, 155, 1631, 3027, 4696, 1781, 147, 57, 1659, 662, 3672, 4365, 807, 4251, 3777, 712, 171, 1660, 2288, 70, 264, 3871, 3594, 247, 1434, 664, 967, 1893, 2635, 4085, 1754, 143, 720, 2497, 725, 654, 4733, 69, 4322, 4612, 4706, 453, 2251, 1365, 2509, 2609, 443, 971, 2929, 4409, 356, 1075, 3955, 3554, 2987, 1110, 1048, 2088, 3055, 2229, 2540, 4747, 3387, 4605, 2760, 4399, 855, 4324, 1320, 1456, 1994, 900, 1553, 4648, 426, 4243, 445, 2384, 390, 4246, 2706, 3709, 1731, 2424, 4643, 652, 1102, 3821, 3785, 1142, 4049, 4037, 860]. statistical [2560, 4325, 4134, 4334, 3244, 3915, 2827, 3803, 530, 4148, 3719, 25, 3795, 3358, 1840, 24, 75, 880, 27, 849, 4506, 3842, 2408, 632, 3102, 2375, 2643, 919, 512, 4274, 4327, 22, 2029, 864, 1503, 1774, 820, 2889, 4607, 3544, 2798, 1884, 4326, 2948, 2891, 2567, 4535, 2081, 4618, 1491, 1152, 2016, 2283, 3352, 263, 3126, 4116, 319, 2426, 2425, 4533, 3628, 1197, 514, 517, 4639, 1146, 2650, 4425, 3201, 4443, 2951, 2818]. statistical-test-complemented [530]. Statistician [99, 705, 762, 59, 47]. Statisticians [38, 43, 149, 445, 1147, 2502]. Statistics [81, 3156, 188, 4, 2858, 17, 1135, 138, 4699, 4693, 4730, 4692, 4700, 4701, 4704, 72, 1387, 90, 4740, 1437, 181, 4705, 127, 163, 15, 4714, 162, 4742, 115, 160, 235, 2192, 4734, 3367, 222, 651, 1407, 1445, 209, 233, 2158, 2320, 4744, 1431, 2156, 2504, 4745, 77, 7, 108, 4708, 4709, 132, 1420, 3173, 5, 157, 2741, 4126, 1155, 2618, 335, 668, 968, 1729, 1008, 46, 39, 771, 774, 3153, 3728, 1019, 239, 1084, 1863, 2326, 2501, 4340, 31, 1149, 3292, 2890, 97, 1185, 2048, 511, 66, 972, 1088, 44, 607, 93, 3118, 340, 2746, 33]. statistics [3162, 1022, 2196, 4172, 2108, 830, 89, 76, 3097, 2770, 3026, 1882, 2821, 552, 18, 1459, 100, 2456, 95, 1055, 4599, 45, 3, 4003, 3990, 484, 460, 496, 2213, 74, 526, 2902, 3216, 11, 879, 2089, 1715, 2193, 2157, 2816, 2273, 2747, 2748, 1961, 2952]. stature [2251]. Status [1689, 3340, 3987, 3798, 4515, 1931, 3986, 4297, 4673, 3632, 4627, 2786, 3713, 4513]. stay [4503]. stayer [843]. stays [4026]. STD [252, 621]. Steady [2829, 2285, 4635]. Steady-state [2829, 2285, 4635]. Steck [4275]. steel [1827]. Stein [4122]. stellar [3248]. Stem [1383, 609, 1105]. STENO [522]. step [2665, 4376, 2161, 4178, 2398, 4064, 1945, 4416, 2189, 1375, 4349, 2582]. step-by-step [2161]. step-stress [4376, 4064]. steps [2881]. Stepwise [1324, 2983, 1259, 1079, 1076, 610]. stick [1793]. stigmatized [3758]. stigmatizing [3685]. Stochastic [3447, 4175, 4143, 1651, 422, 2404, 959, 3280, 2910, 2581, 281, 4512, 187, 1632, 3919, 4114, 4682, 343, 3526, 2649, 3329, 4483, 2294, 1508, 2991, 1170, 4179, 2004, 4683, 3197, 1514, 2455, 3428, 4601, 3422, 3901, 2179, 3276, 4486, 3346, 2499, 2478, 3308, 2703, 2119, 4399]. Stock [1252, 4171, 3427, 4169, 4287, 3252, 2035, 64, 1037, 2675, 2976, 3929, 2263, 870, 2020, 2215, 3840, 2727, 3342, 1814, 3721, 2201, 2680, 1890, 1586, 2382]. stock-based [2215]. stomach [1924]. stone [4568, 2785, 2091]. stop [4525]. stop-loss [4525]. Stopping [4333]. storage [1481]. storms [4520]. Story [57, 18]. Strain [1699]. Strain-specific [1699]. strata [4597, 3906]. strata-bootstrap [3906]. Strategies [1550, 3455, 2476, 1496, 3437, 1101, 2923, 2028, 3236]. Strategy [186, 1503, 4100, 3010, 4049, 1972, 4485, 2861, 4630, 2894, 4095]. stratified [1596, 1214, 3987, 3034, 3015, 2945, 787, 3106, 2009, 3141, 1511, 4474, 1283, 3839, 4597, 1215, 3610, 3409, 3377]. Streamflow [1444, 1607]. Streaming [4283, 4563]. streams [4413, 4569]. strength [4305, 3867, 4215, 4497, 4213, 4261]. Stress [1422, 4261, 4376, 3923, 4354, 3011, 4305, 3867, 4215, 3958, 4064, 881, 4497, 4616, 4213, 2832]. Stress-strength [4261, 4305, 3867, 4215, 4497, 4213]. strictly [3415]. strike [914, 2001]. strings [4326]. strip [1896, 3146]. strip-plot [1896]. stripping [1956]. strips [1827]. stroke [977, 2536]. strong [3199, 4057]. strongly [4077]. Structural [711, 2149, 2307, 2835, 1286, 2079, 845, 915, 3306, 88, 720, 1835, 2737, 1318, 1063, 2910, 3263, 916, 2375, 3204, 2040, 4073, 3597, 1865, 4265, 3628, 3393, 2980, 689, 1552, 2642, 4346, 1540, 1570]. Structurally [1345]. structure [3816, 4304, 3452, 3113, 1571, 1078, 3739, 1061, 2234, 4408, 3360, 3909, 2885, 2777, 3948, 1343, 2443, 4120, 4476, 3691, 4632, 4000, 356, 2636, 2514, 3078, 2791, 3019, 4053, 3669, 1468, 3202]. Structured [4264, 1979, 2294, 2364, 670, 1119]. Structures [1404, 1712, 856, 979, 1794, 661, 2796, 4017, 3428, 3802, 984, 1740, 2975]. student [3157, 4259, 3320, 3345, 3841, 3158, 3161, 3160, 4189, 4028, 1971, 3163, 2551, 3165, 3673, 2855, 2921, 1286, 1509, 3276, 2418]. Student-t [1286]. Studentized [1490, 1477]. Students [209, 2262, 27, 3162, 7, 2173]. Studies [4377, 153, 3641, 900, 2429, 2706, 3694, 3555, 1910, 1232, 2355, 3229, 2335, 2766, 759, 4075, 1755, 2787, 341, 585, 3114, 2457, 764, 1938, 3191, 2010, 3842, 3313, 762, 2899, 4468, 673, 3326, 4322, 3905, 4558, 753, 4573, 3209, 4585, 1231, 2362, 2012, 4491, 1262, 2482, 3492, 4522, 1831, 2047, 2918, 1204, 2946, 4453, 3619, 4129, 2527, 1572]. Study [1692, 1064, 1633, 1431, 1956, 1113, 2245, 2804, 4648, 1125, 4245, 1873, 2969, 929, 2262, 1948, 1525, 3138, 3942, 2995, 84, 2545, 3279, 3069, 615, 977, 2807, 3128, 3207, 793, 1596, 4143, 2011, 523, 2332, 1247, 3941, 2967, 1784, 3511, 2296, 2538, 732, 4329, 2532, 3518, 2414, 4184, 2130, 2048, 3330, 754, 1020, 3487, 71, 849, 724, 1835, 8, 2221, 4371, 636, 254, 4269, 635, 4203, 2613, 1374, 3241, 3770, 3513, 4281, 3632, 3657, 2897, 3162, 2661, 782, 924, 3807, 3502, 1717, 1574, 3910, 2554, 1233, 355, 1908, 4020, 1814, 345, 3159, 696, 4597, 2081, 4618, 1499, 1229, 2662]. study [4394, 1196, 3614, 907, 4136, 1943, 1234, 3093, 1227, 4332, 3721, 2839, 2638, 2785, 3919, 752, 2264, 3553, 4123, 3288, 2095, 4513, 3013, 3914, 4185, 3605, 1481, 1498, 2801, 3900, 4385, 3729, 3294, 4611, 3340, 4396, 2911]. study-specific [3807]. Studying [1288, 2256, 1891, 2242, 2251, 690]. style [1461, 1592]. styles [2070]. stylized [1309]. Stylometric [1711]. Sub [1639, 3383, 4511, 4103]. sub-cohort [4103]. sub-districts [4511]. Sub-pixel [1639]. sub-region [3383]. subdiagonal [591]. subdistribution [4299]. Subgroup [3692, 4504, 4526]. subgrouping [3960]. subject [2881, 4052, 1587, 2833, 849, 462, 1759, 2197, 4229, 2449]. subject-specific [4229]. Subjective [150, 2368, 4608]. Subjective-Probability [150]. Subjects [210, 4112]. submerged [2671]. subpopulations [2358]. Subroutine [149]. subsampling [4660]. subscription [2063]. subscription-based [2063]. Subsemble [2847]. subsequent [3102]. subset [2670, 3999, 302, 4012, 4514, 1779, 1473, 2847]. subset-specific [2847]. subsets [4159, 495]. subsetting [1723]. subspace [4454]. Substantial [3305]. substitute [3737]. substitution [2301]. subsystem [3405]. subtype [4063]. success [1235, 3841, 3160, 3204]. successful [1136, 1138, 2589]. Successive [393, 4384, 3898]. Sudden [1574]. sufficient [4267, 3733]. sugarcane [3702]. Suggestions [759, 1232]. suicide [3091]. suicides [3906]. suitable [1031, 3989, 2672]. sulfide [3322]. sulphate [604]. Sultan [1970]. Sum [1391, 2941, 1589, 2421, 3154, 1055, 2940, 3741, 4029]. Summary [1115, 3292, 4491]. Summed [1150]. Summer [1735]. SUN [3341]. Sunspot [1606]. Sunter [4349]. super [4462, 3824, 2188, 4459]. super-populations [2188]. super-quantile [3696]. superclassification [991]. superimposed [2811]. Superiority [1654]. supervised [1364, 4584, 2443, 2792, 3656]. supervised-weighted [2792]. supplement [1290]. supplemental [3659]. supply [4418, 2758]. Supplying [1651]. Support [1958, 2797, 3080, 3391, 3356, 3381, 3011, 4264, 3989, 2133, 4455, 3004, 4281, 4389, 4338, 2600, 3607, 2792]. support-vector [2133]. supportive [519]. suppression [3695]. Sure [3722, 3778, 2906]. Surface [1613, 1441, 1691, 1393, 2244, 3949, 1888, 4302, 3066, 2775, 590, 1061, 3568, 686, 1161, 1035, 443, 580, 2648, 1565, 1849, 3732].

Surfaces [1609, 208, 3382]. surgeon [966]. surrogate [4231, 1307]. surrounding [392]. Surveillance [781, 1436, 1731, 4413, 2086, 880, 1774, 551, 2075, 3091, 4621]. Survey [119, 69, 3166, 795, 161, 867, 212, 2684, 4549, 2713, 4460, 2346, 4411, 3025, 2296, 2656, 332, 1500, 2597, 4583, 2803, 1717, 3522, 4419, 3579, 4555, 2907, 1919, 331, 3409, 526, 2865, 3832, 2153, 925, 2907, 547]. Surveys [3740, 1627, 1364, 2995, 3685, 2752, 1563, 2297, 3106, 3603, 4623, 2523, 593, 1812, 4055, 3330]. Survival [4239, 2435, 2319, 1615, 1208, 766, 146, 3269, 739, 755, 775, 2761, 757, 1236, 2034, 3439, 3510, 1221, 763, 3344, 2098, 3194, 728, 734, 4655, 2649, 3301, 2912, 1209, 1210, 1937, 758, 4435, 732, 769, 1243, 1219, 3863, 3014, 2870, 3843, 3295, 4241, 3759, 4102, 744, 733, 564, 4363, 3831, 1033, 1226, 4511, 4479, 2917, 3374, 4572, 2282, 2151, 1220, 3484, 2511, 2738, 2850, 2943, 1230, 1205, 3559, 1222, 738, 747, 2220, 2584, 3396, 4482, 1004, 3731, 768, 907, 2204, 1234, 3796, 1227, 1358, 4671, 2047, 1204, 4124, 4557, 2253, 4621, 4628]. survival-rate [758]. surviving [3194]. Survivor [1453]. survivors [1997, 4043, 4241]. sustainable [786]. SVM [3866, 4450]. SVR [3825]. SVS [3754]. swarm [4159, 4162]. Sweden [1748]. Switching [1191, 707, 697, 1941, 3029, 910, 2263, 3945, 567, 2020, 32, 466, 1090, 646, 1183]. Switching-one-column [1191]. Switzerland [4599]. Symmetric [1598, 1416, 3769, 2779, 3121, 4418, 2487, 3639, 1185, 3280, 4408, 4358, 1122, 2188, 1308]. symmetrical [4060]. Symmetry [1395, 1445, 1415, 1482, 2144, 1023, 939, 2160, 2255, 2352, 591, 909, 2657, 3018, 3668]. symptoms [4543, 3817, 2513]. synchrotron [4221]. Syndrome [1574, 2597, 4531, 1733]. Syndromic [1436]. synergism [442]. Syntax [1382]. Synthetic [2227, 1633, 4716, 662, 2346, 3316, 2907, 1282, 2834]. System [1657, 144, 1704, 1642, 828, 522, 748, 697, 1917, 1317, 3349, 1725, 261, 799, 970, 3705, 565, 910, 511, 4503, 4335, 2886, 521, 2431, 3225, 2837, 2708, 1013, 1090, 589, 514, 3149, 394, 525, 2382, 4308]. Systematic [3330, 3942, 507, 1927]. systems [1320, 63, 707, 780, 1040, 1731, 4427, 3167, 524, 3311, 523, 1031, 247, 1323, 3528, 3389, 4240, 567, 2597, 743, 3249, 1109, 466, 867, 646, 1884]. systemwise [1722].

t [4031, 3567, 1286]. table [2307, 3844, 1865, 286]. Tables [1395, 1445, 1394, 1482, 2573, 3970, 528, 1092, 2715, 4211, 2827, 341, 530, 2348, 567, 3652, 3506, 1549, 2179, 2041, 544, 556, 645, 709, 826, 284, 3938, 3444, 1943, 575, 3670, 3781, 2378, 2979, 909, 776, 1484, 2657, 3018, 4044]. Tabulated [128]. tactics [1211]. tag [28, 1226]. tag-return [1226]. tagged [738]. tagging [1230]. Taguchi [1270, 3322, 4184, 1818, 777, 969]. TAIEX [2722]. Tail [4151, 3000, 2083, 4615, 1486, 4065, 4383, 3806, 1048, 3974, 4342]. Tail-weighted [3000]. tailed [1722, 2430, 2868, 3121, 1033, 3280, 3983, 1335, 1432, 1564, 4108, 3074]. tails [2179]. Taiwan [3906]. Tales [78, 268]. Tamil [3524]. Tang [2822]. tangency [4058]. tangent [2295]. Tango [2554]. tank [2886]. taper [2176]. TARCH [1806]. Target [1492, 4102, 2621, 1515]. targeted [3644]. targeting [4496]. tau [3071, 543]. taught [44]. taut [4326]. tawny [1243]. tax [4395]. taxes [3250]. taxonomic [3686]. Taylor [4066, 2213]. TCSP [1180]. TCSP-1 [1180]. Teaching [238, 163, 4714, 222, 209, 172, 2193, 77, 5, 2305, 19, 445, 3170, 79, 340, 10, 49, 359, 1055]. team [1985, 1137, 4304, 3735, 3241, 1911, 1072]. teams [3727]. technical [2807, 684, 1573, 2727, 3827]. Technique [211, 1607, 2280, 254, 1021, 4249, 2960, 3699, 551]. Techniques [1301, 4715, 1883, 2292, 3193, 1250, 2237, 934, 562, 4461, 1725, 3381, 1362, 3023, 1840, 21, 4371, 2408, 1361, 3686, 1960, 3126, 2717, 3003]. technological [3857]. technology [17, 4451, 3566, 659, 1197]. teenage [3561]. teeth [4051, 3686]. Tehran [2246]. Teissier [4225]. tele [3714]. tele-interpersonal [3714]. Temperature [1682, 1651, 3933, 1691, 3949, 3066, 3109, 3430, 2353, 3482, 1735, 2648, 1850, 2306, 4330]. temperatures [1362, 1152, 768, 287]. tempered [3511]. Temporal [4743, 2716, 2152, 2145, 2521, 1682, 2180, 1372, 4582, 4553, 1966, 4037, 1778, 3016, 3695, 2105, 3927, 4642, 4566, 3715, 2399, 3910, 2959, 2206, 4535, 2246, 3375, 3136, 2388, 1740, 2232, 4502, 1817, 4467, 3919, 3524, 2855, 4308, 4521, 2819]. temporary [745, 911]. tennis [981]. tensor [4642, 4585, 4291, 4571, 3216]. term [3179, 2235, 4043, 1219, 4012, 4241, 3594, 2490, 3146, 4073, 1990, 1850, 3731, 2555, 2791, 3019, 903]. terminal [3507, 4069, 3339, 3733]. terms [1722, 4323, 4292, 3538]. tern [1205]. Terns [1227]. territorial [1208]. territory [4384]. Terry [2135, 3499]. tessellations [1794]. Test [1443, 1507, 1406, 144, 914, 163, 1391, 1698, 1418, 1773, 1687, 184, 1390, 1205, 1686, 2579, 3249, 1416, 1628, 170, 3149, 1081, 3907, 3072, 1722, 528, 1092, 1291, 1125, 372, 1906, 449, 968, 588, 278, 3935, 599, 3039, 4550, 3767, 3345, 2021, 3369, 4655, 2347, 2654, 2941, 3203, 1551, 464, 1913, 2301, 579, 1581, 408, 3662, 2901, 999, 2481, 3844, 3318, 4594, 3803, 3578, 530, 1158, 79, 3891, 931, 4081, 3071, 797, 1078, 918, 1588, 1872, 2440, 4309, 1992, 2436, 395, 607, 2673, 1555, 3191, 851, 636, 846, 1062, 889, 2522, 2917, 4686, 2266, 2419, 3096, 3186, 2111, 3909, 4104, 2386]. test [3258, 1095, 1355, 4602, 3371, 4074, 2043, 3855, 2550, 2633, 304, 1989, 1315, 1023, 939, 4573, 4474, 2040, 1541, 1591, 3154, 2397, 3272, 3098, 3343, 1174, 1233, 1993, 2289, 4464, 1567, 1194, 4236, 840, 4008, 3385, 3817, 1982, 2160, 2231, 4242, 4394, 3192, 1066, 1109, 3099, 677, 2255, 4658, 1501, 1741, 2217, 2839, 4003, 484, 460, 496, 4094, 2638, 2939, 3833, 286, 1723, 2062, 2720, 710, 2271, 2902, 4016, 698, 1199, 2940, 3741, 3053, 4185, 4406, 4613, 4446, 2700, 899, 986, 1738, 4444, 2778, 3262, 2155, 4029, 253, 185, 1650, 1635]. test-retest [3262]. Testing [1081, 1625, 1748, 1487, 811, 929, 3239, 4530, 599, 501, 1235, 1929, 3039, 4427, 1292, 1595, 2186, 3552, 2033, 3117, 2367, 4674, 3875, 1851, 602, 2198, 4165, 1087, 1410, 2148, 1401, 2312, 1323, 1544, 2163, 4578, 2126, 1952, 4422, 4468, 4017, 1853, 2178, 924, 4293, 172, 2179, 3058, 3783, 1701, 1529, 1415, 747, 1431, 1916, 3817, 1632, 3610, 1428, 4265, 593, 3370, 4138, 1628, 3078, 614, 1622, 2046, 4247, 3668, 1933, 3383, 4473, 1128, 3935, 3665, 2171, 1932, 1009, 2965, 1309, 3030, 1718, 408, 2326, 2507, 3108, 3083, 3282, 469, 3172, 1785, 3793, 2200, 2057, 2552, 539, 2931, 4506]. testing [3084, 3144, 2207, 480, 3617, 4567, 2121, 4437, 61, 49, 3209, 2798, 3119, 4260, 844, 4476, 2648, 1497, 4056, 2058, 4616, 3898, 677, 3670, 1501, 3540, 399, 808, 2197, 4320, 1340, 526, 689, 298, 3583, 4449, 2936, 2700, 2444, 4029]. Tests [2144, 1637, 1413, 4487, 1665, 1699, 4092, 1672, 2211, 1445, 1390, 3183, 3971, 1439, 3839, 3641, 1422, 1678, 898, 1482, 3086, 4224, 2968, 3903, 445, 929, 1250, 2853, 1752, 2049, 393, 2243, 1901, 988, 1265, 1780, 1843, 3274, 1781, 4475, 3122, 3821, 1756, 238, 3923, 4354, 1783, 4435, 4166, 2161, 3987, 3415, 825, 330, 1093, 264, 3956, 3738, 2486, 1460, 1892, 2200, 1258, 1969, 1121, 725, 379, 457, 1864, 2116, 2431, 3285, 1172, 3414, 1306, 2661, 505, 2598, 3905, 1870, 2534, 583, 1351, 701, 1076, 4009, 2366, 2072, 4272, 1013, 4208, 881, 1335, 2183, 2030, 961, 1187, 1025, 2764, 4614]. tests [838, 2016, 1122, 4574, 1816, 2928, 110, 882, 1055, 1097, 2217, 866, 1580, 1337, 3968, 4275, 491, 965, 338, 1815, 3516, 4617, 2354, 2801, 3463, 4129, 3262, 2556]. tests-theory [393]. tetrahedralization [3036]. Texas [2652, 3737, 4737]. text [60]. Textile [199, 2408]. texture [666]. textures [2929]. TGCnA [3927]. th [4348]. Thailand [3088]. their [829, 652, 1598, 2560, 1517, 3470, 1508, 377, 3528, 2581, 2613, 3829, 252, 4383, 1403, 755, 3462]. theorem [1966]. Theoretical [374, 4675]. Theory [4728, 4701, 4724, 2158, 4736, 3698, 2903, 2994, 426, 2479, 1820, 106, 393, 970, 2288, 2758, 3862, 489, 2950, 2526, 661, 2534, 786, 2891, 1881, 2704, 1955, 4452, 2394]. therapeutic [442]. therapy [3049, 3664, 3050, 635]. there [3350, 1193, 3250, 4042, 947, 3284, 1648, 226]. Thermal [3322, 1642, 2835, 3219, 1481]. thermal-hydraulics [3219]. Thick [1432]. Thick-tailed [1432]. thickness [836]. thing [1107]. Things [2667]. thinning [4006]. Third [2898, 1899, 1531, 2817, 4276, 1988, 2823, 1921, 4070, 4406]. third-generation [1921]. Third-order [2898]. third-variable [4070, 4406]. Thirty [25]. thoracic [3441]. Thought [69]. Thoughts [5, 207, 35, 51]. Three [2096, 1404, 4584, 87, 2620, 2537, 1634, 1445, 208, 4252, 3460, 1830, 584, 763, 693, 3167, 3794, 948, 2432, 1840, 1768, 960, 2886, 1794, 1324, 782, 4464, 4120, 4476, 1943, 3087, 3670, 1337]. three-arm [4464]. Three-Dimensional [208, 2096, 1445, 2432, 1794, 782]. three-fold [1768]. Three-level [1404, 1634, 3460, 3087]. three-parameter [584]. Three-part [2620, 4252]. three-phase [1324]. three-sample [4476]. three-state [3794]. three-way [3670]. Threshold [1384, 3237, 1422, 4180, 4046, 3128, 3915, 4040, 2069, 2991, 4012, 4021, 1538, 4603, 4122, 3568, 3144, 3707, 2904, 2996]. thresholding [3388]. throat [261]. throughput [3457, 2650, 4570]. thyroid [3474]. Tianjin [4108]. tidal [904]. tiddlywinks [1317]. Tie [170]. Ties [3499, 3142, 2241, 4430, 2839, 4261]. Tightened [913, 1029, 1059, 565, 617, 1543, 1963, 576, 776]. tightened-normal-tightened [565, 1543, 1963, 576, 776]. tightening [697]. tilting [4459]. Time [4710, 4726, 103, 1977, 4727, 1999, 2321, 121, 4077, 1614, 4713, 4721, 116, 1664, 3848, 182, 2226, 4232, 2612, 1880, 164, 1402, 4395, 1439, 2191, 4722, 1448, 4336, 2119, 4539, 4111, 2975, 3383, 3988, 4046, 4072, 62, 2337, 3480, 2032, 2896, 3510, 811, 1406, 2498, 2117, 2697, 3997, 2239, 3138, 3786, 3355, 4268, 3199, 1359, 1305, 1531, 3423, 2998, 4331, 3794, 555, 2866, 2667, 1595, 3057, 1951, 1181, 1209, 3334, 2851, 1010, 2011, 2284, 938, 2165, 4645, 2432, 4321, 4168, 2294, 2562, 825, 2285, 3437, 3226, 3863, 3880, 3014, 1043, 1469, 3337, 2422, 3390, 3259, 2360, 1778, 3945]. time [601, 1261, 2635, 1544, 3327, 2277, 2400, 1318, 2910, 3508, 2604, 4673, 3838, 521, 2644, 1171, 2722, 3392, 3770, 3933, 4489, 4679, 3840, 3354, 4397, 4567, 3037, 4651, 4602, 3095, 4552, 3650, 2982, 3231, 2745, 4173, 3910, 2040, 3544, 1462, 2276, 2708, 3792, 1132, 1255, 2608, 2428, 1908, 2260, 1983, 1945, 2356, 4084, 2546, 2911, 4073, 2919, 527, 1307, 1973, 2601, 4482, 2510, 1850, 994, 3976, 3006, 2555, 1005, 2152, 869, 1075, 3796, 2208, 1834, 4138, 4463, 3953, 2653, 481, 4649, 3070, 3856, 4629, 3800, 3765, 1015, 1016, 1164, 2927, 3553, 3897, 4105, 4229, 699, 4688, 1831, 4123, 4054, 3091, 4405, 3669, 4180, 2022, 4390]. time [2137, 4034, 4425, 4500, 3132, 3432, 4110, 1570, 4628, 2395]. time-between-events-and-amplitude [4054]. time-course [2337, 4123]. time-dependent [4673, 3838, 1908, 4482, 3796, 4628]. time-dynamic [4489]. time-frequency [3006]. time-inhomogeneous [2927]. time-invariant [2285]. time-limited [2697]. Time-Series [2321, 121, 2226, 62, 3480, 2239, 2851, 3226, 2360, 521, 1983]. time-specific [1209]. time-to-event [4321, 3770, 4084, 1307, 3856, 3132]. time-to-funding [4173]. time-trends [1075]. time-truncated [2032]. time-use [2667]. Time-varying [1999, 4232, 4336, 2119, 2975, 3997, 3786, 3437, 2604, 2982, 2608, 2510, 1834, 3953, 2927, 2022]. Times [1651, 1615, 1320, 1795, 4039, 2113, 3301, 2912, 894, 3819, 3090, 3339, 4486, 2584, 3642, 883, 4557]. TIMET [1651]. timing [1210]. tissue [2223]. tit [763]. Titanium [1651]. Title [770, 822, 946]. Title-page [770, 822, 946]. TNT [776, 576]. TNT- [776, 576]. TNTVSS [1963]. tobacco [4648]. Tobago [3555]. Tobit [3229, 3341, 4418, 3582, 2148, 3569, 3534, 1932, 1410, 3770]. tobit-Birnbaum [3582]. Today [125]. toddlers [3481]. tolerable [3912]. Tolerance [1403, 1533, 3879, 695, 3895, 1111]. Tolerances [1618, 1702]. Tomographic [1455]. tomography [3847, 3636, 418, 663, 2625]. Tool [163, 372, 2873, 1138, 4159, 20, 661, 3874, 670, 359]. Tools [4715, 2889, 4120, 3393]. top [3297, 3934]. top- [3934]. topic [4450]. Topics [1862]. Topp [1508, 4497, 1802]. torsion [974]. Total [1388, 2018, 1571, 3051, 931, 2434, 4172, 982, 1079]. totally [1463]. totals [1511]. toughness [3497]. Tour [4728, 1251]. tourism [3513]. Tourist [3210, 4042]. tournament [1272]. tournaments [908]. tower [4231]. toxicology [3912]. trace [2733, 3254, 4680]. tracing [4472]. track [304]. tracking [4648, 892, 4510]. tractor [3566]. trade [3061, 959, 4679, 752]. trade-offs [752]. trading [2727, 1890]. traditional [870, 2594]. traffic [2548, 4582, 4012, 2203, 4258, 699, 4641]. train [2538, 3392]. trained [657]. Training [1707, 1353, 1357, 27, 1992, 2385]. trains [4057]. Trait [1644, 4549, 3334, 3169, 3300, 4468, 2704, 2587, 2482, 4229, 3202, 2448]. traits [3108, 1228, 4663, 2850, 3516]. trajectories [3374]. trajectory [1211, 3313, 3860, 2455, 2862, 1990, 2955, 4308]. transaction [1983, 3721, 1586]. transcranial [2536]. Transcription [2792]. transect [1188]. transfer [3455, 601]. transferring [4286]. transform [961, 3070, 2743, 4521, 1336]. transform-based [4521]. Transformation [1418, 1045, 470, 798, 4195, 2492, 702, 1043, 973, 801, 1129, 2722, 2433, 2916, 1095, 1104, 4627, 2446, 3855, 461, 3995, 2666, 685, 2143, 1196, 4141, 4106, 1044, 2525, 3713, 1112, 1077, 1253, 1257]. Transformation- [1045]. transformations [3728, 1321, 2488, 1054, 3787, 283]. transformed [1039, 4139, 1010, 4168, 3705, 4140, 3245, 4145]. transformer [3180]. Transforming [1589]. Transient [741, 4564]. transiting [3040]. transition [4367, 3215, 2239, 3785, 2153, 744, 2190, 3509, 3272, 3726, 3973, 2992, 3410]. transition-specific [744]. transitional [2656]. transitions [3164, 1246, 4014]. transitive [3724]. transmission [4551]. transmissions [2589]. transmitted [2113]. transmuted [4431]. transplant [3184, 2912, 4424, 2372]. transplant-free [2912]. transplantation [3986, 3139, 3796]. Transport [154, 3797, 254, 33, 3013]. transportation [1140]. trapping [3441]. Traven [1729]. trawl [2865]. Treating [2028]. Treatment [1443, 123, 1081, 4681, 4655, 4339, 3795, 4400, 4230, 2364, 3624, 344, 4477, 4646, 3752, 2784, 4634, 2402, 1741, 2315, 2638, 1831, 4406]. Treatments [1443, 1626, 3584, 3719, 2607, 842, 796, 804, 3752, 3209, 1194, 4380, 4634]. Tree [2364, 2176, 1792, 1247, 4435, 3557, 3768, 1601, 3017, 3691, 4626, 2373]. Tree-structured [2364]. TreeKDE [4626]. trees [3085, 794, 515, 1792, 656, 788, 4266, 2427, 3984, 3887, 4521]. Trend [1665, 1342, 1691, 1143, 2808, 379, 457, 3151, 3502, 1541, 2397, 1876, 4406]. trend-free [2808]. trend-projecting [1143]. Trends [1417, 1691, 2336, 2998, 1219, 3066, 2445, 3774, 869, 1075, 2232, 4618]. Trial [175, 1615, 125, 146, 3795, 3644, 3434, 2084, 3056, 4379, 976, 4112, 2264, 4522]. Triallel [1248]. Trials [1398, 3107, 1975, 4655, 2129, 3217, 3224, 2739, 1115, 1082, 3259, 2457, 1872, 2314, 4230, 3624, 1034, 1118, 2598, 3807, 3677, 4084, 4333, 1297, 4099, 1740, 1759, 3087, 3589, 2402, 2209, 4658, 655, 3133, 1600, 1281]. triangular [3513]. triangulation [4582]. tribal [3110]. Trimmed [1646, 1418, 3751, 1095, 1104, 1044]. trimming [4172]. trimodal [4661]. Trinidad [3555]. trios [3981]. triple [4647, 4456, 589]. triple-negative [4456]. trips [2076]. tristeza [315, 441]. Trojan [1248]. tropical [4520, 678]. trough [1762]. trucks [1956]. True [57, 2853, 4388, 4659, 2693]. truncated [3907, 1906, 2032, 1977, 3122, 482, 3071, 3947, 3371, 3650, 266, 1989, 4488, 4011, 3545, 2270, 2663, 2978, 3218, 4094, 1580, 4258, 4106, 4628]. truncation [1997, 2052]. Truthful [1414]. TS [1802]. TS-GTL [1802]. Tsallis [3907]. Tu [2822]. tuberculosis [3664, 3614]. Tucker [2653, 3506]. Tukey [170]. tumor [3719]. tumors [3315]. tumour [843, 4063, 344]. Tuning [4086, 4059, 3104]. Tunisia [1948]. turbines [4231]. turbulent [3814]. Turkey [4538, 3517, 3069, 2789, 3204, 4330, 4258, 4170]. Turkish [4171, 2789, 4162]. turning [1553, 4043, 3423, 2575, 874]. turnover [1985, 764, 3432]. TV [1664]. twin [1258, 2012]. twittering [483]. Two [2881, 3616, 1381, 2868, 175, 1392, 2358, 1578, 4228, 495, 75, 3096, 2205, 2398, 1418, 1858, 1390, 1431, 3976, 1647, 4141, 1802, 965, 4444, 1077, 3400, 856, 2805, 1507, 3503, 2733, 2430, 1487, 584, 3751, 1873, 1029, 1059, 1917, 932, 4376, 4273, 3239, 3665, 599, 3254, 1265, 2347, 3301, 2965, 2659, 3841, 2715, 3889, 2957, 2725, 3063, 341, 3875, 4598, 1516, 530, 799, 772, 4081, 3259, 3793, 2200, 2607, 1871, 964, 2147, 2635, 827, 2440, 4178, 617, 2485, 846, 1062, 2735, 4297, 4673, 3879, 2126, 4610, 1519, 2151, 1012, 2773, 3909, 2803, 1104, 280, 2024, 505, 2598, 4558, 2483, 4263]. two [3910, 2040, 4556, 1735, 3718, 2356, 1194, 669, 3148, 4416, 3946, 4380, 2707, 1459, 4559, 4295, 265, 3067, 2058, 2784, 2764, 695, 984, 2867, 3291, 3476, 2189, 4242, 3610, 4662, 3352, 3444, 1075, 3758, 4353, 581, 1579, 1834, 4138, 1006, 3822, 2638, 3188, 1817, 2694, 3149, 2712, 2588, 3808, 2047, 2199, 2533, 689, 1051, 2728, 1804, 2940, 4349, 1484, 4689, 4446, 2278, 2448, 3262, 2104, 2582, 1389]. two- [584]. two-armed [1487]. two-by-two [3793]. two-component [2725]. two-course [1075]. two-dimensional [3444]. two-group [2440]. two-level [856, 2733, 1029, 1059, 3254, 772, 964, 617, 2803, 4263, 984, 3188]. two-parameter [2805, 932, 2659, 3889, 1516, 846, 3879, 3291, 3476, 2712, 1804]. two-part [2483, 4559]. two-phase [2147, 1834]. two-piece [4273, 695]. two-plan [1917, 2867]. Two-Sample [1390, 1431, 4444, 1265, 2347, 4081, 3909, 1104, 1459, 2638, 2728, 4446]. Two-sided [1802, 1871, 3718, 3822, 2694]. Two-stage [2358, 1507, 4376, 3665, 4598, 3259, 1062, 2598, 2356, 3758, 4353, 2047]. Two-step [2398, 4178, 4416, 2189, 4349, 2582]. Two-tailed [2868, 2430]. two-vote [799]. Two-way [1578, 1418, 599, 2715, 341, 3875, 530, 505, 4556, 4295, 265, 2058, 1051, 1484, 4689]. Type [3503, 4581, 1391, 2770, 4662, 4426, 4359, 4004, 3816, 4013, 3837, 968, 3784, 988, 2021, 1807, 3057, 3541, 4354, 4365, 4590, 2481, 2068, 3965, 1981, 842, 851, 2672, 4035, 2644, 3527, 3770, 2454, 3947, 4315, 4322, 4612, 2251, 3493, 4672, 2630, 2710, 3497, 3699, 3916, 3820, 2515, 2924, 4162, 4255, 1963, 576, 2799, 909, 1539, 4350, 1099, 3856, 3553, 3757, 965, 4319, 2805, 463, 4097, 4305, 3247, 1828, 3418, 3397, 1698, 1344, 3748, 4236, 3178, 3291, 1866, 1110, 1339, 1302, 1198]. Type-1 [3247]. type-2 [1981]. type-I [4426, 3784, 2672, 2454, 2630, 4097, 1828]. Type-II [4581, 4662, 4359, 4004, 2021, 4354, 4612, 4672, 3916, 2515, 2799, 2805, 4305, 3418, 1344, 3748, 4236, 3291]. types [3369, 3841, 3690, 3507, 4297, 4673, 3093]. typing [410]. typologies [2701]. typology [226].

U [3773, 1538]. U-shaped [1538].

U-uniform [3773]. U.S [4552]. U.S. [2435]. Uganda [346]. UK [4169, 1700, 144, 1472, 701, 4395]. ULM [750]. ultra [4301, 3778, 3198]. ultra-high [4301, 3778, 3198]. Ultrahigh [4267, 4124]. Ultrahigh-dimensional [4267, 4124]. ultrasound [2536]. Ultrastructural [4473]. umbrella [1355]. UMP [1482]. unaided [591]. unascertained [4545]. Unavailability [1386, 944, 990]. Unbalanced [3004, 253, 2879, 2476, 2780]. Unbiased [3171, 1303, 1482, 872, 4122, 400, 2138, 1308]. Uncertain [4720, 2274]. uncertainties [2613]. Uncertainty [150, 3299, 4720, 1912, 2107, 927, 1512, 3015, 2593, 3652, 2617, 3124, 4014, 3352, 3219, 3789, 4307]. unconditional [4211, 3844, 1082, 4195, 1097]. unconfoundedness [2884]. under-dependence [1571]. under-five [4511]. ‘Under-reporting [1167]. underdispersed [2589, 1821, 2938]. underdispersion [3708, 3411, 4116, 2692]. undergoing [3050]. undergraduate [849]. underground [789]. Underlying [1637, 2247, 3221, 971]. undernourishment [3126]. underreported [1103]. Underreporting [983]. Understand [4720].

Understanding [188, 2387, 4189, 4747, 420, 2530, 4488, 2948]. undetectable [1295]. undirected [3724]. undocumented [2841]. Unemployment [1707, 1311]. Unequal [103, 1396, 1105, 4681, 2538, 990, 2207, 358, 800]. Unexpected [4723, 2651, 1485]. Unified [3194, 1434, 4590, 4081, 3497, 3886, 1501, 4003, 2095]. unifier [26]. uniform [2732, 2025, 3298, 3773, 4035, 2105, 3542, 1122, 2839, 985]. uniformity [2466, 838, 710]. unify [106]. unilateral [4380]. unimodal [3147, 3635, 930, 3996]. union [2198, 2994]. union-intersection [2198]. unique [4598]. Unit [4156, 845, 1246, 1870, 1416, 1428, 1836, 4119, 1309, 1756, 2033, 4166, 4214, 825, 4625, 4165, 3956, 4200, 1460, 1318, 4466, 3743, 3921, 3922, 3920, 3727, 2040, 2072, 1335, 1025, 1497, 1816, 4619, 4675, 1340, 491, 688, 3053]. unit- [4119]. unit-level [4214]. Unit-Lindley [4156, 3743]. unit-Weibull [3922]. United [4415, 4553, 3930, 303, 3621, 4629, 4539, 4360]. units [1781, 2016]. unity [1551]. Univariate [4749, 4326, 929, 4268, 2083, 2592, 2289, 3120, 859, 773]. Universal [1708, 1141]. universe [268]. universitie [3450]. university [2305, 2262, 3345, 3158, 3160, 849, 3162, 1971, 3159, 2551, 3171, 157, 1931, 3163]. unknown [828, 811, 620, 4550, 2021, 4339, 2944, 2163, 3398, 894, 2234, 2308, 2040, 779, 838, 4062, 3335, 4117, 4250, 3442]. unless [3807]. Unmasking [899]. unmeasured [4459]. unmixing [1778]. unobserved [3418, 4520, 4357, 4056, 2028]. unpaired [4278]. unrelated [2212, 3389, 3768, 3469, 727, 4153]. unreplicated [1736, 2654]. Unseen [1673]. unstable [1760]. unstoppable [1467]. Unsupervised [3890, 4450, 2513]. Untangle [3628]. untargeted [3644]. unusual [4511]. unweighted [1279]. up-down [1074]. update [4636]. Updating [3503, 2494, 2341]. upgrading [3360]. upland [794, 786]. upon [471, 1781, 611, 3878, 2016]. upper [1842, 2419, 3806, 3476, 898, 899]. ups [1896]. upside [3520]. upside-down [3520]. upward [1536]. Urban [3013, 3555, 4582, 4370]. Urdu [234]. USA [2099, 3250, 4535, 4618]. usability [3258]. usage [2960, 2063, 1373]. Use [997, 468, 154, 150, 195, 3738, 182, 162, 102, 222, 3421, 1734, 1234, 903, 2103, 1354, 19, 421, 1372, 2667, 2003, 1321, 2602, 523, 1783, 1241, 2967, 860, 2739, 4184, 1159, 3848, 880, 745, 512, 2215, 3481, 33, 2534, 751, 461, 1206, 423, 685, 2907, 4202, 4275, 1207]. used [977, 1142, 3621, 2803, 2962, 982, 1056, 1148, 768, 2078, 2836]. Useful [188, 163, 775, 769, 3750, 3031]. user [519]. users [903, 610]. Uses [760, 122]. Using [1444, 499, 1124, 3663, 2424, 865, 211, 1414, 2582, 3049, 1084, 1645, 181, 1020, 1184, 1053, 1395, 1860, 4735, 419, 3626, 3586, 2325, 1689, 1024, 2826, 375, 551, 1150, 1055, 1477, 4709, 1683, 1425, 2089, 1039, 1451, 957, 2304, 3213, 4710, 996, 2380, 3805, 3193, 2804, 1330, 1748, 1478, 4681, 3698, 2337, 1830, 2082, 1278, 3134, 3357, 2921, 1681, 1693, 4549, 2896, 4687, 968, 4376, 1238, 1190, 4387, 2664, 929, 2423, 2387, 3355, 2995, 2920, 711, 4233, 1792, 4433, 1280, 3784, 3517, 3660, 2484, 1169, 1242, 3234, 4575, 3526, 4655, 1060, 1962, 1413, 2772, 2900, 3294, 3456, 4544, 4458]. using [2912, 3182, 3203, 3835, 442, 482, 3322, 1126, 657, 4316, 2713, 1141, 2942, 1675, 4007, 1944, 2222, 2675, 3174, 2249, 3662, 1863, 3804, 1228, 3570, 1571, 1582, 3639, 4321, 2464, 4168, 2887, 2922, 2562, 4207, 2018, 2890, 2689, 1219, 2561, 2198, 2656, 3678, 3437, 578, 4369, 1362, 4165, 1639, 3705, 1512, 2690, 3050, 3014, 2758, 4189, 931, 3337, 1127, 1264, 3795, 3390, 2406, 1914, 2200, 2488, 744, 1239, 4113, 733, 3315, 604, 2566, 3567, 2312, 1251, 1538, 972, 3152, 2348, 4635, 3514, 3537, 2944, 1875, 4421, 4462, 721, 4328, 3824, 1518, 3771, 2226, 4063, 3857, 1737, 2621, 344, 2408]. using [3981, 3774, 4477, 2581, 2375, 3204, 382, 456, 762, 1054, 1864, 2214, 2722, 3004, 2020, 2830, 3927, 3467, 2597, 3129, 3936, 817, 4583, 2493, 2087, 4204, 2029, 3206, 3245, 3430, 3371, 4552, 2405, 3855, 3958, 2465, 4051, 1569, 1957, 1465, 3043, 3361, 1803, 3424, 4607, 2837, 2410, 874, 1747, 1827, 3948, 1326, 2509, 4210, 3544, 4067, 3881, 1237, 1462, 3097, 2630, 3272, 299, 1163, 1766, 1528, 1255, 4244, 3542, 3253, 3699, 4128, 4667, 4254, 2972, 765, 1983, 4084, 2559, 1222, 731, 1567, 4512, 1814, 4338, 4589, 930, 961, 4559, 3702, 3248, 1881, 1432, 4186, 4497, 1939, 1982, 3394, 3136]. using [3312, 4626, 1004, 994, 1960, 2189, 3044, 2841, 3126, 3219, 2027, 4136, 356, 3087, 3278, 4402, 3589, 3906, 2539, 621, 4330, 2583, 2523, 1151, 3464, 2551, 3076, 4138, 3721, 2703, 2717, 893, 2514, 3060, 3176, 2693, 2232, 2517, 4146, 4649, 3296, 2535, 4467, 1475, 1521, 3409, 3815, 3897, 4229, 2047, 698, 2996, 4261, 4349, 4605, 1281, 1349, 3864, 4027, 1889, 2173, 2303, 986, 4453, 4521, 1308, 3216, 1955, 2820]. uterine [4043]. utilise [392]. utilities [1354, 4317]. utility [1084, 2559]. utilization [3859, 4633, 235, 2628]. utilize [4509]. Utilizing [1421, 3311]. Uttar [2280].

V [2819, 681, 826, 3993]. Vaccine [1699, 999, 4379]. vacuum [721]. vacuum-formed [721]. Valen [2902]. Valencian [3919]. Valid [3521]. Validating [2514, 2316, 4506]. Validation [3891, 4371, 1307, 2995, 4049, 4137, 1719, 1878, 3776, 3995, 4133]. Validity [1637, 1092, 733]. Valuation [4716, 1452, 2813, 2227]. Valuations [3621, 3415]. Value [3825, 775, 372, 1820, 4415, 3355, 3780, 2637, 4287, 3835, 3450, 3174, 1367, 37, 2758, 502, 918, 1754, 1922, 3536, 3206, 3954, 1366, 3422, 2972, 1157, 4020, 4659, 4008, 2793, 2306, 1955, 4452, 4154, 4150, 1308]. Value-at-risk [3825, 3355, 4287, 3835, 1922]. valued [4438, 4645, 4223, 4455, 3101, 2055, 3321, 3964, 4006, 2692, 4685, 4346, 4425]. Values [1687, 3086, 588, 1923, 3350, 728, 4461, 3717, 825, 42, 3306, 565, 2293, 3775, 2126, 3418, 567, 3975, 2396, 2989, 4074, 1176, 2534, 3422, 1567, 3385, 3476, 1065, 3403, 3955, 411, 2693, 399, 3744, 2980, 2996, 4247]. Vancouver [3767]. VAR [1760, 1935, 1109, 1955]. Variability [1450, 4005, 1655, 4523, 3616, 2023, 1084, 494, 2489, 1765, 1850, 1499, 2231, 1327, 4190]. Variable [2383, 3786, 4587, 4118, 3010, 1388, 3590, 3525, 2736, 302, 2427, 1654, 4496, 3881, 4644, 2846, 2703, 2786, 2781, 3208, 2525, 3979, 4027, 2812, 2763, 3693, 3679, 2034, 2732, 4495, 1819, 1941, 979, 3007, 3447, 4458, 3329, 2582, 1186, 2166, 3402, 4434, 3244, 2294, 1287, 4369, 644, 1347, 4635, 4045, 1041, 787, 1522, 1737, 3674, 2134, 3129, 4440, 2871, 4381, 3967, 3123, 461, 4666, 4403, 956, 3851, 3293, 2177, 2283, 4099, 1874, 4619, 4071, 2006, 2939, 3364, 3335, 3188, 4562, 3230, 969, 3453, 4429, 1474, 4319, 3914, 4392, 3864, 2441, 4070, 4406, 4250, 1593, 3486, 3001, 3198, 3629, 3900, 2571, 2698, 2878]. variable-based [2571]. variable-lot-size [1041]. Variables [145, 156, 4052, 1483, 1646, 1407, 1393, 1636, 1697, 1291, 2988, 1558, 1917, 3270, 3205, 3665, 3865, 3685, 2869, 4198, 2101, 1126, 1829, 2222, 3941, 2620, 4598, 1516, 1286, 3989, 885, 2771, 3255, 2407, 1833, 3362, 2004, 1969, 1063, 4092, 3593, 1361, 2087, 2256, 3612, 3145, 1324, 3586, 1543, 3101, 4011, 3111, 3273, 1473, 1206, 975, 685, 359, 1764, 3312, 3200, 1963, 2867, 1065, 2451, 1943, 389, 692, 839, 1090, 2315, 1207, 3531, 321, 1552, 2449, 2569, 3475, 3584, 4398, 2170]. Variance [1682, 1396, 145, 156, 2790, 4694, 168, 196, 2065, 4133, 2103, 1128, 847, 2367, 4284, 4166, 2298, 4674, 1578, 83, 3002, 3473, 3738, 2731, 296, 378, 4030, 396, 640, 2436, 2163, 2008, 851, 853, 873, 2634, 2184, 2661, 1773, 1870, 3502, 1056, 1071, 3411, 297, 318, 696, 458, 265, 3271, 3352, 3278, 2013, 3370, 550, 1513, 2719, 1494, 4386, 2525, 3741, 4154, 3475, 2385, 2848, 2556]. variance-ratio [2436]. variances [4681, 542, 2965, 1170, 1852, 940, 2707, 844, 3898, 965, 3583, 3741]. variant [3190, 3068, 4342]. Variate [117, 450, 1791]. Variates [1608]. Variation [3468, 2804, 4652, 4220, 1487, 532, 734, 3894, 1829, 1219, 3373, 1053, 1220, 1858, 3749, 2132, 2539, 4003, 4146, 3524, 3861, 3868, 3262, 3693]. Variational [4266]. Variations [182, 2224, 849, 2783, 287]. varietal [2254, 1163]. variety [2739]. variograms [802]. various [3050, 2933]. VARMA [3337]. Varying [1778, 1449, 3651, 3997, 3786, 532, 1999, 4077, 4316, 3030, 2736, 4207, 3437, 3567, 2187, 723, 959, 2973, 2604, 2433, 4232, 4489, 3130, 3461, 3601, 3090, 4552, 2982, 1574, 2608, 2510, 4336, 1834, 3953, 2119, 3462, 2854, 2927, 3560, 2046, 4586, 2022, 4111, 3198, 2571, 2698, 2763, 2975, 4447, 3378]. Varying-coefficient [3651, 2187, 2973, 2433, 3130, 3461, 3560, 2571]. Varying-time [1778]. VCs [4173]. Vector [656, 923, 3637, 2478, 2797, 949, 3080, 3391, 3381, 3011, 4264, 1958, 2133, 2282, 3004, 2020, 4281, 4496, 4067, 2055, 4338, 3044, 3607, 2792, 2119, 2927, 491]. Vector-borne [2478]. vector-valued [2055]. vectors [1857, 2773, 2591, 688]. vegetation [2671]. velocities [3623]. velocity [2536, 1874]. vendors [3942]. venture [4173]. Verification [3528, 1686, 2836]. versatile [2347, 2299, 1284]. version [1125, 3323, 3536]. Versus [1633, 1536, 1019, 1581, 1243, 1660, 3306, 3255, 1522, 597, 1390, 4644, 3200, 907, 4142, 3779, 3262]. vertically [2113]. Very [4705, 2156, 2555]. vessel [3180]. veteran [3091]. veterans [3043]. veterinary [907, 2441]. via [2383, 1510, 4441, 2034, 4682, 1820, 2239, 836, 929, 2026, 1139, 927, 1967, 3010, 4728, 1951, 3322, 1294, 472, 2097, 4424, 3525, 4485, 440, 2489, 4192, 3891, 3419, 3061, 2528, 1597, 4045, 4076, 3661, 4517, 2111, 2256, 2379, 4566, 3162, 2178, 3787, 3502, 3036, 4067, 2591, 3964, 4464, 2443, 3495, 3100, 1798, 2846, 4162, 904, 2792, 3924, 2362, 2703, 1712, 2533, 4591, 3940, 3041, 2812, 3629, 4557, 4291, 2170]. Viability [920]. view [2994, 20]. viewpoint [1492]. VIF [2210, 4434]. vigorous [4531]. vine [4157]. vineyards [3016]. vinyl [3833]. violations [1916, 882]. Violent [2141, 2173]. viral [2729, 2398, 3547, 3132]. virtual [3415, 1159]. virus [4422, 300, 315, 441]. viscoelastic [3322]. visibility [4097]. vision [1717, 591]. visit [2360]. visits [2795]. visual [4120, 2929, 1499]. visualisation [3216]. Visualization [1251]. Visualizing [2487, 2070]. vital [1040]. vocabulary [1592]. Vol [1007]. Volatility [4721, 2277, 2608, 996, 1806, 3212, 2991, 3930, 2581, 3926, 3428, 3901, 3276, 4512, 1814, 2703, 2201, 2119, 1890, 2382]. volatility-trading [1890]. volleyball [4048]. Volume [822, 876, 878, 946, 788, 1105, 4395, 1890, 770, 877]. voluntary [3383]. volunteer [4583]. Voronoi [1582]. vote [799, 931]. voter [799, 2708, 4014]. votes [2920, 4360]. voting [4588, 4360, 4465]. voxel [4462, 1846]. voxel-based [1846]. voxel-wise [4462]. vs [2771, 3304]. VSSI [2645]. vulnerable [4510].

W [2822, 2818]. wage [4195]. Wagtails [1208]. waist [2251]. waiting [4486, 4463]. wake [4608]. Wales [2141, 1521]. walk [1913, 1785]. Walker [52]. wall [1336]. wander [6]. want [1]. war [1462]. warbler [756]. Waring [3870]. warming [1748]. warning [2075, 610]. warping [2772, 890, 4629]. Warsaw [1252]. Was [134]. Washington [2452]. Water [1619, 789, 1717]. waterfowl [2591]. waterpipe [4648]. waters [1840]. Watson [1578, 1465]. wave [3256, 1874]. waved [4548]. Wavelet [2658, 1606, 4122, 1607, 3070, 4386, 3388, 1330, 1333, 3199, 4286, 3083, 2422, 1786, 1952, 3084, 2510, 4521, 1495]. wavelet-based [3199, 2510]. Wavelet-converted [1606]. Wavelets [4712, 660, 1748, 1124, 1020, 3006, 1328, 2066]. Way [1387, 3267, 1128, 599, 3982, 541, 2468, 2715, 341, 3875, 1578, 530, 1087, 2348, 4204, 1418, 505, 2041, 4556, 4295, 265, 2058, 3670, 550, 1513, 1051, 1484, 4689]. ways [1943]. WCDANM [3993]. weakly [3596]. wealth [4373]. Wearing [165]. weather [3663, 4413]. web [3480, 4583, 4450, 2684]. Website [1378]. websites [3513]. wedding [761]. weed [3702]. Weibull [4205, 2820, 2472, 3319, 4469, 1906, 2741, 3834, 3671, 668, 1169, 3520, 4356, 2725, 1530, 4200, 2870, 3566, 2300, 4215, 1763, 1981, 1969, 894, 3062, 846, 1828, 1537, 4554, 2330, 1457, 3922, 4735, 3928, 3493, 4672, 2630, 2412, 2892, 3548, 4064, 3563, 4254, 4677, 4560, 956, 3853, 2660, 4298, 4417, 994, 1532, 2669, 883, 1110, 1339, 3744, 4362, 2936, 3220, 2253]. Weibull- [4356]. Weibull-generalized [2870]. Weibull-Negative-Binomial [3220]. weighing [3718]. Weight [147, 3805, 2627, 925, 2931, 4172, 1803]. Weighted [2299, 1306, 1416, 1368, 2127, 4364, 2051, 4523, 3028, 4423, 1847, 2882, 3297, 3520, 807, 3818, 4081, 3366, 4045, 2931, 3000, 4580, 1864, 4667, 4018, 2255, 2792, 2259, 3208, 3801, 4405, 3741, 1327, 4613, 4446]. Weighting [1833, 3362, 3007, 4061]. Weights [18, 2135, 35, 4058, 1976, 2803, 1026, 893, 3483, 4342]. Weiß [1541]. welfare [3167, 3153]. well [3203, 2491, 2837]. were [3537]. west [4237, 3383, 1472]. Western [1517, 3689]. wet [3180]. wet-on-wet [3180]. wheels [2538, 2863]. where [1201, 931, 1202, 4597]. Which [74, 103, 1587, 3637, 2220, 3531]. whiskers [46]. white [1236, 1218]. white-fronted [1236]. Whitening [1424]. Whither [6]. Whitney [2941, 968, 2486]. Whittle [1782]. Who [2505]. Whose [42, 447]. wide [758, 2560]. widely [2078]. widespread [2293]. width [4142]. Wiener [2116]. Wilcoxon [774, 2941, 2486, 3154]. Wild [2436]. Wild-bootstrapped [2436]. Wilde [1008]. wildfires [3919]. Wiley [4669]. Wilk [840]. William [2653]. willingness [2813]. willow [756]. Wimbledon [981]. win [3883, 1911]. WinBUGS [2458]. wind [3179, 2880, 2002, 2180, 490, 250, 2308, 2626, 4231, 443, 3800]. wind-blown [2626]. Window [1437, 3077, 2165, 3003]. window-wise [3003]. windows [3129]. wine [3234, 2254, 983, 1167]. winners [824]. winter [3016, 1472]. Winters [1700]. wireless [4451]. wise [2771, 4462, 4403, 3003]. Withdrawal [2229, 500]. withdrawals [301]. within [494, 2828, 2518, 2872, 2077, 3853, 282, 2022]. within-group [494]. without [4100, 1396, 4019, 2162, 3960, 1105, 4074, 2108, 4556, 3489, 4430, 4095, 3078]. women [1311, 2740, 2832]. wood [4311]. Wool [1424]. word [3888, 4499, 3682]. Work [183, 2966, 1140, 3356]. Work-In-Progress [183]. workers [4370]. workforce [4496]. working [3739]. Works [1711]. workstations [1841]. World [2664, 4720, 1317, 823, 233, 867, 2274]. world-class [823]. worrying [3340]. wrapped [3400, 381, 1122]. Wynn [138].

X [2822, 1733, 3766, 4221, 902]. X-ray [3766, 4221]. X-S [902]. XII [3541, 1269, 3891, 2350, 2770, 2204, 2572, 1996]. XY = [1102].

yard [85]. year [2304, 3345, 3910, 1234, 339]. Years [111, 369, 879, 3930, 981, 1924, 1556].

Yeast [1659]. Yi [4599]. Yichuan [4599]. Yield [1684, 1641, 470, 2026, 2603, 3639, 790, 3774, 4090, 3375, 2939]. yields [3814]. Yor [2824]. York [2002]. Youden [917]. young [732, 3602]. yuan [1955]. Yule [3417].

z} [1392]. Zambia [619]. Zap [1664]. Zealand [4542, 2865]. Zenga [2563]. Zero [3888, 3836, 4207, 910, 3884, 2624, 1058, 1842, 4098, 4072, 4441, 2307, 4331, 3647, 3859, 3574, 3788, 2766, 4347, 3600, 3009, 3042, 3989, 2360, 3114, 1285, 4635, 2596, 4069, 3981, 3323, 3896, 1601, 4352, 1925, 2661, 3451, 3505, 1808, 3489, 1575, 3969, 2473, 4255, 1749, 1865, 3992, 3932, 2791, 4258, 339, 963, 1164, 2046, 2856, 2122, 2700, 3885, 2812, 1933, 4368]. Zeroadjusted [3836]. zero-bound [1842]. Zeroinflated [3888, 4207, 3884, 2624, 1058, 4098, 4072, 4331, 3859, 3574, 3788, 2766, 4347, 3600, 3009, 3042, 3989, 2360, 3114, 2596, 4069, 3323, 3896, 1601, 4352, 3451, 3505, 1808, 3489, 2473, 3992, 3932, 1164, 2856, 2122, 2700, 3885, 2812, 4368]. zero-inflation [4441, 2046, 1933]. zero-order [1285]. zero-state [4635]. zerotruncated [4258]. zeroes [4252]. zeros [4678, 1808, 3972, 4116, 3416, 3628, 3393, 2056, 1164]. Zhang [1635, 1315]. Zhao [4599]. Zhezhen [4599]. Zipf [852]. Zonation [1610, 1502]. zone [3524]. zoonotic [1948].

References

Geweke:1974:WDY

[1] S. E. Geweke. Why do you want to know? Bulletin in Applied Statistics (BIAS), 1(1):3–5, 1974. CODEN ???? ISSN ????

Kapadia:1974:SP

[2] R. Kapadia. A school poll. Bulletin in Applied Statistics (BIAS), 1(1):7–10, 1974. CODEN ???? ISSN ????

Till:1974:BSN

[3] M. D. Till. On bringing statistics to the non-numerate. Bulletin in Applied Statistics (BIAS), 1(1):11–13, 1974. CODEN ???? ISSN ????

Bedford:1974:SBG

[4] S. E. Bedford. Statistics at the British Glass Industry Research Association. Bulletin in Applied Statistics (BIAS), 1(1):14–18, 1974. CODEN ???? ISSN ????

Ainslie:1974:TTS

[5] Alan C. Ainslie. Thoughts on the teaching of statistics. Bulletin in Applied Statistics (BIAS), 1(1):19–20, 1974. CODEN ???? ISSN ????

TAG:1974:WSY

[6] TAG. Whither should you wander? Bulletin in Applied Statistics (BIAS), 1(1):21–22, 1974. CODEN ???? ISSN ????

Stone:1974:SSE

[7] J. A. R. Stone. Statistics for students of engineering. Bulletin in Applied Statistics (BIAS), 1(1):23–25, 1974. CODEN ???? ISSN ????

Kanji:1974:PSR

[8] G. K. Kanji and W. Lean. A preliminary study of residential prices in Sheffield. Bulletin in Applied Statistics (BIAS), 1(1):26–31, 1974. CODEN ???? ISSN ????

Gilchrist:1974:P

[9] Warren Gilchrist. Polystats. Bulletin in Applied Statistics (BIAS), 1(1):32– 33, 1974. CODEN ???? ISSN ????

Langner:1974:IDW

[10] A. J. Langner. Interlinking of disciplines when teaching mathematics. Bulletin in Applied Statistics (BIAS), 1(1):34–39, 1974. CODEN ???? ISSN ????

Anonymous:1974:RMS

[11] Anonymous. Report from the Manchester–Sheffield School of Probability and Statistics. Bulletin in Applied Statistics (BIAS), 1(1):40–41, 1974. CODEN ???? ISSN ????

Anonymous:1974:PP

[12] Anonymous. Puzzle page. Bulletin in Applied Statistics (BIAS), 1(1):42–43, 1974. CODEN ???? ISSN ????

Anonymous:1974:NN

[13] Anonymous. Notes and news. Bulletin in Applied Statistics (BIAS), 1(1):44– 47, 1974. CODEN ???? ISSN ????

Anonymous:1974:E

[14] Anonymous. Editorial. Bulletin in Applied Statistics (BIAS), 1(1):2–??, 1974. CODEN ???? ISSN ????

Huitson:1974:SDD

[15] A. Huitson. Statistics degree with a difference. Bulletin in Applied Statistics (BIAS), 1(1):6–??, 1974. CODEN ???? ISSN ????

Anonymous:1975:E

[16] Anonymous. Editorial. Bulletin in Applied Statistics (BIAS), 2(1):1–2, 1975. CODEN ???? ISSN ????

Bissell:1975:STC

[17] A. F. Bissell. Statistics and technology I.C.I. (Fibres) Pontypool. Bulletin in Applied Statistics (BIAS), 2(1):3–5, 1975. CODEN ???? ISSN ????

Pridmore:1975:WMS

[18] W. A. Pridmore. Weights & measures — a story of European statistics. Bulletin in Applied Statistics (BIAS), 2(1): 6–16, 1975. CODEN ???? ISSN ????

Back:1975:NUC

[19] H. L. Back and A. Riley. A note on the use of a computer in generating examples for practical classes in biometry teaching to non-mathematicians. Bulletin in Applied Statistics (BIAS), 2(1): 17–18, 1975. CODEN ???? ISSN ????

Dixie:1975:RTJ

[20] J. M. Dixie. The right tool for the job — a personal view of course design. Bulletin in Applied Statistics (BIAS), 2(1):19–27, 1975. CODEN ???? ISSN ????

Holmes:1975:GTD

[21] D. I. Holmes. Graphical techniques in data analysis. Bulletin in Applied Statistics (BIAS), 2(1):27–31, 1975. CODEN ???? ISSN ????

Lomnicki:1975:SRH

[22] Z. A. Lomnicki. Some remarks on how to find problems deserving statistical investigation. Bulletin in Applied Statistics (BIAS), 2(1):33–38, 1975. CODEN ???? ISSN ????

Barber:1975:PCA

[23] D. C. Barber, P. J. Howlett, and R. C. Smart. Principal component analysis in medical research. Bulletin in Applied Statistics (BIAS), 2(1):39–43, 1975. CODEN ???? ISSN ????

Harris:1975:SSA

[24] P. T. Harris. Some statistical aspects of opinion polling and election forecasting. Bulletin in Applied Statistics (BIAS), 2(1):44–51, 1975. CODEN ???? ISSN ????

Goldsmith:1975:TSQ

[25] P. L. Goldsmith. Thirty statistical quotations. Bulletin in Applied Statistics (BIAS), 2(1):52–58, 1975. CODEN ???? ISSN ????

Jackson:1975:ROU

[26] P. H. Jackson. Remarks on‘connectors’ and ‘unifiers’. Bulletin in Applied Statistics (BIAS), 2(1):59–60, 1975. CODEN ???? ISSN ????

Huitson:1975:ITS

[27] A. Huitson. Industrial training for statistical students. Bulletin in Applied Statistics (BIAS), 2(1):61–62, 1975. CODEN ???? ISSN ????

Holmes:1975:RT

[28] D. I. Holmes. A reply to tag. Bulletin in Applied Statistics (BIAS), 2(1):63– 64, 1975. CODEN ???? ISSN ????

Anonymous:1975:QN

[29] Anonymous. Quiz no. 1. Bulletin in Applied Statistics (BIAS), 2(1):65–66, 1975. CODEN ???? ISSN ????

Ward:1975:LE

[30] D. H. Ward, R. Anthony, J. B. Copas, and L. S. Goddard. Letters to the Editor. Bulletin in Applied Statistics (BIAS), 2(1):67–69, 1975. CODEN ???? ISSN ????

Ehrenberg:1975:MS

[31] A. S. C. Ehrenberg. Mathematics and statistics. Bulletin in Applied Statistics (BIAS), 2(2):1–6, 1975. CODEN ???? ISSN ????

Sasieni:1975:EBS

[32] M. W. Sasieni. Entropy and brand switching. Bulletin in Applied Statistics (BIAS), 2(2):7–13, 1975. CODEN ???? ISSN ????

Maher:1975:USR

[33] M. J. Maher. The use of statistics and O.R. in transport planning and engineering. Bulletin in Applied Statistics (BIAS), 2(2):14–19, 1975. CODEN ???? ISSN ????

Bedford:1975:SN

[34] S. E. Bedford. Something for nothing. Bulletin in Applied Statistics (BIAS), 2(2):20–23, 1975. CODEN ???? ISSN ????

Greenman:1975:SCT

[35] Greenman. Some critical thoughts on weights and measures. Bulletin in Applied Statistics (BIAS), 2(2):24–28, 1975. CODEN ???? ISSN ????

Walker:1975:CSO

[36] Monica A. Walker. Choosing a sample for opinion polls and election forecasting. Bulletin in Applied Statistics (BIAS), 2(2):33–38, 1975. CODEN ???? ISSN ????

Griffiths:1975:SV

[37] J. D. Griffiths. Simulating the value of ϕ. Bulletin in Applied Statistics (BIAS), 2(2):39–45, 1975. CODEN ???? ISSN ????

Anderson:1975:SRC

[38] O. D. Anderson. Statisticians in the red corner, social scientists in the blue. Bulletin in Applied Statistics (BIAS), 2(2):46–51, 1975. CODEN ???? ISSN ????

Bligh:1975:MCC

[39] A. A. Bligh. On the mathematical content of courses in applied statistics: calculating the correlation coefficient. Bulletin in Applied Statistics (BIAS), 2(2):52–64, 1975. CODEN ???? ISSN ????

Homles:1975:Q

[40] D. I. Homles. Quiz. Bulletin in Applied Statistics (BIAS), 2(2):65–67, 1975. CODEN ???? ISSN ????

Nicholson:1975:LE

[41] D. Nicholson. Letters to the Editor. Bulletin in Applied Statistics (BIAS), 2(2):67–??, 1975. CODEN ???? ISSN ????

Goldstein:1976:WVE

[42] Harvey Goldstein. Whose values in educational research? Bulletin in Applied Statistics (BIAS), 3(1):1–7, 1976. CO-DEN ???? ISSN ????

Cobby:1976:SCO

[43] J. M. Cobby. Statisticians in clover (and other forage crops). Bulletin in Applied Statistics (BIAS), 3(1):8–14, 1976. CODEN ???? ISSN ????

Jeffers:1976:CAS

[44] J. N. R. Jeffers. Can applied statistics be taught? Bulletin in Applied Statistics (BIAS), 3(1):17–29, 1976. CODEN ???? ISSN ????

Swires-Hennessy:1976:SPC

[45] E. Swires-Hennessy. Some problems of collecting statistics at the BSO. Bulletin in Applied Statistics (BIAS), 3(1): 30–38, 1976. CODEN ???? ISSN ????

Bissell:1976:MKS

[46] A. F. Bissell. My kind of statistics: raindrops on roses and whiskers on kittens. Bulletin in Applied Statistics (BIAS), 3(1):39–43, 1976. CODEN ???? ISSN ????

Smith:1976:SLG

[47] G. W. Smith. The statistician in local government. Bulletin in Applied Statistics (BIAS), 3(1):44–51, 1976. CODEN ???? ISSN ????

Anderson:1976:SAJ

[48] O. D. Anderson. Is seasonally adjusting justified? Bulletin in Applied Statistics (BIAS), 3(1):52–54, 1976. CODEN ???? ISSN ????

Neave:1976:THT

[49] Henry R. Neave. The teaching of hypothesis-testing. Bulletin in Applied Statistics (BIAS), 3(1):55–63, 1976. CODEN ???? ISSN ????

Kapadia:1976:PI

[50] R. Kapadia. The politics of inflation. Bulletin in Applied Statistics (BIAS), 3(1):64–74, 1976. CODEN ???? ISSN ????

Pridmore:1976:SCT

[51] W. A. Pridmore. Some critical thoughts — all squared? Bulletin in Applied Statistics (BIAS), 3(1):75–83, 1976. CODEN ???? ISSN ????

Harris:1976:RMM

[52] P. T. Harris. A reply to Mrs. M. Walker. Bulletin in Applied Statistics (BIAS), 3(1):84–85, 1976. CODEN ???? ISSN ????

Anonymous:1976:QNa

[53] Anonymous. Quiz no. 3. Bulletin in Applied Statistics (BIAS), 3(1):86–87, 1976. CODEN ???? ISSN ????

Sprent:1976:LE

[54] P. Sprent, S. C. Pearce, and Denis H. Ward. Letters to the Editor. Bulletin in Applied Statistics (BIAS), 3(1):88– 90, 1976. CODEN ???? ISSN ????

Anonymous:1976:QNb

[55] Anonymous. QSN newsletter. Bulletin in Applied Statistics (BIAS), 3(1):91– 92, 1976. CODEN ???? ISSN ????

Ehrenberg:1976:BR

[56] A. S. C. Ehrenberg. Book reviews. Bulletin in Applied Statistics (BIAS), 3(1): 93–94, 1976. CODEN ???? ISSN ????

Chatfield:1976:STS

[57] C. Chatfield. A statistical true story. Bulletin in Applied Statistics (BIAS), 3(2):1–18, 1976. CODEN ???? ISSN ????

Haws:1976:SDA

[58] A. P. Haws and A. Hedge. Some developments in the analysis of public attitude data. Bulletin in Applied Statistics (BIAS), 3(2):19–29, 1976. CODEN ???? ISSN ????

Smale:1976:RIS

[59] D. Smale. The role of the industrial statistician. Bulletin in Applied Statistics (BIAS), 3(2):30–38, 1976. CODEN ???? ISSN ????

Bee:1976:MCT

[60] R. E. Bee. The mode of composition of a text. Bulletin in Applied Statistics (BIAS), 3(2):39–48, 1976. CODEN ???? ISSN ????

Neave:1976:NTW

[61] Henry R. Neave. Nonparametric testing — why and how. Bulletin in Applied Statistics (BIAS), 3(2):49–58, 1976. CODEN ???? ISSN ????

Anderson:1976:NDT

[62] O. D. Anderson. Naive discrimination in time-series analysis. Bulletin in Applied Statistics (BIAS), 3(2):59–66, 1976. CODEN ???? ISSN ????

Alderson:1976:ISH

[63] M. R. Alderson and L. Rushton. Information systems in the health service. Bulletin in Applied Statistics (BIAS), 3(2):67–77, 1976. CODEN ???? ISSN ????

Cox:1976:OPS

[64] T. F. Cox. Optimal policies in the stock market. Bulletin in Applied Statistics (BIAS), 3(2):78–88, 1976. CODEN ???? ISSN ????

LeRoux:1976:NME

[65] A. A. Le Roux. The need for mathematics in estimating bias. Bulletin in Applied Statistics (BIAS), 3(2):89–95, 1976. CODEN ???? ISSN ????

Harrison:1976:MKS

[66] P. J. Harrison. My kind of statistics. Bulletin in Applied Statistics (BIAS), 3(2):96–98, 1976. CODEN ???? ISSN ????

Anderson:1976:SJ

[67] O. D. Anderson. Statistical jemmies. Bulletin in Applied Statistics (BIAS), 3(2):99–102, 1976. CODEN ???? ISSN ????

Kanji:1976:LE

[68] Kanji. Letters to the Editor. Bulletin in Applied Statistics (BIAS), 3(2):103– 104, 1976. CODEN ???? ISSN ????

Lindley:1977:SST

[69] D. V. Lindley. A survey of statistical thought. Bulletin in Applied Statistics (BIAS), 4(1):1–17, 1977. CODEN ???? ISSN ????

Goldstein:1977:SE

[70] Harvey Goldstein and Gopal Kanji. Statistical ethics. Bulletin in Applied Statistics (BIAS), 4(1):18–23, 1977. CODEN ???? ISSN ????

Jeffers:1977:DFC

[71] J. N. R. Jeffers. Discriminant functions: a case study. Bulletin in Applied Statistics (BIAS), 4(1):24–38, 1977. CODEN ???? ISSN ????

Finney:1977:SSS

[72] D. J. Finney. Statistics, schools and society. Bulletin in Applied Statistics (BIAS), 4(1):39–54, 1977. CODEN ???? ISSN ????

Bissell:1977:J

[73] A. F. Bissell. The jacknife. Bulletin in Applied Statistics (BIAS), 4(1):55–64, 1977. CODEN ???? ISSN ????

Weatherall:1977:WS

[74] Ms. Rosamund Weatherall. Which? statistics. Bulletin in Applied Statistics (BIAS), 4(1):65–74, 1977. CODEN ???? ISSN ????

Hills:1977:TSP

[75] M. Hills. Two statistical problems of importance in physical anthropology. Bulletin in Applied Statistics (BIAS), 4(1):75–82, 1977. CODEN ???? ISSN ????

Newbold:1977:MKS

[76] Paul Newbold. My kind of statistics. Bulletin in Applied Statistics (BIAS), 4(1):83–87, 1977. CODEN ???? ISSN ????

Smith:1977:TBS

[77] D. J. Smith. Teaching basic statistics by mean of simulations. Bulletin in Applied Statistics (BIAS), 4(1):88–94, 1977. CODEN ???? ISSN ????

Anderson:1977:TLB

[78] O. D. Anderson. Tales from the life of Bath: Dr Chatfield’s cover. Bulletin in Applied Statistics (BIAS), 4(1):95–102, 1977. CODEN ???? ISSN ????

Gilohrist:1977:SCT

[79] R. Gilohrist. Some comments on the teaching of the likelihood ratio test. Bulletin in Applied Statistics (BIAS), 4 (1):103–107, 1977. CODEN ???? ISSN ????

Norcliffe:1977:SCN

[80] A. Norcliffe and H. Porteous. Some comments on the need for mathematics in estimating bias. Bulletin in Applied Statistics (BIAS), 4(1):108–110, 1977. CODEN ???? ISSN ????

Anonymous:1977:CIS

[81] Anonymous. 1975 current index to statistics. Bulletin in Applied Statistics (BIAS), 4(1):117–??, 1977. CO-DEN ???? ISSN ????

Cane:1977:WM

[82] R. Cane. What are models for? Bulletin in Applied Statistics (BIAS), 4(2): 1–8, 1977. CODEN ???? ISSN ????

Freeman:1977:AVI

[83] G. H. Freeman. The analysis of variance — has it had its day? Bulletin in Applied Statistics (BIAS), 4(2):9–15, 1977. CODEN ???? ISSN ????

Bissell:1977:IIS

[84] A. F. Bissell. Inconsistency of inspection standards — a case study. Bulletin in Applied Statistics (BIAS), 4(2):16– 27, 1977. CODEN ???? ISSN ????

Porteous:1977:SMY

[85] H. L. Porteous. The saga of the megalithic yard. Bulletin in Applied Statistics (BIAS), 4(2):28–39, 1977. CODEN ???? ISSN ????

Jeffers:1977:LDS

[86] J. N. R. Jeffers. Lake district soils. Bulletin in Applied Statistics (BIAS), 4(2): 40–52, 1977. CODEN ???? ISSN ????

Ehrenberg:1977:TED

[87] A. S. C. Ehrenberg. Three exercises in data presentation. Bulletin in Applied Statistics (BIAS), 4(2):53–69, 1977. CODEN ???? ISSN ????

Harvey:1977:CSC

[88] R. A. Harvey. The contribution of structural change to manufacturing labour productivity. Bulletin in Applied Statistics (BIAS), 4(2):70–76, 1977. CODEN ???? ISSN ????

Nelder:1977:MKS

[89] J. A. Nelder. My kind of statistics. Bulletin in Applied Statistics (BIAS), 4(2): 77–81, 1977. CODEN ???? ISSN ????

Gauvain:1977:ISE

[90] Suzette Gauvain. The importance of statistics and epidemiology in occupational health. Bulletin in Applied Statistics (BIAS), 4(2):82–101, 1977. CODEN ???? ISSN ????

Anderson:1977:GCF

[91] O. R. Anderson. Generalised cost function regression analysis when the errors are not normally distributed. Bulletin in Applied Statistics (BIAS), 4(2):102– 115, 1977. CODEN ???? ISSN ????

Cooper:1977:SMC

[92] I. Cooper. Seminar on ‘Mini Computers’, Sheffield 8th December 1976. Bulletin in Applied Statistics (BIAS), 4(2): 116–119, 1977. CODEN ???? ISSN ????

Kanji:1977:RBS

[93] Kanji. Report of a business statistics seminar. Bulletin in Applied Statistics (BIAS), 4(2):120–121, 1977. CODEN ???? ISSN ????

Chanter:1977:LE

[94] D. O. Chanter. Letters to the Editor. Bulletin in Applied Statistics (BIAS), 4 (2):122–123, 1977. CODEN ???? ISSN ????

Sprent:1978:FAS

[95] P. Sprent. Fashions in applied statistics. Bulletin in Applied Statistics (BIAS), 5(1):1–7, 1978. CODEN ???? ISSN ????

Smith:1978:REL

[96] Gordon L. Smith, Christine Michie, and James A. Pope Marine. Robust estimators of location. Bulletin in Applied Statistics (BIAS), 5(1):8–21, 1978. CODEN ???? ISSN ????

Freeman:1978:CRS

[97] G. H. Freeman. Consultation and research in a statistics section. Bulletin in Applied Statistics (BIAS), 5(1):22– 32, 1978. CODEN ???? ISSN ????

Bissell:1978:N

[98] A. F. Bissell. Nomograms. Bulletin in Applied Statistics (BIAS), 5(1):33–48, 1978. CODEN ???? ISSN ????

Chanter:1978:CS

[99] D. O. Chanter. Communication and the statistician. Bulletin in Applied Statistics (BIAS), 5(1):49–56, 1978. CODEN ???? ISSN ????

Sampford:1978:MKS

[100] M. R. Sampford. My kind of statistics. Bulletin in Applied Statistics (BIAS), 5(1):57–62, 1978. CODEN ???? ISSN ????

Pridmore:1978:MP

[101] W. A. Pridmore. The motion picture. Bulletin in Applied Statistics (BIAS), 5(1):63–81, 1978. CODEN ???? ISSN ????

Jones:1978:SEU

[102] B. Jones. Some examples of the use of an algorithm for constructing optimal block designs. Bulletin in Applied Statistics (BIAS), 5(1):82–92, 1978. CODEN ???? ISSN ????

Anderson:1978:NAT

[103] O. D. Anderson. A note on analysing time series data which accumulate over unequal observation intervals. Bulletin in Applied Statistics (BIAS), 5(1):91– 94, 1978. CODEN ???? ISSN ????

Harvey:1978:FAP

[104] R. A. Harvey. Factors affecting productivity. Bulletin in Applied Statistics (BIAS), 5(1):95–110, 1978. CODEN ???? ISSN ????

Anonymous:1978:LEM

[105] Anonymous. Letters to the Editor: Mr. Kanji. Bulletin in Applied Statistics (BIAS), 5(1):111–112, 1978. CODEN ???? ISSN ????

Bissell:1978:AUT

[106] A. F. Bissell. An attempt to unify the theory of quality control procedures. Bulletin in Applied Statistics (BIAS), 5 (2):113–128, 1978. CODEN ???? ISSN ????

Cheek:1978:PER

[107] Peter Cheek. Problems with estimating the rate of inflation. Bulletin in Applied Statistics (BIAS), 5(2):129–136, 1978. CODEN ???? ISSN ????

Thomson:1978:SR

[108] D. M. Thomson. Statistics and the ratemeter. Bulletin in Applied Statistics (BIAS), 5(2):137–145, 1978. CO-DEN ???? ISSN ????

Springall:1978:RMS

[109] A. Springall. A review of multidimensional scaling. Bulletin in Applied Statistics (BIAS), 5(2):146–192, 1978. CODEN ???? ISSN ????

Slater:1978:STB

[110] M. Slater. Significance tests or Bayesian analysis: an elementary example. Bulletin in Applied Statistics (BIAS), 5(2):193–217, 1978. CODEN ???? ISSN ????

Dimming:1978:YBE

[111] I. G. Dimming. 50 years of biometry at East Malling Q. Bulletin in Applied Statistics (BIAS), 5(2):218–227, 1978. CODEN ???? ISSN ????

Ehrenberg:1978:RED

[112] A. S. C. Ehrenberg. A revised exercise in data processing. Bulletin in Applied Statistics (BIAS), 5(2):228–230, 1978. CODEN ???? ISSN ????

Wells:1978:PAF

[113] I. S. Wells. Probabilistic approach to factual questions: a real case. Bulletin in Applied Statistics (BIAS), 5(2):231– 233, 1978. CODEN ???? ISSN ????

Jackson:1978:LE

[114] Paul Jackson. Letter to the editor. Bulletin in Applied Statistics (BIAS), 5(2): 234–237, 1978. CODEN ???? ISSN ????

Kanchi:1978:BAS

[115] Gopal Kanchi and A. Huitson. Bulletin in Applied Statistics (BIAS). Bulletin in Applied Statistics (BIAS), 5(2):112– ??, 1978. CODEN ???? ISSN ????

Freeman:1979:EDT

[116] G. H. Freeman. Experimental design from the time of Fisher to the 1970’s. Bulletin in Applied Statistics (BIAS), 6 (1):1–6, 1979. CODEN ???? ISSN ????

Campbell:1979:SPA

[117] Norm A. Campbell. Some practical aspects of canonical variate analysis. Bulletin in Applied Statistics (BIAS), 6(1): 7–18, 1979. CODEN ???? ISSN ????

Goldstein:1979:APC

[118] Harvey Goldstein. Age, period and cohort effects — a confounded confusion. Bulletin in Applied Statistics (BIAS), 6(1):19–24, 1979. CODEN ???? ISSN ????

Jones:1979:CAS

[119] B. Jones. Cluster analysis of some social survey data. Bulletin in Applied Statistics (BIAS), 6(1):25–56, 1979. CODEN ???? ISSN ????

Cox:1979:PSR

[120] Trevor F. Cox. Probability and Special Relativity. Bulletin in Applied Statistics (BIAS), 6(1):57–65, 1979. CODEN ???? ISSN ????

Chatfield:1979:RCT

[121] C. Chatfield. A report on the 1978 I.O.S. Conference on Time-Series Analysis. Bulletin in Applied Statistics (BIAS), 6(1):66–72, 1979. CODEN ???? ISSN ????

Mitchell:1979:CPU

[122] B. Mitchell and E. Swires-Hennessy. The census of production and its uses. Bulletin in Applied Statistics (BIAS), 6(1):73–88, 1979. CODEN ???? ISSN ????

Phillips:1979:TM

[123] C. I. Phillips. Treatment of myopia. Bulletin in Applied Statistics (BIAS), 6(1):89–100, 1979. CODEN ???? ISSN ????

Chatfield:1979:LEa

[124] C. Chatfield. Letter to the Editor. Bulletin in Applied Statistics (BIAS), 6(1): 101–102, 1979. CODEN ???? ISSN ????

Pearce:1979:FFT

[125] S. C. Pearce. The Fisherian field trial today. Bulletin in Applied Statistics (BIAS), 6(2):103–111, 1979. CODEN ???? ISSN ????

Bissell:1979:LDE

[126] A. F. Bissell. Life distribution of expendable components. Bulletin in Applied Statistics (BIAS), 6(2):112–136, 1979. CODEN ???? ISSN ????

Healy:1979:DMS

[127] M. J. R. Healy. Does medical statistics exist? Bulletin in Applied Statistics (BIAS), 6(2):137–182, 1979. CODEN ???? ISSN ????

Jeffers:1979:EDA

[128] J. N. R. Jeffers. Exploratory data analysis of tabulated information. Bulletin in Applied Statistics (BIAS), 6(2):183– 194, 1979. CODEN ???? ISSN ????

Jones:1979:PLG

[129] K. F. Jones and N. Kay. Pearl’s Law and the growth of the Scottish National Party. Bulletin in Applied Statistics (BIAS), 6(2):195–214, 1979. CODEN ???? ISSN ????

Harvey:1979:OMP

[130] R. A. Harvey and S. Morris. Output measures for productivity comparisons. Bulletin in Applied Statistics (BIAS), 6 (2):215–221, 1979. CODEN ???? ISSN ????

Chatfield:1979:LEb

[131] C. Chatfield. Letter to the Editor. Bulletin in Applied Statistics (BIAS), 6(2): 222–??, 1979. CODEN ???? ISSN ????

Wynn:1980:SSP

[132] H. P. Wynn. Should statistics be professionalized? Bulletin in Applied Statistics (BIAS), 7(1):1–9, 1980. CO-DEN ???? ISSN ????

Dyke:1980:CFE

[133] G. V. Dyke. Covariance and field experiments. Bulletin in Applied Statistics (BIAS), 7(1):10–21, 1980. CODEN ???? ISSN ????

Ryder:1980:HWE

[134] K. Ryder. But how was the experiment done? Bulletin in Applied Statistics (BIAS), 7(1):22–33, 1980. CODEN ???? ISSN ????

Avery:1980:DD

[135] J. D. Avery and J. S. Baron. A degree of difficulty. Bulletin in Applied Statistics (BIAS), 7(1):34–53, 1980. CODEN ???? ISSN ????

Boyd:1980:MND

[136] K. T. Boyd. Multi-nomial distributions and ratio-estimates. Bulletin in Applied Statistics (BIAS), 7(1):54–58, 1980. CODEN ???? ISSN ????

Stone:1980:S

[137] Richard Stone. Sigmoids. Bulletin in Applied Statistics (BIAS), 7(1):59–119, 1980. CODEN ???? ISSN ????

Boyd:1980:SSP

[138] K. T. Boyd. Should statistics be professionalized? A reply to Dr. Wynn.

Bulletin in Applied Statistics (BIAS), 7 (1):120–123, 1980. CODEN ???? ISSN ????

Phillips:1980:LE

[139] C. I. Phillips. Letter to the Editor. Bulletin in Applied Statistics (BIAS), 7(1): 124–??, 1980. CODEN ???? ISSN ????

Freeman:1980:PDE

[140] G. H. Freeman. Practical designs for eliminating positional effects. Bulletin in Applied Statistics (BIAS), 7(2):125– 140, 1980. CODEN ???? ISSN ????

Prescott:1980:RSR

[141] P. Prescott. A review of some robust data analysis and multiple outlier detection procedures. Bulletin in Applied Statistics (BIAS), 7(2):141–158, 1980. CODEN ???? ISSN ????

Lewis:1980:FER

[142] S. M. Lewis and A. M. Dean. Factorial experiments in resolvable generalized cyclic designs. Bulletin in Applied Statistics (BIAS), 7(2):159–167, 1980. CODEN ???? ISSN ????

Jones:1980:CSC

[143] Byron Jones. The computer as a statistical consultant. Bulletin in Applied Statistics (BIAS), 7(2):168–195, 1980. CODEN ???? ISSN ????

Bissell:1980:RTU

[144] A. F. Bissell. The reference test in the UK average quantity system. Bulletin in Applied Statistics (BIAS), 7(2):196– 202, 1980. CODEN ???? ISSN ????

Bravo:1980:SIV

[145] P. C. Bravo. Sampling inspection by variables with known variance. Bulletin in Applied Statistics (BIAS), 7(2):203– 222, 1980. CODEN ???? ISSN ????

Read:1980:SAS

[146] K. L. Q. Read, A. A. Noura, and R. R. Harris. Survival analysis for a sequenced clinical trial with covariates. Bulletin in Applied Statistics (BIAS), 7 (2):223–266, 1980. CODEN ???? ISSN ????

Chapman:1980:BWL

[147] P. F. Chapman and J. G. Fryer. Birth weight and the linear statistical model. Bulletin in Applied Statistics (BIAS), 7 (2):267–306, 1980. CODEN ???? ISSN ????

Thompson:1981:SRD

[148] R. Thompson. Some recent developments in the analysis and design of quantitative genetic experiments. Bulletin in Applied Statistics (BIAS), 8(1): 1–11, 1981. CODEN ???? ISSN ????

Healy:1981:SLS

[149] M. J. R. Healy. A subroutine library for statisticians. Bulletin in Applied Statistics (BIAS), 8(1):12–25, 1981. CODEN ???? ISSN ????

Dickey:1981:EUU

[150] J. M. Dickey. Expert uncertainty and the use of subjective-probability models. Bulletin in Applied Statistics (BIAS), 8(1):26–56, 1981. CODEN ???? ISSN ????

Represent:1981:ADG

[151] P. T. Represent. Anagramatic decomposition or I go to mop Canadian matrices. Bulletin in Applied Statistics (BIAS), 8(1):57–61, 1981. CODEN ???? ISSN ????

Ashford:1981:AIH

[152] J. R. Ashford and Alison Wearing. The assessment of industrial health hazards on the basis of observational data — some statistical pitfalls. Bulletin in Applied Statistics (BIAS), 8(1):62–81, 1981. CODEN ???? ISSN ???? See comments [165].

Shirley:1981:PRE

[153] Eryl Shirley. Problems relating to experimental design and procedures in laboratory studies on animals. Bulletin in Applied Statistics (BIAS), 8(1):82– 90, 1981. CODEN ???? ISSN ????

Cockayne:1981:CUS

[154] G. Cockayne, G. Kanji, and W. Tyson. The collection and use of sample data on public transport demand. Bulletin in Applied Statistics (BIAS), 8(1):91– 108, 1981. CODEN ???? ISSN ????

Bercai:1981:PRP

[155] Andrew Bercai. A proposal to resolve a potential procedural inconsistency affecting some parameters in statistical inference. Bulletin in Applied Statistics (BIAS), 8(1):109–114, 1981. CODEN ???? ISSN ???? See note [174].

Bravo:1981:ACS

[156] P. C. Bravo. Amendments and corrections sampling inspection by variables with known variance. Bulletin in Applied Statistics (BIAS), 8(1):116–117, 1981. CODEN ???? ISSN ????

Anonymous:1981:NOU

[157] Anonymous. A new Open University statistics course. Bulletin in Applied Statistics (BIAS), 8(1):118–119, 1981. CODEN ???? ISSN ????

English:1981:PSF

[158] Brian English and Peter Freeman. Prehistoric Scottish fish bones. Bulletin in Applied Statistics (BIAS), 8(2):120– 133, 1981. CODEN ???? ISSN ????

Yates:1981:R

[159] Frank Yates. Rothamsted in the 1930s. Bulletin in Applied Statistics (BIAS), 8 (2):134–156, 1981. CODEN ???? ISSN ????

Kanji:1981:HDS

[160] G. K. Kanji. History of the development of statistics in India. Bulletin in Applied Statistics (BIAS), 8(2):157– 167, 1981. CODEN ???? ISSN ????

Sharot:1981:PSD

[161] T. Sharot. A problem in survey design — changing the sampling frame. Bulletin in Applied Statistics (BIAS), 8(2): 168–178, 1981. CODEN ???? ISSN ????

Hutchison:1981:USG

[162] Dougal Hutchison. The use of statistics in government decision-making with particular reference to the Reports of Royal Commissions. Bulletin in Applied Statistics (BIAS), 8(2):179–223, 1981. CODEN ???? ISSN ????

Hope:1981:CSS

[163] R. P. Hope. Computer simulation of a simple hypothesis test a useful tool when teaching a first statistics course. Bulletin in Applied Statistics (BIAS), 8 (2):224–230, 1981. CODEN ???? ISSN ????

Oliver:1981:PTP

[164] F. R. Oliver. Process time, Pearl’s Law, and the Scottish National Party. Bulletin in Applied Statistics (BIAS), 8(2): 231–237, 1981. CODEN ???? ISSN ????

Darby:1981:AIH

[165] S. C. Darby. The assessment of industrial health hazards — some comments on the paper by J. R. Ashford and Alison Wearing. Bulletin in Applied Statistics (BIAS), 8(2):238–244, 1981. CODEN ???? ISSN ???? See [152].

Huitson:1981:BR

[166] A. Huitson. Book review. Bulletin in Applied Statistics (BIAS), 8(2):245– 247, 1981. CODEN ???? ISSN ????

Bissell:1982:ORF

[167] A. F. Bissell. Ordered random fractions for chronological sampling. Bulletin in Applied Statistics (BIAS), 9(1): 1–4, 1982. CODEN ???? ISSN ????

Kirby:1982:AVD

[168] Andrew Kirby. Analysis of variance in a descriptive context: a geographic example. Bulletin in Applied Statistics (BIAS), 9(1):5–15, 1982. CODEN ???? ISSN ????

Titterington:1982:IB

[169] D. M. Titterington. Irreverent Bayes. Bulletin in Applied Statistics (BIAS), 9(1):16–18, 1982. CODEN ???? ISSN ????

Vegelius:1982:NTP

[170] Jan Vegelius. A note on the tie problem for the Siegel–Tukey non-parametric dispersion test. Bulletin in Applied Statistics (BIAS), 9(1):19–23, 1982. CODEN ???? ISSN ????

Gadsden:1982:SAC

[171] R. J. Gadsden. Statistical analysis of cometary data. Bulletin in Applied Statistics (BIAS), 9(1):24–36, 1982. CODEN ???? ISSN ????

Neave:1982:THT

[172] Henry R. Neave. Teaching hypothesistesting by simulation. Bulletin in Applied Statistics (BIAS), 9(1):37–40, 1982. CODEN ???? ISSN ????

Thunhurst:1982:MHC

[173] Colin Thunhurst. Measuring the health consequence of social and economic policy. Bulletin in Applied Statistics (BIAS), 9(1):41–59, 1982. CODEN ???? ISSN ????

Shoesmith:1982:NBP

[174] E. Shoesmith. A note on Bercai’s ‘Proposal to Resolve A Potential Procedural Inconsistency’. Bulletin in Applied Statistics (BIAS), 9(1):60–66, 1982. CODEN ???? ISSN ???? See [155].

Barker:1982:TPC

[175] N. Barker, R. J. Hews, A. Huitson, and J. Poloniecki. The two period cross over trial. Bulletin in Applied Statistics (BIAS), 9(1):67–115, 1982. CO-DEN ???? ISSN ????

Barron:1982:LE

[176] J. S. Barron. Letters to the Editor. Bulletin in Applied Statistics (BIAS), 9(1): 117–118, 1982. CODEN ???? ISSN ????

Darekar:1982:BRa

[177] B. Darekar. Book review. Bulletin in Applied Statistics (BIAS), 9(1):119– 121, 1982. CODEN ???? ISSN ????

Reid:1982:BR

[178] Alan Reid. Book review. Bulletin in Applied Statistics (BIAS), 9(1):122– 124, 1982. CODEN ???? ISSN ????

Huitson:1982:BR

[179] Alan Huitson. Book review. Bulletin in Applied Statistics (BIAS), 9(1):125– 126, 1982. CODEN ???? ISSN ????

Jeffers:1982:BR

[180] J. N. R. Jeffers. Book review. Bulletin in Applied Statistics (BIAS), 9(1):127– 128, 1982. CODEN ???? ISSN ????

Goldsmith:1982:USM

[181] P. L. Goldsmith. Using statistics in man-made fibres research. Bulletin in Applied Statistics (BIAS), 9(2):129– 145, 1982. CODEN ???? ISSN ????

Herzberg:1982:EUA

[182] Agnee M. Herzberg. An example of the use of Andrews’ plots to detect time variations in model parameters and to detect outlying observations. Bulletin in Applied Statistics (BIAS), 9(2):146– 154, 1982. CODEN ???? ISSN ????

Bissell:1982:WPA

[183] A. F. Bissell. Work-in-progress audit. Bulletin in Applied Statistics (BIAS), 9 (2):155–164, 1982. CODEN ???? ISSN ????

Neave:1982:NLO

[184] Henry R. Neave. A new look at an old test. Bulletin in Applied Statistics (BIAS), 9(2):165–178, 1982. CODEN ???? ISSN ????

Croucher:1982:AAT

[185] John S. Croucher. Anglo–Australian Test cricket dismissals 1946–1980. Bulletin in Applied Statistics (BIAS), 9(2): 179–193, 1982. CODEN ???? ISSN ????

France:1982:MDR

[186] J. France and A. P. Haws. A model for developing a recreation strategy. Bulletin in Applied Statistics (BIAS), 9(2): 194–211, 1982. CODEN ???? ISSN ????

Raggett:1982:SME

[187] G. F. Raggett. A stochastic model of the Eyam Plague. Bulletin in Applied Statistics (BIAS), 9(2):212–225, 1982. CODEN ???? ISSN ????

Back:1982:MCA

[188] H. L. Back. Multiple choice answers generated by computer for service courses in statistics — a useful compromise on the chore of ‘calculation to aid the understanding’. Bulletin in Applied Statistics (BIAS), 9(2):226– 230, 1982. CODEN ???? ISSN ????

James:1982:BR

[189] M. C. James. Book review. Bulletin in Applied Statistics (BIAS), 9(2):231– 232, 1982. CODEN ???? ISSN ????

Buitson:1982:BR

[190] Alan Buitson. Book review. Bulletin in Applied Statistics (BIAS), 9(2):233– 235, 1982. CODEN ???? ISSN ????

Holmes:1982:BR

[191] D. I. Holmes. Book review. Bulletin in Applied Statistics (BIAS), 9(2):236– 237, 1982. CODEN ???? ISSN ????

Hoare:1982:BR

[192] R. C. Hoare. Book review. Bulletin in Applied Statistics (BIAS), 9(2):238– 239, 1982. CODEN ???? ISSN ????

Darekar:1982:BRb

[193] B. Darekar. Book review. Bulletin in Applied Statistics (BIAS), 9(2):240– 241, 1982. CODEN ???? ISSN ????

Rao:1982:BR

[194] T. Subba Rao. Book review. Bulletin in Applied Statistics (BIAS), 9(2):242– 245, 1982. CODEN ???? ISSN ????

Games:1983:UCD

[195] Paul A. Games. Use of contrasts in a design of experiments — Anova course. Bulletin in Applied Statistics (BIAS), 10(1):1–17, 1983. CODEN ???? ISSN ????

Riley:1983:HAV

[196] Janet Riley, I. Bekele, and Brenda Shrewsbury. How an analysis of variance is affected by the degree of precision of the data. Bulletin in Applied Statistics (BIAS), 10(1):18–43, 1983. CODEN ???? ISSN ????

Oliver:1983:IES

[197] F. R. Oliver. Imports and the EEC — some regression estimates compared. Bulletin in Applied Statistics (BIAS), 10(1):44–50, 1983. CODEN ???? ISSN ????

Holt:1983:LM

[198] D. Holt. Living with a model. Bulletin in Applied Statistics (BIAS), 10(1):51– 86, 1983. CODEN ???? ISSN ????

Bissell:1983:GBT

[199] A. F. Bissell. Grouping of breaks on a textile machine. Bulletin in Applied Statistics (BIAS), 10(1):87–96, 1983. CODEN ???? ISSN ????

Ubogu:1983:FNG

[200] Roland E. Ubogu. Forecasting Nigerian governments expenditures: a multiple regression approach. Bulletin in Applied Statistics (BIAS), 10(1):97–124, 1983. CODEN ???? ISSN ????

Raggett:1983:BR

[201] G. F. Raggett. Book reviews. Bulletin in Applied Statistics (BIAS), 10 (1):125–127, 1983. CODEN ???? ISSN ????

Kanji:1983:BR

[202] G. K. Kanji. Book reviews. Bulletin in Applied Statistics (BIAS), 10(1):127– 129, 1983. CODEN ???? ISSN ????

Huitson:1983:BR

[203] Alan Huitson. Book reviews. Bulletin in Applied Statistics (BIAS), 10 (1):129–131, 1983. CODEN ???? ISSN ????

Anonymous:1983:BRa

[204] Anonymous. Book reviews. Bulletin in Applied Statistics (BIAS), 10(1):131– 132, 1983. CODEN ???? ISSN ????

Holmes:1983:BR

[205] D. I. Holmes. Book reviews. Bulletin in Applied Statistics (BIAS), 10(1):132– 133, 1983. CODEN ???? ISSN ????

Reid:1983:BR

[206] Alan Reid. Book reviews. Bulletin in Applied Statistics (BIAS), 10(1):133– 135, 1983. CODEN ???? ISSN ????

Freeman:1983:CRB

[207] G. H. Freeman. Consultancy and research in biometrics — some thoughts inspired by reading the biography of R. A. Fisher. Bulletin in Applied Statistics (BIAS), 10(2):138–157, 1983. CODEN ???? ISSN ????

Morgan:1983:ITD

[208] Byron J. T. Morgan. Illustration of three-dimensional surfaces. Bulletin in Applied Statistics (BIAS), 10(2):158– 166, 1983. CODEN ???? ISSN ????

Morgan:1983:TSA

[209] D. D. V. Morgan and J. L. Zambrano R. Teaching statistics to agricultural engineering students in Mexico. Bulletin in Applied Statistics (BIAS), 10 (2):167–174, 1983. CODEN ???? ISSN ????

Campbell:1983:SAA

[210] M. J. Campbell. Spectral analysis applied to physiological signals in human subjects. Bulletin in Applied Statistics (BIAS), 10(2):175–193, 1983. CODEN ???? ISSN ????

Buckland:1983:MCM

[211] Stephen T. Buckland. Monte Carlo methods for confidence interval estimation using the bootstrap technique. Bulletin in Applied Statistics (BIAS), 10(2):194–212, 1983. CODEN ???? ISSN ????

Tuan:1983:SSM

[212] Chyau Tuan. Sample survey methods and family income data. Bulletin in Applied Statistics (BIAS), 10(2):213– 230, 1983. CODEN ???? ISSN ????

Back:1983:NDM

[213] H. L. Back and P. W. Murphy. A note on derived means of scaled data. Bulletin in Applied Statistics (BIAS), 10 (2):231–239, 1983. CODEN ???? ISSN ????

Liddell:1983:D

[214] Douglas Liddell. On the distributions of X^{2} and F. Bulletin in Applied Statistics (BIAS), 10(2):240–242, 1983. CO-DEN ???? ISSN ????

Matthews:1983:BR

[215] L. R. Matthews. Book reviews. Bulletin in Applied Statistics (BIAS), 10 (2):244–247, 1983. CODEN ???? ISSN ????

Anonymous:1983:BRb

[216] Anonymous. Book reviews. Bulletin in Applied Statistics (BIAS), 10(2):247– 249, 1983. CODEN ???? ISSN ????

Reed:1983:BR

[217] Alan Reed. Book reviews. Bulletin in Applied Statistics (BIAS), 10(2):249– 251, 1983. CODEN ???? ISSN ????

Haigh:1983:BR

[218] John Haigh. Book reviews. Bulletin in Applied Statistics (BIAS), 10(2):251– 253, 1983. CODEN ???? ISSN ????

Ahmad:1983:BR

[219] R. Ahmad. Book reviews. Bulletin in Applied Statistics (BIAS), 10(2):253– 255, 1983. CODEN ???? ISSN ????

Chanter:1984:CSD

[220] D. O. Chanter. Curtailed sigmoid doseresponse models for fungicide experiments. Journal of Applied Statistics, 11(1):2–11, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Clarke:1984:AAR

[221] M. R. Clarke. An analytical approach to rounding. Journal of Applied Statistics, 11(1):12–20, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mangles:1984:UMT

[222] T. H. Mangles. The use of microcomputers in the teaching of statistics. Journal of Applied Statistics, 11(1):21– 28, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bishop:1984:RHO

[223] Simon H. Bishop and Byron Jones. A review of higher order crossover designs. Journal of Applied Statistics, 11(1):29–50, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hill:1984:ICE

[224] I. D. Hill. Interpreting ‘crazy’ estimates. Journal of Applied Statistics, 11(1):51–53, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tricker:1984:ERD

[225] Tony Tricker. Effects of rounding data sampled from the exponential distribution. Journal of Applied Statistics, 11(1):54–87, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yeomans:1984:TAE

[226] K. A. Yeomans. Is there an alternative to the European farm typology? Journal of Applied Statistics, 11(1):88–119, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRa

[227] Anonymous. Book reviwes. Journal of Applied Statistics, 11(1):120–121, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRb

[228] Anonymous. Book reviews. Journal of Applied Statistics, 11(1):123–125, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRc

[229] Anonymous. Book reviews. Journal of Applied Statistics, 11(1):121–??, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRd

[230] Anonymous. Book reviews. Journal of Applied Statistics, 11(1):122–??, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRe

[231] Anonymous. Book reviews. Journal of Applied Statistics, 11(1):123–??, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:E

[232] Anonymous. Editorial. Journal of Applied Statistics, 11(1):1–??, 1984. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mullins:1984:SPR

[233] Eamonn Mullins. Statistics, Plato and the real world. Journal of Applied Statistics, 11(2):136–141, 1984. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khan:1984:EII

[234] M. Z. Khan, M. Ilyas, and I. Saud. Entropy and internal information of printed Urdu. Journal of Applied Statistics, 11(2):142–154, 1984. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kiregyera:1984:SSS

[235] Ben Kiregyera. Statistics for smallscale industries in East Africa: problems of collection and utilization. Journal of Applied Statistics, 11(2):155– 162, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Loukas:1984:DPP

[236] S. Loukas. On the delay in publication of papers in scientific journals. Journal of Applied Statistics, 11(2):163– 169, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harris:1984:FAH

[237] R. R. Harris and E. F. Harding. The fiducial argument and Hacking’s Principle of Irrelevance. Journal of Applied Statistics, 11(2):170–181, 1984. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Crocker:1984:THT

[238] G. R. Crocker and D. E. Wright. Teaching hypothesis tests by simulation: a problem oriented approach. Journal of Applied Statistics, 11(2):182–187, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Davidson:1984:SSS

[239] Michael J. Davidson and Tony Bendell. Some statistics of statistics. Journal of Applied Statistics, 11(2):188–224, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dhanoa:1984:BPE

[240] M. S. Dhanoa and J. France. A BA-SIC program for the evaluation of the gamma functions. Journal of Applied Statistics, 11(2):225–228, 1984. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:LE

[241] Anonymous. Letters to the Editor. Journal of Applied Statistics, 11(2): 229–230, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRf

[242] Anonymous. Book review. Journal of Applied Statistics, 11(2):231–232, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRg

[243] Anonymous. Book review. Journal of Applied Statistics, 11(2):233–234, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1984:BRh

[244] Anonymous. Book review. Journal of Applied Statistics, 11(2):234–235, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Freeman:1985:AII

[245] G. H. Freeman. The analysis and interpretation of interactions. Journal of Applied Statistics, 12(1):3–10, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goldsmith:1985:OPM

[246] Peter L. Goldsmith. On opinion polls and multi-party elections. Journal of Applied Statistics, 12(1):11–17, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:1985:SES

[247] D. J. Hand. Statistical expert systems: necessary attributes. Journal of Applied Statistics, 12(1):19–27, 1985. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Howard:1985:ARA

[248] P. J. A. Howard and D. M. Howard. The application of ridit analysis to phenological observations. Journal of Applied Statistics, 12(1):29–35, 1985. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silver:1985:PSN

[249] M. S. Silver. Parameter stability and normative market segmentation. Journal of Applied Statistics, 12(1):37–47, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kanji:1985:MMW

[250] G. K. Kanji. A mixture model for wind shear data. Journal of Applied Statistics, 12(1):49–58, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Duckworth:1985:BBB

[251] Frank C. Duckworth and Andrew C. Hill. Brain of Britain: brain or chance? Journal of Applied Statistics, 12(1):59– 69, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Osanaiye:1985:MCD

[252] P. A. Osanaiye. Multiple chain-deferred inspection plans and their compatibility with the multiple plans in MIL-STD-105D and equivalent schemes. Journal of Applied Statistics, 12(1):71– 81, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aleong:1985:EDM

[253] J. Aleong and D. Howard. Extensions of the Duncan’s Multiple Range Test for unbalanced data. Journal of Applied Statistics, 12(1):83–90, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kiregyera:1985:ARS

[254] B. Kiregyera, S. O. Akach, and Hani Afifi. Application of recapturesampling technique to the study of Matatu transport operations in Nairobi (Kenya). Journal of Applied Statistics, 12(1):91–100, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRa

[255] Anonymous. Book reviews. Journal of Applied Statistics, 12(1):102–103, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRb

[256] Anonymous. Book reviews. Journal of Applied Statistics, 12(1):101–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRc

[257] Anonymous. Book reviews. Journal of Applied Statistics, 12(1):102–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRd

[258] Anonymous. Book reviews. Journal of Applied Statistics, 12(1):103–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bound:1985:LE

[259] John Bound. Letter to the Editor. Journal of Applied Statistics, 12(1): 104–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jeffers:1985:DLB

[260] J. N. R. Jeffers. Data logging in the biological sciences. Journal of Applied Statistics, 12(2):107–112, 1985. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cox:1985:OQS

[261] Trevor F. Cox, John P. Birchall, and Henry Wong. Optimising the queuing system for an ear, nose and throat outpatient clinic. Journal of Applied Statistics, 12(2):113–126, 1985. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gilliland:1985:MCL

[262] Dennis C. Gilliland. On the MacMahon cube law for election results. Journal of Applied Statistics, 12(2):127–135, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shoesmith:1985:BDD

[263] Eddie Shoesmith. Birth-date distribution and statistical significance. Journal of Applied Statistics, 12(2):137– 145, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gordon:1985:SSB

[264] T. F. Gordon. Statistical simulations on the BBC microcomputer: significance tests of the pseudo-random number generator. Journal of Applied Statistics, 12(2):147–155, 1985. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rowell:1985:CAA

[265] J. G. Rowell and D. E. Walters. A critical appraisal of the analysis of variance for two-way classifications. Journal of Applied Statistics, 12(2):157–169, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marcus:1985:TSP

[266] Ruth Marcus. A truncated sequential procedure for estimating the means of insect populations. Journal of Applied Statistics, 12(2):171–176, 1985. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilkie:1985:SEF

[267] D. Wilkie. Simple exponential fitting. Journal of Applied Statistics, 12(2):

177–179, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bryant:1985:ORC

[268] Jim Bryant. Operational research changes: tales from a parallel universe. Journal of Applied Statistics, 12(2):181–187, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRe

[269] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):189–190, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRf

[270] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):189–190, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRg

[271] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):190–191, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRh

[272] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):191–192, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRi

[273] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):191–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRj

[274] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):192–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRk

[275] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):192–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1985:BRl

[276] Anonymous. Book reviews. Journal of Applied Statistics, 12(2):193–??, 1985. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1986:LRP

[277] A. F. Bissell. Not-so-linear regression (part 1). Journal of Applied Statistics, 13(1):3–11, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bera:1986:AFP

[278] Anil K. Bera and Colin R. Mckenzie. Alternative forms and properties of the score test. Journal of Applied Statistics, 13(1):13–25, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fresen:1986:EPP

[279] J. L. Fresen and J. W. Fresen. Estimating the parameter in the Pauling equation. Journal of Applied Statistics, 13(1):27–37, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Madry:1986:ACA

[280] Wieslaw Madry and Walerian Piotrowski. Analysis of the combining ability of two parental groups in incomplete diallel crosses. Journal of Applied Statistics, 13(1):39–48, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ogallo:1986:SMR

[281] Laban A. J. Ogallo. Stochastic modelling of regional annual rainfall anomalies in East Africa. Journal of Applied Statistics, 13(1):49–56, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilkie:1986:OCF

[282] D. Wilkie. Orthogonal comparisons for Fourier analysis within mixed level factorial experiments. Journal of Applied Statistics, 13(1):57–66, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oxley:1986:BCP

[283] Leslie T. Oxley. Box–Cox power transformations and the demand for money functions in France, Germany, Italy and The Netherlands. Journal of Applied Statistics, 13(1):67–75, 1986. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ryder:1986:NTG

[284] K. Ryder. Numerical tables: guidelines for presentation. Journal of Applied Statistics, 13(1):77–88, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Heinrich:1986:AED

[285] I. Heinrich, B. Sweetman, and J. A. D. Anderson. Application of an empirically derived classification of ‘nonspecific’ low back pain. Journal of Applied Statistics, 13(1):89–96, 1986. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Walters:1986:NEA

[286] David Eurof Walters. A note on the economical application of the ‘exact’ test for a contingency table. Journal of Applied Statistics, 13(1):97–104, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilkie:1986:PFF

[287] D. Wilkie. Polynomial and Fourier fits to diurnal variations in ambient temperatures in Bechuanaland. Journal of Applied Statistics, 13(1):105–110, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRa

[288] Anonymous. Book reviews. Journal of Applied Statistics, 13(1):113–114, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRb

[289] Anonymous. Book reviews. Journal of Applied Statistics, 13(1):114–115, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:A

[290] Anonymous. Announcement. Journal of Applied Statistics, 13(1):111–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRc

[291] Anonymous. Book reviews. Journal of Applied Statistics, 13(1):113–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRd

[292] Anonymous. Book reviews. Journal of Applied Statistics, 13(1):114–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRe

[293] Anonymous. Book reviews. Journal of Applied Statistics, 13(1):116–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRf

[294] Anonymous. Book reviews. Journal of Applied Statistics, 13(1):117–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRg

[295] Anonymous. Book reviews. Journal of Applied Statistics, 13(1):117–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harrison:1986:AVC

[296] D. Harrison, G. K. Kanji, and R. J. Gadsden. Analysis of variance for circular data. Journal of Applied Statistics, 13(2):123–138, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearce:1986:DDA

[297] S. C. Pearce. Defective data in the analysis of variance. Journal of Applied Statistics, 13(2):139–147, 1986. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wonnacott:1986:BCH

[298] Thomas H. Wonnacott. Bayesian and Classical hypothesis testing. Journal of Applied Statistics, 13(2):149–157, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Novotny:1986:MSU

[299] Timothy J. Novotny and Lyman L. Mcdonald. Model selection using discriminant analysis. Journal of Applied Statistics, 13(2):159–165, 1986. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marcus:1986:MSN

[300] Ruth Marcus and B. Raccah. Model for the spread of non-persistent virus diseases. Journal of Applied Statistics, 13(2):167–175, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hog:1986:EPD

[301] Esben Høg. Estimation of a pure death process with an application to withdrawals and resumptions of automobile insurances. Journal of Applied Statistics, 13(2):177–186, 1986. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Galpin:1986:VSS

[302] Jacqueline S. Galpin and Douglas M. Hawkins. Variable subset selection for optimal regression prediction at a specified point. Journal of Applied Statistics, 13(2):187–198, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Herbert:1986:DAS

[303] John H. Herbert. Data analysis of sales of natural gas to households in the United States. Journal of Applied Statistics, 13(2):199–211, 1986. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mcculloch:1986:DTH

[304] Brian Mcculloch and Tony Van Zijl. Direct test of Harville’s multi-entry competitions model on race-track betting data. Journal of Applied Statistics, 13(2):213–220, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oliver:1986:DLS

[305] F. R. Oliver. Direct least squares estimation of Phillips’s function. Journal of Applied Statistics, 13(2):221– 225, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRh

[306] Anonymous. Book reviews. Journal of Applied Statistics, 13(2):227–228, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRi

[307] Anonymous. Book reviews. Journal of Applied Statistics, 13(2):229–230, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRj

[308] Anonymous. Book reviews. Journal of Applied Statistics, 13(2):228–??, 1986.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRk

[309] Anonymous. Book reviews. Journal of Applied Statistics, 13(2):229–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRl

[310] Anonymous. Book reviews. Journal of Applied Statistics, 13(2):230–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRm

[311] Anonymous. Book reviews. Journal of Applied Statistics, 13(2):231–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:BRn

[312] Anonymous. Book reviews. Journal of Applied Statistics, 13(2):231–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1986:E

[313] Anonymous. Editorial. Journal of Applied Statistics, 13(2):121–??, 1986. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1987:LRPa

[314] A. F. Bissell. Not-so-linear regression (part 2). Journal of Applied Statistics, 14(1):3–13, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marcus:1987:DCT

[315] Ruth Marcus, R. Salomon, and M. Bar-Joseph. The distribution of citrus tristeza virus in citrus groves in Israel. Journal of Applied Statistics, 14(1):15– 21, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Okafor:1987:MLE

[316] R. O. Okafor. Maximum likelihood estimation from incomplete data. Journal of Applied Statistics, 14(1):23–33, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hahn:1987:ACC

[317] Gerald J. Hahn and Michael B. Cockrum. Adapting control charts to meet practical needs: a chemical processing application. Journal of Applied Statistics, 14(1):35–52, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearce:1987:SSA

[318] S. C. Pearce. Some simple algorithms for the analysis of variance with nonorthogonal block data. Journal of Applied Statistics, 14(1):53–59, 1987. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stefani:1987:ASM

[319] Raymond T. Stefani. Applications of statistical methods to American football. Journal of Applied Statistics, 14 (1):61–73, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deligonul:1987:DBC

[320] Z. Seyda Deligonul and A. Erhan Mergen. Dependence bias in conventional p-charts and its correction with an approximate lot quality distribution.

Journal of Applied Statistics, 14(1):75– 81, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilkie:1987:AFE

[321] D. Wilkie. Analysis of factorial experiments and least squares polynomial fitting by the method of orthogonal polynomials for any spacing of the levels of the independent variables. Journal of Applied Statistics, 14(1):83–89, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:1987:NNC

[322] M. Singh. A non-normal class of distribution function for dose-binary response curve*. Journal of Applied Statistics, 14(1):91–97, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRa

[323] Anonymous. Book reviews. Journal of Applied Statistics, 14(1):99–100, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:A

[324] Anonymous. Announcement. Journal of Applied Statistics, 14(1):102– 103, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRb

[325] Anonymous. Book reviews. Journal of Applied Statistics, 14(1):100–??, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRc

[326] Anonymous. Book reviews. Journal of Applied Statistics, 14(1):101–??, 1987.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRd

[327] Anonymous. Book reviews. Journal of Applied Statistics, 14(1):101–??, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barnard:1987:WII

[328] G. A. Barnard. What influences influence? Journal of Applied Statistics, 14(2):107–116, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cook:1987:IA

[329] R. Dennis Cook. Influence assessment. Journal of Applied Statistics, 14(2):117–131, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Geisser:1987:IOD

[330] Seymour Geisser. Influential observations, diagnostics and discovery tests. Journal of Applied Statistics, 14(2): 133–142, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Smith:1987:IOS

[331] T. M. F. Smith. Influential observations in survey sampling. Journal of Applied Statistics, 14(2):143–152, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Greenacre:1987:IAP

[332] M. J. Greenacre. Influential analysis and presentation of survey data. Journal of Applied Statistics, 14(2):153– 164, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shewry:1987:MES

[333] M. C. Shewry and H. P. Wynn. Maximum entropy sampling. Journal of Applied Statistics, 14(2):165–170, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Munoz-Garcia:1987:DOM

[334] J. Munoz-Garcia, J. L. Moreno-Rebollo, and A. Pascual-Acosta. Detecting outliers in s-multinomial distributions through adjusted residuals. Journal of Applied Statistics, 14(2): 171–176, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balasooriya:1987:ECS

[335] U. Balasooriya, Y. K. Tse, and Y. S. Liew. An empirical comparison of some statistics for identifying outliers and influential observations in linear regression models. Journal of Applied Statistics, 14(2):177–184, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Greenacre:1987:D

[336] Michael Greenacre. I.D.A. Journal of Applied Statistics, 14(2):185–??, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1987:LRPb

[337] A. F. Bissell. Not-so-linear regression (part 3). Journal of Applied Statistics, 14(3):189–194, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wonnacott:1987:CIH

[338] Thomas Wonnacott. Confidence intervals or hypothesis tests? Journal of Applied Statistics, 14(3):195–201, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

VanDerLaan:1987:LMP

[339] Bob Van Der Laan and Johan Koerts. A logit model for the probability of having non-zero expenses for medical services during a year. Journal of Applied Statistics, 14(3):203–217, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kenett:1987:SET

[340] Ron S. Kenett and David M. Steinberg. Some experiences teaching factorial design in introductory statistics courses. Journal of Applied Statistics, 14(3):219–227, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Farley:1987:MAT

[341] T. M. M. Farley and I. T. Jolliffe. Measures of association for two-way contingency tables: two case studies. Journal of Applied Statistics, 14(3):229– 237, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:1987:FME

[342] W. K. Li, H. Wong, and Y. V. Hui. Fractional modelling on emergency room attendance data. Journal of Applied Statistics, 14(3):239– 249, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bouhaddou:1987:PCA

[343] O. Bouhaddou, C. H. Obled, and Tuan Pham Dinh. Principal component analysis and interpolation of stochastic processes: methods and simulation. Journal of Applied Statistics, 14(3):251–267, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kinsella:1987:QAL

[344] A. Kinsella and A. Dunne. Quantitative assessment of a lung tumour treatment using a multiple regression spline model. Journal of Applied Statistics, 14(3):269–273, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rao:1987:ESS

[345] U. L. Gouranga Rao and Alan Yuk-Pang Tam. An empirical study of selection and estimation of alternative models of the Lorenz curve. Journal of Applied Statistics, 14(3):275–280, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kmietowicz:1987:SPC

[346] Z. W. Kmietowicz and H. M. K. Rwamasaka. Some problems in the calculation of indices of industrial production in developing countries: the case of Uganda, 1963–71. Journal of Applied Statistics, 14(3):281–295, 1987. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRe

[347] Anonymous. Book reviews. Journal of Applied Statistics, 14(3):298–299, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRf

[348] Anonymous. Book reviews. Journal of Applied Statistics, 14(3):299–300, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRg

[349] Anonymous. Book reviews. Journal of Applied Statistics, 14(3):297–??, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRh

[350] Anonymous. Book reviews. Journal of Applied Statistics, 14(3):297–??, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRi

[351] Anonymous. Book reviews. Journal of Applied Statistics, 14(3):298–??, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1987:BRj

[352] Anonymous. Book reviews. Journal of Applied Statistics, 14(3):300–??, 1987. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jeffers:1988:BAP

[353] J. N. R. Jeffers. Biometrics: achievement and potential. Journal of Applied Statistics, 15(1):5–17, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chatfield:1988:WBM

[354] C. Chatfield. What is the ‘best’ method of forecasting? Journal of Applied Statistics, 15(1):19–38, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pack:1988:IOP

[355] P. Pack, I. T. Jolliffe, and B. J. T. Morgan. Influential observations in principal component analysis: a case study. Journal of Applied Statistics, 15(1):39– 52, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:1988:SAI

[356] M. Singh and B. Gilliver. Statistical analysis of intercropping data using a correlated error structure. Journal of Applied Statistics, 15(1):53–61, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonett:1988:END

[357] Douglas G. Bonett. Estimating the number of defects under imperfect inspection. Journal of Applied Statistics, 15(1):63–67, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rudolph:1988:EUP

[358] Patricia M. Rudolph and Marvin Karson. The effect of unequal priors and unequal misclassification costs on MDA. Journal of Applied Statistics, 15(1):69–83, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roller:1988:AVI

[359] William F. Roller. Adjusted variables: an important tool for teaching regression in the applications curricula. Journal of Applied Statistics, 15(1):85–95, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1988:CCL

[360] A. F. Bissell. Control chart limits for attributes and events. Journal of Applied Statistics, 15(1):97–105, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huitson:1988:BRa

[361] A. Huitson. Book reviews. Journal of Applied Statistics, 15(1):107–108, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kirby:1988:BR

[362] Simon Kirby. Book reviews. Journal of Applied Statistics, 15(1):108–109, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huitson:1988:BRb

[363] A. Huitson. Book reviews. Journal of Applied Statistics, 15(1):109–110, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1988:ASI

[364] Anonymous. Announcement of special issue: ‘statistical methods in digital image processing’. Journal of Applied Statistics, 15(1):111–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Castle:1988:BR

[365] Richard Castle. Book reviews. Journal of Applied Statistics, 15(1):107–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:1988:BR

[366] A. Watson. Book reviews. Journal of Applied Statistics, 15(1):109–??, 1988.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zielinski:1988:BR

[367] Andrzej Zielinski. Book reviews. Journal of Applied Statistics, 15(1):110–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1988:C

[368] Anonymous. Corrigenda. Journal of Applied Statistics, 15(1):111–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1988:EFY

[369] Anonymous. Editorial: Fifteen years of being biased. Journal of Applied Statistics, 15(1):3–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:1988:DDA

[370] K. V. Mardia. Directional data analysis: an overview. Journal of Applied Statistics, 15(2):115–122, 1988. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:1988:LDH

[371] G. S. Watson. The Langevin distribution on high dimensional spheres. Journal of Applied Statistics, 15(2):123– 130, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arsham:1988:KVM

[372] H. Arsham. Kuiper’s P -value as a measuring tool and decision procedure for the goodness-of-fit test. Journal of Applied Statistics, 15(2):131–135, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jupp:1988:RDD

[373] P. E. Jupp. Residuals for directional data. Journal of Applied Statistics, 15(2):137–147, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bagchi:1988:TCM

[374] Parthasarathy Bagchi and Irwin Guttman. Theoretical considerations of the multivariate von Mises–Fisher distribution. Journal of Applied Statistics, 15(2):149–169, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prentice:1988:USP

[375] Michael J. Prentice. Using sample principal axes to estimate small circles and ellipses on the sphere. Journal of Applied Statistics, 15(2):171–177, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Smeeton:1988:DAC

[376] Nigel Smeeton and Greg Wilkinson. The detection of annual clusters in individual patterns of parasuicide. Journal of Applied Statistics, 15(2):179– 182, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haskey:1988:ROA

[377] John C. Haskey. The relative orientation of addresses of spouses before their marriage: an analysis of circular data. Journal of Applied Statistics, 15(2):183–195, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harrison:1988:DAV

[378] D. Harrison and G. K. Kanji. The development of analysis of variance for circular data. Journal of Applied Statistics, 15(2):197–223, 1988. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lawson:1988:TST

[379] Andrew Lawson. On tests for spatial trend in a non-homogeneous Poisson process. Journal of Applied Statistics, 15(2):225–234, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Clark:1988:ESA

[380] R. M. Clark. An evaluation by simulation of alternative estimators for the marginal Fisher distribution. Journal of Applied Statistics, 15(2):235– 246, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kent:1988:MLE

[381] John T. Kent and David E. Tyler. Maximum likelihood estimation for the wrapped Cauchy distribution. Journal of Applied Statistics, 15(2):247– 254, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lawson:1988:FMD

[382] Andrew Lawson. Fitting the von Mises distribution using glim. Journal of Applied Statistics, 15(2):255–260, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miller:1988:SR

[383] G. Miller. Software review. Journal of Applied Statistics, 15(2):263–264, 1988.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bishop:1988:BRa

[384] C. Bishop. Book reviews. Journal of Applied Statistics, 15(2):261–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bishop:1988:BRb

[385] C. Bishop. Book reviews. Journal of Applied Statistics, 15(2):261–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huitson:1988:BRc

[386] A. Huitson. Book reviews. Journal of Applied Statistics, 15(2):262–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patel:1988:BR

[387] Sumit Patel. Book reviews. Journal of Applied Statistics, 15(2):262–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hiutson:1988:BR

[388] Alan Hiutson. Book reviews. Journal of Applied Statistics, 15(2):263–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:1988:FNL

[389] M. Singh and G. K. Kanji. Fitting a non-linear model with errors in both variables and its application. Journal of Applied Statistics, 15(3):267–274, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barry:1988:ISM

[390] J. T. Barry. An investigation of statistical matching. Journal of Applied Statistics, 15(3):275–283, 1988. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chuang-Stein:1988:SSL

[391] Christy Chuang-Stein and Steven F. Francom. Some simple log-linear models for the analysis of ordinal response data. Journal of Applied Statistics, 15(3):285–294, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Osanaiye:1988:EII

[392] P. A. Osanaiye and S. A. Alebiosu. Effects of industrial inspection errors on some plans that utilise the surrounding lot information. Journal of Applied Statistics, 15(3):295–304, 1988. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1988:SDT

[393] A. F. Bissell and R. J. Williamson. Successive difference tests-theory and interpretation. Journal of Applied Statistics, 15(3):305–323, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Williams:1988:ISE

[394] M. K. Williams. An intelligent system for experimental design-some ideas. Journal of Applied Statistics, 15(3): 325–333, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Johnsson:1988:BMT

[395] Tommy Johnsson. Bootstrap multiple test procedure. Journal of Applied Statistics, 15(3):335–339, 1988. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Herbert:1988:RVE

[396] John H. Herbert and Phillip S. Kott. Robust variance estimation in linear regression. Journal of Applied Statistics, 15(3):341–345, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baxter:1988:DBP

[397] Mike Baxter and Richard Stevenson. Discriminating between the Poisson and negative binomial distributions: an application to goal scoring in Association football. Journal of Applied Statistics, 15(3):347–354, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chowdhury:1988:AOL

[398] A. Chowdhury and U. Balasooriya. Atypical observations in linear models. Journal of Applied Statistics, 15 (3):355–362, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tse:1988:ALC

[399] Y. K. Tse. Assessing Lund’s critical values for testing for outliers in linear regression models. Journal of Applied Statistics, 15(3):363–366, 1988. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rashed:1988:SUR

[400] A. F. Rashed and M. Metwally. Simulated and unbiased reliability sampling plans. Journal of Applied Statistics, 15(3):367–379, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1988:BRa

[401] Anonymous. Book reviews. Journal of Applied Statistics, 15(3):381–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1988:BRb

[402] Anonymous. Book reviews. Journal of Applied Statistics, 15(3):381–??, 1988. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearce:1989:SCE

[403] S. C. Pearce. The size of a comparative experiment. Journal of Applied Statistics, 16(1):3–6, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1989:IMS

[404] A. F. Bissell. Interpreting mean squares in saturated fractional designs. Journal of Applied Statistics, 16(1):7–18, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Preece:1989:FES

[405] D. A. Preece. Factorial experimentation in second-order Latin cubes. Journal of Applied Statistics, 16(1):19–24, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hossain:1989:DIO

[406] A. Hossain and D. N. Naik. Detection of influential observations in multivariate regression. Journal of Applied Statistics, 16(1):25–37, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huitson:1989:PPA

[407] Alan Huitson. Problems with Procrustes analysis. Journal of Applied Statistics, 16(1):39–45, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeWet:1989:TLO

[408] Tertius De Wet, John Fresen, and Jill Fresen. The test for logical order in English language testing. Journal of Applied Statistics, 16(1):47–54, 1989. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haines:1989:AMB

[409] L. M. Haines, W. P. Munoz, and C. J. Van Gelderen. ARIMA modelling of birth data. Journal of Applied Statistics, 16(1):55–67, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tai:1989:ITH

[410] J. J. Tai, Keh-Min Chiou, Chun-Jean Lee, Lung-An Li, Po-Chang Lee, and Shan-Yie Liu. Index typing of HLA serological data. Journal of Applied Statistics, 16(1):69–75, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Subramani:1989:NIL

[411] J. Subramani and K. N. Ponnuswamy. A non-iterative least squares estimation of missing values in experimental designs. Journal of Applied Statistics, 16(1):77–86, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soundararajan:1989:DMD

[412] V. Soundararajan and R. Vijayaraghavan. On designing multiple deferred state sampling (MDS-1 (0, 2)) plans involving minimum risks. Journal of Applied Statistics, 16(1):87–94, 1989. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berny:1989:NCD

[413] J. Berny. New concepts in deterministic growth curve forecasting. Journal of Applied Statistics, 16(1):95–120, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:Aa

[414] Anonymous. Announcement. Journal of Applied Statistics, 16(1):121– 122, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:1989:MMB

[415] K. V. Mardia. Markov models and Bayesian methods in image analysis. Journal of Applied Statistics, 16(2): 125–130, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See erratum [570].

Dubes:1989:RFM

[416] Richard C. Dubes and Anil K. Jain. Random field models in image analysis. Journal of Applied Statistics, 16 (2):131–164, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brillinger:1989:CNI

[417] David R. Brillinger, Kenneth H. Downing, Robert M. Glaeser, and Guy Perkins. Combining noisy images of small crystalline domains in high resolution electron microscopy. Journal of Applied Statistics, 16(2):165–175, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jones:1989:OSD

[418] M. C. Jones and B. W. Silverman. An orthogonal series density estimation approach to reconstructing positron emission tomography images. Journal of Applied Statistics, 16(2):177– 191, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Molina:1989:USM

[419] R. Molina and B. D. Ripley. Using spatial models as priors in astronomical image analysis. Journal of Applied Statistics, 16(2):193–206, 1989. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Grenander:1989:TAI

[420] Ulf Grenander and Daniel Macrae Keenan. Towards automated image understanding. Journal of Applied Statistics, 16(2):207–221, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barnett:1989:USS

[421] Glen Barnett, Susan Crowe, Malcolm Hudson, Pui-Lam Leung, Khairil Notodiputro, Richard Proudfoot, and John Sims. The use of small scale prototypes in image reconstruction from projections. Journal of Applied Statistics, 16(2):223–242, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Frieden:1989:SAI

[422] B. Roy Frieden. Stochastic approaches to inversion problems in optics. Journal of Applied Statistics, 16(2):243– 266, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qian:1989:UGM

[423] W. Qian and D. M. Titterington. On the use of Gibbs Markov chain models in the analysis of images based on second-order pairwise interactive distributions. Journal of Applied Statistics, 16(2):267–281, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pieczynski:1989:ECR

[424] Wojciech Pieczynski. Estimation of context in random fields. Journal of Applied Statistics, 16(2):283–290, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:Ab

[425] Anonymous. Abstracts. Journal of Applied Statistics, 16(2):291–292, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:ANR

[426] Anonymous. Abstracts news: rankings for research in statistical theory. Journal of Applied Statistics, 16(2):293– 294, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRa

[427] Anonymous. Book reviews. Journal of Applied Statistics, 16(2):295–296, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRb

[428] Anonymous. Book review. Journal of Applied Statistics, 16(2):298–299, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRc

[429] Anonymous. Book review. Journal of Applied Statistics, 16(2):299–300, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRd

[430] Anonymous. Book review. Journal of Applied Statistics, 16(2):299–300, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRe

[431] Anonymous. Book review. Journal of Applied Statistics, 16(2):300–301, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRf

[432] Anonymous. Book review. Journal of Applied Statistics, 16(2):301–302, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRg

[433] Anonymous. Book review. Journal of Applied Statistics, 16(2):295–??, 1989.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRh

[434] Anonymous. Book review. Journal of Applied Statistics, 16(2):296–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRi

[435] Anonymous. Book review. Journal of Applied Statistics, 16(2):297–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRj

[436] Anonymous. Book reivew. Journal of Applied Statistics, 16(2):297–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRk

[437] Anonymous. Book review. Journal of Applied Statistics, 16(2):298–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRl

[438] Anonymous. Book review. Journal of Applied Statistics, 16(2):299–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRm

[439] Anonymous. Book review. Journal of Applied Statistics, 16(2):301–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Erto:1989:RAR

[440] P. Erto. Reliability assessments by repair shops via maintenance data. Journal of Applied Statistics, 16(3):303– 313, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). This is the revised version (containing further applicative examples) of the paper published in the proceedings of the Reliability ‘87 Conference, 14–16 April, Birmingham, U.K., as anticipated during the oral presentation given then.

Marcus:1989:MSC

[441] Ruth Marcus, Hovav Talpaz, and Moshe Bar-Joseph. A model for the spread and control of citrus tristeza virus disease. Journal of Applied Statistics, 16(3):315–320, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chuang-Stein:1989:ETS

[442] Christy Chuang-Stein and R. John Weaver. An evaluation of therapeutic synergism using linear logit models. Journal of Applied Statistics, 16 (3):321–329, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oludhe:1989:SCS

[443] C. Oludhe and L. Ogallo. Statistical characteristics of the surface wind speeds over Kenya. Journal of Applied Statistics, 16(3):331–344, 1989. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Veevers:1989:EAM

[444] A. Veevers and M. C. K. Tweedie. Examples of the application of matrix factors of a design matrix. Journal of Applied Statistics, 16(3):345–354, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bangdiwala:1989:TCS

[445] Shrikant I. Bangdiwala. The teaching of the concepts of statistical tests of hypotheses to non-statisticians. Journal of Applied Statistics, 16(3):355– 361, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:1989:ELE

[446] M. Singh. Estimation of land equivalent ratios. Journal of Applied Statistics, 16(3):363–371, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Srinivasan:1989:AIM

[447] M. R. Srinivasan. An algorithm for inverting matrices whose elements are polynomials. Journal of Applied Statistics, 16(3):373–375, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ojo:1989:ASP

[448] M. O. Ojo. Analysis of some prison data. Journal of Applied Statistics, 16(3):377–383, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baczkowski:1989:TSI

[449] A. J. Baczkowski and K. V. Mardia. A test of spatial independence. Journal of Applied Statistics, 16(3):385–389, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dagpunar:1989:CPP

[450] J. S. Dagpunar. A compact and portable Poisson random variate generator. Journal of Applied Statistics, 16(3):391–393, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Besag:1989:DIP

[451] Julian Besag. Digital image processing: Towards Bayesian image analysis. Journal of Applied Statistics, 16 (3):395–407, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Clifford:1989:RPI

[452] Peter Clifford and R. D. Middleton. Reconstruction of polygonal images. Journal of Applied Statistics, 16(3):409– 422, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:1989:DIP

[453] K. V. Mardia and T. J. Hainsworth. Digital image processing: Statistical aspects of moment invariants in image analysis. Journal of Applied Statistics, 16(3):423–435, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:LE

[454] Anonymous. Letter to the Editor. Journal of Applied Statistics, 16(3): 437–438, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1989:BRn

[455] Anonymous. Book review. Journal of Applied Statistics, 16(3):439–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lawson:1989:FMD

[456] Andrew Lawson. Fitting the von Mises distribution using GLIM. Journal of Applied Statistics, 16(3):442–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lawson:1989:TST

[457] Andrew Lawson. On tests for spatial trend in a non-homogeneous Poisson process. Journal of Applied Statistics, 16(3):441–??, 1989. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ross:1990:IVF

[458] G. J. S. Ross. Incomplete variance functions. Journal of Applied Statistics, 17(1):3–8, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kmietowicz:1990:SLO

[459] Z. W. Kmietowicz. The significance of the lead in opinion polls. Journal of Applied Statistics, 17(1):9–30, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tricker:1990:ERSa

[460] A. R. Tricker. The effect of rounding on the significance level of certain normal test statistics. Journal of Applied Statistics, 17(1):31–38, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ohtani:1990:UBC

[461] Kazuhiro Ohtani. On the use of the Box–Cox transformation in a linear regression with a lagged dependent variable. Journal of Applied Statistics, 17(1):39–47, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McEntee:1990:ERP

[462] Edward F. McEntee. An economic rectification plan for a process subject to a random shift into a persistent defective state. Journal of Applied Statistics, 17(1):49–54, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balakrishnan:1990:MLE

[463] N. Balakrishnan. On the maximum likelihood estimation of the location and scale parameters of exponential distribution based on multiply Type II censored samples. Journal of Applied Statistics, 17(1):55–61, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cox:1990:GLR

[464] Trevor F. Cox and Khalifa H. Jaber. The generalized likelihood ratio test for the Behrens–Fisher problem. Journal of Applied Statistics, 17(1):63–71, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Okafor:1990:BAE

[465] Raymond Okafor. A biweight approach to estimate currency exchange rate: the Nigerian example. Journal of Applied Statistics, 17(1):73–82, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soundararajan:1990:CSMa

[466] V. Soundararajan and S. Devaraj Arumainayagam. Construction and selection of modified quick switching systems. Journal of Applied Statistics, 17(1):83–114, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilson:1990:MMC

[467] Jeffrey R. Wilson and David L. Turner. A multinomial model for cross classified data with a measure of dependency. Journal of Applied Statistics, 17(1):115–124, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Agarwal:1990:UAB

[468] S. C. Agarwal and M. N. Das. Use of n-ary block designs in dialled crosses evaluation. Journal of Applied Statistics, 17(1):125–131, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gill:1990:IPL

[469] P. S. Gill, M. L. Tiku, and David C. Vaughan. Inference problems in life testing under multivariate normality. Journal of Applied Statistics, 17(1): 133–147, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Najjar:1990:RTN

[470] S. N. Al-Najjar. Reparametrization and transformation of a non-linear model relating plant density and crop yield. Journal of Applied Statistics, 17(1):149–158, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baczkowski:1990:PBU

[471] A. J. Baczkowski and K. V. Mardia. Prediction based upon maximizing squared correlation for stationary processes and simple kriging. Journal of Applied Statistics, 17(1):159–164, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dagpunar:1990:SMD

[472] J. S. Dagpunar. Sampling from the von Mises distribution via a comparison of random numbers. Journal of Applied Statistics, 17(1):165–168, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:1990:BRa

[473] D. J. Hand. Book reviews. Journal of Applied Statistics, 17(1):172–173, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Blinkhorn:1990:BR

[474] S. F. Blinkhorn. Book reviews. Journal of Applied Statistics, 17(1):173– 174, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huitson:1990:BRa

[475] Alan Huitson. Book review. Journal of Applied Statistics, 17(1):171–??, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reid:1990:BR

[476] Alan Reid. Book reviews. Journal of Applied Statistics, 17(1):172–??, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:1990:BRb

[477] D. J. Hand. Book reviews. Journal of Applied Statistics, 17(1):173–??, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huitson:1990:BRb

[478] A. Huitson. Book reviews. Journal of Applied Statistics, 17(1):174–??, 1990.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1990:O

[479] Anonymous. Obituary. Journal of Applied Statistics, 17(1):169–??, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Logothetis:1990:MAA

[480] N. Logothetis. Minute-accumulating analysis and life testing. Journal of Applied Statistics, 17(2):175–185, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tong:1990:CTS

[481] H. Tong and P. Dabas. Cluster of time series models: an example. Journal of Applied Statistics, 17(2):187–198, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Clarke:1990:ELF

[482] Ralp T. Clarke, J. S. Welton, and M. Ladle. Estimation of larval feeding rates using normal linear models involving censored and truncated observations. Journal of Applied Statistics, 17(2):199–209, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1990:EPT

[483] A. F. Bissell. The eccentrics of political twittering. Journal of Applied Statistics, 17(2):211–217, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tricker:1990:ERP

[484] A. R. Tricker. The effect of rounding on the power level of certain normal test statistics. Journal of Applied Statistics, 17(2):219–228, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baxter:1990:PCC

[485] M. J. Baxter, H. E. M. Cool, and M. P. Heyworth. Principal component and correspondence analysis of compositional data: some similarities. Journal of Applied Statistics, 17(2):229– 235, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalirajan:1990:ERM

[486] K. P. Kalirajan. On the estimation of a regression model with fixed and random coefficients. Journal of Applied Statistics, 17(2):237–244, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Naik:1990:PIG

[487] D. N. Naik. Prediction intervals for growth curves. Journal of Applied Statistics, 17(2):245–254, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roller:1990:POA

[488] William F. Roller. Partial-order analysis of fiscal capacity of states. Journal of Applied Statistics, 17(2):255– 262, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kapoor:1990:ACT

[489] S. Kapoor and G. K. Kanji. Application of the characterization theory to the mixture model. Journal of Applied Statistics, 17(2):263–270, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jones:1990:LNM

[490] P. N. Jones and G. J. Mclachlan. Laplace-normal mixtures fitted to wind shear data. Journal of Applied Statistics, 17(2):271–276, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:1990:PTI

[491] Geoffrey S. Watson. Permutation tests for the independence of a scalar and a unit vector. Journal of Applied Statistics, 17(2):277–281, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Everitt:1990:CMM

[492] B. S. Everitt and C. Merette. The clustering of mixed-mode data: a comparison of possible approaches. Journal of Applied Statistics, 17(3):283–297, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ross:1990:EPN

[493] G. J. S. Ross. Estimation problems of non-linear functional relationships. Journal of Applied Statistics, 17(3):299–306, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Digby:1990:MAW

[494] P. G. N. Digby and K. L. Moore. Multivariate analysis of within-group variability. Journal of Applied Statistics, 17(3):307–312, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hadi:1990:TGD

[495] Ali S. Hadi. Two graphical displays for the detection of potentially influential subsets in regression. Journal of Applied Statistics, 17(3):313–327, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tricker:1990:ERSb

[496] A. R. Tricker. The effect of rounding on the significance level and power of certain test statistics for non-normal data. Journal of Applied Statistics, 17(3):329–340, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bagchi:1990:SOD

[497] Parthasarathy Bagchi and Irwin Guttman. Spuriosity and outliers in directional data. Journal of Applied Statistics, 17(3):341–350, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ipinyomi:1990:RRC

[498] R. A. Ipinyomi. Resolvable row-column designs. Journal of Applied Statistics, 17(3):351–355, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alemayehu:1990:UBC

[499] Demissie Alemayehu and Kjell Doksum. Using the bootstrap in correlation analysis, with application to a longitudinal data set. Journal of Applied Statistics, 17(3):357–368, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gribbin:1990:MDS

[500] J. O. Gribbin and S. I. McClean. Modelling the duration of spells of withdrawal from a female dominated labour force. Journal of Applied Statistics, 17(3):369–381, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonett:1990:TRN

[501] Douglas G. Bonett and J. Arthur Woodward. Testing residual normality in the ANOVA model. Journal of Applied Statistics, 17(3):383–387, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hunt:1990:SIM

[502] D. N. Hunt and M. J. Bell. Semiiterative missing value estimation. Journal of Applied Statistics, 17(3): 389–396, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soundararajan:1990:CSMb

[503] V. Soundararajan and R. Vijayaraghavan. Construction and selection of multiple dependent (deferred) state sampling plan. Journal of Applied Statistics, 17(3):397–409, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Murty:1990:BPP

[504] Vedula N. Murty. Binomial and Poisson probabilities with an HP-28S Scientific Advanced Calculator. Journal of Applied Statistics, 17(3):411–415, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mansouri:1990:CRT

[505] H. Mansouri and Z. Govindarajulu. A class of rank tests for interaction in two-way layouts. Journal of Applied Statistics, 17(3):417–426, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Govindaraju:1990:SMD

[506] K. Govindaraju and K. Subramani. Selection of multiple deferred state MDS-1 sampling plans for given acceptable quality level and limiting quality level involving minimum risks. Journal of Applied Statistics, 17(3):427–434, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Herbert:1990:ENC

[507] John H. Herbert and Khoan Dinh. An empirical note concerning systematic measurement errors in regression analysis. Journal of Applied Statistics, 17(3):435–441, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:1990:LLL

[508] K. V. Mardia. On longevity and life-line. Journal of Applied Statistics, 17(3):443–448, 1990. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arnold:1990:PPA

[509] Gillian M. Arnold. Problems with Procrustes analysis. Journal of Applied Statistics, 17(3):449–451, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See comments

Huitson:1990:CPP

[510] Alan Huitson. Comments on: ‘Problems with Procrustes analysis’ by Gillian M. Arnold. Journal of Applied Statistics, 17(3):451–452, 1990. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [509].

Hand:1991:ESS

[511] D. J. Hand. Expert system and statistics. Journal of Applied Statistics, 18 (1):1–2, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lawson:1991:USK

[512] K. W. Lawson and P. A. Golder. The use of semantic knowledge to enhance statistical software. Journal of Applied Statistics, 18(1):3–21, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cox:1991:KBR

[513] Louis Anthony Cox, Jr. Knowledgebased resolution of conflicting expert opinions. Journal of Applied Statistics, 18(1):23–34, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tung:1991:ESS

[514] S. T. Young Tung and J. H. Schuenemeyer. An expert system for statistical consulting. Journal of Applied Statistics, 18(1):35–47, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Biggs:1991:MCM

[515] David Biggs, Barry De Ville, and Ed Suen. A method of choosing multiway partitions for classification and decision trees. Journal of Applied Statistics, 18(1):49–62, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Almond:1991:BBG

[516] Russell Almond. Building blocks for graphical belief models. Journal of Applied Statistics, 18(1):63–76, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

VanDenBerg:1991:ISE

[517] Gerda M. Van Den Berg. Ideas of statistical experts on the choice of an analysis method. Journal of Applied Statistics, 18(1):77–86, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hoschka:1991:ECI

[518] Peter Hoschka and Willi Klösgen. EXPLORA: content interpretation of data. Journal of Applied Statistics, 18 (1):87–97, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Visser:1991:KAS

[519] Ronald Visser and Nelleke Slooff. Knowledge acquisition for a supportive user interface for correspondence analysis. Journal of Applied Statistics, 18(1):99–106, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Molenaar:1991:ACM

[520] Ivo W. Molenaar and Paul Passchier. The Agree Consultant: more agreeable software? Journal of Applied Statistics, 18(1):107–120, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Krtolica:1991:STE

[521] Radovan Krtolica, Ivan Obradovic, and Radmilo Bozinovic. SFILTEX: toward an expert system for time-series filter design. Journal of Applied Statistics, 18(1):121–137, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andersen:1991:SES

[522] Lars Rude Andersen, Jens Herman Krebs, and Jens Damgaard Andersen. STENO: an expert system for medical diagnosis based on graphical models and model search. Journal of Applied Statistics, 18(1):139–153, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deutsch-Mcleish:1991:SUB

[523] Mary Deutsch-Mcleish, Paulyn Yao, and Tatiana Stirtzinger. A study on the use of belief functions for medical expert systems. Journal of Applied Statistics, 18(1):155–174, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeGreef:1991:ACO

[524] Paul De Greef. Analysis of co-operation for consultation systems. Journal of Applied Statistics, 18(1):175–184, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Williams:1991:CAE

[525] Calvin L. Williams. A clinical application of expert system methodology. Journal of Applied Statistics, 18(1):185–201, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilson:1991:CCS

[526] Jeffrey R. Wilson and Patrici M. Wilson. A comparison of chi-squared statistics for testing homogeneity of survey data: High School and Beyond survey. Journal of Applied Statistics, 18(2):203–213, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rao:1991:ATS

[527] Shobha Rao, Asawari Kanade, Smita Joshi, and Sharayu Paranjape. Application of time series models to detect regulatory patterns in nitrogen output of adult rats. Journal of Applied Statistics, 18(2):215–232, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andres:1991:NCR

[528] A. Martín Andrés, J. D. Luna Del Castillo, and I. Herranz Tejedor. New critical region tables for Fisher’s exact test. Journal of Applied Statistics, 18(2):233–254, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Woodward:1991:SME

[529] J. Arthur Woodward and Douglas G. Bonett. Simple main effects in factorial designs. Journal of Applied Statistics, 18(2):255–264, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fung:1991:STC

[530] Wing-Kam Fung. A statistical-test-complemented graphical method for detecting multiple outliers in two-way tables. Journal of Applied Statistics, 18(2):265–274, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Abeysinghe:1991:SAC

[531] Tilak Abeysinghe. A seasonal analysis of Chinese births. Journal of Applied Statistics, 18(2):275–286, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1991:EVD

[532] A. F. O. Bissell. Estimating variation from data with varying element sizes. Journal of Applied Statistics, 18(2):287–295, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1991:BRa

[533] Anonymous. Book reviews. Journal of Applied Statistics, 18(2):297–298, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1991:BRb

[534] Anonymous. Book reviews. Journal of Applied Statistics, 18(2):298–299, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1991:BRc

[535] Anonymous. Book reviews. Journal of Applied Statistics, 18(2):299–300, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1991:BRd

[536] Anonymous. Book reviews. Journal of Applied Statistics, 18(2):297–??, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1991:BRe

[537] Anonymous. Book reviews. Journal of Applied Statistics, 18(2):299–??, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1991:BRf

[538] Anonymous. Book reviews. Journal of Applied Statistics, 18(2):300–??, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalirajan:1991:STR

[539] K. P. Kalirajan and S. K. W. Jayasuriya. Simultaneous testing of regression disturbances for heteroscedasticity and non-normality. Journal of Applied Statistics, 18(3):307–312, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Srivastava:1991:RSD

[540] R. Srivastava, V. K. Gupta, and A. Dey. Robustness of some designs against missing data. Journal of Applied Statistics, 18(3):313–318, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chakraborti:1991:GCM

[541] S. Chakraborti. Graphical comparison of means with a standard in the one-way layout. Journal of Applied Statistics, 18(3):319–329, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chadoeuf:1991:AVM

[542] Joël Chadoeuf and Jean Baptiste Denis. Asymptotic variances for the multiplicative interaction model. Journal of Applied Statistics, 18(3):331–353, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:1991:TNA

[543] Ann-Lee Wang. Tau Ngau: an analysis and some applications of a Chinese dominoes game. Journal of Applied Statistics, 18(3):355–360, 1991. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Raju:1991:PTC

[544] C. Raju. Procedures and tables for the construction and selection of chain sampling plans ChSP-4A(c1, c2)r — part 1. Journal of Applied Statistics, 18(3):361–381, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1991:BRg

[545] Anonymous. Book review. Journal of Applied Statistics, 18(3):383–??, 1991. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Box:1992:SPD

[546] George Box and Stephen Jones. Splitplot designs for robust product experimentation. Journal of Applied Statistics, 19(1):3–26, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sen:1992:LDB

[547] Amode R. Sen and John P. Thompson. Lattice design in the Bow River Recreation Survey. Journal of Applied Statistics, 19(1):27–39, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barnett:1992:BAG

[548] V. Barnett and D. E. Wright. Biomedical applications for a generalized linear functional Poisson model. Journal of Applied Statistics, 19(1):41–47, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Watkins:1992:MLE

[549] A. J. Watkins and K. V. Mardia. Maximum likelihood estimation and prediction mean square error in the spatial linear model. Journal of Applied Statistics, 19(1):49–59, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tricker:1992:HOW

[550] A. R. Tricker. How the one-way analysis of variance model is affected by the degree of precision of the data. Journal of Applied Statistics, 19(1):61–73, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Radaelli:1992:UCT

[551] Giovanni Radaelli. Using the Cuscore technique in the surveillance of rare health events. Journal of Applied Statistics, 19(1):75–81, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Posa:1992:IFS

[552] D. Posa. The indicator formalism in spatial statistics. Journal of Applied Statistics, 19(1):83–101, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bansal:1992:SBS

[553] Ashok K. Bansal and Pankaj Sinha. Sensitivity of Bayesian sampling inspection schemes to a non-normal prior distribution. Journal of Applied Statistics, 19(1):103–109, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pettit:1992:PCN

[554] L. I. Pettit. Probabilities that contaminated normal observations are extreme. Journal of Applied Statistics, 19(1):111–116, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chan:1992:NTS

[555] Wai-Sum Chan. A note on time series model specification in the presence of outliers. Journal of Applied Statistics, 19(1):117–124, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Raju:1992:PTC

[556] C. Raju. Procedures and tables for the construction and selection of chain sampling plans ChSP-4A(c1,c2)r-Part 2. Journal of Applied Statistics, 19(1):125–140, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soundararajan:1992:SSP

[557] V. Soundararajan and S. Devaraj Arumainayagam. Some sampling plans with identical operating characteristic curves. Journal of Applied Statistics, 19(1):141–153, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shi:1992:NLC

[558] Sheng G. Shi and Winson Taam. Non-linear canonical correlation analysis with a simulated annealing solution. Journal of Applied Statistics, 19(1):155–165, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Howard:1992:CAS

[559] D. M. Howard. Canonical analysis of soil data. Journal of Applied Statistics, 19(2):171–181, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Danaher:1992:ACE

[560] Peter J. Danaher. An application of canonical expansions to media exposure distributions. Journal of Applied Statistics, 19(2):183–192, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1992:LTO

[561] A. F. Bissell. Lines through the origin — is NO INT the answer? Journal of Applied Statistics, 19(2):193–210, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Buxton:1992:SCA

[562] J. R. Buxton. Some comments on the application of diagnostic techniques in empirical modelling. Journal of Applied Statistics, 19(2):211–221, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hadzivukovic:1992:CPF

[563] S. Hadzivukovic, E. Nikolic-Djoric, and K. Cobanovic. The choice of perturbation factor in ridge regression. Journal of Applied Statistics, 19(2):223– 230, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Howlader:1992:LLS

[564] H. A. Howlader and G. Weiss. Loglogistic survival estimation based on failure-censored data. Journal of Applied Statistics, 19(2):231–240, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Govindaraju:1992:STN

[565] K. Govindaraju and K. Subramani. Selection of a tightened-normal-tightened system for given values of the acceptable quality level and limiting quality level. Journal of Applied Statistics, 19(2):241–250, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stefani:1992:PHA

[566] Raymond Stefani and Stephen Clarke. Predictions and home advantage for Australian rules football. Journal of Applied Statistics, 19(2):251–261, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kuralmani:1992:MTS

[567] V. Kuralmani and K. Govindaraju. Modified tables for the selection of quick switching systems for values of the given acceptable quality level and limiting quality level. Journal of Applied Statistics, 19(2):263–272, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yamamoto:1992:MEB

[568] E. Yamamoto and T. Yanagimoto. Moment estimators for the beta-binomial distribution. Journal of Applied Statistics, 19(2):273–283, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Constable:1992:SSA

[569] P. D. L. Constable and K. V. Mardia. Size and shape analysis of palmar interdigital areas. Journal of Applied Statistics, 19(2):285–292, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:1992:EMM

[570] K. V. Mardia. Erratum: Markov models and Bayesian methods in image analysis. Journal of Applied Statistics, 19(2):293–??, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [415].

Farnsworth:1992:ICL

[571] David L. Farnsworth. r-invariant curves for linear regression. Journal of Applied Statistics, 19(3):299–303, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kmietowicz:1992:SSL

[572] Z. W. Kmietowicz. The significance of the sample lead: alternative proof. Journal of Applied Statistics, 19(3): 305–308, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baines:1992:SAD

[573] A. Baines. A simplified approach to determination of optimum conditions experiments. Journal of Applied Statistics, 19(3):309–321, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Joanes:1992:IEM

[574] Derrick Joanes and Christine Gill. Interval estimation of a measure of diversity. Journal of Applied Statistics, 19(3):323–327, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soundararajan:1992:PTC

[575] V. Soundararajan and R. Vijayaraghavan. Procedures and tables for the construction and selection of conditional double sampling plans. Journal of Applied Statistics, 19(3):329–338, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soundararajan:1992:CST

[576] V. Soundararajan and R. Vijayaraghavan. Construction and selection of tightened-normal-tightened sampling inspection scheme of type TNT-(n1, n2; c). Journal of Applied Statistics, 19(3):339–349, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bissell:1992:MSS

[577] A. F. Bissell. Mean squares in saturated fractional designs revisited. Journal of Applied Statistics, 19(3):351– 366, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:1992:DOE

[578] Subir Ghosh and Yi-Jing Duh. Determination of optimum experimental conditions using dispersion main effects and interactions of factors in replicated factorial experiments. Journal of Applied Statistics, 19(3):367–378, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:1992:SSP

[579] Biswanath Das. Small sample properties of a sign test carried out under a Curtailed Inverse Sampling Scheme. Journal of Applied Statistics, 19(3): 379–390, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:1992:MSR

[580] Sung H. Park and Hyuk J. Kim. A measure of slope-rotatability for second order response surface experimental designs. Journal of Applied Statistics, 19(3):391–404, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Srivastava:1992:ASD

[581] Deo Kumar Srivastava, Govind S. Mudholkar, and Anil Mudholkar. Assessing the significance of difference between two quick estimates of location. Journal of Applied Statistics, 19(3):405–416, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1992:BR

[582] Anonymous. Book review. Journal of Applied Statistics, 19(3):417–418, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mcaleer:1992:RTO

[583] Michael Mcaleer and Y. K. Tse. The robustness of tests of outliers and functional form. Journal of Applied Statistics, 19(4):427–436, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bai:1992:DBN

[584] Jun Bai, Anthony J. Jakeman, and Michael McAleer. Discrimination between nested two- and three-parameter air pollutant frequency distributions. Journal of Applied Statistics, 19(4): 437–453, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ganguly:1992:PLM

[585] S. S. Ganguly and U. Naik Nimbalkar. Polytomous logistic model for matched pairs case-control design: studies with covariates. Journal of Applied Statistics, 19(4):455–464, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tricker:1992:EPR

[586] A. R. Tricker. Estimation of parameters for rounded data from nonnormal distributions. Journal of Applied Statistics, 19(4):465–471, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:1992:NIE

[587] M. Singh, G. K. Kanji, and K. S. El-Bizri. A note on inverse estimation in non-linear models. Journal of Applied Statistics, 19(4):473–477, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Basak:1992:SAC

[588] Indrani Basak, William R. Balch, and Prasanta Basak. Skewness: asymptotic critical values for a test related to Pearson’s measure. Journal of Applied Statistics, 19(4):479–487, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Subbarayan:1992:APC

[589] A. Subbarayan. On the application of a pure cyclic triple system for plant breeding experiments. Journal of Applied Statistics, 19(4):489–500, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huda:1992:MDE

[590] S. Huda and M. Shafiq. Minimax designs for estimating the slope of a second-order response surface in a cubic region. Journal of Applied Statistics, 19(4):501–507, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tomizawa:1992:MSE

[591] Sadao Tomizawa. A model of symmetry with exponents along every subdiagonal and its application to data on unaided vision of pupils at Japanese elementary schools. Journal of Applied Statistics, 19(4):509–512, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Segerstedt:1992:CPG

[592] Bo Segerstedt and Hans Nyquist. On the conditioning problem in generalized linear models. Journal of Applied Statistics, 19(4):513–526, 1992. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stonehouse:1992:THB

[593] John M. Stonehouse. Testing for notat-home bias in social surveys. Journal of Applied Statistics, 19(4):527– 531, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Irony:1992:BED

[594] Telba Z. Irony. Bayesian estimation for discrete distributions. Journal of Applied Statistics, 19(4):533–549, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nelder:1992:MCT

[595] J. A. Nelder. Models for curves through the origin. Journal of Applied Statistics, 19(4):551–552, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deligonul:1993:QSA

[596] Seyda Deligonul. Quick specification of acceptance sampling plans for dependent attribute sampling. Journal of Applied Statistics, 20(1):3–11, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mui:1993:FSP

[597] H. W. Mui and C. W. Chu. Forecasting the spot price of gold: combined forecast approaches versus a composite forecast approach. Journal of Applied Statistics, 20(1):13–23, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Broemeling:1993:BEM

[598] Lyle D. Broemeling and Peyton Cook. Bayesian estimation of the mean of an autoregressive process. Journal of Applied Statistics, 20(1):25–39, 1993. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boik:1993:TAT

[599] Robert J. Boik. Testing additivity in two-way classifications with no replications: the locally best invariant test. Journal of Applied Statistics, 20(1):41– 55, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gibbings:1993:GIM

[600] J. C. Gibbings. A generalized iteration method for deriving multiple regression curves of non-linear functions. Journal of Applied Statistics, 20(1):57–67, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hoglund:1993:MIT

[601] Rune Höglund and Ralf Östermark. Multiple input transfer function noise modelling in the time and frequency domain: empirical evidence from Monte Carlo simulations. Journal of Applied Statistics, 20(1):69–93, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gates:1993:TCS

[602] J. Gates. Testing for circularity of spatially located objects. Journal of Applied Statistics, 20(1):95–103, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prescott:1993:MDF

[603] P. Prescott, N. R. Draper, A. M. Dean, and S. M. Lewis. Mixture designs for five components in orthogonal blocks. Journal of Applied Statistics, 20(1):105–117, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Howard:1993:AEA

[604] P. J. A. Howard. An analysis of the effects of atmospheric sulphate and ammonium on soil solution chemistry using median polishing. Journal of Applied Statistics, 20(1):119–126, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Storvik:1993:DRS

[605] Geir Storvik. Data reduction by separation of signal and noise components for multivariate spatial images. Journal of Applied Statistics, 20(1):127–136, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garcia-Heras:1993:CDO

[606] J. García-Heras, J. Muñoz-García, and A. Pascual-Acosta. Criteria for the detection of outliers in the multinomial model. Journal of Applied Statistics, 20(1):137–142, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jolliffe:1993:ISO

[607] Ian T. Jolliffe and Stephen G. Lukudu. The influence of a single observation on some standard test statistics. Journal of Applied Statistics, 20(1):143– 151, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

South:1993:HHG

[608] M. C. South, G. B. Wetherill, and M. T. Tham. Hitch-hiker’s guide to genetic algorithms. Journal of Applied Statistics, 20(1):153–175, 1993. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chadoeuf:1993:GPP

[609] J. Chadoeuf, M. Goulard, and S. Pellerin. A Gibbs point process on a finite series of circles: the insertion of the primary roots of maize around the stem. Journal of Applied Statistics, 20(1):177–185, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

White:1993:JKS

[610] A. P. White and W. Z. Liu. The jack-knife with a stepwise discriminant algorithm — a warning to BMDP users. Journal of Applied Statistics, 20(1):187–190, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deadman:1993:EAC

[611] D. F. Deadman and D. J. Pyle. The effect of the abolition of capital punishment upon homicides in Great Britain: an application of intervention analysis. Journal of Applied Statistics, 20(1):191–206, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Asher:1993:BR

[612] J. M. Asher. Book review. Journal of Applied Statistics, 20(1):207–??, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Poon:1993:MLE

[613] Wai-Yin Poon, Sik-Yum Lee, and P. M. Bentler. Maximum likelihood estimation in a model with interval data.

Journal of Applied Statistics, 20(2): 219–227, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Upton:1993:TOC

[614] Graham Upton. Testing for outliers in circular data. Journal of Applied Statistics, 20(2):229–235, 1993. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:1993:SSP

[615] Yu-Min Chen and Evan E. Anderson. Small sample properties of the logit maximum likelihood estimator: a simulation study. Journal of Applied Statistics, 20(2):237–250, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Greenacre:1993:BCA

[616] Michael J. Greenacre. Biplots in correspondence analysis. Journal of Applied Statistics, 20(2):251–269, 1993. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kandasamy:1993:STT

[617] C. Kandasamy and K. Govindaraju. Selection of tightened two-level continuous sampling plans. Journal of Applied Statistics, 20(2):271–284, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Percy:1993:PGL

[618] David F. Percy. Prediction for generalized linear models. Journal of Applied Statistics, 20(2):285–291, 1993. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kmietowicz:1993:RII

[619] Z. W. Kmietowicz. Reliability of indices of industrial production in developing countries: the case of Zambia, 1980–89. Journal of Applied Statistics, 20(2):293–309, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Burrows:1993:GND

[620] Peter M. Burrows. Grouping normal distributions with unknown parameters. Journal of Applied Statistics, 20(2):311–318, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soundararajan:1993:CMS

[621] V. Soundararajan and S. Devaraj Arumainayagam. Characterization of MIL-STD-105D double sampling plans using normalized OC curves. Journal of Applied Statistics, 20(2):319–329, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Everitt:1993:BR

[622] B. S. Everitt. Book reviews. Journal of Applied Statistics, 20(2):331–332, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Buckland:1993:BR

[623] Stephen Buckland. Book reviews. Journal of Applied Statistics, 20(2):332– 333, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Smith:1993:BR

[624] Peter W. F. Smith. Book reviews. Journal of Applied Statistics, 20(2): 333–334, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ray:1993:BR

[625] W. D. Ray. Book reviews. Journal of Applied Statistics, 20(2):334–335, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dewey:1993:BR

[626] Michael Dewey. Book reviews. Journal of Applied Statistics, 20(2):335– 336, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Blackwell:1993:BR

[627] Paul Blackwell. Book reviews. Journal of Applied Statistics, 20(2):336– 337, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kilpatrick:1993:BR

[628] D. Kilpatrick. Book reviews. Journal of Applied Statistics, 20(2):337–338, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:1993:BR

[629] David J. Hand. Book reviews. Journal of Applied Statistics, 20(2):331–??, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Taam:1993:NMC

[630] Winson Taam, Perla Subbaiah, and James W. Liddy. A note on multivariate capability indices. Journal of Applied Statistics, 20(3):339–351, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barry:1993:MPP

[631] Jon Barry and Mike Nicholson. Measuring the probability of patch detection for four spatial sampling designs. Journal of Applied Statistics, 20(3):353–362, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kocherlakota:1993:SPA

[632] Kathleen Kocherlakota and Subrahmaniam Kocherlakota. A statistical problem arising in ascertainment. Journal of Applied Statistics, 20(3):363– 374, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonett:1993:SSD

[633] Douglas G. Bonett. Sample size determination for p-charts. Journal of Applied Statistics, 20(3):375–379, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prentice:1993:AR

[634] Michael J. Prentice. Axial regression. Journal of Applied Statistics, 20(3): 381–391, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Koziol:1993:ACS

[635] James A. Koziol, Milton Erman, Boris Pasche, Roza Hajdukovic, and Merrill M. Mitler. Assessing a changepoint in a sequence of repeated measurements with application to a low-energy emission therapy sleep study. Journal of Applied Statistics, 20(3):393– 400, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khamis:1993:CSC

[636] H. J. Khamis. A comparative study of the δ-corrected Kolmogorov–Smirnov test. Journal of Applied Statistics, 20 (3):401–421, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Govindaraju:1993:SMD

[637] K. Govindaraju and K. Subramani. Selection of multiple deferred (dependent) state sampling plans for given acceptable quality level and limiting quality level. Journal of Applied Statistics, 20(3):423–428, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1993:BRa

[638] Anonymous. Book reviews. Journal of Applied Statistics, 20(3):429–430, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1993:BRb

[639] Anonymous. Book reviews. Journal of Applied Statistics, 20(3):429–??, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jansen:1993:SMF

[640] J. Jansen. A simple method for fitting a linear model involving variance components. Journal of Applied Statistics, 20(4):435–444, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gan:1993:ODC

[641] F. F. Gan. An optimal design of CUSUM control charts for binomial counts. Journal of Applied Statistics, 20(4):445–460, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gore:1993:IMD

[642] A. P. Gore and V. R. Prayag. Index of marine degradation: log-linear model approach. Journal of Applied Statistics, 20(4):461–466, 1993. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kuralmani:1993:SCS

[643] V. Kuralmani and K. Govindaraju. Selection of conditional sampling plans for given AQL and LQL. Journal of Applied Statistics, 20(4):467–479, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haque:1993:PEN

[644] Mohammed Ohidul Haque. Problems in estimating non-linear Engel parameters in the absence of appropriate grouped means and average elasticity of variable elasticity Engel functions. Journal of Applied Statistics, 20(4):481–494, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Raju:1993:PTC

[645] C. Raju and M. N. Narasimha Murthy. Procedures and tables for the construction and selection of chain sampling plans ChSP-4A(c1, c2)r — part 3. Journal of Applied Statistics, 20(4): 495–511, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Subramani:1993:SMQ

[646] K. Subramani and K. Govindaraju. Selection of modified quick switching systems for given acceptable and limiting quality levels. Journal of Applied Statistics, 20(4):513–524, 1993. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berridge:1993:BRa

[647] Damon Berridge. Book reviews. Journal of Applied Statistics, 20(4):525– 526, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pakes:1993:BR

[648] Anthony G. Pakes. Book reviews. Journal of Applied Statistics, 20(4):526– 527, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chikuse:1993:BR

[649] Yasuko Chikuse. Book reviews. Journal of Applied Statistics, 20(4):527– 528, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berridge:1993:BRb

[650] Damon Berridge. Book reviews. Journal of Applied Statistics, 20(4):528– 529, 1993. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:1994:SIO

[651] K. V. Mardia. Statistics and images 2: overview. Journal of Applied Statistics, 21(1–2):1–5, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cheng:1994:SSN

[652] Bing Cheng and D. M. Titterington. A small selection of neural network methods and their statistical connections. Journal of Applied Statistics, 21(1–2): 9–37, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ripley:1994:NNF

[653] B. D. Ripley. Neural networks and flexible regression and discrimination. Journal of Applied Statistics, 21(1–2): 39–57, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kittler:1994:SPR

[654] J. Kittler. Statistical pattern recognition in image analysis. Journal of Applied Statistics, 21(1–2):61–75, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Taylor:1994:CTC

[655] C. C. Taylor and R. J. Henery. Comparative trials in classification of image data. Journal of Applied Statistics, 21(1–2):77–91, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cosman:1994:VQC

[656] Pamela C. Cosman, Robert M. Gray, and Richard A. Olshen. Vector quantization: clustering and classification trees. Journal of Applied Statistics, 21(1–2):93–108, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cootes:1994:ISU

[657] T. F. Cootes, C. J. Taylor, D. H. Cooper, and J. Graham. Image search using trained flexible shape models. Journal of Applied Statistics, 21(1–2): 111–139, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miller:1994:MMA

[658] Michael I. Miller, Sarang Joshi, David R. Maffitt, James G. Mcnally, and Ulf Grenander. Membranes, mitochondria and amoebae: shape models. Journal of Applied Statistics, 21(1–2): 141–163, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Subrahmonia:1994:ABT

[659] Jayashree Subrahmonia, Daniel Keren, and David B. Cooper. An algebraic and Bayesian technology for object recognition. Journal of Applied Statistics, 21(1–2):165–205, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kay:1994:W

[660] Jim Kay. Wavelets. Journal of Applied Statistics, 21(1–2):209–224, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lindeberg:1994:SST

[661] Tony Lindeberg. Scale-space theory: a basic tool for analyzing structures at different scales. Journal of Applied Statistics, 21(1–2):225–270, 1994. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Der:1994:SCF

[662] S. Z. Der, R. Chellappa, and E. J. M. Rignot. Statistical characterization of forward looking infra-red, laser radar and synthetic aperture radar imagery. Journal of Applied Statistics, 21(1–2): 273–312, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weir:1994:MDS

[663] Iain S. Weir and Peter J. Green. Modelling data from single-photon emission computerized tomography. Journal of Applied Statistics, 21(1–2):313– 337, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haralick:1994:SM

[664] R. M. Haralick, E. Dougherty, J. Ha, T. Kanungo, S. Karasu, C. K. Lee, L. Rystrom, V. Ramesh, and I. Phillips. Statistical morphology. Journal of Applied Statistics, 21(1–2):341–354, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:1994:EML

[665] D. Lee and Y. Vardi. Experiments with maximum likelihood method for image motion deblurring. Journal of Applied Statistics, 21(1–2):355–383, 1994. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rabe:1994:MMT

[666] S. Rabe, K. V. Mardia, and J. F. Boyce. Modelling micro-texture. Journal of Applied Statistics, 21(1–2):385– 406, 1994. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:1994:ACR

[667] D. J. Hand. Assessing classification rules. Journal of Applied Statistics, 21 (3):3–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Balasooriya:1994:SBG

[668] U. Balasooriya and T. Abeysinghe. Selecting between gamma and Weibull distributions — an approach based on predictions of order statistics. Journal of Applied Statistics, 21(3):17–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Preisler:1994:AJA

[669] H. K. Preisler. Analysis of the joint action of two biological compounds on the gypsy moth. Journal of Applied Statistics, 21(3):29–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Luijtens:1994:LNL

[670] K. Luijtens, F. Symons, and M. Vuylsteke-Wauters. Linear and non-linear canonical correlation analysis: an exploratory tool for the analysis of groupstructured data. Journal of Applied Statistics, 21(3):43–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Everitt:1994:EMD

[671] B. S. Everitt. Exploring multivariate data graphically: a brief review with examples. Journal of Applied Statistics, 21(3):63–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Reed:1994:REL

[672] J. F. Reed and D. B. Stark. Robustness estimators of location: a classification of linear and adaptive estimators. Journal of Applied Statistics, 21(3):95– ??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lovison:1994:LLM

[673] G. Lovison. Log-linear modelling of data from matched case-control studies. Journal of Applied Statistics, 21 (3):125–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ridout:1994:RCG

[674] M. S. Ridout. A response curve that generalizes the rectangular hypberbola. Journal of Applied Statistics, 21(3): 143–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Raju:1994:DCS

[675] C. Raju and J. Jothikumar. A design of chain sampling plan ChSP-1 based on Kullback–Leibler information. Journal of Applied Statistics, 21(3):153–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Berny:1994:NCA

[676] J. Berny. New concepts and applications in growth phenomena. Journal of Applied Statistics, 21(3):161–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Smeeton:1994:MST

[677] N. C. Smeeton. Marking scheme for the test for logical order in English language testing: is it fair? Journal of Applied Statistics, 21(3):191–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lund:1994:AAC

[678] R. Lund. Annual arrival cycle of Atlantic tropical cyclones. Journal of Applied Statistics, 21(4):195–??, ????

1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Edwards:1994:AFI

[679] L. J. Edwards and P. K. Sen. Approximating Fisher’s information for the functional measurement error model. Journal of Applied Statistics, 21(4): 205–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cologne:1994:SPL

[680] J. Cologne and R. Sposto. Smooth piecewise linear regression splines with hyperbolic covariates. Journal of Applied Statistics, 21(4):221–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kandasamy:1994:SCV

[681] C. Kandasamy and K. Govindaraju. Selection of CSP-V continuous sampling plans. Journal of Applied Statistics, 21(4):235–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Koesmarno:1994:HFA

[682] H. K. Koesmarno. Heuristic fitting algorithm and analytical approximation for estimating linear phase of logistic data. Journal of Applied Statistics, 21 (4):255–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Raju:1994:DCC

[683] C. Raju and M. N. Narasimha Murthy. Designing ChSP-4(c1, c2) plans. Journal of Applied Statistics, 21(4):267–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalirajan:1994:ETE

[684] K. P. Kalirajan and R. T. Shand. Estimating technical efficiency with nonneutral shift of the frontier. Journal of Applied Statistics, 21(4):279–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramirez:1994:EUI

[685] O. A. Ramirez, C. B. Moss, and W. G. Boggess. Estimation and use of the inverse hyperbolic sine transformation to model non-normal correlated random variables. Journal of Applied Statistics, 21(4):289–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Khuri:1994:EBE

[686] A. I. Khuri. Effect of blocking on the estimation of a response surface. Journal of Applied Statistics, 21(4):305–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bagchi:1994:EBE

[687] P. Bagchi. Empirical Bayes estimation in directional data. Journal of Applied Statistics, 21(4):317–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:1994:DAB

[688] G. S. Watson. Distribution of angles between unit vectors and the multiple comparison problem for unit vectors. Journal of Applied Statistics, 21 (4):327–??, ???? 1994. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wong:1995:RTE

[689] M. Y. Wong. Robustness of testing the equality of two slopes in simple structural relationships. Journal of Applied Statistics, 22(1):3–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

OBrien:1995:DIM

[690] Ralph G. O’Brien and Marc W. Heft. Discrimination indexes and models for studying sensory functioning in aging. Journal of Applied Statistics, 22(1):9– ??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Jolliffe:1995:RPC

[691] Ian T. Jolliffe. Rotation of principal components: Choice of normalization constraints. Journal of Applied Statistics, 22(1):29–35, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:1995:EPA

[692] M. Singh, N. P. Saxena, and M. C. Saxena. Estimation of pinnule area in chickpea leaves: An errors-in-variables approach. Journal of Applied Statistics, 22(1):37–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bunge:1995:CTE

[693] J. Bunge, M. Fitzpatrick, and J. Handley. Comparison of three estimators of the number of species. Journal of Applied Statistics, 22(1):45–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Jolliffe:1995:IIO

[694] I. T. Jolliffe, B. Jones, and B. J. T. Morgan. Identifying influential observations in hierarchical cluster analysis. Journal of Applied Statistics, 22(1):61– 80, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Scallan:1995:ATD

[695] A. J. Scallan. Asymmetric tolerance distributions: An application of two-piece distributions. Journal of Applied Statistics, 22(1):81–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Reed:1995:RAV

[696] J. F. Reed and D. B. Stark. Robust analysis of variance: a simulation study. Journal of Applied Statistics, 22 (1):87–104, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Arumainayagam:1995:CSQ

[697] S. Devaraj Arumainayagam and V. Soundararajan. Construction and selection of quick switching doublesampling system: Sample size tightening. Journal of Applied Statistics, 22 (1):105–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wong:1995:PTC

[698] H. Wong and W. K. Li. Portmanteau test for conditional heteroscedasticity, using ranks of squared residuals. Journal of Applied Statistics, 22(1):121–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:1995:IMO

[699] S. M. Watson, E. Redfern, S. Clark, M. Tight, and N. Davies. An influence method for outlier detection applied to time series traffic data. Journal of Applied Statistics, 22(1):135–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kmietowicz:1995:EDS

[700] Z. W. Kmietowicz. Exact distribution of the sample lead. Journal of Applied Statistics, 22(1):151–161, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mudambi:1995:SNP

[701] Ram Mudambi and Larry W. Taylor. Some non-parametric tests for duration dependence: An application to UK business cycle data. Journal of Applied Statistics, 22(1):163–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Gordon:1995:LTF

[702] A. D. Gordon. Local transformation of facial features. Journal of Applied Statistics, 22(1):179–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cadima:1995:LCI

[703] Jorge Cadima and Ian T. Jolliffe. Loadings and correlations in the interpretation of principal components. Journal of Applied Statistics, 22(2):203– 214, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:1995:AIN

[704] Yuejen Zhao and Andy H. Lee. Assessment of influence in non-linear measurement error models. Journal of Applied Statistics, 22(2):215–225, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Jeffers:1995:SC

[705] J. N. R. Jeffers. The statistician and the computer. Journal of Applied Statistics, 22(2):227–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tricker:1995:EBI

[706] A. R. Tricker. Effect on Bayesian inference of the degree of precision of recorded data. Journal of Applied Statistics, 22(2):235–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Arumainayagam:1995:CEM

[707] S. D. Arumainayagam and V. Soundararajan. Construction and evaluation of matched quick switching systems. Journal of Applied Statistics, 22 (2):245–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

John:1995:PRD

[708] Peter W. M. John. Patterns in residuals for 2^{n} designs. Journal of Applied Statistics, 22(2):253–259, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Raju:1995:PTC

[709] C. Raju and M. N. Narasimha Murthy. Procedures and tables for the construction and selection of chain sampling plans ChSP-4A(c1, c2)r — part 4. Journal of Applied Statistics, 22(2): 261–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:1995:TUD

[710] Geoffrey S. Watson. U_{n}^{2} test for uniformity of discrete distributions. Journal of Applied Statistics, 22(2):273– 276, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bolfarine:1995:SCC

[711] Heleno Bolfarine and Manuel Galea Rojas. Structural comparative calibration using the EM algorithm. Journal of Applied Statistics, 22(2):277–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Finney:1995:SSE

[712] David J. Finney. Statistical science and effective scientific communication. Journal of Applied Statistics, 22(2): 293–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hunt:1995:LES

[713] D. N. Hunt and E. Li. Loss of efficiency in spoilt experiments. Journal of Applied Statistics, 22(2):309–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ying:1995:SRAa

[714] Lisa H. Ying, Friedrich Pukelsheim, and Norman R. Draper. Slope rotatability over all directions designs. Journal of Applied Statistics, 22(3): 331–341, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ying:1995:SRAb

[715] Lisa H. Ying, Friedrich Pukelsheim, and Norman R. Draper. Slope rotatability over all directions designs for k ≥ 4. Journal of Applied Statistics, 22(3): 343–354, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearce:1995:CL

[716] S. C. Pearce. Cubic lattices. Journal of Applied Statistics, 22(3):355–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:1995:AAD

[717] R. Chen, O. Hawaleshka, and D. R. Strong. Analytical approach for determining the cost optimum raw material quality level. Journal of Applied Statistics, 22(3):375–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Raju:1995:MRC

[718] C. Raju and M. N. Narasimha Murthy. Minimum risks chain sampling plans (ChSP-4(c1, c2)) indexed by acceptable quality level and limiting quality level. Journal of Applied Statistics, 22 (3):389–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kumar:1995:CIM

[719] V. S. Sampath Kumar and P. Vasantha. Continuous inspection of Markov processes with a clearance interval. Journal of Applied Statistics, 22(3):427–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Jones:1995:SAR

[720] K. F. Jones, N. W. Taylor, and E. B. Fosam. Statistical approaches for the reliability of structural designs. Journal of Applied Statistics, 22(4):443–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Jones:1995:IIP

[721] K. F. Jones, I. Marrs, J. Young, and C. Townend. Investigation and improvement of a process for vacuumformed ceramic fibre composites, using a fractional 2” experimental design. Journal of Applied Statistics, 22 (4):459–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Vazquez-Polo:1995:BPE

[722] F.-J. Vázquez-Polo and A. Hernández Bastida. Behaviour of the posterior error rate with partial prior information in auditing. Journal of Applied Statistics, 22(4):469–476, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalirajan:1995:EIR

[723] K. P. Kalirajan and M. B. Obwona. Estimating individual response coefficients in varying coefficients regression models. Journal of Applied Statistics, 22(4):477–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Joshi:1995:PFC

[724] B. H. Joshi, U. V. Naik-Nimbalkar, S. R. Deshmukh, and A. D. Dharmadhikari. Prediction of failure — a case study. Journal of Applied Statistics, 22 (4):485–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Khattree:1995:STR

[725] Ravindra Khattree and Dayanand N. Naik. Statistical tests for random effects in staggered nested designs. Journal of Applied Statistics, 22(4):495–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wallach:1995:EBO

[726] D. Wallach. Empirical Bayes optimal fertilizer decisions. Journal of Applied Statistics, 22(4):507–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Takada:1995:ESU

[727] Hirokazu Takada, Aman Ullah, and Yu-Min Chen. Estimation of the seemingly unrelated regression model when the error covariance matrix is singular. Journal of Applied Statistics, 22 (4):517–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Carling:1995:SVI

[728] Kenneth Carling. Starting values for the iterative maximum likelihood estimator in survival analysis. Journal of Applied Statistics, 22(4):531–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Gana:1995:RRE

[729] Rajaram Gana. Ridge regression estimation of the linear probability model. Journal of Applied Statistics, 22(4): 537–??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:1995:E

[730] Anonymous.  Editorial.  Journal

of Applied Statistics, 22(5/6):553–??, November 1995. ISSN 0266-4763 (print), 1360-0532 (electronic). URL

http://www.catchword.co.uk/cgi-
bin/cgi?ini=carfax&body=linker&
reqidx=/catchword/carfax/13600532/
v22n5/s1/p553.

Pollock:1995:SHN

[731] Kenneth H. Pollock, Michael J. Conroy, and William S. Hearn. Separation of hunting and natural mortality using ring-return models: An overview. Journal of Applied Statistics, 22(5/6): 557–566, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s2/p557.

Francis:1995:ESR

[732] Charles M. Francis. Estimating survival rates from recoveries of birds ringed as young: a case study. Journal of Applied Statistics, 22(5/6): 567–578, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s3/p567.

Hoffmann:1995:IPI

[733] Annette Hoffmann and John R. Skalski. Inferential properties of an individual-based survival model using release-recapture data: Sample size, validity and power. Journal of Applied Statistics, 22(5/6): 579–596, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s4/p579.

Catchpole:1995:MAV

[734] E. A. Catchpole, S. N. Freeman, and B. J. T. Morgan. Modelling age variation in survival and reporting rates for recovery models. Journal of Applied Statistics, 22(5/6): 597–610, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s5/p597.

Burnham:1995:SAO

[735] K. P. Burnham, D. R. Anderson, and G. C. White. Selection among open population capture-recapture models when capture probabilities are heterogeneous. Journal of Applied Statistics, 22(5/6):611–624, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s6/p611.

Pugesek:1995:MRA

[736] Bruce H. Pugesek, Chris Nations, Kenneth L. Diem, and Roger Pradel. Mark-resighting analysis of a California gull population. Journal of Applied Statistics, 22(5/6):625– 640, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s7/p625.

Piper:1995:MRR

[737] S. E. Piper. A model of the ringrecovery reporting process for the Cape Griffon Gyps coprotheres. Journal of Applied Statistics, 22(5/6): 641–660, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

                                        catchword/carfax/13600532/v22n5/
                                        s8/p641.

Pollock:1995:CRS

[738] Kenneth H. Pollock, Christine M. Bunck, Scott R. Winterstein, and Chiu-Lan Chen. A capture-recapture survival analysis model for radio-tagged animals. Journal of Applied Statistics, 22(5/6):661–672, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s9/p661.

Schmutz:1995:SEE

[739] Joel A. Schmutz, David H. Ward, James S. Sedinger, and Eric A. Rexstad. Survival estimation and the effects of dependency among animals. Journal of Applied Statistics, 22(5/6): 673–682, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s10/p673.

vanNoordwijk:1995:BDO

[740] Arie J. van Noordwijk. On bias due to observer distribution in the analysis of data on natal dispersal in birds. Journal of Applied Statistics, 22(5/6): 683–694, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s11/p683.

Pradel:1995:TAR

[741] Roger Pradel, Evan Cooch, and Fred Cooke. Transient animals in a resident population of snow geese: Local emigration or heterogeneity? Journal of Applied Statistics, 22(5/6): 695–710, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s12/p695.

Schwarz:1995:EMA

[742] Carl James Schwarz, Barbara Ganter, and Branta Leucopsis. Estimating the movement among staging areas of the barnacle goose (branta leucopsis). Journal of Applied Statistics, 22 (5/6):711–724, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s13/p711.

Lindberg:1995:ENS

[743] Mark S. Lindberg, James S. Sedinger, and Eric A. Rexstad. Estimating nest site fidelity of adult female black brant with multi-state modeling and geographic information systems. Journal of Applied Statistics, 22(5/6): 725–736, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s14/p725.

Hestbeck:1995:BTS

[744] Jay B. Hestbeck. Bias in transitionspecific survival and movement probabilities estimated using capturerecapture data. Journal of Applied Statistics, 22(5/6):737–750, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL

                                        http://www.catchword.co.uk/cgi-
                                        bin/cgi?ini=carfax&body=linker&
                                        reqidx=/catchword/carfax/13600532/
                                        v22n5/s15/p737.

Kendall:1995:USC

[745] William L. Kendall and James D. Nichols. On the use of secondary capture-recapture samples to estimate temporary emigration and breeding proportions. Journal of Applied Statistics, 22(5/6): 751–762, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s16/p751.

Conroy:1995:CPM

[746] Michael J. Conroy. Comparison of programs MULT, ESTI-MATE and BROWNIE. Journal of Applied Statistics, 22(5/6):763– 774, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s17/p763.

Pradel:1995:THE

[747] R. Pradel, A.-M. Reboulet, and A. Viallefont. Testing hypotheses and estimating survival from capture histories with CR. Journal of Applied Statistics, 22(5/6):775– 784, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s18/p775.

Arnason:1995:PES

[748] A. Neil Arnason and Carl J. Schwarz. POPAN-4: Enhancements to a system for the analysis of mark-recapture data from open populations. Journal of Applied Statistics, 22(5/6): 785–800, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s19/p785.

Catchpole:1995:MEA

[749] E. A. Catchpole. MATLAB: An environment for analyzing ringrecovery and recapture data. Journal of Applied Statistics, 22(5/6): 801–816, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s20/p801.

Legendre:1995:USC

[750] S. Legendre and J. Clobert. ULM, a software for conservation and evolutionary biologists. Journal of Applied Statistics, 22(5/6):817– 834, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s21/p817.

Nichols:1995:UMS

[751] James D. Nichols and William L. Kendall. The use of multi-state capture-recapture models to address questions in evolutionary ecology. Journal of Applied Statistics, 22(5/6): 835–846, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s22/p835.

Viallefont:1995:ETO

[752] Anne Viallefont, Evan G. Cooch, and Fred Cooke. Estimation of tradeoffs with capture-recapture models: a case study on the lesser snow goose. Journal of Applied Statistics, 22(5/6): 847–862, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s23/p847.

Martin:1995:RRS

[753] Thomas E. Martin, Jean Clobert, and David R. Anderson. Return rates in studies of life history evolution: Are biases large? Journal of Applied Statistics, 22(5/6): 863–876, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s24/p863.

Hestbeck:1995:PSM

[754] Jay B. Hestbeck. Population study and management of Atlantic Flyway Canada geese. Journal of Applied Statistics, 22(5/6):877– 890, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s25/p877.

Szep:1995:SRH

[755] Tibor Szep. Survival rates of Hungarian sand martins and their relationship with Sahel rainfall. Journal of Applied Statistics, 22(5/6): 891–904, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s26/p891.

Peach:1995:DDB

[756] William J. Peach, Humphrey Q. P. Crick, and John H. Marchant. The demography of the decline in the British willow warbler population. Journal of Applied Statistics, 22(5/6): 905–922, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s27/p905.

Aebischer:1995:IEH

[757] N. J. Aebischer. Investigating the effects of hunting on the survival of British pigeons and doves by analysis of ringing recoveries. Journal of Applied Statistics, 22(5/6): 923–934, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s28/p923.

DeSante:1995:PIS

[758] David F. DeSante, Kenneth M. Burton, James F. Saracco, and Brett L. Walker. Productivity indices and survival-rate estimates from MAPS, a continent-wide programme of constanteffort mist-netting in North America. Journal of Applied Statistics, 22(5/6): 935–948, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s29/p935.

Desante:1995:SFD

[759] David F. Desante. Suggestions for future directions for studies of marked migratory landbirds from the perspective of a practitioner in population management and conservation. Journal of Applied Statistics, 22(5/6): 949–966, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s30/p949.

Baillie:1995:URD

[760] Stephen R. Baillie. Uses of ringing data for the conservation and management of bird populations: a ringing scheme perspective. Journal of Applied Statistics, 22(5/6): 967–988, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s31/p967.

Clobert:1995:CRE

[761] J. Clobert. Capture-recapture and evolutionary ecology: a difficult wedding? Journal of Applied Statistics, 22(5/6): 989–1008, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s32/p989.

Lebreton:1995:FPD

[762] Jean-Dominique Lebreton. The future of population dynamic studies using marked individuals: a statistician’s perspective. Journal of Applied Statistics, 22(5/6):1009– 1030, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s33/p1009.

Bauchau:1995:CSE

[763] V. Bauchau and A. J. Van Noordwijk. Comparison of survival estimates obtained from three different methods of recapture in the same population of the great tit. Journal of Applied Statistics, 22(5/6):1031–1038, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s34/p1031.

Kaiser:1995:ETM

[764] Andreas Kaiser. Estimating turnover, movements and capture parameters of resting passerines in standardized capture-recapture studies. Journal of Applied Statistics, 22(5/6):1039– 1048, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s35/p1039.

Pendleton:1995:DBP

[765] Grey W. Pendleton and John R. Sauer. Delineating bird populations using ring recoveries. Journal of Applied Statistics, 22(5/6):

1049–1056, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s36/p1049.

Rattiste:1995:SRB

[766] Kalev Rattiste and Vilju Lilleleht. Survival rates of breeding common gulls in Estonia. Journal of Applied Statistics, 22(5/6):1057–1062, November 1, 1995. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s37/p1057.

Dawson:1995:EBS

[767] Deanna K. Dawson, John R. Sauer, Paul A. Wood, Mauro Berlanga, Marcia H. Wilson, and Chandler S. Robbins. Estimating bird species richness from capture and count data. Journal of Applied Statistics, 22(5/6): 1063–1068, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v22n5/
s38/p1063.

Senar:1995:MHS

[768] J. C. Senar and J. L. Copete. Mediterranean house sparrows (passer domesticus) are not used to freezing temperatures: an analysis of survival rates. Journal of Applied Statistics, 22(5/6): 1069–1074, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s39/p1069.

Francis:1995:HUR

[769] Charles M. Francis. How useful are recoveries of North American passerines for survival analyses? Journal of Applied Statistics, 22(5/6): 1075–1081, November 1, 1995. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v22n5/
                                        s40/p1075.

Anonymous:1995:TPV

[770] Anonymous. Title-page, Volume Contents and Author Index, 1995. Journal of Applied Statistics, 22(5/6):1083– ??, ???? 1995. ISSN 0266-4763 (print), 1360-0532 (electronic).

Box:1996:RSQ

[771] George E. P. Box. Role of statistics in quality and productivity improvement. Journal of Applied Statistics, 23(1):3–20, February 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n1/
                                        s1/p3.

Gilmour:1996:FFS

[772] S. G. Gilmour and R. Mead. Fixing a factor in the sequential design of two-level fractional factorial experiments. Journal of Applied Statistics, 23(1):21–30, February 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s2/p21.

Walton:1996:FMD

[773] Todd L. Walton, Jr. Fill-in of missing data in univariate coastal data. Journal of Applied Statistics, 23(1): 31–40, February 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s3/p31.

Castagliola:1996:OAC

[774] Philippe Castagliola. Optimized algorithms for computing Wilcoxon’s T_{n}, Wilcoxon’s W_{m,n} and Ansari-Bradley’s A_{m,n} statistics when m and n are small. Journal of Applied Statistics, 23(1):41–57, February 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s4/p41.

Achcar:1996:EVR

[775] Jorge Alberto Achcar and Vera Lucia Damasceno. Extreme value regression models: a useful reparameterization for the survival function. Journal of Applied Statistics, 23(1): 59–68, February 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n1/
                                        s5/p59.

Vijayaraghavan:1996:PTS

[776] R. Vijayaraghavan and V. Soundararajan. Procedures and tables for the selection of tightened-normal-tightened (TNT-(n; c_{1}, c_{2})) sampling schemes. Journal of Applied Statistics, 23(1): 69–80, February 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n1/
                                        s6/p69.

Tsui:1996:CLT

[777] Kwok-Leung Tsui. A critical look at Taguchi’s modelling approach for robust design. Journal of Applied Statistics, 23(1):81–96, February 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n1/
                                        s7/p81.

Caudill:1996:MLE

[778] Steven B. Caudill. Maximum likelihood estimation in a model with interval data: a comment and extension. Journal of Applied Statistics, 23 (1):97–104, February 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n1/
                                        s8/p97.

Radaelli:1996:DUI

[779] Giovanni Radaelli. Detection of an unknown increase in the rate of a rare event. Journal of Applied Statistics, 23(1):105–114, February 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s9/p105.

Arumainayagam:1996:SSI

[780] S. Devaraj Arumainayagam and V. Soundararajan. Sampling systems indexed by the crossover point. Journal of Applied Statistics, 23(1): 115–124, February 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s10/p115.

Kenett:1996:DAA

[781] Ron S. Kenett and Moshe Pollak. Data-analytic aspects of the Shiryayev-Roberts control chart: Surveillance of a non-homogeneous Poisson process. Journal of Applied Statistics, 23 (1):125–137, February 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s11/p125.

Mardia:1996:STD

[782] K. V. Mardia, A. J. Baczkowski, X. Feng, and P. A. Millner. A study of three-dimensional curves. Journal of Applied Statistics, 23(1):139– 148, February 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s12/p139.

Anonymous:1996:BRa

[783] Anonymous. Book reviews. Journal of Applied Statistics, 23(1):149–??, February 1996. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n1/
s13/p149.

Jeffers:1996:E

[784] J. N. R. Jeffers. Editorial. Journal of Applied Statistics, 23(2/3): 163–164, June 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n2/
s1/p163.

Young:1996:SCE

[785] Peter Young, Stuart Parkinson, and Matthew Lees. Simplicity out of complexity in environmental modelling: Occam’s razor revisited. Journal of Applied Statistics, 23(2/3): 165–210, June 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v23n2/
s2/p165.

Moss:1996:ADT

[786] R. Moss, D. C. Catt, N. G. Bayfield, and D. D. French. The application of decision theory to sustainable management of an upland Scottish estate. Journal of Applied Statistics, 23(2/3):211–230, June 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n2/
s3/p211.

Kaur:1996:ESC

[787] Amarjot Kaur, G. P. Patil, Susan J. Shirk, and Charles Taillie. Environmental sampling with a concomitant variable: a comparison between ranked set sampling and stratified simple random sampling. Journal of Applied Statistics, 23(2/3): 231–255, June 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n2/
s4/p231.

Gregoire:1996:NLM

[788] Timothy G. Gregoire and Oliver Schabenberger. A non-linear mixedeffects model to predict cumulative bole volume of standing trees. Journal of Applied Statistics, 23(2/3): 257–272, June 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?

                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n2/
                                        s5/p257.

Freeman:1996:MMU

[789] G. H. Freeman, S. D. Jacka, J. E. H. Shaw, and J. Q. Smith. Modelling the management of underground water assets. Journal of Applied Statistics, 23(2/3):273–284, June 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n2/
                                        s6/p273.

Green:1996:BGY

[790] Edwin J. Green and William E. Strawderman. A Bayesian growth and yield model for slash pine plantations. Journal of Applied Statistics, 23(2/3):285–300, June 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n2/
                                        s7/p285.

Lawson:1996:APE

[791] Andrew B. Lawson and Natalie W. Harrington. The analysis of putative environmental pollution gradients in spatially correlated epidemiological data. Journal of Applied Statistics, 23(2/3):301–310, June 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n2/
                                        s8/p301.

Clarke:1996:DBQ

[792] R. T. Clarke, M. T. Furse, J. F. Wright, and D. Moss. Derivation of a biological quality index for river sites: Comparison of the observed with the expected fauna. Journal of Applied Statistics, 23(2/3):311–332, June 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n2/
s9/p311.

Coles:1996:AEP

[793] Stuart G. Coles and Francesca Pan. The analysis of extreme pollution levels: a case study. Journal of Applied Statistics, 23(2/3): 333–348, June 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n2/
s10/p333.

Bell:1996:ACT

[794] John F. Bell. Application of classification trees to the habitat preference of upland birds. Journal of Applied Statistics, 23(2/3):349–360, June 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n2/
s11/p349.

Nicholson:1996:SDD

[795] Mike Nicholson and Jon Barry. Survey design for detecting patches. Journal of Applied Statistics, 23(2/3):

361–368, June 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n2/
s12/p361.

Liu:1996:SGSa

[796] W. Liu. Some group sequential procedures for comparing several treatments with a control. Journal of Applied Statistics, 23(4):371– 383, August 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n4/
s1/p371.

Huerta:1996:BAS

[797] Gabriel Huerta and Manuel Mendoza. A Bayesian approach to sample size selection in a multiple hypotheses test for the Bernoulli model. Journal of Applied Statistics, 23(4): 385–394, August 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n4/
s2/p385.

Chinn:1996:CT

[798] S. Chinn. Choosing a transformation. Journal of Applied Statistics, 23(4):395–404, August 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v23n4/
s3/p395.

Fung:1996:DVS

[799] Wing K. Fung. The dependence of voter shares under a two-vote system. Journal of Applied Statistics, 23(4):405–412, August 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n4/
s4/p405.

Sugden:1996:CLS

[800] R. A. Sugden, T. M. F. Smith, and R. P. Brown. Chao’s list sequential scheme for unequal probability sampling. Journal of Applied Statistics, 23(4):413–421, August 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n4/
s5/p413.

Koesmarno:1996:CSP

[801] H. K. Koesmarno. Class-size percentile transformation for reconstructing a distribution function. Journal of Applied Statistics, 23(4):423– 434, August 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n4/
s6/p423.

Cherry:1996:ENP

[802] S. Cherry, J. Banfield, and W. F. Quimby. An evaluation of a nonparametric method of estimating semivariograms of isotropic spatial processes. Journal of Applied Statistics, 23(4):435–449, August 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n4/
                                        s7/p435.

Anonymous:1996:BRb

[803] Anonymous. Book reviews. Journal of Applied Statistics, 23(4):451–??, August 1996. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n4/
                                        s8/p451.

Liu:1996:SGSb

[804] W. Liu. Some group sequential procedures for comparing several treatments with a control. Journal of Applied Statistics, 23(4):1000–1001, August 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Evans:1996:AML

[805] Mark Evans. Aspects of maximum likelihood estimation of asymmetric growth curves. Journal of Applied Statistics, 23(5):467– 492, October 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v23n5/
s1/p467.

Arnoldsson:1996:OAL

[806] Göran Arnoldsson. Optimal allocation for linear combinations of linear predictors in generalized linear models. Journal of Applied Statistics, 23 (5):493–505, October 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n5/
s2/p493.

ElBarmi:1996:SIC

[807] Hammou El Barmi, Ian R. Harris, and Ayanendranath Basu. Statistical inference concerning weighted Poisson rates under some natural order restrictions. Journal of Applied Statistics, 23(5):507–514, October 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n5/
s3/p507.

Tse:1996:EML

[808] S. K. Tse and Geoffrey Tso. Efficiencies of maximum likelihood estimators under censoring in life testing. Journal of Applied Statistics, 23 (5):515–524, October 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n5/
s4/p515.

Hammond:1996:ACS

[809] P. S. Hammond and P. Rothery. Application of computer sampling in the estimation of seal diet. Journal of Applied Statistics, 23(5):525– 534, October 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n5/
                                        s5/p525.

Kim:1996:LIM

[810] Myung Geun Kim. Local influence in multivariate normal data. Journal of Applied Statistics, 23(5):535– 542, October 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n5/
                                        s6/p535.

Baker:1996:TST

[811] Rose D. Baker. Testing for space– time clusters of unknown size. Journal of Applied Statistics, 23(5):543– 554, October 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v23n5/
                                        s7/p543.

Finney:1996:NHR

[812] David J. Finney. A note on the history of regression. Journal of Applied Statistics, 23(5):555–558, October 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n5/
s8/p555.

Anonymous:1996:BRc

[813] Anonymous. Book reviews. Journal of Applied Statistics, 23(5):559–??, October 1996. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n5/
s9/p559.

Jeffers:1996:MAR

[814] J. N. R. Jeffers. Multivariate analysis of a reference collection of elm leaves. Journal of Applied Statistics, 23(6):571–588, December 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s1/p571.

Lee:1996:EAM

[815] Jong Hyup Lee. Estimation for the autoregressive moving average process observed with noise. Journal of Applied Statistics, 23(6):589– 599, December 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s2/p589.

Jeong:1996:DHS

[816] Hyean-Seok Jeong, Jong-In Park, and Bong-Jin Yum. Development of (r, T) hybrid sampling plans for exponential lifetime distributions. Journal of Applied Statistics, 23(6):601– 608, December 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s3/p601.

Lindsay:1996:MDD

[817] S. R. Lindsay, G. R. Wood, and R. C. Woollons. Modelling the diameter distribution of forest stands using the Burr distribution. Journal of Applied Statistics, 23(6):609– 620, December 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s4/p609.

Scarf:1996:EEC

[818] Philip A. Scarf and Patrick J. Laycock. Estimation of extremes in corrosion engineering. Journal of Applied Statistics, 23(6):621–644, December 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s5/p621.

Suresh:1996:SSP

[819] K. K. Suresh and T. B. Ramkumar. Selection of a sampling plan indexed with maximum allowable average outgoing quality. Journal of Applied Statistics, 23(6):645–654, December 1, 1996. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s6/p645.

Mardia:1996:SPF

[820] K. V. Mardia, A. Coombes, J. Kirkbride, A. Linney, and J. L. Bowie. On statistical problems with face identification from photographs. Journal of Applied Statistics, 23(6):655– 676, December 1, 1996. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s7/p655.

Anonymous:1996:BRd

[821] Anonymous. Book reviews. Journal of Applied Statistics, 23(6):677–??, December 1996. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v23n6/
s8/p677.

Anonymous:1996:TPV

[822] Anonymous. Title-page, volume contents and author index, 1996. Journal of Applied Statistics, 23(6):681–??, December 1996. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chatterjee:1997:GPW

[823] Sangit Chatterjee and Matthew Laudato. Gender and performance of world-class athletes. Journal of Applied Statistics, 24(1):3– 10, February 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n1/
                                        s1/p3.

Yu:1997:HPE

[824] P. L. H. Yu and K. Lam. How to predict election winners from a poll. Journal of Applied Statistics, 24(1):11–23, February 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n1/
                                        s2/p11.

Franses:1997:CVU

[825] Philip Hans Franses and Bart Hobijn. Critical values for unit root tests in seasonal time series. Journal of Applied Statistics, 24(1):25– 47, February 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n1/
                                        s3/p25.

Raju:1997:PTC

[826] C. Raju and J. Jothikumar. Procedures and tables for the construction and selection of chain sampling plans ChSP-4A(c_{1}, c_{2})r. V. Journal of Applied Statistics, 24(1):49– 76, February 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n1/
s4/p49.

Hung:1997:CTL

[827] Ken Hung. A comparison of two large sample confidence intervals for a proportion: A Monte Carlo simulation. Journal of Applied Statistics, 24(1):77–83, February 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n1/
s5/p77.

Al-Mutairi:1997:DLS

[828] Dhaifalla K. Al-Mutairi and Satish K. Agarwal. Distributions of the lifetimes of system components operating under an unknown common environment. Journal of Applied Statistics, 24(1):85–96, February 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n1/
s6/p85.

Agrawal:1997:HPR

[829] M. C. Agrawal and A. B. Sthapit. Hierarchic predictive ratio-based and product-based estimators and their efficiency. Journal of Applied Statistics, 24(1):97–104, February 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n1/
                                        s7/p97.

Munoz:1997:IKS

[830] Sergio R. Munoz and Shrikant I. Bangdiwala. Interpretation of Kappaand B statistics measures of agreement. Journal of Applied Statistics, 24 (1):105–112, February 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n1/
                                        s8/p105.

Gilchrist:1997:MQD

[831] Warren Gilchrist. Modelling with quantile distribution functions. Journal of Applied Statistics, 24(1):113– 122, February 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n1/
                                        s9/p113.

Anonymous:1997:BRa

[832] Anonymous. Book reviews. Journal of Applied Statistics, 24(1):123–??, February 1997. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n1/
                                        s10/p123.

Pettit:1997:SPF

[833] L. I. Pettit and J. L. Palmer. Seasonal patterns of fertility measures: a Bayesian approach. Journal of Applied Statistics, 24(2):139–146, April 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s1/p139.

Prescott:1997:FPM

[834] P. Prescott, N. R. Draper, S. M. Lewis, and A. M. Dean. Further properties of mixture designs for five components in orthogonal blocks. Journal of Applied Statistics, 24(2):147–156, April 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s2/p147.

Nummi:1997:ERE

[835] Tapio Nummi. Estimation in a random effects growth curve model. Journal of Applied Statistics, 24(2): 157–168, April 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s3/p157.

Basford:1997:MDS

[836] K. E. Basford, G. J. Mclachlan, and M. G. York. Modelling the distribution of stamp paper thickness via finite normal mixtures: The 1872 Hidalgo stamp issue of Mexico revisited. Journal of Applied Statistics, 24(2):169–180, April 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s4/p169.

Shao:1997:MIA

[837] Y. Eric Shao. Multiple intervention analysis with application to sales promotion data. Journal of Applied Statistics, 24(2):181–192, April 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s5/p181.

Schader:1997:PTU

[838] Martin Schader and Friedrich Schmid. Power of tests for uniformity when limits are unknown. Journal of Applied Statistics, 24(2):193–205, April 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s6/p193.

Soundararajan:1997:SSS

[839] V. Soundararajan and A. L. Christina. Selection of single sampling variables plans based on the minimum angle. Journal of Applied Statistics, 24(2):207–218, April 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?

ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s7/p207.

Rahman:1997:MTS

[840] M. Mahibbur Rahman and Z. Govindarajulu. A modification of the test of Shapiro and Wilk for normality. Journal of Applied Statistics, 24(2):219–235, April 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s8/p219.

Anonymous:1997:BRb

[841] Anonymous. Book reviews. Journal of Applied Statistics, 24(2):237–??, April 1997. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n2/
s9/p237.

Kao:1997:OSS

[842] L. H. Kao and S. Chakraborti. Onesided sign-type non-parametric procedures for comparing treatments with a control in a randomized complete block design. Journal of Applied Statistics, 24(3):251–264, June 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n3/
s1/p251.

Chen:1997:MSM

[843] H. H. Chen, S. W. Duffy, and L. Tabár. A mover-stayer mixture of Markov chain models for the assessment of dedifferentiation and tumour progression in breast cancer. Journal of Applied Statistics, 24(3):265–278, June 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n3/
                                        s2/p265.

Rao:1997:GMT

[844] C. V. Rao and S. Hari Krishna. A graphical method for testing the equality of several variances. Journal of Applied Statistics, 24(3): 279–288, June 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n3/
                                        s3/p279.

Ghatak:1997:URS

[845] Anita Ghatak. Unit roots and structural breaks: The case of India 1900– 1988. Journal of Applied Statistics, 24(3):289–300, June 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n3/
                                        s4/p289.

Khamis:1997:CKS

[846] H. J. Khamis. The δ-corrected Kolmogorov–Smirnov test for the two-parameter Weibull distribution. Journal of Applied Statistics, 24(3): 301–317, June 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n3/
s5/p301.

Buning:1997:RAV

[847] Herbert Büning. Robust analysis of variance. Journal of Applied Statistics, 24(3):319–332, June 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n3/
s6/p319.

Munoz-Garcia:1997:IOB

[848] J. Muñoz-García, R. Pino-Mejías, J. M. Muñoz-Pichardo, and M. D. Cubiles-De-La-Vega. Identification of outlier bootstrap samples. Journal of Applied Statistics, 24(3):333–342, June 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n3/
s7/p333.

Johnes:1997:IUV

[849] Jill Johnes. Inter-university variations in undergraduate non-completion rates: a statistical analysis by subject of study. Journal of Applied Statistics, 24(3):343–362, June 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?

                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n3/
                                        s8/p343.

Anonymous:1997:BRc

[850] Anonymous. Book reviews. Journal of Applied Statistics, 24(3):363–??, June 1997. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n3/
                                        s9/p363.

Katholi:1997:AVT

[851] Charles R. Katholi, Anthony Merriweather, and Thomas R. Unnasch. An analysis of variance type test for comparing clusters of DNA sequences based on randomization test methodologies. Journal of Applied Statistics, 24(4):371–382, August 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n4/
                                        s1/p371.

Irmay:1997:RBZ

[852] Shragga Irmay. The relationship between Zipf’s law and the distribution of first digits. Journal of Applied Statistics, 24(4):383–393, August 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n4/
                                        s2/p383.

Khattree:1997:EVC

[853] Ravindra Khattree, Dayanand N. Naik, and Robert L. Mason. Estimation of variance components in staggered nested designs. Journal of Applied Statistics, 24(4):395– 408, August 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n4/
s3/p395.

Schoen:1997:DAF

[854] Eric D. Schoen and Kirsten Wolff. Design and analysis of a fractional 41 31 25 split-plot experiment. Journal of Applied Statistics, 24(4):409– 420, August 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n4/
s4/p409.

Acuna:1997:SAR

[855] Carmen O. Acuña and Joseph Horowitz. A statistical approach to the resolution of point sources. Journal of Applied Statistics, 24(4): 421–435, August 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n4/
s5/p421.

Aggarwal:1997:CGS

[856] M. L. Aggarwal, A. Goel, and S. R. Chowdhury. Catalogue of group structures for two-level fractional factorial designs. Journal of Applied Statistics, 24(4):437–452, August 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n4/
s6/p437.

Salagame:1997:FHD

[857] Raviprakash R. Salagame and Russell R. Barton. Factorial hypercube designs for spatial correlation regression. Journal of Applied Statistics, 24(4):453–474, August 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n4/
s7/p453.

Zhang:1997:DCR

[858] Nien Fan Zhang. Detection capability of residual control chart for stationary process data. Journal of Applied Statistics, 24(4):475–492, August 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n4/
s8/p475.

Tarpey:1997:EPP

[859] Thaddeus Tarpey. Estimating principal points of univariate distributions. Journal of Applied Statistics, 24(5):499–512, October 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?

ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n5/
s1/p499.

Dryden:1997:RUC

[860] I. L. Dryden, K. V. Mardia, and A. N. Walder. Review of the use of context in statistical image analysis. Journal of Applied Statistics, 24 (5):513–538, October 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n5/
s2/p513.

Wang:1997:MAF

[861] Reay-Chen Wang and Chung-Ho Chen. Minimum average fraction inspected for continuous sampling plan CSP-1 under inspection error. Journal of Applied Statistics, 24(5):539– 548, October 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n5/
s3/p539.

Ducharme:1997:CSA

[862] Gilles R. Ducharme. Consistent selection of the actual model in regression analysis. Journal of Applied Statistics, 24(5):549–558, October 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n5/
s4/p549.

Tellez:1997:ENP

[863] E. Ayuga Tellez, C. Ayuga Tellez, C. Gonzalez Garcia, and E. Martinez Falero. Estimation of, nonparametric regression in the analysis of the anti-inflammatory activity of diverse extracts of Sideritis foetens. Journal of Applied Statistics, 24(5):559–572, October 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n5/
                                        s5/p559.

Madry:1997:SAM

[864] Wieslaw Madry. A statistical approach to multivariate evaluation of diversity with respect to quantitative characteristics in cereal germplasm collections. Journal of Applied Statistics, 24(5):573–588, October 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n5/
                                        s6/p573.

Boyles:1997:UCS

[865] Russell A. Boyles. Using the chi-square statistic to monitor compositional process data. Journal of Applied Statistics, 24(5):589–602, October 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v24n5/
                                        s7/p589.

Tanizaki:1997:PCN

[866] Hisashi Tanizaki. Power comparison of non-parametric tests: Small-sample properties from Monte Carlo experiments. Journal of Applied Statistics, 24(5):603–632, October 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n5/
s8/p603.

Stefani:1997:SMW

[867] Raymond T. Stefani. Survey of the major world sports rating systems. Journal of Applied Statistics, 24(6): 635–646, December 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s1/p635.

Qannari:1997:NAB

[868] E. M. Qannari, E. Vigneau, and M. Semenou. New approach in biased regression. Journal of Applied Statistics, 24 (6):647–658, December 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s2/p647.

Singh:1997:ETD

[869] Murari Singh and Michael Jones. Estimating time to detect time trends in continuous cropping. Journal of Applied Statistics, 24(6):659– 670, December 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s3/p659.

Kaboudan:1997:NTA

[870] M. A. Kaboudan. Non-traditional analysis of stock returns. Journal of Applied Statistics, 24(6):671– 688, December 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s4/p671.

Balamurali:1997:DAS

[871] S. Balamurali and M. Kalyanasundaram. Determination of an attribute single sampling scheme. Journal of Applied Statistics, 24(6):689– 696, December 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s5/p689.

Barnett:1997:BLU

[872] Vic Barnett and Karen Moore. Best linear unbiased estimates in rankedset sampling with particular reference to imperfect ordering. Journal of Applied Statistics, 24(6):697– 710, December 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v24n6/
s6/p697.

Khuri:1997:QDG

[873] A. I. Khuri. Quantile dispersion graphs for analysis of variance estimates of variance components. Journal of Applied Statistics, 24(6):711– 722, December 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s7/p711.

Mudambi:1997:ETP

[874] Ram Mudambi. Estimating turning points using polynomial regression. Journal of Applied Statistics, 24 (6):723–732, December 1, 1997. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s8/p723.

ODonnell:1997:MSE

[875] Eileen M. O’Donnell and G. Geoffrey Vining. Mean squared error of prediction approach to the analysis of a combined array. Journal of Applied Statistics, 24(6):733–746, December 1, 1997. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s9/p733.

Anonymous:1997:CAI

[876] Anonymous. Contents and author index: Volume 24, 1997. Journal of Applied Statistics, 24(6):747–??, December 1997. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s10/p747.

Anonymous:1997:CV

[877] Anonymous. Contents, Volume 24, 1997. Journal of Applied Statistics, 24(6):748–??, December 1997. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s11/p748.

Anonymous:1997:IAV

[878] Anonymous. Index of authors, volume 24, 1997. Journal of Applied Statistics, 24(6):751–??, December 1997. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v24n6/
s12/p751.

Anonymous:1998:EYA

[879] Anonymous. Editorial: 25 years of Applied Statistics. Journal of Applied Statistics, 25(1):3–22, February 1998. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s1/p3.

Hillson:1998:SSM

[880] Eric M. Hillson, Jack H. Reeves, and Charlotte A. McMillan. A statistical signalling model for use in surveillance of adverse drug reaction data. Journal of Applied Statistics, 25(1):23–40, February 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s2/p23.

Park:1998:ODA

[881] Sang-Jun Park and Bong-Jin Yum. Optimal design of accelerated life tests under modified stress loading methods. Journal of Applied Statistics, 25(1):41–62, February 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s3/p41.

Stonehouse:1998:RTU

[882] John M. Stonehouse and Guy J. Forrester. Robustness of the t and U tests under combined assumption violations. Journal of Applied Statistics, 25(1):63–74, February 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s4/p63.

Tse:1998:EET

[883] Siu-Keung Tse and Hak-Keung Yuen. Expected experiment times for the Weibull distribution under progressive censoring with random removals. Journal of Applied Statistics, 25(1): 75–83, February 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s5/p75.

Tan:1998:EHS

[884] Wai-Yuan Tan, Si Chin Tang, and Sho Rong Lee. Estimation of HIV seroconversion and effects of age in the San Francisco homosexual population. Journal of Applied Statistics, 25(1):85–102, February 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s6/p85.

Govindaraju:1998:CSP

[885] K. Govindaraju and S. Balamurali. Chain sampling plan for variables inspection. Journal of Applied Statistics, 25(1):103–109, February 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s7/p103.

Qiao:1998:BEE

[886] Hongzhu Qiao and Chris P. Tsokos. Best efficient estimates of the intensity function of the power law process. Journal of Applied Statistics, 25 (1):111–120, February 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s8/p111.

Yahyah:1998:GAM

[887] M. Yahyah, A. Baines, and D. N. Joanes. Graphical approach to model adequacy based on exact and near replicates. Journal of Applied Statistics, 25(1):121–129, February 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s9/p121.

Toler:1998:GPE

[888] J. E. Toler and P. M. Burrows. Genotypic performance over environmental arrays: a non-linear grouping protocol. Journal of Applied Statistics, 25 (1):131–143, February 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n1/
s10/p131.

Kim:1998:LIT

[889] Myung Geun Kim. Local influence on a test of linear hypothesis in multiple regression model. Journal of Applied Statistics, 25(1):145– 152, February 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

                                        catchword/carfax/13600532/v25n1/
                                        s11/p145.

Glasbey:1998:RIW

[890] C. A. Glasbey and K. V. Mardia. A review of image-warping methods. Journal of Applied Statistics, 25(2):155–171, April 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n2/
                                        s1/p155.

Bredrup:1998:ISD

[891] Edvin Bredrup and Li-Chun Zhang. Imperfectly shuffled decks in bridge. Journal of Applied Statistics, 25(2): 173–179, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n2/
                                        s2/p173.

Koziol:1998:NPI

[892] James A. Koziol. A non-parametric index of tracking. Journal of Applied Statistics, 25(2):181–191, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n2/
                                        s3/p181.

Taylor:1998:CFQ

[893] James W. Taylor and Derek W. Bunn. Combining forecast quantiles using quantile regression: Investigating the derived weights, estimator bias and imposing constraints. Journal of Applied Statistics, 25(2):193–206, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n2/
s4/p193.

Kabir:1998:EWD

[894] A. B. M. Zohrul Kabir. Estimation of Weibull distribution parameters for irregular interval group failure data with unknown failure times. Journal of Applied Statistics, 25(2): 207–219, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n2/
s5/p207.

Ali:1998:PMH

[895] Mukhtar M. Ali. Probability models on horse-race outcomes. Journal of Applied Statistics, 25(2):221– 229, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n2/
s6/p221.

Mardia:1998:FRN

[896] K. V. Mardia. Fisher’s repeated normal integral function and shape distributions. Journal of Applied Statistics, 25(2):231–235, April 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n2/
                                        s7/p231.

Andrade:1998:PAR

[897] Dalton F. Andrade and Julio M. Singer. Profile analysis for randomized complete block experiments. Journal of Applied Statistics, 25(2): 237–244, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n2/
                                        s8/p237.

Zhang:1998:TMU

[898] Jin Zhang. Tests for multiple upper or lower outliers in an exponential sample. Journal of Applied Statistics, 25(2):245–255, April 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n2/
                                        s9/p245.

Zhang:1998:UTM

[899] Jin Zhang and Xueren Wang. Unmasking test for multiple upper or lower outliers in normal samples. Journal of Applied Statistics, 25(2): 257–261, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n2/
                                        s10/p257.

Anderson:1998:CAI

[900] D. R. Anderson, K. P. Burnham, and G. C. White. Comparison of Akaike information criterion and consistent Akaike information criterion for model selection and statistical inference from capture-recapture studies. Journal of Applied Statistics, 25(2):263–282, April 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n2/
s11/p263.

Guerrero:1998:FEC

[901] Victor M. Guerrero and Edmundo Berumen. Forecasting electricity consumption with extra-model information provided by consumers. Journal of Applied Statistics, 25(2): 283–299, April 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n2/
s12/p283.

Mittag:1998:GIE

[902] Hans-Joachim Mittag and Dietmar Stemann. Gauge imprecision effect on the performance of the X-S control chart. Journal of Applied Statistics, 25(3):307–317, June 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n3/
s1/p307.

Wang:1998:UMM

[903] Duolao Wang and Mike Murphy. Use of a mixture model for the analysis of contraceptive-use duration among long-term users. Journal of Applied Statistics, 25(3):319–332, June 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n3/
                                        s2/p319.

Sherman:1998:ATD

[904] Michael Sherman, F. Michael Speed, Jr., and F. Michael Speed. Analysis of tidal data via the blockwise bootstrap. Journal of Applied Statistics, 25(3):333–340, June 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n3/
                                        s3/p333.

Vijayaraghavan:1998:DES

[905] R. Vijayaraghavan and V. Soundararajan. Design and evaluation of skip-lot sampling inspection plans with double-sampling plan as the reference plan. Journal of Applied Statistics, 25(3):341–348, June 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n3/
                                        s4/p341.

Ghosh:1998:ICD

[906] D. K. Ghosh and S. C. Bagui. Identification of confounded design and its interactions. Journal of Applied Statistics, 25(3):349–356, June 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n3/
s5/p349.

Shoukri:1998:PVS

[907] M. M. Shoukri, M. Attanasio, and J. M. Sargeant. Parametric versus semi-parametric models for the analysis of correlated survival data: a case study in veterinary epidemiology. Journal of Applied Statistics, 25(3):357–374, June 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n3/
s6/p357.

Edwards:1998:NPP

[908] Christopher Todd Edwards. Nonparametric procedure for knockout tournaments. Journal of Applied Statistics, 25(3):375–385, June 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n3/
s7/p375.

Tomizawa:1998:PDT

[909] Sadao Tomizawa, Takashi Seo, and Hideharu Yamamoto. Power-divergence-type measure of departure from symmetry for square contingency tables that have nominal categories. Journal of Applied Statistics, 25(3):

387–398, June 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n3/
                                        s8/p387.

Govindaraju:1998:ZAN

[910] K. Govindaraju and S. Ganesalingam. Zero acceptance number quick switching system for compliance sampling. Journal of Applied Statistics, 25(3): 399–407, June 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n3/
                                        s9/p399.

Trivez:1998:AEL

[911] F. Javier Trivez and Javier Nievas. Analyzing the effects of level shifts and temporary changes on the identification of ARIMA models. Journal of Applied Statistics, 25(3): 409–424, June 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n3/
                                        s10/p409.

Li:1998:BOM

[912] Stan Z. Li. Bayesian object matching. Journal of Applied Statistics, 25(3):425–443, June 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v25n3/
s11/p425.

Govindaraju:1998:TSL

[913] K. Govindaraju and S. Balamurali. Tightened single-level continuous sampling plan. Journal of Applied Statistics, 25(4):451–461, August 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s1/p451.

Efaw:1998:TAS

[914] Fritz Efaw. Test of alternative strike settlement models. Journal of Applied Statistics, 25(4):463– 474, August 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s2/p463.

Ghatak:1998:ACF

[915] Anita Ghatak. Aggregate consumption functions for India: a cointegration analysis under structural changes, 1919–86. Journal of Applied Statistics, 25(4):475–488, August 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s3/p475.

Koop:1998:CDE

[916] Gary Koop. Carbon dioxide emissions and economic growth: a structural approach. Journal of Applied Statistics, 25(4):489–515, August 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s4/p489.

Dmitrienko:1998:DPY

[917] Alexei Dmitrienko and Z. Govindarajulu. The ‘demon’ problem of Youden: Exponential case. Journal of Applied Statistics, 25(4):517– 523, August 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s5/p517.

Hutson:1998:DEP

[918] Alan D. Hutson. Direct estimation of the percentile ‘p-value’ for the one-sample median test. Journal of Applied Statistics, 25(4):525– 533, August 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s6/p525.

Lai:1998:ECF

[919] C. D. Lai, K. Govindaraju, and M. Xie. Effects of correlation on fraction nonconforming statistical process control procedures. Journal of Applied Statistics, 25(4):535–543, August 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s7/p535.

Veevers:1998:VCI

[920] Alan Veevers. Viability and capability indexes for multi-response processes. Journal of Applied Statistics, 25(4):545–558, August 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s8/p545.

Zhang:1998:EPC

[921] Nien Fan Zhang. Estimating process capability indexes for autocorrelated data. Journal of Applied Statistics, 25(4):559–574, August 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s9/p559.

Anonymous:1998:BRa

[922] Anonymous. Book reviews. Journal of Applied Statistics, 25(4):575–??, August 1998. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n4/
s10/p575.

Ghatak:1998:VAM

[923] Anita Ghatak. Vector autoregression modelling and forecasting growth of South Korea. Journal of Applied Statistics, 25(5):579–592, October 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s1/p579.

Matejic:1998:TBA

[924] Larisa Matejic. Testing for brain anomalies: a hippocampus study. Journal of Applied Statistics, 25(5): 593–600, October 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s2/p593.

Hosseini:1998:IOH

[925] M. Hosseini, R. G. Carpenter, and K. Mohammad. Identification of outlying height and weight data in the Iranian National Health Survey 1990– 92. Journal of Applied Statistics, 25 (5):601–612, October 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s3/p601.

Prescott:1998:MDC

[926] Philip Prescott and Norman R. Draper. Mixture designs for constrained components in orthogonal blocks. Journal of Applied Statistics, 25(5):613– 638, October 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s4/p613.

Carpenter:1998:APU

[927] James Carpenter. Assessing parameter uncertainty via bootstrap likelihood ratio confidence regions. Journal of Applied Statistics, 25(5):639– 649, October 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s5/p639.

Reed:1998:CAE

[928] James F. Reed III. Contributions to adaptive estimation. Journal of Applied Statistics, 25(5):651– 669, October 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s6/p651.

Baxter:1998:TMN

[929] M. J. Baxter and N. H. Gale. Testing for multivariate normality via univariate tests: a case study using lead isotope ratio data. Journal of Applied Statistics, 25(5):671– 683, October 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s7/p671.

Ribic:1998:EAM

[930] Christine A. Ribic and Thomas W. Miller. Evaluation of alternative model selection criteria in the analysis of unimodal response curves using CART. Journal of Applied Statistics, 25(5): 685–698, October 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n5/
                                        s8/p685.

Gunter:1998:CTV

[931] Christine Gunter, Colin Rallings, and Michael Thrasher. Calculating the total vote where the district magnitude is greater than one: a test of some algorithms using British local election data. Journal of Applied Statistics, 25(5):699–706, October 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n5/
                                        s9/p699.

Balasooriya:1998:RSP

[932] Uditha Balasooriya and Sutaip L. C. Saw. Reliability sampling plans for the two-parameter exponential distribution under progressive censoring. Journal of Applied Statistics, 25 (5):707–714, October 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n5/
                                        s10/p707.

Anonymous:1998:BRb

[933] Anonymous. Book reviews. Journal of Applied Statistics, 25(5):715–??, October 1998. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n5/
s11/p715.

Brajkovic:1998:MME

[934] Vladimir Brajkovic. Mechanics of microelectronics examined by design of experiments techniques. Journal of Applied Statistics, 25(6):723– 731, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s1/p723.

Wang:1998:MAF

[935] Reay-Chen Wang and Chung-Ho Chen. Minimum average fraction inspected for short-run CSP-1 plan. Journal of Applied Statistics, 25(6):733– 738, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s2/p733.

Chen:1998:IPM

[936] Ling Chen. Improved penalized mean for estimating the mean concentration of contaminants. Journal of Applied Statistics, 25(6):739– 750, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s3/p739.

Clarke:1998:OPD

[937] Brenton R. Clarke and Toby Lewis. An outlier problem in the determination of ore grade. Journal of Applied Statistics, 25(6):751–762, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s4/p751.

DeLuna:1998:PPA

[938] Xavier De Luna. Projected polynomial autoregression for prediction of stationary time series. Journal of Applied Statistics, 25(6):763– 775, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s5/p763.

Modarres:1998:HTH

[939] Reza Modarres and Joseph L. Gastwirth. Hybrid test for the hypothesis of symmetry. Journal of Applied Statistics, 25(6):777–783, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s6/p777.

Ojima:1998:GFE

[940] Yoshikazu Ojima. General formulae for expectations, variances and covariances of the mean squares for staggered nested designs. Journal of Applied Statistics, 25(6):785– 799, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s7/p785.

Pearn:1998:NGP

[941] W. L. Pearn and K. S. Chen. New generalization of process capability index C_{pk}. Journal of Applied Statistics, 25 (6):801–810, December 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s8/p801.

Upton:1998:RH

[942] Graham Upton. Rounding halves. Journal of Applied Statistics, 25(6): 811–816, December 1, 1998. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v25n6/
s9/p811.

Montepiedra:1998:AGA

[943] Grace Montepiedra, Danny Myers, and Arthur B. Yeh. Application of genetic algorithms to the construction of exact D-optimal designs. Journal of Applied Statistics, 25(6):817– 826, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n6/
                                        s10/p817.

Ghosh:1998:RCD

[944] D. K. Ghosh and Naimesh R. Desai. Robustness of complete diallel crosses plans to the unavailability of one block. Journal of Applied Statistics, 25(6):827–837, December 1, 1998. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n6/
                                        s11/p827.

Anonymous:1998:BRc

[945] Anonymous. Book reviews. Journal of Applied Statistics, 25(6):839–??, December 1998. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v25n6/
                                        s12/p839.

Anonymous:1998:TPV

[946] Anonymous. Title-page, volume contents and author index, 1998. Journal of Applied Statistics, 25(6):843–??, December 1998. ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:1999:MCM

[947] Saling Huang and Morton B. Brown. A Markov chain model for longitudinal categorical data when there may be non-ignorable non-response. Journal of Applied Statistics, 26(1): 5–18, January 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n1/
s1/p5.

Chan:1999:OOB

[948] L. Y. Chan and M. K. Sandhu. Optimal orthogonal block designs for a quadratic mixture model for three components. Journal of Applied Statistics, 26(1):19–34, January 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n1/
s2/p19.

Chan:1999:EJF

[949] Wai-Sum Chan, Siu Hung Cheung, and Ka Ho Wu. Exact joint forecast regions for vector autoregressive models. Journal of Applied Statistics, 26(1):35–44, January 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n1/
s3/p35.

Gomez:1999:NPE

[950] Guadalupe Gómez and M. Luz Calle. Non-parametric estimation with doubly censored data. Journal of Applied Statistics, 26(1):45– 58, January 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n1/
                                        s4/p45.

Krzanowski:1999:AMA

[951] W. J. Krzanowski. Antedependence models in the analysis of multigroup high-dimensional data. Journal of Applied Statistics, 26(1):59– 67, January 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n1/
                                        s5/p59.

Martin:1999:PCM

[952] Donald E. K. Martin and James F. Troendle. Paired comparison models applied to the design of the Major League baseball play-offs. Journal of Applied Statistics, 26(1):69– 80, January 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n1/
                                        s6/p69.

Meeden:1999:IEP

[953] Glen Meeden. Interval estimators for the population mean for skewed distributions with a small sample size. Journal of Applied Statistics, 26(1): 81–96, January 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v26n1/
s7/p81.

Geisser:1999:RBM

[954] Seymour Geisser and Teddy Seidenfeld. Remarks on the ‘Bayesian’ method of moments. Journal of Applied Statistics, 26(1):97–101, January 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n1/

s8/p97. See remarks [1178].

Young:1999:RBA

[955] Peter Young and Diego Pedregal. Recursive and en-bloc approaches to signal extraction. Journal of Applied Statistics, 26(1):103– 128, January 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n1/
s9/p103.

Ramalhoto:1999:SCC

[956] M. F. Ramalhoto and M. Morais. Shewart control charts for the scale parameter of a Weibull control variable with fixed and variable sampling intervals. Journal of Applied Statistics, 26(1):129–160, January 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n1/
s10/p129.

Adams:1999:PPH

[957] Joseph Brian Adams. Predicting pickle harvests using a parametric feed-forward neural network. Journal of Applied Statistics, 26(2):165– 176, February 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n2/
                                        s1/p165.

Billor:1999:ALI

[958] Nedret Billor and R. M. Loynes. An application of the local influence approach to ridge regression. Journal of Applied Statistics, 26(2):177– 183, February 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n2/
                                        s2/p177.

Kalirajan:1999:SVC

[959] K. Kalirajan. Stochastic varying coefficients gravity model: an application in trade analysis. Journal of Applied Statistics, 26(2):185– 193, February 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n2/
                                        s3/p185.

Kaushik:1999:PDC

[960] L. S. Kaushik. Partial diallel crosses based on three associate class association schemes. Journal of Applied Statistics, 26(2):195–201, February 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n2/
s4/p195.

Richter:1999:NET

[961] Scott J. Richter and Mark E. Payton. Nearly exact tests in factorial experiments using the aligned rank transform. Journal of Applied Statistics, 26 (2):203–217, February 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n2/
s5/p203.

Hughes:1999:ACD

[962] James W. Hughes and Elizabeth Savoca. Accounting for censoring in duration data: an application to estimating the effect of legal reforms on the duration of medical malpractice disputes. Journal of Applied Statistics, 26(2):219–228, February 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n2/
s6/p219.

Vijayaraghavan:1999:PSC

[963] R. Vijayaraghavan. Procedure for the selection of CSP-M one-level skip-lot sampling inspection plans that have a single sampling plan with acceptance number zero as the reference plan. Journal of Applied Statistics, 26 (2):229–233, February 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n2/
                                        s7/p229.

Holm:1999:SEL

[964] Sture Holm and Kerstin Wiklander. Simultaneous estimation of location and dispersion in two-level fractional factorial designs. Journal of Applied Statistics, 26(2):235–242, February 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n2/
                                        s8/p235.

Wludyka:1999:TNP

[965] Peter S. Wludyka and Peter R. Nelson. Two non-parametric, analysis-of-means-type tests for homogeneity of variances. Journal of Applied Statistics, 26(2):243–256, February 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n2/
                                        s9/p243.

Yau:1999:ESE

[966] K. K. W. Yau, S. K. Ng, M. T. Cheung, and M. C. Tung. Estimation of surgeon effects in the analysis of post-operative colorectal cancer patients data. Journal of Applied Statistics, 26(2):257–272, February 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n2/
s10/p257.

Harrison:1999:SPC

[967] P. J. Harrison and I. C. H. Lai. Statistical process control and model monitoring. Journal of Applied Statistics, 26(2):273–292, February 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n2/
s11/p273.

Bandyopadhyay:1999:STN

[968] Uttam Bandyopadhyay and Atanu Biswas. Sequential-type nonparametric test using Mann–Whitney statistics. Journal of Applied Statistics, 26(3):301–308, March 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n3/
s1/p301.

Wang:1999:DVS

[969] Reay-Chen Wang and Chung-Ho Chen. Designing a variable sampling plan based on Taguchi’s loss function. Journal of Applied Statistics, 26(3): 309–313, March 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n3/
s2/p309.

Galliani:1999:QTB

[970] G. Galliani, F. Filippini, and F. Screpanti. A queuing-theory-based approach to evaluate the efficiency of a network of automated stations and of a communication system. Journal of Applied Statistics, 26(3):315– 326, March 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n3/
                                        s3/p315.

Pasaribu:1999:SAU

[971] U. S. Pasaribu, A. G. Hawkes, and S. J. Wainwright. Statistical assumptions underlying the fitting of the Michaëlis-Menten equation. Journal of Applied Statistics, 26(3):327– 341, March 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n3/
                                        s4/p327.

Hutson:1999:CNC

[972] Alan D. Hutson. Calculating nonparametric confidence intervals for quantiles using fractional order statistics. Journal of Applied Statistics, 26(3):343–353, March 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n3/
                                        s5/p343.

Isogai:1999:PTD

[973] Takafumi Isogai. Power transformation of the F distribution and a power normal family. Journal of Applied Statistics, 26(3):355–371, March 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n3/
s6/p355.

Mardia:1999:ET

[974] K. V. Mardia, R. J. Morris, A. N. Walder, and J. J. Koenderink. Estimation of torsion. Journal of Applied Statistics, 26(3):373– 381, March 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n3/
s7/p373.

Poon:1999:SHD

[975] Wai-Yin Poon and Chi-Ming Lee. Sources of heterogeneity in distributions with ordered categorical variables. Journal of Applied Statistics, 26(3):383–392, March 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n3/
s8/p383.

Rodriguez-Carvajal:1999:MCT

[976] Luis A. Rodriguez-Carvajal and Geoff H. Freeman. Multivariate AB − BA crossover trial. Journal of Applied Statistics, 26(3):393– 403, March 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n3/
s9/p393.

Chen:1999:MCM

[977] Pai-Lien Chen, Estrada J. Bernard, and Pranab K. Sen. A Markov chain model used in analyzing disease history applied to a stroke study. Journal of Applied Statistics, 26(4): 413–422, May 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n4/
s1/p413.

Ghosh:1999:CPP

[978] Kaushik Ghosh, S. Rao Jammalamadaka, and Mangalam Vasudaven. Change-point problems for the von Mises distribution. Journal of Applied Statistics, 26(4):423–434, May 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n4/
s2/p423.

Cassel:1999:RPL

[979] Claes Cassel, Peter Hackl, and Anders H. Westlund. Robustness of partial least-squares method for estimating latent variable quality structures. Journal of Applied Statistics, 26(4):435–446, May 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n4/
s3/p435.

Lillegard:1999:ECI

[980] Magnar Lillegard and Steinar Engen. Exact confidence intervals generated by conditional parametric bootstrapping. Journal of Applied Statistics, 26(4):447–459, May 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n4/
s4/p447.

Magnus:1999:FST

[981] Jan R. Magnus and Franc J. G. M. Klaassen. The final set in a tennis match: four years at Wimbledon. Journal of Applied Statistics, 26(4):461–468, May 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n4/
s5/p461.

Mingoti:1999:BET

[982] Sueli A. Mingoti. Bayesian estimator for the total number of distinct species when quadrat sampling is used. Journal of Applied Statistics, 26(4):469–483, May 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

                                        catchword/carfax/13600532/v26n4/
                                        s6/p469.

RibeiroRamos:1999:UPP

[983] Francisco Fernando Ribeiro Ramos. Underreporting of purchases of port wine. Journal of Applied Statistics, 26(4):485–494, May 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n4/

s7/p485. See correction [1167].

Schoen:1999:DFT

[984] Eric D. Schoen. Designing fractional two-level experiments with nested error structures. Journal of Applied Statistics, 26(4):495–508, May 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n4/
                                        s8/p495.

Walker:1999:UPD

[985] Stephen Walker. The uniform power distribution. Journal of Applied Statistics, 26(4):509–517, May 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n4/
                                        s9/p509.

Zhang:1999:OTN

[986] Paul Zhang. Omnibus test of normality using the Q statistic. Journal of Applied Statistics, 26(4): 519–528, May 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n4/
s10/p519.

Anonymous:1999:BRa

[987] Anonymous. Book reviews. Journal of Applied Statistics, 26(4):529–??, May 1999. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n4/
s11/p529.

Buning:1999:AJT

[988] Herbert Büning. Adaptive Jonckheeretype tests for ordered alternatives. Journal of Applied Statistics, 26(5): 541–551, July 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n5/
s1/p541.

Dechert:1999:EMC

[989] Jerry D. Dechert and Kenneth E. Case. An economic model for clinical quality control. Journal of Applied Statistics, 26(5):553–562, July 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n5/
s2/p553.

Ghosh:1999:RCD

[990] D. K. Ghosh and Naimesh R. Desai. Robustness of a complete diallel crosses plan with an unequal number of crosses to the unavailability of one block. Journal of Applied Statistics, 26(5):563–577, July 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n5/
                                        s3/p563.

Taylor:1999:FSA

[991] Paul C. Taylor and David J. Hand. Finding ‘superclassifications’ with an acceptable misclassification rate. Journal of Applied Statistics, 26(5): 579–590, July 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n5/
                                        s4/p579.

Midi:1999:PER

[992] Habshah Midi. Preliminary estimators for robust non-linear regression estimation. Journal of Applied Statistics, 26(5):591–600, July 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n5/
                                        s5/p591.

RuizdeVilla:1999:NPR

[993] M. Carme Ruiz de Villa, M. Salomé E. Cabral, Eduardo Escrich Escriche, and Montse Solanas. A nonparametric regression approach to repeated measures analysis in cancer experiments. Journal of Applied Statistics, 26(5):601–611, July 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n5/
s6/p601.

Scallan:1999:RMI

[994] A. J. Scallan. Regression modelling of interval-censored failure time data using the Weibull distribution. Journal of Applied Statistics, 26(5): 613–618, July 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n5/
s7/p613.

Schmidt:1999:ARS

[995] Alexandra Mello Schmidt, Dani Gamerman, Ajax R. B. Moreira, Pnpe-Rj, Im-Ufrj, and Ipea-Rj. An adaptive resampling scheme for cycle estimation. Journal of Applied Statistics, 26(5):619–641, July 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n5/
s8/p619.

Aidid:1999:MMS

[996] Sharifah Sakinah Aidid and Mick Silver. Modelling market shares by segments using volatility. Journal of Applied Statistics, 26(5): 643–660, July 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n5/
                                        s9/p643.

Achcar:1999:UEP

[997] Jorge Alberto Achcar and Gilberto de Arañjo Pereira. Use of exponential power distributions for mixture models in the presence of covariates. Journal of Applied Statistics, 26(6):669–679, August 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n6/
                                        s1/p669.

Dewoody:1999:ARR

[998] Yssa Dewoody, V. T. Gururaj, and Clyde Martin. Assessing risk for rare events. Journal of Applied Statistics, 26(6):681–687, August 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n6/
                                        s2/p681.

Duchateau:1999:DRS

[999] L. Duchateau, D. L. Berkvens, and G. J. Rowlands. Decision rules for small vaccine experiments with binary outcomes based on conditional and expected power and size of the Fisherexact test. Journal of Applied Statistics, 26(6):689–699, August 1, 1999.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n6/
s3/p689.

Fang:1999:PCC

[1000] Yue Fang and John Zhang. Performance of control charts for autoregressive conditional heteroscedastic processes. Journal of Applied Statistics, 26(6):701–714, August 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n6/
s4/p701.

Harris:1999:IEM

[1001] Peter Harris, Mark Hann, Simon P. J.

Kirby, and John C. Dearden. Interval estimation of the median effective dose for a logistic dose-response curve. Journal of Applied Statistics, 26(6):715–722, August 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n6/
s5/p715.

Lund:1999:LCD

[1002] Ulric Lund. Least circular distance regression for directional data. Journal of Applied Statistics, 26(6): 723–733, August 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

                                        catchword/carfax/13600532/v26n6/
                                        s6/p723.

Mardia:1999:ASS

[1003] K. V. Mardia, A. N. Walder, E. Berry, D. Sharples, P. A. Millner, and R. A. Dickson. Assessing spinal shape. Journal of Applied Statistics, 26(6): 735–745, August 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n6/
                                        s7/p735.

Scallan:1999:FMD

[1004] A. J. Scallan. Fitting a mixture distribution to complex censored survival data using generalized linear models. Journal of Applied Statistics, 26 (6):747–753, August 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n6/
                                        s8/p747.

Shao:1999:IAT

[1005] Yuehjen E. Shao, Yue-Fa Lin, and Soe-Tsyr Yuan. Integrated application of time series multiple-interventions analysis and knowledge-based reasoning. Journal of Applied Statistics, 26 (6):755–766, August 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n6/
                                        s9/p755.

Tajuddin:1999:CBT

[1006] I. H. Tajuddin. A comparison between two simple measures of skewness. Journal of Applied Statistics, 26(6):767–774, August 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n6/
s10/p767.

Anonymous:1999:VNS

[1007] Anonymous. Vol. 26 no. 7 September 1999 contents. Journal of Applied Statistics, 26(7):1–??, September 1999. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

catchword.co.uk/cgi-bin/cgi?ini=
carfax&body=linker&reqidx=1360-
0532%28%290%3A0L.1.

Binongo:1999:BBS

[1008] José Nilo G. Binongo and M. W. A.

Smith. A bridge between statistics and literature: the graphs of Oscar Wilde’s literary genres. Journal of Applied Statistics, 26(7):781– 787, September 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s1/p781.

Chen:1999:SEM

[1009] Zhenmin Chen. A simple exact method for testing hypotheses about the shape parameter of a log-normal distribution. Journal of Applied Statistics, 26 (7):789–805, September 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s2/p789.

DeBruin:1999:FPT

[1010] Paul De Bruin and Philip Hans Franses. Forecasting power-transformed time series data. Journal of Applied Statistics, 26(7):807–815, September 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s3/p807.

Dodge:1999:SGB

[1011] Yadolah Dodge and Ali S. Hadi. Simple graphs and bounds for the elements of the hat matrix. Journal of Applied Statistics, 26(7):817– 823, September 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s4/p817.

Lehtonen:1999:DTC

[1012] T. Lehtonen and J.-O. Malmberg. Do two competing frequencies differ significantly? Journal of Applied Statistics, 26(7):825–830, September 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s5/p825.

Ostermark:1999:SCC

[1013] Ralf Østermark and Rune Höglund.

Simulating competing cointegration tests in a bivariate system. Journal of Applied Statistics, 26(7):831– 846, September 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s6/p831.

Pigeon:1999:CNA

[1014] Joseph G. Pigeon and Joseph F. Heyse.

A cautionary note about assessing the fit of logistic regression models. Journal of Applied Statistics, 26(7): 847–853, September 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s7/p847.

Wang:1999:MPRa

[1015] Peiming Wang and Martin L. Puterman. Markov Poisson regression models for discrete time series. Part 1: Methodology. Journal of Applied Statistics, 26(7):855– 869, September 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s8/p855.

Wang:1999:MPRb

[1016] Peiming Wang and Martin L. Puterman. Markov Poisson regression models for discrete time series. Part 2: Applications. Journal of Applied Statistics, 26(7):871–882, September 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n7/
s9/p871.

Balamurali:1999:DCD

[1017] S. Balamurali and M. Kalyanasundaram. Determination of conditional double sampling scheme. Journal of Applied Statistics, 26(8):893– 902, December 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s1/p893.

Chen:1999:SLU

[1018] Chung-Ho Chen. Specification limit under a quality loss function. Journal of Applied Statistics, 26(8):903– 908, December 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s2/p903.

Cooper:1999:ANN

[1019] John C. B. Cooper. Artificial neural networks versus multivariate statistics: an application from economics. Journal of Applied Statistics, 26(8): 909–921, December 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s3/p909.

Horgan:1999:UWD

[1020] Graham W. Horgan. Using wavelets for data smoothing: a simulation study. Journal of Applied Statistics, 26(8): 923–932, December 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s4/p923.

Lee:1999:GTC

[1021] Juneyoung Lee and Andre I. Khuri.

Graphical technique for comparing designs for random models. Journal of Applied Statistics, 26(8):933– 947, December 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s5/p933.

Mardia:1999:DSS

[1022] K. V. Mardia. Directional statistics and shape analysis. Journal of Applied Statistics, 26(8):949– 957, December 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s6/p949.

Mira:1999:DFT

[1023] Antonietta Mira. Distribution-free test for symmetry based on Bonferroni’s measure. Journal of Applied Statistics, 26(8):959–972, December 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n8/
                                        s7/p959.

Okafor:1999:UEB

[1024] Ray Okafor. Using an empirical Bayes model to estimate currency exchange rate. Journal of Applied Statistics, 26 (8):973–983, December 1, 1999. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n8/
                                        s8/p973.

Rodrigues:1999:PSU

[1025] Paulo M. M. Rodrigues and Denise R. Osborn. Performance of seasonal unit root tests for monthly data. Journal of Applied Statistics, 26(8):985– 1004, December 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v26n8/
                                        s9/p985.

Papalia:1999:LGM

[1026] Rosa Bernardini Papalia. Local generalized method of moments estimation based on kernel weights: an application to panel data. Journal of Applied Statistics, 26(8):1005– 1015, December 1, 1999. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s10/p1005.

Anonymous:1999:BRb

[1027] Anonymous. Book reviews. Journal of Applied Statistics, 26(8):1017–??, December 1999. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v26n8/
s11/p1017.

Baker:2000:AND

[1028] R. D. Baker. Application of a new discrete distribution. Journal of Applied Statistics, 27(1):5–21, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s1/p5.

Balamurali:2000:GTT

[1029] S. Balamurali and M. Kalyanasundaram. Generalized tightened twolevel continuous sampling plans. Journal of Applied Statistics, 27(1):23– 38, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s2/p23.

Arslan:2000:RLE

[1030] Olcay Arslan and Nedret Billor. Robust Liu estimator for regression based on an M-estimator. Journal of Applied Statistics, 27(1):39– 47, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n1/
                                        s3/p39.

Divecha:2000:SSI

[1031] Jyoti Divecha. Search for suitable incomplete block designs for complete diallel cross systems. Journal of Applied Statistics, 27(1):49– 62, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n1/
                                        s4/p49.

Simon:2000:IEE

[1032] Yangxin Huang Simon, P. J. Kirby Peter, Harris John, and C. Dearden. Interval estimation of the 90% effective dose: a comparison of bootstrap resampling methods with some large-sample approaches. Journal of Applied Statistics, 27(1):63–73, January 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n1/
                                        s5/p63.

Isogai:2000:AFE

[1033] Takafumi Isogai. Analysis of factorial experiments for survival data with long-tailed distributions. Journal of Applied Statistics, 27(1):75– 101, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s6/p75.

Lai:2000:FBM

[1034] Dejian Lai, Barry R. Davis, and Robert J. Hardy. Fractional Brownian motion and clinical trials. Journal of Applied Statistics, 27(1):103– 108, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s7/p103.

Nelder:2000:FMR

[1035] J. A. Nelder. Functional marginality and response-surface fitting. Journal of Applied Statistics, 27(1):109– 112, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s8/p109.

Reschenhofer:2000:MAF

[1036] Erhard Reschenhofer. Modification of autoregressive fractionally integrated moving average models for the estimation of persistence. Journal of Applied Statistics, 27(1):113– 118, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s9/p113.

David:2000:ELE

[1037] Jon Vilasuso David and Katz. Estimates of the likelihood of extreme returns in international stock markets. Journal of Applied Statistics, 27 (1):119–130, January 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s10/p119.

Anonymous:2000:BRa

[1038] Anonymous. Book reviews. Journal of Applied Statistics, 27(1):131–??, January 2000. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n1/
s11/p131.

Aadland:2000:DIU

[1039] David M. Aadland. Distribution and interpolation using transformed data. Journal of Applied Statistics, 27(2): 141–156, February 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v27n2/
s1/p141.

Ayhan:2000:EVE

[1040] H. Öztay Ayhan. Estimators of vital events in dual-record systems. Journal of Applied Statistics, 27(2): 157–169, February 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n2/
s2/p157.

Kalyanasundaram:2000:DVL

[1041] M. Kalyanasundaram and S. Balamurali. Determination of variable-lot-size attribute sampling plan indexed by the acceptable quality level and average outgoing quality level for continuous production. Journal of Applied Statistics, 27(2):171–175, February 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n2/
s3/p171.

George:2000:ISI

[1042] Alberto Luceno George and E. P.

Box. Influence of the sampling interval, decision limit and autocorrelation on the average run length in Cusum charts. Journal of Applied Statistics, 27(2):177–183, February 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n2/
s4/p177.

Guerrero:2000:SLP

[1043] Victor M. Guerrero. Selecting a linearizing power transformation for time series. Journal of Applied Statistics, 27(2):185–195, February 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n2/
                                        s5/p185.

Wei-Ming:2000:NJT

[1044] Jiin-Huarng Guo Wei-Ming and Luh. Normalized Johnson’s transformation one-sample trimmed t for nonnormality. Journal of Applied Statistics, 27(2):197–203, February 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n2/
                                        s6/p197.

Southworth:2000:TLI

[1045] R. Southworth, K. V. Mardia, and C. C. Taylor. Transformation- and label-invariant neural network for the classification of landmark data. Journal of Applied Statistics, 27(2):205– 215, February 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n2/
                                        s7/p205.

Myers:2000:RCA

[1046] Leann Myers and Stephanie T. Broyles. Regression coefficient analysis for correlated binomial outcomes. Journal of Applied Statistics, 27(2):217– 234, February 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n2/
s8/p217.

Nummi:2000:AGC

[1047] Tapio Nummi. Analysis of growth curves under measurement errors. Journal of Applied Statistics, 27(2): 235–243, February 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n2/
s9/p235.

TyrchaR:2000:SMS

[1048] J. Tyrcha R., Sundberg P., Lindskog B., and Sundstrom. Statistical modelling and saddle-point approximation of tail probabilities for accumulated splice loss in fibre-optic networks. Journal of Applied Statistics, 27(2):245–256, February 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n2/
s10/p245.

Anonymous:2000:BRb

[1049] Anonymous. Book reviews. Journal of Applied Statistics, 27(2):257–??, February 2000. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?

                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n2/
                                        s11/p257.

Amaral:2000:DPH

[1050] J. A. Amaral, M. B. Rosario, and M. T. Paixao. Data and projections of HIV and AIDS in Portugal. Journal of Applied Statistics, 27(3):269–279, March 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n3/
                                        s1/p269.

Wong:2000:SPM

[1051] Ying Wang Wong and Siu Hung Cheung. Simultaneous pairwise multiple comparisons in a two-way analysis of covariance model. Journal of Applied Statistics, 27(3):281– 291, March 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n3/
                                        s2/p281.

Jung:2000:LIA

[1052] Kang-Mo Jung. Local influence assessment in canonical correlation analysis. Journal of Applied Statistics, 27(3):293–301, March 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n3/
                                        s3/p293.

Kvist:2000:UCR

[1053] Trine Kvist, Henrik Gislason, and Poul Thyregod. Using continuationratio logits to analyze the variation of the age composition of fish catches. Journal of Applied Statistics, 27(3): 303–319, March 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n3/
s4/p303.

Lee:2000:BAG

[1054] Jack C. Lee and Kuo-Ching Liu.

Bayesian analysis of a general growth curve model with predictions using power transformations and AR(1) autoregressive dependence. Journal of Applied Statistics, 27(3):321– 336, March 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n3/
s5/p321.

Su:2000:UMR

[1055] Yeong-Tzay Su and Chyi-Lyi Kathleen Liang. Using multivariate rank sum tests to evaluate effectiveness of computer applications in teaching business statistics. Journal of Applied Statistics, 27(3):337–345, March 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n3/
s6/p337.

Mvoi:2000:CEV

[1056] Sifa Mvoi and Yan-Xia Lin. Criteria for estimating the variance function used in the asymptotic quasi-likelihood approach. Journal of Applied Statistics, 27(3):347–362, March 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n3/
                                        s7/p347.

Sawiris:2000:OGB

[1057] Milad Sawiris. Optimum grouping and the boundary problem. Journal of Applied Statistics, 27(3):363– 371, March 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n3/
                                        s8/p363.

Vieira:2000:ZIP

[1058] A. M. C. Vieira, J. P. Hinde, and C. G. B. Demetrio. Zero-inflated proportion data models applied to a biological control assay. Journal of Applied Statistics, 27(3):373– 389, March 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n3/
                                        s9/p373.

Balamurali:2000:MTT

[1059] S. Balamurali and K. Govindaraju. Modified tightened two-level continuous sampling plans. Journal of Applied Statistics, 27(4):397–409, May 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n4/
s1/p397.

Chang:2000:LRM

[1060] Bei-Hung Chang, Stuart Lipsitz, and Christine Waternaux. Logistic regression in meta-analysis using aggregate data. Journal of Applied Statistics, 27(4):411–424, May 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n4/
s2/p411.

Johansson:2000:MSS

[1061] Jan-Olof Johansson. Modelling the surface structure of newsprint. Journal of Applied Statistics, 27(4): 425–438, May 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n4/
s3/p425.

Khamis:2000:TSC

[1062] H. J. Khamis. The two-stage δcorrected Kolmogorov–Smirnov test. Journal of Applied Statistics, 27(4): 439–450, May 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/

                                        catchword/carfax/13600532/v27n4/
                                        s4/p439.

Kim:2000:OIO

[1063] Myung Geun Kim. Outliers and influential observations in the structural errors-in-variables model. Journal of Applied Statistics, 27(4): 451–460, May 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n4/
                                        s5/p451.

Lai:2000:SMM

[1064] C. D. Lai, M. Xie, and K. Govindaraju. Study of a Markov model for a high-quality dependent process. Journal of Applied Statistics, 27(4):461–473, May 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n4/
                                        s6/p461.

Shalabh:2000:PVV

[1065] Shalabh. Prediction of values of variables in linear measurement error model. Journal of Applied Statistics, 27(4):475–482, May 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n4/
                                        s7/p475.

Shoukri:2000:SSR

[1066] M. M. Shoukri and O. Demirkaya.

Sample size requirements to test the equality of raters’ precision. Journal of Applied Statistics, 27(4): 483–494, May 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n4/
s8/p483.

Wang:2000:AGM

[1067] Zheng Wang. An algorithm for generalized monotonic smoothing. Journal of Applied Statistics, 27(4): 495–507, May 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n4/
s9/p495.

Woodroof:2000:BE

[1068] Jon Woodroof. Bootstrapping: as easy as 1–2–3. Journal of Applied Statistics, 27(4):509–517, May 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n4/
s10/p509.

Anonymous:2000:BRc

[1069] Anonymous. Book reviews. Journal of Applied Statistics, 27(4):519–??, May 2000. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?

                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n4/
                                        s11/p519.

Hand:2000:DAS

[1070] David J. Hand and Niall M. Adams. Defining attributes for scorecard construction in credit scoring. Journal of Applied Statistics, 27(5): 527–540, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n5/
                                        s1/p527.

Ojima:2000:GSN

[1071] Yoshikazu Ojima. Generalized staggered nested designs for variance components estimation. Journal of Applied Statistics, 27(5):541–553, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n5/
                                        s2/p541.

Yilmaz:2000:PNT

[1072] Mustafa R. Yilmaz and Sangit Chatterjee. Patterns of NBA team performance from 1950 to 1998. Journal of Applied Statistics, 27(5):555–566, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n5/
                                        s3/p555.

Hutson:2000:CQF

[1073] Alan D. Hutson. A composite quantile function estimator with applications in bootstrapping. Journal of Applied Statistics, 27(5):567–577, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s4/p567.

Li:2000:ABM

[1074] Hui Li and Robert A. Malkin. An approximate Bayesian up-down method for estimating a percentage point on a dose-response curve. Journal of Applied Statistics, 27(5):579–587, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s5/p579.

Singh:2000:SET

[1075] Murari Singh and Michael J. Jones.

Statistical estimation of time-trends in two-course crop rotations. Journal of Applied Statistics, 27(5): 589–597, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s6/p589.

Mudholkar:2000:CRS

[1076] Govind S. Mudholkar and Deo Kumar Srivastava. A class of robust stepwise alternatives to Hotelling’s T^{2} tests. Journal of Applied Statistics, 27(5):599–619, July 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s7/p599.

Zou:2000:TTM

[1077] Kelly H. Zou and W. J. Hall. Two transformation models for estimating an ROC curve derived from continuous data. Journal of Applied Statistics, 27(5):621–631, July 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s8/p621.

Huffer:2000:TMS

[1078] Fred W. Huffer and Cheolyong Park. A test for multivariate structure. Journal of Applied Statistics, 27(5): 633–650, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s9/p633.

Mingoti:2000:SBE

[1079] Sueli A. Mingoti. A stepwise Bayesian estimator for the total number of distinct species in finite populations: sampling by elements. Journal of Applied Statistics, 27(5):651–670, July 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?

ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s10/p651.

Anonymous:2000:BRd

[1080] Anonymous. Book reviews. Journal of Applied Statistics, 27(5):671–??, July 2000. ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n5/
s11/p671.

Abrevaya:2000:TTE

[1081] Jason Abrevaya. Testing for a treatment effect in a heterogeneous population: a modified sign-test statistic and a leapfrog statistic. Journal of Applied Statistics, 27(6):679– 687, August 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n6/
s1/p679.

Garcia:2000:OUC

[1082] J. M. Tapia García and A. Martín Andrés. Optimal unconditional critical regions for 2 × 2 multinomial trials. Journal of Applied Statistics, 27 (6):689–695, August 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n6/
s2/p689.

Chen:2000:DCP

[1083] Chung-Ho Chen and Chao-Yu Chou. Design of a CSP-1 plan based on regret-balanced criterion. Journal of Applied Statistics, 27(6):697– 701, August 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n6/
                                        s3/p697.

DeCorte:2000:UOS

[1084] Wilfried De Corte. Using order statistics to assess the sampling variability of personnel selection utility estimates. Journal of Applied Statistics, 27(6):703–713, August 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n6/
                                        s4/p703.

Ghosh:2000:RDD

[1085] D. K. Ghosh and P. C. Biswas. Robust designs for diallel crosses against the missing of one block. Journal of Applied Statistics, 27(6):715– 723, August 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n6/
                                        s5/p715.

Govindaraju:2000:CCL

[1086] K. Govindaraju and M. Bebbington. Combined continuous lot by lot acceptance sampling plan. Journal of Applied Statistics, 27(6):725– 730, August 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n6/
s6/p725.

Guo:2000:TMO

[1087] Jiin-Huarng Guo and Wei-Ming Luh.

Testing methods for the one-way fixed effects ANOVA models of log-normal samples. Journal of Applied Statistics, 27(6):731–738, August 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n6/
s7/p731.

Ingman:2000:LBO

[1088] Dov Ingman and Boris Lipnik. Lossbased optimal control statistics for control charts. Journal of Applied Statistics, 27(6):739–756, August 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n6/
s8/p739.

Kanji:2000:MRC

[1089] G. K. Kanji and Osama Hasan Arif. Median rankit control chart by the quantile approach. Journal of Applied Statistics, 27(6):757– 770, August 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n6/
                                        s9/p757.

Soundararajan:2000:QSV

[1090] V. Soundararajan and M. Palanivel. Quick switching variables single sampling (QSVSS) system indexed by AQL and AOQL. Journal of Applied Statistics, 27(6):771–778, August 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n6/
                                        s10/p771.

Simar:2000:GMB

[1091] Léopold Simar and Paul W. Wilson. A general methodology for bootstrapping in non-parametric frontier models. Journal of Applied Statistics, 27 (6):779–802, August 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n6/
                                        s11/p779.

Andres:2000:MEQ

[1092] A. Martín Andrés and I. Herranz Tejedor. On the minimum expected quantity for the validity of the chisquared test in 2 × 2 tables. Journal of Applied Statistics, 27(7):807– 820, September 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/

catchword/carfax/13600532/v27n7/
s1/p807.

Gonzalo:2000:RCT

[1093] Jesús Gonzalo and Tae-Hwy Lee. On the robustness of cointegration tests when series are fractionally integrated. Journal of Applied Statistics, 27(7): 821–827, September 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n7/
s2/p821.

Govindaraju:2000:DGC

[1094] K. Govindaraju and C. Kandasamy.

Design of generalized CSP-C continuous sampling plan. Journal of Applied Statistics, 27(7):829– 841, September 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n7/
s3/p829.

Luh:2000:ATT

[1095] Wei-Ming Luh and Jiin-Huarng Guo.

Approximate transformation trimmed mean methods to the test of simple linear regression slope equality. Journal of Applied Statistics, 27(7): 843–857, September 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n7/
s4/p843.

Pewsey:2000:PIA

[1096] Arthur Pewsey. Problems of inference for Azzalini’s skewnormal distribution. Journal of Applied Statistics, 27(7):859–870, September 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n7/
                                        s5/p859.

Tang:2000:TLD

[1097] Man-Lai Tang. On tests of linearity for dose response data: asymptotic, exact conditional and exact unconditional tests. Journal of Applied Statistics, 27 (7):871–880, September 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n7/
                                        s6/p871.

Salter:2000:SBP

[1098] Stephen Salter and Neville Topham. Side betting and playing the National Lottery: an exercise in policy design. Journal of Applied Statistics, 27(7): 881–899, September 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

                                        www.catchword.co.uk/cgi-bin/cgi?
                                        ini=carfax&body=linker&reqidx=/
                                        catchword/carfax/13600532/v27n7/
                                        s7/p881.

Vijayaraghavan:2000:DES

[1099] R. Vijayaraghavan. Design and evaluation of skip-lot sampling plans of type SkSP-3. Journal of Applied Statistics, 27(7):901–908, September 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n7/
s8/p901.

Wesolowsky:2000:DES

[1100] George O. Wesolowsky. Detecting excessive similarity in answers on multiple choice exams. Journal of Applied Statistics, 27(7):909– 921, September 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n7/
s9/p909.

Giglio:2000:GBS

[1101] Beatrice Giglio, Eva Riccomagno, and Henry P. Wynn. Gröbner basis strategies in regression. Journal of Applied Statistics, 27(7):923– 938, September 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://

www.catchword.co.uk/cgi-bin/cgi?
ini=carfax&body=linker&reqidx=/
catchword/carfax/13600532/v27n7/
s10/p923.

Chou:2000:ESD

[1102] Chao-Yu Chou, Chung-Ho Chen, and Hui-Rong Liu. Economic-statistical design of XY = charts for non-normal data by considering quality loss. Journal of Applied Statistics, 27(8):939– 951, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fader:2000:NMU

[1103] Peter S. Fader and Bruce G. S. Hardie. A note on modelling underreported Poisson counts. Journal of Applied Statistics, 27(8):953–964, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luh:2000:JTT

[1104] Wei-Ming Luh and Jiin-Huarng Guo. Johnson’s transformation two-sample trimmed t and its bootstrap method for heterogeneity and non-normality. Journal of Applied Statistics, 27(8): 965–973, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Magnussen:2000:UPS

[1105] Steen Magnussen. Unequal probability sampling in fixed area plots of stem volume with and without prior inclusion probabilities. Journal of Applied Statistics, 27(8):975–990, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mate:2000:ECR

[1106] Carlos Maté and Rafael Calderón. Exploring the characteristics of rotating electric machines with factor analysis. Journal of Applied Statistics, 27(8): 991–1006, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nelder:2000:QLP

[1107] J. A. Nelder. Quasi-likelihood and pseudo-likelihood are not the same thing. Journal of Applied Statistics, 27 (8):1007–1011, November 1, 2000. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sharma:2000:APD

[1108] M. K. Sharma. Application of PBIB designs in CDC Method IV. Journal of Applied Statistics, 27(8):1013– 1019, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shukur:2000:SIG

[1109] Ghazi Shukur and Panagiotis Mantalos. A simple investigation of the Granger-causality test in integratedcointegrated VAR systems. Journal of Applied Statistics, 27(8):1021– 1031, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tse:2000:SAW

[1110] Siu Keung Tse, Chunyan Yang, and Hak-Keung Yuen. Statistical analysis of Weibull distributed lifetime data under Type II progressive censoring with binomial removals. Journal of Applied Statistics, 27(8):1033–1043, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wright:2000:CLS

[1111] Peter A. Wright. Choosing a lower specification limit for an exponential process with ‘the larger the better’ tolerance: a simple, exact solution. Journal of Applied Statistics, 27(8):1045– 1049, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2000:PME

[1112] Zhenlin Yang and Min Xie. Process monitoring of exponentially distributed characteristics through an optimal normalizing transformation. Journal of Applied Statistics, 27(8): 1051–1063, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Awadhi:2001:PDA

[1113] S. A. Al-Awadhi and P. H. Garthwaite. Prior distribution assessment for a multivariate normal distribution: an experimental study. Journal of Applied Statistics, 28(1):5–23, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsay:2001:PSE

[1114] P. K. Tsay and A. Chao. Population size estimation for capture-recapture models with applications to epidemiological data. Journal of Applied Statistics, 28(1):25–36, January 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fuchs:2001:SMS

[1115] Camil Fuchs and Morton B. Brown. Summary measurements and screening in clinical trials with replicate observations. Journal of Applied Statistics, 28 (1):37–51, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gregori:2001:EFF

[1116] D. Gregori, C. Rocco, S. Miocic, and L. Mestroni. Estimating the frequency of familial dilated cardiomyopathy in the presence of misclassification errors. Journal of Applied Statistics, 28 (1):53–62, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lal:2001:RDE

[1117] Krishan Lal, V. K. Gupta, and Lalmohan Bhar. Robustness of designed experiments against missing data. Journal of Applied Statistics, 28(1):63–79, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lindsey:2001:ASS

[1118] P. J. Lindsey. Adapting sample size calculations to repeated measurements in clinical trials. Journal of Applied Statistics, 28(1):81–89, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Naik:2001:AMR

[1119] Dayanand N. Naik and Shantha S.

Rao. Analysis of multivariate repeated measures data with a Kronecker product structured covariance matrix. Journal of Applied Statistics, 28(1): 91–105, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Piccarreta:2001:NMN

[1120] Raffaella Piccarreta. A new measure of nominal—ordinal association. Journal of Applied Statistics, 28(1): 107–120, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kantam:2001:ASB

[1121] R. R. L. Kantam, K. Rosaiah, and G. Srinivasa Rao. Acceptance sampling based on life tests: log-logistic model. Journal of Applied Statistics, 28(1):121–128, January 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sengupta:2001:OTI

[1122] Ashis Sengupta and Chandranath Pal. On optimal tests for isotropy against the symmetric wrapped stable— circular uniform mixture family. Journal of Applied Statistics, 28(1):129– 143, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2001:BRa

[1123] Anonymous. Book reviews. Journal of Applied Statistics, 28(1):145–148, January 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bayazit:2001:UWD

[1124] Mehmetcik Bayazit and Hafzullah Aksoy. Using wavelets for data generation. Journal of Applied Statistics, 28 (2):157–166, February 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aparicio:2001:AEV

[1125] Teresa Aparicio and Inmaculada Villanua. The asymptotically efficient version of the information matrix test in binary choice models. A study of size and power. Journal of Applied Statistics, 28(2):167–182, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Congdon:2001:PAI

[1126] Peter Congdon. Predicting adverse infant health outcomes using routine screening variables: modelling the impact of interdependent risk factors. Journal of Applied Statistics, 28(2): 183–197, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gupta:2001:OPP

[1127] Abhijit Gupta. Optimization of product performance of a paint formulation using a mixture experiment. Journal of Applied Statistics, 28(2):199– 213, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Argac:2001:NTN

[1128] Dogan Argaç, Kepher H. Makambi, and Joachim Hartung. A note on testing the nullity of the between group variance in the one-way random effects model under variance heterogeneity. Journal of Applied Statistics, 28(2): 215–222, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2001:BAG

[1129] Jack C. Lee and W. H. Lien. Bayesian analysis of a growth curve model with power transformation, random effects and AR(1) dependence. Journal of Applied Statistics, 28(2):223–238, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2001:CWR

[1130] Youngjo Lee. Can we recover information from concordant pairs in binary matched pairs? Journal of Applied Statistics, 28(2):239–246, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martin:2001:IFA

[1131] Donald E. K. Martin. Influence functions applied to the estimation of mean rain rate. Journal of Applied Statistics, 28(2):247–258, February 1, 2001.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2001:BAT

[1132] Man-Suk Oh and Yong Bin Lim. Bayesian analysis of time series Poisson data. Journal of Applied Statistics, 28(2):259–271, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roddam:2001:AML

[1133] Andrew W. Roddam. An approximate maximum likelihood procedure for parameter estimation in multivariate discrete data regression models. Journal of Applied Statistics, 28(2): 273–279, February 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arvidsson:2001:IFI

[1134] M. Arvidsson, P. Kammerlind, A. Hynen, and B. Bergman. Identification of factors influencing dispersion in splitplot experiments. Journal of Applied Statistics, 28(3–4):269–283, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Box:2001:SD

[1135] George Box. Statistics for discovery. Journal of Applied Statistics, 28(3–4): 285–299, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Caulcutt:2001:WSS

[1136] Roland Caulcutt. Why is Six Sigma so successful? Journal of Applied Statistics, 28(3–4):301–306, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chambers:2001:STE

[1137] P. R. G. Chambers, J. L. Piggott, and S. Y. Coleman. SPC — a team effort for process improvement across four Area Control Centres. Journal of Applied Statistics, 28(3–4):307–324, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coleman:2001:STC

[1138] S. Y. Coleman, G. Arunakumar, F. Foldvary, and R. Feltham. SPC as a tool for creating a successful business measurement framework. Journal of Applied Statistics, 28(3–4):325–334, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bruce:2001:ICQ

[1139] David Bruce and Shirley Y. Coleman.

Improving communication via quantitative management. Journal of Applied Statistics, 28(3–4):335–341, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coleman:2001:SMI

[1140] S. Y. Coleman, A. Gordon, and P. R.

Chambers. SPC–making it work for the gas transportation business. Journal of Applied Statistics, 28(3–4):343–351, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cox:2001:TIU

[1141] M. A. A. Cox. Towards the implementation of a universal control chart and estimation of its average run length using a spreadsheet: an artificial neural network is employed to model the parameters in a special case. Journal of Applied Statistics, 28(3–4): 353–364, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cox:2001:MSU

[1142] Trevor F. Cox. Multidimensional scaling used in multivariate statistical process control. Journal of Applied Statistics, 28(3–4):365–378, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Godolphin:2001:OTP

[1143] E. J. Godolphin. Observable trendprojecting state-space models. Journal of Applied Statistics, 28(3–4):379–389, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goh:2001:PAE

[1144] T. N. Goh. A pragmatic approach to experimental design in industry. Journal of Applied Statistics, 28(3–4): 391–398, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Henderson:2001:EIA

[1145] G. Robin Henderson. EWMA and industrial applications to feedback adjustment and control. Journal of Applied Statistics, 28(3–4):399–407, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weighell:2001:SMC

[1146] M. Weighell, E. B. Martin, and A. J. Morris. The statistical monitoring of a complex manufacturing process.

Journal of Applied Statistics, 28(3–4): 409–425, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Montgomery:2001:OCI

[1147] Douglas C. Montgomery. Opportunities and challenges for industrial statisticians. Journal of Applied Statistics, 28(3–4):427–439, March 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pongcharoen:2001:ADE

[1148] P. Pongcharoen, D. J. Stewardson, C. Hicks, and P. M. Braiden. Applying designed experiments to optimize the performance of genetic algorithms used for scheduling complex products in the capital goods industry. Journal of Applied Statistics, 28 (3–4):441–455, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elamir:2001:CCB

[1149] Elsayed A. H. Elamir and Allan H. Seheult. Control charts based on linear combinations of order statistics. Journal of Applied Statistics, 28(3–4): 457–468, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stewardson:2001:USR

[1150] Dave Stewardson and Shirley Coleman.

Using the Summed Rank Cusum for monitoring environmental data from industrial processes. Journal of Applied Statistics, 28(3–4):469–484, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stewardson:2001:DPS

[1151] Dave Stewardson, David Porter, and Tony Kelly. The dangers posed by saddle points, and other problems, when using central composite designs. Journal of Applied Statistics, 28(3–4): 485–495, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scarrott:2001:BSM

[1152] Carl Scarrott and Granville Tunnicliffe Wilson. Building a statistical model to predict reactor temperatures. Journal of Applied Statistics, 28(3–4): 497–511, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2001:RCAa

[1153] Seung-Ho Kang and Chul Ahn. Regression coefficient analysis for correlated binomial outcomes. Journal of Applied Statistics, 28(3–4):513–514, March 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Myers:2001:R

[1154] Leann Myers and Stephanie T. Broyles. Response. Journal of Applied Statistics, 28(3–4):??, March 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bairamov:2001:NGF

[1155] I. Bairamov, S. Kotz, and M. Bekçi. New generalized Farlie-Gumbel-Morgenstern distributions and concomitants of order statistics. Journal of Applied Statistics, 28(5):521–536, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chan:2001:ODL

[1156] Ling-Yau Chan and Ying-Nan Guan.

A- and D-optimal designs for a log contrast model for experiments with mixtures. Journal of Applied Statistics, 28(5):537–546, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pauli:2001:PLI

[1157] Francesco Pauli and Stuart Coles. Penalized likelihood inference in extreme value analyses. Journal of Applied Statistics, 28(5):547–560, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garren:2001:BGF

[1158] Steven T. Garren, Richard L. Smith, and Walter W. Piegorsch. Bootstrap goodness-of-fit test for the betabinomial model. Journal of Applied Statistics, 28(5):561–571, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghatak:2001:UAA

[1159] Subrata Ghatak and Jalal U. Siddiki.

The use of the ARDL approach in estimating virtual exchange rates in India. Journal of Applied Statistics, 28 (5):573–583, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Krzanowski:2001:DBI

[1160] W. J. Krzanowski. Data-based interval estimation of classification error rates. Journal of Applied Statistics, 28 (5):585–595, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lim:2001:NON

[1161] Yong B. Lim and B. S. So. A note on the optimal number of centre runs in a second phase design of response surface methods. Journal of Applied Statistics, 28(5):597–602, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Montenegro:2001:SSP

[1162] Alvaro Montenegro. On sample size and precision in ordinary least squares. Journal of Applied Statistics, 28(5): 603–605, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nuel:2001:PDU

[1163] G. Nuel, S. Robin, and C. P. Baril. Predicting distances using a linear model: the case of varietal distinctness. Journal of Applied Statistics, 28 (5):607–621, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2001:MZI

[1164] Peiming Wang. Markov zero-inflated Poisson regression models for a time series of counts with excess zeros. Journal of Applied Statistics, 28(5):623–632, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2001:RCAb

[1165] Seung-Ho Kang and Chul Ahn. Regression coefficient analysis for correlated binomial outcomes. Journal of Applied Statistics, 28(5):633–634, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Myers:2001:AR

[1166] Leann Myers and Stephanie T. Broyles.

Authors’ reply. Journal of Applied Statistics, 28(5):??, July 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Winkelmann:2001:URP

[1167] Rainer Winkelmann. ‘Under-reporting of purchases of port wine’: a correction. Journal of Applied Statistics, 28(5):??, July 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [983].

Bagui:2001:EBD

[1168] S. C. Bagui and D. K. Ghosh. Efficiency balanced designs through reinforcement. Journal of Applied Statistics, 28(6):649–658, August 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cancho:2001:MPI

[1169] Vicente G. Cancho and Heleno Bolfarine. Modeling the presence of immunes by using the exponentiated-Weibull model. Journal of Applied Statistics, 28(6):659–671, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Glickman:2001:DPC

[1170] Mark E. Glickman. Dynamic paired comparison models with stochastic variances. Journal of Applied Statistics, 28(6):673–689, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2001:SPS

[1171] Andy H. Lee, John S. Yick, and Yer Van Hui. Sensitivity of the portmanteau statistic in time series modeling. Journal of Applied Statistics, 28(6):691–702, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Louzada-Neto:2001:ERA

[1172] Francisco Louzada-Neto and Juan Carlos Pardo-Fernández. The effect of reparametrization on accelerated lifetime tests. Journal of Applied Statistics, 28(6):703–711, August 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mills:2001:BCA

[1173] Terence C. Mills. Business cycle asymmetry and duration dependence: an international perspective. Journal of Applied Statistics, 28(6):713–724, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oberrath:2001:SMT

[1174] Reik Oberrath and Katrin Böhning-Gaese. The Signed Mantel test to cope with autocorrelation in comparative analyses. Journal of Applied Statistics, 28(6):725–736, August 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Poon:2001:IMM

[1175] Wai-Yin Poon and Man-Lai Tang. Influence measure in maximum likelihood estimate for models of lifetime data. Journal of Applied Statistics, 28 (6):737–742, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mansson:2001:MVR

[1176] Ralph A. Mansson and Philip Prescott.

Missing values in replicated Latin squares. Journal of Applied Statistics, 28(6):743–757, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wenzel:2001:HME

[1177] Thomas Wenzel. Hits-and-misses for the evaluation and combination of forecasts. Journal of Applied Statistics, 28(6):759–773, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zellner:2001:RCB

[1178] Arnold Zellner. Remarks on a ‘critique’ of the Bayesian Method of Moments. Journal of Applied Statistics, 28(6):775–778, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [954].

Anonymous:2001:BRb

[1179] Anonymous. Book reviews. Journal of Applied Statistics, 28(6):779–785, August 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2001:MAF

[1180] Chung-Ho Chen, Te-Shiang Cheng, and Chao-Yu Chou. Minimum average fraction inspected for TCSP-1 plan. Journal of Applied Statistics, 28(7): 793–799, September 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chiang:2001:PPG

[1181] Jeongwen Chiang, Ching-Fan Chung, and Emily T. Cremers. Promotions and the pattern of grocery shopping time. Journal of Applied Statistics, 28 (7):801–819, September 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2001:CBD

[1182] D. K. Ghosh and S. B. Shrivastava. A class of BIB designs with repeated blocks. Journal of Applied Statistics, 28 (7):821–833, September 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ho:2001:FSP

[1183] Tsung-Wu Ho. Finite-sample properties of the bootstrap estimator in a Markov-switching model. Journal of Applied Statistics, 28(7):835– 842, September 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huettmann:2001:UPS

[1184] F. Huettmann and A. W. Diamond. Using PCA scores to classify species communities: an example for pelagic seabird distribution. Journal of Applied Statistics, 28(7):843–853, September 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garcia-Perez:2001:SSC

[1185] Miguel A. García-Pérez and Vicente Núñez-Antón. Small-sample comparisons for powerdivergence goodness-of-fit statistics for symmetric and skewed simple null hypotheses. Journal of Applied Statistics, 28(7):855–874, September 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Costa:2001:EDC

[1186] Antonio F. B. Costa and M. A. Rahim.

Economic design of X charts with variable parameters: the Markov chain approach. Journal of Applied Statistics, 28(7):875–885, September 1, 2001. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roberts:2001:MTH

[1187] John M. Roberts and Devon D. Brewer.

Measures and tests of heaping in discrete quantitative distributions. Journal of Applied Statistics, 28(7):887– 896, September 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rothery:2001:AGA

[1188] Peter Rothery and David B. Roy. Application of generalized additive models to butterfly transect count data. Journal of Applied Statistics, 28(7): 897–909, September 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Walker:2001:GRR

[1189] Stephen G. Walker and Christopher J.

Page. Generalized ridge regression and a generalization of the C P statistic. Journal of Applied Statistics, 28(7): 911–922, September 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barnett:2001:EPP

[1190] Vic Barnett and Maria Cecilia Mendes Barreto. Estimators for a Poisson parameter using ranked set sampling. Journal of Applied Statistics, 28(8): 929–941, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bisgaard:2001:SOC

[1191] Søren Bisgaard and Murat Kulahci. Switching-one-column follow-up experiments for Plackett-Burman designs. Journal of Applied Statistics, 28(8): 943–949, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bourke:2001:GCC

[1192] Patrick D. Bourke. The geometric CUSUM chart with sampling inspection for monitoring fraction defective. Journal of Applied Statistics, 28(8): 951–972, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Box:2001:EDW

[1193] George Box and Ian Hau. Experimental designs when there are one or more factor constraints. Journal of Applied Statistics, 28(8):973–989, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prayag:2001:MTC

[1194] V. R. Prayag and S. A. Chiplonkar. A multiple test for comparing two treatments with control: interval hypotheses approach. Journal of Applied Statistics, 28(8):991–1001, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Muttlak:2001:REE

[1195] Hassen A. Muttlak. Regression estimators in extreme and median ranked set samples. Journal of Applied Statistics, 28(8):1003–1017, November 1, 2001.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shin:2001:SEP

[1196] Key-Il Shin and Hee-Jeong Kang. A study on the effect of power transformation in the ARMA(p, q) model. Journal of Applied Statistics, 28(8): 1019–1028, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tu:2001:DRR

[1197] X. M. Tu, J. Kowalski, A. Begley, P. Houck, S. Mazumdar, J. Miewald, D. J. Buysse, and D. J. Kupfer. Data recycling: a response to the changing technology from the statistical perspective with application to psychiatric sleep research. Journal of Applied Statistics, 28(8):1029–1049, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Adham:2001:MGT

[1198] Samia A. Adham and Stephen G.

Walker. A multivariate Gompertztype distribution. Journal of Applied Statistics, 28(8):1051–1065, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wong:2001:DFL

[1199] Man-Yu Wong and Shuanglin Zhang.

Degrees of freedom and the likelihood ratio test for the generalized Behrens– Fisher problem. Journal of Applied Statistics, 28(8):1067–1074, November 1, 2001. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Seber:2002:CRB

[1200] George A. F. Seber and Carl J. Schwarz. Capture-recapture: before and after EURING 2000. Journal of Applied Statistics, 29(1–4):5–18, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cooch:2002:OSL

[1201] Evan G. Cooch, Emmanuelle Cam, and William Link. Occam’s shadow: levels of analysis in evolutionary ecology– where to next? Journal of Applied Statistics, 29(1–4):19–48, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nichols:2002:DCO

[1202] James D. Nichols. Discussion comments on: “Occam’s shadow: levels of analysis in evolutionary ecology — where to next?” by Cooch, Cam and Link. Journal of Applied Statistics, 29 (1–4):49–52, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Clobert:2002:CRE

[1203] Jean Clobert. Capture-recapture and evolutionary ecology: further comments. Journal of Applied Statistics, 29(1–4):53–56, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yoccoz:2002:CRC

[1204] Nigel G. Yoccoz, Kjell E. Erikstad, Jan O. Bustnes, et al. Costs of reproduction in common eiders (somateria mollissima): an assessment of relationships between reproductive effort and future survival and reproduction based on observational and experimental studies. Journal of Applied Statistics, 29(1–4):57–64, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nisbet:2002:TAS

[1205] Ian C. T. Nisbet and Emmanuelle Cam.

Test for age-specificity in survival of the common tern. Journal of Applied Statistics, 29(1–4):65–83, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pollock:2002:UAV

[1206] Kenneth H. Pollock. The use of auxiliary variables in capture-recapture modelling: an overview. Journal of Applied Statistics, 29(1–4):85–102, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

White:2002:DCU

[1207] Gary C. White. Discussion comments on: the use of auxiliary variables in capture-recapture modelling. An overview. Journal of Applied Statistics, 29(1–4):103–106, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Piper:2002:SAT

[1208] S. E. Piper. Survival of adult, territorial Longtailed Wagtails Motacilla clara: the effects of environmental factors and individual covariates. Journal of Applied Statistics, 29(1–4):107– 124, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Conroy:2002:AIT

[1209] Michael J. Conroy, Juan Carlos Senar, and Jordi Domènech. Analysis of individual- and time-specific covariate effects on survival of Serinus serinus in north-eastern Spain. Journal of Applied Statistics, 29(1–4):125–142, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cooch:2002:FSS

[1210] Evan G. Cooch. Fledging size and survival in snow geese: timing is everything (or is it?). Journal of Applied Statistics, 29(1–4):143–162, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cam:2002:IBT

[1211] Emmanuelle Cam, Bernard Cadiou, James E. Hines, et al. Influence of behavioural tactics on recruitment and reproductive trajectory in the kittiwake. Journal of Applied Statistics, 29(1–4): 163–185, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brooks:2002:BMA

[1212] S. P. Brooks, E. A. Catchpole, B. J. T. Morgan, et al. Bayesian methods for analysing ringing data. Journal of Applied Statistics, 29(1–4):187–206, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Link:2002:MBE

[1213] William A. Link, Evan G. Cooch, and Emmanuelle Cam. Model-based estimation of individual fitness. Journal of Applied Statistics, 29(1–4):207– 224, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dupuis:2002:PDS

[1214] J. A. Dupuis. Prior distributions for stratified capture-recapture models. Journal of Applied Statistics, 29(1–4): 225–237, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schwarz:2002:DCP

[1215] C. J. Schwarz. Discussion comments on ’prior distributions for stratified capture-recapture models’. Journal of Applied Statistics, 29(1–4):239–240, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dupuis:2002:RCS

[1216] J. A. Dupuis. Response to Carl Schwarz. Journal of Applied Statistics, 29(1–4):241–244, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Burnham:2002:ESR

[1217] Kenneth P. Burnham and Gary C.

White. Evaluation of some random effects methodology applicable to bird ringing data. Journal of Applied Statistics, 29(1–4):245–264, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Johnson:2002:DCE

[1218] Douglas H. Johnson. Discussion comments on ‘Evaluation of some random effects methodology applicable to bird ringing data’ by Burnham & White. Journal of Applied Statistics, 29(1–4): 265–266, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Franklin:2002:ELT

[1219] Alan B. Franklin, David R. Anderson, and Kenneth P. Burnham. Estimation of long-term trends and variation in avian survival probabilities using random effects models. Journal of Applied Statistics, 29(1–4):267–287, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Loison:2002:DSV

[1220] Anne Loison, Bernt-Erik Sæther, Kurt Jerstad, et al. Disentangling the sources of variation in the survival of the European dipper. Journal of Applied Statistics, 29(1–4):289–304, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barker:2002:MDD

[1221] Richard Barker, David Fletcher, and Paul Scofield. Measuring density dependence in survival from markrecapture data. Journal of Applied Statistics, 29(1–4):305–313, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pledger:2002:MHS

[1222] Shirley Pledger and Carl J. Schwarz. Modelling heterogeneity of survival in band-recovery data using mixtures. Journal of Applied Statistics, 29(1–4): 315–327, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Royle:2002:RES

[1223] J. Andrew Royle and William A. Link. Random effects and shrinkage estimation in capture-recapture models. Journal of Applied Statistics, 29(1–4): 329–351, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lebreton:2002:MRM

[1224] J. D. Lebreton and R. Pradel Cefe.

Multistate recapture models: modelling incomplete individual histories. Journal of Applied Statistics, 29(1–4): 353–369, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dhondt:2002:DCM

[1225] Andréa. Dhondt. Discussion comments ’multistate recapture models: modelling incomplete individual histories’– why are we doing all this? Journal of Applied Statistics, 29(1–4):371– 372, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Joe:2002:SSM

[1226] Mijeom Joe and Kenneth H. Pollock Biomathematics. Separation of survival and movement rates in multi-state tagreturn and capture-recapture models. Journal of Applied Statistics, 29(1–4): 373–384, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Spendelow:2002:MPS

[1227] Jeffrey A. Spendelow, James D.

Nichols, James E. Hines, et al. Modelling postfledging survival and age-specific breeding probabilities in species with delayed maturity: a case study of Roseate Terns at Falkner Island, Connecticut. Journal of Applied Statistics, 29(1–4):385–405, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Doligez:2002:CRA

[1228] Blandine Doligez, Jean Clobert, Richard A. Pettifor, et al. Costs of reproduction: assessing responses to brood size manipulation on lifehistory and behavioural traits using multi-state capture-recapture models. Journal of Applied Statistics, 29(1–4): 407–423, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Senar:2002:AEB

[1229] Juan Carlos Senar, Michael J. Conroy, and Antoni Borras. Asymmetric exchange between populations differing in habitat quality: a metapopulation study on the citril finch. Journal of Applied Statistics, 29(1–4):425–441, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nasution:2002:OAS

[1230] Marlina D. Nasution, Cavell Brownie, and Kenneth H. Pollock. Optimal allocation of sample sizes between regular banding and radio-tagging for estimating annual survival and emigration rates. Journal of Applied Statistics, 29 (1–4):443–457, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schwarz:2002:RQE

[1231] Carl James Schwarz. Real and quasiexperiments in capture-recapture studies. Journal of Applied Statistics, 29 (1–4):459–473, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Conroy:2002:RQE

[1232] Michael J. Conroy. Real and quasiexperiments in capture-recapture studies: suggestions for advancing the state of the art. Journal of Applied Statistics, 29(1–4):475–477, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Otis:2002:RAB

[1233] David L. Otis and Gary C. White.

Re-analysis of a banding study to test the effects of an experimental increase in bag limits of mourning doves. Journal of Applied Statistics, 29(1–4): 479–495, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Slattery:2002:UBM

[1234] S. M. Slattery and R. T. Alisauskas.

Use of the Barker model in an experiment examining covariate effects on first-year survival in Ross’s Geese (chen rossii): a case study. Journal of Applied Statistics, 29(1–4):497–508, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boulinier:2002:TEC

[1235] Thierry Boulinier, Nigel G. Yoccoz, Karen D. McCoy, et al. Testing the effect of conspecific reproductive success on dispersal and recruitment decisions in a colonial bird: design issues. Journal of Applied Statistics, 29(1–4): 509–520, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alisauskas:2002:ENS

[1236] R. T. Alisauskas and M. S. Lindberg. Effects of neckbands on survival and fidelity of white-fronted and Canada geese captured as non-breeding adults. Journal of Applied Statistics, 29(1–4): 521–537, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nichols:2002:ADE

[1237] James D. Nichols and James E. Hines. Approaches for the direct estimation of u and demographic contributions to u using capture-recapture data. Journal of Applied Statistics, 29(1–4): 539–568, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barker:2002:DCA

[1238] Richard Barker, Evan Cooch, and Carl Schwarz. Discussion comments on: ‘Approaches for the direct estimation of u and demographic contributions to u using capture-recapture data’. Journal of Applied Statistics, 29(1–4): 569–572, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hines:2002:IPB

[1239] James E. Hines and James D. Nichols. Investigations of potential bias in the estimation of u using Pradel’s (1996) model for capture-recapture data. Journal of Applied Statistics, 29 (1–4):573–587, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sandercock:2002:ERP

[1240] Brett K. Sandercock and Steven R.

Beissinger. Estimating rates of population change for a neotropical parrot with ratio, mark-recapture and matrix methods. Journal of Applied Statistics, 29(1–4):589–607, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dreitz:2002:URD

[1241] Victoria J. Dreitz, James D. Nichols, James E. Hines, et al. The use of resighting data to estimate the rate of population growth of the snail kite in Florida. Journal of Applied Statistics, 29(1–4):609–623, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Catchpole:2002:SPP

[1242] E. A. Catchpole, B. J. T. Morgan, and A. Viallefont. Solving problems in parameter redundancy using computer algebra. Journal of Applied Statistics, 29 (1–4):625–636, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Francis:2002:EAS

[1243] Charles M. Francis and Pertti Saurola.

Estimating age-specific survival rates of tawny owls–recaptures versus recoveries. Journal of Applied Statistics, 29 (1–4):637–647, January 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arnason:2002:PEC

[1244] A. Neil Arnason and Carl J. Schwarz.

POPAN-6: exploring convergence and estimate properties with SIMULATE.

Journal of Applied Statistics, 29(1–4): 649–668, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ugarte:2002:BR

[1245] M. Dolores Ugarte. Book Review. Journal of Applied Statistics, 29(1– 4):??, January 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harvey:2002:URD

[1246] David I. Harvey and Terence C. Mills. Unit roots and double smooth transitions. Journal of Applied Statistics, 29 (5):675–683, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Diaz:2002:GLM

[1247] M. P. Díaz, A. H. Barchuk, S. Luque, et al. Generalized linear models to study spatial distribution of tree species in Argentinean arid Chaco. Journal of Applied Statistics, 29(5): 685–694, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dharmalingam:2002:CPT

[1248] M. Dharmalingam. Construction of Partial Triallel Crosses based on Trojan Square Design. Journal of Applied Statistics, 29(5):695–702, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahn:2002:MRC

[1249] Chul Ahn, Sin-Ho Jung, and Seung-Ho Kang. Modified regression coefficient analysis for repeated binary measurements. Journal of Applied Statistics, 29(5):703–710, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Belaire-Franch:2002:ICC

[1250] Jorge Belaire-Franch and Dulce Contreras-Bayarri. Improving crosscorrelation tests through re-sampling techniques. Journal of Applied Statistics, 29(5):711–720, July 1, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huh:2002:VMD

[1251] Moon Yul Huh and Kiyeol Kim. Visualization of multidimensional data using modifications of the Grand Tour. Journal of Applied Statistics, 29(5): 721–728, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kominek:2002:MCS

[1252] Zbigniew Kominek. Minimum chisquared estimation of stable distributions parameters: an application to the Warsaw Stock Exchange. Journal of Applied Statistics, 29(5):729–744, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zou:2002:ETC

[1253] Kelly H. Zou and W. J. Hall. On estimating a transformation correlation coefficient. Journal of Applied Statistics, 29(5):745–760, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kepner:2002:OEP

[1254] James Kepner and Dennis Wackerly.

Observations on the effect of the prior distribution on the predictive distribution in Bayesian inferences. Journal of Applied Statistics, 29(5):761–769, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2002:BMS

[1255] Man-Suk Oh and Dong Wan Shin. Bayesian model selection and parameter estimation for possibly asymmetric and non-stationary time series using a reversible jump Markov chain Monte Carlo approach. Journal of Applied Statistics, 29(5):771–789, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2002:BRa

[1256] Anonymous. Book reviews. Journal of Applied Statistics, 29(5):791–794, July 1, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zou:2002:SPT

[1257] Kelly H. Zou and W. J. Hall. Semiparametric and parametric transformation models for comparing diagnostic markers with paired design. Journal of Applied Statistics, 29(6):803–816, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hatemi-J:2002:MBC

[1258] Abdulnasser Hatemi-J and Ghazi Shukur. Multivariate-based causality tests of twin deficits in the US. Journal of Applied Statistics, 29(6): 817–824, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Agostinelli:2002:RSR

[1259] C. Agostinelli. Robust stepwise regression. Journal of Applied Statistics, 29 (6):825–840, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gamero:2002:RDM

[1260] M. D. Jiménez Gamero, J. M. Muñoz Pichardo, J. Muñoz García, and A. Pascual Acosta. Rao distance as a measure of influence in the multivariate linear model. Journal of Applied Statistics, 29(6):841–854, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Houseman:2002:ART

[1261] E. Andres Houseman, Louise Ryan, Jonathan I. Levy, and John D. Spengler. Autocorrelation in real-time continuous monitoring of microenvironments. Journal of Applied Statistics, 29(6):855–872, August 01, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2002:AMR

[1262] Jixian Wang and Peter T. Donnan. Adjusting for missing record linkage in outcome studies. Journal of Applied Statistics, 29(6):873–884, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2002:CCC

[1263] Sung H. Park and Kiho Kim. Construction of central composite designs for balanced orthogonal blocks. Journal of Applied Statistics, 29(6):885– 893, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hafeez:2002:DOU

[1264] K. Hafeez, H. Rowlands, G. Kanji, and S. Iqbal. Design optimization using ANOVA. Journal of Applied Statistics, 29(6):895–906, August 01, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Buning:2002:RPM

[1265] Herbert Büning. Robustness and power of modified Lepage, Kolmogorov– Smirnov and Cramér–von Mises twosample tests. Journal of Applied Statistics, 29(6):907–924, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Causey:2002:PEN

[1266] Beverley D. Causey. Parametric estimation of the number of classes in a population. Journal of Applied Statistics, 29(6):925–934, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patrangenaru:2002:BAP

[1267] Vic Patrangenaru and Kanti V. Mardia. A bootstrap approach to Pluto’s origin. Journal of Applied Statistics, 29 (6):935–943, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2002:BRb

[1268] Anonymous. Book reviews. Journal of Applied Statistics, 29(6):945– 946, August 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghitany:2002:MLE

[1269] M. E. Ghitany and S. Al-Awadhi. Maximum likelihood estimation of Burr XII distribution parameters under random censoring. Journal of Applied Statistics, 29(7):955–965, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balamurali:2002:CGR

[1270] S. Balamurali and M. Kalyanasundaram. Construction of a generalized robust Taguchi capability index. Journal of Applied Statistics, 29(7):967– 971, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andersson:2002:MCP

[1271] E. Andersson. Monitoring cyclical processes. A non-parametric approach. Journal of Applied Statistics, 29(7): 973–990, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Caudill:2002:HSB

[1272] Steven B. Caudill and Norman H. Godwin. Heterogeneous skewness in binary choice models: predicting outcomes in the men’s NCAA basketball tournament. Journal of Applied Statistics, 29 (7):991–1001, September 01, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2002:EDC

[1273] Chung-Ho Chen and Chao-Yu Chou.

Economic design of continuous sampling plan under linear inspection cost. Journal of Applied Statistics, 29(7):

1003–1009, September 01, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Verzilli:2002:MCE

[1274] Claudio J. Verzilli and James R. Carpenter. A Monte Carlo EM algorithm for random-coefficient-based dropout models. Journal of Applied Statistics, 29(7):1011–1021, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mays:2002:SSS

[1275] James E. Mays and Jeffrey B. Birch. Smoothing for small samples with model misspecification: nonparametric and semiparametric concerns. Journal of Applied Statistics, 29(7):1023– 1045, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Climov:2002:SES

[1276] Daniela Climov, Michel Delecroix, and Léopold Simar. Semiparametric estimation in single index Poisson regression: a practical approach. Journal of Applied Statistics, 29(7):1047– 1070, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2002:RIE

[1277] Yangxin Huang. Robustness of interval estimation of the 90% effective dose: bootstrap resampling and some largesample parametric methods. Journal of Applied Statistics, 29(7):1071– 1081, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Antoniadou:2002:EOF

[1278] T. Antoniadou and D. Wallach. Evaluating optimal fertilizer rates using plant measurements. Journal of Applied Statistics, 29(7):1083–1099, September 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khuri:2002:GEA

[1279] André I. Khuri. Graphical evaluation of the adequacy of the method of unweighted means. Journal of Applied Statistics, 29(8):1107–1119, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bourke:2002:CSP

[1280] Patrick D. Bourke. A continuous sampling plan using CUSUMs. Journal of Applied Statistics, 29(8):1121– 1133, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yanez:2002:ACC

[1281] N. David Yanez III, Richard A. Kronmal, Jennifer C. Nelson, and Todd A. Alonzo. Analysing change in clinical trials using quasi-likelihoods. Journal of Applied Statistics, 29(8):1135– 1145, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2002:ESM

[1282] S. P. Singh, A. K. Srivastava, and B. V. S. Sisodia. Evaluation of a synthetic method of estimation for small areas. Journal of Applied Statistics, 29(8):1147–1151, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Natarajan:2002:PCS

[1283] Loki Natarajan and John O’Quigley. Predictive capability of stratified proportional hazards models. Journal of Applied Statistics, 29(8):1153– 1163, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kotz:2002:VBD

[1284] Samuel Kotz and J. Renévan Dorp. A versatile bivariate distribution on a bounded domain: another look at the product moment correlation. Journal of Applied Statistics, 29(8):1165– 1179, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hutson:2002:APL

[1285] Alan D. Hutson, Lauren Fishbein, Patricia O’Brien, and Peter W. Stacpoole. Accounting for plasma levels below detection limits in a one-compartment zero-order absorption pharmacokinetics model. Journal of Applied Statistics, 29(8):1181–1190, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Galea:2002:IDS

[1286] Manuel Galea, Heleno Bolfarine, and Filidor Vilcalabra. Influence diagnostics for the structural errors-in-variables model under the Student-t distribution. Journal of Applied Statistics, 29(8):1191–1204, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Friedl:2002:EGV

[1287] Herwig Friedl and Erwin Stampfer. Estimating general variable acceptance sampling plans by bootstrap methods. Journal of Applied Statistics, 29(8): 1205–1217, November 01, 2002. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chatterjee:2002:SIP

[1288] Sangit Chatterjee, Frederick Wiseman, and Robert Perez. Studying improved performance in golf. Journal of Applied Statistics, 29(8):1219–1227, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bar-Hen:2002:IMD

[1289] Avner Bar-Hen. Influence of missing data on compact designs for spacing experiments. Journal of Applied Statistics, 29(8):1229–1240, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Box:2002:FSF

[1290] George E. P. Box and Alberto Luceño.

Feedforward as a supplement to feedback adjustment in allowing for feedstock changes. Journal of Applied Statistics, 29(8):1241–1254, November 01, 2002. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aoki:2003:BAN

[1291] Reiko Aoki, Jorge A. Achcar, Heleno Bolfarine, and Julio M. Singer. Bayesian analysis of null intercept errors-in-variables regression for pretest/post-test data. Journal of Applied Statistics, 30(1):3–12, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Carapeto:2003:THR

[1292] Maria Carapeto and William A. Holt. Testing for heteroscedasticity in regression models. Journal of Applied Statistics, 30(1):13–20, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Panda:2003:RCD

[1293] D. K. Panda, Rajender Parsad, and V. K. Sharma. Robustness of complete diallel crossing plans against exchange of one cross. Journal of Applied Statistics, 30(1):21–35, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Corrente:2003:CBC

[1294] José Eduardo Corrente, Liciana V. A. S. Chalita, and Jeanete Alves Moreira. Choosing between Cox proportional hazards and logistic models for interval- censored data via bootstrap. Journal of Applied Statistics, 30(1):37–47, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vounatsou:2003:SMG

[1295] Penelope Vounatsou, Tom Smith, and Alan E. Gelfand. Spatial modelling of gene frequencies in the presence of undetectable alleles. Journal of Applied Statistics, 30(1):49–62, ???? 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karlis:2003:EAM

[1296] Dimitris Karlis. An EM algorithm for multivariate Poisson distribution and related models. Journal of Applied Statistics, 30(1):63–77, ???? 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reed:2003:ABR

[1297] James F. Reed III. Adjusting for bias in randomized cluster trials. Journal of Applied Statistics, 30(1):79–85, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tjelmeland:2003:BMS

[1298] Håkon Tjelmeland and Kjetill Vassmo Lund. Bayesian modelling of spatial compositional data. Journal of Applied Statistics, 30(1):87–100, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2003:BRa

[1299] Anonymous. Book reviews. Journal of Applied Statistics, 30(1):101–105, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Illert:2003:LAS

[1300] Christopher R. Illert. Lexigenesis in ancestral south-east Australian Aboriginal language. Journal of Applied Statistics, 30(2):113–143, ???? 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2003:SBT

[1301] Kaushik Ghosh, Rao Jammalamadaka, and Ram C. Tiwari. Semiparametric Bayesian techniques for problems in circular data. Journal of Applied Statistics, 30(2):145–161, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2003:IPD

[1302] Shuo-Jye Wu and Chun-Tao Chang. Inference in the Pareto distribution based on progressive Type II censoring with random removals. Journal of Applied Statistics, 30(2):163–172, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wan:2003:UEM

[1303] Alan T. K. Wan, Anoop Chaturvedi, and Guohuazou Zou. Unbiased estimation of the MSE matrices of improved estimators in linear regression. Journal of Applied Statistics, 30(2):173–189, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2003:BIM

[1304] Man-Suk Oh, Jung Whan Choi, and Dai-Gyoung Kim. Bayesian inference and model selection in latent class logit models with parameter constraints: an application to market segmentation. Journal of Applied Statistics, 30(2):191–204, ???? 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cai:2003:SDM

[1305] Yuzhi Cai and Neville Davies. A simple diagnostic method of outlier detection for stationary Gaussian time series. Journal of Applied Statistics, 30(2):205–223, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Makambi:2003:WIC

[1306] Kepher H. Makambi. Weighted inverse chi-square method for correlated significance tests. Journal of Applied Statistics, 30(2):225–234, ???? 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Renard:2003:VLM

[1307] Didier Renard, Helena Geys, Geert Molenberghs, Tomasz Burzykowski, Marc Buyse, Tony Vangeneugden, and Luc Bijnens. Validation of a longitudinally measured surrogate marker for a time-to-event endpoint. Journal of Applied Statistics, 30(2):235–247, ???? 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zheng:2003:IBL

[1308] Gang Zheng and Mohammad F. Al-Saleh. Improving the best linear unbiased estimator for the scale parameter of symmetric distributions by using the absolute value of ranked set samples. Journal of Applied Statistics, 30(3):253–265, April 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cook:2003:SAU

[1309] Steven Cook. The stylized approach to unit root testing: neglected contributions and the cost of simplicity. Journal of Applied Statistics, 30(3):267–272, April 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dasgupta:2003:EII

[1310] Tirthankar Dasgupta. An economic inspection interval for control of defective items in a hot rolling mill. Journal of Applied Statistics, 30(3):273–282, April 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeUna-Alvarez:2003:EUL

[1311] Jacobo De Uña-Álvarez, M. Soledad Otero-GiráLdez, and Gema álvarez Llorente. Estimation under lengthbias and right-censoring: an application to unemployment duration analysis for married women. Journal of Applied Statistics, 30(3):283–291, April 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gustafson:2003:EDP

[1312] Paul Gustafson and Lawrence J. Walker. An extension of the Dirichlet prior for the analysis of longitudinal multinomial data. Journal of Applied Statistics, 30(3):293–310, April 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2003:MSJ

[1313] Saralees Nadarajah and Samuel Kotz. Moments of some J-shaped distributions. Journal of Applied Statistics, 30(3):311–317, April 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gonzalez:2003:BAS

[1314] Camino González and Gabriel Palomo. Bayesian acceptance sampling plans following economic criteria: an application to paper pulp manufacturing. Journal of Applied Statistics, 30(3):319–333, April 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mingoti:2003:NZO

[1315] Sueli A. Mingoti and Otaviano F. Neves. A note on the Zhang omnibus test for normality based on the Q statistic. Journal of Applied Statistics, 30(3):335–341, April 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiang:2003:ALC

[1316] Liming Xiang, Andy H. Lee, and Siu-Keung Tse. Assessing local cluster influence in generalized linear mixed models. Journal of Applied Statistics, 30(4):349–359, May 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barrie:2003:NSR

[1317] Patrick J. Barrie. A new sports ratings system: the tiddlywinks world ratings. Journal of Applied Statistics, 30(4):361–372, May 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kawasaki:2003:DSU

[1318] Yoshinori Kawasaki and Philip Hans Franses. Detecting seasonal unit roots in a structural time series model. Journal of Applied Statistics, 30(4):373– 387, May 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tarpey:2003:EAS

[1319] Thaddeus Tarpey. Estimating the average slope. Journal of Applied Statistics, 30(4):389–395, May 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Oraini:2003:ESD

[1320] Hajaj A. Al-Oraini and M. A. Rahim.

Economic statistical design of xy control charts for systems with gamma(5, 2) in-control times. Journal of Applied Statistics, 30(4):397–409, May 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dagne:2003:UPT

[1321] Getachew Asfaw Dagne. The use of power transformations in small area estimation. Journal of Applied Statistics, 30(4):411–423, May 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Corrente:2003:OMG

[1322] José Eduardo Corrente and Maria Del Pilar Díaz. Ordinal models and generalized estimating equations to evaluate disease severity. Journal of Applied Statistics, 30(4):425–439, May 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hussain:2003:TAN

[1323] Shakir Hussain and Ghazi Shukur. Testing for autocorrelation in nonstationary dynamic systems of equations. Journal of Applied Statistics, 30(4):441–454, May 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Magnussen:2003:SET

[1324] Steen Magnussen. Stepwise estimators for three-phase sampling of categorical variables. Journal of Applied Statistics, 30(5):461–475, June 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2003:FPD

[1325] Youngjo Lee and John A. Nelder. False parsimony and its detection with GLMs. Journal of Applied Statistics, 30(5):477–483, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mvoi:2003:AEU

[1326] Sifa Mvoi and Yan-Xia Lin. Analysis of experiments using the asymptotic quasi-likelihood approach. Journal of Applied Statistics, 30(5):485–505, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yeh:2003:MEW

[1327] Arthur B. Yeh, Dennis K. J.

Lin, Honghong Zhou, and Chandramouliswaran Venkataramani. A multivariate exponentially weighted moving average control chart for monitoring process variability. Journal of Applied Statistics, 30(5):507–536, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wong:2003:MFW

[1328] H. Wong, Wai-Cheung Ip, Zhongjie Xie, and Xueli Lui. Modelling and forecasting by wavelets, and the application to exchange rates. Journal of Applied Statistics, 30(5):537–553, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weigand:2003:EOI

[1329] Christopher Weigand. Economically optimal inspection policy with geometric adaptation. Journal of Applied Statistics, 30(5):555–569, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Almasri:2003:ICR

[1330] Abdullah Almasri and Ghazi Shukur.

An illustration of the causality relation between government spending and revenue using wavelet analysis on Finnish data. Journal of Applied Statistics, 30(5):571–584, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Leary:2003:OOA

[1331] Stephen Leary, Atul Bhaskar, and Andy Keane. Optimal orthogonalarray-based latin hypercubes. Journal of Applied Statistics, 30(5):585–598, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2003:BRb

[1332] Anonymous. Book reviews. Journal of Applied Statistics, 30(5):599–602, June 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aykroyd:2003:WAS

[1333] R. G. Aykroyd and K. V. Mardia. A wavelet approach to shape analysis for spinal curves. Journal of Applied Statistics, 30(6):605–623, July 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2003:SDO

[1334] Ren-Fen Lee and Deng-Yuan Huang. On some data oriented robust estimation procedures for means. Journal of Applied Statistics, 30(6):625–634, July 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patterson:2003:IFT

[1335] K. D. Patterson and S. M. Heravi. The impact of fat-tailed distributions on some leading unit roots tests. Journal of Applied Statistics, 30(6):635–667, July 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Robin:2003:PCC

[1336] S. Robin, M. Lecomte, H. Höfte, and G. Mouille. A procedure for the clustering of cell wall mutants in the model plant Arabidopsis based on Fourier-transform infrared (FT-IR) spectroscopy. Journal of Applied Statistics, 30(6):669–681, July 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2003:ETL

[1337] Daniel X. Wang and Michael D. Conerly. Evaluation of three lack of fit tests in linear regression models. Journal of Applied Statistics, 30(6):683–696, July 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2003:CDC

[1338] D. K. Ghosh and P. C. Biswas. Complete diallel crosses plans through balanced incomplete block designs. Journal of Applied Statistics, 30(6):697– 708, July 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tse:2003:RSP

[1339] Siu-Keung Tse and Chunyan Yang. Reliability sampling plans for the Weibull distribution under Type II progressive censoring with binomial removals. Journal of Applied Statistics, 30(6): 709–718, July 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vougas:2003:RLU

[1340] Dimitrios V. Vougas. Reconsidering LM unit root testing. Journal of Applied Statistics, 30(7):727–741, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Krzanowski:2003:NPE

[1341] W. J. Krzanowski. Non-parametric estimation of distance between groups. Journal of Applied Statistics, 30(7): 743–750, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kowalski:2003:TAR

[1342] J. Kowalski and X. M. Tu. Trend analysis with response incompatible formats and measurement error. Journal of Applied Statistics, 30(7):751–770, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pan:2003:LIA

[1343] Jian-Xin Pan and Peng Bai. Local influence analysis in the growth curve model with Rao’s simple covariance structure. Journal of Applied Statistics, 30(7):771–781, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2003:EPI

[1344] Chien-Tai Lin and N. Balakrishnan. Exact prediction intervals for exponential distributions based on doubly Type-II censored samples. Journal of Applied Statistics, 30(7):783–801, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tideman:2003:SCP

[1345] Nicolaus Tideman and Reza Kheirandish. Structurally consistent probabilities of selecting answers. Journal of Applied Statistics, 30(7):803–811, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2003:RFC

[1346] Duolao Wang, Panuwat Lertsithichai, Kiran Nanchahal, and Mohammed Yousufuddin. Risk factors of coronary heart disease: a Bayesian model averaging approach. Journal of Applied Statistics, 30(7):813–826, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Imon:2003:RDA

[1347] A. H. M. Rahmatullah Imon. Residuals from deletion in added variable plots. Journal of Applied Statistics, 30(7):827–841, August 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2003:ERE

[1348] Youngjo Lee and John A. Nelder.

Extended-REML estimators. Journal of Applied Statistics, 30(8):845–856, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yau:2003:MRI

[1349] Kelvin K. W. Yau, Karen C. H. Yip, and H. K. Yuen. Modelling repeated insurance claim frequency data using the generalized linear mixed model. Journal of Applied Statistics, 30(8): 857–865, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sharma:2003:MBR

[1350] V. K. Sharma, Seema Jaggi, and Cini Varghese. Minimal balanced repeated measurements designs. Journal of Applied Statistics, 30(8):867–872, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mingoti:2003:NSS

[1351] Sueli A. Mingoti. A note on the sample size required in sequential tests for the generalized binomial distribution. Journal of Applied Statistics, 30 (8):873–879, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2003:BAA

[1352] Hyun Sook Oh, Seoung gon Ko, and Man-Suk Oh. A Bayesian approach to assessing population bioequivalence in a 2×2 crossover design. Journal of Applied Statistics, 30(8):881–891, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bellio:2003:IEJ

[1353] R. Bellio and E. Gori. Impact evaluation of job training programmes: selection bias in multilevel models. Journal of Applied Statistics, 30(8):893– 907, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Austin:2003:UFM

[1354] Peter C. Austin and Michael D. Escobar. The use of finite mixture models to estimate the distribution of the health utilities index in the presence of a ceiling effect. Journal of Applied Statistics, 30(8):909–923, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Magel:2003:NPT

[1355] Rhonda C. Magel and Li Qin. A non-parametric test for umbrella alternatives based on ranked-set sampling. Journal of Applied Statistics, 30 (8):925–937, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arnoldsson:2003:ODB

[1356] Göran Arnoldsson. Optimal designs for beta-binomial logistic regression models. Journal of Applied Statistics, 30 (8):939–951, October 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Farewell:2003:PDE

[1357] Vernon T. Farewell and Agnes M.

Herzberg. Plaid designs for the evaluation of training for medical practitioners. Journal of Applied Statistics, 30 (9):957–965, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stevens:2003:EMI

[1358] Richard J. Stevens. Evaluation of methods for interval estimation of model outputs, with application to survival models. Journal of Applied Statistics, 30(9):967–981, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cai:2003:MPC

[1359] Yuzhi Cai and Neville Davies. Monitoring the parameter changes in general ARIMA time series models. Journal of Applied Statistics, 30(9):983–1001, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Thomson:2003:EAD

[1360] David L. Thomson, Arie van Noordwijk, and Ward Hagemeijer. Estimating avian dispersal distances from data on ringed birds. Journal of Applied Statistics, 30(9):1003–1008, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Laaksonen:2003:AIT

[1361] Seppo Laaksonen. Alternative imputation techniques for complex metric variables. Journal of Applied Statistics, 30 (9):1009–1020, November 2003. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gil-Alana:2003:EDD

[1362] Luis A. Gil-Alana. Estimation of the degree of dependence in the temperatures in the northern hemisphere using semi-parametric techniques. Journal of Applied Statistics, 30(9):1021–1031, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Walhin:2003:BHD

[1363] J. F. Walhin. Bivariate Hofmann distributions. Journal of Applied Statistics, 30(9):1033–1046, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ayhan:2003:MRE

[1364] H. Öztay Ayhan. Models of response error components in supervised interview-reinterview surveys. Journal of Applied Statistics, 30(9):1047–1054, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Muttlak:2003:SQC

[1365] Hassen A. Muttlak and Walid S. Al-Sabah. Statistical quality control based on ranked set sampling. Journal of Applied Statistics, 30(9):1055–1078, November 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2003:LDF

[1366] Saralees Nadarajah, Kosto Mitov, and Samuel Kotz. Local dependence functions for extreme value distributions. Journal of Applied Statistics, 30(10): 1081–1100, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Greenacre:2003:SVD

[1367] Michael Greenacre. Singular value decomposition of matched matrices. Journal of Applied Statistics, 30(10): 1101–1113, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sonesson:2003:ESE

[1368] Christian Sonesson. Evaluations of some Exponentially Weighted Moving Average methods. Journal of Applied Statistics, 30(10):1115–1133, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2003:SND

[1369] Yangxin Huang. Selection of number of dose levels and its robustness for binary response data. Journal of Applied Statistics, 30(10):1135–1146, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gil-Alana:2003:FIA

[1370] L. A. Gil-Alana. A fractional integration analysis of the population in some OECD countries. Journal of Applied Statistics, 30(10):1147–1159, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Salvador:2003:ASI

[1371] Manuel Salvador and Pilar Gargallo. Automatic selective intervention in dynamic linear models. Journal of Applied Statistics, 30(10):1161–1184, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cardot:2003:FAP

[1372] Hervé Cardot, Robert Faivre, and Michel Goulard. Functional approaches for predicting land use with the temporal evolution of coarse resolution remote sensing data. Journal of Applied Statistics, 30(10):1185–1199, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Till:2003:BMC

[1373] Robert J. Till and David J. Hand. Behavioural models of credit card usage. Journal of Applied Statistics, 30(10): 1201–1220, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lai:2003:CSC

[1374] Dejian Lai, Rakesh Sharma, Jerry S. Wolinsky, and Ponnada A. Narayana. A comparative study of correlation coefficients in spatially MRSI-observed neurochemicals from multiple sclerosis patients. Journal of Applied Statistics, 30(10):1221–1229, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2003:MCB

[1375] Rand R. Wilcox. Multiple comparisons based on a modified one-step Mestimator. Journal of Applied Statistics, 30(10):1231–1241, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2003:BRc

[1376] Anonymous. Book reviews. Journal of Applied Statistics, 30(10):1243–1247, December 2003. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ismail:2004:SES

[1377] Sanaa Ismaïl. A simple estimator for the shape parameter of the Pareto distribution with economics and medical applications. Journal of Applied Statistics, 31(1):3–13, January 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DiScala:2004:BHM

[1378] L. Di Scala, L. L. La Rocca, and G. Consonni. A Bayesian hierarchical model for the evaluation of a website. Journal of Applied Statistics, 31(1):15–27, January 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pace:2004:ERL

[1379] L. Pace, A. Salvan, and L. Ventura.

The effects of rounding on likelihood procedures. Journal of Applied Statistics, 31(1):29–48, January 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2004:SSP

[1380] Aiyi Liu, Enrique F. Schisterman, and Eric Teoh. Sample size and power calculation in comparing diagnostic accuracy of biomarkers with pooled assessments. Journal of Applied Statistics, 31(1):49–59, January 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andres:2004:ETP

[1381] A. Martín Andrés and I. Herranz Tejedor. The equivalence of two proportions revisited. Journal of Applied Statistics, 31(1):61–72, January 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Illert:2004:PGO

[1382] C. Illert and A. Allison. Phono-genesis and the origin of accusative syntax in proto-Australian language. Journal of Applied Statistics, 31(1):73–104, January 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nummi:2004:PSM

[1383] T. Nummi and J. Möttönen. Prediction of stem measurements of Scots pine. Journal of Applied Statistics, 31 (1):105–114, January 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gouveia:2004:TCP

[1384] Pedro M. D. C. B. Gouveia and Paulo M. M. Rodrigues. Threshold cointegration and the PPP hypothesis. Journal of Applied Statistics, 31 (1):115–127, January 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2004:IBS

[1385] Rand R. Wilcox. Inferences based on a skipped correlation coefficient. Journal of Applied Statistics, 31(2): 131–143, February 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mansson:2004:RCP

[1386] Ralph Mansson and Philip Prescott.

Robustness of a class of partial dialled cross designs to the unavailability of a complete block of observations. Journal of Applied Statistics, 31 (2):145–160, February 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garcia-Perez:2004:SSC

[1387] Miguel A. García-Pérez and Vicente Núñez-Antón. Small-sample comparisons for goodness-of-fit statistics in one-way multinomials with composite hypotheses. Journal of Applied Statistics, 31(2):161–181, February 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2004:MAT

[1388] Chung-Ho Chen and Chao-Yu Chou.

Minimum average total inspection of variable lot-size sampling plan for continuous production. Journal of Applied Statistics, 31(2):183–189, February 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lesty:2004:GMG

[1389] Claude Lesty, Jocelyne Pleau-Varet, and Michèle Kujas. Geometric method and generalized linear models. Two opposite multiparametric approaches illustrated on a sample of pituitary adenomas. Journal of Applied Statistics, 31(2):191–213, February 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Neuhauser:2004:MTV

[1390] Markus Neuhäuser, Herbert Büning, and Ludwig A. Hothorn. Maximum test versus adaptive tests for the twosample location problem. Journal of Applied Statistics, 31(2):215–227, February 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karrison:2004:RST

[1391] Theodore G. Karrison and Peter C. O’Brien. A rank-sum-type test for paired data with multiple endpoints. Journal of Applied Statistics, 31(2): 229–238, February 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Benton:2004:TZB

[1392] Tom Benton, David Hand, and Martin Crowder. Two z’s are better than one. Journal of Applied Statistics, 31 (2):239–247, February 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miro-Quesada:2004:BAM

[1393] Guillermo Miró-Quesada, Enrique Del Castillo, and John J. Peterson. A Bayesian approach for multiple response surface optimization in the presence of noise variables. Journal of Applied Statistics, 31(3):251–270, April 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tomizawa:2004:CDA

[1394] Sadao Tomizawa, Nobuko Miyamoto, and Ryo Funato. Conditional difference asymmetry model for square contingency tables with nominal categories. Journal of Applied Statistics, 31(3):271–277, April 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lawal:2004:UGD

[1395] H. Bayo Lawal. Using a GLM to decompose the symmetry model in square contingency tables with ordered categories. Journal of Applied Statistics, 31(3):279–303, April 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berger:2004:SVE

[1396] Yves G. Berger. A simple variance estimator for unequal probability sampling without replacement. Journal of Applied Statistics, 31(3):305–315, April 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalgonda:2004:MQC

[1397] A. A. Kalgonda and S. R. Kulkarni.

Multivariate quality control chart for autocorrelated processes. Journal of Applied Statistics, 31(3):317–327, April 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deardon:2004:MAC

[1398] R. Deardon, S. G. Gilmour, N. A. Butler, K. Phelps, and R. Kennedy. A method for ascertaining and controlling representation bias in field trials for airborne plant pathogens. Journal of Applied Statistics, 31(3):329–343, April 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Albers:2004:ENP

[1399] Willem Albers and Wilbert C. M. Kallenberg. Empirical non-parametric control charts: Estimation effects and corrections. Journal of Applied Statistics, 31(3):345–360, April 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tarpey:2004:LCE

[1400] Thaddeus Tarpey and Richard D. Sanders. Linear conditional expectation for discretized distributions. Journal of Applied Statistics, 31(3):361– 372, April 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Holgersson:2004:TMA

[1401] H. E. T. Holgersson. Testing for multivariate autocorrelation. Journal of Applied Statistics, 31(4):379–395, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pan:2004:ASS

[1402] Xia Pan and Jeffrey Jarrett. Applying state space to SPC: Monitoring multivariate time series. Journal of Applied Statistics, 31(4):397–418, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sa:2004:OSC

[1403] Ping Sa and Luminita Razaila. Onesided continuous tolerance limits and their accompanying sample size problem. Journal of Applied Statistics, 31(4):419–434, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chowdhury:2004:CGS

[1404] Sanjoy Roy Chowdhury. Catalogue of group structures for three-level fractional factorial designs. Journal of Applied Statistics, 31(4):435–444, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maravelakis:2004:ECM

[1405] Petros E. Maravelakis, John Panaretos, and Stelios Psarakis. EWMA chart and measurement error. Journal of Applied Statistics, 31(4):445–455, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baker:2004:MKT

[1406] Rose D. Baker. A modified Knox test of space–time clustering. Journal of Applied Statistics, 31(4):457–463, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:2004:SSD

[1407] K. V. Mardia, J. Kirkbride, and F. L.

Bookstein. Statistics of shape, direction and cylindrical variables. Journal of Applied Statistics, 31(4):465–479, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balamurali:2004:MCC

[1408] S. Balamurali and K. Subramani. Modified CSP-C continuous sampling plan for consumer protection. Journal of Applied Statistics, 31(4):481–494, May 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Colosimo:2004:SMC

[1409] B. M. Colosimo, R. Pan, and E. del Castillo. A sequential Markov chain Monte Carlo approach to set-up adjustment of a process over a set of lots. Journal of Applied Statistics, 31(5):499–520, June 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Holden:2004:TNA

[1410] Darryl Holden. Testing the normality assumption in the Tobit model. Journal of Applied Statistics, 31(5):521–532, June 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cheung:2004:DGC

[1411] Kui-Yin Cheung and Elspeth Thomson. The demand for gasoline in China: a cointegration analysis. Journal of Applied Statistics, 31(5):533–544, June 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Russell:2004:FFD

[1412] K. G. Russell, S. M. Lewis, and A. Dean. Fractional factorial designs for the detection of interactions between design and noise factors. Journal of Applied Statistics, 31(5):545–552, June 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chau:2004:NPB

[1413] P. H. Chau and Paul S. F. Yip. Nonparametric back-projection of HIV positive tests using multinomial and Poisson settings. Journal of Applied Statistics, 31(5):553–564, June 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2004:URR

[1414] Horng-Jinh Chang, Chih-Li Wang, and Kuo-Chung Huang. Using randomized response to estimate the proportion and truthful reporting probability in a dichotomous finite population. Journal of Applied Statistics, 31(5):565–573, June 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pewsey:2004:TCR

[1415] Arthur Pewsey. Testing for circular reflective symmetry about a known median axis. Journal of Applied Statistics, 31(5):575–585, June 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shin:2004:MUR

[1416] Key-Il Shin. A multivariate unit root test based on the modified weighted symmetric estimator for VAR(p). Journal of Applied Statistics, 31(5):587– 596, June 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Congdon:2004:MTI

[1417] Peter Congdon. Modelling trends and inequality in small area mortality. Journal of Applied Statistics, 31(6):603–622, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luh:2004:IRT

[1418] Wei-Ming Luh and Jiin-Huarng Guo.

Improved robust test statistic based on trimmed means and Hall’s transformation for two-way ANOVA models under non-normality. Journal of Applied Statistics, 31(6):623–643, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Mutairi:2004:BCR

[1419] D. K. Al-Mutairi. Bayesian computations for random environment models. Journal of Applied Statistics, 31(6):645–659, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zheng:2004:MFO

[1420] Gang Zheng. Maximizing a family of optimal statistics over a nuisance parameter with applications to genetic data analysis. Journal of Applied Statistics, 31(6):661–671, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hutson:2004:UFE

[1421] Alan D. Hutson. Utilizing the flexibility of the epsilon-skew-normal distribution for common regression problems. Journal of Applied Statistics, 31(6):673–683, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tojeiro:2004:BAA

[1422] Cynthia A. V. Tojeiro, Francisco Louzada-Neto, and Heleno Bolfarine. A Bayesian analysis for accelerated lifetime tests under an exponential power law model with threshold stress. Journal of Applied Statistics, 31(6):685– 691, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2004:DOP

[1423] Chung-Ho Chen. Determining the optimum process mean of a one-sided specification limit with the linear quality loss function of product. Journal of Applied Statistics, 31(6):693–703, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mukhopadhyay:2004:EOB

[1424] Arup Ranjan Mukhopadhyay. Establishing optimum bleaching conditions for whitening canary wool. Journal of Applied Statistics, 31(6):705–738, July 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yoo:2004:NAM

[1425] Seung-Hoon Yoo. A note on an approximation of the mobile communications expenditures distribution function using a mixture model. Journal of Applied Statistics, 31(7):747–752, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakir:2004:OGA

[1426] M. A. Bakir and B. Altunkaynak. The optimization with the genetic algorithm approach of the multi-objective, joint economical design of the x and R control charts. Journal of Applied Statistics, 31(7):753–772, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ren:2004:FME

[1427] Haobo Ren, Xiao-Hua Zhou, and Hua Liang. A flexible method for estimating the ROC curve. Journal of Applied Statistics, 31(7):773–784, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shively:2004:TUR

[1428] Philip A. Shively. Testing for a unit root in ARIMA processes. Journal of Applied Statistics, 31(7):785–798, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ferrari:2004:BRM

[1429] Silvia L. P. Ferrari and Francisco Cribari-Neto. Beta regression for modelling rates and proportions. Journal of Applied Statistics, 31(7):799–815, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miceli:2004:RPI

[1430] Rosalba Miceli, Lara Lusa, and Luigi Mariani. Revising a prognostic index developed for classification purposes: an application to gastric cancer data. Journal of Applied Statistics, 31(7):817–830, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reed:2004:RTS

[1431] James F. Reed III and David B. Stark. Robust two-sample statistics for testing equality of means: a simulation study. Journal of Applied Statistics, 31(7):831–854, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rosa:2004:BLD

[1432] G. J. M. Rosa, D. Gianola, and C. R. Padovani. Bayesian longitudinal data analysis with mixed models and thick-tailed distributions using MCMC. Journal of Applied Statistics, 31(7):855–873, August 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:2004:PD

[1433] David J. Hand. Pattern discovery.

Journal of Applied Statistics, 31(8): 883–884, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:2004:PDD

[1434] David J. Hand and Richard J. Bolton.

Pattern discovery and detection: a unified statistical methodology. Journal of Applied Statistics, 31(8):885–924, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Navarro:2004:PM

[1435] Gonzalo Navarro. Pattern matching. Journal of Applied Statistics, 31 (8):925–949, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lawson:2004:DGS

[1436] Andrew Lawson, Allan Clark, and Carmen L. Vidal Rodeiro. Developments in general and syndromic surveillance for small area health data. Journal of Applied Statistics, 31(8):951–966, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Glaz:2004:MWD

[1437] Joseph Glaz and Zhenkui Zhang. Multiple window discrete scan statistics. Journal of Applied Statistics, 31(8): 967–980, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

VonEye:2004:CFA

[1438] Alexander Von Eye and Eduardo Gutiérrez Peña. Configural frequency analysis: the search for extreme cells. Journal of Applied Statistics, 31(8):981–997, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2004:DSG

[1439] Cheolwoo Park, J. S. Marron, and Vitaliana Rondonotti. Dependent SiZer: Goodness-of-fit tests for time series models. Journal of Applied Statistics, 31(8):999–1017, October 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Padmanabhan:2004:IPP

[1440] Balaji Padmanabhan. The interestingness paradox in pattern discovery. Journal of Applied Statistics, 31(8): 1019–1035, October 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gilmour:2004:IFL

[1441] Steven G. Gilmour. Irregular fourlevel response surface designs. Journal of Applied Statistics, 31(9):1043–1048, November 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Galea:2004:IDL

[1442] Manuel Galea, Víctor Leiva-Sánchez, and Gilberto A. Paula. Influence diagnostics in log-Birnbaum–Saunders regression models. Journal of Applied Statistics, 31(9):1049–1064, November 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aggarwal:2004:SNR

[1443] M. L. Aggarwal, Lih-Yuan Deng, and Mithilesh Kumar Jha. Some new residual treatment effects designs for comparing test treatments with a control. Journal of Applied Statistics, 31(9): 1065–1081, November 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aksoy:2004:UMC

[1444] Hafzullah Aksoy. Using Markov chains for non-perennial daily streamflow data generation. Journal of Applied Statistics, 31(9):1083–1094, November 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Menendez:2004:TST

[1445] M. L. Menéndez, J. A. Pardo, and L. Pardo. Tests of symmetry in threedimensional contingency tables based on phi-divergence statistics. Journal of Applied Statistics, 31(9):1095–1114, November 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearn:2004:DIP

[1446] W. L. Pearn, Y. C. Chang, and Chien-Wei Wu. Distributional and inferential properties of the process loss indices. Journal of Applied Statistics, 31 (9):1115–1135, November 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2004:BR

[1447] Anonymous. Book reviews. Journal of Applied Statistics, 31(9):1137–1139, November 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shi:2004:AQL

[1448] Genming Shi and N. Rao Chaganty.

Application of quasi-least squares to analyse replicated autoregressive time series regression models. Journal of Applied Statistics, 31(10):1147–1156, December 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2004:VDD

[1449] Jin-Guan Lin, Bo-Cheng Wei, and Nan-Song Zhang. Varying dispersion diagnostics for inverse Gaussian regression models. Journal of Applied Statistics, 31(10):1157–1170, December 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Costa:2004:MPM

[1450] A. F. B. Costa and M. A. Rahim. Monitoring process mean and variability with one non-central chi-square chart. Journal of Applied Statistics, 31(10): 1171–1183, December 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Abujiya:2004:QCC

[1451] Mu’azu R. Abujiya and Hassen A. Muttlak. Quality control chart for the mean using double ranked set sampling. Journal of Applied Statistics, 31 (10):1185–1201, December 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yoo:2004:NBA

[1452] Seung-Hoon Yoo. A note on a Bayesian approach to a dichotomous choice environmental valuation model. Journal of Applied Statistics, 31(10): 1203–1209, December 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reineke:2004:EHD

[1453] David M. Reineke and John S. Crown.

Estimation of hazard, density and survivor functions for randomly censored data. Journal of Applied Statistics, 31 (10):1211–1225, December 2004. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Granger:2004:FPI

[1454] Clive W. J. Granger and Yongil Jeon.

Forecasting performance of information criteria with many macro series. Journal of Applied Statistics, 31(10): 1227–1240, December 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Derado:2004:SBG

[1455] Gordana Derado, Kanti V. Mardia, Vic Patrangenaru, and Hilary W. Thompson. A shape-based glaucoma index for tomographic images. Journal of Applied Statistics, 31(10):1241–1248, December 2004. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Amos:2005:SAH

[1456] Don Amos, T. Randolph Beard, and Steven B. Caudill. A statistical analysis of the handling characteristics of certain sporting arms: frontier regression, the moment of inertia, and the radius of gyration. Journal of Applied Statistics, 32(1):3–16, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marks:2005:EWP

[1457] Neil B. Marks. Estimation of Weibull parameters from common percentiles.

Journal of Applied Statistics, 32(1):17– 24, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chou:2005:ACC

[1458] Chao-Yu Chou, Chung-Ho Chen, and Hui-Rong Liu. Acceptance control charts for non-normal data. Journal of Applied Statistics, 32(1):25–36, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reed:2005:CTS

[1459] James F. Reed III. Contributions to two-sample statistics. Journal of Applied Statistics, 32(1):37–44, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Halkos:2005:CAU

[1460] George E. Halkos and Ilias S. Kevork. A comparison of alternative unit root tests. Journal of Applied Statistics, 32 (1):45–60, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Riba:2005:CPE

[1461] Alex Riba and Josep Ginebra. Changepoint estimation in a multinomial sequence and homogeneity of literary style. Journal of Applied Statistics, 32 (1):61–74, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nielsen:2005:CRC

[1462] Jens Nielsen, Henrik Jensen, and Per K. Andersen. Creating a reference to a complex emergency situation using time series methods: war in Guinea– Bissau 1998–1999. Journal of Applied Statistics, 32(1):75–86, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tomar:2005:TBB

[1463] Jitendra Singh Tomar, Seema Jaggi, and Cini Varghese. On totally balanced block designs for competition effects. Journal of Applied Statistics, 32(1):87– 97, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2005:BRa

[1464] Anonymous. Book reviews. Journal of Applied Statistics, 32(1):99–104, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Micheas:2005:ASD

[1465] Athanasios C. Micheas and Dipak K.

Dey. Assessing shape differences in populations of shapes using the complex Watson shape distribution. Journal of Applied Statistics, 32(2):105– 116, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hall:2005:ICR

[1466] Matthew A. Hall and Matthew S.

Mayo. The impact of correlated readings on the estimation of the average area under readers’ ROC curves. Journal of Applied Statistics, 32(2):117– 125, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Amaral:2005:DPH

[1467] J. A. Amaral, E. P. Pereira, and M. T.

Paixão. Data and projections of HIV/ AIDS cases in Portugal: an unstoppable epidemic? Journal of Applied Statistics, 32(2):127–140, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2005:MIM

[1468] Paul Zhang. Multiple imputation of missing data with ante-dependence covariance structure. Journal of Applied Statistics, 32(2):141–155, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guerrero:2005:REA

[1469] Victor M. Guerrero. Restricted estimation of an adjusted time series: application to Mexico’s industrial production index. Journal of Applied Statistics, 32(2):157–172, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Graffelman:2005:EBC

[1470] Jan Graffelman. Enriched biplots for canonical correlation analysis. Journal of Applied Statistics, 32(2):173– 188, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2005:BRb

[1471] Anonymous. Book reviews. Journal of Applied Statistics, 32(2):189–190, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hussain:2005:EFH

[1472] S. Hussain, R. Harrison, J. Ayres, S. Walter, J. Hawker, R. Wilson, and G. Shukur. Estimation and forecasting hospital admissions due to influenza: Planning for winter pressure. the case of the West Midlands, UK. Journal of Applied Statistics, 32(3):191–205, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2005:JIM

[1473] Sung H. Park, Bum S. Lee, and Hyang S. Jung. Joint impact of multiple observations on a subset of variables in multiple linear regression. Journal of Applied Statistics, 32(3):207–219, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2005:ALF

[1474] Zhang Wu, Yu Tian, and Sheng Zhang. Adjusted-loss-function charts with variable sample sizes and sampling intervals. Journal of Applied Statistics, 32(3):221–242, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2005:INR

[1475] Duolao Wang and Michael Murphy.

Identifying nonlinear relationships in regression using the ACE Algorithm. Journal of Applied Statistics, 32(3): 243–258, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brown:2005:GEP

[1476] Ulysses J. Brown III and Dharam S.

Rana. Generalized exchange and propensity for military service: The moderating effect of prior military exposure. Journal of Applied Statistics, 32(3):259–270, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tracy:2005:USR

[1477] Ronald L. Tracy and David P. Doane.

Using the studentized range to assess kurtosis. Journal of Applied Statistics, 32(3):271–280, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alonso:2005:FEN

[1478] Andrés M. Alonso and Juan J. Romo. Forecast of the expected non-epidemic morbidity of acute diseases using resampling methods. Journal of Applied Statistics, 32(3):281–295, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeLeonAdams:2005:EOF

[1479] Guillermo De León Adams, Pere Grima Cintas, and Xavier Tort-Martorell Llabrés. Experimentation order in factorial designs with 8 or 16 runs. Journal of Applied Statistics, 32(3):297– 313, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2005:BRc

[1480] Anonymous. Book reviews. Journal of Applied Statistics, 32(3):315–320, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zalba:2005:EST

[1481] Belén Zalba, Belén Sánchez-valverde, and José M. Marín. An experimental study of thermal energy storage with phase change materials by design of experiments. Journal of Applied Statistics, 32(4):321–332, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aaberge:2005:CEU

[1482] Rolf Aaberge and Li chun Zhang. A class of exact UMP unbiased tests for conditional symmetry in small-sample square contingency tables. Journal of Applied Statistics, 32(4):333–340, 2005.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Husson:2005:SPA

[1483] François Husson and Jérôme Pagès.

Scatter plot and additional variables. Journal of Applied Statistics, 32(4): 341–350, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yakovchuk:2005:MCC

[1484] Natasha Yakovchuk and Thomas R.

Willemain. Monte Carlo comparison of estimation methods for additive twoway tables. Journal of Applied Statistics, 32(4):351–374, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zeifman:2005:MUS

[1485] Michael I. Zeifman and Dov Ingman. Modelling of unexpected shift in SPC. Journal of Applied Statistics, 32(4):375–386, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dobric:2005:NEL

[1486] Jadran Dobrić and Friedrich Schmid.

Nonparametric estimation of the lower tail dependence λ_{L} in bivariate copulas. Journal of Applied Statistics, 32 (4):387–407, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Argac:2005:THC

[1487] Dogan Argaç. Testing hypotheses on coefficients of variation from a series of two-armed experiments. Journal of Applied Statistics, 32(4):409–419, 2005.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Isogai:2005:ANP

[1488] Takafumi Isogai. Applications of a new power normal family. Journal of Applied Statistics, 32(4):421–436, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2005:BRd

[1489] Anonymous. Book reviews. Journal of Applied Statistics, 32(4):437–441, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kuonen:2005:SBC

[1490] Diego Kuonen. Studentized bootstrap confidence intervals based on Mestimates. Journal of Applied Statistics, 32(5):443–460, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ryan:2005:MCS

[1491] Thomas P. Ryan and William H. Woodall. The most-cited statistical papers. Journal of Applied Statistics, 32(5):461–474, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nicholson:2005:TDC

[1492] Mike Nicholson and Jon Barry. Target detection from a classical and a Bayesian viewpoint. Journal of Applied Statistics, 32(5):475–482, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dancer:2005:RSP

[1493] Diane Dancer and Andrew Tremayne. R-squared and prediction in regression with ordered quantitative response. Journal of Applied Statistics, 32(5):483–493, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2005:ECV

[1494] Rand R. Wilcox. Estimating the conditional variance of Y, given X, in a simple regression model. Journal of Applied Statistics, 32(5):495–502, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huerta:2005:MBW

[1495] Gabriel Huerta. Multivariate Bayes Wavelet shrinkage and applications. Journal of Applied Statistics, 32(5): 529–542, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dietrich:2005:ESS

[1496] Jason Dietrich. The effects of sampling strategies on the small sample properties of the logit estimator. Journal of Applied Statistics, 32(6):543–554, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodrigues:2005:SAT

[1497] Paulo M. M. Rodrigues and Philip Hans Franses. A sequential approach to testing seasonal unit roots in high frequency data. Journal of Applied Statistics, 32(6):555–569, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2005:SSS

[1498] John Zhang and Mahmud Ibrahim. A simulation study on SPSS ridge regression and ordinary least squares regression procedures for multicollinearity data. Journal of Applied Statistics, 32(6):571–588, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schrevens:2005:CSB

[1499] E. Schrevens, H. Coppenolle, and K. M. Portier. A comparative study between latent class binomial segmentation and mixed-effects logistic regression to explore between-respondent variability in visual preference for horticultural products. Journal of Applied Statistics, 32(6):589–605, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Greer:2005:CFD

[1500] Mark R. Greer. Combination forecasting for directional accuracy: an application to survey interest rate forecasts. Journal of Applied Statistics, 32(6):607–615, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2005:UST

[1501] Man Lai Tang and Ka Ho Wu. A unified sequential test procedure for simultaneous testing the equality of several binomial proportions to a specified standard. Journal of Applied Statistics, 32(6):617–624, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Illert:2005:OLZ

[1502] Christopher R. Illert. Origins of linguistic zonation in the Australian Alps. Part 1 — Huygens’ Principle. Journal of Applied Statistics, 32(6):625– 659, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahfouz:2005:SSA

[1503] Mariam Mahfouz, Pascale Giraudet, and Michel Chaput. Strategy for a statistical analysis of odour influence on the mammalian olfactory bulb responsiveness. Journal of Applied Statistics, 32(6):661–679, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2005:BRe

[1504] Anonymous. Book reviews. Journal of Applied Statistics, 32(6):681–682, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2005:GND

[1505] Saralees Nadarajah. A generalized normal distribution. Journal of Applied Statistics, 32(7):685–694, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pan:2005:AAM

[1506] Xia Pan. An alternative approach to multivariate EWMA control chart. Journal of Applied Statistics, 32(7): 695–705, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Ballaa:2005:TCB

[1507] Norah R. Al-Ballaa. Test for cointegration based on two-stage least squares. Journal of Applied Statistics, 32(7):707–713, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghitany:2005:SRM

[1508] M. E. Ghitany, S. Kotz, and M. Xie. On some reliability measures and their stochastic orderings for the Topp– Leone distribution. Journal of Applied Statistics, 32(7):715–722, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Labra:2005:LIN

[1509] Filidor V. Labra, Reiko Aoki, and Heleno Bolfarine. Local influence in null intercept measurement error regression under a Student t model. Journal of Applied Statistics, 32(7):723–740, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alkhamisi:2005:BAL

[1510] M. A. Alkhamisi and Ghazi Shukur. Bayesian analysis of a linear mixed model with AR(p) errors via MCMC. Journal of Applied Statistics, 32(7): 741–755, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mohr:2005:CLE

[1511] Donna L. Mohr. Confidence limits for estimates of totals from stratified samples, with application to Medicare Part B overpayment audits. Journal of Applied Statistics, 32(7):757–769, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gomez-deniz:2005:MUI

[1512] Emilio Gómez-déniz and Francisco J. Vázquez-polo. Modelling uncertainty in insurance Bonus–Malus premium principles by using a Bayesian robustness approach. Journal of Applied Statistics, 32(7):771–784, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsou:2005:IVF

[1513] Tsung-Shan Tsou. Inferences of variance function — a parametric robust way. Journal of Applied Statistics, 32(8):785–796, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marin:2005:FSS

[1514] Juan Miguel Marín, Lluis M. Plà, and David Ríos-Insua. Forecasting for some stochastic process models related to sow farm management. Journal of Applied Statistics, 32(8):797–812, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shao:2005:NDO

[1515] Yuehjen E. Shao, John W. Fowler, and George C. Runger. A note on determining an optimal target by considering the dependence of holding costs and the quality characteristics. Journal of Applied Statistics, 32(8):813–822, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fernandez:2005:PCV

[1516] Arturo J. Fernández. Progressively censored variables sampling plans for twoparameter exponential distributions. Journal of Applied Statistics, 32(8): 823–829, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Durbec:2005:FDA

[1517] C. MantéJ. P. Durbec and J. C. Dauvin. A functional data-analytic approach to the classification of species according to their spatial dispersion. Application to a marine macrobenthic community from the Bay of Morlaix (Western English Channel). Journal of Applied Statistics, 32(8):831–840, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jung:2005:BAP

[1518] Yoon Young Jung, Dong Wan Shin, and Man-Suk Oh. Bayesian analysis of panel data using an MTAR model. Journal of Applied Statistics, 32(8):841–854, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lai:2005:CTC

[1519] Dejian Lai and Shyang-Yun Pamela K. Shiao. Comparing two clinical measurements: a linear mixed model approach. Journal of Applied Statistics, 32(8):855–860, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Adams:2005:LNN

[1520] Joseph Brian Adams and Yijin Wert. Logistic and neural network models for predicting a hospital admission. Journal of Applied Statistics, 32(8):861– 869, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2005:MFM

[1521] Duolao Wang and Pengjun Lu. Modelling and forecasting mortality distributions in England and Wales using the Lee–Carter model. Journal of Applied Statistics, 32(9):873–885, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khuri:2005:SVM

[1522] André I. Khuri. Slack-variable models versus Scheffé’s mixture models. Journal of Applied Statistics, 32(9):887– 908, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2005:SRE

[1523] Yoeng-Kuan Chang and Deng-Yuan Huang. On some robust estimation procedures for quantiles based on data. Journal of Applied Statistics, 32(9):909–927, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Imon:2005:IMI

[1524] A. H. M. Rahmatullah Imon. Identifying multiple influential observations in linear regression. Journal of Applied Statistics, 32(9):929–946, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Benasseni:2005:CSP

[1525] Jacques Bénasséni. A concentration study of principal components. Journal of Applied Statistics, 32(9):947– 957, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bedrick:2005:GMM

[1526] Edward J. Bedrick. Graphical modelling and the Mahalanobis distance. Journal of Applied Statistics, 32(9): 959–967, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chipman:2005:IDR

[1527] Hugh A. Chipman and Hong Gu. Interpretable dimension reduction. Journal of Applied Statistics, 32(9):969– 987, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nunez-Antonio:2005:BAD

[1528] Gabriel Nuñez-Antonio and Eduardo Gutiérrez-Peña. A Bayesian analysis of directional data using the projected normal distribution. Journal of Applied Statistics, 32(10):995–1001, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearn:2005:TPC

[1529] W. L. Pearn, M. H. Shu, and B. M. Hsu. Testing process capability based on C_{pm} in the presence of random measurement errors. Journal of Applied Statistics, 32(10):1003–1024, 2005. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghitany:2005:MOE

[1530] M. E. Ghitany, E. K. Al-Hussaini, and R. A. Al-Jarallah. Marshall–Olkin extended Weibull distribution and its application to censored data. Journal of Applied Statistics, 32(10):1025–1034, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Callens:2005:IET

[1531] Marc Callens and Christophe Croux. The impact of education on third births. A multilevel discrete-time hazard analysis. Journal of Applied Statistics, 32(10):1035–1050, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Senoglu:2005:RFD

[1532] Bırdal Şenoğlu. Robust 2^{k} factorial design with Weibull error distributions. Journal of Applied Statistics, 32(10):1051–1066, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shirke:2005:TIE

[1533] D. T. Shirke, R. R. Kumbhar, and D. Kundu. Tolerance intervals for exponentiated scale family of distributions. Journal of Applied Statistics, 32(10):1067–1074, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lopez-Fidalgo:2005:ODC

[1534] J. López-Fidalgo, J. M. Rodríguez-Díaz, G. Sánchez, and M. T. Santos-Martín. Optimal designs for compartmental models with correlated observations. Journal of Applied Statistics, 32(10):1075–1088, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonett:2005:RCI

[1535] Douglas G. Bonett. Robust confidence interval for a residual standard deviation. Journal of Applied Statistics, 32(10):1089–1094, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bourke:2006:RCV

[1536] Patrick D. Bourke. The RL_{2} chart versus the np chart for detecting upward shifts in fraction defective. Journal of Applied Statistics, 33(1):1–15, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Komori:2006:PWC

[1537] Yoshio Komori. Properties of the Weibull cumulative exposure model.

Journal of Applied Statistics, 33(1):17– 34, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hunt:2006:MDD

[1538] Daniel Hunt and Dale Bowman. Modeling developmental data using U-shaped threshold dose-response curves. Journal of Applied Statistics, 33(1):35–47, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tong:2006:ATI

[1539] Lee-Ing Tong and Chien-Hui Yang. Analyzing type II censored data obtained from repetitious experiments. Journal of Applied Statistics, 33(1):49–63, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhu:2006:NIA

[1540] Li-Ping Zhu, Li-Xing Zhu, and Shi-Song Mao. A non-iterative approach to estimating parameters in a linear structural equation model. Journal of Applied Statistics, 33(1):65–78, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Neuhauser:2006:ETT

[1541] Markus Neuhäuser. An exact test for trend among binomial proportions based on a modified Baumgartner– Weiß–Schindler statistic. Journal of Applied Statistics, 33(1):79–88, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeAlba:2006:BIM

[1542] Enrique De Alba, Juan J. Fernández-Durán, and M. Mercedes Gregorio-Domínguez. Bayesian inference for the mean and standard deviation of a normal population when only the sample size, mean and range are observed. Journal of Applied Statistics, 33(1):89– 99, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Muthuraj:2006:DCT

[1543] D. Muthuraj and D. Senthilkumar. Designing and construction of tightenednormal-tightened variables sampling scheme. Journal of Applied Statistics, 33(1):101–111, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jasic:2006:TEN

[1544] Teo Jašić and Douglas Wood. Testing for efficiency and non-linearity in market and natural time series. Journal of Applied Statistics, 33(2):113–138, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balamurali:2006:AOQ

[1545] S. Balamurali and Chi-Hyuck Jun. Average outgoing quality of CSP-C continuous sampling plan under short run production processes. Journal of Applied Statistics, 33(2):139–154, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2006:BAE

[1546] Yangxin Huang and Hulin Wu. A Bayesian approach for estimating antiviral efficacy in HIV dynamic models. Journal of Applied Statistics, 33 (2):155–174, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sengupta:2006:MMS

[1547] Ashis Sengupta and Fidelis I. Ugwuowo. Modelling multi-stage processes through multivariate distributions. Journal of Applied Statistics, 33(2):175–188, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luceno:2006:RIF

[1548] Alberto Luceño and Jaime Puig-Pey. The random intrinsic fast initial response of one-sided CUSUM charts. Journal of Applied Statistics, 33(2): 189–201, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miyamoto:2006:MIM

[1549] Nobuko Miyamoto, Kouji Tahata, Hirokazu Ebie, and Sadao Tomizawa. Marginal inhomogeneity models for square contingency tables with nominal categories. Journal of Applied Statistics, 33(2):203–215, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Spiegelhalter:2006:SIR

[1550] D. J. Spiegelhalter and E. C. Marshall. Strategies for inference robustness in focused modelling. Journal of Applied Statistics, 33(2):217–232, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cook:2006:FSS

[1551] Steven Cook. A finite-sample sensitivity analysis of the Dickey–Fuller test under local-to-unity detrending. Journal of Applied Statistics, 33(2):233– 240, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Woodhouse:2006:GSS

[1552] Richard Woodhouse. Graphical solutions for structural regression assist errors-in-variables modelling. Journal of Applied Statistics, 33(3):241– 255, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andersson:2006:SSA

[1553] E. Andersson, D. Bock, and M. Frisén.

Some statistical aspects of methods for detection of turning points in business cycles. Journal of Applied Statistics, 33(3):257–278, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2006:ODP

[1554] Fong jung Yu and Jiang liang Hou.

Optimization of design parameters for control charts with multiple assignable causes. Journal of Applied Statistics, 33(3):279–290, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kantam:2006:ERT

[1555] R. R. L. Kantam, G. Srinivasa Rao, and B. Sriram. An economic reliability test plan: Log-logistic distribution. Journal of Applied Statistics, 33(3):291–296, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wade:2006:CDE

[1556] A. M. Wade, K. Lawrence, W. Mandy, and D. Skuse. Charting the development of emotion recognition from 6 years of age. Journal of Applied Statistics, 33(3):297–315, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2006:CIP

[1557] Rand R. Wilcox. Confidence intervals for prediction intervals. Journal of Applied Statistics, 33(3):317–326, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balamurali:2006:RGS

[1558] S. Balamurali and Chi hyuck Jun. Repetitive group sampling procedure for variables inspection. Journal of Applied Statistics, 33(3):327–338, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:2006:PIA

[1559] Kanti V. Mardia, Paul McDonnell, and Alf D. Linney. Penalized image averaging and discrimination with facial and fishery applications. Journal of Applied Statistics, 33(3):339–371, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barone:2006:BAN

[1560] Stefano Barone and Alberto Lombardo. Balanced Asymmetrical Nearly Orthogonal Designs for first and second order effect estimation. Journal of Applied Statistics, 33(4):373–386, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kulahci:2006:GAM

[1561] Murat Kulahci and Søren Bisgaard. A generalization of the alias matrix. Journal of Applied Statistics, 33(4):387– 395, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Piepel:2006:NCC

[1562] Greg F. Piepel. A note comparing component-slope, Scheffé; and Cox parameterizations of the linear mixture experiment model. Journal of Applied Statistics, 33(4):397–403, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kenett:2006:PDS

[1563] Ron S. Kenett. On the planning and design of sample surveys. Journal of Applied Statistics, 33(4):405–415, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yun:2006:MEH

[1564] Sungcheol Yun, So Young Sohn, and Youngjo Lee. Modelling and estimating heavy-tailed non-homogeneous correlated queues: Pareto-inverse gamma HGLM with covariates. Journal of Applied Statistics, 33(4):417–425, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samarah:2006:RSC

[1565] Issam K. Samarah, Gamal S. Weheba, and Thomas E. Lacy. Response surface characterization of the mechanical behavior of impact-damaged sandwich composites. Journal of Applied Statistics, 33(4):427–437, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chakraborti:2006:PED

[1566] Subha Chakraborti. Parameter estimation and design considerations in prospective applications of the X^{−} chart. Journal of Applied Statistics, 33(4):439–459, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pomory:2006:NCV

[1567] Christopher M. Pomory. A note on calculating P values from 0.15–0.005 for the Anderson–Darling normality test using the F distribution. Journal of Applied Statistics, 33(4):461–462, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2006:BRa

[1568] Anonymous. Book reviews. Journal of Applied Statistics, 33(4):463–466, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McSweeney:2006:MPP

[1569] L. A. McSweeney. Monitoring paper production using a spectral control chart designed to detect in the presence of multiple cycles. Journal of Applied Statistics, 33(5):467–480, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zizza:2006:MOG

[1570] Roberta Zizza. A measure of output gap for Italy through structural time series models. Journal of Applied Statistics, 33(5):481–496, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dorea:2006:ETN

[1571] Chang C. Y. Dorea and Sueli A. Mingoti. Estimating the total number of distinct species using quadrat sampling and under-dependence structure. Journal of Applied Statistics, 33(5):497– 512, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ganguly:2006:CLM

[1572] S. S. Ganguly. Cumulative logit models for matched pairs case-control design: Studies with covariates. Journal of Applied Statistics, 33(5):513–522, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalirajan:2006:BFM

[1573] Kaliappa Kalirajan and Shashanka Bhide. Bias free measurement of technical efficiency. Journal of Applied Statistics, 33(5):523–533, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mooney:2006:MSV

[1574] Jennifer A. Mooney, Ian T. Jolliffe, and Peter J. Helms. Modelling seasonally varying data: a case study for Sudden Infant Death Syndrome (SIDS). Journal of Applied Statistics, 33(5):535– 547, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peng:2006:MMM

[1575] J. Peng, C. I. C. Lee, and L. Liu. Maxmin multiple comparison procedure for comparing several dose levels with a zero dose control. Journal of Applied Statistics, 33(5):549–555, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khazaee:2006:RMB

[1576] M. Khazaee and K. Shafie. Regression models for Boolean random sets. Journal of Applied Statistics, 33(5):557– 567, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2006:BRb

[1577] Anonymous. Book reviews. Journal of Applied Statistics, 33(5):569–574, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Figueiredo:2006:TWA

[1578] Adelaide Figueiredo. Two-way analysis of variance for data from a concentrated bipolar Watson distribution. Journal of Applied Statistics, 33(6): 575–581, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stamey:2006:BSS

[1579] James D. Stamey, Dean M. Young, and Tom L. Bratcher. Bayesian sample-size determination for one and two Poisson rate parameters with applications to quality control. Journal of Applied Statistics, 33(6):583–594, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsai:2006:ASB

[1580] Tzong-Ru Tsai and Shuo-Jye Wu. Acceptance sampling based on truncated life tests for generalized Rayleigh distribution. Journal of Applied Statistics, 33(6):595–600, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2006:ODC

[1581] Ashish Das, Sudhir Gupta, and Sanpei Kageyama. A-optimal diallel crosses for test versus control comparisons. Journal of Applied Statistics, 33(6):601–608, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dryden:2006:ISU

[1582] Ian L. Dryden, Rahman Farnoosh, and Charles C. Taylor. Image segmentation using Voronoi polygons and MCMC, with application to muscle fibre images. Journal of Applied Statistics, 33(6):609–622, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arvidsson:2006:IDE

[1583] Martin Arvidsson, Ida Gremyr, and Bo Bergman. Interpretation of dispersion effects in a robust design context. Journal of Applied Statistics, 33(6):623–627, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hadjicostas:2006:MPC

[1584] Petros Hadjicostas. Maximizing proportions of correct classifications in binary logistic regression. Journal of Applied Statistics, 33(6):629–640, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2006:BRc

[1585] Anonymous. Book reviews. Journal of Applied Statistics, 33(6):641–644, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yi:2006:CMP

[1586] G. Yi, S. Coleman, and Q. Ren.

CUSUM method in predicting regime shifts and its performance in different stock markets allowing for transaction fees. Journal of Applied Statistics, 33(7):647–661, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deardon:2006:DFE

[1587] Rob Deardon, Steven G. Gilmour, Neil A. Butler, Kath Phelps, and Roy Kennedy. Designing field experiments which are subject to representation bias. Journal of Applied Statistics, 33(7):663–678, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hutson:2006:MET

[1588] Alan D. Hutson. Modifying the exact test for a binomial proportion and comparisons with other approaches. Journal of Applied Statistics, 33(7):679– 690, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kao:2006:TEM

[1589] Shih-Chou Kao, Chuan-Ching Ho, and Ying-Chin Ho. Transforming the exponential by minimizing the sum of the absolute differences. Journal of Applied Statistics, 33(7):691–702, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martin:2006:EBM

[1590] Michael A. Martin and Steven Roberts. An evaluation of bootstrap methods for outlier detection in least squares regression. Journal of Applied Statistics, 33(7):703–720, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Neuhauser:2006:RMO

[1591] Markus Neuhäuser and Ludwig A. Hothorn. A robust modification of the ordered-heterogeneity test. Journal of Applied Statistics, 33(7):721–727, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Riba:2006:DVH

[1592] Alex Riba and Josep Ginebra. Diversity of vocabulary and homogeneity of literary style. Journal of Applied Statistics, 33(7):729–741, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zayeri:2006:LVR

[1593] Farid Zayeri and Anoshirvan Kazemnejad. A latent variable regression model for asymmetric bivariate ordered categorical data. Journal of Applied Statistics, 33(7):743–753, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2006:BRd

[1594] Anonymous. Book reviews. Journal of Applied Statistics, 33(7):755–758, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2006:TSR

[1595] Wen-Den Chen. Testing for spurious regression in a panel data model with the individual number and time length growing. Journal of Applied Statistics, 33(8):759–772, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cooper:2006:FEA

[1596] Tony Cooper and Mary G. Leitnaker.

Further exploratory analysis of splitplot experiments to study certain stratified effects. Journal of Applied Statistics, 33(8):773–786, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ismail:2006:BFI

[1597] Sanaa A. Ismail and Hesham A. Auda. Bayesian and fiducial inference for the inverse Gaussian distribution via Gibbs sampler. Journal of Applied Statistics, 33(8):787–805, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chiang:2006:SQT

[1598] Yuang-Chin Chiang, Lin-An Chen, and Hsien-Chueh Peter Yang. Symmetric quantiles and their applications. Journal of Applied Statistics, 33(8):807– 817, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tellez:2006:ENP

[1599] E. Ayuga Téllez, A. J. Martín Fernández, C. González García, and E. Martínez Falero. Estimation of nonparametric regression for dasometric measures. Journal of Applied Statistics, 33(8):819–836, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2006:OPD

[1600] Jixian Wang. Optimal parametric design with applications to pharmacokinetic and pharmacodynamic trials. Journal of Applied Statistics, 33(8): 837–852, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2006:DTA

[1601] Seong-Keon Lee and Seohoon Jin. Decision tree approaches for zero-inflated count data. Journal of Applied Statistics, 33(8):853–865, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Escarela:2006:BMC

[1602] Drgabriel Escarela and Jacques F.

Carriére. A bivariate model of claim frequencies and severities. Journal of Applied Statistics, 33(8):867–883, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2006:BRe

[1603] Anonymous. Book reviews. Journal of Applied Statistics, 33(8):885–890, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hoff:2006:BMI

[1604] Ayoe Hoff. Bootstrapping Malmquist Indices for Danish Seiners in the North Sea and Skagerrak. Journal of Applied Statistics, 33(9):891–907, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Iwasaki:2006:BNB

[1605] Masakazu Iwasaki and Hiroe Tsubaki. Bivariate negative binomial generalized linear models for environmental count data. Journal of Applied Statistics, 33(9):909–923, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kaboudan:2006:CFW

[1606] Mak Kaboudan. Computational forecasting of wavelet-converted monthly sunspot numbers. Journal of Applied Statistics, 33(9):925–941, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kucuk:2006:WRT

[1607] Murat Küçük and Necati Ağiralioğlu.

Wavelet regression technique for streamflow prediction. Journal of Applied Statistics, 33(9):943–960, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liang:2006:DIL

[1608] Kuo-Yuan Liang, Jack C. Lee, and Kurt S. H. Shao. On the distribution of the inverted linear compound of dependent F -variates and its application to the combination of forecasts. Journal of Applied Statistics, 33(9):961–973, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2006:MDS

[1609] Ho-Seog Kang, Kee-Hoon Kang, and Sung H. Park. Minimax designs for the stability of slope estimation on secondorder response surfaces. Journal of Applied Statistics, 33(9):975–988, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Illert:2006:OLZ

[1610] Christopher R. Illert. Origins of linguistic zonation in the Australian Alps. Part 2 — Snell’s Law. Journal of Applied Statistics, 33(9):989–1030, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2006:APG

[1611] Saralees Nadarajah. Acknowledgement of priority: the generalized normal distribution. Journal of Applied Statistics, 33(9):1031–1032, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2006:BRf

[1612] Anonymous. Book reviews. Journal of Applied Statistics, 33(9):1033–1035, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

George:2006:CED

[1613] Pradeep George and Madara M. Ogot.

A compromise experimental design method for parametric polynomial response surface approximations. Journal of Applied Statistics, 33(10):1037– 1050, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cook:2006:IBC

[1614] Steven Cook and Alan Speight. International business cycle asymmetry and time irreversible nonlinearities. Journal of Applied Statistics, 33(10):1051–1065, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kudus:2006:PMI

[1615] Kamziah Abd Kudus, A. C. Kimber, and J. Lapongan. A parametric model for the interval censored survival times of Acacia Mangium plantation in a spacing trial. Journal of Applied Statistics, 33(10):1067–1074, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McHale:2006:AGS

[1616] Ian G. McHale and Patrick J. Laycock.

Applications of a general stable law regression model. Journal of Applied Statistics, 33(10):1075–1084, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Minemura:2006:IRM

[1617] Eiji Minemura. An interest-rate model analysis based on data augmentation Bayesian forecasting. Journal of Applied Statistics, 33(10):1085–1104, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearn:2006:MPP

[1618] W. L. Pearn, Y. C. Chang, and Chien-Wei Wu. Measuring process performance based on expected loss with asymmetric tolerances. Journal of Applied Statistics, 33(10):1105–1120, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ewing:2006:DFR

[1619] Bradley T. Ewing, Teresa Kerr, and Mark A. Thompson. Do flow rates respond asymmetrically to water level? Evidence from the Edwards Aquifer. Journal of Applied Statistics, 33(10): 1121–1129, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rice:2006:SDP

[1620] Kenneth Rice and David Spiegelhalter. A simple diagnostic plot connecting robust estimation, outlier detection, and false discovery rates. Journal of Applied Statistics, 33(10):1131–1147, 2006. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jones:2006:BR

[1621] M. C. Jones. Book review. Journal of Applied Statistics, 33(10):1149–1151, 2006. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Whittemore:2007:BFD

[1622] Alice S. Whittemore. A Bayesian false discovery rate for multiple testing. Journal of Applied Statistics, 34(1):1– 9, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yen:2007:MMS

[1623] Amy Ming-Fang Yen and Tony Hsiu-Hsi Chen. Mixture multi-state Markov regression model. Journal of Applied Statistics, 34(1):11–21, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patterson:2007:BRT

[1624] K. D. Patterson. Bias reduction through first-order mean correction, bootstrapping and recursive mean adjustment. Journal of Applied Statistics, 34(1):23–45, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Kandari:2007:TMI

[1625] Noriah M. Al-Kandari, Sana S.

BuHamra, and S. E. Ahmed. Testing and merging information for effect size estimation. Journal of Applied Statistics, 34(1):47–60, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Herzberg:2007:OBD

[1626] Agnes M. Herzberg and Richard G.

Jarrett. A-optimal block designs with additional singly replicated treatments. Journal of Applied Statistics, 34(1):61– 70, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2007:DSR

[1627] Housila P. Singh and Mariano Ruiz Espejo. Double sampling ratio-product estimator of a finite population mean in sample surveys. Journal of Applied Statistics, 34(1):71–85, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Thadewald:2007:JBT

[1628] Thorsten Thadewald and Herbert Büning. Jarque–Bera test and its competitors for testing normality — a power comparison. Journal of Applied Statistics, 34(1):87–105, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2007:NPE

[1629] Zhiguo Wang, Jinde Wang, and Xue Liang. Non-parametric estimation for NHPP software reliability models. Journal of Applied Statistics, 34(1): 107–119, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRa

[1630] Anonymous. Book reviews. Journal of Applied Statistics, 34(1):121–127, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bondell:2007:SIB

[1631] Howard D. Bondell, Aiyi Liu, and Enrique F. Schisterman. Statistical inference based on pooled data: a moment-based estimating equation approach. Journal of Applied Statistics, 34(2):129–140, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Senoglu:2007:REH

[1632] Bırdal Şenoğlu. Robust estimation and hypothesis testing of linear contrasts in analysis of covariance with stochastic covariates. Journal of Applied Statistics, 34(2):141–151, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Militino:2007:BSV

[1633] A. F. Militino, M. D. Ugarte, and T. Goicoa. A BLUP synthetic versus an EBLUP estimator: An empirical study of a small area estimation problem. Journal of Applied Statistics, 34(2):153–165, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kolaiti:2007:CTL

[1634] E. Kolaiti and C. Koukouvinos. A comparison of three-level orthogonal arrays in the presence of a possible correlation in observations. Journal of Applied Statistics, 34(2):167–175, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hwang:2007:RZO

[1635] Yi-Ting Hwang and Peir-Feng Wei. A remark on the Zhang Omnibus Test for normality. Journal of Applied Statistics, 34(2):177–184, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2007:LCL

[1636] Saralees Nadarajah and Samuel Kotz.

On the linear combination of Laplace and logistic random variables. Journal of Applied Statistics, 34(2):185– 194, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Amin:2007:TVA

[1637] Zeinab H. Amin. Tests for the validity of the assumption that the underlying distribution of life is Pareto. Journal of Applied Statistics, 34(2):195–201, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Corain:2007:NPM

[1638] L. Corain and L. Salmaso. A nonparametric method for defining a global preference ranking of industrial products. Journal of Applied Statistics, 34 (2):203–216, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Glasbey:2007:ASP

[1639] C. A. Glasbey and G. W. A. M. Van Der Heijden. Alignment and sub-pixel interpolation of images using Fourier methods. Journal of Applied Statistics, 34(2):217–230, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRb

[1640] Anonymous. Book reviews. Journal of Applied Statistics, 34(2):231–238, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2007:AMA

[1641] Arup Kumar Das. Application of multivariate analysis to increase the yield of dry cell batteries. Journal of Applied Statistics, 34(3):239–248, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mukhopadhyay:2007:TEE

[1642] Indranil Mukhopadhyay, Sudipta Chatterjee, and Aditya Chatterjee. Towards enhancement of the economy of a thermal power generating system through prediction of plant efficiency. Journal of Applied Statistics, 34(3):249–259, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boutahar:2007:EML

[1643] Mohamed Boutahar, Vêlayoudom Marimoutou, and Leíla Nouira. Estimation methods of the long memory parameter: Monte Carlo analysis and application. Journal of Applied Statistics, 34(3):261–301, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeMenezes:2007:CFL

[1644] Lilian M. De Menezes and Ana Lasaosa. Comparing fits of latent trait and latent class models applied to sparse binary data: An illustration with human resource management data. Journal of Applied Statistics, 34(3):303–319, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Diekmann:2007:FDU

[1645] Andreas Diekmann. Not the first digit!

using Benford’s Law to detect fraudulent scientific data. Journal of Applied Statistics, 34(3):321–329, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jung:2007:LTS

[1646] Kang-Mo Jung. Least trimmed squares estimator in the errors-in-variables model. Journal of Applied Statistics, 34(3):331–338, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sithole:2007:BLM

[1647] Jabu S. Sithole and Peter W. Jones. Bivariate longitudinal model for detecting prescribing change in two drugs simultaneously with correlated errors. Journal of Applied Statistics, 34(3):339– 352, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2007:RAS

[1648] Rand R. Wilcox. Robust ANCOVA: Some small-sample results when there are multiple groups and multiple covariates. Journal of Applied Statistics, 34(3):353–364, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRc

[1649] Anonymous. Book reviews. Journal of Applied Statistics, 34(3):365–369, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeGooijer:2007:PNS

[1650] Jan G. De Gooijer. Power of the Neyman Smooth Test for evaluating multivariate forecast densities. Journal of Applied Statistics, 34(4):371–381, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Evans:2007:SMT

[1651] Mark Evans and Richard E. Johnston. Stochastic modelling of times to temperature for furnaces supplying titanium blooms to a rolling mill at TIMET. Journal of Applied Statistics, 34(4):383–397, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsangari:2007:AMC

[1652] Haritini Tsangari. An alternative methodology for combining different forecasting models. Journal of Applied Statistics, 34(4):403–421, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Taylor:2007:NED

[1653] Larry W. Taylor. Nonparametric estimation of duration dependence in militarized interstate disputes. Journal of Applied Statistics, 34(4):423–441, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahadik:2007:SVS

[1654] Shashibhushan B. Mahadik and Digambar T. Shirke. On the superiority of a variable sampling interval control chart. Journal of Applied Statistics, 34(4):443–458, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jarrett:2007:MVA

[1655] Jeffrey E. Jarrett and Xia Pan. Monitoring variability and analyzing multivariate autocorrelated processes. Journal of Applied Statistics, 34(4):459– 469, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Subbaiah:2007:RIO

[1656] Perla Subbaiah and George Xia. Robustness of inference for one-sample problem with correlated observations. Journal of Applied Statistics, 34(4): 471–486, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andreev:2007:CEN

[1657] Andriy Andreev, Antti Kanto, and Pekka Malo. Computational examples of a new method for distribution selection in the Pearson system. Journal of Applied Statistics, 34(4):487–506, 2007.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRd

[1658] Anonymous. Book reviews. Journal of Applied Statistics, 34(4):507–509, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2007:SPN

[1659] Prasun Das and Sasadhar Bera. Standardization of process norms in Baker’s yeast fermentation through statistical models in comparison with neural networks. Journal of Applied Statistics, 34(5):511–527, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gembris:2007:EAR

[1660] Daniel Gembris, John G. Taylor, and Dieter Suter. Evolution of athletic records: Statistical effects versus real improvements. Journal of Applied Statistics, 34(5):529–545, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Magnussen:2007:SBM

[1661] S. Magnussen and R. Reeves. Samplebased maximum likelihood estimation of the autologistic model. Journal of Applied Statistics, 34(5):547–561, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chien:2007:RDU

[1662] Li-Chu Chien and Tsung-Shan Tsou. Regression diagnostic under model misspecification. Journal of Applied Statistics, 34(5):563–575, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jaggi:2007:OCB

[1663] Seema Jaggi, Cini Varghese, and V. K.

Gupta. Optimal circular block designs for neighbouring competition effects. Journal of Applied Statistics, 34(5):577–584, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gustafson:2007:DDA

[1664] Paul Gustafson and S. Siddarth. Describing the dynamics of attention to TV commercials: a hierarchical Bayes analysis of the time to zap an ad. Journal of Applied Statistics, 34(5):585– 609, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Felipe:2007:ORD

[1665] A. Felipe, M. L. Menéndez, and L. Pardo. Order-restricted dose-related trend phi-divergence tests for generalized linear models. Journal of Applied Statistics, 34(5):611–623, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Allen:2007:EMS

[1666] Shannon E. Allen and Burt Holland.

Expected mean squares for hierarchical factorial layouts with population imbalance. Journal of Applied Statistics, 34(5):625–637, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRe

[1667] Anonymous. Book reviews. Journal of Applied Statistics, 34(5):639–643, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roger:2007:EBM

[1668] Patrick Roger and Marie-Helene Broihanne. Efficiency of betting markets and rationality of players: Evidence from the French 6/49 Lotto. Journal of Applied Statistics, 34(6):645–662, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arellano-Valle:2007:BIS

[1669] R. B. Arellano-Valle, H. Bolfarine, and V. H. Lachos. Bayesian inference for skew-normal linear mixed models. Journal of Applied Statistics, 34(6):663–682, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pozueta:2007:IDE

[1670] Lourdes Pozueta, Xavier Tort-Martorell, and Lluis Marco. Identifying dispersion effects in robust design experiments — issues and improvements. Journal of Applied Statistics, 34(6):683–699, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chacko:2007:EPB

[1671] Manoj Chacko and P. Yageen Thomas. Estimation of a parameter of bivariate Pareto distribution by ranked set sampling. Journal of Applied Statistics, 34(6):703–714, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2007:EDF

[1672] Chien-Tai Lin and Cheng-Chieh Chou. Empirical-distribution-function tests for the beta-binomial model. Journal of Applied Statistics, 34(6):715–724, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2007:INU

[1673] Hongmei Zhang. Inferences on the number of unseen species and the number of abundant/rare species. Journal of Applied Statistics, 34(6):725– 740, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sun:2007:IOC

[1674] Juan Sun, Lifu Bi, Yaojun Chi, Guowei Huang, Chun Fan, Kazuo Aoki, Akihiro Kono, Tian Hui, and Junichi Misumi. The impact of ovarian cancer on life expectancy in Japan. Journal of Applied Statistics, 34(6):741–747, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cuesta-Albertos:2007:CSS

[1675] Juan Antonio Cuesta-Albertos, Ricardo Fraiman, Antonio Galves, Jesús Garcia, and Marcela Svarc. Classifying speech sonority functional data using a projected Kolmogorov–Smirnov approach. Journal of Applied Statistics, 34(6):749–761, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRf

[1676] Anonymous. Book reviews. Journal of Applied Statistics, 34(6):763–769, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2007:APF

[1677] Trevor Park. Alternative penalty functions for penalized likelihood principal components. Journal of Applied Statistics, 34(7):767–777, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wong:2007:SLN

[1678] Man Yu Wong and D. R. Cox. On the screening of large numbers of significance tests. Journal of Applied Statistics, 34(7):779–783, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

ElGhourabi:2007:RRC

[1679] Mohamed El Ghourabi and Mohamed Limam. Residual responses to change patterns of autocorrelated processes. Journal of Applied Statistics, 34(7): 785–798, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vijayaraghavan:2007:MSD

[1680] R. Vijayaraghavan. Minimum size double sampling plans for large isolated lots. Journal of Applied Statistics, 34(7):799–806, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Avalos:2007:SMH

[1681] Carlos Díaz Avalos. Spatial modeling of habitat preferences of biological species using Markov random fields. Journal of Applied Statistics, 34(7):807–821, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Benth:2007:STM

[1682] Jūraté šaltyté Benth, Fred Espen Benth, and Paulius Jalinskas. A spatial-temporal model for temperature with seasonal variance. Journal of Applied Statistics, 34(7):823–841, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2007:NCC

[1683] Zhang Wu and Qinan Wang. An NP control chart using double inspections. Journal of Applied Statistics, 34(7):843–855, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2007:MSC

[1684] Tee Chin Chang and Fah Fatt Gan.

Modified Shewhart charts for high yield processes. Journal of Applied Statistics, 34(7):857–877, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRg

[1685] Anonymous. Book reviews. Journal of Applied Statistics, 34(7):879–885, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nofuentes:2007:REK

[1686] J. A. Roldán Nofuentes and J. D. Luna Del Castillo. Risk of error and the kappa coefficient of a binary diagnostic test in the presence of partial verification. Journal of Applied Statistics, 34(8):887–898, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marks:2007:KST

[1687] Neil B. Marks. Kolmogorov–Smirnov test statistic and critical values for the Erlang-3 and Erlang-4 distributions. Journal of Applied Statistics, 34(8):899–906, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xie:2007:CDI

[1688] Feng-Chang Xie, Bo-Cheng Wei, and Jin-Guan Lin. Case-deletion influence measures for the data from multivariate t distributions. Journal of Applied Statistics, 34(8):907–921, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Namata:2007:EFI

[1689] Harriet Namata, Ziv Shkedy, Christel Faes, Marc Aerts, Geert Molenberghs, Heide Theeten, Pierre Van Damme, and Philippe Beutels. Estimation of the force of infection from current status data using generalized linear mixed models. Journal of Applied Statistics, 34(8):923–939, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearn:2007:MCI

[1690] W. L. Pearn, F. K. Wang, and C. H. Yen. Multivariate capability indices: Distributional and inferential properties. Journal of Applied Statistics, 34 (8):941–962, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mills:2007:NTD

[1691] Terence C. Mills. A note on trend decomposition: The “classical” approach revisited with an application to surface temperature trends. Journal of Applied Statistics, 34(8):963–972, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Escalante-Vazquez:2007:SSB

[1692] Edgardo J. Escalante-Vázquez. SPC study of a brewing process. Journal of Applied Statistics, 34(8):973–984, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Babak:2007:ELD

[1693] Olena Babak, Birgir Hrafnkelsson, and Olafur P. Palsson. Estimation of the length distribution of marine populations in the Gaussian-multinomial setting using the method of moments. Journal of Applied Statistics, 34(8): 985–991, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Poon:2007:LCI

[1694] Wai-Yin Poon and Yat Sun Poon.

Local conditional influence. Journal of Applied Statistics, 34(8):997–1009, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRh

[1695] Anonymous. Book reviews. Journal of Applied Statistics, 34(8):1011–1017, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ren:2007:BAC

[1696] Dianxu Ren and Roslyn A. Stone.

A Bayesian adjustment for covariate misclassification with correlated binary outcome data. Journal of Applied Statistics, 34(9):1019–1034, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See corrigendum [2091].

Sarnacchiaro:2007:EDA

[1697] Pasquale Sarnacchiaro and Antonello D’ambra. Explorative data analysis and CATANOVA for ordinal variables: An integrated approach. Journal of Applied Statistics, 34(9):1035–1050, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lim:2007:CKL

[1698] Jonggun Lim and Sangun Park. Censored Kullback–Leibler information and goodness-of-fit test with Type II censored data. Journal of Applied Statistics, 34(9):1051–1064, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hyun:2007:HTS

[1699] Seunggeun Hyun and Yanqing Sun. Hypotheses tests of strain-specific vaccine efficacy adjusted for covariate effects. Journal of Applied Statistics, 34(9):1065–1073, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bermudez:2007:HWF

[1700] J. D. Bermúdez, J. V. Segura, and E. Vercher. Holt–Winters forecasting: An alternative formulation applied to UK air passenger data. Journal of Applied Statistics, 34(9):1075–1090, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Otrano:2007:TEP

[1701] Edoardo Otrano and Umberto Triacca. Testing for equal predictability of stationary ARMA processes. Journal of Applied Statistics, 34(9):1091–1108, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2007:BPA

[1702] Chien-Wei Wu and M. H. Shu. A Bayesian procedure for assessing process performance based on expected relative loss with asymmetric tolerances. Journal of Applied Statistics, 34(9):1109–1123, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2007:SCM

[1703] C. Xu, P. A. Dowd, K. V. Mardia, R. J. Fowell, and C. C. Taylor. Simulating correlated marked-point processes. Journal of Applied Statistics, 34(9):1125–1134, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kanji:2007:PSH

[1704] Gopal K. Kanji and Parvesh K.

Chopra. Poverty as a system: Human contestability approach to poverty measurement. Journal of Applied Statistics, 34(9):1135–1158, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRi

[1705] Anonymous. Book reviews. Journal of Applied Statistics, 34(9):1159–1164, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vidal:2007:IOF

[1706] Ignacio Vidal, Pilar Iglesias, and Manuel Galea. Influential observations in the functional measurement error model. Journal of Applied Statistics, 34(10):1165–1183, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Olave:2007:SPB

[1707] Pilar Olave and Manuel Salvador.

Semi-parametric Bayesian analysis of the proportional hazard rate model an application to the effect of training programs on graduate unemployment. Journal of Applied Statistics, 34(10):1185–1205, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luceno:2007:UPC

[1708] Alberto Luceño. A universal QQ-plot for continuous non-homogeneous populations. Journal of Applied Statistics, 34(10):1207–1223, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Clark:2007:MSP

[1709] David E. Clark, Louise M. Ryan, and F. L. Lucas. A multi-state piecewise exponential model of hospital outcomes after injury. Journal of Applied Statistics, 34(10):1225–1239, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Caroni:2007:RDM

[1710] Chrys Caroni and Nedret Billor. Robust detection of multiple outliers in grouped multivariate data. Journal of Applied Statistics, 34(10):1241–1250, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gill:2007:SAK

[1711] Paramjit S. Gill, Tim B. Swartz, and Michael Treschow. A stylometric analysis of King Alfred’s literary works. Journal of Applied Statistics, 34(10):1251–1258, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Takahashi:2007:CIB

[1712] Daniel Yasumasa Takahashi, Luiz Antonio Baccal, and Koichi Sameshima. Connectivity inference between neural structures via partial directed coherence. Journal of Applied Statistics, 34(10):1259–1273, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Langsrud:2007:ADE

[1713] Øyvind Langsrud, Kjetil Jørgensen, Ragni Ofstad, and Tormod Næs. Analyzing designed experiments with multiple responses. Journal of Applied Statistics, 34(10):1275–1296, 2007. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2007:BRj

[1714] Anonymous. Book reviews. Journal of Applied Statistics, 34(10):1297–1301, 2007. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2008:PGK

[1715] Anonymous. Professor Gopal Kanji’s retirement as editor of Journal of Applied Statistics. Journal of Applied Statistics, 35(1):1–8, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2008:PHA

[1716] Rand R. Wilcox. Post-hoc analyses in multiple regression based on prediction error. Journal of Applied Statistics, 35(1):9–17, January 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Monness:2008:CSC

[1717] Erik Mønness, Kim Pearce, and Shirley Coleman. Comparing a survey and a conjoint study: the future vision of water intermediaries. Journal of Applied Statistics, 35(1):19–30, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dai:2008:OTG

[1718] Hongying Dai and Richard Charnigo. Omnibus testing and gene filtration in microarray data analysis. Journal of Applied Statistics, 35(1):31–47, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2008:BVA

[1719] Xiaomo Jiang and Sankaran Mahadevan. Bayesian validation assessment of multivariate computational models. Journal of Applied Statistics, 35 (1):49–65, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Koksal:2008:EPS

[1720] Gülser Köksal, Burcu Kantar, Taylan Ali Ula, and Murat Caner Testik. The effect of Phase I sample size on the run length performance of control charts for autocorrelated data. Journal of Applied Statistics, 35(1):67–87, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chakraborti:2008:PPC

[1721] S. Chakraborti and S. W. Human. Properties and performance of the cchart for attributes data. Journal of Applied Statistics, 35(1):89–100, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alkhamisi:2008:EFT

[1722] Mahdi A. Alkhamisi, Ghadban Khalaf, and Ghazi Shukur. The effect of fat-tailed error terms on the properties of systemwise RESET test. Journal of Applied Statistics, 35(1):101–113, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2008:EPM

[1723] Daniel X. Wang and Michael D. Conerly. Evaluating the power of Minitab’s data subsetting lack of fit test in multiple linear regression. Journal of Applied Statistics, 35(1):115–124, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2008:BRa

[1724] Anonymous. Book reviews. Journal of Applied Statistics, 35(1):125–129, January 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coetzer:2008:OCS

[1725] R. L. J. Coetzer, D. H. Morgan, and H. Maumela. Optimization of a catalyst system through the sequential application of experimental design techniques. Journal of Applied Statistics, 35 (2):131–147, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vijayaraghavan:2008:PSG

[1726] R. Vijayaraghavan, K. Rajagopal, and A. Loganathan. A procedure for selection of a gamma-Poisson single sampling plan by attributes. Journal of Applied Statistics, 35(2):149–160, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pearce:2008:MDP

[1727] Kim F. Pearce and Shirley Y. Coleman. Modern-day perception of historic footwear and its links to preference. Journal of Applied Statistics, 35 (2):161–178, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Galea:2008:IDC

[1728] Manuel Galea, JoséA. Díaz-García, and Filidor Vilca. Influence diagnostics in the capital asset pricing model under elliptical distributions. Journal of Applied Statistics, 35(2):179–192, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berendt:2008:CLB

[1729] Emil B. Berendt. Contracts, livestock, and the Bernoulli process: an application of statistics to B. Traven’s “cattle drive”. Journal of Applied Statistics, 35 (2):193–202, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sinha:2008:BOA

[1730] Pankaj Sinha and Ashok K. Bansal. Bayesian optimization analysis with ML-II ϵ-contaminated prior. Journal of Applied Statistics, 35(2):203–211, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bock:2008:ACF

[1731] David Bock. Aspects on the control of false alarms in statistical surveillance and the impact on the return of financial decision systems. Journal of Applied Statistics, 35(2):213–227, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2008:BRb

[1732] Anonymous. Book reviews. Journal of Applied Statistics, 35(2):229–232, February 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pino-Mejias:2008:CCM

[1733] Rafael Pino-Mejías, Mercedes Carrasco-Mairena, Antonio Pascual-Acosta, María-Dolores Cubiles-De-La-Vega, and Joaquín Muñoz-García. A comparison of classification models to identify the Fragile X Syndrome. Journal of Applied Statistics, 35(3):233–244, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sengupta:2008:UAL

[1734] Raghu Nandan Sengupta. Use of asymmetric loss functions in sequential estimation problems for multiple linear regression. Journal of Applied Statistics, 35(3):245–261, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pauli:2008:STE

[1735] Francesco Pauli and Laura Rizzi. Summer temperature effects on deaths and hospital admissions among the elderly population in two Italian cities. Journal of Applied Statistics, 35(3):263–276, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Angelopoulos:2008:DAE

[1736] P. Angelopoulos and C. Koukouvinos.

Detecting active effects in unreplicated designs. Journal of Applied Statistics, 35(3):277–281, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2008:SVS

[1737] Sung-Soo Kim, Sung H. Park, and W. J. Krzanowski. Simultaneous variable selection and outlier identification in linear regression using the meanshift outlier model. Journal of Applied Statistics, 35(3):283–291, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2008:ILC

[1738] Tonglin Zhang and Ge Lin. Identification of local clusters for count data: a model-based Moran’s I test. Journal of Applied Statistics, 35(3):293–306, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roy:2008:CAP

[1739] Anuradha Roy. Computation aspects of the parameter estimates of linear mixed effects model in multivariate repeated measures set-up. Journal of Applied Statistics, 35(3):307–320, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2008:MST

[1740] Murari Singh and Michael J. Jones. Modelling spatial–temporal covariance structures in monocropping barley trials. Journal of Applied Statistics, 35 (3):321–333, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2008:BQT

[1741] Man-Lai Tang, Maozai Tian, and Ping-Shing Chan. On the bootstrap quantile-treatment-effect test. Journal of Applied Statistics, 35(3):335–350, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2008:BRc

[1742] Anonymous. Book reviews. Journal of Applied Statistics, 35(3):351–356, March 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gonzalez-Davila:2008:SSD

[1743] Enrique González-Dávila, Josep Ginebra, and Roberto Dorta-Guerra. Sample size determination for 2^{k−r} experiments with a binomial response. Journal of Applied Statistics, 35(4):357– 367, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ruiz:2008:BMM

[1744] M. Ruiz, F. J. Girón, C. J. Pérez, J. Martín, and C. Rojano. A Bayesian model for multinomial sampling with misclassified data. Journal of Applied Statistics, 35(4):369–382, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pavia-Miralles:2008:QCN

[1745] Jose Manuel Pavía-Miralles and Beatriz Larraz-Iribas. Quick counts from non-selected polling stations. Journal of Applied Statistics, 35(4):383–405, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Espinheira:2008:BRR

[1746] Patrícia L. Espinheira, Silvia L. P. Ferrari, and Francisco Cribari-Neto. On beta regression residuals. Journal of Applied Statistics, 35(4):407–419, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mukhopadhyay:2008:MAC

[1747] Arup Ranjan Mukhopadhyay. Multivariate attribute control chart using Mahalanobis D^{2} statistic. Journal of Applied Statistics, 35(4):421–429, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Almasri:2008:TCW

[1748] Abdullah Almasri, Håkan Locking, and Ghazi Shukur. Testing for climate warming in Sweden during 1850–1999, using wavelets analysis. Journal of Applied Statistics, 35(4):431–443, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shankar:2008:BAM

[1749] Viswanathan Shankar and Shrikant I. Bangdiwala. Behavior of agreement measures in the presence of zero cells and biased marginal distributions. Journal of Applied Statistics, 35(4):445–464, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2008:RRM

[1750] Hsiuying Wang. Ranking responses in multiple-choice questions. Journal of Applied Statistics, 35(4):465–474, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2008:BRd

[1751] Anonymous. Book reviews. Journal of Applied Statistics, 35(4):475–479, April 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Best:2008:CTN

[1752] D. J. Best, J. C. W. Rayner, and O. Thas. X^{2} and its components as tests of normality for grouped data. Journal of Applied Statistics, 35(5):481–492, May 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nachane:2008:FIR

[1753] Dilip Nachane and Jose G. Clavel.

Forecasting interest rates: a comparative assessment of some secondgeneration nonlinear models. Journal of Applied Statistics, 35(5):493–514, May 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Isogai:2008:SME

[1754] Takafumi Isogai, Hiroaki Uchida, Susumu Miyama, and Sadao Nishiyama. Statistical modeling of enamel rater value data. Journal of Applied Statistics, 35(5):515–535, May 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Duffy:2008:SOB

[1755] Stephen W. Duffy, Anne-Helene Olsen, Rhian Gabe, László Tabár, Jane Warwick, Hilary Fielder, Laufey Tryggvadóttir, and Olorunsola F. Agbaje. Screening opportunity bias in case-control studies of cancer screening. Journal of Applied Statistics, 35(5):537–546, May 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cook:2008:SRU

[1756] Steven Cook. The sensitivity of robust unit root tests. Journal of Applied Statistics, 35(5):547–557, May 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gupta:2008:IEP

[1757] Sat Gupta and Javid Shabbir. On improvement in estimating the population mean in simple random sampling. Journal of Applied Statistics, 35(5):559–566, May 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Branscum:2008:BNP

[1758] Adam J. Branscum, Timothy E. Hanson, and Ian A. Gardner. Bayesian nonparametric models for regional prevalence estimation. Journal of Applied Statistics, 35(5):567–582, May 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2008:SSS

[1759] Bahadur Singh, Susan Halabi, and Michael J. Schell. Sample size selection in clinical trials when population means are subject to a partial order: onesided ordered alternatives. Journal of Applied Statistics, 35(5):583–600, May 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hacker:2008:OLL

[1760] R. Scott Hacker and Abdulnasser Hatemi-J. Optimal lag-length choice in stable and unstable VAR models under situations of homoscedasticity and ARCH. Journal of Applied Statistics, 35(6):601–615, June 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lopez-Fidalgo:2008:MRO

[1761] J. López-Fidalgo, R. Martín-Martín, and M. Stehlík. Marginally restricted D-optimal designs for correlated observations. Journal of Applied Statistics, 35(6):617–632, June 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mazumder:2008:FFP

[1762] Rahul Mazumder. Fluid flow pattern analysis in a trough region: a nonparametric approach. Journal of Applied Statistics, 35(6):633–645, June 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2008:MCM

[1763] H. Jiang, M. Xie, and L. C. Tang.

Markov chain Monte Carlo methods for parameter estimation of the modified Weibull distribution. Journal of Applied Statistics, 35(6):647–658, June 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rousson:2008:MFD

[1764] Valentin Rousson. Monotone fitting for developmental variables. Journal of Applied Statistics, 35(6):659–670, June 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lai:2008:RCC

[1765] C. D. Lai and K. Govindaraju. Reduction of control-chart signal variability for high-quality processes. Journal of Applied Statistics, 35(6):671–679, June 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nummi:2008:AGC

[1766] Tapio Nummi and Laura Koskela. Analysis of growth curve data by using cubic smoothing splines. Journal of Applied Statistics, 35(6):681–691, June 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sengupta:2008:BAC

[1767] Ashis Sengupta and Arnab Kumar Laha. A Bayesian analysis of the change-point problem for directional data. Journal of Applied Statistics, 35(6):693–700, June 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jung:2008:CDT

[1768] Byoung Cheol Jung, AndréI. Khuri, and Juneyoung Lee. Comparison of designs for the three-fold nested random model. Journal of Applied Statistics, 35(6):701–715, June 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alpuim:2008:ERA

[1769] Teresa Alpuim and Abdel El-Shaarawi. On the efficiency of regression analysis with AR(p) errors. Journal of Applied Statistics, 35(7):717–737, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prescott:2008:OMC

[1770] Philip Prescott and Norman R. Draper. D-optimal mixture component-amount designs for quadratic and cubic models. Journal of Applied Statistics, 35(7):739–749, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Meintanis:2008:NIP

[1771] Simos G. Meintanis. New inference procedures for generalized Poisson distributions. Journal of Applied Statistics, 35(7):751–762, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mingoti:2008:ESC

[1772] Sueli A. Mingoti, Júlia P. De Carvalho, and Joab De Oliveira Lima. On the estimation of serial correlation in Markov-dependent production processes. Journal of Applied Statistics, 35(7):763–771, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mark:2008:TPF

[1773] Sigyn Mark and Sture Holm. Test and prediction in factorial models with independent variance estimates. Journal of Applied Statistics, 35(7):773–782, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahmoud:2008:PCS

[1774] Mahmoud A. Mahmoud, William H.

Woodall, and Robert E. Davis. Performance comparison of some likelihood ratio-based statistical surveillance methods. Journal of Applied Statistics, 35(7):783–798, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boone:2008:BHR

[1775] Edward L. Boone, Susan J. Simmons, Haikun Bao, and Ann E. Stapleton. Bayesian hierarchical regression models for detecting QTLs in plant experiments. Journal of Applied Statistics, 35(7):799–808, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Costa:2008:BCC

[1776] A. F. B. Costa and M. A. G. Machado. Bivariate control charts with double sampling. Journal of Applied Statistics, 35(7):809–822, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Robinson:2008:BR

[1777] Andrew Robinson. Book review. Journal of Applied Statistics, 35(7):823– 824, July 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

HerveCardot:2008:VTR

[1778] HervéCardot, Philippe Maisongrande, and Robert Faivre. Varying-time random effects models for longitudinal data: unmixing and temporal interpolation of remote-sensing data. Journal of Applied Statistics, 35(8):827–846, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hamilton:2008:BSS

[1779] Cody Hamilton, Tom L. Bratcher, and James D. Stamey. Bayesian subset selection approach to ranking normal means. Journal of Applied Statistics, 35(8):847–851, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Buning:2008:ABT

[1780] Herbert Büning and Michael Rietz. Adaptive bootstrap tests and its competitors in the c-sample scale problem. Journal of Applied Statistics, 35 (8):853–866, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chan:2008:SIS

[1781] P. S. Chan, H. K. T. Ng, and N. Balakrishnan. Statistical inference for start-up demonstration tests with rejection of units upon observing d failures. Journal of Applied Statistics, 35 (8):867–878, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2008:DII

[1782] Wen-Den Chen. Detecting and identifying interventions with the Whittle spectral approach in a long memory panel data model. Journal of Applied Statistics, 35(8):879–892, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Djira:2008:EEU

[1783] Gemechis D. Djira, Volker Guiard, and Frank Bretz. Efficient and easy-to-use sample size formulas in ratio-based non-inferiority tests. Journal of Applied Statistics, 35(8):893–900, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elevli:2008:CAR

[1784] Sermin Elevli, Nevin Uzgören, and Birol Elevli. Correspondence analysis of repair data: a case study for electric cable shovels. Journal of Applied Statistics, 35(8):901–908, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Halkos:2008:SPT

[1785] George E. Halkos and Ilias S. Kevork.

A sequential procedure for testing the existence of a random walk model in finite samples. Journal of Applied Statistics, 35(8):909–925, August 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2008:FWA

[1786] Donghoh Kim, Youngjo Lee, and Hee-Seok Oh. A fast wavelet approach for recovering damaged images. Journal of Applied Statistics, 35(8):927–938, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Makambi:2008:BR

[1787] Kepher H. Makambi. Book review. Journal of Applied Statistics, 35(8): 939–940, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakouch:2008:BR

[1788] Hassan S. Bakouch. Book review. Journal of Applied Statistics, 35(8):941– 942, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wooff:2008:BR

[1789] David Wooff. Book review. Journal of Applied Statistics, 35(8):943–944, August 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aykroyd:2008:E

[1790] Robert G. Aykroyd. Editorial. Journal of Applied Statistics, 35(9):945–946, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gardner-Lubbe:2008:MFP

[1791] Sugnet Gardner-Lubbe, Niēl J. Le Roux, and John C. Gowers. Measures of fit in principal component and canonical variate analyses. Journal of Applied Statistics, 35(9):947–965, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boone:2008:SCM

[1792] Edward L. Boone and Bronson P. Bullock. Spatial correlation matrix selection using Bayesian model averaging to characterize inter-tree competition in loblolly pine trees. Journal of Applied Statistics, 35(9):967–977, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Almorza:2008:RED

[1793] David Almorza and M. Hortensia García. Results of exploratory data analysis in the broken stick model. Journal of Applied Statistics, 35(9): 979–983, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lautensack:2008:FTD

[1794] Claudia Lautensack. Fitting threedimensional Laguerre tessellations to foam structures. Journal of Applied Statistics, 35(9):985–995, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Armero:2008:BAT

[1795] Carmen Armero, Antonio López-Quílez, and Rut López-Sánchez. Bayesian assessment of times to diagnosis in breast cancer screening.

Journal of Applied Statistics, 35(9): 997–1009, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bousquet:2008:DPD

[1796] Nicolas Bousquet. Diagnostics of priordata agreement in applied Bayesian analysis. Journal of Applied Statistics, 35(9):1011–1029, September 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Georgievska:2008:SRP

[1797] A. Georgievska, L. Georgievska, A. Stojanovic, and N. Todorovic. Sovereign rescheduling probabilities in emerging markets: a comparison with credit rating agencies’ ratings. Journal of Applied Statistics, 35(9):1031–1051, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rueda-Sabater:2008:AAS

[1798] Cristina Rueda-Sabater and Pedro C. Alvarez-Esteban. The analysis of agespecific fertility patterns via logistic models. Journal of Applied Statistics, 35(9):1053–1070, September 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goethals:2008:FMC

[1799] Klara Goethals, Paul Janssen, and Luc Duchateau. Frailty models and copulas: similarities and differences. Journal of Applied Statistics, 35(9): 1071–1079, September 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2008:GFM

[1800] Honghu Liu, Yan Zheng, and Jie Shen.

Goodness-of-fit measures of R^{2} for repeated measures mixed effect models. Journal of Applied Statistics, 35(10): 1081–1092, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Raffalovich:2008:MSP

[1801] Lawrence E. Raffalovich, Glenn D.

Deane, David Armstrong, and Hui-Shien Tsao. Model selection procedures in social research: Monte-Carlo simulation results. Journal of Applied Statistics, 35(10):1093–1114, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vicari:2008:TSG

[1802] Donatella Vicari, Johan Rene Van Dorp, and Samuel Kotz. Two-sided generalized Topp and Leone (TS-GTL) distributions. Journal of Applied Statistics, 35(10):1115–1129, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mills:2008:PBF

[1803] Terence C. Mills. Predicting body fat using weight–height indices. Journal of Applied Statistics, 35(10):1131–1138, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2008:ETP

[1804] Shuo-Jye Wu. Estimation of the twoparameter bathtub-shaped lifetime distribution with progressive censoring. Journal of Applied Statistics, 35(10): 1139–1150, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramasamy:2008:SRG

[1805] Subburaj Ramasamy and Gopal Govindasamy. A software reliability growth model addressing learning. Journal of Applied Statistics, 35(10):1151–1168, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Degiannakis:2008:AAT

[1806] Stavros Degiannakis. ARFIMAX and ARFIMAX-TARCH realized volatility modeling. Journal of Applied Statistics, 35(10):1169–1180, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2008:MTQ

[1807] Lin-An Chen, Hsien-Chueh Peter Yang, and Chau-Shyun Tang. Mode type quasi-range and its applications. Journal of Applied Statistics, 35(10): 1181–1192, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Moghimbeigi:2008:MZI

[1808] Abbas Moghimbeigi, Mohammed Reza Eshraghian, Kazem Mohammad, and Brian Mcardle. Multilevel zero-inflated negative binomial regression modeling for over-dispersed count data with extra zeros. Journal of Applied Statistics, 35(10):1193–1202, October 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Amin:2008:BIP

[1809] Zeinab H. Amin. Bayesian inference for the Pareto lifetime model under progressive censoring with binomial removals. Journal of Applied Statistics, 35(11):1203–1217, November 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2008:BAI

[1810] Eun Sug Park, Roger E. Smith, Thomas J. Freeman, and Clifford H. Spiegelman. A Bayesian approach for improved pavement performance prediction. Journal of Applied Statistics, 35(11):1219–1238, November 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cancho:2008:BAS

[1811] V. G. Cancho, Reiko Aoki, and V. H.

Lachos. Bayesian analysis for a skew extension of the multivariate null intercept measurement error model. Journal of Applied Statistics, 35(11):1239–1251, November 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ugarte:2008:AEE

[1812] M. D. Ugarte, A. F. Militino, and T. Goicoa. Adjusting economic estimates in business surveys. Journal of Applied Statistics, 35(11):1253–1265, November 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2008:BCC

[1813] Zhang Wu, Jianxin Jiao, and Ying Liu.

A binomial CUSUM chart for detecting large shifts in fraction nonconforming. Journal of Applied Statistics, 35 (11):1267–1276, November 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rahman:2008:ISV

[1814] Md. Mostafizur Rahman, Jian-Ping Zhu, and M. Sayedur Rahman. Impact study of volatility modelling of Bangladesh stock index using nonnormal density. Journal of Applied Statistics, 35(11):1277–1292, November 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2008:PPG

[1815] Shuo-Jye Wu, Chun-Tao Chang, Kang-Jun Liao, and Syuan-Rong Huang. Planning of progressive groupcensoring life tests with cost considerations. Journal of Applied Statistics, 35 (11):1293–1304, November 2008. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shin:2008:DUR

[1816] Dong Wan Shin, Yoon Young Jung, and Man-Suk Oh. Double unit root tests for cross-sectionally dependent panel data. Journal of Applied Statistics, 35(11):1305–1321, November 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vallejos:2008:AAB

[1817] Ronny Vallejos. Assessing the association between two spatial or temporal sequences. Journal of Applied Statistics, 35(12):1323–1343, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jahanshahi:2008:OPF

[1818] M. Jahanshahi, M. H. Sanati, and Z. Babaei. Optimization of parameters for the fabrication of gelatin nanoparticles by the Taguchi robust design method. Journal of Applied Statistics, 35(12):1345–1353, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Austin:2008:LSP

[1819] Peter C. Austin. The large-sample performance of backwards variable elimination. Journal of Applied Statistics, 35(12):1355–1370, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bar-Lev:2008:PCI

[1820] Shaul K. Bar-Lev. Point and confidence interval estimates for a global maximum via extreme value theory. Journal of Applied Statistics, 35(12): 1371–1381, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saha:2008:SED

[1821] Krishna K. Saha. Semiparametric estimation for the dispersion parameter in the analysis of over- or underdispersed count data. Journal of Applied Statistics, 35(12):1383–1397, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Angelopoulos:2008:SRP

[1822] P. Angelopoulos and C. Koukouvinos.

Some robust parameter designs from orthogonal arrays. Journal of Applied Statistics, 35(12):1399–1408, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maity:2008:ALR

[1823] Arnab Maity and Michael Sherman.

On adaptive linear regression. Journal of Applied Statistics, 35(12):1409–1422, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pienaar:2008:BRa

[1824] Jacques A. Pienaar. Book review. Journal of Applied Statistics, 35(12): 1423–1424, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pienaar:2008:BRb

[1825] Jacques A. Pienaar. Book review. Journal of Applied Statistics, 35(12): 1425–1426, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barber:2008:BR

[1826] Stuart Barber. Book review. Journal of Applied Statistics, 35(12):1427–1428, December 2008. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mukhopadhyay:2009:PMP

[1827] A. Mukhopadhyay and A. Iqbal. Prediction of mechanical property of steel strips using multivariate adaptive regression splines. Journal of Applied Statistics, 36(1):1–9, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2009:RAS

[1828] Min Kim and Bong-Jin Yum. Reliability acceptance sampling plans for the Weibull distribution under accelerated Type-I censoring. Journal of Applied Statistics, 36(1):11–20, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Curto:2009:CVA

[1829] JoséDias Curto and JoséCastro Pinto.

The coefficient of variation asymptotic distribution in the case of non-iid random variables. Journal of Applied Statistics, 36(1):21–32, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Angelopoulos:2009:MIU

[1830] P. Angelopoulos, H. Evangelaras, and C. Koukouvinos. Model identification using 27 runs three level orthogonal arrays. Journal of Applied Statistics, 36(1):33–38, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weston:2009:DST

[1831] Claire L. Weston and John R. Thompson. The definition of start time in cancer treatment studies analysed by nonmixture cure models. Journal of Applied Statistics, 36(1):39–52, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Economou:2009:FPF

[1832] P. Economou and C. Caroni. Fitting parametric frailty and mixture models under biased sampling. Journal of Applied Statistics, 36(1):53–66, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huh:2009:WVM

[1833] Myung-Hoe Huh and Yong B. Lim.

Weighting variables in K-means clustering. Journal of Applied Statistics, 36(1):67–78, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Su:2009:TPA

[1834] Yih Su and Jing-Shiang Hwang. A two-phase approach to estimating timevarying parameters in the capital asset pricing model. Journal of Applied Statistics, 36(1):79–89, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jouini:2009:ASB

[1835] Jamel Jouini. Analysis of structural break models based on the evolutionary spectrum: Monte Carlo study and application. Journal of Applied Statistics, 36(1):91–110, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brown:2009:RBU

[1836] Ulysses J. Brown III, Stephen B. Knouse, James B. Stewart, and Ruby L. Beale. The relationship between unit diversity and perceptions of organizational performance in the military. Journal of Applied Statistics, 36 (1):111–120, January 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bebbington:2009:MLC

[1837] M. Bebbington, C. D. Lai, and R. Zitikis. Modeling lactation curves: classical parametric models re-examined and modified. Journal of Applied Statistics, 36(2):121–133, February 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sheu:2009:EGC

[1838] Shey-Huei Sheu and Yu-Tai Hsieh. The extended GWMA control chart. Journal of Applied Statistics, 36(2):

135–147, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vilar:2009:NAA

[1839] J. M. Vilar, R. Cao, M. C. Ausín, and C. González-Fragueiro. Nonparametric analysis of aggregate loss models. Journal of Applied Statistics, 36 (2):149–166, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haputhantri:2009:EGC

[1840] S. S. K. Haputhantri, J. Moreau, and S. Lek. Exploring gillnet catch efficiency of sardines in the coastal waters of Sri Lanka by means of three statistical techniques: a comparison of linear and nonlinear modelling techniques. Journal of Applied Statistics, 36 (2):167–179, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Celano:2009:RDA

[1841] Giovanni Celano. Robust design of adaptive control charts for manual manufacturing/inspection workstations. Journal of Applied Statistics, 36 (2):181–203, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Albing:2009:SZB

[1842] Malin Albing and Kerstin Vännman.

Skewed zero-bound distributions and process capability indices for upper specifications. Journal of Applied Statistics, 36(2):205–221, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cabras:2009:DBG

[1843] S. Cabras and M. E. Castellanos. Default Bayesian goodness-of-fit tests for the skew-normal model. Journal of Applied Statistics, 36(2):223–232, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luo:2009:AED

[1844] Weiqi Luo. Analysing ecological data. Journal of Applied Statistics, 36(2): 233–234, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barber:2009:B

[1845] Stuart Barber. Bioinformatics. Journal of Applied Statistics, 36(2):235–236, February 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kornak:2009:NVB

[1846] John Kornak, Bruce Dunham, Deborah A. Hall, and Mark P. Haggard. Nonlinear voxel-based modelling of the haemodynamic response in fMRI. Journal of Applied Statistics, 36(3):237–253, March 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chiu:2009:GWM

[1847] Wen-Chih Chiu. Generally weighted moving average control charts with fast initial response features. Journal of Applied Statistics, 36(3):255–275, March 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2009:BDG

[1848] Saralees Nadarajah. A bivariate distribution with gamma and beta marginals with application to drought data. Journal of Applied Statistics, 36(3):277– 301, March 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2009:OBR

[1849] Li Wang, Scott M. Kowalski, and G. Geoffrey Vining. Orthogonal blocking of response surface split-plot designs. Journal of Applied Statistics, 36(3):303–321, March 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Santos:2009:LTV

[1850] João Santos and Solange Leite. Longterm variability of the temperature time series recorded in Lisbon. Journal of Applied Statistics, 36(3):323–337, March 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Franses:2009:THR

[1851] Philip Hans Franses, Bert de Groot, and Rianne Legerstee. Testing for harmonic regressors. Journal of Applied Statistics, 36(3):339–346, March 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Imon:2009:DRD

[1852] A. H. M. Rahmatullah Imon. Deletion residuals in the detection of heterogeneity of variances in linear regression. Journal of Applied Statistics, 36(3):347–358, March 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez-Camblor:2009:TED

[1853] Pablo Martínez-Camblor and Aina Yañez. Testing the equality of diagnostic effectiveness of one measure with respect to k different features. Journal of Applied Statistics, 36(4):359–367, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cankaya:2009:QMR

[1854] E. Çankaya and N. R. J. Fieller. Quantal models: a review with additional methodological development. Journal of Applied Statistics, 36(4):369–384, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bettendorf:2009:DDR

[1855] L. Bettendorf, S. A. van der Geest, and G. H. Kuper. Do daily retail gasoline prices adjust asymmetrically? Journal of Applied Statistics, 36(4):385–397, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weiss:2009:MCP

[1856] Christian H. Weiß. Monitoring correlated processes with binomial marginals. Journal of Applied Statistics, 36(4):399–414, April 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2009:ILC

[1857] S. H. Lin and R. S. Wang. Inferences on a linear combination of K multivariate normal mean vectors. Journal of Applied Statistics, 36(4):415–428, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahmoudvand:2009:TNC

[1858] Rahim Mahmoudvand and Hossein Hassani. Two new confidence intervals for the coefficient of variation in a normal distribution. Journal of Applied Statistics, 36(4):429–442, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ozaki:2009:BRP

[1859] Vitor A. Ozaki and Ralph S. Silva.

Bayesian ratemaking procedure of crop insurance contracts with skewed distribution. Journal of Applied Statistics, 36(4):443–452, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liang:2009:UBM

[1860] Qiwei Liang, Huajiang Li, Paul Mendes, Hans Roethig, and Kim Frost-Pineda. Using bootstrap method to evaluate the estimates of nicotine equivalents from linear mixed model and generalized estimating equation. Journal of Applied Statistics, 36(4): 453–463, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2009:BBD

[1861] Z. Q. John Lu. Bayesian biostatistics and diagnostic medicine. Journal of Applied Statistics, 36(4):465–466, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Srivastava:2009:TB

[1862] Mukesh Srivastava and M. Abbas. Topics in biostatistics. Journal of Applied Statistics, 36(4):467–468, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deviere:2009:ALD

[1863] Elisabeth Dévière. Analyzing linguistic data: a practical introduction to statistics using R. Journal of Applied Statistics, 36(4):471–472, April 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2009:SET

[1864] Seung-Hwan Lee. Some estimators and tests for accelerated hazards model using weighted cumulative hazard difference. Journal of Applied Statistics, 36(5):473–482, May 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shi:2009:BCI

[1865] Lei Shi, Hongyuan Sun, and Peng Bai. Bayesian confidence interval for difference of the proportions in a 2 × 2 table with structural zero. Journal of Applied Statistics, 36(5):483–494, May 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shah:2009:BEP

[1866] J. B. Shah and M. N. Patel. Bayesian estimation of parameters of mixed geometric failure models from Type I group censored sample. Journal of Applied Statistics, 36(5):495–506, May 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Habshah:2009:PDR

[1867] M. Habshah, M. R. Norazan, and A. H. M. Rahmatullah Imon. The performance of diagnostic-robust generalized potentials for the identification of multiple high leverage points in linear regression. Journal of Applied Statistics, 36(5):507–520, May 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2009:ICS

[1868] Wanbo Lu and Tzong-Ru Tsai. Interval censored sampling plans for the loglogistic lifetime distribution. Journal of Applied Statistics, 36(5):521–536, May 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Antonakis:2009:ANB

[1869] A. C. Antonakis and M. E. Sfakianakis.

Assessing naïve Bayes as a method for screening credit applicants. Journal of Applied Statistics, 36(5):537–545, May 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martins:2009:URT

[1870] Luis F. Martins. Unit root tests and dramatic shifts with infinite variance processes. Journal of Applied Statistics, 36(5):547–571, May 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Herrerias-Velasco:2009:GTS

[1871] JoséManuel Herrerías-Velasco, Rafael Herrerías-Pleguezuelo, and Johan Renévan Dorp. The generalized two-sided power distribution. Journal of Applied Statistics, 36(5):573–587, May 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hutson:2009:PTA

[1872] Alan D. Hutson, Maurie Markman, and Mark F. Brady. A permutation test approach to phase II historical control trials. Journal of Applied Statistics, 36(6):589–599, June 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baklizi:2009:OTS

[1873] Ayman Baklizi and B. M. Golam Kibria. One and two sample confidence intervals for estimating the mean of skewed populations: an empirical comparative study. Journal of Applied Statistics, 36(6):601–609, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shetty:2009:MCS

[1874] Veena Shetty, Christopher H. Morrell, and Samer S. Najjar. Modeling a cross-sectional response variable with longitudinal predictors: an example of pulse pressure and pulse wave velocity. Journal of Applied Statistics, 36(6):611–619, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jammalamadaka:2009:AMD

[1875] S. Rao Jammalamadaka, Md. Aleemuddin Siddiqi, and Kaushik Ghosh. Analysis of microtubule dynamics using growth curve models. Journal of Applied Statistics, 36(6):621–631, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Thakkinstian:2009:CBP

[1876] Ammarin Thakkinstian, John R. Thompson, Cosetta Minelli, and John Attia. Choosing between per-genotype, per-allele, and trend approaches for initial detection of gene–disease association. Journal of Applied Statistics, 36(6):633–646, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Escarela:2009:CBM

[1877] Gabriel Escarela, Luis Carlos Pérez-Ruíz, and Russell J. Bowater. A copula-based Markov chain model for the analysis of binary longitudinal data. Journal of Applied Statistics, 36(6):647–657, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2009:BIM

[1878] Xiaomo Jiang and Sankaran Mahadevan. Bayesian inference method for model validation and confidence extrapolation. Journal of Applied Statistics, 36(6):659–677, June 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Trivez:2009:DLS

[1879] F. Javier Trívez and Beatriz Catalán.

Detecting level shifts in ARMA– GARCH (1,1) Models. Journal of Applied Statistics, 36(6):679–697, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Militino:2009:BRT

[1880] Ana F. Militino. Book review: Time Series Analysis. Journal of Applied Statistics, 36(6):699–700, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Robinson:2009:BRA

[1881] Andrew Robinson. Book review: Adaptive design theory and implementation using SAS and R. Journal of Applied Statistics, 36(6):701–702, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pienaar:2009:BRP

[1882] Jacques A. Pienaar. Book review: Probability and statistics with R. Journal of Applied Statistics, 36(6):703– 704, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4708].

Pienaar:2009:BRC

[1883] Jacques A. Pienaar. Book review: Control Techniques for Complex Networks. Journal of Applied Statistics, 36(6):707–708, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pienaar:2009:BRR

[1884] Jacques A. Pienaar. Book review: Random networks for communication: from statistical physics to information systems. Journal of Applied Statistics, 36(6):707–708, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maruotti:2009:FNH

[1885] Antonello Maruotti. Fairness of the national health service in Italy: a bivariate correlated random effects model. Journal of Applied Statistics, 36(7): 709–722, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dumangane:2009:DIS

[1886] Montezuma Dumangane, Nicoletta Rosati, and Anna Volossovitch. Departure from independence and stationarity in a handball match. Journal of Applied Statistics, 36(7):723–741, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Correa:2009:EOG

[1887] Alexander A. Correa, Pere Grima, and Xavier Tort-Martorell. Experimentation order with good properties for 2^{k} factorial designs. Journal of Applied Statistics, 36(7):743–754, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2009:MRS

[1888] Rabindra Nath Das and Sung H. Park.

A measure of robust slope-rotatability for second-order response surface experimental designs. Journal of Applied Statistics, 36(7):755–767, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2009:ARS

[1889] Binbing Yu. Approximating the risk score for disease diagnosis using MARS. Journal of Applied Statistics, 36(7):769–778, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ureche-Rangau:2009:MVT

[1890] Loredana Ureche-Rangau and Quiterie de Rorthays. More on the volatilitytrading volume relationship in emerging markets: The Chinese stock market. Journal of Applied Statistics, 36(7):779–799, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ciavolino:2009:SEM

[1891] E. Ciavolino and J. J. Dahlgaard. Simultaneous Equation Model based on the generalized maximum entropy for studying the effect of management factors on enterprise performance. Journal of Applied Statistics, 36(7):801–815, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hanck:2009:CSC

[1892] Christoph Hanck. Cross-sectional correlation robust tests for panel cointegration. Journal of Applied Statistics, 36(7):817–833, July 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

He:2009:SAE

[1893] Bo He and Clyde F. Martin. Statistical analysis of error for fourthorder ordinary differential equation solvers. Journal of Applied Statistics, 36(8):835–852, August 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bastida:2009:BRC

[1894] Agustín Hernández Bastida, Emilio Gómez Déniz, and JoséMaría Pérez Sánchez. Bayesian robustness of the compound Poisson distribution under bidimensional prior: an application to the collective risk model. Journal of Applied Statistics, 36(8):853–869, August 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Machado:2009:DAB

[1895] Carla Machado, Carlos Daniel Paulino, and Francisco Nunes. Deprivation analysis based on Bayesian latent class models. Journal of Applied Statistics, 36(8):871–891, August 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dutta:2009:OCD

[1896] Ganesh Dutta, Premadhis Das, and Nripes K. Mandal. Optimum covariate designs in split-plot and strip-plot design set-ups. Journal of Applied Statistics, 36(8):893–906, August 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeLuca:2009:ACR

[1897] Giovanni De Luca and Giorgia Rivieccio. Archimedean copulae for risk measurement. Journal of Applied Statistics, 36(8):907–924, August 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vonta:2009:FM

[1898] Filia Vonta. The frailty model. Journal of Applied Statistics, 36(8):927–928, August 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baxter:2009:IGL

[1899] Paul D. Baxter. An introduction to generalised linear models, third edition. Journal of Applied Statistics, 36 (8):929–930, August 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gajewski:2009:EDD

[1900] B. J. Gajewski, R. Lee, M. Bott, U. Piamjariyakul, and R. L. Taunton. On estimating the distribution of data envelopment analysis efficiency scores: an application to nursing homes’ care planning process. Journal of Applied Statistics, 36(9):933–944, September 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boutahar:2009:CNP

[1901] Mohamed Boutahar. Comparison of non-parametric and semi-parametric tests in detecting long memory. Journal of Applied Statistics, 36(9):945–972, September 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Muggeo:2009:SRM

[1902] Vito M. R. Muggeo, Massimo Attanasio, and Mariano Porcu. A segmented regression model for event history data: an application to the fertility patterns in Italy. Journal of Applied Statistics, 36(9):973–988, September 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2009:MCS

[1903] P. Singh. Multiple comparisons of several populations with more than one control with respect to scale parameter. Journal of Applied Statistics, 36 (9):989–998, September 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pal:2009:ODE

[1904] Manisha Pal and Nripes K. Mandal. Optimum designs for estimation of optimum point under cost constraint. Journal of Applied Statistics, 36(9): 999–1008, September 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2009:FMD

[1905] Prasun Das, Arup Kumar Das, and Saddam Hossain. Forecasting models for developing control scheme to improve furnace run length. Journal of Applied Statistics, 36(9):1009–1019, September 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aslam:2009:GAS

[1906] Muhammad Aslam and Chi-Hyuck Jun. A group acceptance sampling plan for truncated life test having Weibull distribution. Journal of Applied Statistics, 36(9):1021–1027, September 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mazucheli:2009:IMM

[1907] J. Mazucheli, J. A. Achcar, E. A.

Coelho-Barros, and F. Louzada-Neto. Infant mortality model for lifetime data. Journal of Applied Statistics, 36 (9):1029–1036, September 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Parsons:2009:ARD

[1908] N. R. Parsons, S. G. Gilmour, and R. N. Edmondson. Analysis of robust design experiments with timedependent ordinal response characteristics: a quality improvement study from the horticulture industry. Journal of Applied Statistics, 36(9):1037–1054, September 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hurtgen:2009:SSI

[1909] Holger Hürtgen and Daniel Gervini.

Semiparametric shape-invariant models for periodic data. Journal of Applied Statistics, 36(10):1055–1065, October 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Conlon:2009:HBM

[1910] E. M. Conlon, B. L. Postier, B. A. Methé, K. P. Nevin, and D. R. Lovley. Hierarchical Bayesian metaanalysis models for cross-platform microarray studies. Journal of Applied Statistics, 36(10):1067–1085, October 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Skinner:2009:SME

[1911] G. K. Skinner and G. H. Freeman. Soccer matches as experiments: how often does the ‘best’ team win? Journal of Applied Statistics, 36(10):1087–1095, October 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alghalith:2009:ECS

[1912] Moawia Alghalith. Empirical comparative statics under price and output uncertainty. Journal of Applied Statistics, 36(10):1097–1100, October 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Crocetta:2009:MLE

[1913] Corrado Crocetta and Nicola Loperfido. Maximum likelihood estimation of correlation between maximal oxygen consumption and the 6-min walk test in patients with chronic heart failure. Journal of Applied Statistics, 36 (10):1101–1108, October 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harel:2009:EAI

[1914] Ofer Harel. The estimation of R^{2} and adjusted R^{2} in incomplete data sets using multiple imputation. Journal of Applied Statistics, 36(10):1109–1118, October 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martin:2009:ADC

[1915] Nirian Martín and Leandro Pardo. On the asymptotic distribution of Cook’s distance in logistic regression models. Journal of Applied Statistics, 36(10): 1119–1146, October 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rieger:2009:TVH

[1916] R. H. Rieger and C. R. Weinberg. Testing for violations of the homogeneity needed for conditional logistic regression. Journal of Applied Statistics, 36 (10):1147–1157, October 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balamurali:2009:DVT

[1917] S. Balamurali and Chi-Hyuck Jun. Designing of a variables two-plan system by minimizing the average sample number. Journal of Applied Statistics, 36 (10):1159–1172, October 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ugarte:2009:LDA

[1918] M. Dolores Ugarte. Longitudinal data analysis. Journal of Applied Statistics, 36(10):1175–1176, October 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Salini:2009:BNC

[1919] Silvia Salini and Ron S. Kenett.

Bayesian networks of customer satisfaction survey data. Journal of Applied Statistics, 36(11):1177–1189, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

He:2009:PDH

[1920] Angela W. W. He and Alan T. K. Wan. Predicting daily highs and lows of exchange rates: a cointegration analysis. Journal of Applied Statistics, 36(11): 1191–1204, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2009:BPA

[1921] Chien-Wei Wu and Tsai-Yu Lin. A Bayesian procedure for assessing process performance based on the thirdgeneration capability index. Journal of Applied Statistics, 36(11):1205–1223, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jeong:2009:NEV

[1922] Seok-Oh Jeong and Kee-Hoon Kang. Nonparametric estimation of value-at-risk. Journal of Applied Statistics, 36 (11):1225–1238, November 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berry:2009:ERP

[1923] Kenneth J. Berry, Janis E. Johnston, and Paul W. Mielke, Jr. Exact and resampling probability values for the Piccarreta nominal–ordinal index of association. Journal of Applied Statistics, 36(11):1239–1249, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sun:2009:PRS

[1924] Juan Sun, Xiaohui Ouyang, Hidekatsu Yoshioka, Wenli Wang, Chun Fan, Hongwei Li, Jianru Wang, Yalin Liu, Li Su, Heping Ma, Ying liu, Yuwen Zhang, Xingguang Zhang, Xuemei Wang, and Yueling Hu. A progressive rise in stomach cancer-related mortality rate during 1970–1995 in Japanese individuals over 85 years of age. Journal of Applied Statistics, 36 (11):1251–1258, November 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Longford:2009:AAZ

[1925] Nicholas T. Longford. Analysis of allzero binomial outcomes with borderline and equilibrium priors. Journal of Applied Statistics, 36(11):1259–1265, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dagne:2009:BHD

[1926] Getachew A. Dagne and James Snyder. Bayesian hierarchical duration model for repeated events: an application to behavioral observations. Journal of Applied Statistics, 36(11): 1267–1279, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sampath:2009:EUS

[1927] S. Sampath, V. Varalakshmi, and R. Geetha. Estimation under systematic sampling schemes for parabolic populations. Journal of Applied Statistics, 36(11):1281–1292, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mukhopadhyay:2009:QDG

[1928] S. Mukhopadhyay and S. W. Looney. Quantile dispersion graphs to compare the efficiencies of cluster randomized designs. Journal of Applied Statistics, 36(11):1293–1305, November 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bradley:2009:TME

[1929] Jonathan R. Bradley and David L. Farnsworth. Testing for mutual exclusivity. Journal of Applied Statistics, 36 (11):1307–1314, November 2009. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DAmbra:2009:CCA

[1930] Luigi D’Ambra, Onur Köksoy, and Biagio Simonetti. Cumulative correspondence analysis of ordered categorical data from industrial experiments. Journal of Applied Statistics, 36(12): 1315–1328, December 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Giancristofaro:2009:ESE

[1931] Rosa Arboretti Giancristofaro, Stefano Bonnini, and Luigi Salmaso. Employment status and education/ employment relationship of PhD graduates from the University of Ferrara. Journal of Applied Statistics, 36(12): 1329–1344, December 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Caudill:2009:MTN

[1932] S. B. Caudill and F. G. Mixon, Jr. More on testing the normality assumption in the Tobit Model. Journal of Applied Statistics, 36(12):1345–1352, December 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2009:TZI

[1933] Y. Zhao, A. H. Lee, V. Burke, and K. K. W. Yau. Testing for zero-inflation in count series: application to occupational health. Journal of Applied Statistics, 36(12):1353–1359, December 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gjestvang:2009:IRR

[1934] Christopher R. Gjestvang and Sarjinder Singh. An improved randomized response model: estimation of mean. Journal of Applied Statistics, 36(12): 1361–1367, December 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lindstrom:2009:FMS

[1935] J. Fredrik Lindström and H. E. T. Holgersson. Forecast mean squared error reduction in the VAR(1) process. Journal of Applied Statistics, 36(12): 1369–1384, December 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alfaro:2009:CRA

[1936] J. L. Alfaro and J. Fco. Ortega. A comparison of robust alternatives to Hotelling’s T^{2} control chart. Journal of Applied Statistics, 36(12):1385–1396, December 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cooray:2010:ASD

[1937] Kahadawala Cooray and Malwane M. A. Ananda. Analyzing survival data with highly negatively skewed distribution: The Gompertz–sinh family. Journal of Applied Statistics, 37(1):1–11, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karagrigoriou:2010:ANC

[1938] A. Karagrigoriou, C. Koukouvinos, and K. Mylona. On the advantages of the non-concave penalized likelihood model selection method with minimum prediction errors in large-scale medical studies. Journal of Applied Statistics, 37(1):13–24, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saleem:2010:BAM

[1939] M. Saleem, M. Aslam, and P. Economou. On the Bayesian analysis of the mixture of power function distribution using the complete and the censored sample. Journal of Applied Statistics, 37(1):25–40, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Micheas:2010:BPA

[1940] Athanasios C. Micheas and Yuqiang Peng. Bayesian Procrustes analysis with applications to hydrology. Journal of Applied Statistics, 37(1):41–55, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beccarini:2010:EOV

[1941] Andrea Beccarini. Eliminating the omitted variable bias by a regimeswitching approach. Journal of Applied Statistics, 37(1):57–75, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Parchami:2010:NGP

[1942] A. Parchami and M. Mashinchi. A new generation of process capability indices. Journal of Applied Statistics, 37(1):77–89, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Simonetti:2010:ADT

[1943] Biagio Simonetti, Eric J. Beh, and Luigi D’Ambra. The analysis of dependence for three ways contingency tables with ordinal variables: a case study of patient satisfaction data. Journal of Applied Statistics, 37(1):91–103, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DalBello:2010:OPP

[1944] L. H. A. Dal Bello and A. F. C. Vieira.

Optimization of a product performance using mixture experiments. Journal of Applied Statistics, 37(1):105–117, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Perry:2010:ITS

[1945] Marcus B. Perry and Joseph J. Pignatiello Jr. Identifying the time of step change in the mean of autocorrelated processes. Journal of Applied Statistics, 37(1):119–136, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fuentes-Garcia:2010:NAC

[1946] R. Fuentes-García and S. G. Walker. A new approach to classification. Journal of Applied Statistics, 37(1):137– 146, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scagliarini:2010:IAD

[1947] Michele Scagliarini. Inference on C_{pk} for autocorrelated data in the presence of random measurement errors. Journal of Applied Statistics, 37(1): 147–158, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ben-Ahmed:2010:GLS

[1948] K. Ben-Ahmed, A. Bouratbine, and M.-A. El-Aroui. Generalized linear spatial models in epidemiology: a case study of zoonotic cutaneous leishmaniasis in Tunisia. Journal of Applied Statistics, 37(1):159–170, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cooray:2010:GGD

[1949] Kahadawala Cooray. Generalized Gumbel distribution. Journal of Applied Statistics, 37(1):171–179, January 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2010:BAD

[1950] Yangxin Huang. A Bayesian approach in differential equation dynamic models incorporating clinical factors and covariates. Journal of Applied Statistics, 37(2):181–199, February 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cheong:2010:EHP

[1951] Chin Wen Cheong. Estimating the Hurst parameter in financial time series via heuristic approaches. Journal of Applied Statistics, 37(2):201–214, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Leong:2010:TSC

[1952] Chee Kian Leong and Weihong Huang.

Testing for spurious and cointegrated regressions: a wavelet approach. Journal of Applied Statistics, 37(2):215– 233, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bouguila:2010:SMC

[1953] Nizar Bouguila, Jian Han Wang, and A. Ben Hamza. Software modules categorization through likelihood and Bayesian analysis of finite Dirichlet mixtures. Journal of Applied Statistics, 37(2):235–252, February 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baio:2010:BHM

[1954] Gianluca Baio and Marta Blangiardo.

Bayesian hierarchical model for the prediction of football results. Journal of Applied Statistics, 37(2):253–264, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2010:ERR

[1955] Zongrun Wang, Weitao Wu, Chao Chen, and Yanju Zhou. The exchange rate risk of Chinese yuan: Using VaR and ES based on extreme value theory. Journal of Applied Statistics, 37 (2):265–282, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akcakoca:2010:IDC

[1956] Hamdi Akcakoca and Nurhak Sutcu.

Investigation of the diesel consumption for trucks at an overburden stripping area by SPC study. Journal of Applied Statistics, 37(2):283–298, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mergen:2010:AAS

[1957] A. Erhan Mergen and Z. Seyda Deligonul. Assessment of acceptance sampling plans using posterior distribution for a dependent process. Journal of Applied Statistics, 37(2):299–307, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gani:2010:SVR

[1958] Walid Gani, Hassen Taleb, and Mohamed Limam. Support vector regression based residual control charts. Journal of Applied Statistics, 37(2): 309–324, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2010:GSD

[1959] Jing Wang. Gibbs sampling in DPbased nonlinear mixed effects models. Journal of Applied Statistics, 37 (2):325–340, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Serroyen:2010:CPD

[1960] Jan Serroyen, Liesbeth Bruckers, Geert Rogiers, and Geert Molenberghs. Characterizing persistent disturbing behavior using longitudinal and multivariate techniques. Journal of Applied Statistics, 37(2):341–355, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Militino:2010:BRS

[1961] Ana F. Militino. Book review: Statistics and data with R. Journal of Applied Statistics, 37(2):357–358, February 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4699].

Chang:2010:LSC

[1962] Jung Hsien Chang and Mao Wei Hung.

Liquidity spreads in the corporate bond market: Estimation using a semiparametric model. Journal of Applied Statistics, 37(3):359–374, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Senthilkumar:2010:CST

[1963] D. Senthilkumar and D. Muthuraj. Construction and selection of tightened-normal-tightened variables sampling scheme of type TNTVSS(n_{1}, n_{2}; k). Journal of Applied Statistics, 37(3):375–390, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alshunnar:2010:CFI

[1964] Fawziah S. Alshunnar, Mohammad Z.

Raqab, and Debasis Kundu. On the comparison of the Fisher information of the log-normal and generalized Rayleigh distributions. Journal of Applied Statistics, 37(3):391–404, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aslam:2010:DAS

[1965] Muhammad Aslam and Chi-Hyuck Jun. A double acceptance sampling plan for generalized log-logistic distributions with known shape parameters. Journal of Applied Statistics, 37(3):405–414, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demattei:2010:CTI

[1966] Christophe Demattei. Le Cam theorem on interval division by randomly chosen points: Pedagogical explanations and application to temporal cluster detection. Journal of Applied Statistics, 37(3):415–424, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chan:2010:IES

[1967] Ling-Yau Chan and Rahul Mukerjee. Interval estimation of a small proportion via inverse sampling. Journal of Applied Statistics, 37(3):425–433, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Figueiredo:2010:SNC

[1968] C. C. Figueiredo, H. Bolfarine, M. C. Sandoval, and C. R. O. P. Lima. On the skew-normal calibration model. Journal of Applied Statistics, 37(3):435– 451, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jun:2010:VRG

[1969] Chi-Hyuck Jun, Hyeseon Lee, Sang-Ho Lee, and S. Balamurali. A variables repetitive group sampling plan under failure-censored reliability tests for Weibull distribution. Journal of Applied Statistics, 37(3):453–460, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2010:NSM

[1970] Chung-Ho Chen. A note on some modified Pulak and Al-Sultan’s model. Journal of Applied Statistics, 37(3): 461–472, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Montanari:2010:SNF

[1971] Angela Montanari and Cinzia Viroli.

A skew-normal factor model for the analysis of student satisfaction towards university courses. Journal of Applied Statistics, 37(3):473–487, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demirtas:2010:DBR

[1972] Hakan Demirtas. A distance-based rounding strategy for post-imputation ordinal data. Journal of Applied Statistics, 37(3):489–500, March 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rios:2010:AMP

[1973] Martín Ríos and Toni Monleón-Getino.

Application of a Markovian process to the calculation of mean time equilibrium in a genetic drift model. Journal of Applied Statistics, 37(3):501–513, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:2010:BAB

[1974] Kanti V. Mardia. Bayesian analysis for bivariate von Mises distributions. Journal of Applied Statistics, 37(3):515–528, March 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baxter:2010:CPR

[1975] Paul D. Baxter and Paul R. Marchant. The cross-product ratio in bivariate lognormal and gamma distributions, with an application to non-randomized trials. Journal of Applied Statistics, 37(4):529–536, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kourlaba:2010:DAC

[1976] Georgia Kourlaba and Demosthenes Panagiotakos. The diagnostic accuracy of a composite index increases as the number of partitions of the components increases and when specific weights are assigned to each component. Journal of Applied Statistics, 37(4):537–554, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aslam:2010:TTA

[1977] Muhammad Aslam, Debasis Kundu, and Munir Ahmad. Time truncated acceptance sampling plans for generalized exponential distribution. Journal of Applied Statistics, 37(4):555–566, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nassar:2010:GSS

[1978] M. M. Nassar, S. M. Khamis, and S. S. Radwan. Geometric sample size determination in Bayesian analysis. Journal of Applied Statistics, 37(4):567–575, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cadima:2010:EBS

[1979] Jorge Cadima, Francisco Lage Calheiros, and Isabel P. Preto. The eigenstructure of block-structured correlation matrices and its implications for principal component analysis. Journal of Applied Statistics, 37(4):577–589, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cribari-Neto:2010:AIH

[1980] Francisco Cribari-Neto and Maria da Glória A. Lima. Approximate inference in heteroskedastic regressions: a numerical evaluation. Journal of Applied Statistics, 37(4):591–615, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2010:MPM

[1981] H. Jiang, M. Xie, and L. C. Tang. On MLEs of the parameters of a modified Weibull distribution for progressively type-2 censored samples. Journal of Applied Statistics, 37(4):617–627, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samawi:2010:OST

[1982] Hani M. Samawi, Mohammed Al-Haj Ebrahem, and Noha Al-Zubaidin. An optimal sign test for one-sample bivariate location model using an alternative bivariate ranked-set sample. Journal of Applied Statistics, 37(4):629–650, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pereira:2010:SAE

[1983] Luís Nobre Pereira and Pedro Simões Coelho. Small area estimation of mean price of habitation transaction using time-series and cross-sectional arealevel models. Journal of Applied Statistics, 37(4):651–666, April 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Majeske:2010:QBP

[1984] Karl D. Majeske, Terri Lynch-Caris, and Janet Brelin-Fornari. Quantifying R^{2} bias in the presence of measurement error. Journal of Applied Statistics, 37(4):667–677, April 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balduck:2010:ECT

[1985] Anne-Line Balduck, Anita Prinzie, and Marc Buelens. The effectiveness of coach turnover and the effect on home team advantage, team quality and team ranking. Journal of Applied Statistics, 37(4):679–689, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Adeyemi:2010:REM

[1986] Shola Adeyemi, Thierry J. Chaussalet, Haifeng Xie, and Md Asaduzaman. Random effects models for operational patient pathways. Journal of Applied Statistics, 37(4):691–701, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nielsen:2010:BRG

[1987] Sören Feodor Nielsen. Book review: Generalized linear models for insurance data. Journal of Applied Statistics, 37(4):703–??, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4702].

Lu:2010:BMD

[1988] Z. Q. John Lu. Bayesian methods for data analysis, third edition. Journal of Applied Statistics, 37(4):705–706, April 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Michlin:2010:SBT

[1989] Y. H. Michlin and G. Grabarnik.

Search boundaries of the truncated discrete sequential test. Journal of Applied Statistics, 37(5):707–724, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rolfe:2010:LCP

[1990] M. I. Rolfe, K. Mengersen, G. Beadle, K. Vearncombe, B. Andrew, H. L. Johnson, and C. Walsh. Latent class piecewise linear trajectory modelling for short-term cognition responses after chemotherapy for breast cancer patients. Journal of Applied Statistics, 37(5):725–738, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Morales:2010:LOS

[1991] Marco Morales. Lag order selection for an optimal autoregressive covariance matrix estimator. Journal of Applied Statistics, 37(5):739–748, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jacques:2010:EMB

[1992] J. Jacques and C. Biernacki. Extension of model-based classification for binary data when training and test populations differ. Journal of Applied Statistics, 37(5):749–766, May 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2010:PTM

[1993] Hyo-Il Park and Seung-Man Hong. A permutation test for multivariate data with grouped components. Journal of Applied Statistics, 37(5):767–778, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

An:2010:SAD

[1994] Hyonggin An, Roderick J. A. Little, and Andrea Bozoki. A statistical algorithm for detecting cognitive plateaus in Alzheimer’s disease. Journal of Applied Statistics, 37(5):779–789, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiao:2010:GLR

[1995] Zhiguo Xiao, Jun Shao, and Mari Palta. GMM in linear regression for longitudinal data with multiple covariates measured with error. Journal of Applied Statistics, 37(5):791–805, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2010:RRE

[1996] Fu-Kwun Wang and Yung-Fu Cheng. Robust regression for estimating the Burr XII parameters with outliers. Journal of Applied Statistics, 37(5): 807–819, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anzures-Cabrera:2010:CRL

[1997] J. Anzures-Cabrera and J. L. Hutton. Competing risks, left truncation and late entry effect in A-bomb survivors cohort. Journal of Applied Statistics, 37(5):821–831, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Serroyen:2010:FES

[1998] Jan Serroyen, Geert Molenberghs, Marc Aerts, Ellen Vloeberghs, Peter Paul De Deyn, and Geert Verbeke.

Flexible estimation of serial correlation in nonlinear mixed models. Journal of Applied Statistics, 37(5):833–846, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bortoluzzo:2010:TVA

[1999] Adriana B. Bortoluzzo, Pedro A.

Morettin, and Clelia M. C. Toloi. Timevarying autoregressive conditional duration model. Journal of Applied Statistics, 37(5):847–864, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coelho-Barros:2010:LPM

[2000] Emílio Augusto Coelho-Barros, Jorge Alberto Achcar, and Josmar Mazucheli. Longitudinal Poisson modeling: an application for CD4 counting in HIVinfected patients. Journal of Applied Statistics, 37(5):865–880, May 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McCullough:2010:BPI

[2001] B. D. McCullough and Thomas P.

McWilliams. Baseball players with the initial “K” do not strike out more often. Journal of Applied Statistics, 37(6):881–891, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Benth:2010:AMW

[2002] Jūraté Šaltyté Benth and Fred Espen Benth. Analysis and modelling of wind speed in New York. Journal of Applied Statistics, 37(6):893–909, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Copas:2010:BUG

[2003] Andrew J. Copas and Shaun R. Seaman. Bias from the use of generalized estimating equations to analyze incomplete longitudinal binary data. Journal of Applied Statistics, 37(6):911–922, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Islam:2010:MLR

[2004] M. Qamarul Islam and Moti L. Tiku. Multiple linear regression model with stochastic design variables. Journal of Applied Statistics, 37(6):923–943, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Matawie:2010:BDE

[2005] K. M. Matawie and A. Assaf. Bayesian and DEA efficiency modelling: an application to hospital foodservice operations. Journal of Applied Statistics, 37(6):945–953, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Torng:2010:NNC

[2006] Chau-Chen Torng, Chun-Chieh Tseng, and Pei-Hsi Lee. Non-normality and combined double sampling and variable sampling interval control charts. Journal of Applied Statistics, 37(6):955– 967, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Famoye:2010:BNB

[2007] Felix Famoye. On the bivariate negative binomial regression model. Journal of Applied Statistics, 37(6):969–981, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Juutilainen:2010:HCI

[2008] Ilmari Juutilainen and Juha Röning.

How to compare interpretatively different models for the conditional variance function. Journal of Applied Statistics, 37(6):983–997, June 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Koyuncu:2010:IEP

[2009] Nursel Koyuncu and Cem Kadilar.

On improvement in estimating population mean in stratified random sampling. Journal of Applied Statistics, 37(6):999–1013, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2010:SAC

[2010] Young-Ju Kim. Semiparametric analysis for case-control studies: a partial smoothing spline approach. Journal of Applied Statistics, 37(6):1015–1025, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeIaco:2010:STC

[2011] Sandra De Iaco. Space–time correlation analysis: a comparative study. Journal of Applied Statistics, 37(6):1027–1041, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsai:2010:EBM

[2012] Miao-Yu Tsai. Extended Bayesian model averaging for heritability in twin studies. Journal of Applied Statistics, 37(6):1043–1058, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Srivastava:2010:AVC

[2013] Mukesh Srivastava. Analysis of variance and covariance: how to choose and construct models for the life sciences. Journal of Applied Statistics, 37(6):1059–1060, June 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kao:2010:NOS

[2014] Shih-Chou Kao. Normalization of the origin-shifted exponential distribution for control chart construction. Journal of Applied Statistics, 37(7):1067–1087, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Han:2010:DCR

[2015] Dong Han, Fugee Tsung, Yanting Li, and Jinguo Xian. Detection of changes in a random financial sequence with a stable distribution. Journal of Applied Statistics, 37(7):1089–1111, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scollnik:2010:BSI

[2016] David P. M. Scollnik. Bayesian statistical inference for start-up demonstration tests with rejection of units upon observing d failures. Journal of Applied Statistics, 37(7):1113–1121, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fukuda:2010:PCG

[2017] Kosei Fukuda. Parameter changes in GARCH model. Journal of Applied Statistics, 37(7):1123–1135, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fiserova:2010:TLS

[2018] Eva Fišerová and Karel Hron. Total least squares solution for compositional data using linear models. Journal of Applied Statistics, 37(7):1137–1152, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Heron:2010:BDL

[2019] Elizabeth A. Heron and Cathal D.

Walsh. Bayesian discrete latent spatial modeling of crack initiation in orthopaedic hip replacement bone cement. Journal of Applied Statistics, 37(7):1153–1171, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2010:EID

[2020] Ming-Yuan Leon Li and Chun-Nan Chen. Examining the interrelation dynamics between option and stock markets using the Markov-switching vector error correction model. Journal of Applied Statistics, 37(7):1173–1191, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Castro-Kuriss:2010:GFT

[2021] Claudia Castro-Kuriss, Diana M. Kelmansky, Víctor Leiva, and Elena J. Martínez. On a goodness-of-fit test for normality with unknown parameters and type-II censored data. Journal of Applied Statistics, 37(7):1193–1211, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2010:MSS

[2022] Qin Yu, Wan Tang, Sue Marcus, Yan Ma, Hui Zhang, and Xin Tu. Modeling sensitivity and specificity with a time-varying reference standard within a longitudinal setting. Journal of Applied Statistics, 37(7):1213–1230, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Antzoulakos:2010:RRS

[2023] Demetrios L. Antzoulakos and Athanasios C. Rakitzis. Runs rules schemes for monitoring process variability. Journal of Applied Statistics, 37(7):1231–1247, July 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahmoud:2010:MSL

[2024] Mahmoud A. Mahmoud, J. P. Morgan, and William H. Woodall. The monitoring of simple linear regression profiles with two observations per sample. Journal of Applied Statistics, 37 (8):1249–1263, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bagheban:2010:GUA

[2025] Alireza Akbarzadeh Bagheban and Farid Zayeri. A generalization of the uniform association model for assessing rater agreement in ordinal scales. Journal of Applied Statistics, 37(8): 1265–1273, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhattacharya:2010:ECY

[2026] A. Bhattacharya. Estimating crop yield via Gaussian quadrature. Journal of Applied Statistics, 37(8):1275–1281, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shu:2010:MIF

[2027] Ming-Hung Shu and Hsien-Chung Wu.

Monitoring imprecise fraction of nonconforming items using p control charts. Journal of Applied Statistics, 37 (8):1283–1297, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sarstedt:2010:TUH

[2028] Marko Sarstedt and Christian M.

Ringle. Treating unobserved heterogeneity in PLS path modeling: a comparison of FIMIX-PLS with different data analysis strategies. Journal of Applied Statistics, 37(8):1299–1318, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lyu:2010:MBA

[2029] JrJung Lyu and MingNan Chen. Measurement of bivariate attributes using a novel statistical model. Journal of Applied Statistics, 37(8):1319–1334, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Principato:2010:IGS

[2030] Fabio Principato, Angela Vullo, and Domenica Matranga. On implementation of the Gibbs sampler for estimating the accuracy of multiple diagnostic tests. Journal of Applied Statistics, 37 (8):1335–1354, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Souza:2010:BOA

[2031] Aparecida D. P. Souza and Helio S.

Migon. Bayesian outlier analysis in binary regression. Journal of Applied Statistics, 37(8):1355–1368, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aslam:2010:DTT

[2032] M. Aslam, C.-H. Jun, and M. Ahmad. Design of a time-truncated double sampling plan for a general life distribution. Journal of Applied Statistics, 37 (8):1369–1379, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demetrescu:2010:TSL

[2033] Matei Demetrescu, Uwe Hassler, and Adina I. Tarcolea. Testing for stationarity in large panels with crossdependence, and US evidence on unit labor cost. Journal of Applied Statistics, 37(8):1381–1397, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Androulakis:2010:RSA

[2034] Emmanouil Androulakis, Christos Koukouvinos, Kalliopi Mylona, and Filia Vonta. A real survival analysis application via variable selection methods for Cox’s proportional hazards model. Journal of Applied Statistics, 37(8):1399–1406, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chong:2010:IDD

[2035] Terence Tai-Leung Chong, Zimu Li, Haiqiang Chen, and Melvin J. Hinich. An investigation of duration dependence in the American stock market cycle. Journal of Applied Statistics, 37 (8):1407–1416, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hutchison:2010:BRH

[2036] Dougal Hutchison. Book review: Handbook of multilevel analysis. Journal of Applied Statistics, 37(8):1417–1418, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4703].

Vonta:2010:BRM

[2037] Ilia Vonta. Book review: Model selection and model averaging. Journal of Applied Statistics, 37(8):1419–1420, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4698].

Shade:2010:SDA

[2038] John Shade. Software for data analysis. Journal of Applied Statistics, 37 (8):1421–1422, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4697].

Anonymous:2010:C

[2039] Anonymous. Corrigendum. Journal of Applied Statistics, 37(8):1423, August 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Narayan:2010:NUR

[2040] Paresh Kumar Narayan and Stephan Popp. A new unit root test with two structural breaks in level and slope at unknown time. Journal of Applied Statistics, 37(9):1425–1438, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2010:BAI

[2041] Yousung Park and Bo-Seung Choi.

Bayesian analysis for incomplete multiway contingency tables with nonignorable nonresponse. Journal of Applied Statistics, 37(9):1439–1453, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aerts:2010:MSR

[2042] Marc Aerts, Niel Hens, and Jeffrey S. Simonoff. Model selection in regression based on pre-smoothing. Journal of Applied Statistics, 37(9):1455–1472, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mantalos:2010:ESG

[2043] Panagiotis Mantalos and Ghazi Shukur. The effect of spillover on the Granger causality test. Journal of Applied Statistics, 37(9):1473–1486, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ren:2010:NBH

[2044] Shiquan Ren, Hong Lai, Wenjing Tong, Mostafa Aminzadeh, Xuezhang Hou, and Shenghan Lai. Nonparametric bootstrapping for hierarchical data. Journal of Applied Statistics, 37(9): 1487–1498, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haghighi:2010:LDM

[2045] Firoozeh Haghighi and Mikhail Nikulin. On the linear degradation model with multiple failure modes. Journal of Applied Statistics, 37(9):1499–1507, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xie:2010:TVZ

[2046] Feng-Chang Xie, Jin-Guan Lin, and Bo-Cheng Wei. Testing for varying zero-inflation and dispersion in generalized Poisson regression models. Journal of Applied Statistics, 37(9): 1509–1522, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weston:2010:MSC

[2047] Claire L. Weston and John R. Thompson. Modeling survival in childhood cancer studies using two-stage non-mixture cure models. Journal of Applied Statistics, 37(9):1523–1535, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gomez-Deniz:2010:SBL

[2048] Emilio Gómez-Déniz and Enrique Calderín-Ojeda. A study of Bayesian local robustness with applications in actuarial statistics. Journal of Applied Statistics, 37(9):1537–1546, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Best:2010:FTF

[2049] D. J. Best, J. C. W. Rayner, and O. Thas. Four tests of fit for the beta-binomial distribution. Journal of Applied Statistics, 37(9):1547–1554, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nikoloulopoulos:2010:RCM

[2050] Aristidis K. Nikoloulopoulos and Dimitris Karlis. Regression in a copula model for bivariate count data.

Journal of Applied Statistics, 37(9): 1555–1568, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Agostinelli:2010:APO

[2051] Claudio Agostinelli and Luisa Bisaglia. ARFIMA processes and outliers: a weighted likelihood approach. Journal of Applied Statistics, 37(9): 1569–1584, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Puig:2010:SMM

[2052] Xavier Puig, Josep Ginebra, and Marti Font. The Sichel model and the mixing and truncation order. Journal of Applied Statistics, 37(9):1585–1603, September 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nurunnabi:2010:IMI

[2053] A. A. M. Nurunnabi, A. H. M. Rahmatullah Imon, and M. Nasser. Identification of multiple influential observations in logistic regression. Journal of Applied Statistics, 37(10):1605–1624, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kannan:2010:GEC

[2054] Nandini Kannan, Debasis Kundu, P. Nair, and R. C. Tripathi. The generalized exponential cure rate model with covariates. Journal of Applied Statistics, 37(10):1625–1636, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ostermark:2010:CPH

[2055] Ralf Östermark. Concurrent processing of heteroskedastic vector-valued mixture density models. Journal of Applied Statistics, 37(10):1637–1659, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Todem:2010:RER

[2056] D. Todem, Y. Zhang, A. Ismail, and W. Sohn. Random effects regression models for count data with excess zeros in caries research. Journal of Applied Statistics, 37(10):1661–1679, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hwang:2010:EFC

[2057] Yi-Ting Hwang, Jia-Jung Lai, and Shyh-Tyan Ou. Evaluations of FWERcontrolling methods in multiple hypothesis testing. Journal of Applied Statistics, 37(10):1681–1694, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ryden:2010:EIR

[2058] Jesper Rydén and Sven Erick Alm. The effect of interaction and rounding error in two-way ANOVA: example of impact on testing for normality. Journal of Applied Statistics, 37(10):1695–1701, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jones:2010:BMQ

[2059] Douglas H. Jones and Francis A. Méndez Mediavilla. A Bayesian method for query approximation. Journal of Applied Statistics, 37(10):1703–1715, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chiu:2010:CCF

[2060] Jing-Er Chiu and Tsen-I Kuo. Control charts for fraction nonconforming in a bivariate binomial process. Journal of Applied Statistics, 37(10): 1717–1728, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pryseley:2010:EPR

[2061] Assam Pryseley, Koen Mintiens, Katia Knapen, Yves Van der Stede, and Geert Molenberghs. Estimating precision, repeatability, and reproducibility from Gaussian and non-Gaussian data: a mixed models approach. Journal of Applied Statistics, 37(10):1729–1747, October 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2010:BAS

[2062] Jie Wang and James D. Stamey. A Bayesian algorithm for sample size determination for equivalence and noninferiority test. Journal of Applied Statistics, 37(10):1749–1759, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samimi:2010:MHS

[2063] Y. Samimi and A. Aghaie. Monitoring heterogeneous serially correlated usage behavior in subscription-based services. Journal of Applied Statistics, 37(10): 1761–1777, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ober:2010:BRM

[2064] Pieter Bastiaan Ober. Book review: Modeling with Data. Journal of Applied Statistics, 37(10):1779–??, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4715].

Srivastava:2010:BRA

[2065] Mukesh Srivastava. Book review:

Analysis of Variance and Covariance. Journal of Applied Statistics, 37(10): 1781–1782, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4694].

Klapper:2010:BRD

[2066] Jennifer Klapper. Book review: Discrete Fourier Analysis and Wavelets. Journal of Applied Statistics, 37(10): 1783–1784, October 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4712].

Lipovetsky:2010:DLC

[2067] Stan Lipovetsky. Double logistic curve in regression modeling. Journal of Applied Statistics, 37(11):1785–1793, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Graham:2010:PNS

[2068] M. A. Graham, S. W. Human, and S. Chakraborti. A Phase I nonparametric Shewhart-type control chart based on the median. Journal of Applied Statistics, 37(11):1795–1813, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Figini:2010:TBA

[2069] Silvia Figini, Paolo Giudici, and Pierpaolo Uberti. A threshold based approach to merge data in financial risk management. Journal of Applied Statistics, 37(11):1815–1824, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Conversano:2010:AMF

[2070] Claudio Conversano and Domenico Vistocco. Analysis of mutual funds’ management styles: a modeling, ranking and visualizing approach. Journal of Applied Statistics, 37(11):1825–1845, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Unkel:2010:IEF

[2071] Steffen Unkel, Nickolay T. Trendafilov, Abdel Hannachi, and Ian T. Jolliffe. Independent exploratory factor analysis with application to atmospheric science data. Journal of Applied Statistics, 37 (11):1847–1862, November 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2010:BTU

[2072] Man-Suk Oh and Dong Wan Shin. Bayesian tests for unit root and multiple breaks. Journal of Applied Statistics, 37(11):1863–1874, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Diana:2010:NSR

[2073] Giancarlo Diana and Pier Francesco Perri. New scrambled response models for estimating the mean of a sensitive quantitative character. Journal of Applied Statistics, 37(11):1875–1890, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Van:2010:CPD

[2074] Tai Vo Van and T. Pham-Gia. Clustering probability distributions. Journal of Applied Statistics, 37(11):1891–1910, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sparks:2010:EWC

[2075] Ross S. Sparks, Tim Keighley, and David Muscatello. Early warning CUSUM plans for surveillance of negative binomial daily disease counts. Journal of Applied Statistics, 37(11): 1911–1929, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bijwaard:2010:RIS

[2076] Govert E. Bijwaard. Regularity in individual shopping trips: implications for duration models in marketing. Journal of Applied Statistics, 37 (11):1931–1945, November 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Papalia:2010:DDP

[2077] Rosa Bernardini Papalia. Data disaggregation procedures within a maximum entropy framework. Journal of Applied Statistics, 37(11):1947–1959, November 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tanco:2010:WDE

[2078] Martín Tanco, Elisabeth Viles, Maria Jesus Álvarez, and Laura Ilzarbe. Why is not design of experiments widely used by engineers in Europe? Journal of Applied Statistics, 37(12):1961–1977, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Galea:2010:SSM

[2079] Manuel Galea, David Cademartori, and Filidor Vilca. The structural Sharpe model under t-distributions. Journal of Applied Statistics, 37(12): 1979–1990, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2010:CCS

[2080] Yafen Liu, Zhen He, M. Shamsuzzaman, and Zhang Wu. A combined control scheme for monitoring the frequency and size of an attribute event. Journal of Applied Statistics, 37(12): 1991–2013, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ross:2010:RNS

[2081] Gavin J. S. Ross, Prajneshu, and C. Sarada. Reparameterization of nonlinear statistical models: a case study. Journal of Applied Statistics, 37(12): 2015–2026, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Antoch:2010:ECP

[2082] Jaromír Antoch, Lubos Prchal, Maria Rosaria De Rosa, and Pascal Sarda. Electricity consumption prediction with functional linear regression using spline estimators. Journal of Applied Statistics, 37(12):2027–2041, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cheong:2010:OCS

[2083] Chin Wen Cheong. Optimal choice of sample fraction in univariate financial tail index estimation. Journal of Applied Statistics, 37(12):2043–2056, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lui:2010:NOR

[2084] Kung-Jong Lui and Kuang-Chao Chang. Notes on odds ratio estimation for a randomized clinical trial with noncompliance and missing outcomes. Journal of Applied Statistics, 37(12): 2057–2071, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gadre:2010:SSM

[2085] M. P. Gadre, K. A. Joshi, and R. N.

Rattihalli. A side sensitive modified group runs control chart to detect shifts in the process mean. Journal of Applied Statistics, 37(12):2073–2087, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Frisen:2010:EMS

[2086] Marianne Frisén, Eva Andersson, and Linus Schiöler. Evaluation of multivariate surveillance. Journal of Applied Statistics, 37(12):2089–2100, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lombardo:2010:SMC

[2087] Rosaria Lombardo and Eric J. Beh.

Simple and multiple correspondence analysis for ordinal-scale variables using orthogonal polynomials. Journal of Applied Statistics, 37(12):2101–2116, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Verbeke:2010:BRS

[2088] Tobias Verbeke. Book reviews: Statistical design and Design of comparative experiments. Journal of Applied Statistics, 37(12):2117–2118, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4696, 4695].

Yunus:2010:SUS

[2089] Faisel Yunus. Statistics Using SPSS: An Integrative Approach, second edition. Journal of Applied Statistics, 37 (12):2119–2120, December 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4709].

Brooms:2010:BRD

[2090] A. C. Brooms. Book review: Data Manipulation with R. Journal of Applied Statistics, 37(12):2121, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4707].

Anonymous:2010:CBA

[2091] Anonymous.  Corrigendum:  A

Bayesian adjustment for covariate misclassification with correlated binary outcome data, D. Ren and R. A. Stone. Journal of Applied Statistics, 37(12): 2123–2124, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [1696].

Anonymous:2010:EB

[2092] Anonymous. Editorial board. Journal of Applied Statistics, 37(12):ebi–??, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aykroyd:2011:E

[2093] Robert G. Aykroyd. Editorial. Journal of Applied Statistics, 38(1):1–??, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gosoniu:2011:NSP

[2094] Laura Gosoniu and Penelope Vounatsou.

Non-stationary partition modeling of geostatistical data for malaria risk mapping. Journal of Applied Statistics, 38(1):3–13, January 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiong:2011:UAM

[2095] Chengjie Xiong, Gerald van Belle, Kejun Zhu, J. Philip Miller, and John C. Morris. A unified approach of metaanalysis: application to an antecedent biomarker study in Alzheimer’s disease. Journal of Applied Statistics, 38(1):15–27, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Awadhi:2011:TDB

[2096] Fahimah Al-Awadhi, Merrilee Hurn, and Christopher Jennison. Threedimensional Bayesian image analysis and confocal microscopy. Journal of Applied Statistics, 38(1):29–46, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demirtas:2011:GMC

[2097] Hakan Demirtas and Donald Hedeker.

Generating multivariate continuous data via the notion of nearest neighbors. Journal of Applied Statistics, 38(1):47–55, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cancho:2011:FMS

[2098] Vicente G. Cancho, Josemar Rodrigues, and Mário de Castro. A flexible model for survival data with a cure rate: a Bayesian approach. Journal of Applied Statistics, 38(1):57–70, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Caporale:2011:FII

[2099] Guglielmo Maria Caporale and Luis A. Gil-Alana. Fractional integration and impulse responses: a bivariate application to real output in the USA and four Scandinavian countries. Journal of Applied Statistics, 38(1):71–85, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

AparecidaDeAraujoRodrigues:2011:DSC

[2100] Aurélia Aparecida De Araújo Rodrigues, Eugenio Kahn Epprecht, and Maysa Sacramento De Magalhães. Double-sampling control charts for attributes. Journal of Applied Statistics, 38(1):87–112, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chien:2011:RDP

[2101] Li-Chu Chien. A robust diagnostic plot for explanatory variables under model mis-specification. Journal of Applied Statistics, 38(1):113–126, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balka:2011:BLI

[2102] Jeremy Balka, Anthony F. Desmond, and Paul D. McNicholas. Bayesian and likelihood inference for cure rates based on defective inverse Gaussian regression models. Journal of Applied Statistics, 38(1):127–144, January 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Allen:2011:UVC

[2103] E. J. Allen and V. T. Farewell. The use of variance components for the assessment of outcome measures in rheumatology. Journal of Applied Statistics, 38(1):145–159, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deLeon:2011:CNP

[2104] Guillermo de León, Pere Grima, and Xavier Tort-Martorell. Comparison of normal probability plots and dot plots in judging the significance of effects in two level factorial designs. Journal of Applied Statistics, 38(1):161–174, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2011:STA

[2105] Bin Li, R. S. Sanderlin, Rebecca A.

Melanson, and Qingzhao Yu. Spatiotemporal analysis of a plant disease in a non-uniform crop: a Monte Carlo approach. Journal of Applied Statistics, 38(1):175–182, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andrade:2011:PRS

[2106] J. A. A. Andrade and J. P. Gosling.

Predicting rainy seasons: quantifying the beliefs of prophets. Journal of Applied Statistics, 38(1):183–193, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baharumshah:2011:IUE

[2107] Ahmad Zubaidi Baharumshah, Nor Aishah Hamzah, and Shamsul Rijal Muhammad Sabri. Inflation uncertainty and economic growth: evidence from the LAD ARCH model. Journal of Applied Statistics, 38(1):195–206, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mielke:2011:RRO

[2108] Paul W. Mielke, Jr., Kenneth J. Berry, and Janis E. Johnston. Robustness without rank order statistics. Journal of Applied Statistics, 38 (1):207–214, January 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vangeneugden:2011:MCE

[2109] Tony Vangeneugden, Geert Molenberghs, Geert Verbeke, and Clarice G. B. Demétrio. Marginal correlation from an extended random-effects model for repeated and overdispersed counts. Journal of Applied Statistics, 38(2):215–232, February 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Costa:2011:CCB

[2110] A. F. B. Costa and M. A. G. Machado. A control chart based on sample ranges for monitoring the covariance matrix of the multivariate processes. Journal of Applied Statistics, 38(2):233–245, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2011:ACL

[2111] Kuo-Chin Lin. Assessing cumulative logit models via a score test in random effect models. Journal of Applied Statistics, 38(2):247–259, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Crato:2011:SLN

[2112] N. Crato, R. R. Linhares, and S. R. C.

Lopes. α-stable laws for noncoding regions in DNA sequences. Journal of Applied Statistics, 38(2):261–271, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Banerjee:2011:EHD

[2113] T. Banerjee, G. Grover, T. Pensi, and D. Banerjee. Estimation of hazard of death in vertically transmitted HIV-1-infected children for doubly censored failure times and fixed covariates. Journal of Applied Statistics, 38 (2):273–285, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalatzis:2011:MFC

[2114] Aquiles E. G. Kalatzis, Camila F. Bassetto, and Carlos R. Azzoni. Multicollinearity and financial constraint in investment decisions: a Bayesian generalized ridge regression. Journal of Applied Statistics, 38(2):287–299, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiaosi:2011:CEB

[2115] Xu Xiaosi, Chen Ying, and Zheng Haitao. The comparison of enterprise bankruptcy forecasting method. Journal of Applied Statistics, 38(2): 301–308, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lim:2011:ODA

[2116] Heonsang Lim and Bong-Jin Yum. Optimal design of accelerated degradation tests based on Wiener process models. Journal of Applied Statistics, 38 (2):309–325, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balestrassi:2011:MDM

[2117] P. P. Balestrassi, A. P. Paiva, A. C. Zambroni de Souza, J. B. Turrioni, and Elmira Popova. A multivariate descriptor method for changepoint detection in nonlinear time series. Journal of Applied Statistics, 38 (2):327–342, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zeller:2011:LIA

[2118] C. B. Zeller, V. H. Lachos, and F. E. Vilca-Labra. Local influence analysis for regression models with scale mixtures of skew-normal distributions. Journal of Applied Statistics, 38(2): 343–368, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Triantafyllopoulos:2011:TVV

[2119] K. Triantafyllopoulos. Time-varying vector autoregressive models with stochastic volatility. Journal of Applied Statistics, 38(2):369–382, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weiss:2011:DMI

[2120] Christian H. Weiß. Detecting mean increases in Poisson INAR(1) processes with EWMA control charts. Journal of Applied Statistics, 38(2):383–398, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mukherjee:2011:SSS

[2121] Amitava Mukherjee and Barendra Purkait. Simultaneous semi-sequential testing of dual alternatives for pattern recognition. Journal of Applied Statistics, 38(2):399–419, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2011:OIR

[2122] Jun Yang, Min Xie, and Thong Ngee Goh. Outlier identification and robust parameter estimation in a zeroinflated Poisson model. Journal of Applied Statistics, 38(2):421–430, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Robinson:2011:BR

[2123] Andrew Robinson. Book review. Journal of Applied Statistics, 38(2):431–??, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brooms:2011:BR

[2124] A. C. Brooms. Book review. Journal of Applied Statistics, 38(2):433–434, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2011:BR

[2125] Kassim S. Mwitondi. Book review.

Journal of Applied Statistics, 38(2): 435–??, February 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Krzanowski:2011:TDB

[2126] Wojtek J. Krzanowski and David J. Hand. Testing the difference between two Kolmogorov–Smirnov values in the context of receiver operating characteristic curves. Journal of Applied Statistics, 38(3):437–450, March 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2011:WBM

[2127] Qingzhao Yu. Weighted bagging: a modification of AdaBoost from the perspective of importance sampling. Journal of Applied Statistics, 38(3):451– 463, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Golosnoy:2011:ISE

[2128] Vasyl Golosnoy and Roman Liesenfeld. Interval shrinkage estimators. Journal of Applied Statistics, 38(3):465–477, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2011:CCI

[2129] Shande Chen. A class of confidence intervals for sequential phase II clinical trials with binary outcome. Journal of Applied Statistics, 38(3):479–489, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2011:BSP

[2130] Himadri Ghosh, M. A. Iquebal, and Prajneshu. Bootstrap study of parameter estimates for nonlinear Richards growth model through genetic algorithm. Journal of Applied Statistics, 38(3):491–500, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fidalgo:2011:DIP

[2131] J. López Fidalgo, I. M. Ortiz Rodríguez, and Weng Kee Wong. Design issues for population growth models. Journal of Applied Statistics, 38 (3):501–512, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Riquelme:2011:IDC

[2132] Marco Riquelme, Víctor Leiva, Manuel Galea, and Antonio Sanhueza. Influence diagnostics on the coefficient of variation of elliptically contoured distributions. Journal of Applied Statistics, 38(3):513–532, March 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hossain:2011:CFM

[2133] Altaf Hossain and Mohammed Nasser.

Comparison of the finite mixture of ARMA-GARCH, back propagation neural networks and support-vector machines in forecasting financial returns. Journal of Applied Statistics, 38(3):533–551, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2011:REC

[2134] Yu-Chang Lin and Chao-Yu Chou. Robustness of the EWMA and the combined X-EWMA control charts with variable sampling intervals to nonnormality. Journal of Applied Statistics, 38(3):553–570, March 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Abbas:2011:EBT

[2135] Nasir Abbas and Muhammad Aslam.

Extending the Bradley–Terry model for paired comparisons to accommodate weights. Journal of Applied Statistics, 38(3):571–580, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pakyari:2011:NMA

[2136] Reza Pakyari. Nonparametric mixture analysis of rock crab of the genus Leptograpsus. Journal of Applied Statistics, 38(3):581–589, March 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2011:BPA

[2137] Jiajia Zhang and Andrew B. Lawson. Bayesian parametric accelerated failure time spatial model and its application to prostate cancer. Journal of Applied Statistics, 38(3):591–603, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2011:RAU

[2138] Jianwen Xu and Hu Yang. On the restricted almost unbiased estimators in linear regression. Journal of Applied Statistics, 38(3):605–617, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Giolo:2011:FMA

[2139] Suely Ruiz Giolo and Clarice Garcia Borges Demétrio. A frailty modeling approach for parental effects in animal breeding. Journal of Applied Statistics, 38(3):619–629, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mishra:2011:NAR

[2140] R. K. Mishra and Zillur Rahman. Nonparametric approach to rank global petroleum business opportunities. Journal of Applied Statistics, 38 (3):631–646, March 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saridakis:2011:VCI

[2141] George Saridakis. Violent crime and incentives in the long-run: evidence from England and Wales. Journal of Applied Statistics, 38(4):647–660, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Costa:2011:EME

[2142] Antonio F. B. Costa and Philippe Castagliola. Effect of measurement error and autocorrelation on the X^{−} chart. Journal of Applied Statistics, 38(4):661–673, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2011:SAT

[2143] Pao-Sheng Shen. Semiparametric analysis of transformation models with doubly censored data. Journal of Applied Statistics, 38(4):675–682, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Abd-Elfattah:2011:TSR

[2144] Ehab F. Abd-Elfattah and Ronald W.

Butler. Tests for symmetry with right censoring. Journal of Applied Statistics, 38(4):683–693, April 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bel:2011:STF

[2145] Liliane Bel, Avner Bar-Hen, Rémy Petit, and Rachid Cheddadi. Spatiotemporal functional regression on paleoecological data. Journal of Applied Statistics, 38(4):695–704, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Terrera:2011:RCP

[2146] G. Muniz Terrera, A. van den Hout, and F. E. Matthews. Random change point models: investigating cognitive decline in the presence of missing data. Journal of Applied Statistics, 38(4):705–716, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hsu:2011:TPM

[2147] Bi-Min Hsu and Ming-Hung Shu. A two-phase method for controlling Erlang-failure processes with high reliability. Journal of Applied Statistics, 38(4):717–734, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Holden:2011:THT

[2148] Darryl Holden. Testing for heteroskedasticity in the tobit and probit models. Journal of Applied Statistics, 38(4):735–744, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Congdon:2011:SEM

[2149] Peter Congdon. Structural equation models for area health outcomes with model selection. Journal of Applied Statistics, 38(4):745–767, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vexler:2011:NLI

[2150] Albert Vexler, Shuling Liu, and Enrique F. Schisterman. Nonparametriclikelihood inference based on costeffectively-sampled-data. Journal of Applied Statistics, 38(4):769–783, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2011:CBD

[2151] Jimin Lee and Seunggeun Hyun. Confidence bands for the difference of two survival functions under the additive risk model. Journal of Applied Statistics, 38(4):785–797, April 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silva:2011:TDR

[2152] E. Silva, V. M. Guerrero, and D. Peña.

Temporal disaggregation and restricted forecasting of multiple population time series. Journal of Applied Statistics, 38(4):799–815, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghahroodi:2011:BTM

[2153] Z. Rezaei Ghahroodi, M. Ganjali, F. Harandi, and D. Berridge. Bivariate transition model for analysing ordinal and nominal categorical responses: an application to the Labour Force Survey data. Journal of Applied Statistics, 38(4):817–832, April 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schorgendorfer:2011:CAC

[2154] A. Schörgendorfer, L. V. Madden, and A. C. Bathke. Choosing appropriate covariance matrices in a nonparametric analysis of factorials in block designs. Journal of Applied Statistics, 38(4):833–850, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deBernardini:2011:FBS

[2155] Diego F. de Bernardini and Laura L. R. Rifo. Full Bayesian significance test for extremal distributions. Journal of Applied Statistics, 38(4):851–863, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ristic:2011:BRS

[2156] Miroslav M. Ristić. Book review: Statistics: A Very Short Introduction. Journal of Applied Statistics, 38(4):865–??, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4705].

Rosenblad:2011:BRC

[2157] Andreas Rosenblad. Book review: The Concise Encyclopedia of Statistics. Journal of Applied Statistics, 38(4):867–868, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4704].

Ober:2011:BRA

[2158] Pieter Bastiaan Ober. Book review: Asymptotic Theory of Statistics and Probability. Journal of Applied Statistics, 38(4):869–??, April 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4701].

Lemonte:2011:IDB

[2159] Artur J. Lemonte and Alexandre G. Patriota. Influence diagnostics in Birnbaum–Saunders nonlinear regression models. Journal of Applied Statistics, 38(5):871–884, May 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samawi:2011:NTS

[2160] Hani M. Samawi, Amal Helu, and Robert Vogel. A nonparametric test of symmetry based on the overlapping coefficient. Journal of Applied Statistics, 38(5):885–898, May 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Esteban:2011:SSA

[2161] Luis Mariano Esteban, Gerardo Sanz, and Angel Borque. A step-by-step algorithm for combining diagnostic tests. Journal of Applied Statistics, 38(5):899–911, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Farall:2011:MMD

[2162] Andrés Farall, Ricardo Maronna, and Tomás Tetzlaff. A mixture model for the detection of Neosporosis without a gold standard. Journal of Applied Statistics, 38(5):913–926, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jin:2011:TVC

[2163] Baisuo Jin, Mong-Na Lo Huang, and Baiqi Miao. Testing for variance changes in autoregressive models with unknown order. Journal of Applied Statistics, 38(5):927–936, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

VanSanden:2011:ELD

[2164] Suzy Van Sanden and Tomasz Burzykowski. Evaluation of Laplace distribution-based ANOVA models applied to microarray data. Journal of Applied Statistics, 38(5):937–950, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demir:2011:CRA

[2165] Eren Demir and Thierry Chaussalet. Capturing the re-admission process: focus on time window. Journal of Applied Statistics, 38(5):951–960, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Curtis:2011:VSA

[2166] S. McKay Curtis and Sujit K. Ghosh. A variable selection approach to monotonic regression with Bernstein polynomials. Journal of Applied Statistics, 38(5):961–976, May 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Taraldsen:2011:ARE

[2167] Gunnar Taraldsen. Analysis of rounded exponential data. Journal of Applied Statistics, 38(5):977–986, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yuan:2011:BFH

[2168] Ao Yuan, Guanjie Chen, Juan Xiong, Wenqing He, Wen Jin, and Charles Rotimi. Bayesian-frequentist hybrid model with application to the analysis of gene copy number changes. Journal of Applied Statistics, 38(5):987–1005, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2011:MMC

[2169] Guoqing Wu, Chao Chen, and Xuefeng Yan. Modified minimum covariance determinant estimator and its application to outlier detection of chemical process data. Journal of Applied Statistics, 38(5):1007–1020, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deLeon:2011:CDC

[2170] A. R. de Leon, A. Soo, and T. Williamson. Classification with discrete and continuous variables via general mixed-data models. Journal of Applied Statistics, 38(5):1021–1032, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cammarota:2011:DSR

[2171] Camillo Cammarota. The differencesign runs length distribution in testing for serial independence. Journal of Applied Statistics, 38(5):1033–1043, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nassar:2011:BSS

[2172] M. M. Nassar, S. M. Khamis, and S. S.

Radwan. On Bayesian sample size determination. Journal of Applied Statistics, 38(5):1045–1054, May 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yusuf:2011:IPV

[2173] Bidemi Yusuf, Olayinka Omigbodun, Babatunde Adedokun, and Odunayo Akinyemi. Identifying predictors of violent behaviour among students using the conventional logistic and multilevel logistic models. Journal of Applied Statistics, 38(5):1055–1061, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nobre:2011:LAL

[2174] Juvêncio S. Nobre and Julio M. Singer.

Leverage analysis for linear mixed models. Journal of Applied Statistics, 38 (5):1063–1072, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pewsey:2011:EMD

[2175] Arthur Pewsey, Kunio Shimizu, and Rolando de la Cruz. On an extension of the von Mises distribution due to Batschelet. Journal of Applied Statistics, 38(5):1073–1085, May 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berhe:2011:ODK

[2176] Leakemariam Berhe and Göran Arnoldsson. D_{s}-optimal designs for Kozak’s tree taper model. Journal of Applied Statistics, 38(5):1087–1102, May 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samani:2011:BLV

[2177] E. Bahrami Samani and M. Ganjali. Bayesian latent variable model for mixed continuous and ordinal responses with possibility of missing responses. Journal of Applied Statistics, 38(6):1103–1116, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez-Camblor:2011:TEA

[2178] Pablo Martínez-Camblor. Testing the equality among distribution functions from independent and right censored samples via Cramér–von Mises criterion. Journal of Applied Statistics, 38 (6):1117–1131, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ng:2011:TSO

[2179] Chi Tim Ng, Johan Lim, and Kyu S.

Hahn. Testing stochastic orders in tails of contingency tables. Journal of Applied Statistics, 38(6):1133–1149, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Benth:2011:STM

[2180] Jurate Saltyte Benth and Laura Saltyte. Spatial-temporal model for wind speed in Lithuania. Journal of Applied Statistics, 38(6):1151–1168, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Albing:2011:ESR

[2181] Malin Albing and Kerstin Vännman.

Elliptical safety region plots for C_{pk}. Journal of Applied Statistics, 38(6): 1169–1187, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Villa:2011:MAS

[2182] Guillermo Villa, Isabel Molina, and Roland Fried. Modeling attendance at Spanish professional football league. Journal of Applied Statistics, 38(6): 1189–1206, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Poleto:2011:CDT

[2183] Frederico Z. Poleto, Julio M. Singer, and Carlos Daniel Paulino. Comparing diagnostic tests with missing data. Journal of Applied Statistics, 38(6):1207–1222, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Magnussen:2011:MBP

[2184] Steen Magnussen and Ron McRoberts. A modified bootstrap procedure for cluster sampling variance estimation of species richness. Journal of Applied Statistics, 38(6):1223–1238, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Louzada-Neto:2011:PED

[2185] Francisco Louzada-Neto, Vicente G. Cancho, and Gladys D. C. Barriga. The Poisson-exponential distribution: a Bayesian approach. Journal of Applied Statistics, 38(6):1239–1248, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Crawford:2011:TQI

[2186] Sara B. Crawford and John J. Hanfelt. Testing for qualitative interaction of multiple sources of informative dropout in longitudinal data. Journal of Applied Statistics, 38(6):1249–1264, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2011:ELG

[2187] Zhensheng Huang. Empirical likelihood for generalized partially linear varyingcoefficient models. Journal of Applied Statistics, 38(6):1265–1275, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yavuz:2011:CEM

[2188] Arzu Altin Yavuz and Bırdal Şenoğlu. Comparison of estimation methods for the finite population mean in simple random sampling: symmetric superpopulations. Journal of Applied Statistics, 38(6):1277–1288, June 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shahriari:2011:TSR

[2189] Hamid Shahriari, Orod Ahmadi, and Amir H. Shokouhi. A two-step robust estimation of the process mean using M-estimator. Journal of Applied Statistics, 38(6):1289–1301, June 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez:2011:BME

[2190] Edson Zangiacomi Martinez, Davi Casale Aragon, and Jorge Alberto Achcar. A Bayesian model for estimating the malaria transition probabilities considering individuals lost to followup. Journal of Applied Statistics, 38 (6):1303–1309, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pemberton:2011:BRT

[2191] John Pemberton. Book review: Time Series Analysis with Applications in R, Second edition. Journal of Applied Statistics, 38(6):1311–1312, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Klapper:2011:BRI

[2192] Jennifer H. Klapper. Book review: Introductory Statistics with R, second edition. Journal of Applied Statistics, 38 (6):1312–1313, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4700].

Ristic:2011:BRG

[2193] Miroslav M. Ristić. Book review:

Guide to Teaching Statistics. Journal of Applied Statistics, 38(6):1313–1314, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4714].

Gazi:2011:BRC

[2194] Boran Gazi. Book review: Credit Risk Management. Journal of Applied Statistics, 38(6):1314–??, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4691].

Ober:2011:BRM

[2195] Pieter Bastiaan Ober. Book review: Modern Regression Methods. Journal of Applied Statistics, 38(6):1315–??, June 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4690, 4718].

Martinez-Camblor:2011:PNS

[2196] Pablo Martínez-Camblor, Carlos Carleos, and Norberto Corral. Powerful nonparametric statistics to compare k independent ROC curves. Journal of Applied Statistics, 38(7):1317–1332, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vexler:2011:LTP

[2197] Albert Vexler, Jihnhee Yu, and Alan D. Hutson. Likelihood testing populations modeled by autoregressive process subject to the limit of detection in applications to longitudinal biomedical data. Journal of Applied Statistics, 38(7):1333–1346, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gerard:2011:TCD

[2198] Patrick D. Gerard and Julia L. Sharp. Testing for co-directional interactions using union-intersection and intersection-union methods. Journal of Applied Statistics, 38(7):1347–1358, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2011:MES

[2199] Rand R. Wilcox and Tian S. Tian.

Measuring effect size: a robust heteroscedastic approach for two or more groups. Journal of Applied Statistics, 38(7):1359–1368, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harri:2011:NTT

[2200] Ardian Harri and Keith H. Coble. Normality testing: two new tests using Lmoments. Journal of Applied Statistics, 38(7):1369–1379, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Takahashi:2011:EAV

[2201] Katsuyuki Takahashi and Isao Shoji.

An empirical analysis of the volatility of the Japanese stock price index: a non-parametric approach. Journal of Applied Statistics, 38(7):1381–1394, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Perdona:2011:GHM

[2202] Gleici Castro Perdoná and Francisco Louzada-Neto. A general hazard model for lifetime data in the presence of cure rate. Journal of Applied Statistics, 38(7):1395–1405, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2011:LRD

[2203] Cheolwoo Park, Félix Hernández-Campos, Long Le, J. S. Marron, Juhyun Park, Vladas Pipiras, F. D. Smith, Richard L. Smith, Michele Trovero, and Zhengyuan Zhu. Long-range dependence analysis of Internet traffic. Journal of Applied Statistics, 38 (7):1407–1433, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silva:2011:RLB

[2204] Giovana O. Silva, Edwin M. M. Ortega, and Gilberto A. Paula. Residuals for log-Burr XII regression models in survival analysis. Journal of Applied Statistics, 38(7):1435–1445, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2011:TKR

[2205] Yalian Li and Hu Yang. Two kinds of restricted modified estimators in linear regression model. Journal of Applied Statistics, 38(7):1447–1454, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Proietti:2011:MTD

[2206] Tommaso Proietti. Multivariate temporal disaggregation with crosssectional constraints. Journal of Applied Statistics, 38(7):1455–1466, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2011:CIP

[2207] Shih-Chia Liu, Kuo-Szu Chiang, Cheng-Hsiang Lin, and Ting-Chin Deng. Confidence interval procedures for proportions estimated by group testing with groups of unequal size adjusted for overdispersion. Journal of Applied Statistics, 38(7):1467–1482, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stogiannis:2011:CFH

[2208] D. Stogiannis, C. Caroni, C. E. Anagnostopoulos, and I. K. Toumpoulis. Comparing first hitting time and proportional hazards regression models. Journal of Applied Statistics, 38(7): 1483–1492, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Su:2011:CDP

[2209] Zheng Su. A class of designs for Phase I cancer clinical trials combining Bayesian and likelihood approaches. Journal of Applied Statistics, 38(7):1493–1498, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Curto:2011:CVC

[2210] José Dias Curto and José Castro Pinto.

The corrected VIF (CVIF). Journal of Applied Statistics, 38(7):1499–1507, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2011:THC

[2211] Jin-Guan Lin, Li-Xing Zhu, Chun-Zheng Cao, and Yong Li. Tests of heteroscedasticity and correlation in multivariate t regression models with AR and ARMA errors. Journal of Applied Statistics, 38(7):1509–1531, July 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Creemers:2011:RAS

[2212] An Creemers, Marc Aerts, Niel Hens, Ziv Shkedy, Frank De Smet, and Philippe Beutels. Revealing agespecific past and future unrelated costs of pneumococcal infections by flexible generalized estimating equations. Journal of Applied Statistics, 38(8): 1533–1547, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Verma:2011:TLS

[2213] Vijay Verma and Gianni Betti. Taylor linearization sampling errors and design effects for poverty measures and other complex statistics. Journal of Applied Statistics, 38(8):1549–1576, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2011:LBA

[2214] Keunbaik Lee, Sanggil Kang, Xuefeng Liu, and Daekwan Seo. Likelihoodbased approach for analysis of longitudinal nominal data using marginalized random effects models. Journal of Applied Statistics, 38(8):1577–1590, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2011:DLF

[2215] Ming-Yuan Leon Li and Shang-En Shine Yu. Do large firms overly use stock-based incentive compensation? Journal of Applied Statistics, 38 (8):1591–1606, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chien:2011:DPB

[2216] Li-Chu Chien. Diagnostic plots in beta-regression models. Journal of Applied Statistics, 38(8):1607–1622, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2011:SCP

[2217] Liansheng Tang, Ming Tan, and Xiao-Hua Zhou. A sequential conditional probability ratio test procedure for comparing diagnostic tests. Journal of Applied Statistics, 38(8):1623–1632, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Santana:2011:IAS

[2218] Lucia Santana, Filidor Vilca, and Víctor Leiva. Influence analysis in skew-Birnbaum–Saunders regression models and applications. Journal of Applied Statistics, 38(8):1633–1649, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rahimzadeh:2011:NMC

[2219] Mitra Rahimzadeh, Ebrahim Hajizadeh, and Farzad Eskandari. Nonmixture cure correlated frailty models in Bayesian approach. Journal of Applied Statistics, 38(8):1651–1663, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reed:2011:FPS

[2220] William J. Reed. A flexible parametric survival model which allows a bathtubshaped hazard rate function. Journal of Applied Statistics, 38(8):1665–1680, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kantar:2011:MCS

[2221] Yeliz Mert Kantar, Ilhan Usta, and Sükrü Acitas. A Monte Carlo simulation study on partially adaptive estimators of linear regression models. Journal of Applied Statistics, 38(8): 1681–1699, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DalBello:2011:OPP

[2222] L. H. A. Dal Bello and A. F. C. Vieira. Optimization of a product performance using mixture experiments including process variables. Journal of Applied Statistics, 38(8):1701–1715, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vieira:2011:DGL

[2223] Afrânio M. C. Vieira, Roseli A. Leandro, Clarice G. B. Demétrio, and Geert Molenberghs. Double generalized linear model for tissue culture proportion data: a Bayesian perspective. Journal of Applied Statistics, 38(8):1717–1731, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2011:JMA

[2224] Chun-Shu Chen and Hong-Ding Yang. A joint modeling approach for spatial earthquake risk variations. Journal of Applied Statistics, 38(8):1733–1741, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nielsen:2011:BRS

[2225] Søren Feodor Nielsen. Book review: SAS for data analysis. Journal of Applied Statistics, 38(8):1743–1744, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4706].

Kang:2011:BRT

[2226] Long Kang. Book review: Time-series data analysis using EViews. Journal of Applied Statistics, 38(8):1744–1745, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4710].

Hamidieh:2011:BRS

[2227] Kam Hamidieh. Book review: Synthetic CDOs modelling, valuation and risk management. Journal of Applied Statistics, 38(8):1745–1746, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4716].

Bakouch:2011:BRP

[2228] Hassan S. Bakouch. Book review:

Probability, Markov chains, queues, and simulation. Journal of Applied Statistics, 38(8):1746–??, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4719].

Wei:2011:SWS

[2229] Yinghui Wei and Peter Neal. Statement of withdrawal: Statistical analysis of an endemic disease from a capturerecapture experiment. Journal of Applied Statistics, 38(8):1747–??, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). The Editor and Publishers Taylor & Francis are withdrawing the above article from publication in Journal of Applied Statistics at the request of the authors. This paper was previously published in online only form as part of Taylor & Francis’ iFirst service. Date of online publication 12 April 2011, DOI 10.1080/02664763.2011.567249.

Anonymous:2011:GKP

[2230] Anonymous. Gopal Kanji Prize 2010.

Journal of Applied Statistics, 38(9): 1749–??, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

SenGupta:2011:NTH

[2231] Ashis SenGupta and Hon Keung Tony Ng. Nonparametric test for the homogeneity of the overall variability. Journal of Applied Statistics, 38(9): 1751–1768, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Torabi:2011:STM

[2232] Mahmoud Torabi and Rhonda J. Rosychuk. Spatio-temporal modelling using B-spline for disease mapping: analysis of childhood cancer trends. Journal of Applied Statistics, 38(9):1769–1781, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Collignon:2011:DAP

[2233] O. Collignon, J.-M. Monnez, P. Vallois, F. Codreanu, J.-M. Renaudin, G. Kanny, M. Brulliard, B. E. Bihain, S. Jacquenet, and D. Moneret-Vautrin. Discriminant analyses of peanut allergy severity scores. Journal of Applied Statistics, 38(9):1783–1799, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kazianka:2011:BAE

[2234] Hannes Kazianka, Michael Mulyk, and Jürgen Pilz. A Bayesian approach to estimating linear mixtures with unknown covariance structure. Journal of Applied Statistics, 38(9):1801–1817, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Atem:2011:FBM

[2235] Folefac D. Atem, Ravi K. Sharma, and Stewart J. Anderson. Fitting bivariate multilevel models to assess longterm changes in body mass index and cigarette smoking. Journal of Applied Statistics, 38(9):1819–1831, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2011:AOC

[2236] Q. Li, G. Zheng, and R. Tiwari.

Analysis of ordered categorical data with score averaging: with applications to case-control genetic associations. Journal of Applied Statistics, 38 (9):1833–1843, September 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Black:2011:MDT

[2237] Anne C. Black, Ofer Harel, and D. Betsy McCoach. Missing data techniques for multilevel data: implications of model misspecification. Journal of Applied Statistics, 38(9): 1845–1865, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2011:SMB

[2238] John J. Chen, Guangxiang Zhang, Chen Ji, and George F. Steinhardt. Simple moment-based inferences of generalized concordance correlation. Journal of Applied Statistics, 38(9): 1867–1882, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bartimote-Aufflick:2011:AOT

[2239] Kathryn Bartimote-Aufflick and Peter C. Thomson. The analysis of ordinal time-series data via a transition (Markov) model. Journal of Applied Statistics, 38(9):1883–1897, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2011:BAI

[2240] Jie Chen, Ayten Yiğiter, and Kuang-Chao Chang. A Bayesian approach to inference about a change point model with application to DNA copy number experimental data. Journal of Applied Statistics, 38(9):1899–1913, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garthwaite:2011:IBP

[2241] Paul H. Garthwaite and John R. Crawford. Inference for a binomial proportion in the presence of ties. Journal of Applied Statistics, 38(9):1915–1934, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Majumdar:2011:NSS

[2242] Anandamayee Majumdar, Corinna Gries, and Jason Walker. A nonstationary spatial generalized linear mixed model approach for studying plant diversity. Journal of Applied Statistics, 38(9):1935–1950, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bordin:2011:HTA

[2243] Tatiana B. Bordin, Hildete P. Pinheiro, and Aluísio Pinheiro. Homogeneity tests among groups for microsatellite data. Journal of Applied Statistics, 38 (9):1951–1962, September 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2011:BHA

[2244] Younan Chen and Keying Ye. A Bayesian hierarchical approach to dual response surface modelling. Journal of Applied Statistics, 38(9):1963–1975, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Nasser:2011:ASN

[2245] Amjad D. Al-Nasser, Mohammad Y.

Al-Rawwash, and Anas S. Alakhras. An approach to setting up a national customer satisfaction index: the Jordan case study. Journal of Applied Statistics, 38(9):1977–1993, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rivaz:2011:STM

[2246] Firoozeh Rivaz, Mohsen Mohammadzadeh, and Majid Jafari Khaledi. Spatio-temporal modeling and prediction of CO concentrations in Tehran City. Journal of Applied Statistics, 38 (9):1995–2007, September 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hajian-Tilaki:2011:EPR

[2247] Karimollah Hajian-Tilaki, James A.

Hanley, and Vahid Nassiri. An extension of parametric ROC analysis for calculating diagnostic accuracy when underlying distributions are mixture of Gaussian. Journal of Applied Statistics, 38(9):2009–2022, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stockhammar:2011:PDE

[2248] Pär Stockhammar and Lars-Erik Öller. On the probability distribution of economic growth. Journal of Applied Statistics, 38(9):2023–2041, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deniz:2011:MSU

[2249] Eylem Deniz, Oguz Akbilgic, and J. Andrew Howe. Model selection using information criteria under a new estimation method: least squares ratio. Journal of Applied Statistics, 38 (9):2043–2050, September 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fallah:2011:ANN

[2250] Nader Fallah, Arnold Mitnitski, and Kenneth Rockwood. Applying neural network Poisson regression to predict cognitive score changes. Journal of Applied Statistics, 38(9):2051–2062, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Murthy:2011:SIS

[2251] Bhamidipati Narasimha Murthy, Ngianga-Bakwin Kandala, Radhakrishnan Ezhil, Prabhdeep Kaur, and Ramachandra Sudha. Statistical issues in studying the relative importance of body mass index, waist circumference, waist hip ratio and waist stature ratio to predict type 2 diabetes. Journal of Applied Statistics, 38(9):2063–2070, September 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Human:2011:REC

[2252] S. W. Human, P. Kritzinger, and S. Chakraborti. Robustness of the EWMA control chart for individual observations. Journal of Applied Statistics, 38(10):2071–2087, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2011:AAW

[2253] Yun Zhao, Andy H. Lee, Kelvin K. W.

Yau, and Geoffrey J. McLachlan. Assessing the adequacy of Weibull survival models: a simulated envelope approach. Journal of Applied Statistics, 38(10):2089–2097, October 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gutierrez:2011:MBD

[2254] Luis Gutiérrez, Fernando A. Quintana, Dietrich von Baer, and Claudia Mardones. Multivariate Bayesian discrimination for varietal authentication of Chilean red wine. Journal of Applied Statistics, 38(10):2099–2109, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Smith:2011:WTI

[2255] Thomas J. Smith and Cornelius M.

McKenna. A weighted test of internal symmetry. Journal of Applied Statistics, 38(10):2111–2118, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lombardo:2011:SDB

[2256] Rosaria Lombardo, Eric J. Beh, and Antonello D’Ambra. Studying the dependence between ordinal-nominal categorical variables via orthogonal polynomials. Journal of Applied Statistics, 38(10):2119–2132, October 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barbato:2011:FPS

[2257] G. Barbato, E. M. Barini, G. Genta, and R. Levi. Features and performance of some outlier detection methods. Journal of Applied Statistics, 38 (10):2133–2149, October 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Albarran:2011:NMD

[2258] I. Albarrán, P. J. Alonso, and J. M. Marin. Nonlinear models of disability and age applied to census data. Journal of Applied Statistics, 38(10): 2151–2163, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sparks:2011:OEW

[2259] R. S. Sparks, T. Keighley, and D. Muscatello. Optimal exponentially weighted moving average (EWMA) plans for detecting seasonal epidemics when faced with non-homogeneous negative binomial counts. Journal of Applied Statistics, 38(10):2165–2181, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patterson:2011:MSS

[2260] Kerry Patterson, Hossein Hassani, Saeed Heravi, and Anatoly Zhigljavsky. Multivariate singular spectrum analysis for forecasting revisions to real-time data. Journal of Applied Statistics, 38 (10):2183–2211, October 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

dosSantos:2011:BAB

[2261] Carlos Aparecido dos Santos and Jorge Alberto Achcar. A Bayesian analysis for the Block and Basu bivariate exponential distribution in the presence of covariates and censored data. Journal of Applied Statistics, 38(10): 2213–2223, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Belloc:2011:HIC

[2262] F. Belloc, A. Maruotti, and L. Petrella. How individual characteristics affect university students drop-out: a semiparametric mixed-effects model for an Italian case study. Journal of Applied Statistics, 38(10):2225–2239, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2011:SRS

[2263] Beibei Guo, Yuehua Wu, Hong Xie, and Baiqi Miao. A segmented regimeswitching model with its application to stock market indices. Journal of Applied Statistics, 38(10):2241–2252, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Walters:2011:ADC

[2264] Stephen J. Walters, C. Jane Morrell, and Pauline Slade. Analysing data from a cluster randomized trial (cRCT) in primary care: a case study. Journal of Applied Statistics, 38(10): 2253–2269, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shi:2011:CRM

[2265] Peng Shi and Wei Zhang. A copula regression model for estimating firm efficiency in the insurance industry. Journal of Applied Statistics, 38(10): 2271–2287, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lai:2011:PTA

[2266] Yinglei Lai, Baolin Wu, and Hongyu Zhao. A permutation test approach to the choice of size k for the nearest neighbors classifier. Journal of Applied Statistics, 38(10):2289–2302, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Snoussi:2011:SSR

[2267] A. Snoussi. SPC for short-run multivariate autocorrelated processes. Journal of Applied Statistics, 38(10):2303– 2312, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hubbard:2011:CNH

[2268] R. A. Hubbard and X. H. Zhou. A comparison of non-homogeneous Markov regression models with application to Alzheimer’s disease progression. Journal of Applied Statistics, 38(10):2313– 2326, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coondoo:2011:DLP

[2269] Dipankor Coondoo, Amita Majumder, and Somnath Chattopadhyay. Districtlevel poverty estimation: a proposed method. Journal of Applied Statistics, 38(10):2327–2343, October 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2011:ELR

[2270] Pao-Sheng Shen. Empirical likelihood ratio with doubly truncated data. Journal of Applied Statistics, 38(10): 2345–2353, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilding:2011:LMB

[2271] Gregory E. Wilding, Xueya Cai, Alan Hutson, and Zhangsheng Yu. A linear model-based test for the heterogeneity of conditional correlations. Journal of Applied Statistics, 38(10): 2355–2366, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2011:BCR

[2272] Kassim S. Mwitondi. Bayesian computation with R. Journal of Applied Statistics, 38(10):2367–2368, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4711].

Ober:2011:BRB

[2273] Pieter Bastiaan Ober. Book review:

Basic statistics. Journal of Applied Statistics, 38(10):2367, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kostenko:2011:BRP

[2274] Andrey Kostenko. Book review: Picturing the uncertain world. Journal of Applied Statistics, 38(10):2368–2369, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4720].

Shang:2011:BRD

[2275] Han Lin Shang. Book review: Dynamic linear models with R. Journal of Applied Statistics, 38(10):2369–2370, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4717].

Nielsen:2011:BRI

[2276] Søren Feodor Nielsen. Book review: Introductory time series with R. Journal of Applied Statistics, 38(10):2370–2371, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4713].

Kang:2011:BRV

[2277] Long Kang. Book review: Volatility and time series econometrics: essays in honor of Robert Engle. Journal of Applied Statistics, 38(10):2371–2372, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4721].

Zhang:2011:MLB

[2278] H. Zhang, Y. Xia, R. Chen, D. Gunzler, W. Tang, and Xin Tu. Modeling longitudinal binomial responses: implications from two dueling paradigms. Journal of Applied Statistics, 38(11): 2373–2390, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gijbels:2011:SEM

[2279] I. Gijbels and I. Prosdocimi. Smooth estimation of mean and dispersion function in extended generalized additive models with application to Italian induced abortion data. Journal of Applied Statistics, 38(11):2391–2411, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chandra:2011:DLE

[2280] Hukum Chandra, Nicola Salvati, and U. C. Sud. Disaggregate-level estimates of indebtedness in the state of Uttar Pradesh in India: an application of small-area estimation technique. Journal of Applied Statistics, 38(11): 2413–2432, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ehlers:2011:CBM

[2281] Ricardo S. Ehlers. Comparison of Bayesian models for production efficiency. Journal of Applied Statistics, 38 (11):2433–2443, November 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lama:2011:PLR

[2282] Nicola Lama, Patrizia Boracchi, and Elia Biganzoli. Partial logistic relevance vector machines in survival analysis. Journal of Applied Statistics, 38 (11):2445–2458, November 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Seif:2011:MES

[2283] Asghar Seif, Alireza Faraz, Cédric Heuchenne, Erwin Saniga, and M. B. Moghadam. A modified economicstatistical design of the T^{2} control chart with variable sample sizes and control limits. Journal of Applied Statistics, 38 (11):2459–2469, November 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeIaco:2011:NST

[2284] Sandra De Iaco. A new space–time multivariate approach for environmental data analysis. Journal of Applied Statistics, 38(11):2471–2483, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gazioglu:2011:SAL

[2285] Suzan Gazioğlu and E. Marian Scott. Sensitivity analysis of linear timeinvariant compartmental models with steady-state constraint. Journal of Applied Statistics, 38(11):2485–2509, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Longford:2011:MMI

[2286] N. T. Longford and Pierpaolo D’Urso. Mixture models with an improper component. Journal of Applied Statistics, 38(11):2511–2521, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goodman:2011:DMC

[2287] Melody S. Goodman, Yi Li, and Ram C. Tiwari. Detecting multiple change points in piecewise constant hazard functions. Journal of Applied Statistics, 38(11):2523–2532, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2011:SLT

[2288] Himadri Ghosh and Prajneshu. Statistical learning theory for fitting multimodal distribution to rainfall data: an application. Journal of Applied Statistics, 38(11):2533–2545, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patel:2011:GTL

[2289] Rakhee Dinubhai Patel and Frederic Paik Schoenberg. A graphical test for local self-similarity in univariate data. Journal of Applied Statistics, 38 (11):2547–2562, November 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2011:HJI

[2290] Chia-Lin Chang, Michael McAleer, and Les Oxley. How are journal impact, prestige and article influence related? An application to neuroscience. Journal of Applied Statistics, 38(11): 2563–2573, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2011:SFE

[2291] Chi-Shuan Liu and Fang-Chih Tien. A single-featured EWMA-X control chart for detecting shifts in process mean and standard deviation. Journal of Applied Statistics, 38(11):2575–2596, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akay:2011:AMT

[2292] Kadri Ulas Akay and Müjgan Tez.

Alternative modeling techniques for the quantal response data in mixture experiments. Journal of Applied Statistics, 38(11):2597–2616, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hubbard:2011:WMV

[2293] Raymond Hubbard. The widespread misinterpretation of p-values as error probabilities. Journal of Applied Statistics, 38(11):2617–2626, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Farcomeni:2011:JAO

[2294] Alessio Farcomeni, Alessandra Nardi, and Elena Fabrizi. Joint analysis of occurrence and time to stability after entrance into the Italian labour market: an approach based on a Bayesian cure model with structured stochastic search variable selection. Journal of Applied Statistics, 38(11):2627–2646, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jeng:2011:ATD

[2295] Shuen-Lin Jeng and Yu-Te Liu. Adaptive tangent distance classifier on recognition of handwritten digits. Journal of Applied Statistics, 38(11):2647–2659, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fernandez-Aguirre:2011:KEL

[2296] K. Fernández-Aguirre, M. I. Landaluce-Calvo, A. Martín-Arroyuelos, and J. I. Modroño-Herrán. Knowledge extraction from a large on-line survey: a case study for a higher education corporate marketing. Journal of Applied Statistics, 38(11):2661–2679, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kenett:2011:PDS

[2297] Ron S. Kenett. On the planning and design of sample surveys. Journal of Applied Statistics, 38(11):2681–??, November 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Faddy:2011:ACD

[2298] M. J. Faddy and D. M. Smith. Analysis of count data with covariate dependence in both mean and variance. Journal of Applied Statistics, 38(12): 2683–2694, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gupta:2011:WIG

[2299] Ramesh C. Gupta and Debasis Kundu.

Weighted inverse Gaussian — a versatile lifetime model. Journal of Applied Statistics, 38(12):2695–2708, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hung:2011:CBM

[2300] Wen-Liang Hung and Yen-Chang Chang. Comparison between method of moments and entropy regularization algorithm applied to parameter estimation for mixed-Weibull distribution. Journal of Applied Statistics, 38(12): 2709–2722, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cybis:2011:PLR

[2301] G. B. Cybis, S. R. C. Lopes, and H. P.

Pinheiro. Power of the likelihood ratio test for models of DNA base substitution. Journal of Applied Statistics, 38 (12):2723–2737, December 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ma:2011:CED

[2302] Chang-Xing Ma, Albert Vexler, Enrique F. Schisterman, and Lili Tian.

Cost-efficient designs based on linearly associated biomarkers. Journal of Applied Statistics, 38(12):2739–2750, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zagar:2011:LSI

[2303] Tina Zagar, Vesna Zadnik, and Maja Primic Zakelj. Local standardized incidence ratio estimates and comparison with other mapping methods for small geographical areas using Slovenian breast cancer data. Journal of Applied Statistics, 38(12):2751–2761, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Afroughi:2011:ACS

[2304] Solaiman Afroughi, Soghrat Faghihzadeh, Majid Jafari Khaledi, Mehdi Ghandehari Motlagh, and Ebrahim Hajizadeh. Analysis of clustered spatially correlated binary data using autologistic model and Bayesian method with an application to dental caries of 3– 5-year-old children. Journal of Applied Statistics, 38(12):2763–2774, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bacci:2011:MIM

[2305] Silvia Bacci and Valeria Caviezel. Multilevel IRT models for the university teaching evaluation. Journal of Applied Statistics, 38(12):2775–2791, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scotto:2011:EVC

[2306] Manuel G. Scotto, Susana M. Barbosa, and Andrés M. Alonso. Extreme value and cluster analysis of European daily temperature series. Journal of Applied Statistics, 38(12):2793–2804, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bai:2011:BCI

[2307] Peng Bai, Wen Gan, and Lei Shi.

Bayesian confidence interval for the risk ratio in a correlated 2 × 2 table with structural zero. Journal of Applied Statistics, 38(12):2805–2817, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Leiva:2011:MWE

[2308] Víctor Leiva, Emilia Athayde, Cecilia Azevedo, and Carolina Marchant. Modeling wind energy flux by a Birnbaum–Saunders distribution with an unknown shift parameter. Journal of Applied Statistics, 38(12):2819–2838, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rakitzis:2011:IOS

[2309] Athanasios C. Rakitzis and Demetrios L.

Antzoulakos. On the improvement of one-sided S control charts. Journal of Applied Statistics, 38(12):2839–2858, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sankaran:2011:SFM

[2310] P. G. Sankaran and P. Anisha. Shared frailty model for recurrent event data with multiple causes. Journal of Applied Statistics, 38(12):2859–2868, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2011:ANR

[2311] Jin Zhang. Adaptive normal reference bandwidth based on quantile for kernel density estimation. Journal of Applied Statistics, 38(12):2869–2880, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huber:2011:TCB

[2312] Martin Huber. Testing for covariate balance using quantile regression and resampling methods. Journal of Applied Statistics, 38(12):2881–2899, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2011:RLI

[2313] Chien-Hung Chen and Tsung-Shan Tsou. Robust likelihood inferences for multivariate correlated data. Journal of Applied Statistics, 38(12):2901–2910, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jia:2011:AAB

[2314] Nan Jia and Thomas M. Braun. The adaptive accelerated biased coin design for phase I clinical trials. Journal of Applied Statistics, 38(12):2911–2924, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2011:CTR

[2315] Xinyu Tang and Abdus S. Wahed. Comparison of treatment regimes with adjustment for auxiliary variables. Journal of Applied Statistics, 38(12): 2925–2938, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haupt:2011:CVF

[2316] Harry Haupt, Kathrin Kagerer, and Joachim Schnurbus. Cross-validating fit and predictive accuracy of nonlinear quantile regressions. Journal of Applied Statistics, 38(12):2939–2954, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Langrock:2011:SAN

[2317] Roland Langrock. Some applications of nonlinear and non-Gaussian statespace modelling by means of hidden Markov models. Journal of Applied Statistics, 38(12):2955–2970, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Domma:2011:MLE

[2318] Filippo Domma, Sabrina Giordano, and Mariangela Zenga. Maximum likelihood estimation in Dagum distribution with censored samples. Journal of Applied Statistics, 38(12):2971–2985, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karagrigoriou:2011:FMS

[2319] Alex Karagrigoriou. Frailty models in survival analysis. Journal of Applied Statistics, 38(12):2988–2989, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ober:2011:RPF

[2320] Pieter Bastiaan Ober. Random phenomena: Fundamentals of probability and statistics for engineers. Journal of Applied Statistics, 38(12):2989–2990, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Castagliola:2011:ITS

[2321] Philippe Castagliola. Introduction to time-series modelling. Journal of Applied Statistics, 38(12):2993–2994, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2011:IES

[2322] Kassim S. Mwitondi. Interpreting economic and social data. Journal of Applied Statistics, 38(12):2994–2995, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Laberge:2011:ARM

[2323] Yves Laberge. Advising on research methods: a consultant’s companion. Journal of Applied Statistics, 38(12): 2991–??, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shang:2011:BN

[2324] Han Lin Shang. Bayesian nonparametrics. Journal of Applied Statistics, 38 (12):2990–??, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2011:DAU

[2325] Kassim S. Mwitondi. Data analysis using SAS ENTERPRISE GUIDE. Journal of Applied Statistics, 38(12): 2993–??, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dialsingh:2011:MTP

[2326] Isaac Dialsingh. Multiple testing problems in pharmaceutical statistics. Journal of Applied Statistics, 38(12): 2987–??, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shang:2011:NPE

[2327] Han Lin Shang. Non-parametric econometrics. Journal of Applied Statistics, 38(12):2992–??, December 2011. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2011:EB

[2328] Anonymous. Editorial board. Journal of Applied Statistics, 38(12):ebi, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brito:2012:MID

[2329] Paula Brito and A. Pedro Duarte Silva.

Modelling interval data with normal and skew-normal distributions. Journal of Applied Statistics, 39(1):3–20, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2012:NCL

[2330] Wanbo Lu and Daimin Shi. A new compounding life distribution: the Weibull–Poisson distribution. Journal of Applied Statistics, 39(1):21–38, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zahran:2012:MPS

[2331] Sammy Zahran, Michael A. Long, and Kenneth J. Berry. Measures of predictor sensitivity for order-insensitive partitioning of multiple correlation. Journal of Applied Statistics, 39(1):39–51, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Diana:2012:CBA

[2332] Giancarlo Diana and Pier Francesco Perri. A calibration-based approach to sensitive data: a simulation study. Journal of Applied Statistics, 39(1):53– 65, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demidenko:2012:CIB

[2333] Eugene Demidenko. Confidence intervals and bands for the binormal ROC curve revisited. Journal of Applied Statistics, 39(1):67–79, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Masala:2012:EOE

[2334] Giovanni Masala. Earthquakes occurrences estimation through a parametric semi-Markov approach. Journal of Applied Statistics, 39(1):81–96, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2012:DRE

[2335] Rabindra Nath Das and Jeong-Soo Park. Discrepancy in regression estimates between log-normal and gamma: some case studies. Journal of Applied Statistics, 39(1):97–111, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Minard:2012:TSC

[2336] Charles G. Minard, Wenyaw Chan, David W. Wetter, and Carol J. Etzel. Trends in smoking cessation: a Markov model approach. Journal of Applied Statistics, 39(1):113–127, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Angelini:2012:CTC

[2337] Claudia Angelini, Daniela De Canditiis, and Marianna Pensky. Clustering time-course microarray data using functional Bayesian infinite mixture model. Journal of Applied Statistics, 39(1):129–149, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2012:BCR

[2338] Guoyi Zhang and Yan Lu. Biascorrected random forests in regression. Journal of Applied Statistics, 39 (1):151–160, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zahid:2012:ORR

[2339] Faisal M. Zahid and Shahla Ramzan.

Ordinal ridge regression with categorical predictors. Journal of Applied Statistics, 39(1):161–171, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsai:2012:AII

[2340] Miao-Yu Tsai. Assessing inter- and intra-agreement for dependent binary data: a Bayesian hierarchical correlation approach. Journal of Applied Statistics, 39(1):173–187, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jacques:2012:CUP

[2341] R. M. Jacques, N. R. J. Fieller, and E. K. Ainscow. A classification updating procedure motivated by highcontent screening data. Journal of Applied Statistics, 39(1):189–198, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Congdon:2012:SRE

[2342] P. Congdon and C. D. Lloyd. A spatial random-effects model for interzone flows: commuting in Northern Ireland. Journal of Applied Statistics, 39 (1):199–213, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schuller:2012:MHM

[2343] Jan C. Schuller. The malicious host: a minimax solution of the Monty Hall problem. Journal of Applied Statistics, 39(1):215–221, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aykroyd:2012:E

[2344] Robert G. Aykroyd. Editorial. Journal of Applied Statistics, 39(1):1–??, January 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schioler:2012:MOD

[2345] Linus Schiöler and Marianne Frisén. Multivariate outbreak detection. Journal of Applied Statistics, 39(2):223– 242, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Drechsler:2012:NDD

[2346] Jörg Drechsler. New data dissemination approaches in old Europe — synthetic datasets for a German establishment survey. Journal of Applied Statistics, 39(2):243–265, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2012:JBV

[2347] Yu-Mei Chang, Chun-Shu Chen, and Pao-Sheng Shen. A jackknife-based versatile test for two-sample problems with right-censored data. Journal of Applied Statistics, 39(2):267–277, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Iki:2012:MDM

[2348] Kiyotaka Iki, Kouji Tahata, and Sadao Tomizawa. Measure of departure from marginal homogeneity using marginal odds for multi-way tables with ordered categories. Journal of Applied Statistics, 39(2):279–295, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McIntyre:2012:BDS

[2349] Julie McIntyre and Ronald P. Barry.

Bivariate deconvolution with SIMEX: an application to mapping Alaska earthquake density. Journal of Applied Statistics, 39(2):297–308, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lio:2012:EBX

[2350] Y. L. Lio and Tzong-Ru Tsai. Estimation of δ = P (X < Y) for Burr XII distribution based on the progressively first failure-censored samples. Journal of Applied Statistics, 39 (2):309–322, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maravelakis:2012:MEE

[2351] Petros E. Maravelakis. Measurement error effect on the CUSUM control chart. Journal of Applied Statistics, 39 (2):323–336, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sorensen:2012:QSF

[2352] Helle Sørensen, Anders Tolver, Maj Halling Thomsen, and Pia Haubro Andersen. Quantification of symmetry for functional data with application to equine lameness classification. Journal of Applied Statistics, 39(2):337–360, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mills:2012:SPM

[2353] Terence C. Mills. Semi-parametric modelling of temperature records. Journal of Applied Statistics, 39(2): 361–383, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2012:SMG

[2354] Caiya Zhang, Yanbiao Xiang, and Xinmei Shen. Some multivariate goodnessof-fit tests based on data depth. Journal of Applied Statistics, 39(2):385– 397, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Crager:2012:GSH

[2355] Michael R. Crager. Generalizing the standardized hazard ratio to multivariate proportional hazards regression, with an application to clinical genomic studies. Journal of Applied Statistics, 39(2):399–417, February 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Perry:2012:ITC

[2356] Marcus B. Perry and Joseph J. Pignatiello, Jr. Identifying the time of change in the mean of a twostage nested process. Journal of Applied Statistics, 39(2):419–433, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dixon:2012:AMF

[2357] Stephanie N. Dixon, Gerarda A. Darlington, and Victoria Edge. Applying a marginalized frailty model to competing risks. Journal of Applied Statistics, 39(2):435–443, February 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chervoneva:2012:TSH

[2358] Inna Chervoneva, Tingting Zhan, Boris Iglewicz, Walter W. Hauck, and David E. Birk. Two-stage hierarchical modeling for analysis of subpopulations in conditional distributions. Journal of Applied Statistics, 39(2): 445–460, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2012:AR

[2359] Anonymous. Acknowledgement of referees, 2010–2011. Journal of Applied Statistics, 39(2):461–465, February 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hasan:2012:RAZ

[2360] M. Tariqul Hasan, Gary Sneddon, and Renjun Ma. Regression analysis of zero-inflated time-series counts: application to air pollution related emergency room visit data. Journal of Applied Statistics, 39(3):467– 476, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Caudill:2012:FAP

[2361] Steven B. Caudill, James E. Long, and Franklin G. Mixon Jr. Female athletic participation and income: evidence from a latent class model. Journal of Applied Statistics, 39(3):477– 488, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Song:2012:EFD

[2362] Jie Song, Herman W. Raadsma, and Peter C. Thomson. Evaluation of false discovery rate and power via sample size in microarray studies. Journal of Applied Statistics, 39(3):489–500, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Barbiero:2012:IER

[2363] Alessandro Barbiero. Interval estimators for reliability: the bivariate normal case. Journal of Applied Statistics, 39(3):501–512, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2012:TSA

[2364] Joseph Kang, Xiaogang Su, Brian Hitsman, Kiang Liu, and Donald Lloyd-Jones. Tree-structured analysis of treatment effects with large observational data. Journal of Applied Statistics, 39(3):513–529, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lachos:2012:NIS

[2365] Victor H. Lachos, Celso R. B. Cabral, and Carlos A. Abanto-Valle. A noniterative sampling Bayesian method for linear mixed models with normal independent distributions. Journal of Applied Statistics, 39(3):531– 549, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Nunes:2012:TRP

[2366] Célia Nunes, Dário Ferreira, Sandra S.

Ferreira, and João T. Mexia. F - tests with a rare pathology. Journal of Applied Statistics, 39(3):551– 561, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Drikvandi:2012:TVC

[2367] Reza Drikvandi, Ahmad Khodadadi, and Geert Verbeke. Testing variance components in balanced linear growth curve models. Journal of Applied Statistics, 39(3):563–572, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Miller:2012:QFS

[2368] Richard W. Miller. Quad folding: a simple idea for the subjective property characterization of large sample sets. Journal of Applied Statistics, 39(3): 573–580, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Weiss:2012:FOI

[2369] Christian H. Weiß. Fully observed INAR(1) processes. Journal of Applied Statistics, 39(3):581–598, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2012:AAD

[2370] Hsiu-Wen Chen, Weng Kee Wong, and Hongquan Xu. An augmented approach to the desirability function. Journal of Applied Statistics, 39(3):

599–613, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Uribe-Opazo:2012:IDG

[2371] Miguel Angel Uribe-Opazo, Joelmir André Borssoi, and Manuel Galea. Influence diagnostics in Gaussian spatial linear models. Journal of Applied Statistics, 39(3):615–630, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Jaffa:2012:SEC

[2372] Miran A. Jaffa, Ayad A. Jaffa, and Stuart R. Lipsitz. Slope estimation of covariates that influence renal outcome following renal transplant adjusting for informative right censoring. Journal of Applied Statistics, 39(3): 631–642, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2012:PTS

[2373] Hsiuying Wang and Shan-Lin Hung. Phylogenetic tree selection by the adjusted k-means approach. Journal of Applied Statistics, 39(3):643–655, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Holgersson:2012:EMS

[2374] H. E. T. Holgersson, Peter S. Karlsson, and Rashid Mansoor. Estimating mean-standard deviation ratios of financial data. Journal of Applied Statistics, 39(3):657–671, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kuroki:2012:OCP

[2375] Manabu Kuroki. Optimizing a control plan using a structural equation model with an application to statistical process analysis. Journal of Applied Statistics, 39(3):673–694, 2012. ISSN 0266-4763 (print), 1360-0532 (electronic).

deWinter:2012:FRP

[2376] J. C. F. de Winter and D. Dodou. Factor recovery by principal axis factoring and maximum likelihood factor analysis as a function of factor pattern and sample size. Journal of Applied Statistics, 39(4):695–710, April 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rocchetti:2012:MDD

[2377] Irene Rocchetti, Domenica Taruscio, and Marco Alfò. Modeling delay in diagnosis of NF: under reporting, incidence and prevalence estimates. Journal of Applied Statistics, 39(4):711– 721, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tahata:2012:QAM

[2378] Kouji Tahata. Quasi-asymmetry model for square tables with nominal categories. Journal of Applied Statistics, 39(4):723–729, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lombardo:2012:SAM

[2379] Rosaria Lombardo, Pietro Amenta, Myrtille Vivien, and Robert Sabatier. Sensory analysis via multi-block multivariate additive PLS splines. Journal of Applied Statistics, 39(4):731–743, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Omari:2012:IQC

[2380] Amer Ibrahim Al-Omari and Abdul Haq. Improved quality control charts for monitoring the process mean, using double-ranked set sampling methods. Journal of Applied Statistics, 39(4):745–763, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saha:2012:PLB

[2381] Krishna K. Saha, Debaraj Sen, and Chun Jin. Profile likelihood-based confidence interval for the dispersion parameter in count data. Journal of Applied Statistics, 39(4):765–783, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2012:LOP

[2382] Yao Yu and Jun Wang. Latticeoriented percolation system applied to volatility behavior of stock market. Journal of Applied Statistics, 39(4):785–797, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alhamzawi:2012:VSQ

[2383] Rahim Alhamzawi and Keming Yu. Variable selection in quantile regression via Gibbs sampling. Journal of Applied Statistics, 39(4):799–813, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barnum:2012:EDC

[2384] Darold T. Barnum, John M. Gleason, Matthew G. Karlaftis, Glen T. Schumock, Karen L. Shields, Sonali Tandon, and Surrey M. Walton. Estimating DEA confidence intervals with statistical panel data analysis. Journal of Applied Statistics, 39(4):815–828, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2012:EBV

[2385] Chun-Xia Zhang, Guan-Wei Wang, and Jiang-She Zhang. An empirical biasvariance analysis of DECORATE ensemble method at different training sample sizes. Journal of Applied Statistics, 39(4):829–850, April 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lombard:2012:MRT

[2386] F. Lombard and C. J. Potgieter. A multivariate rank test for comparing mass size distributions. Journal of Applied Statistics, 39(4):851–865, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Berchialla:2012:UEF

[2387] P. Berchialla, S. Snidero, A. Stancu, C. Scarinzi, R. Corradetti, and D. Gregori. Understanding the epidemiology of foreign body injuries in children using a data-driven Bayesian network. Journal of Applied Statistics, 39(4):867–874, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sharmin:2012:STM

[2388] Sifat Sharmin and Md. Israt Rayhan.

Spatio-temporal modeling of infectious disease dynamics. Journal of Applied Statistics, 39(4):875–882, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Neustifter:2012:MPP

[2389] Benjamin Neustifter, Stephen L. Rathbun, and Saul Shiffman. Mixed-Poisson point process with partially observed covariates: ecological momentary assessment of smoking. Journal of Applied Statistics, 39(4):883–899, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Othus:2012:CPC

[2390] Megan Othus, Yi Li, and Ram Tiwari. Change-point cure models with application to estimating the changepoint effect of age of diagnosis among prostate cancer patients. Journal of Applied Statistics, 39(4):901–911, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2012:AMS

[2391] Yun-Huan Lee and Chun-Shu Chen. Autoregressive model selection based on a prediction perspective. Journal of Applied Statistics, 39(4):913–922, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nielsen:2012:IGG

[2392] Søren Feodor Nielsen. Introduction to general and generalized linear models. Journal of Applied Statistics, 39(4):923–924, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2012:BRO

[2393] Long Kang. Book review: The Oxford Handbook of Credit Derivatives. Journal of Applied Statistics, 39(4):924– 925, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4725].

Espejo:2012:BRE

[2394] Mariano Ruiz Espejo. Book review: Exercises and solutions in biostatistical theory. Journal of Applied Statistics, 39(4):925–926, April 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4724].

Bakouch:2012:BRT

[2395] Hassan S. Bakouch. Book review: Time series: Modeling, Computation, and Inference. Journal of Applied Statistics, 39(4):923–??, April 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4722].

Lagona:2012:MBC

[2396] Francesco Lagona and Marco Picone.

Model-based clustering of multivariate skew data with circular components and missing values. Journal of Applied Statistics, 39(5):927–945, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ning:2012:ELR

[2397] Wei Ning. Empirical likelihood ratio test for a mean change point model with a linear trend followed by an abrupt change. Journal of Applied Statistics, 39(5):947–961, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2012:TSL

[2398] Wei Liu and Lang Wu. Two-step and likelihood methods for HIV viral dynamic models with covariate measurement errors and missing data. Journal of Applied Statistics, 39(5):963–978, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mendes:2012:ECC

[2399] Susana Mendes, M. José Fernández-Gómez, Mário Jorge Pereira, Ulisses Miranda Azeiteiro, and M. Purificación Galindo-Villardón. An empirical comparison of Canonical Correspondence Analysis and STATICO in the identification of spatio-temporal ecological relationships. Journal of Applied Statistics, 39(5):979–994, May 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kauermann:2012:MMF

[2400] Göran Kauermann and Nina Westerheide. To move or not to move to find a new job: spatial duration time model with dynamic covariate effects. Journal of Applied Statistics, 39(5):995–1009, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Quintero-Sarmiento:2012:EIM

[2401] Adrián Quintero-Sarmiento, Edilberto Cepeda-Cuervo, and Vicente Núñez-Antón. Estimating infant mortality in Colombia: some overdispersion modelling approaches. Journal of Applied Statistics, 39(5):1011–1036, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soltanian:2012:GGM

[2402] Ali Reza Soltanian and Soghrat Faghihzadeh. A generalization of the Grizzle model to the estimation of treatment effects in crossover trials with non-compliance. Journal of Applied Statistics, 39(5):1037–1048, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Russo:2012:IDH

[2403] Cibele M. Russo, Gilberto A. Paula, Francisco José A. Cysneiros, and Reiko Aoki. Influence diagnostics in heteroscedastic and/or autoregressive nonlinear elliptical models for correlated data. Journal of Applied Statistics, 39(5):1049–1067, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Israel:2012:SMG

[2404] Vinicius P. Israel and Hélio S. Migon.

Stochastic models for greenhouse gas emission rate estimation from hydroelectric reservoirs: a Bayesian hierarchical approach. Journal of Applied Statistics, 39(5):1069–1086, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mansourian:2012:BAL

[2405] Marjan Mansourian, Anoshirvan Kazemnejad, Iraj Kazemi, Farid Zayeri, and Masoud Soheilian. Bayesian analysis of longitudinal ordered data with flexible random effects using McMC: application to diabetic macular edema data. Journal of Applied Statistics, 39(5):1087–1100, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hanchane:2012:SEP

[2406] Saïd Hanchane and Tarek Mostafa.

Solving endogeneity problems in multilevel estimation: an example using education production functions. Journal of Applied Statistics, 39(5):1101–1114, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hron:2012:LRC

[2407] K. Hron, P. Filzmoser, and K. Thompson. Linear regression with compositional explanatory variables. Journal of Applied Statistics, 39(5):1115–1128, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kiruthika:2012:CTF

[2408] C. Kiruthika and R. Chandrasekaran. Classification of textile fabrics using statistical multivariate techniques. Journal of Applied Statistics, 39(5): 1129–1138, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deardon:2012:SME

[2409] Rob Deardon, Babak Habibzadeh, and Hau Yi Chung. Spatial measurement error in infectious disease models. Journal of Applied Statistics, 39 (5):1139–1150, May 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Morrell:2012:SPC

[2410] Christopher H. Morrell, Larry J. Brant, Shan Sheng, and E. Jeffrey Metter. Screening for prostate cancer using multivariate mixed-effects models. Journal of Applied Statistics, 39 (6):1151–1175, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2012:BPC

[2411] Liuxia Wang. Bayesian principal component regression with data-driven component selection. Journal of Applied Statistics, 39(6):1177–1189, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ortega:2012:NBB

[2412] Edwin M. M. Ortega, Gauss M. Cordeiro, and Michael W. Kattan. The negative binomial-beta Weibull regression model to predict the cure of prostate cancer. Journal of Applied Statistics, 39(6):1191–1210, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kulczycki:2012:CGC

[2413] Piotr Kulczycki, Małgorzata Charytanowicz, Piotr A. Kowalski, and Szymon Lukasik. The Complete Gradient Clustering Algorithm: properties in practical applications. Journal of Applied Statistics, 39(6):1211–1224, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

George:2012:CIE

[2414] Florence George and B. M. Golam Kibria. Confidence intervals for estimating the population signal-to-noise ratio: a simulation study. Journal of Applied Statistics, 39(6):1225–1240, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2012:GRM

[2415] Zhiyong Zhang, John J. McArdle, and John R. Nesselroade. Growth rate models: emphasizing growth rate analysis through growth curve modeling. Journal of Applied Statistics, 39 (6):1241–1262, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cutillo:2012:IRP

[2416] Andrea Cutillo and Claudio Ceccarelli.

The internal relocation premium: are migrants positively or negatively selected? Evidence from Italy. Journal of Applied Statistics, 39(6):1263–1278, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Voudouris:2012:MSK

[2417] Vlasios Voudouris, Robert Gilchrist, Robert Rigby, John Sedgwick, and Dimitrios Stasinopoulos. Modelling skewness and kurtosis with the BCPE density in GAMLSS. Journal of Applied Statistics, 39(6):1279–1293, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vrontos:2012:EHF

[2418] Ioannis Vrontos. Evidence for hedge fund predictability from a multivariate Student’s t full-factor GARCH model. Journal of Applied Statistics, 39(6):1295–1321, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lalitha:2012:MOT

[2419] S. Lalitha and Nirpeksh Kumar. Multiple outlier test for upper outliers in an exponential sample. Journal of Applied Statistics, 39(6):1323–1330, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bardet:2012:NPM

[2420] Jean-Marc Bardet, Imen Kammoun, and Veronique Billat. A new process for modeling heartbeat signals during exhaustive run with an adaptive estimator of its fractal parameters. Journal of Applied Statistics, 39(6):1331–1351, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2012:GLE

[2421] Xu-Qing Liu and Bo Li. General linear estimators under the prediction error sum of squares criterion in a linear regression model. Journal of Applied Statistics, 39(6):1353–1361, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hafner:2012:CCW

[2422] Christian M. Hafner. Cross-correlating wavelet coefficients with applications to high-frequency financial time series. Journal of Applied Statistics, 39 (6):1363–1379, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beh:2012:EMA

[2423] Eric J. Beh. Exploratory multivariate analysis by example using R. Journal of Applied Statistics, 39(6):1381–1382, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boshnakov:2012:URD

[2424] Georgi N. Boshnakov. Using R for data management, statistical analysis, and graphics. Journal of Applied Statistics, 39(6):1382–1383, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stoimenova:2012:RNS

[2425] Eugenia Stoimenova. Robust nonparametric statistical methods. Journal of Applied Statistics, 39(6):1383–1384, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stoimenova:2012:NSI

[2426] Eugenia Stoimenova. Nonparametric statistical inference. Journal of Applied Statistics, 39(6):1384–1385, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kuo:2012:VSF

[2427] Shu-Fu Kuo and Yu-Shan Shih. Variable selection for functional density trees. Journal of Applied Statistics, 39(7):1387–1395, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2012:NAI

[2428] Jin-Hong Park. Nonparametric approach to intervention time series modeling. Journal of Applied Statistics, 39(7):1397–1408, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ardoino:2012:FPM

[2429] I. Ardoino, E. M. Biganzoli, C. Bajdik, P. J. Lisboa, P. Boracchi, and F. Ambrogi. Flexible parametric modelling of the hazard function in breast cancer studies. Journal of Applied Statistics, 39(7):1409–1421, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andres:2012:ATT

[2430] A. Martín Andrés, M. Álvarez Hernández, and I. Herranz Tejedor. Asymptotic two-tailed confidence intervals for the difference of proportions. Journal of Applied Statistics, 39(7):1423–1435, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2012:LAC

[2431] Carol Y. Lin, Lance A. Waller, and Robert H. Lyles. The likelihood approach for the comparison of medical diagnostic system with multiple binary tests. Journal of Applied Statistics, 39(7):1437–1454, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Donald:2012:CTD

[2432] Margaret R. Donald, Chris Strickland, Clair L. Alston, Rick Young, and Kerrie L. Mengersen. Comparison of threedimensional profiles over time. Journal of Applied Statistics, 39(7):1455–1474, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2012:GPL

[2433] Jianbo Li, Minggao Gu, and Tao Hu. General partially linear varyingcoefficient transformation models for ranking data. Journal of Applied Statistics, 39(7):1475–1488, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li-Thiao-Te:2012:BMA

[2434] Sébastien Li-Thiao-Té, Daudin Jean-Jacques, and Robin Stéphane. Bayesian model averaging for estimating the number of classes: applications to the total number of species in metagenomics. Journal of Applied Statistics, 39(7):1489–1504, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2012:SUP

[2435] Shuyi Jiang. Survival in the U.S.

petroleum refining industry. Journal of Applied Statistics, 39(7):1505–1530, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jeong:2012:WBV

[2436] Jinook Jeong and Byunguk Kang.

Wild-bootstrapped variance-ratio test for autocorrelation in the presence of heteroskedasticity. Journal of Applied Statistics, 39(7):1531–1542, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Palacios-Gonzalez:2012:ICD

[2437] Federico Palacios-González and Rosa María García-Fernández. Interpretation of the coefficient of determination of an ANOVA model as a measure of polarization. Journal of Applied Statistics, 39(7):1543–1555, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Longford:2012:PIE

[2438] Nicholas T. Longford, Maria Grazia Pittau, Roberto Zelli, and Riccardo Massari. Poverty and inequality in European regions. Journal of Applied Statistics, 39(7):1557–1576, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Correa:2012:EOF

[2439] Alexander A. Correa, Pere Grima, and Xavier Tort-Martorell. Experimentation order in factorial designs: new findings. Journal of Applied Statistics, 39(7):1577–1591, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hutson:2012:MEA

[2440] Alan D. Hutson and Gregory E. Wilding. Maintaining the exchangeability assumption for a two-group permutation test in the non-randomized setting. Journal of Applied Statistics, 39(7):1593–1603, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yoon:2012:GVS

[2441] Young Joo Yoon, Cheolwoo Park, Erik Hofmeister, and Sangwook Kang. Group variable selection in cardiopulmonary cerebral resuscitation data for veterinary patients. Journal of Applied Statistics, 39(7):1605–1621, July 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lwin:2012:MCP

[2442] Thaung Lwin. Modelling of connected processes. Journal of Applied Statistics, 39(8):1623–1641, August 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pavlenko:2012:CSA

[2443] Tatjana Pavlenko, Anders Björkström, and Annika Tillander. Covariance structure approximation via gLasso in high-dimensional supervised classification. Journal of Applied Statistics, 39 (8):1643–1666, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2012:HTR

[2444] Yueqin Zhao and Dayanand N.

Naik. Hypothesis testing with Rao’s quadratic entropy and its application to dinosaur biodiversity. Journal of Applied Statistics, 39(8):1667–1680, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Klimova:2012:CFA

[2445] Anna Klimova and Tamás Rudas.

Coordinate-free analysis of trends in British social mobility. Journal of Applied Statistics, 39(8):1681–1691, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mameli:2012:LSC

[2446] Valentina Mameli, Monica Musio, Erik Sauleau, and Annibale Biggeri. Large sample confidence intervals for the skewness parameter of the skew-normal distribution based on Fisher’s transformation. Journal of Applied Statistics, 39(8):1693–1702, August 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Benyelles:2012:MDE

[2447] Wafaa Benyelles and Tahar Mourid. On a minimum distance estimate of the period in functional autoregressive processes. Journal of Applied Statistics, 39 (8):1703–1718, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhu:2012:IQT

[2448] Lan Zhu, Su Chen, Zhuoxin Jiang, Zhiwu Zhang, Hung-Chih Ku, Xuesong Li, Melinda McCann, Steve Harris, George Lust, Pual Jones, and Rory Todhunter. Identification of quantitative trait loci for canine hip dysplasia by two sequential multipoint linkage analyses. Journal of Applied Statistics, 39(8):1719–1731, August 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2012:BAG

[2449] Huiyun Wu, Qingxia Chen, Lorraine B. Ware, and Tatsuki Koyama. A Bayesian approach for generalized linear models with explanatory biomarker measurement variables subject to detection limit: an application to acute lung injury. Journal of Applied Statistics, 39(8):1733–1747, August 2012.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Teles:2012:CCM

[2450] Júlia Teles. Concordance coefficients to measure the agreement among several sets of ranks. Journal of Applied Statistics, 39(8):1749–1764, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shukur:2012:MRM

[2451] Ghazi Shukur and Zangin Zeebari. Median regression for SUR models with the same explanatory variables in each equation. Journal of Applied Statistics, 39(8):1765–1779, August 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dasgupta:2012:MOC

[2452] Nairanjana Dasgupta and Monte J.

Shaffer. Many-to-one comparison of nonlinear growth curves for Washington’s Red Delicious apple. Journal of Applied Statistics, 39(8):1781–1795, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gonzalez-Manteiga:2012:PLM

[2453] Wenceslao Gonzalez-Manteiga, Guillermo Henry, and Daniela Rodriguez. Partly linear models on Riemannian manifolds. Journal of Applied Statistics, 39 (8):1797–1809, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2012:BIU

[2454] Yu-Jau Lin and Y. L. Lio. Bayesian inference under progressive type-I interval censoring. Journal of Applied Statistics, 39(8):1811–1824, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maroulas:2012:EAS

[2455] Vasileios Maroulas. Error analysis of stochastic flight trajectory prediction models. Journal of Applied Statistics, 39(8):1825–1841, August 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shang:2012:GSD

[2456] Han Lin Shang. Graphics for statistics and data analysis with R. Journal of Applied Statistics, 39(8):1843–1844, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Holloway:2012:DAQ

[2457] Ivana Holloway. Design and analysis of quality of life studies in clinical trials. Journal of Applied Statistics, 39 (8):1844–1845, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gosling:2012:BAM

[2458] John Paul Gosling. Bayesian analysis made simple: An excel GUI for Win-BUGS. Journal of Applied Statistics, 39(8):1845–1846, August 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2012:LMM

[2459] Zhengfeng Guo. Linear model methodology. Journal of Applied Statistics, 39 (8):1846–1847, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nielsen:2012:EIM

[2460] Søren Feodor Nielsen. An elementary introduction to mathematical finance. Journal of Applied Statistics, 39 (8):1847–1848, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Young:2012:OED

[2461] Derek S. Young. Optimal experimental design with R. Journal of Applied Statistics, 39(8):1848–1849, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mosammam:2012:OHN

[2462] A. M. Mosammam. The Oxford Handbook of Nonlinear Filtering. Journal of Applied Statistics, 39(8):1849–1850, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beeley:2012:ABH

[2463] Chris Beeley. Applied Bayesian hierarchical methods. Journal of Applied Statistics, 39(8):1845–??, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Er:2012:MGL

[2464] Sebnem Er. Multivariate generalized linear models using R. Journal of Applied Statistics, 39(8):1851–??, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martella:2012:CMD

[2465] Francesca Martella and Maurizio Vichi.

Clustering microarray data using model-based double K-means. Journal of Applied Statistics, 39(9):1853–1869, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lombard:2012:CPD

[2466] F. Lombard and R. K. Maxwell. A cusum procedure to detect deviations from uniformity in angular data. Journal of Applied Statistics, 39(9): 1871–1880, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2012:SME

[2467] Pao-Sheng Shen. Semiparametric mixed-effects models for clustered doubly censored data. Journal of Applied Statistics, 39(9):1881–1892, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2012:SWD

[2468] Gikuang Jeff Chen. A simple way to deal with multicollinearity. Journal of Applied Statistics, 39(9):1893–1909, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2012:SSD

[2469] Chung-I Li and Jeh-Nan Pan. Sample size determination for estimating multivariate process capability indices based on lower confidence limits. Journal of Applied Statistics, 39(9): 1911–1920, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tovar:2012:APP

[2470] Ainhoa Oguiza Tovar, Inmaculada Gallastegui Zulaica, and Vicente Núñez-Antón. Analysis of pseudo-panel data with dependent samples. Journal of Applied Statistics, 39(9):1921–1937, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coetzer:2012:EMD

[2471] R. L. J. Coetzer, R. F. Rossouw, and N. J. Le Roux. Efficient maximin distance designs for experiments in mixtures. Journal of Applied Statistics, 39 (9):1939–1951, September 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Achcar:2012:WGE

[2472] Jorge Alberto Achcar, Edilberto Cepeda-Cuervo, and Eliane R. Rodrigues. Weibull and generalised exponential overdispersion models with an application to ozone air pollution. Journal of Applied Statistics, 39(9): 1953–1963, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samani:2012:LEL

[2473] E. Bahrami Samani, Y. Amirian, and M. Ganjali. Likelihood estimation for longitudinal zero-inflated power series regression models. Journal of Applied Statistics, 39(9):1965–1974, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Meligkotsidou:2012:BPD

[2474] Loukia Meligkotsidou, Elias Tzavalis, and Ioannis D. Vrontos. A Bayesian panel data framework for examining the economic growth convergence hypothesis: do the G7 countries converge? Journal of Applied Statistics, 39(9):1975–1990, September 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2012:CMD

[2475] Hongxia Yang, Jun Wang, and Alexsandra Mojslovic. Cascade model with Dirichlet process for analyzing multiple dyadic matrices. Journal of Applied Statistics, 39(9):1991–2003, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Colosimo:2012:PMS

[2476] Enrico A. Colosimo, Maria Arlene Fausto, Marta Afonso Freitas, and Jorge Andrade Pinto. Practical modeling strategies for unbalanced longitudinal data analysis. Journal of Applied Statistics, 39(9):2005–2013, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2012:NEM

[2477] Bhim Singh, B. V. S. Sisodia, Anupam Singh, and R. P. Kaushal. A note on the estimation methods of crop production at the block level. Journal of Applied Statistics, 39(9): 2015–2027, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samat:2012:VBI

[2478] N. A. Samat and D. F. Percy. Vectorborne infectious disease mapping with stochastic difference equations: an analysis of dengue disease in Malaysia. Journal of Applied Statistics, 39(9): 2029–2046, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bacci:2012:LDD

[2479] Silvia Bacci. Longitudinal data: different approaches in the context of item-response theory models. Journal of Applied Statistics, 39(9):2047–2065, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2012:NLD

[2480] H. Zhang, Q. Yu, C. Feng, D. Gunzler, P. Wu, and X. M. Tu. A new look at the difference between the GEE and the GLMM when modeling longitudinal count responses. Journal of Applied Statistics, 39(9):2067–2079, September 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ebegil:2012:TSC

[2481] Meral Ebegil and Fikri Gökpinar. A test statistic to choose between liu-type and least-squares estimator based on mean square error criteria. Journal of Applied Statistics, 39(10):2081–2096, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2012:RMD

[2482] Jiun-Yi Wang, Li-Ching Chen, and Hui-Min Lin. Robust methods for detecting familial aggregation of a quantitative trait in matched case-control family studies. Journal of Applied Statistics, 39(10):2097–2111, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McPherson:2012:ETP

[2483] Sterling McPherson and Celestina Barbosa-Leiker. An example of a twopart latent growth curve model for semicontinuous outcomes in the health sciences. Journal of Applied Statistics, 39(10):2113–2128, October 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Calis:2012:NPF

[2484] Nazif Çalis and Hamza Erol. A new per-field classification method using mixture discriminant analysis. Journal of Applied Statistics, 39(10):2129–2140, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kawasaki:2012:BIT

[2485] Youhei Kawasaki and Etsuo Miyaoka. A Bayesian inference of P (λ_{1} < λ_{2}) for two Poisson parameters. Journal of Applied Statistics, 39(10):2141–2152, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2012:OSS

[2486] Jiin-Huarng Guo. Optimal sample size planning for the Wilcoxon–Mann– Whitney and van Elteren tests under cost constraints. Journal of Applied Statistics, 39(10):2153–2164, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DAmbra:2012:VME

[2487] Luigi D’Ambra, Antonello D’Ambra, and Pasquale Sarnacchiaro. Visualizing main effects and interaction in multiple non-symmetric correspondence analysis. Journal of Applied Statistics, 39 (10):2165–2175, October 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

He:2012:MIU

[2488] Yulei He and Trivellore E. Raghunathan. Multiple imputation using multivariate gh transformations. Journal of Applied Statistics, 39(10): 2177–2198, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fotouhi:2012:EVD

[2489] H. Fotouhi and M. Golalizadeh.

Exploring the variability of DNA molecules via principal geodesic analysis on the shape space. Journal of Applied Statistics, 39(10):2199–2207, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Louzada:2012:NLT

[2490] Francisco Louzada, Vicente G. Cancho, Mari Roman, and José G. Leite. A new long-term lifetime distribution induced by a latent complementary risk framework. Journal of Applied Statistics, 39(10):2209–2222, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kasparian:2012:OBL

[2491] Jérôme Kasparian and Antoine Rolland. OECD’s ‘Better Life Index’: can any country be well ranked? Journal of Applied Statistics, 39(10):2223–2230, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gillard:2012:GBC

[2492] Jonathan Gillard. A generalised Box– Cox transformation for the parametric estimation of clinical reference intervals. Journal of Applied Statistics, 39(10):2231–2245, October 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lo:2012:EPL

[2493] Yungtai Lo. Estimating the prevalence of low-lumbar spine bone mineral density in older men with or at risk for HIV infection using normal mixture models. Journal of Applied Statistics, 39 (10):2247–2258, October 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hung:2012:SUC

[2494] Wen-Liang Hung, Shou-Jen Chang-Chien, and Miin-Shen Yang. Selfupdating clustering algorithm for estimating the parameters in mixtures of von Mises distributions. Journal of Applied Statistics, 39(10):2259–2274, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2012:MCC

[2495] Hong Li and Wei Ning. Multiple comparisons with a control under heteroscedasticity. Journal of Applied Statistics, 39(10):2275–2283, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2012:BSP

[2496] Taeyoung Park, Robert T. Krafty, and Alvaro I. Sánchez. Bayesian semiparametric analysis of Poisson changepoint regression models: application to policy-making in Cali, Colombia. Journal of Applied Statistics, 39(10): 2285–2298, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karagrigoriou:2012:SIM

[2497] Alex Karagrigoriou. Statistical inference: The minimum distance approach. Journal of Applied Statistics, 39(10): 2299–2300, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakouch:2012:ATS

[2498] Hassan S. Bakouch. Applied time series analysis. Journal of Applied Statistics, 39(10):2300–2301, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodrigues:2012:RSN

[2499] Eliane R. Rodrigues. Reversibility and stochastic networks. Journal of Applied Statistics, 39(10):2301–2302, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kamarianakis:2012:OHE

[2500] Yiannis Kamarianakis. The Oxford Handbook of Economic Forecasting. Journal of Applied Statistics, 39(10): 2303–2304, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dialsingh:2012:EAS

[2501] Isaac Dialsingh. Expansions and asymptotics for statistics. Journal of Applied Statistics, 39(10):2304–2305, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

delaCruz:2012:BID

[2502] Rolando de la Cruz. Bayesian ideas and data analysis: An introduction for scientists and statisticians. Journal of Applied Statistics, 39(10):2306–2307, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Laberge:2012:OHQ

[2503] Yves Laberge. The Oxford Handbook of Quantitative Asset Management. Journal of Applied Statistics, 39(10):2307–2308, October 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sapatinas:2012:SHD

[2504] Theofanis Sapatinas. Statistics for high-dimensional data. Journal of Applied Statistics, 39(10):2308–2309, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:2012:WSR

[2505] David J. Hand. Who’s #1? The science of rating and ranking. Journal of Applied Statistics, 39(10):2309–2310, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hughes:2012:BCA

[2506] P. William Hughes. Biostatistics: a computing approach. Journal of Applied Statistics, 39(10):2306–??, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dialsingh:2012:LSI

[2507] Isaac Dialsingh. Large-scale inference: empirical Bayes methods for estimation, testing, and prediction. Journal of Applied Statistics, 39(10):2305–??, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Keeble:2012:RP

[2508] Claire Keeble. The R primer. Journal of Applied Statistics, 39(10):2303– ??, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2012:SDM

[2509] Kassim S. Mwitondi. Statistical data mining using SAS applications. Journal of Applied Statistics, 39(10):2302– ??, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Salcedo:2012:WBT

[2510] G. E. Salcedo, R. F. Porto, S. Y. Roa, and F. R. Momo. A wavelet-based time-varying autoregressive model for non-stationary and irregular time series. Journal of Applied Statistics, 39 (11):2313–2325, November 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahabadi:2012:ILS

[2511] S. Eftekhari Mahabadi and M. Ganjali. An index of local sensitivity to non-ignorability for parametric survival models with potential nonrandom missing covariate: an application to the SEER cancer registry data. Journal of Applied Statistics, 39(11): 2327–2348, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pardo:2012:GPL

[2512] M. C. Pardo and R. Alonso. A generalized Q–Q plot for longitudinal data. Journal of Applied Statistics, 39(11): 2349–2362, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

White:2012:DPM

[2513] Nicole White, Helen Johnson, Peter Silburn, and Kerrie Mengersen. Dirichlet process mixture models for unsupervised clustering of symptoms in Parkinson’s disease. Journal of Applied Statistics, 39(11):2363–2377, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Taylor:2012:VPS

[2514] Charles C. Taylor, Kanti V. Mardia, Marco Di Marzio, and Agnese Panzera. Validating protein structure using kernel density estimates. Journal of Applied Statistics, 39(11): 2379–2388, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rastogi:2012:EPB

[2515] Manoj Kumar Rastogi, Yogesh Mani Tripathi, and Shuo-Jye Wu. Estimating the parameters of a bathtubshaped distribution under progressive type-II censoring. Journal of Applied Statistics, 39(11):2389–2411, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Iddi:2012:JMM

[2516] Samuel Iddi and Geert Molenberghs. A joint marginalized multilevel model for longitudinal outcomes. Journal of Applied Statistics, 39(11):2413–2430, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Torabi:2012:SMU

[2517] Mahmoud Torabi. Spatial modeling using frequentist approach for disease mapping. Journal of Applied Statistics, 39(11):2431–2439, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ma:2012:IKW

[2518] Yan Ma. On inference for Kendall’s τ within a longitudinal data setting. Journal of Applied Statistics, 39(11): 2441–2452, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goncalves:2012:SSE

[2519] Luzia Gonçalves, M. Rosário de Oliveira, Cláudia Pascoal, and Ana Pires. Sample size for estimating a binomial proportion: comparison of different methods. Journal of Applied Statistics, 39 (11):2453–2473, November 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mardia:2012:MCM

[2520] Kanti V. Mardia, John T. Kent, Zhengzheng Zhang, Charles C. Taylor, and Thomas Hamelryck. Mixtures of concentrated multivariate sine distributions with applications to bioinformatics. Journal of Applied Statistics, 39 (11):2475–2492, November 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Benigni:2012:STI

[2521] Matthew Benigni and Reinhard Furrer. Spatio-temporal improvised explosive device monitoring: improving detection to minimise attacks. Journal of Applied Statistics, 39(11):2493–2508, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kimber:2012:CBT

[2522] Alan Kimber and Shah-Jalal Sarker. A covariance-based test for shared frailty in multivariate lifetime data. Journal of Applied Statistics, 39(11): 2509–2522, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ster:2012:ISS

[2523] Irina Chis Ster. Inference for serological surveys investigating past exposures to infections resulting in long-lasting immunity — an approach using finite mixture models with concomitant information. Journal of Applied Statistics, 39 (11):2523–2542, November 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2012:GKP

[2524] Anonymous. Gopal Kanji Prize 2011. Journal of Applied Statistics, 39(11): 2311–??, November 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2012:VSJ

[2525] Liu-Cang Wu, Zhong-Zhan Zhang, and Deng-Ke Xu. Variable selection in joint mean and variance models of Box– Cox transformation. Journal of Applied Statistics, 39(12):2543–2555, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2012:AHB

[2526] Chang-Shuai Li. Analysis of hedging based on co-persistence theory. Journal of Applied Statistics, 39(12): 2557–2567, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zintzaras:2012:PGO

[2527] Elias Zintzaras. The power of generalized odds ratio in assessing association in genetic studies with known mode of inheritance. Journal of Applied Statistics, 39(12):2569–2581, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hirose:2012:CFA

[2528] Kei Hirose and Tomoyuki Higuchi. Creating facial animation of characters via MoCap data. Journal of Applied Statistics, 39(12):2583–2597, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hofer:2012:BSR

[2529] Vera Hofer and Johannes Leitner. A bivariate Sarmanov regression model for count data with generalised Poisson marginals. Journal of Applied Statistics, 39(12):2599–2617, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Landram:2012:CUP

[2530] Suzanne V. Landram and Frank G.

Landram. A computational understanding of partial and part determination coefficients. Journal of Applied Statistics, 39(12):2619–2626, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zeebari:2012:DRE

[2531] Zangin Zeebari. Developing ridge estimation method for median regression. Journal of Applied Statistics, 39(12): 2627–2638, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gajewski:2012:DEA

[2532] Byron J. Gajewski, Robert Lee, and Nancy Dunton. Data envelopment analysis in the presence of measurement error: case study from the National Database of Nursing Quality Indicators^{(}R) (NDNQI^{(}R)). Journal of Applied Statistics, 39(12):2639–2653, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2012:CTD

[2533] Rand R. Wilcox and David M. Erceg-Hurn. Comparing two dependent groups via quantiles. Journal of Applied Statistics, 39(12):2655–2664, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Matteucci:2012:UPV

[2534] Mariagiulia Matteucci, Stefania Mignani, and Bernard P. Veldkamp. The use of predicted values for item parameters in item response theory models: an application in intelligence tests. Journal of Applied Statistics, 39(12): 2665–2683, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Venegas:2012:RMU

[2535] Osvaldo Venegas, Francisco Rodríguez, Héctor W. Gómez, Juan F. Olivares-Pacheco, and Heleno Bolfarine. Robust modeling using the generalized epsilonskew-t distribution. Journal of Applied Statistics, 39(12):2685–2698, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mukhopadhyay:2012:SSR

[2536] S. Mukhopadhyay, I. Das, and K. Das.

Selection of a stroke risk model based on transcranial Doppler ultrasound velocity. Journal of Applied Statistics, 39 (12):2699–2712, December 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Holgersson:2012:TEM

[2537] H. E. T. Holgersson and Peter S. Karlsson. Three estimators of the Mahalanobis distance in high-dimensional data. Journal of Applied Statistics, 39 (12):2713–2720, December 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ferreira:2012:DDA

[2538] Julio C. Ferreira, Marta A. Freitas, and Enrico A. Colosimo. Degradation data analysis for samples under unequal operating conditions: a case study on train wheels. Journal of Applied Statistics, 39(12):2721–2739, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soliman:2012:ECV

[2539] Ahmed A. Soliman, A. H. Abd Ellah, N. A. Abou-Elheggag, and A. A. Modhesh. Estimation of the coefficient of variation for non-normal model using progressive first-failure-censoring data. Journal of Applied Statistics, 39(12): 2741–2758, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wei:2012:SAE

[2540] Yinghui Wei, Peter Neal, Sandra Telfer, and Mike Begon. Statistical analysis of an endemic disease from a capture-recapture experiment. Journal of Applied Statistics, 39(12): 2759–2773, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2012:C

[2541] Anonymous. Corrigendum. Journal of Applied Statistics, 39(12):2777–2778, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2012:E

[2542] Anonymous. Erratum. Journal of Applied Statistics, 39(12):2775–??, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2012:EB

[2543] Anonymous. Editorial Board. Journal of Applied Statistics, 39(12):ebi–??, December 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aykroyd:2013:E

[2544] Robert G. Aykroyd. Editorial. Journal of Applied Statistics, 40(1):1, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bowater:2013:HSS

[2545] Russell J. Bowater and Gabriel Escarela. Heterogeneity and study size in random-effects meta-analysis. Journal of Applied Statistics, 40(1):2–16, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Pike:2013:JMC

[2546] Francis Pike and Lisa Weissfeld. Joint modeling of censored longitudinal and event time data. Journal of Applied Statistics, 40(1):17–27, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fernandez-Ponce:2013:BID

[2547] J. M. Fernández-Ponce, F. Palacios-Rodríguez, and M. R. Rodríguez-Griñolo. Bayesian influence diagnostics in radiocarbon dating. Journal of Applied Statistics, 40(1):28–39, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Arroyo:2013:LPD

[2548] Antonio S. M. Arroyo, Antonio García-Ferrer, Aránzazu de Juan Fernández, and Rocío Sánchez-Mangas. Lower posterior death probabilities from a quick medical response in road traffic accidents. Journal of Applied Statistics, 40(1):40–58, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Garthwaite:2013:PDE

[2549] Paul H. Garthwaite, Shafeeqah A. Al-Awadhi, Fadlalla G. Elfadaly, and David J. Jenkinson. Prior distribution elicitation for generalized linear and piecewise-linear models. Journal of Applied Statistics, 40(1):59–75, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez-Camblor:2013:HTP

[2550] Pablo Martínez-Camblor, Norberto Corral, and Jesus María de la Hera.

Hypothesis test for paired samples in the presence of missing data. Journal of Applied Statistics, 40(1):76–87, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Sulis:2013:BAI

[2551] Isabella Sulis and Vincenza Capursi. Building up adjusted indicators of students’ evaluation of university courses using generalized item response models. Journal of Applied Statistics, 40(1): 88–102, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kachiashvili:2013:QOB

[2552] K. J. Kachiashvili, M. A. Hashmi, and A. Mueed. Quasi-optimal Bayesian procedures of many hypotheses testing. Journal of Applied Statistics, 40 (1):103–122, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bao:2013:JMA

[2553] Yanchun Bao, Hongsheng Dai, Tao Wang, and Sung-Kiang Chuang. A joint modelling approach for clustered recurrent events and death events. Journal of Applied Statistics, 40(1): 123–140, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Nissi:2013:SSH

[2554] Eugenia Nissi and Annalina Sarra. A simulation study on the hybrid nature of Tango’s index. Journal of Applied Statistics, 40(1):141–151, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Shang:2013:FTS

[2555] Han Lin Shang. Functional time series approach for forecasting very short-term electricity demand. Journal of Applied Statistics, 40(1):152– 168, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zimmerman:2013:HVB

[2556] Donald W. Zimmerman. Heterogeneity of variance and biased hypothesis tests. Journal of Applied Statistics, 40(1):169–193, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2013:MMM

[2557] S. H. Lin and R. S. Wang. Modified method on the means for several log-normal distributions. Journal of Applied Statistics, 40(1):194– 208, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2013:IRE

[2558] Xu-Qing Liu, Feng Gao, and Zhen-Feng Yu. Improved ridge estimators in a linear regression model. Journal of Applied Statistics, 40(1):209– 220, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Pires:2013:EUF

[2559] Cesaltina Pires, Andreia Dionísio, and Luís Coelho. Estimating utility functions using generalized maximum entropy. Journal of Applied Statistics, 40(1):221–234, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Duan:2013:PCA

[2560] Faping Duan, Daniel Ogden, Ling Xu, Kang Liu, George Lust, Jody Sandler, Nathan L. Dykes, Lan Zhu, Steven Harris, Paul Jones, Rory J. Todhunter, and Zhiwu Zhang. Principal component analysis of canine hip dysplasia phenotypes and their statistical power for genome-wide association mapping. Journal of Applied Statistics, 40(2): 235–251, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fu:2013:PEN

[2561] Liyong Fu, Yuancai Lei, Ram P. Sharma, and Shouzheng Tang. Parameter estimation of nonlinear mixedeffects models using first-order conditional linearization and the EM algorithm. Journal of Applied Statistics, 40(2):252–265, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2013:IPA

[2562] Yuanhua Feng. An iterative plug-in algorithm for decomposing seasonal time series using the Berlin Method. Journal of Applied Statistics, 40(2):266– 281, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Greselin:2013:CGZ

[2563] Francesca Greselin, Leo Pasquazzi, and Ricardas Zitikis. Contrasting the Gini and Zenga indices of economic inequality. Journal of Applied Statistics, 40(2): 282–297, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Baldi:2013:MHC

[2564] Ileana Baldi, Eva Pagano, Paola Berchialla, Alessandro Desideri, Alberto Ferrando, Franco Merletti, and Dario Gregori. Modeling healthcare costs in simultaneous presence of asymmetry, heteroscedasticity and correlation. Journal of Applied Statistics, 40(2):298–310, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wei:2013:CSS

[2565] Lai Wei and Alan D. Hutson. A comment on sample size calculations for binomial confidence intervals. Journal of Applied Statistics, 40(2):311– 319, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Howe:2013:RMM

[2566] J. Andrew Howe and Hamparsum Bozdogan. Robust mixture model cluster analysis using adaptive kernels. Journal of Applied Statistics, 40(2):320– 336, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Quatember:2013:FMS

[2567] Andreas Quatember and Monika Cornelia Hausner. A family of methods for statistical disclosure control. Journal of Applied Statistics, 40(2):337– 346, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cao:2013:END

[2568] Xiting Cao, Baolin Wu, and Marshall I.

Hertz. Empirical null distributionbased modeling of multi-class differential gene expression detection. Journal of Applied Statistics, 40(2):347– 357, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2013:SCA

[2569] Baolin Wu. Sparse cluster analysis of large-scale discrete variables with application to single nucleotide polymorphism data. Journal of Applied Statistics, 40(2):358–367, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2013:JCR

[2570] Dongliang Wang and Alan D. Hutson. Joint confidence region estimation of L-moment ratios with an extension to right censored data. Journal of Applied Statistics, 40(2):368– 379, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2013:IVB

[2571] Peixin Zhao and Liugen Xue. Instrumental variable-based empirical likelihood inferences for varying-coefficient models with error-prone covariates. Journal of Applied Statistics, 40(2): 380–396, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Usta:2013:DEM

[2572] Ilhan Usta. Different estimation methods for the parameters of the extended Burr XII distribution. Journal of Applied Statistics, 40(2):397– 414, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Abascal:2013:EDM

[2573] Elena Abascal, Vidal Díaz de Rada, Ignacio García Lautre, and M. Isabel Landaluce. Extending dual multiple factor analysis to categorical tables. Journal of Applied Statistics, 40(2): 415–428, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Angelopoulos:2013:OAA

[2574] P. Angelopoulos, K. Drosou, and C. Koukouvinos. An orthogonal arrays approach to robust parameter designs methodology. Journal of Applied Statistics, 40(2):429–437, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Golosnoy:2013:SNT

[2575] Vasyl Golosnoy and Jens Hogrefe. Signaling NBER turning points: a sequential approach. Journal of Applied Statistics, 40(2):438–448, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2013:OAC

[2576] Hea-Jung Kim. Optimal asymmetric classification procedures for intervalscreened normal data. Journal of Applied Statistics, 40(2):449–462, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2013:EML

[2577] Shuangzhe Liu. Econometric methods for labour economics. Journal of Applied Statistics, 40(2):463–464, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2013:DMR

[2578] Kassim S. Mwitondi. Data mining with Rattle and R. Journal of Applied Statistics, 40(2):464, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Parkinson:2013:TSM

[2579] Michael Parkinson. z test for the significance of the mean of a stable probability distribution with 1 < α ≤ 2. Journal of Applied Statistics, 40(3): 465–482, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mehmood:2013:CCL

[2580] Rashid Mehmood, Muhammad Riaz, and Ronald J. M. M. Does. Control charts for location based on different sampling schemes. Journal of Applied Statistics, 40(3):483–494, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kulikova:2013:SVM

[2581] M. V. Kulikova and D. R. Taylor. Stochastic volatility models for exchange rates and their estimation using quasi-maximum-likelihood methods: an application to the South African Rand. Journal of Applied Statistics, 40(3):495–507, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ciavolino:2013:UHT

[2582] Enrico Ciavolino and Mariangela Nitti. Using the Hybrid Two-Step estimation approach for the identification of second-order latent variable models. Journal of Applied Statistics, 40(3): 508–526, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

StClair:2013:CMU

[2583] Katherine St. Clair, Eric Dunton, and John Giudice. A comparison of models using removal effort to estimate animal abundance. Journal of Applied Statistics, 40(3):527–545, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Rogers:2013:JMP

[2584] J. K. Rogers and J. L. Hutton. Joint modelling of pre-randomisation event counts and multiple postrandomisation survival times with cure rates: application to data for early epilepsy and single seizures. Journal of Applied Statistics, 40(3):546–562, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Arostegui:2013:RCB

[2585] Inmaculada Arostegui, Vicente Núñez-Antón, and José M. Quintana. On the recoding of continuous and bounded indexes to a binomial form: an application to quality-of-life scores. Journal of Applied Statistics, 40(3):563– 582, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Enki:2013:CAI

[2586] Doyo G. Enki, Nickolay T. Trendafilov, and Ian T. Jolliffe. A clustering approach to interpretable principal components. Journal of Applied Statistics, 40(3):583–599, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Subedi:2013:SSP

[2587] S. Subedi, Z. Feng, R. Deardon, and F. S. Schenkel. SNP selection for predicting a quantitative trait. Journal of Applied Statistics, 40(3):600– 613, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2013:SBB

[2588] Dongliang Wang and Alan D. Hutson. Smooth bootstrap-based confidence intervals for one binomial proportion and difference of two proportions. Journal of Applied Statistics, 40(3):614–625, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Feirer:2013:MUF

[2589] Verena Feirer, Herwig Friedl, and Ulrich Hirn. Modelling over- and underdispersed frequencies of successful ink transmissions onto paper. Journal of Applied Statistics, 40(3):626– 643, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kan:2013:RRR

[2590] Betül Kan, Özlem Alpu, and Berna Yazici. Robust ridge and robust Liu estimator for regression based on the LTS estimator. Journal of Applied Statistics, 40(3):644–655, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Oleson:2013:PID

[2591] Jacob J. Oleson and Christopher K. Wikle. Predicting infectious disease outbreak risk via migratory waterfowl vectors. Journal of Applied Statistics, 40(3):656–673, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Melo:2013:DBA

[2592] Sandra E. Melo and Oscar O. Melo. Distance-based approach in univariate longitudinal data analysis. Journal of Applied Statistics, 40(3):674–692, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Heydarbeygie:2013:NME

[2593] Akbar Heydarbeygie and Nima Ahmadi. Nonparametric methods for the estimation of imputation uncertainty. Journal of Applied Statistics, 40(3): 693–698, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2013:GPE

[2594] Yang Li, Yichen Qin, Yanming Xie, and Feng Tian. Grouped penalization estimation of the osteoporosis data in the traditional Chinese medicine. Journal of Applied Statistics, 40(4):699– 711, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fescioglu-Unver:2013:CAN

[2595] Nilgun Fescioglu-Unver and Basak Tanyeri. A comparison of artificial neural network and multinomial logit models in predicting mergers. Journal of Applied Statistics, 40(4):712– 720, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kibria:2013:SRR

[2596] B. M. Golam Kibria, Kristofer Månsson, and Ghazi Shukur. Some ridge regression estimators for the zero-inflated Poisson model. Journal of Applied Statistics, 40(4):721–735, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2013:CDR

[2597] Hui Lin, Chong Wang, Peng Liu, and Derald J. Holtkamp. Construction of disease risk scoring systems using logistic group lasso: application to porcine reproductive and respiratory syndrome survey data. Journal of Applied Statistics, 40(4):736–746, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Marozzi:2013:ACS

[2598] Marco Marozzi. Adaptive choice of scale tests in flexible two-stage designs with applications in experimental ecology and clinical trials. Journal of Applied Statistics, 40(4):747– 762, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Gorecki:2013:SCL

[2599] T. Górecki. Sequential correction of linear classifiers. Journal of Applied Statistics, 40(4):763–776, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Shaddick:2013:LSB

[2600] Gavin Shaddick, Haojie Yan, Ruth Salway, Danielle Vienneau, Daphne Kounali, and David Briggs. Largescale Bayesian spatial modelling of air pollution for policy support. Journal of Applied Statistics, 40(4):777– 794, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ruggieri:2013:FPC

[2601] Mariantonietta Ruggieri, Antonella Plaia, Francesca Di Salvo, and Gianna Agró. Functional principal component analysis for the explorative analysis of multisite-multivariate air pollution time series with long gaps. Journal of Applied Statistics, 40(4):795–807, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

DeIaco:2013:UDM

[2602] Sandra De Iaco. On the use of different metrics for assessing complex pattern reproductions. Journal of Applied Statistics, 40(4):808–822, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chandra:2013:ESD

[2603] Hukum Chandra. Exploring spatial dependence in area-level random effect model for disaggregate-level crop yield estimation. Journal of Applied Statistics, 40(4):823–842, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Klungsoyr:2013:TVM

[2604] Ole Klungsøyr, Joe Sexton, Inger Sandanger, and Jan F. Nygård. A timevarying measurement error model for age of onset of a psychiatric diagnosis: applied to first depressive episode diagnosed by the Composite International Diagnostic Interview (CIDI). Journal of Applied Statistics, 40(4):843– 861, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Faisal:2013:NAC

[2605] Muhammad Faisal, Andreas Futschik, and Ijaz Hussain. A new approach to choose acceptance cutoff for approximate Bayesian computation. Journal of Applied Statistics, 40(4):862– 869, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Yan:2013:ENR

[2606] Zaizai Yan, Miaomiao Li, and Yalu Yan. An efficient non-rejective implementation of the πps sampling designs. Journal of Applied Statistics, 40(4):870–886, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hayter:2013:IDB

[2607] A. J. Hayter. Inferences on the difference between future observations for comparing two treatments. Journal of Applied Statistics, 40(4):887–900, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Otranto:2013:VCP

[2608] Edoardo Otranto. Volatility clustering in the presence of time-varying model parameters. Journal of Applied Statistics, 40(4):901–915, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2013:SCC

[2609] Kassim S. Mwitondi. Statistical computing in C++ and R. Journal of Applied Statistics, 40(4):916, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ober:2013:PGR

[2610] Pieter Bastiaan Ober. A practitioner’s guide to resampling for data analysis, data mining, and modeling. Journal of Applied Statistics, 40(4):917, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ober:2013:DDB

[2611] Pieter Bastiaan Ober. Data driven business decisions. Journal of Applied Statistics, 40(4):917–918, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mabrouk:2013:TSM

[2612] Anouar Ben Mabrouk. Time series modelling of neuroscience data. Journal of Applied Statistics, 40(4):918– 919, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Laberge:2013:SNP

[2613] Yves Laberge. Simulating nature: a philosophical study of computersimulation uncertainties and their role in climate science and policy advice. Journal of Applied Statistics, 40(4): 919–920, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Espejo:2013:S

[2614] Mariano Ruiz Espejo. Sampling. Journal of Applied Statistics, 40(4):920– 921, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kamarianakis:2013:ECD

[2615] Yiannis Kamarianakis. Ergodic control of diffusion processes. Journal of Applied Statistics, 40(4):921–922, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Beeley:2013:BRD

[2616] Chris Beeley. Behavioural research data analysis with R. Journal of Applied Statistics, 40(4):922, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mosammam:2013:GMS

[2617] A. M. Mosammam. Geostatistics:

modeling spatial uncertainty, second edition. Journal of Applied Statistics, 40(4):923, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakouch:2013:RS

[2618] Hassan S. Bakouch. R for statistics.

Journal of Applied Statistics, 40(4): 924, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2013:Ea

[2619] Anonymous. Erratum. Journal of Applied Statistics, 40(4):v, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fang:2013:TPM

[2620] Kuangnan Fang and Shuangge Ma.

Three-part model for fractional response variables with application to Chinese household health insurance coverage. Journal of Applied Statistics, 40(5):925–940, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2013:ITC

[2621] Seung-Gu Kim, Jeong-Soo Park, and Yung-Seop Lee. Identification of target clusters by using the restricted normal mixture model. Journal of Applied Statistics, 40(5):941–960, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Houlding:2013:SPC

[2622] B. Houlding and J. Haslett. Scheduling parallel conference sessions: an application of a novel hybrid clustering algorithm for ensuring constrained cardinality. Journal of Applied Statistics, 40(5):961–971, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2013:MBM

[2623] I-Tang Yu. A modification of the Box– Meyer method for finding the active factors in screening experiments. Journal of Applied Statistics, 40(5):972– 984, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Stewart:2013:ZIB

[2624] Connie Stewart. Zero-inflated beta distribution for modeling the proportions in quantitative fatty acid signature analysis. Journal of Applied Statistics, 40(5):985–992, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhou:2013:BMC

[2625] Yan Zhou, John A. D. Aston, and Adam M. Johansen. Bayesian model comparison for compartmental models with applications in positron emission tomography. Journal of Applied Statistics, 40(5):993–1016, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Nakamura:2013:MRL

[2626] Nagatomo Nakamura and Takahiro Tsuchiya. A model of regression lines through a common point: estimation of the focal point in wind-blown sand phenomena. Journal of Applied Statistics, 40(5):1017–1031, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Goncalves:2013:CCA

[2627] Antonio Carlos Gonçalves, Renan M.

V. R. Almeida, Marcos Pereira Estellita Lins, and Carlos P. Samanez. Canonical correlation analysis in the definition of weight restrictions for data envelopment analysis. Journal of Applied Statistics, 40(5):1032–1043, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zimmer:2013:IPH

[2628] David M. Zimmer. Intertemporal persistence in healthcare spending and utilization: the role of insurance. Journal of Applied Statistics, 40(5):1044–1063, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Islam:2013:GBB

[2629] M. Ataharul Islam, Abdulhamid A.

Alzaid, Rafiqul I. Chowdhury, and Khalaf S. Sultan. A generalized bivariate Bernoulli model with covariate dependence. Journal of Applied Statistics, 40(5):1064–1075, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Noor:2013:BII

[2630] Farzana Noor and Muhammad Aslam.

Bayesian inference of the inverse Weibull mixture distribution using type-I censoring. Journal of Applied Statistics, 40(5):1076–1089, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Nassiri:2013:GQR

[2631] Vahid Nassiri and Ignace Loris. A generalized quantile regression model. Journal of Applied Statistics, 40(5): 1090–1105, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Danish:2013:BER

[2632] Muhammad Yameen Danish and Muhammad Aslam. Bayesian estimation for randomly censored generalized exponential distribution under asymmetric loss functions. Journal of Applied Statistics, 40(5):1106–1119, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

McAssey:2013:EGF

[2633] Michael P. McAssey. An empirical goodness-of-fit test for multivariate distributions. Journal of Applied Statistics, 40(5):1120–1131, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Luo:2013:NVE

[2634] Jiong Luo and Zheng Su. A note on variance estimation in the Cox proportional hazards model. Journal of Applied Statistics, 40(5):1132–1139, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hua:2013:SCB

[2635] Steven Y. Hua, D. L. Hawkins, and Jihao Zhou. Statistical considerations in bioequivalence of two area under the concentration-time curves obtained from serial sampling data. Journal of Applied Statistics, 40(5):1140–1154, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Stahlschmidt:2013:BNS

[2636] Stephan Stahlschmidt, Helmut Tausendteufel, and Wolfgang K. Härdle. Bayesian networks for sex-related homicides: structure learning and prediction. Journal of Applied Statistics, 40(6):1155–1171, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Calabrese:2013:MSM

[2637] Raffaella Calabrese and Silvia Angela Osmetti. Modelling small and medium enterprise loan defaults as rare events: the generalized extreme value regression model. Journal of Applied Statistics, 40(6):1172–1188, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsai:2013:EPE

[2638] Wan-Min Tsai, Albert Vexler, and Gregory Gurevich. An extensive power evaluation of a novel two-sample density-based empirical likelihood ratio test for paired data with an application to a treatment study of attentiondeficit/hyperactivity disorder and severe mood dysregulation. Journal of Applied Statistics, 40(6):1189–1208, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Vicari:2013:MLR

[2639] Donatella Vicari and Maurizio Vichi.

Multivariate linear regression for heterogeneous data. Journal of Applied Statistics, 40(6):1209–1230, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hong:2013:MIM

[2640] Hyokyoung Grace Hong and Jianhui Zhou. A multi-index model for quantile regression with ordinal data. Journal of Applied Statistics, 40(6):1231–1245, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Jin:2013:HFP

[2641] Hua Jin and Qi Mo. Hip fracture prediction from a new classification algorithm based on recursive partitioning methods. Journal of Applied Statistics, 40(6):1246–1253, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Yanuar:2013:BSE

[2642] Ferra Yanuar, Kamarulzaman Ibrahim, and Abdul Aziz Jemain. Bayesian structural equation modeling for the health index. Journal of Applied Statistics, 40(6):1254–1269, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kurtek:2013:SAS

[2643] Sebastian Kurtek, Wei Wu, Gary E. Christensen, and Anuj Srivastava. Segmentation, alignment and statistical analysis of biosignals with application to disease classification. Journal of Applied Statistics, 40(6):1270–1288, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kundu:2013:BAP

[2644] Debasis Kundu. Bayesian analysis for partially complete time and type of failure data. Journal of Applied Statistics, 40(6):1289–1300, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Niaki:2013:EDM

[2645] Seyed Taghi Akhavan Niaki and Paravaneh Jahani. The economic design of multivariate binomial EWMA VSSI control charts. Journal of Applied Statistics, 40(6):1301–1318, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Vera:2013:DRO

[2646] Miler Jerkovic Vera, Bojanic Dubravka, Jorgovanovic Nikola, Ilic Vojin, and Petrovacki-Balj Bojana. Detecting and removing outlier(s) in electromyographic gait-related patterns. Journal of Applied Statistics, 40(6):1319–1332, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bercu:2013:SCM

[2647] Sophie Bercu and Frédéric Proïa. A SARIMAX coupled modelling applied to individual load curves intraday forecasting. Journal of Applied Statistics, 40(6):1333–1348, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Reschenhofer:2013:RTS

[2648] Erhard Reschenhofer. Robust testing for stationarity of global surface temperature. Journal of Applied Statistics, 40(6):1349–1361, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chandra:2013:SFA

[2649] K. Suresh Chandra, G. Gopal, and M. Ramadurai. A stochastic frontier approach to survival analysis. Journal of Applied Statistics, 40(6):1362–1371, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2013:NSM

[2650] Jiaqi Yang, Wei Zhang, and Baolin Wu.

A note on statistical method for genotype calling of high-throughput singlenucleotide polymorphism arrays. Journal of Applied Statistics, 40(6):1372– 1381, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2013:BRE

[2651] Jing Xu. Book review: Expect the unexpected: a first course in biostatistics, by Raluca Balan and Gilles Lamothe, ISBN-13 978-981-4291-32-3. Journal of Applied Statistics, 40(6):1382–1383, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2013:BRI

[2652] Jing Xu. Book review: Introduction to probability with Texas Hold’em examples, by Frederic Paik Schoenberg, ISBN-13 978-1-4398-27680. Journal of Applied Statistics, 40(6):1383–1384, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tiwari:2013:BRE

[2653] Abhay Kumar Tiwari. Book review: Economic time series, edited by William R. Bell, Scott H. Holan, and Tucker S. McElroy, ISBN-13 978-1-4398-4657-5. Journal of Applied Statistics, 40(6):1384–1385, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2013:PTM

[2654] Ying Chen. A powerful test method for analyzing unreplicated factorials. Journal of Applied Statistics, 40(7):1387– 1401, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Datta:2013:BFM

[2655] Saheli Datta, Raquel Prado, and Abel Rodríguez. Bayesian factor models in characterizing molecular adaptation. Journal of Applied Statistics, 40(7):

1402–1424, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghahroodi:2013:BAA

[2656] Z. Rezaei Ghahroodi and M. Ganjali.

A Bayesian approach for analysing longitudinal nominal outcomes using random coefficients transitional generalized logit model: an application to the labour force survey data. Journal of Applied Statistics, 40(7):1425–1445, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Yamamoto:2013:PSM

[2657] Kouji Yamamoto, Shota Murakami, and Sadao Tomizawa. Point-symmetry models and decomposition for collapsed square contingency tables. Journal of Applied Statistics, 40(7):1446–1452, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

BenMabrouk:2013:WFH

[2658] Anouar BenMabrouk and Olfa Zaafrane. Wavelet fuzzy hybrid model for physico-financial signals. Journal of Applied Statistics, 40(7):1453–1463, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cheng:2013:EIT

[2659] Conghua Cheng, Jinyuan Chen, and Jianming Bai. Exact inferences of the two-parameter exponential distribution and Pareto distribution with censored data. Journal of Applied Statistics, 40 (7):1464–1479, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Raqab:2013:DBG

[2660] Mohammad Z. Raqab. Discriminating between the generalized Rayleigh and Weibull distributions. Journal of Applied Statistics, 40(7):1480–1493, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Makambi:2013:ETN

[2661] K. H. Makambi. Extended tests for non-zero between-study variance. Journal of Applied Statistics, 40(7):1494– 1505, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Sertdemir:2013:CMM

[2662] Y. Sertdemir, H. R. Burgut, Z. N. Alparslan, I. Unal, and S. Gunasti. Comparing the methods of measuring multi-rater agreement on an ordinal rating scale: a simulation study with an application to real data. Journal of Applied Statistics, 40(7):1506–1519, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2013:AHM

[2663] Pao-Sheng Shen. Additive hazards model with truncated and doubly censored data. Journal of Applied Statistics, 40(7):1520–1532, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bastos:2013:PPF

[2664] Leonardo Soares Bastos and Joel Mauricio Correa da Rosa. Predicting probabilities for the 2010 FIFA World Cup games using a Poisson–gamma model. Journal of Applied Statistics, 40(7): 1533–1544, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Acitas:2013:OSE

[2665] Sukru Acitas, Pelin Kasap, Birdal Senoğlu, and Olcay Arslan. Onestep M-estimators: Jones and Faddy’s skewed t-distribution. Journal of Applied Statistics, 40(7):1545–1560, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Pettit:2013:EIO

[2666] L. I. Pettit and N. Sothinathan. Effect of individual observations on the Box– Cox transformation. Journal of Applied Statistics, 40(7):1561–1571, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chapela:2013:TMU

[2667] Jorge González Chapela. Things that make us different: analysis of deviance with time-use data. Journal of Applied Statistics, 40(7):1572–1585, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodrigues-Motta:2013:MMC

[2668] Mariana Rodrigues-Motta, Hildete P.

Pinheiro, Eduardo G. Martins, Márcio S. Araújo, and Sérgio F. dos Reis. Multivariate models for correlated count data. Journal of Applied Statistics, 40 (7):1586–1596, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2013:BEP

[2669] Sanjay Kumar Singh, Umesh Singh, and Dinesh Kumar. Bayesian estimation of parameters of inverse Weibull distribution. Journal of Applied Statistics, 40(7):1597–1607, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakir:2013:SSP

[2670] Saad T. Bakir. A subset selection procedure for multinomial distributions. Journal of Applied Statistics, 40(7):

1608–1618, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mante:2013:AOC

[2671] Claude Manté, Guillaume Bernard, Patrick Bonhomme, and David Nerini. Application of ordinal correspondence analysis for submerged aquatic vegetation monitoring. Journal of Applied Statistics, 40(8):1619–1638, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kazmi:2013:SSP

[2672] Syed Mohsin Ali Kazmi, Muhammad Aslam, Sajid Ali, and Nasir Abbas. Selection of suitable prior for the Bayesian mixture of a class of lifetime distributions under type-I censored datasets. Journal of Applied Statistics, 40(8):1639–1658, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kada:2013:MDT

[2673] Akiko Kada, Zhihong Cai, and Manabu Kuroki. Medical diagnostic test based on the potential test result approach: bounds and identification. Journal of Applied Statistics, 40(8):1659–1672, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Gang:2013:MCC

[2674] Ting-Ting Gang, Jun Yang, and Yu Zhao. Multivariate control chart based on the highest possibility region. Journal of Applied Statistics, 40 (8):1673–1681, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

DeAngelis:2013:DAS

[2675] Luca De Angelis and Leonard J. Paas. A dynamic analysis of stock markets using a hidden Markov model. Journal of Applied Statistics, 40(8):1682–1700, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Barbieri:2013:CIA

[2676] Laura Barbieri. Causality and interdependence analysis in linear econometric models with an application to fertility. Journal of Applied Statistics, 40 (8):1701–1716, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Asgharzadeh:2013:PPL

[2677] A. Asgharzadeh, Hassan S. Bakouch, and L. Esmaeili. Pareto Poisson– Lindley distribution with applications. Journal of Applied Statistics, 40(8): 1717–1734, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ertas:2013:DIO

[2678] Hasan Ertaş, Murat Erisoğlu, and Selahattin Kaçıranlar. Detecting influential observations in Liu and modified Liu estimators. Journal of Applied Statistics, 40(8):1735–1745, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalaylioglu:2013:BSM

[2679] Z. I. Kalaylioğlu and O. Ozturk.

Bayesian semiparametric models for nonignorable missing mechanisms in generalized linear models. Journal of Applied Statistics, 40(8):1746–1763, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Trede:2013:DSR

[2680] Mark Trede and Cornelia Savu. Do stock returns have an Archimedean copula? Journal of Applied Statistics, 40(8):1764–1778, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2013:BGC

[2681] Zhiyong Zhang. Bayesian growth curve models with the generalized error distribution. Journal of Applied Statistics, 40(8):1779–1795, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ferreira:2013:BMA

[2682] Guillermo Ferreira, Luis M. Castro, Victor H. Lachos, and Ronaldo Dias. Bayesian modeling of autoregressive partial linear models with scale mixture of normal errors. Journal of Applied Statistics, 40(8):1796–1816, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Maurya:2013:MCC

[2683] Vishal Maurya, Amar Nath Gill, and Parminder Singh. Multiple comparisons with a control for exponential location parameters under heteroscedasticity. Journal of Applied Statistics, 40 (8):1817–1830, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Annoni:2013:MIW

[2684] Paola Annoni, Dorota Weziak-Bialowolska, and Hania Farhan. Measuring the impact of the Web: Rasch modelling for survey evaluation. Journal of Applied Statistics, 40(8):1831–1851, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2013:C

[2685] Anonymous. Corrigendum. Journal of Applied Statistics, 40(8):v, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Canning:2013:NPE

[2686] David Canning, Declan French, and Michael Moore. Non-parametric estimation of data dimensionality prior to data compression: the case of the human development index. Journal of Applied Statistics, 40(9):1853–1863, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Achcar:2013:BBB

[2687] Jorge Alberto Achcar, Emílio Augusto Coelho-Barros, and Josmar Mazucheli. Block and Basu bivariate lifetime distribution in the presence of cure fraction. Journal of Applied Statistics, 40 (9):1864–1874, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hung:2013:CAP

[2688] Wen-Liang Hung and De-Hua Chen.

Clustering algorithm for proximityrelation matrix and its applications. Journal of Applied Statistics, 40(9): 1875–1892, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fragoso:2013:ECA

[2689] Rui Fragoso and Maria Leonor da Silva Carvalho. Estimation of cost allocation coefficients at the farm level using an entropy approach. Journal of Applied Statistics, 40(9):1893–1906, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Green:2013:LAI

[2690] R. M. Green and M. S. Bebbington.

A longitudinal analysis of infant and senescent mortality using mixture models. Journal of Applied Statistics, 40 (9):1907–1920, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Gupta:2013:NBM

[2691] Arindam Gupta and Samba Siva Rao Pasupuleti. A new behavioural model for fertility schedules. Journal of Applied Statistics, 40(9):1921–1930, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Weiss:2013:IVA

[2692] Christian H. Weiß. Integer-valued autoregressive models for counts showing underdispersion. Journal of Applied Statistics, 40(9):1931–1948, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tong:2013:EPT

[2693] Tiejun Tong, Zeny Feng, Julia S. Hilton, and Hongyu Zhao. Estimating the proportion of true null hypotheses using the pattern of observed pvalues. Journal of Applied Statistics, 40(9):1949–1964, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Vicari:2013:BBT

[2694] Donatella Vicari and Johan René van Dorp. On a bounded bimodal two-sided distribution fitted to the Old-Faithful geyser data. Journal of Applied Statistics, 40(9):1965–1978, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Maboudou-Tchao:2013:DMC

[2695] Edgard M. Maboudou-Tchao and Douglas M. Hawkins. Detection of multiple change-points in multivariate data. Journal of Applied Statistics, 40(9): 1979–1995, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Yi:2013:OAR

[2696] Yanqing Yi and Yuan Yuan. An optimal allocation for response-adaptive designs. Journal of Applied Statistics, 40(9):1996–2008, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Barbieri:2013:BLC

[2697] Laura Barbieri, Mario Faliva, and Maria Grazia Zoia. Band-limited component estimation in time-limited economic series. Journal of Applied Statistics, 40(9):2009–2023, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2013:RVS

[2698] Weihua Zhao, Riquan Zhang, and Jicai Liu. Robust variable selection for the varying coefficient model based on composite L_{1}–L_{2} regression. Journal of Applied Statistics, 40(9):2024–2040, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Longford:2013:SC

[2699] Nicholas T. Longford. Searching for contaminants. Journal of Applied Statistics, 40(9):2041–2055, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zamani:2013:STT

[2700] Hossein Zamani and Noriszura Ismail. Score test for testing zero-inflated Poisson regression against zero-inflated generalized Poisson alternatives. Journal of Applied Statistics, 40(9):2056– 2068, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Borgoni:2013:IRP

[2701] Riccardo Borgoni, Valeria Tritto, and Daniela de Bartolo. Identifying radonprone building typologies by marginal modelling. Journal of Applied Statistics, 40(9):2069–2086, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

dosSantos:2013:CAC

[2702] Marcelo Justus dos Santos and Ana Lúcia Kassouf. A cointegration analysis of crime, economic activity, and police performance in São Paulo city. Journal of Applied Statistics, 40 (10):2087–2109, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Surapaitoolkorn:2013:VDS

[2703] Wantanee Surapaitoolkorn. Variable dimension via stochastic volatility model using FX rates. Journal of Applied Statistics, 40(10):2110–2128, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Santos:2013:MGI

[2704] Jose R. S. Santos, Caio L. N. Azevedo, and Heleno Bolfarine. A multiple group item response theory model with centered skew-normal latent trait distributions under a Bayesian framework. Journal of Applied Statistics, 40(10): 2129–2149, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Faria:2013:FDM

[2705] S. Faria and F. Gonçalves. Financial data modeling by Poisson mixture regression. Journal of Applied Statistics, 40(10):2150–2162, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Benasseni:2013:CAS

[2706] Jacques Bénasséni. A concentration approach to sensitivity studies in statistical estimation problems. Journal of Applied Statistics, 40(10):2163–2180, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Rajic:2013:CIR

[2707] Vesna Ćojbasić Rajić and J. Stanojević. Confidence intervals for the ratio of two variances. Journal of Applied Statistics, 40(10):2181–2187, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Niu:2013:PLS

[2708] Hongli Niu and Jun Wang. Powerlaw scaling behavior analysis of financial time series model by voter interacting dynamic system. Journal of Applied Statistics, 40(10):2188–2203, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Altaf:2013:AAD

[2709] Saima Altaf, Muhammad Aslam, and Muhammad Aslam. Analysis of the amended Davidson model with order effect for paired comparison in the Bayesian paradigm. Journal of Applied Statistics, 40(10):2204–2218, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ozkale:2013:IMA

[2710] M. Revan Özkale. Influence measures in affine combination type regression. Journal of Applied Statistics, 40 (10):2219–2243, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2013:SMC

[2711] Yangxin Huang, Getachew A. Dagne, and Jeong-Gun Park. Segmental modeling of changing immunologic response for CD4 data with skewness, missingness and dropout. Journal of Applied Statistics, 40(10):2244–2258, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2013:REE

[2712] Min Wang, Jing Zhao, Xiaoqian Sun, and Chanseok Park. Robust explicit estimation of the twoparameter Birnbaum–Saunders distribution. Journal of Applied Statistics, 40(10):2259–2274, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Costantini:2013:FIP

[2713] Mauro Costantini. Forecasting the industrial production using alternative factor models and business survey data. Journal of Applied Statistics, 40(10): 2275–2289, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hassani:2013:FBD

[2714] Hossein Hassani, Saeed Heravi, Gary Brown, and Daniel Ayoubkhani. Forecasting before, during, and after recession with singular spectrum analysis. Journal of Applied Statistics, 40 (10):2290–2302, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Demirhan:2013:BEL

[2715] Haydar Demirhan. Bayesian estimation of log odds ratios over two-way contingency tables with intraclass correlated cells. Journal of Applied Statistics, 40 (10):2303–2316, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Seifert:2013:TDI

[2716] Ludovic Seifert, Jean-François Coeurjolly, Romain Hérault, Léo Wattebled, and Keith Davids. Temporal dynamics of inter-limb coordination in ice climbing revealed through change-point analysis of the geodesic mean of circular data. Journal of Applied Statistics, 40(11):2317–2331, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2013:REL

[2717] Chor Foon Tang. A revisitation of the export-led growth hypothesis in Malaysia using the leveraged bootstrap simulation and rolling causality techniques. Journal of Applied Statistics, 40 (11):2332–2340, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chiu:2013:PPO

[2718] Jing-Er Chiu and Chih-Hsin Tsai.

Properties and performance of onesided cumulative count of conforming chart with parameter estimation in high-quality processes. Journal of Applied Statistics, 40(11):2341–2353, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Vilca:2013:VSM

[2719] Filidor Vilca, Mariana Rodrigues-Motta, and Víctor Leiva. On a variance stabilizing model and its application to genomic data. Journal of Applied Statistics, 40(11):2354–2371, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2013:SLS

[2720] Lei Wang, Yukun Liu, Wei Wu, and Xiaolong Pu. Sequential LND sensitivity test for binary response data. Journal of Applied Statistics, 40(11):2372–2384, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fawcett:2013:MSC

[2721] Lee Fawcett and Neil Thorpe. Mobile safety cameras: estimating casualty reductions and the demand for secondary healthcare. Journal of Applied Statistics, 40(11):2385–2406, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2013:ITF

[2722] M. H. Lee, H. J. Sadaei, and Suhartono. Improving TAIEX forecasting using fuzzy time series with Box–Cox power transformation. Journal of Applied Statistics, 40(11):2407–2422, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Graversen:2013:EPD

[2723] Therese Graversen and Steffen Lauritzen. Estimation of parameters in DNA mixture analysis. Journal of Applied Statistics, 40(11):2423–2436, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kus:2013:OPG

[2724] Coskun Kus, Yunus Akdogan, and Shuo-Jye Wu. Optimal progressive group censoring scheme under cost considerations for Pareto distribution. Journal of Applied Statistics, 40(11): 2437–2450, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Erisoglu:2013:HDM

[2725] Ülkü Erisoğlu, Murat Erisoğlu, and Nazif Çalis. Heterogeneous data modeling with two-component Weibull– Poisson distribution. Journal of Applied Statistics, 40(11):2451–2461, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Schall:2013:AFR

[2726] Robert Schall and Dianne Weatherall. Accuracy and fairness of rain rules for interrupted one-day cricket matches. Journal of Applied Statistics, 40(11):2462–2479, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Milionis:2013:DPP

[2727] Alexandros E. Milionis and Evangelia Papanagiotou. Decomposing the predictive performance of the moving average trading rule of technical analysis: the contribution of linear and non-linear dependencies in stock returns. Journal of Applied Statistics, 40 (11):2480–2494, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wooff:2013:RSF

[2728] David A. Wooff and Amin Jamalzadeh.

Robust and scale-free effect sizes for non-normal two-sample comparisons, with applications in e-commerce. Journal of Applied Statistics, 40(11):2495– 2515, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Han:2013:HBI

[2729] Gang Han, Yangxin Huang, Qizhai Li, Lili Chen, and Xi Zhang. Hybrid Bayesian inference on HIV viral dynamic models. Journal of Applied Statistics, 40(11):2516–2532, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2013:NMC

[2730] Yao Zhang, Eric T. Bradlow, and Dylan S. Small. New measures of clumpiness for incidence data. Journal of Applied Statistics, 40(11):2533–2548, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2013:ESS

[2731] Jiin-Huarng Guo and Wei-Ming Luh. Efficient sample size allocation with cost constraints for heterogeneousvariance group comparison. Journal of Applied Statistics, 40(12):2549–2563, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Androulakis:2013:NVS

[2732] E. Androulakis and C. Koukouvinos. A new variable selection method for uniform designs. Journal of Applied Statistics, 40(12):2564–2578, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Alberink:2013:PLR

[2733] Ivo Alberink, Annabel Bolck, and Sonja Menges. Posterior likelihood ratios for evaluation of forensic trace evidence given a two-level model on the data. Journal of Applied Statistics, 40 (12):2579–2600, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Imon:2013:IMH

[2734] A. H. M. Rahmatullah Imon and Ali S. Hadi. Identification of multiple high leverage points in logistic regression. Journal of Applied Statistics, 40(12): 2601–2616, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kibria:2013:PNC

[2735] B. M. Golam Kibria and Shipra Banik.

Parametric and nonparametric confidence intervals for estimating the difference of means of two skewed populations. Journal of Applied Statistics, 40 (12):2617–2636, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2013:VSP

[2736] Sanying Feng and Liugen Xue. Variable selection for partially varying coefficient single-index model. Journal of Applied Statistics, 40(12):2637–2652, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kashikar:2013:IAM

[2737] Akanksha S. Kashikar, Neelabh Rohan, and T. V. Ramanathan. Integer autoregressive models with structural breaks. Journal of Applied Statistics, 40(12):2653–2669, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahmood:2013:FMC

[2738] Sharif Mahmood, Begum Zainab, and A. H. M. Mahbub Latif. Frailty modeling for clustered survival data: an application to birth interval in Bangladesh. Journal of Applied Statistics, 40(12):2670–2680, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Forkman:2013:URV

[2739] Johannes Forkman. The use of a reference variety for comparisons in incomplete series of crop variety trials. Journal of Applied Statistics, 40(12):2681– 2698, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2013:FMC

[2740] Le Kang, Randy Carter, Kathleen Darcy, James Kauderer, and Shu-Yuan Liao. A fast Monte Carlo expectationmaximization algorithm for estimation in latent class model analysis with an application to assess diagnostic accuracy for cervical neoplasia in women with atypical glandular cells. Journal of Applied Statistics, 40(12):2699–2719, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ateya:2013:EUM

[2741] Saieed F. Ateya. Estimation under modified Weibull distribution based on right censored generalized order statistics. Journal of Applied Statistics, 40 (12):2720–2734, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Karia:2013:FIA

[2742] Abdul Aziz Karia, Imbarine Bujang, and Ismail Ahmad. Fractionally integrated ARMA for crude palm oil prices prediction: case of potentially overdifference. Journal of Applied Statistics, 40(12):2735–2748, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2013:MTB

[2743] Gang Wang, Jun Wang, and Mingyu Wang. Modular-transform based clustering. Journal of Applied Statistics, 40 (12):2749–2759, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Deniz:2013:NDD

[2744] Emilio Gómez Déniz. A new discrete distribution: properties and applications in medical care. Journal of Applied Statistics, 40(12):2760–2770, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Moklyachuk:2013:BRA

[2745] Mikhail Moklyachuk. Book review:

Advances in time series forecasting, by Cagdas Hakan Aladag and Erol Eğrioğlu, ISBN-13 978-1-60805-373-5. Journal of Applied Statistics, 40(12): 2771–2772, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Laberge:2013:BRH

[2746] Yves Laberge. Book review: Handbook of statistics in clinical oncology. Journal of Applied Statistics, 40(12):2772– 2773, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Stoimenova:2013:BRM

[2747] Eugenia Stoimenova. Book review:

Methodology in robust and nonparametric statistics, by Jana Jarečová, Pranab K. Sen, and Jan Picek, ISBN-13 978-1-4398-4068-9. Journal of Applied Statistics, 40(12):2773, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Webster:2013:BRB

[2748] Mark Webster. Book review: Bayesian statistics, by Peter M. Lee, ISBN 978-1-1183-3257-3. Journal of Applied Statistics, 40(12):2773–2774, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Shang:2013:BBP

[2749] Han Lin Shang. The BUGS book:

a practical introduction to Bayesian analysis. Journal of Applied Statistics, 40(12):2774–2775, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ober:2013:ILR

[2750] Pieter Bastiaan Ober. Introduction to linear regression analysis. Journal of Applied Statistics, 40(12):2775–2776, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Porzio:2013:RAE

[2751] Giovanni C. Porzio. Regression analysis by example. Journal of Applied Statistics, 40(12):2776–2777, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Keeble:2013:MLE

[2752] Claire Keeble. Maximum-likelihood estimation for sample surveys. Journal of Applied Statistics, 40(12):2777, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Espejo:2013:CIP

[2753] Mariano Ruiz Espejo. Confidence intervals for proportions and related measures of effect size. Journal of Applied Statistics, 40(12):2778, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Espejo:2013:DAE

[2754] Mariano Ruiz Espejo. Design and analysis of experiments in the health sciences. Journal of Applied Statistics, 40 (12):2778–2779, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2013:Eb

[2755] Anonymous. Erratum. Journal of Applied Statistics, 40(12):v, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2013:EBa

[2756] Anonymous. Editorial board. Journal of Applied Statistics, 40(12):ebi, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Editor:2014:E

[2757] Robert G. Aykroyd Editor. Editorial.

Journal of Applied Statistics, 41(1):1, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Gulpinar:2014:MOG

[2758] Nalan Gülpinar and Kabir Katata.

Modelling oil and gas supply disruption risks using extreme-value theory and copula. Journal of Applied Statistics, 41(1):2–25, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Han:2014:MCS

[2759] Sung Wan Han, Rickson C. Mesquita, Theresa M. Busch, and Mary E. Putt. A method for choosing the smoothing parameter in a semi-parametric model for detecting change-points in blood flow. Journal of Applied Statistics, 41(1):26–45, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2014:SPM

[2760] Xiaowei Yang, Bin Peng, Rongqi Chen, Qian Zhang, Dianwen Zhu, Qing J. Zhang, Fuzhong Xue, and Lihong Qi. Statistical profiling methods with hierarchical logistic regression for healthcare providers with binary outcomes. Journal of Applied Statistics, 41(1):46– 59, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Adeniji:2014:IDA

[2761] Abidemi K. Adeniji, Steven H. Belle, and Abdus S. Wahed. Incorporating diagnostic accuracy into the estimation of discrete survival function. Journal of Applied Statistics, 41(1):60–72, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Puig:2014:BCA

[2762] Xavier Puig and Josep Ginebra. A Bayesian cluster analysis of election results. Journal of Applied Statistics, 41(1):73–94, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2014:VSV

[2763] Weihua Zhao, Riquan Zhang, Yazhao Lv, and Jicai Liu. Variable selection for varying dispersion beta regression model. Journal of Applied Statistics, 41(1):95–108, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Samawi:2014:NTS

[2764] Hani M. Samawi and Robert Vogel. Notes on two sample tests for partially correlated (paired) data. Journal of Applied Statistics, 41(1):109– 117, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lozano-Aguilera:2014:PPP

[2765] E. D. Lozano-Aguilera, María Dolores Estudillo-Martínez, and Sonia Castillo-Gutiérrez. A proposal for plotting positions in probability plots. Journal of Applied Statistics, 41(1):118– 126, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

DeSantis:2014:APM

[2766] Stacia M. DeSantis, Christos Lazaridis, Shuang Ji, and Francis G. Spinale. Analyzing propensity matched zeroinflated count outcomes in observational studies. Journal of Applied Statistics, 41(1):127–141, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Laurini:2014:DFD

[2767] Márcio Poletti Laurini. Dynamic functional data analysis with nonparametric state space models. Journal of Applied Statistics, 41(1):142– 163, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Roukema:2014:FDA

[2768] Boudewijn F. Roukema. A first-digit anomaly in the 2009 Iranian Presidential election. Journal of Applied Statistics, 41(1):164–199, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Saghir:2014:CCM

[2769] Aamir Saghir and Zhengyan Lin. Control chart for monitoring multivariate COM–Poisson attributes. Journal of Applied Statistics, 41(1):200– 214, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Panahi:2014:PEP

[2770] Hanieh Panahi and Abdolreza Sayyareh. Parameter estimation and prediction of order statistics for the Burr Type XII distribution with type II censoring. Journal of Applied Statistics, 41(1):215–232, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Holgersson:2014:DVV

[2771] H. E. T. Holgersson, L. Nordström, and Ö. Öner. Dummy variables vs. category-wise models. Journal of Applied Statistics, 41(2):233–241, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2014:IWU

[2772] Ting-Li Chen and Stuart Geman. Image warping using radial basis functions. Journal of Applied Statistics, 41(2):242–258, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2014:CMV

[2773] S. H. Lin. Comparing the mean vectors of two independent multivariate log-normal distributions. Journal of Applied Statistics, 41(2):259– 274, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Nademi:2014:MAA

[2774] Arash Nademi and Rahman Farnoosh. Mixtures of autoregressive-autoregressive conditionally heteroscedastic models: semi-parametric approach. Journal of Applied Statistics, 41(2):275–293, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hokimoto:2014:NHH

[2775] Tsukasa Hokimoto and Kunio Shimizu. A non-homogeneous hidden Markov model for predicting the distribution of sea surface elevation. Journal of Applied Statistics, 41(2):294–319, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fioruci:2014:BMG

[2776] José A. Fioruci, Ricardo S. Ehlers, and Marinho G. Andrade Filho. Bayesian multivariate GARCH models with dynamic correlations and asymmetric error distributions. Journal of Applied Statistics, 41(2):320–331, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

McCullough:2014:RNS

[2777] Michael McCullough, Thomas L.

Marsh, and Ron C. Mittelhammer. Reconstructing nonlinear structure in regression residuals. Journal of Applied Statistics, 41(2):332–350, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhou:2014:PTM

[2778] Ming Zhou and Yongzhao Shao. A powerful test for multivariate normality. Journal of Applied Statistics, 41(2): 351–363, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Atsedeweyn:2014:LRM

[2779] A. Asrat Atsedeweyn and K. Srinivasa Rao. Linear regression model with new symmetric distributed errors. Journal of Applied Statistics, 41(2):364– 381, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Shiu:2014:DPO

[2780] Ji-Liang Shiu and Chia-Hung D. Sun.

The determinants of price in online auctions: more evidence from unbalanced panel data. Journal of Applied Statistics, 41(2):382–392, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2014:VSM

[2781] Xiaoguang Wang and Junhui Fan. Variable selection for multivariate generalized linear models. Journal of Applied Statistics, 41(2):393–406, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Amiri:2014:REA

[2782] Saeid Amiri. The resampling of entropies with the application of biodiversity. Journal of Applied Statistics, 41(2):407–422, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Plummer:2014:BAL

[2783] Paul J. Plummer and Jie Chen. A Bayesian approach for locating change points in a compound Poisson process with application to detecting DNA copy number variations. Journal of Applied Statistics, 41(2):423–438, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Saha:2014:HBC

[2784] Krishna K. Saha, Roger Bilisoly, and Darius M. Dziuda. Hybrid-based confidence intervals for the ratio of two treatment means in the over-dispersed Poisson data. Journal of Applied Statistics, 41(2):439–453, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tutmez:2014:ANS

[2785] Bulent Tutmez. Analyzing nonstationarity in cement stone pit by median polish interpolation: a case study. Journal of Applied Statistics, 41(2): 454–466, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tian:2014:VSH

[2786] Guo-Liang Tian, Mingqiu Wang, and Lixin Song. Variable selection in the high-dimensional continuous generalized linear model with current status data. Journal of Applied Statistics, 41(3):467–483, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Dwivedi:2014:MER

[2787] Alok Kumar Dwivedi, Indika Mallawaarachchi, Soyoung Lee, and Patrick Tarwater. Methods for estimating relative risk in studies of common binary outcomes. Journal of Applied Statistics, 41(3): 484–500, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Niloofar:2014:NMI

[2788] P. Niloofar and M. Ganjali. A new multivariate imputation method based on Bayesian networks. Journal of Applied Statistics, 41(3):501–518, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cilan:2014:LCA

[2789] Çigdem Aricigil Çilan. Latent class analysis for measuring Turkish People’s future expectations for Turkey. Journal of Applied Statistics, 41(3):519– 529, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Dai:2014:VEN

[2790] Wenlin Dai and Tiejun Tong. Variance estimation in nonparametric regression with jump discontinuities. Journal of Applied Statistics, 41(3):530–545, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ullah:2014:DTS

[2791] Wali Ullah, Yasumasa Matsuda, and Yoshihiko Tsukuda. Dynamics of the term structure of interest rates and monetary policy: is monetary policy effective during zero interest rate policy? Journal of Applied Statistics, 41(3):546–572, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Sohn:2014:TFB

[2792] Insuk Sohn, Jooyong Shim, Changha Hwang, Sujong Kim, and Jae Won Lee. Transcription factor-binding site identification and gene classification via fusion of the supervised-weighted discrete kernel clustering and support vector machine. Journal of Applied Statistics, 41(3):573–581, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Schmiedt:2014:MEV

[2793] A. B. Schmiedt, H. H. Dickert, W. Bleck, and U. Kamps. Multivariate extreme value analysis and its relevance in a metallographical application. Journal of Applied Statistics, 41(3):582–595, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Diana:2014:HHE

[2794] Giancarlo Diana, Saba Riaz, and Javid Shabbir. Hansen and Hurwitz estimator with scrambled response on the second call. Journal of Applied Statistics, 41(3):596–611, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Torabi:2014:HBB

[2795] Mahmoud Torabi. Hierarchical Bayesian bivariate disease mapping: analysis of children and adults asthma visits to hospital. Journal of Applied Statistics, 41(3):612–621, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Louzada:2014:MCC

[2796] Francisco Louzada, Mário de Castro, Vera Tomazella, and Jhon F. B. Gonzales. Modeling categorical covariates for lifetime data in the presence of cure fraction by Bayesian partition structures. Journal of Applied Statistics, 41(3):622–634, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chaabane:2014:NAR

[2797] Najeh Chaâbane. A novel autoregressive fractionally integrated moving average-least-squares support vector machine model for electricity spot prices prediction. Journal of Applied Statistics, 41(3):635–651, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Nyssen:2014:STF

[2798] Edgard Nyssen and Wolfgang Jacquet.

A statistical testing framework for evaluating the quality of measurement processes. Journal of Applied Statistics, 41(3):652–659, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Tian:2014:IMG

[2799] Yuzhu Tian, Qianqian Zhu, and Maozai Tian. Inference for mixed generalized exponential distribution under progressively type-II censored samples. Journal of Applied Statistics, 41(3):660– 676, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cepeda-Cuervo:2014:BBR

[2800] Edilberto Cepeda-Cuervo, Jorge Alberto Achcar, and Liliana Garrido Lopera. Bivariate beta regression models: joint modeling of the mean, dispersion and association parameters. Journal of Applied Statistics, 41(3):677– 687, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2014:EIM

[2801] Yi Zhang, Haitao Chu, and Donglin Zeng. Evaluation of incomplete multiple diagnostic tests, with an application in the colon cancer family registry study. Journal of Applied Statistics, 41(3):688–700, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Cron:2014:CFM

[2802] Andrew Cron, Liang Zhang, and Deepak Agarwal. Collaborative filtering for massive multinomial data. Journal of Applied Statistics, 41(4):701– 715, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lucas:2014:MMS

[2803] Jean-Paul Lucas, Véronique Sébille, Alain Le Tertre, Yann Le Strat, and Lise Bellanger. Multilevel modelling of survey data: impact of the twolevel weights used in the pseudolikelihood. Journal of Applied Statistics, 41(4):716–732, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Albatineh:2014:CIE

[2804] Ahmed N. Albatineh, B. M. Golam Kibria, Meredith L. Wilcox, and Bashar Zogheib. Confidence interval estimation for the population coefficient of variation using ranked set sampling: a simulation study. Journal of Applied Statistics, 41(4):733– 751, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahmed:2014:BEB

[2805] Essam A. Ahmed. Bayesian estimation based on progressive Type-II censoring from two-parameter bathtubshaped lifetime model: an Markov chain Monte Carlo approach. Journal of Applied Statistics, 41(4):752– 768, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2014:MAD

[2806] Lingsong Zhang, Zhengyuan Zhu, and J. S. Marron. Multiresolution anomaly detection method for fractional Gaussian noise. Journal of Applied Statistics, 41(4):769–784, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2014:EST

[2807] Show-Lin Chen, Nen-Jing Chen, and Rwei-Ju Chuang. An empirical study on technical analysis: GARCH (1, 1) model. Journal of Applied Statistics, 41(4):785–801, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Hilow:2014:MCL

[2808] Hisham Hilow. Minimum cost linear trend-free 12-run fractional factorial designs. Journal of Applied Statistics, 41(4):802–816, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2014:GMA

[2809] Keya Rani Das and A. H. M. Rahmatullah Imon. Geometric median and its application in the identification of multiple outliers. Journal of Applied Statistics, 41(4):817–831, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ali:2014:GPC

[2810] Sajid Ali and Muhammad Riaz. On the generalized process capability under simple and mixture models. Journal of Applied Statistics, 41(4):832– 852, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mitra:2014:EBR

[2811] Sharmishtha Mitra and Amit Mitra. M-estimator-based robust estimation of the number of components of a superimposed sinusoidal signal model. Journal of Applied Statistics, 41(4): 853–878, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Zeng:2014:VSA

[2812] Ping Zeng, Yongyue Wei, Yang Zhao, Jin Liu, Liya Liu, Ruyang Zhang, Jianwei Gou, Shuiping Huang, and Feng Chen. Variable selection approach for zero-inflated count data via adaptive lasso. Journal of Applied Statistics, 41(4):879–894, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Fackle-Fornius:2014:MOD

[2813] Ellinor Fackle-Fornius and Linda Anna Wänström. Minimax D-optimal designs of contingent valuation experiments: willingness to pay for environmentally friendly clothes. Journal of Applied Statistics, 41(4):895– 908, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Gillard:2014:BRR

[2814] Jonathan Gillard. Book review: The R book, second edition. Journal of Applied Statistics, 41(4):909, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Dialsingh:2014:BRR

[2815] Isaac Dialsingh. Book review: Risk assessment and decision analysis with Bayesian networks, by N. Fenton and M. Neil, ISBN-13 978-1-4398-0910-5. Journal of Applied Statistics, 41(4): 910, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Dialsingh:2014:BRC

[2816] Isaac Dialsingh. Book review: Computational statistics, second edition, by Geof H. Givens and Jennifer A. Hoeting, ISBN-13 978-0-470-53331-4. Journal of Applied Statistics, 41(4):910– 911, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Espejo:2014:BRM

[2817] Mariano Ruiz Espejo. Book review:

Medical biostatistics, third edition, by Adhaya Indrayan, ISBN-13 978-1-4398-8414-0. Journal of Applied Statistics, 41(4):911, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Sahu:2014:BRS

[2818] Anshuman Sahu. Book review: Statistical methods in customer relationship management, edited by J. Mateu and W. G. Müller, ISBN-13 978-0-470-97429-2. Journal of Applied Statistics, 41(4):912, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mosammam:2014:BRS

[2819] A. M. Mosammam. Book review:

Spatio-temporal design: advances in efficient data acquisition, by V. Kumar and J. Andrew Petersen, ISBN-13 978-1-119-99320-9. Journal of Applied Statistics, 41(4):912–913, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakouch:2014:BRU

[2820] Hassan S. Bakouch. Book review: Using the Weibull distribution: Reliability, modeling and inference, by John I. McCool, ISBN-13 978-1-118-21798-6. Journal of Applied Statistics, 41(4): 913–914, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Popovic:2014:BRC

[2821] Bozidar V. Popović. Book review: A course on statistics for finance, by S. L. Sclove, ISBN-13 978-1-4398-9254-1. Journal of Applied Statistics, 41(4): 914–915, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Dialsingh:2014:BRA

[2822] Isaac Dialsingh. Book review: Applied categorical and count data analysis, by W. Tang, H. He, and M. X. Tu, ISBN-13 978-1-4398-0624-1. Journal of Applied Statistics, 41(4):915, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Neves:2014:BRC

[2823] Cláudia Neves. Book review: Categorical data analysis, third edition, by Alan Agresti, ISBN-13 978-0-470-46363-5. Journal of Applied Statistics, 41(4):915–916, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Moklyachuk:2014:BRE

[2824] Mikhail Moklyachuk. Book review: Exercises in probability, second edition, by Loïc Chaumont and Marc Yor, ISBN-13 978-1-107-60655-5. Journal of Applied Statistics, 41(4):916– 917, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2014:A

[2825] Anonymous. Acknowledgements. Journal of Applied Statistics, 41(4):918– 928, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Ou:2014:UGA

[2826] Shang-Ling Ou, Li yu Daisy Liu, and Yih-Chang Ou. Using a genetic algorithm-based RAROC model for the performance and persistence of the funds. Journal of Applied Statistics, 41(5):929–943, May 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Facevicova:2014:LAS

[2827] K. Facevicová, K. Hron, V. Todorov, D. Guo, and M. Templ. Logratio approach to statistical analysis of 2 × 2 compositional tables. Journal of Applied Statistics, 41(5):944–958, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fawcett:2014:EPS

[2828] Lee Fawcett and David Walshaw. Estimating the probability of simultaneous rainfall extremes within a region: a spatial approach. Journal of Applied Statistics, 41(5):959–976, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dunbar:2014:SSG

[2829] Martin X. Dunbar, Hani M. Samawi, Robert Vogel, and Lili Yu. Steady-state Gibbs sampler estimation for lung cancer data. Journal of Applied Statistics, 41(5):977–988, May 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2014:AFA

[2830] Kun Li, YongSheng Qian, and WenBo Zhao. An auxiliary function approach for Lasso in music composition using cellular automata. Journal of Applied Statistics, 41(5):989–997, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ozkale:2014:RER

[2831] M. Revan Özkale. The relative efficiency of the restricted estimators in linear regression models. Journal of Applied Statistics, 41(5):998–1027, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2014:JMB

[2832] Nanhua Zhang, Henian Chen, and Yuanshu Zou. A joint model of binary and longitudinal data with nonignorable missingness, with application to marital stress and late-life major depression in women. Journal of Applied Statistics, 41(5):1028–1039, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goodman:2014:AOO

[2833] Melody S. Goodman, Yi Li, Anne M. Stoddard, and Glorian Sorensen. Analysis of ordinal outcomes with longitudinal covariates subject to missingness. Journal of Applied Statistics, 41(5):1040–1052, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Templ:2014:SQS

[2834] M. Templ and P. Filzmoser. Simulation and quality of a synthetic closeto-reality employer-employee population. Journal of Applied Statistics, 41 (5):1053–1072, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Borowski:2014:ODS

[2835] Matthias Borowski, Nikolaus Rudak, Birger Hussong, Dominik Wied, Sonja Kuhnt, and Wolfgang Tillmann. Onand offline detection of structural breaks in thermal spraying processes. Journal of Applied Statistics, 41(5): 1073–1090, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Unal:2014:VBS

[2836] Ilker Ünal and H. Refik Burgut. Verification bias on sensitivity and specificity measurements in diagnostic medicine: a comparison of some approaches used for correction. Journal of Applied Statistics, 41(5):1091–1104, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mokhtari:2014:SBW

[2837] S. Mostafa Mokhtari, Hamid Alinejad-Rokny, and Hossein Jalalifar. Selection of the best well control system by using fuzzy multiple-attribute decisionmaking methods. Journal of Applied Statistics, 41(5):1105–1121, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Parchami:2014:NGP

[2838] A. Parchami, B. Sadeghpour-Gildeh, M. Nourbakhsh, and M. Mashinchi. A new generation of process capability indices based on fuzzy measurements. Journal of Applied Statistics, 41(5):1122–1136, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tasdan:2014:SSI

[2839] Feridun Tasdan and Meral Cetin. A simulation study on the influence of ties on uniform scores test for circular data. Journal of Applied Statistics, 41(5):1137–1146, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tasdan:2014:SPE

[2840] Feridun Tasdan and Ozgur Yeniay. A shift parameter estimation based on smoothed Kolmogorov–Smirnov. Journal of Applied Statistics, 41(5):1147– 1159, May 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2014:DMU

[2841] Jie Shen, Colin M. Gallagher, and QiQi Lu. Detection of multiple undocumented change-points using adaptive Lasso. Journal of Applied Statistics, 41(6):1161–1173, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brabec:2014:CSD

[2842] Marek Brabec, Viorel Badescu, and Marius Paulescu. Cloud shade by dynamic logistic modeling. Journal of Applied Statistics, 41(6):1174–1188, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Meshkani:2014:ACU

[2843] Mohammad Reza Meshkani, Afshin Fallah, and Amir Kavousi. Analysis of covariance under inverse Gaussian model. Journal of Applied Statistics, 41(6):1189–1202, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kandala:2014:GDC

[2844] Ngianga-Bakwin Kandala, Samuel O. M. Manda, William W. Tigbe, Henry Mwambi, and Saverio Stranges. Geographic distribution of cardiovascular comorbidities in South Africa: a national cross-sectional analysis. Journal of Applied Statistics, 41(6):1203–1216, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akay:2014:GEL

[2845] Kadri Ulas Akay. A graphical evaluation of logistic ridge estimator in mixture experiments. Journal of Applied Statistics, 41(6):1217–1232, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sakate:2014:VSP

[2846] D. M. Sakate and D. N. Kashid. Variable selection via penalized minimum ϕ-divergence estimation in logistic regression. Journal of Applied Statistics, 41(6):1233–1246, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sapp:2014:SEM

[2847] Stephanie Sapp, Mark J. van der Laan, and John Canny. Subsemble: an ensemble method for combining subsetspecific algorithm fits. Journal of Applied Statistics, 41(6):1247–1259, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2014:ICC

[2848] Guoyi Zhang. Improved R and s control charts for monitoring the process variance. Journal of Applied Statistics, 41(6):1260–1273, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dong:2014:CPL

[2849] Ying Dong, Lixin Song, Mingqiu Wang, and Ying Xu. Combined-penalized likelihood estimations with a diverging number of parameters. Journal of Applied Statistics, 41(6):1274–1285, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maia:2014:MSM

[2850] Rafael Pimentel Maia, Per Madsen, and Rodrigo Labouriau. Multivariate survival mixed models for genetic analysis of longevity traits. Journal of Applied Statistics, 41(6):1286–1306, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dawson:2014:ONA

[2851] Peter Dawson, Paul Downward, and Terence C. Mills. Olympic news and attitudes towards the Olympics: a compositional time-series analysis of how sentiment is affected by events. Journal of Applied Statistics, 41(6):1307–1314, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nurunnabi:2014:PIM

[2852] A. A. M. Nurunnabi, Ali S. Hadi, and A. H. M. R. Imon. Procedures for the identification of multiple influential observations in linear regression. Journal of Applied Statistics, 41(6):1315–1331, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beran:2014:HTD

[2853] Laurens Beran. Hypothesis tests to determine if all true positives have been identified on a receiver operating characteristic curve. Journal of Applied Statistics, 41(6):1332–1341, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2014:BGV

[2854] Xiao-Feng Wang, Bo Hu, Bin Wang, and Kuangnan Fang. Bayesian generalized varying coefficient models for longitudinal proportional data with errors-in-covariates. Journal of Applied Statistics, 41(6):1342–1357, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xing:2014:BMT

[2855] Zhengming Xing, Bradley Nicholson, Monica Jimenez, Timothy Veldman, Lori Hudson, Joseph Lucas, David Dunson, Aimee K. Zaas, Christopher W. Woods, Geoffrey S. Ginsburg, and Lawrence Carin. Bayesian modeling of temporal properties of infectious disease in a college student population. Journal of Applied Statistics, 41(6):1358–1382, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xie:2014:BZI

[2856] Feng-Chang Xie, Jin-Guan Lin, and Bo-Cheng Wei. Bayesian zero-inflated generalized Poisson regression model: estimation and case influence diagnostics. Journal of Applied Statistics, 41 (6):1383–1392, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2014:PDE

[2857] Chih-Yueh Wang. Partial differential equations for probabilists. Journal of Applied Statistics, 41(6):1393–1394, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beeley:2014:S

[2858] Chris Beeley. Statistics. Journal of Applied Statistics, 41(6):1394, June 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2014:C

[2859] Anonymous. Corrigendum. Journal of Applied Statistics, 41(7):v, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Munoz:2014:ODB

[2860] J. F. Muñoz, E. Álvarez, and M. Rueda. Optimum design-based ratio estimators of the distribution function. Journal of Applied Statistics, 41 (7):1395–1407, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Franco:2014:NSS

[2861] Bruno Chaves Franco, Philippe Castagliola, Giovanni Celano, and Antonio Fernando Branco Costa. A new sampling strategy to reduce the effect of autocorrelation on a control chart. Journal of Applied Statistics, 41(7):1408–1421, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nummi:2014:TAB

[2862] Tapio Nummi, Tiina Hakanen, Liudmila Lipiäinen, Ulla Harjunmaa, Matti K. Salo, Marja-Terttu Saha, and Nina Vuorela. A trajectory analysis of body mass index for Finnish children. Journal of Applied Statistics, 41(7):1422–1435, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ye:2014:ARR

[2863] Zhi-Sheng Ye, Jian-Guo Li, and Mengru Zhang. Application of ridge regression and factor analysis in design and production of alloy wheels. Journal of Applied Statistics, 41(7):1436–1452, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Heravi:2014:SSP

[2864] Saeed Heravi and Peter Morgan. Sampling schemes for price index construction: a performance comparison across the classification of individual consumption by purpose food groups. Journal of Applied Statistics, 41(7):1453–1470, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2014:GCI

[2865] Wei-Hwa Wu and Hsin-Neng Hsieh.

Generalized confidence interval estimation for the mean of delta-lognormal distribution: an application to New Zealand trawl survey data. Journal of Applied Statistics, 41(7):1471–1485, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chan:2014:PGP

[2866] J. S. K. Chan, W. Y. Wan, and P. L. H. Yu. A Poisson geometric process approach for predicting dropout and committed first-time blood donors. Journal of Applied Statistics, 41(7):1486–1503, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Senthilkumar:2014:DST

[2867] D. Senthilkumar and B. Esha Raffie. Designing and selection of two-plan variables scheme indexed by crossover point. Journal of Applied Statistics, 41(7):1504–1515, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andres:2014:TTA

[2868] Antonio Martín Andrés and María Álvarez Hernández. Two-tailed asymptotic inferences for a proportion. Journal of Applied Statistics, 41(7):1516–1529, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Carrasco:2014:EVB

[2869] Jalmar M. F. Carrasco, Silvia L. P. Ferrari, and Reinaldo B. Arellano-Valle. Errors-in-variables beta regression models. Journal of Applied Statistics, 41(7):1530–1547, July 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gupta:2014:ASD

[2870] Ramesh C. Gupta and Jie Huang.

Analysis of survival data by a Weibullgeneralized Poisson distribution. Journal of Applied Statistics, 41(7):1548– 1564, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lv:2014:QRV

[2871] Yazhao Lv, Riquan Zhang, Weihua Zhao, and Jicai Liu. Quantile regression and variable selection for the single-index model. Journal of Applied Statistics, 41(7):1565–1577, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Malik:2014:ILM

[2872] Rajat Malik, Rob Deardon, Grace P. S. Kwong, and Benjamin J. Cowling. Individual-level modeling of the spread of influenza within households. Journal of Applied Statistics, 41(7):1578–1592, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beyaztas:2014:RBJ

[2873] Ufuk Beyaztas, Aylin Alin, and Michael A. Martin. Robust BCa–JaB method as a diagnostic tool for linear regression models. Journal of Applied Statistics, 41(7):1593–1610, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Angelopoulos:2014:CCC

[2874] P. Angelopoulos, C. Koukouvinos, and A. Skountzou. A cusum control chart approach for screening active effects in orthogonal-saturated experiments. Journal of Applied Statistics, 41(7):1611–1618, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2014:RAR

[2875] Yang-Jin Kim. Regression analysis of recurrent events data with incomplete observation gaps. Journal of Applied Statistics, 41(7):1619–1626, July 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Din:2014:LPB

[2876] Siti Haslinda Mohd Din, Marek Molas, Jolanda Luime, and Emmanuel Lesaffre. Longitudinal profiles of bounded outcome scores as predictors for disease activity in rheumatoid arthritis patients: a joint modeling approach. Journal of Applied Statistics, 41(8): 1627–1644, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2014:RPE

[2877] Tao Lu, Yangxin Huang, Min Wang, and Feng Qian. A refined parameter estimating approach for HIV dynamic model. Journal of Applied Statistics, 41 (8):1645–1657, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2014:SGV

[2878] Weihua Zhao, Riquan Zhang, and Jicai Liu. Sparse group variable selection based on quantile hierarchical Lasso. Journal of Applied Statistics, 41 (8):1658–1677, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Calabrese:2014:OCR

[2879] Raffaella Calabrese. Optimal cut-off for rare events and unbalanced misclassification costs. Journal of Applied Statistics, 41(8):1678–1693, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bensoussan:2014:GLM

[2880] Alain Bensoussan, Pierre Bertrand, and Alexandre Brouste. A generalized linear model approach to seasonal aspects of wind speed modeling. Journal of Applied Statistics, 41(8):1694–1707, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Nasser:2014:TSG

[2881] Amjad D. Al-Nasser. Two steps generalized maximum entropy estimation procedure for fitting linear regression when both covariates are subject to error. Journal of Applied Statistics, 41 (8):1708–1720, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coolen-Maturi:2014:NWR

[2882] Tahani Coolen-Maturi. A new weighted rank coefficient of concordance. Journal of Applied Statistics, 41(8):1721– 1745, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Islam:2014:IML

[2883] M. Qamarul Islam, Fetih Yildirim, and Mehmet Yazici. Inference in multivariate linear regression models with elliptically distributed errors. Journal of Applied Statistics, 41(8):1746–1766, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deLuna:2014:SAU

[2884] Xavier de Luna and Mathias Lundin. Sensitivity analysis of the unconfoundedness assumption with an application to an evaluation of college choice effects on earnings. Journal of Applied Statistics, 41(8):1767–1784, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Louzada:2014:SND

[2885] Francisco Louzada, Paulo H. Ferreira, and Carlos A. R. Diniz. Skew-normal distribution for growth curve models in presence of a heteroscedasticity structure. Journal of Applied Statistics, 41 (8):1785–1798, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kouadri:2014:BSS

[2886] Abdelmalek Kouadri, Karim Baiche, and Mimoun Zelmat. Blind source separation filters-based-fault detection and isolation in a three tank system. Journal of Applied Statistics, 41(8): 1799–1813, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fan:2014:BMS

[2887] Tsai-Hung Fan, Yi-Fu Wang, and Yi-Chen Zhang. Bayesian model selection in linear mixed effects models with autoregressive(p) errors using mixture priors. Journal of Applied Statistics, 41 (8):1814–1829, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ippoliti:2014:CBS

[2888] Luigi Ippoliti, Simone Di Zio, and Arcangelo Merla. Classification of biomedical signals for differential diagnosis of Raynaud’s phenomenon. Journal of Applied Statistics, 41(8):1830– 1847, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez-Camblor:2014:SST

[2889] Pablo Martínez-Camblor, Carlos Carleos, Jesus Á. Baro, and Javier Cañón. Standard statistical tools for the breed allocation problem. Journal of Applied Statistics, 41(8):1848–1856, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fragkos:2014:AMS

[2890] Konstantinos C. Fragkos. Applied medical statistics using SAS. Journal of Applied Statistics, 41(8):1857–1858, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Popovic:2014:ESS

[2891] Bozidar V. Popović. Exercises and solutions in statistical theory. Journal of Applied Statistics, 41(8):1858, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ortega:2014:LLR

[2892] Edwin M. M. Ortega, Gauss M.

Cordeiro, Elizabeth M. Hashimoto, and Kahadawala Cooray. A log-linear regression model for the odd Weibull distribution with censored data. Journal of Applied Statistics, 41(9):1859–1880, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Leiva:2014:CIB

[2893] Víctor Leiva, Carolina Marchant, Helton Saulo, Muhammad Aslam, and Fernando Rojas. Capability indices for Birnbaum–Saunders processes applied to electronic and food industries. Journal of Applied Statistics, 41(9): 1881–1902, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Matsuura:2014:ERC

[2894] Shun Matsuura. Effectiveness of a random compound noise strategy for robust parameter design. Journal of Applied Statistics, 41(9):1903–1918, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zapata-Vazquez:2014:EEJ

[2895] Rita Esther Zapata-Vázquez, Anthony O’Hagan, and Leonardo Soares Bastos. Eliciting expert judgements about a set of proportions. Journal of Applied Statistics, 41(9):1919–1933, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baghfalaki:2014:JMM

[2896] T. Baghfalaki, M. Ganjali, and D. Berridge. Joint modeling of multivariate longitudinal mixed measurements and time to event data using a Bayesian approach. Journal of Applied Statistics, 41(9):1934–1955, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2014:EMC

[2897] Yun Lu, Sridhar Hannenhalli, Tom Cappola, and Mary Putt. An evaluation of Monte–Carlo logic and logicFS motivated by a study of the regulation of gene expression in heart failure. Journal of Applied Statistics, 41 (9):1956–1975, September 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tarng:2014:TOL

[2898] Chwu-Shiun Tarng. Third-order likelihood-based inference for the lognormal regression model. Journal of Applied Statistics, 41(9):1976–1988, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2014:SLE

[2899] W. Liu, J. C. Hsu, F. Bretz, A. J.

Hayter, and Y. Han. Shelf-life and its estimation in drug stability studies. Journal of Applied Statistics, 41 (9):1989–2000, September 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2014:LDA

[2900] Le Chen, Ao Yuan, Aiyi Liu, and Guanjie Chen. Longitudinal data analysis using Bayesian-frequentist hybrid random effects model. Journal of Applied Statistics, 41(9):2001–2010, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Drake:2014:MIB

[2901] Robert Drake and Apratim Guha.

A mutual information-based k-sample test for discrete distributions. Journal of Applied Statistics, 41(9):2011–2027, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wiltshire:2014:GCT

[2902] Jelani Wiltshire, Fred W. Huffer, and William C. Parker. A general class of test statistics for Van Valen’s Red Queen hypothesis. Journal of Applied Statistics, 41(9):2028–2043, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andreis:2014:MIR

[2903] Federico Andreis and Pier Alda Ferrari. Multidimensional item response theory models for dichotomous data in customer satisfaction evaluation. Journal of Applied Statistics, 41(9): 2044–2055, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodrigues:2014:EEN

[2904] Josemar Rodrigues, Gauss M. Cordeiro, and Jorge Bazan. An extended exponentiated-G-negative binomial family with threshold effect. Journal of Applied Statistics, 41(9):2056–2074, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Thorburn:2014:ADI

[2905] Daniel Thorburn and Can Tongur. Assessing direct and indirect seasonal decomposition in state space. Journal of Applied Statistics, 41(9):2075–2091, September 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mu:2014:SNR

[2906] Weiyan Mu and Shifeng Xiong. Some notes on robust sure independence screening. Journal of Applied Statistics, 41(10):2092–2102, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sakshaug:2014:GSD

[2907] Joseph W. Sakshaug and Trivellore E.

Raghunathan. Generating synthetic data to produce public-use microdata for small geographic areas based on complex sample survey data with application to the National Health Interview Survey. Journal of Applied Statistics, 41(10):2103–2122, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rouigueb:2014:BII

[2908] A. Rouigueb, S. Chitroub, and A. Bouridane. Bayesian inference over ICA models: application to multibiometric score fusion with quality estimates. Journal of Applied Statistics, 41(10):2123–2140, October 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haq:2014:MRS

[2909] Abdul Haq, Jennifer Brown, Elena Moltchanova, and Amer Ibrahim Al-Omari. Mixed ranked set sampling design. Journal of Applied Statistics, 41 (10):2141–2156, October 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2014:SAM

[2910] Jaehee Kim and Sooyoung Cheon.

Stochastic approximation Monte Carlo Gibbs sampling for structural change inference in a Bayesian heteroscedastic time series model. Journal of Applied Statistics, 41(10):2157–2177, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pike:2014:JMM

[2911] Francis Pike, Lisa A. Weissfeld, and Chung-Chou H. Chang. Joint modeling of multivariate censored longitudinal and event time data with application to the Genetic Markers of Inflammation Study. Journal of Applied Statistics, 41(10):2178–2191, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Choi:2014:PTF

[2912] Jiin Choi, Stewart J. Anderson, Thomas J. Richards, and Wesley K. Thompson. Prediction of transplantfree survival in idiopathic pulmonary fibrosis patients using joint models for event times and mixed multivariate longitudinal data. Journal of Applied Statistics, 41(10):2192–2205, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bergtold:2014:PRA

[2913] Jason S. Bergtold and Eberechukwu Onukwugha. The probabilistic reduction approach to specifying multinomial logistic regression models in health outcomes research. Journal of Applied Statistics, 41(10):2206–2221, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nitti:2014:DIA

[2914] M. Nitti and E. Ciavolino. A deflated indicators approach for estimating second-order reflective models through PLS–PM: an empirical illustration. Journal of Applied Statistics, 41(10):2222–2239, October 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiong:2014:NMS

[2915] Wei Xiong and Maozai Tian. A new model selection procedure based on dynamic quantile regression. Journal of Applied Statistics, 41(10):2240–2256, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2014:GPL

[2916] Lin Liu, Jianbo Li, and Riquan Zhang.

General partially linear additive transformation model with right-censored data. Journal of Applied Statistics, 41 (10):2257–2269, October 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ko:2014:ILB

[2917] Feng-Shou Ko. Identification of longitudinal biomarkers for survival by a score test derived from a joint model of longitudinal and competing risks data. Journal of Applied Statistics, 41(10): 2270–2281, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Withanage:2014:JED

[2918] N. Withanage, A. R. de Leon, and C. J.

Rudnisky. Joint estimation of diseasespecific sensitivities and specificities in reader-based multi-disease diagnostic studies of paired organs. Journal of Applied Statistics, 41(10):2282–2297, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramirez:2014:IAL

[2919] Octavio Ramirez, Jeff Mullen, and Alba J. Collart. Insights into the appropriate level of disaggregation for efficient time series model forecasting. Journal of Applied Statistics, 41(10): 2298–2311, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Blangiardo:2014:EBE

[2920] Marta Blangiardo and Gianluca Baio. Evidence of bias in the Eurovision song contest: modelling the votes using Bayesian hierarchical models. Journal of Applied Statistics, 41(10): 2312–2322, October 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Assumpcao:2014:ALI

[2921] R. A. B. Assumpção, M. A. Uribe-Opazo, and M. Galea. Analysis of local influence in geostatistics using Student’s t-distribution. Journal of Applied Statistics, 41(11):2323–2341, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Faraway:2014:RNE

[2922] Julian J. Faraway. Regression for non-Euclidean data using distance matrices. Journal of Applied Statistics, 41 (11):2342–2357, November 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Riedl:2014:KIS

[2923] Maximilian Riedl and Ingo Geishecker.

Keep it simple: estimation strategies for ordered response models with fixed effects. Journal of Applied Statistics, 41 (11):2358–2374, November 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rastogi:2014:EIE

[2924] Manoj Kumar Rastogi and Yogesh Mani Tripathi. Estimation for an inverted exponentiated Rayleigh distribution under type II progressive censoring. Journal of Applied Statistics, 41 (11):2375–2405, November 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fe:2014:EIR

[2925] Eduardo Fé. Estimation and inference in regression discontinuity designs with asymmetric kernels. Journal of Applied Statistics, 41(11):2406–2417, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2014:RIG

[2926] Kalyan Das and Angshuman Sarkar.

Robust inference for generalized partially linear mixed models that account for censored responses and missing covariates — an application to Arctic data analysis. Journal of Applied Statistics, 41(11):2418–2436, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2014:LCQ

[2927] Ji-Xia Wang and Qing-Xian Xiao.

Local composite quantile regression estimation of time-varying parameter vector for multidimensional timeinhomogeneous diffusion models. Journal of Applied Statistics, 41(11):2437– 2449, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Simpson:2014:STH

[2928] Sean L. Simpson, Lloyd J. Edwards, Martin A. Styner, and Keith E. Muller. Separability tests for high-dimensional, low-sample size multivariate repeated measures data. Journal of Applied Statistics, 41(11):2450–2461, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sawadogo:2014:SAR

[2929] Amadou Sawadogo, Dominique Lafon, and Simplice Dossou Gbété. Statistical analysis of rank data from a visual matching of colored textures. Journal of Applied Statistics, 41(11): 2462–2482, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mitra:2014:BHM

[2930] Riten Mitra, Peter Müller, Yuan Ji, Yitan Zhu, Gordon Mills, and Yiling Lu. A Bayesian hierarchical model for inference across related reverse phase protein arrays experiments. Journal of Applied Statistics, 41(11): 2483–2492, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karimi:2014:EWF

[2931] Maryam Karimi and Sayed Mohammad Reza Alavi. The effect of weight function on hypothesis testing in weighted sampling. Journal of Applied Statistics, 41(11):2493–2503, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2014:BAJ

[2932] Dengke Xu, Zhongzhan Zhang, and Liucang Wu. Bayesian analysis of joint mean and covariance models for longitudinal data. Journal of Applied Statistics, 41(11):2504–2514, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2014:IIV

[2933] Su Yun Kang, James McGree, Peter Baade, and Kerrie Mengersen. An investigation of the impact of various geographical scales for the specification of spatial dependence. Journal of Applied Statistics, 41(11):2515–2538, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2014:FRM

[2934] N. Lu, T. Chen, P. Wu, D. Gunzler, H. Zhang, H. He, and X. M. Tu. Functional response models for intraclass correlation coefficients. Journal of Applied Statistics, 41(11):2539–2556, November 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dey:2014:PCG

[2935] Sanku Dey and Tanujit Dey. On progressively censored generalized inverted exponential distribution. Journal of Applied Statistics, 41(12):2557–2576, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2014:FTM

[2936] Jiaqing Xu and Cheng Peng. Fitting and testing the Marshall–Olkin extended Weibull model with randomly censored data. Journal of Applied Statistics, 41(12):2577–2595, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Micheas:2014:HBM

[2937] Athanasios Christou Micheas. Hierarchical Bayesian modeling of marked non-homogeneous Poisson processes with finite mixtures and inclusion of covariate information. Journal of Applied Statistics, 41(12):2596–2615, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zeviani:2014:GCD

[2938] Walmes Marques Zeviani, Paulo Justiniano Ribeiro, Jr., Wagner Hugo Bonat, Silvia Emiko Shimakura, and Joel Augusto Muniz. The Gammacount distribution in the analysis of experimental underdispersed data. Journal of Applied Statistics, 41(12):2616– 2626, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsai:2014:GTV

[2939] Shin-Fu Tsai. A generalized test variable approach for grain yield comparisons of rice. Journal of Applied Statistics, 41(12):2627–2638, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2014:ABS

[2940] Yangyi Xu, Inyoung Kim, and Patrick Schaumont. Adaptive Bayes sum test for the equality of two nonparametric functions. Journal of Applied Statistics, 41(12):2639–2657, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2014:EMW

[2941] R. Chen, T. Chen, N. Lu, H. Zhang, P. Wu, C. Feng, and X. M. Tu. Extending the Mann–Whitney–Wilcoxon rank sum test to longitudinal regression analysis. Journal of Applied Statistics, 41(12):2658–2675, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Crager:2014:PSM

[2942] Michael R. Crager and Gong Tang.

Patient-specific meta-analysis for risk assessment using multivariate proportional hazards regression. Journal of Applied Statistics, 41(12):2676–2695, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maucort-Boulch:2014:MIG

[2943] Delphine Maucort-Boulch, Pascal Roy, and Janez Stare. On a measure of information gain for regression models in survival analysis. Journal of Applied Statistics, 41(12):2696–2708, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jaeger:2014:BPS

[2944] Jonathan Jaeger and Philippe Lambert. Bayesian penalized smoothing approaches in models specified using differential equations with unknown error distributions. Journal of Applied Statistics, 41(12):2709–2726, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kafle:2014:BAS

[2945] Ram C. Kafle, Netra Khanal, and Chris P. Tsokos. Bayesian agestratified joinpoint regression model: an application to lung and brain cancer mortality. Journal of Applied Statistics, 41(12):2727–2742, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2014:SBH

[2946] Huaiye Zhang, Inyoung Kim, and Chun Gun Park. Semiparametric Bayesian hierarchical models for heterogeneous population in nonlinear mixed effect model: application to gastric emptying studies. Journal of Applied Statistics, 41(12):2743–2760, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samani:2014:SAI

[2947] E. Bahrami Samani. Sensitivity analysis for the identifiability with application to latent random effect model for the mixed data. Journal of Applied Statistics, 41(12):2761–2776, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Popovic:2014:BRU

[2948] Bozidar V. Popović. Book review: Understanding advanced statistical methods. Journal of Applied Statistics, 41 (12):2777, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nielsen:2014:BRI

[2949] Søren Feodor Nielsen. Book review:

An introduction to analysis of financial data with R. Journal of Applied Statistics, 41(12):2777–2778, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Laberge:2014:BRS

[2950] Yves Laberge. Book review: Selforganised criticality: theory, models and characterisation. Journal of Applied Statistics, 41(12):2778–2779, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Espejo:2014:BRS

[2951] Mariano Ruiz Espejo. Book review:

Statistical methods for handling incomplete data. Journal of Applied Statistics, 41(12):2779–2780, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsagris:2014:BRS

[2952] Michail Tsagris. Book review: Statistics through resampling methods and R, second edition. Journal of Applied Statistics, 41(12):2780–2781, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Laberge:2014:BRR

[2953] Yves Laberge. Book review: Risk modelling in general insurance from principles to practice. Journal of Applied Statistics, 41(12):2781–2782, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2014:EB

[2954] Anonymous. Editorial board. Journal of Applied Statistics, 41(12):ebi, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shah:2015:MDL

[2955] Nilesh H. Shah, Alison E. Hipwell, Stephanie D. Stepp, and Chung-Chou H. Chang. Measures of discrimination for latent group-based trajectory models. Journal of Applied Statistics, 42(1):1–11, January 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gusnanto:2015:SAR

[2956] Arief Gusnanto and Yudi Pawitan. Sparse alternatives to ridge regression: a random effects approach. Journal of Applied Statistics, 42(1):12–26, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elliott:2015:CRT

[2957] P. Elliott and K. Riggs. Confidence regions for two proportions from independent negative binomial distributions. Journal of Applied Statistics, 42(1):27–36, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lipovetsky:2015:ACF

[2958] Stan Lipovetsky. Analytical closedform solution for binary logit regression by categorical predictors. Journal of Applied Statistics, 42(1):37–49, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pasanen:2015:BSS

[2959] Leena Pasanen and Lasse Holmström.

Bayesian scale space analysis of temporal changes in satellite images. Journal of Applied Statistics, 42(1):50–70, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maslennikov:2015:IRF

[2960] Evgeny D. Maslennikov, Alexey V.

Sulimov, Igor A. Savkin, Marina A. Evdokimova, Dmitry A. Zateyshchikov, Valery V. Nosikov, and Vladimir B. Sulimov. An intuitive risk factors search algorithm: usage of the Bayesian network technique in personalized medicine. Journal of Applied Statistics, 42(1):71–87, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hawkins:2015:SCD

[2961] Douglas M. Hawkins and F. Lombard.

Segmentation of circular data. Journal of Applied Statistics, 42(1):88–97, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Malave:2015:BAO

[2962] Peter Malave and Arkadiusz Sitek.

Bayesian analysis of a one-compartment kinetic model used in medical imaging. Journal of Applied Statistics, 42 (1):98–113, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rastegaran:2015:SNR

[2963] Afsane Rastegaran and Mohammad Reza Zadkarami. A skew-normal random effects model for longitudinal ordinal categorical responses with missing data. Journal of Applied Statistics, 42(1):114–126, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dias:2015:AFD

[2964] Ronaldo Dias, Nancy L. Garcia, Guilherme Ludwig, and Marley A. Saraiva. Aggregated functional data model for near-infrared spectroscopy calibration and prediction. Journal of Applied Statistics, 42(1):127–143, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2015:JEL

[2965] Ying-Ju Chen, Wei Ning, and Arjun K. Gupta. Jackknife empirical likelihood method for testing the equality of two variances. Journal of Applied Statistics, 42(1):144–160, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ciavolino:2015:MQW

[2966] Enrico Ciavolino, Maurizio Carpita, and Amjad Al-Nasser. Modelling the quality of work in the Italian social cooperatives combining NPCA–RSM and SEM–GME approaches. Journal of Applied Statistics, 42(1):161–179, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Drosou:2015:CSU

[2967] Krystallenia Drosou, Andreas Artemiou, and Christos Koukouvinos. A comparative study of the use of large margin classifiers on seismic data. Journal of Applied Statistics, 42(1):180–201, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Apergis:2015:CRI

[2968] Nicholas Apergis, Christina Christou, James E. Payne, and James W. Saunoris. The change in real interest rate persistence in OECD countries: evidence from modified panel ratio tests. Journal of Applied Statistics, 42(1):202–213, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barrios:2015:QMS

[2969] Juan Martin Barrios and Eliane R. Rodrigues. A queueing model to study the occurrence and duration of ozone exceedances in Mexico City. Journal of Applied Statistics, 42(1):214–230, January 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Amin:2015:BMH

[2970] Zeinab Amin and Maram Salem.

Bayesian modelling of health insurance losses. Journal of Applied Statistics, 42 (2):231–251, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonat:2015:LAC

[2971] Wagner Hugo Bonat, Paulo Justiniano Ribeiro, Jr., and Walmes Marques Zeviani. Likelihood analysis for a class of beta mixed models. Journal of Applied Statistics, 42(2):252–266, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patino:2015:SPB

[2972] Jairo Alberto Fúquene Patiño. A semiparametric Bayesian extreme value model using a Dirichlet process mixture of gamma densities. Journal of Applied Statistics, 42(2):267–280, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2015:NEV

[2973] Young-Ju Kim. Nonparametric estimation of varying-coefficient single-index models. Journal of Applied Statistics, 42(2):281–291, February 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bao:2015:DFD

[2974] Xinzhong Bao, Qiuyan Tao, and Hongyu Fu. Dynamic financial distress prediction based on Kalman filtering. Journal of Applied Statistics, 42 (2):292–308, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2015:TVC

[2975] Haiyan Zhao, Fred Huffer, and Xu-Feng Niu. Time-varying coefficient models with ARMA–GARCH structures for longitudinal data analysis. Journal of Applied Statistics, 42(2): 309–326, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fang:2015:DDA

[2976] Yan Fang, Ling Liu, and JinZhi Liu. A dynamic double asymmetric copula generalized autoregressive conditional heteroskedasticity model: application to China’s and US stock market. Journal of Applied Statistics, 42 (2):327–346, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Adragni:2015:ISH

[2977] Kofi Placid Adragni. Independent screening in high-dimensional exponential family predictors’ space. Journal of Applied Statistics, 42(2):347–359, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2015:MRM

[2978] Pao-Sheng Shen. Median regression model with doubly truncated data. Journal of Applied Statistics, 42(2): 360–370, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tahata:2015:MAM

[2979] Kouji Tahata and Takuya Yoshimoto.

Marginal asymmetry model for square contingency tables with ordered categories. Journal of Applied Statistics, 42 (2):371–379, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wegmann:2015:BCP

[2980] Bertil Wegmann. Bayesian comparison of private and common values in structural second-price auctions. Journal of Applied Statistics, 42(2):380–397, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gorecki:2015:SCD

[2981] T. Górecki. Sequential combining in discriminant analysis. Journal of Applied Statistics, 42(2):398–408, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McPherson:2015:BCD

[2982] Sterling McPherson and Celestina Barbosa-Leiker. Biomarker classification derived from finite growth mixture modeling with a time-varying covariate: an example with phosphorus and glomerular filtration rate. Journal of Applied Statistics, 42(2):409–427, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shi:2015:SLI

[2983] Lei Shi, Md. Mostafizur Rahman, Wen Gan, and Jianhua Zhao. Stepwise local influence in generalized autoregressive conditional heteroskedasticity models. Journal of Applied Statistics, 42(2):428–444, February 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alhamzawi:2015:MSQ

[2984] Rahim Alhamzawi. Model selection in quantile regression models. Journal of Applied Statistics, 42(2):445–458, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hughes:2015:PSI

[2985] William Hughes. Paradoxes in scientific inference. Journal of Applied Statistics, 42(2):459–460, February 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tiwari:2015:GSB

[2986] Abhay Kumar Tiwari. Getting started with business analytics. Journal of Applied Statistics, 42(2):460–461, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tattar:2015:SMA

[2987] Prabhanjan Tattar. Statistical methods with applications to demography and life insurance. Journal of Applied Statistics, 42(2):461–462, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakouch:2015:GLM

[2988] Hassan S. Bakouch. Generalized linear models for categorical and continuous limited dependent variables. Journal of Applied Statistics, 42(2):462, February 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2015:MIA

[2989] Wei Liu and Shuyou Li. A multiple imputation approach to nonlinear mixed-effects models with covariate measurement errors and missing values. Journal of Applied Statistics, 42(3):463–476, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hossain:2015:ASN

[2990] Ahmed Hossain and Joseph Beyene.

Application of skew-normal distribution for detecting differential expression to microRNA data. Journal of Applied Statistics, 42(3):477–491, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2015:KFB

[2991] Himadri Ghosh, Bishal Gurung, and Prajneshu. Kalman filter-based modelling and forecasting of stochastic volatility with threshold. Journal of Applied Statistics, 42(3):492–507, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Seal:2015:EBE

[2992] Babulal Seal and Sk Jakir Hossain. Empirical Bayes estimation of parameters in Markov transition probability matrix with computational methods. Journal of Applied Statistics, 42 (3):508–519, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ciavolino:2015:GCE

[2993] E. Ciavolino and A. Calcagnì. Generalized cross entropy method for analysing the SERVQUAL model. Journal of Applied Statistics, 42(3):520–534, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Annoni:2015:MVP

[2994] Paola Annoni, Rainer Bruggemann, and Lars Carlsen. A multidimensional view on poverty in the European Union by partial order theory. Journal of Applied Statistics, 42(3):535–554, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beyler:2015:POP

[2995] Nicholas Beyler, Wayne Fuller, Sarah Nusser, and Gregory Welk. Predicting objective physical activity from self-report surveys: a model validation study using estimated generalized least-squares regression. Journal of Applied Statistics, 42(3):555–565, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wong:2015:EVI

[2996] Tong Siu Tung Wong and Wai Keung Li. Extreme values identification in regression using a peaks-over-threshold approach. Journal of Applied Statistics, 42(3):566–576, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yong:2015:DOC

[2997] Luo Yong, Zhu Bo, and Tang Yong. Dynamic optimal capital growth of diversified investment. Journal of Applied Statistics, 42(3):577–588, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Caporale:2015:IMR

[2998] Guglielmo Maria Caporale and Luis A.

Gil-Alana. Infant mortality rates: time trends and fractional integration. Journal of Applied Statistics, 42(3):589–602, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2015:ICC

[2999] Guoyi Zhang and Zhongxue Chen.

Inferences on correlation coefficients of bivariate log-normal distributions. Journal of Applied Statistics, 42(3): 603–613, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Krupskii:2015:TWM

[3000] Pavel Krupskii and Harry Joe. Tailweighted measures of dependence. Journal of Applied Statistics, 42(3): 614–629, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2015:RIR

[3001] Chun-Xia Zhang, Guan-Wei Wang, and Jun-Min Liu. RandGA: injecting randomness into parallel genetic algorithm for variable selection. Journal of Applied Statistics, 42(3):630–647, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garcia:2015:CRV

[3002] C. B. García, J. García, M. M. López Martín, and R. Salmerón. Collinearity: revisiting the variance inflation factor in ridge regression. Journal of Applied Statistics, 42(3):648–661, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tiwari:2015:SLR

[3003] Aviral Kumar Tiwari and Alexander Ludwig. Short- and long-run rolling causality techniques and optimal window-wise lag selection: an application to the export-led growth hypothesis. Journal of Applied Statistics, 42(3):662–675, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2015:UDC

[3004] J. Lee, Y. Wu, and H. Kim. Unbalanced data classification using support vector machines with active learning on scleroderma lung disease patterns. Journal of Applied Statistics, 42(3):676–689, March 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2015:C

[3005] Anonymous. Corrigendum. Journal of Applied Statistics, 42(4):v, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shahbaz:2015:ATF

[3006] Muhammad Shahbaz, Aviral Kumar Tiwari, and Mohammad Iqbal Tahir. Analyzing time-frequency relationship between oil price and exchange rate in Pakistan through wavelets. Journal of Applied Statistics, 42(4):690–704, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Celik:2015:SHB

[3007] Resit Çelik. Stabilizing heteroscedasticity for butterfly-distributed residuals by the weighting absolute centered external variable. Journal of Applied Statistics, 42(4):705–721, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alama:2015:GSB

[3008] L. Alamá, D. Conesa, A. Forte, and E. Tortosa-Ausina. The geography of Spanish bank branches. Journal of Applied Statistics, 42(4):722–744, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fu:2015:BAJ

[3009] Ying zi Fu, Pei xiao Chu, and Li ying Lu. A Bayesian approach of joint models for clustered zero-inflated count data with skewness and measurement errors. Journal of Applied Statistics, 42(4):745–761, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2015:VSM

[3010] Jing Chang and Herbert K. H. Lee.

Variable selection via a multi-stage strategy. Journal of Applied Statistics, 42(4):762–774, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

ElGhourabi:2015:NFS

[3011] Mohamed El Ghourabi, Amira Dridi, and Mohamed Limam. A new financial stress index model based on support vector regression and control chart. Journal of Applied Statistics, 42(4):775–788, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2015:SNF

[3012] M. Liu and T. I. Lin. Skew-normal factor analysis models with incomplete data. Journal of Applied Statistics, 42(4):789–805, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2015:UPT

[3013] Fei Yang, Lin Chen, Yang Cheng, Zhenxing Yao, and Xu Zhang. Urban public transport choice behavior analysis and service improvement policy-making: a case study from the metropolitan city, Chengdu, China. Journal of Applied Statistics, 42(4): 806–816, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Grover:2015:MIC

[3014] Gurprit Grover and Vinay K. Gupta. Multiple imputation of censored survival data in the presence of missing covariates using restricted mean survival time. Journal of Applied Statistics, 42(4):817–827, April 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hamada:2015:UPS

[3015] M. S. Hamada, B. L. Mitchell, and C. T. Necker. On uncertainty of a proportion from a stratified random sample of a small population. Journal of Applied Statistics, 42(4):828–833, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hoegh:2015:STM

[3016] Andrew Hoegh and Scotland Leman.

A spatio-temporal model for assessing winter damage risk to east coast vineyards. Journal of Applied Statistics, 42(4):834–845, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Panagiotakis:2015:MST

[3017] Costas Panagiotakis and Georgios Tziritas. A minimum spanning tree equipartition algorithm for microaggregation. Journal of Applied Statistics, 42(4):846–865, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yamamoto:2015:MDS

[3018] Kouji Yamamoto, Fumika Shimada, and Sadao Tomizawa. Measure of departure from symmetry for the analysis of collapsed square contingency tables with ordered categories. Journal of Applied Statistics, 42(4):866–875, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ullah:2015:GNS

[3019] Wali Ullah, Yasumasa Matsuda, and Yoshihiko Tsukuda. Generalized Nelson–Siegel term structure model: do the second slope and curvature factors improve the in-sample fit and outof-sample forecasts? Journal of Applied Statistics, 42(4):876–904, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jammalamadaka:2015:ADL

[3020] S. Rao Jammalamadaka and Elvynna Leong. Analysis of discrete lifetime data under middle-censoring and in the presence of covariates. Journal of Applied Statistics, 42(4):905–913, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pallmann:2015:AMA

[3021] Philip Pallmann. Applied metaanalysis with R. Journal of Applied Statistics, 42(4):914–915, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andrade:2015:SAH

[3022] Marina A. P. Andrade. Statistical analysis of human growth and development. Journal of Applied Statistics, 42(4):915, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gillard:2015:CPC

[3023] Jonathan Gillard. Constrained principal component analysis and related techniques. Journal of Applied Statistics, 42(4):916, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramsay:2015:HSP

[3024] Paul M. Ramsay. Handbook of spatial point-pattern analysis in ecology. Journal of Applied Statistics, 42(4):916–917, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Espejo:2015:MSS

[3025] Mariano Ruiz Espejo. Modern survey sampling. Journal of Applied Statistics, 42(4):917–918, April 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pandolfo:2015:BRC

[3026] Giuseppe Pandolfo. Book review: Circular statistics in R. Journal of Applied Statistics, 42(4):918–919, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brabec:2015:SMD

[3027] M. Brabec, O. Konár, M. Malý, I. Kasanický, and E. Pelikán. Statistical models for disaggregation and reaggregation of natural gas consumption data. Journal of Applied Statistics, 42(5):921–937, May 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alih:2015:CBM

[3028] Ekele Alih and Hong Choon Ong.

Cluster-based multivariate outlier identification and re-weighted regression in linear models. Journal of Applied Statistics, 42(5):938–955, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Billio:2015:GCM

[3029] Monica Billio and Silvio Di Sanzo.

Granger-causality in Markov switching models. Journal of Applied Statistics, 42(5):956–966, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cribari-Neto:2015:NHT

[3030] Francisco Cribari-Neto and Sadraque E. F. Lucena. Nonnested hypothesis testing in the class of varying dispersion beta regressions. Journal of Applied Statistics, 42(5):967–985, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sinan:2015:UAI

[3031] Alper Sinan and B. Bar Alkan. A useful approach to identify the multicollinearity in the presence of outliers. Journal of Applied Statistics, 42(5):986–993, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Otero:2015:BAI

[3032] Fernando A. Otero, Helcio R. Barreto Orlande, Gloria L. Frontini, and Guillermo E. Eliçabe. Bayesian approach to the inverse problem in a light scattering application. Journal of Applied Statistics, 42(5):994–1016, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lipovetsky:2015:PRI

[3033] Stan Lipovetsky and W. Michael Conklin. Predictor relative importance and matching regression parameters. Journal of Applied Statistics, 42(5):1017– 1031, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghufran:2015:OMS

[3034] Shazia Ghufran, Saman Khowaja, and M. J. Ahsan. Optimum multivariate stratified double sampling design: Chebyshev’s Goal Programming approach. Journal of Applied Statistics, 42(5):1032–1042, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alonso:2015:LLM

[3035] M. Pilar Alonso, Asunción Beamonte, Pilar Gargallo, and Manuel Salvador. Local labour markets delineation: an approach based on evolutionary algorithms and classification methods. Journal of Applied Statistics, 42(5): 1043–1063, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Najibi:2015:BAP

[3036] S. M. Najibi, M. R. Faghihi, M. Golalizadeh, and S. S. Arab. Bayesian alignment of proteins via Delaunay tetrahedralization. Journal of Applied Statistics, 42(5):1064–1079, May 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luo:2015:SIM

[3037] Sheng Luo, Xiao Su, Min Yi, and Kelly K. Hunt. Simultaneous inference of a misclassified outcome and competing risks failure time data. Journal of Applied Statistics, 42(5):1080–1090, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ganguli:2015:DFF

[3038] B. Ganguli, M. Naskar, E. J. Malloy, and E. A. Eisen. Determination of the functional form of the relationship of covariates to the log hazard ratio in a Cox model. Journal of Applied Statistics, 42(5):1091–1105, May 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Branscum:2015:THA

[3039] Adam J. Branscum, Dunlei Cheng, and J. Jack Lee. Testing hypotheses about medical test accuracy: considerations for design and inference. Journal of Applied Statistics, 42(5):1106–1119, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Teng:2015:BMC

[3040] Huei-Wen Teng, Wen-Liang Hung, and Yen-Ju Chao. Bayesian Markov chain Monte Carlo imputation for the transiting exoplanets with an application in clustering analysis. Journal of Applied Statistics, 42(5):1120–1132, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2015:MRL

[3041] Hongxia Yang and Aurelie Lozano. Multi-relational learning via hierarchical nonparametric Bayesian collective matrix factorization. Journal of Applied Statistics, 42(5):1133–1147, May 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garay:2015:BEC

[3042] Aldo M. Garay, Victor H. Lachos, and Heleno Bolfarine. Bayesian estimation and case influence diagnostics for the zero-inflated negative binomial regression model. Journal of Applied Statistics, 42(6):1148–1165, June 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miljkovic:2015:MVH

[3043] Tatjana Miljkovic and Nikita Barabanov. Modeling veterans health benefit grants using the expectation maximization algorithm. Journal of Applied Statistics, 42(6):1166–1182, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shahriari:2015:REM

[3044] Hamid Shahriari and Orod Ahmadi.

Robust estimation of the mean vector for high-dimensional data set using robust clustering. Journal of Applied Statistics, 42(6):1183–1205, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2015:BJM

[3045] Cindy Xin Feng. Bayesian joint modeling of correlated counts data with application to adverse birth outcomes. Journal of Applied Statistics, 42(6):1206–1222, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Serrat:2015:FBA

[3046] Carles Serrat, Montserrat Rué, Carmen Armero, Xavier Piulachs, Hèctor Perpiñán, Anabel Forte, Álvaro Páez, and Guadalupe Gómez. Frequentist and Bayesian approaches for a joint model for prostate cancer risk and longitudinal prostate-specific antigen data. Journal of Applied Statistics, 42(6):1223–1239, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qin:2015:REM

[3047] Guoyou Qin and Zhongyi Zhu. Robust estimation of mean and covariance for longitudinal data with dropouts. Journal of Applied Statistics, 42(6):1240– 1254, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chekouo:2015:PBM

[3048] Thierry Chekouo and Alejandro Murua. The penalized biclustering model and related algorithms. Journal of Applied Statistics, 42(6):1255–1277, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dany:2015:URP

[3049] Antoine Dany, Emmanuelle Dantony, Mad-Hélénie Elsensohn, Emmanuel Villar, Cécile Couchoud, and René Ecochard. Using repeated-prevalence data in multi-state modeling of renal replacement therapy. Journal of Applied Statistics, 42(6):1278–1290, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Grover:2015:AVF

[3050] Gurprit Grover, Ravi Vajala, and Prafulla Kumar Swain. On the assessment of various factors effecting the improvement in CD4 count of aids patients undergoing antiretroviral therapy using generalized Poisson regression. Journal of Applied Statistics, 42(6):1291–1305, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dowd:2015:TBM

[3051] Peter A. Dowd, Eulogio Pardo-Igúzquiza, and Juan José Egozcue. The total bootstrap median: a robust and efficient estimator of location and scale for small samples. Journal of Applied Statistics, 42(6):1306–1321, June 2015.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

dosSantos:2015:MAM

[3052] Thiago Rezende dos Santos and Enrico A. Colosimo. A modified approximate method for analysis of degradation data. Journal of Applied Statistics, 42(6):1322–1331, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2015:BUR

[3053] Xiaorong Yang. Bootstrap unit root test based on least absolute deviation estimation under dependence assumptions. Journal of Applied Statistics, 42(6):1332–1347, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2015:SMR

[3054] Xu Tang, Fah Fatt Gan, and Lingyun Zhang. Standardized mortality ratio for an estimated number of deaths. Journal of Applied Statistics, 42(6): 1348–1366, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2015:SDN

[3055] Shuling Wang, Xiaoyan Wang, and Jiangtao Dai. Statistical diagnosis for non-parametric regression models with random right censorship based on the empirical likelihood method. Journal of Applied Statistics, 42(6):1367–1373, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lui:2015:NEI

[3056] Kung-Jong Lui. Notes on estimation of the intraclass correlation under the AB/BA crossover trial. Journal of Applied Statistics, 42(6):1374–1381, June 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cheng:2015:BJT

[3057] Jung-Yu Cheng and Shinn-Jia Tzeng. Buckley–James-type estimation of quantile regression with recurrent gap time data. Journal of Applied Statistics, 42(7):1383–1401, July 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Niu:2015:TRR

[3058] Cuizhen Niu and Qiang Xia. Testing the rate ratio under inverse sampling based on gradient statistic. Journal of Applied Statistics, 42(7):1402–1420, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Young:2015:RER

[3059] Derek S. Young and David R. Hunter. Random effects regression mixtures for analyzing infant habituation. Journal of Applied Statistics, 42(7):1421–1441, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Thomas:2015:SRM

[3060] Bijamma Thomas, N. N. Midhu, and P. G. Sankaran. A software reliability model using mean residual quantile function. Journal of Applied Statistics, 42(7):1442–1457, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2015:AIT

[3061] Meihui Guo, Yi-Ting Guo, Chi-Jeng Wang, and Liang-Ching Lin. Assessing influential trade effects via highfrequency market reactions. Journal of Applied Statistics, 42(7):1458–1471, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karakoca:2015:CPE

[3062] Aydin Karakoca, Ulku Erisoğlu, and Murat Erisoğlu. A comparison of the parameter estimation methods for bimodal mixture Weibull distribution with complete data. Journal of Applied Statistics, 42(7):1472–1489, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ertas:2015:RTP

[3063] Hasan Ertaş, Selma Toker, and Selahattin Kaçıranlar. Robust two parameter ridge M-estimator for linear regression. Journal of Applied Statistics, 42(7):1490–1502, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hung:2015:ACA

[3064] Wen-Liang Hung and Jenn-Hwai Yang.

Automatic clustering algorithm for fuzzy data. Journal of Applied Statistics, 42(7):1503–1518, July 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gulshan:2015:MPB

[3065] Jahida Gulshan, Md. Mejbahuddin Mina, and Syed Shahadat Hossain. Migration pattern in Bangladesh: a covariate-dependent Markov model. Journal of Applied Statistics, 42(7): 1519–1530, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gil-Alana:2015:LST

[3066] Luis A. Gil-Alana. Linear and segmented trends in sea surface temperature data. Journal of Applied Statistics, 42(7):1531–1546, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roy:2015:EHF

[3067] Sugata Sen Roy and Moumita Chatterjee. Estimating the hazard functions of two alternately occurring recurrent events. Journal of Applied Statistics, 42(7):1547–1555, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yigiter:2015:OCN

[3068] Ayten Yiğiter, Jie Chen, Lingling An, and Nazan Danacioğlu. An online copy number variant detection method for short sequencing reads. Journal of Applied Statistics, 42(7):1556–1571, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cene:2015:AOF

[3069] E. Cene and F. Karaman. Analysing organic food buyers’ perceptions with Bayesian networks: a case study in Turkey. Journal of Applied Statistics, 42(7):1572–1590, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Veer:2015:WST

[3070] Karan Veer and Ravinder Agarwal. Wavelet and short-time Fourier transform comparison-based analysis of myoelectric signals. Journal of Applied Statistics, 42(7):1591–1601, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hsieh:2015:NET

[3071] Jin-Jian Hsieh and Wei-Cheng Huang.

Nonparametric estimation and test of conditional Kendall’s tau under semicompeting risks data and truncated data. Journal of Applied Statistics, 42(7):1602–1616, July 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alih:2015:ORT

[3072] Ekele Alih and Hong Choon Ong. An outlier-resistant test for heteroscedasticity in linear models. Journal of Applied Statistics, 42(8):1617–1634, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2015:CCD

[3073] Guanfu Liu, Xiaolong Pu, Lei Wang, and Dongdong Xiang. CUSUM chart for detecting range shifts when monotonicity of likelihood ratio is invalid. Journal of Applied Statistics, 42(8): 1635–1644, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ziane:2015:ABB

[3074] Y. Ziane, S. Adjabi, and N. Zougab.

Adaptive Bayesian bandwidth selection in asymmetric kernel density estimation for nonnegative heavy-tailed data. Journal of Applied Statistics, 42 (8):1645–1658, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vilca:2015:SNI

[3075] Filidor Vilca, Camila Borelli Zeller, and Gauss M. Cordeiro. The sinhnormal/independent nonlinear regression model. Journal of Applied Statistics, 42(8):1659–1676, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sun:2015:EMC

[3076] Meesun Sun, Kwanghyun Choi, and Sungzoon Cho. Estimating the minority class proportion with the ROC curve using Military Personality Inventory data of the ROK Armed Forces. Journal of Applied Statistics, 42(8): 1677–1689, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coleman:2015:MWD

[3077] Deidra A. Coleman, Donald E. K. Martin, and Brian J. Reich. Multiple window discrete scan statistic for higherorder Markovian sequences. Journal of Applied Statistics, 42(8):1690–1705, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsou:2015:THP

[3078] Tsung-Shan Tsou and Hsiao-Yun Liu. Testing the homogeneity of proportions for clustered binary data without knowing the correlation structure. Journal of Applied Statistics, 42(8): 1706–1715, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alkan:2015:CDP

[3079] B. Baris Alkan, Cemal Atakan, and Nesrin Alkan. A comparison of different procedures for principal component analysis in the presence of outliers. Journal of Applied Statistics, 42 (8):1716–1722, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dhhan:2015:NSI

[3080] Waleed Dhhan, Sohel Rana, and Habshah Midi. Non-sparse ϵ-insensitive support vector regression for outlier detection. Journal of Applied Statistics, 42(8):1723–1739, August 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anderson:2015:MCC

[3081] Aaron Anderson. A Monte Carlo comparison of alternative methods of maximum likelihood ranking in racing sports. Journal of Applied Statistics, 42 (8):1740–1756, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hasler:2015:CMC

[3082] Mario Hasler. Comment on multiple comparisons with a control under heteroscedasticity. Journal of Applied Statistics, 42(8):1757–1758, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fernandez-Macho:2015:CTS

[3083] Javier Fernández-Macho. Comment on testing for spurious and cointegrated regressions: a wavelet approach. Journal of Applied Statistics, 42(8): 1759–1769, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Leong:2015:RCT

[3084] Chee Kian Leong. Response to the comment on testing for spurious and cointegrated regressions: a wavelet approach. Journal of Applied Statistics, 42(8):1770–1772, August 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akbilgic:2015:CTA

[3085] Oguz Akbilgic. Classification trees aided mixed regression model. Journal of Applied Statistics, 42(8):1773–1781, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Abd-Elfattah:2015:SVC

[3086] Ehab F. Abd-Elfattah. Saddlepoint p-values for a class of tests for comparing competing risks with censored data. Journal of Applied Statistics, 42 (8):1782–1791, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2015:CTL

[3087] S. P. Singh, S. Mukhopadhyay, and A. Roy. Comparison of three-level cluster randomized trials using quantile dispersion graphs. Journal of Applied Statistics, 42(8):1792–1812, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yoon:2015:SME

[3088] Sanghoo Yoon, Bungon Kumphon, and Jeong-Soo Park. Spatial modeling of extreme rainfall in northeast Thailand. Journal of Applied Statistics, 42(8): 1813–1828, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2015:IEC

[3089] Chien-Lang Chen, Shang-Ling Ou, and Chen-Tuo Liao. Interval estimation for conformance proportions of multiple quality characteristics. Journal of Applied Statistics, 42(8):1829–1841, August 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2015:EIV

[3090] Yuan Liu, Hongyun Liu, Hang Li, and Qian Zhao. The effects of individually varying times of observations on growth parameter estimations in piecewise growth model. Journal of Applied Statistics, 42(9):1843–1860, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xia:2015:MSD

[3091] Y. Xia, N. Lu, I. Katz, R. Bossarte, J. Arora, H. He, J. X. Tu, B. Stephens, A. Watts, and X. M. Tu. Models for surveillance data under reporting delay: applications to US veteran first-time suicide attempters. Journal of Applied Statistics, 42(9): 1861–1876, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2015:SCA

[3092] Hea-Jung Kim. Segmented classification analysis with a class of rectanglescreened elliptical populations. Journal of Applied Statistics, 42(9):1877–1895, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Spagnoli:2015:MEM

[3093] A. Spagnoli, J. J. Houwing-Duistermaat, and M. Alfó. Mixed-effect models for longitudinal responses with different types of dropout: an application to the Leiden 85-plus study. Journal of Applied Statistics, 42(9): 1896–1910, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2015:RFO

[3094] Rabindra Nath Das, Jinseog Kim, and Youngjo Lee. Robust first-order rotatable lifetime improvement experimental designs. Journal of Applied Statistics, 42(9):1911–1930, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Malehi:2015:JML

[3095] Amal Saki Malehi, Ebrahim Hajizadeh, Kambiz A. Ahmadi, and Parvin Mansouri. Joint modelling of longitudinal biomarker and gap time between recurrent events: copula-based dependence. Journal of Applied Statistics, 42 (9):1931–1945, September 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2015:TST

[3096] Seokho Lee, Johan Lim, Insuk Sohn, Sin-Ho Jung, and Cheol-Keun Park. Two sample test for high-dimensional partially paired data. Journal of Applied Statistics, 42(9):1946–1961, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nodehi:2015:DAP

[3097] A. Nodehi, M. Golalizadeh, and A. Heydari. Dihedral angles principal geodesic analysis using nonlinear statistics. Journal of Applied Statistics, 42 (9):1962–1972, September 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Noughabi:2015:ELR

[3098] Hadi Alizadeh Noughabi. Empirical likelihood ratio-based goodness-of-fit test for the logistic distribution. Journal of Applied Statistics, 42(9): 1973–1983, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silva:2015:APP

[3099] A. R. Silva, C. T. S. Dias, P. R. Cecon, and E. R. Rêgo. An alternative procedure for performing a power analysis of Mantel’s test. Journal of Applied Statistics, 42(9):1984–1992, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rivaz:2015:BSP

[3100] Firoozeh Rivaz and Majid Jafari Khaledi. Bayesian spatial prediction of skew and censored data via a hybrid algorithm. Journal of Applied Statistics, 42(9):1993–2009, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Neto:2015:RMI

[3101] Eufrásio de A. Lima Neto and Ulisses U. dos Anjos. Regression model for interval-valued variables based on copulas. Journal of Applied Statistics, 42(9):2010–2029, September 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kozak:2015:EDI

[3102] Marcin Kozak, Wojtek Krzanowski, Izabela Cichocka, and James Hartley. The effects of data input errors on subsequent statistical inference. Journal of Applied Statistics, 42(9):

2030–2037, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Holden:2015:EIH

[3103] Philip B. Holden, Neil R. Edwards, Paul H. Garthwaite, and Richard D. Wilkinson. Emulation and interpretation of high-dimensional climate model outputs. Journal of Applied Statistics, 42(9):2038–2055, September 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2015:REN

[3104] Abhik Ghosh and Ayanendranath Basu. Robust estimation for nonhomogeneous data and the selection of the optimal tuning parameter: the density power divergence approach. Journal of Applied Statistics, 42(9): 2056–2072, September 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2015:GAD

[3105] Myung Geun Kim. Geometric aspects of deletion diagnostics in multivariate regression. Journal of Applied Statistics, 42(10):2073–2079, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khan:2015:DSS

[3106] M. G. M. Khan, K. G. Reddy, and D. K. Rao. Designing stratified sampling in economic and business surveys. Journal of Applied Statistics, 42(10): 2080–2099, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bandyopadhyay:2015:CDC

[3107] Uttam Bandyopadhyay, Joydeep Basu, and Ganesh Dutta. Crossover design in clinical trials for binary response. Journal of Applied Statistics, 42(10): 2100–2114, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ding:2015:RGT

[3108] Juan Ding and Wenjun Xiong. Robust group testing for multiple traits with misclassification. Journal of Applied Statistics, 42(10):2115–2125, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Keser:2015:SFC

[3109] Istem Koymen Keser and Ipek Deveci Kocakoç. Smoothed functional canonical correlation analysis of humidity and temperature data. Journal of Applied Statistics, 42(10):2126–2140, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Modugno:2015:MMA

[3110] Lucia Modugno, Silvia Cagnone, and Simone Giannerini. A multilevel model with autoregressive components for the analysis of tribal art prices. Journal of Applied Statistics, 42(10):2141–2158, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ortega:2015:NFM

[3111] Edwin M. M. Ortega, Artur J.

Lemonte, Giovana O. Silva, and Gauss M. Cordeiro. New flexible models generated by gamma random variables for lifetime modeling. Journal of Applied Statistics, 42(10):2159–2179, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghebremichael:2015:JMC

[3112] Musie Ghebremichael. Joint modeling of correlated binary outcomes: HIV-1 and HSV-2 co-infection. Journal of Applied Statistics, 42(10):2180–2191, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DAmbra:2015:GAD

[3113] Antonello D’Ambra, Anna Crisci, and Pasquale Sarnacchiaro. A generalized analysis of the dependence structure by means of ANOVA. Journal of Applied Statistics, 42(10):2192–2202, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

He:2015:DFI

[3114] H. He, W. Wang, J. Hu, R. Gallop, P. Crits-Christoph, and Y. Xia. Distribution-free inference of zeroinflated binomial data for longitudinal studies. Journal of Applied Statistics, 42(10):2203–2219, October 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hung:2015:ICA

[3115] Wen-Liang Hung, Shou-Jen Chang-Chien, and Miin-Shen Yang. An intuitive clustering algorithm for spherical data with application to extrasolar planets. Journal of Applied Statistics, 42(10):2220–2232, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Teimourian:2015:JMM

[3116] M. Teimourian, T. Baghfalaki, M. Ganjali, and D. Berridge. Joint modeling of mixed skewed continuous and ordinal longitudinal responses: a Bayesian approach. Journal of Applied Statistics, 42(10):2233–2256, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Donneau:2015:TPO

[3117] A. F. Donneau, M. Mauer, P. Lambert, E. Lesaffre, and A. Albert. Testing the proportional odds assumption in multiply imputed ordinal longitudinal data. Journal of Applied Statistics, 42(10):2257–2279, October 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kemalbay:2015:JDN

[3118] Gulder Kemalbay and Ismihan Bayramoğlu (Bairamov). Joint distribution of new sample rank of bivariate order statistics. Journal of Applied Statistics, 42 (10):2280–2289, October 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ober:2015:SAH

[3119] Pieter Bastiaan Ober. Sequential analysis: hypothesis testing and changepoint detection. Journal of Applied Statistics, 42(10):2290, October 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Popovic:2015:BRH

[3120] Bozidar V. Popović. Book review:

Handbook of univariate and multivariate data analysis with IBM SPSS, second edition. Journal of Applied Statistics, 42(10):2291, October 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balci:2015:RPM

[3121] Sibel Balci and Aysen Dener Akkaya. Robust pairwise multiple comparisons under short-tailed symmetric distributions. Journal of Applied Statistics, 42 (11):2293–2306, November 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1023706.

Chen:2015:CTP

[3122] Wei Chen, Dehui Wang, and Yanfeng Li. A class of tests of proportional hazards assumption for lefttruncated and right-censored data. Journal of Applied Statistics, 42(11): 2307–2320, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1027884.

Nyamundanda:2015:PPL

[3123] Gift Nyamundanda, Avril Hegarty, and Kevin Hayes. Product partition latent variable model for multiple changepoint detection in multivariate data. Journal of Applied Statistics, 42(11): 2321–2334, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1029444.

Pigeot:2015:USG

[3124] Iris Pigeot, Fabian Sobotka, Svend Kreiner, and Ronja Foraita. The uncertainty of a selected graphical model. Journal of Applied Statistics, 42(11):

2335–2352, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1030368.

Louzada:2015:PEM

[3125] Francisco Louzada, Márcia A. C.

Macera, and Vicente G. Cancho. The Poisson-exponential model for recurrent event data: an application to bowel motility data. Journal of Applied Statistics, 42(11):2353– 2366, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1030369.

Shortridge:2015:CAH

[3126] Julie E. Shortridge, Stefanie M. Falconi, Benjamin F. Zaitchik, and Seth D. Guikema. Climate, agriculture, and hunger: statistical prediction of undernourishment using nonlinear regression and data-mining techniques. Journal of Applied Statistics, 42(11): 2367–2390, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1032216.

Ekin:2015:OMM

[3127] Tahir Ekin, R. Muzaffer Musal, and Lawrence V. Fulton. Overpayment models for medical audits: multiple scenarios. Journal of Applied Statistics, 42(11):2391–2405, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1034659.

Chen:2015:EST

[3128] Haiqiang Chen and Yanli Zhu. An empirical study on the threshold cointegration of Chinese A and H cross-listed shares. Journal of Applied Statistics, 42 (11):2406–2419, November 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1034660.

Lin:2015:DCI

[3129] Chen ju Lin and Yi chun Shu. Detecting clusters with increased mean using scan windows with variable radius. Journal of Applied Statistics, 42(11): 2420–2431, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1041013.

Lin:2015:HDV

[3130] Jin-Guan Lin, Yan-Yong Zhao, and Hong-Xia Wang. Heteroscedasticity diagnostics in varying-coefficient partially linear regression models and applications in analyzing Boston housing data. Journal of Applied Statistics, 42 (11):2432–2448, November 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1043623.

Oyedele:2015:CPL

[3131] Opeoluwa F. Oyedele and Sugnet Lubbe. The construction of a partial least-squares biplot. Journal of Applied Statistics, 42(11):2449– 2460, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043858.

Zhao:2015:SML

[3132] Xiaobing Zhao and Xian Zhou. Semiparametric models of longitudinal and time-to-event data with applications to HIV viral dynamics and CD4 counts. Journal of Applied Statistics, 42(11): 2461–2477, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043859.

Varghese:2015:EDO

[3133] Cini Varghese, Eldho Varghese, Seema Jaggi, and Arpan Bhowmik. Experimental designs for open pollination in polycross trials. Journal of Applied Statistics, 42(11):2478– 2484, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043860.

Aslam:2015:NCC

[3134] Muhammad Aslam, Nasrullah Khan, and Chi-Hyuck Jun. A new S^{2} control chart using repetitive sampling. Journal of Applied Statistics, 42(11): 2485–2496, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043861.

Jiang:2015:REP

[3135] Yunlu Jiang. Robust estimation in partially linear regression models. Journal of Applied Statistics, 42(11): 2497–2508, November 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043862.

Sani:2015:RRA

[3136] A. Sani, B. Abapihi, Mukhsar Mukhsar, and Kadir Kadir. Relative risk analysis of dengue cases using convolution extended into spatio-temporal model. Journal of Applied Statistics, 42 (11):2509–2519, November 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1043863.

Hadidoust:2015:MCP

[3137] Zahra Hadidoust, Yaser Samimi, and Hamid Shahriari. Monitoring and change-point estimation for splinemodeled non-linear profiles in phase II. Journal of Applied Statistics, 42(12): 2520–2530, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043864.

Benoit:2015:JCP

[3138] Julia S. Benoit, Wenyaw Chan, and Rachelle S. Doody. Joint coverage probability in a simulation study on continuous-time Markov chain parameter estimation. Journal of Applied Statistics, 42(12):2531– 2538, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043865.

Martinez-Camblor:2015:ERT

[3139] Pablo Martínez-Camblor, Jacobo de Uña-Álvarez, and Carmen Díaz Corte. Expanded renal transplantation: a competing risk model approach. Journal of Applied Statistics, 42(12): 2539–2553, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1043866.

Raffinetti:2015:DBE

[3140] E. Raffinetti and I. Romeo. Dealing with the biased effects issue when handling huge datasets: the case of IN-VALSI data. Journal of Applied Statistics, 42(12):2554–2570, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1043867.

Linder:2015:SBR

[3141] Daniel F. Linder, Hani Samawi, Lili Yu, Arpita Chatterjee, Yisong Huang, and Robert Vogel. On stratified bivariate ranked set sampling for regression estimators. Journal of Applied Statistics, 42(12):2571–2583, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1043868.

Amerise:2015:CMT

[3142] Ilaria L. Amerise and Agostino Tarsitano. Correction methods for ties in rank correlations. Journal of Applied Statistics, 42(12):2584– 2596, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1043870.

Mansourvar:2015:SRR

[3143] Zahra Mansourvar, Torben Martinussen, and Thomas H. Scheike. Semiparametric regression for restricted mean residual life under right censoring. Journal of Applied Statistics, 42 (12):2597–2613, December 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1043871.

Liang:2015:NTM

[3144] Rubing Liang, Cuizhen Niu, Qiang Xia, and Zhiqiang Zhang. Nonlinearity testing and modeling for threshold moving average models. Journal of Applied Statistics, 42(12):2614– 2630, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1043872.

Lucadamo:2015:PHO

[3145] Antonio Lucadamo and Pietro Amenta. A proposal for handling ordinal categorical variables in co-inertia analysis. Journal of Applied Statistics, 42(12): 2631–2638, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1044426.

Plaia:2015:LTE

[3146] Antonella Plaia. Long-term experiments and strip plot designs. Journal of Applied Statistics, 42(12):2639– 2653, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1046821.

Nekoukhou:2015:NFP

[3147] Vahid Nekoukhou and Hamid Bidram.

A new four-parameter discrete distribution with bathtub and unimodal failure rate. Journal of Applied Statistics, 42(12):2654–2670, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1046822.

Pundir:2015:DDA

[3148] Sudesh Pundir and R. Amala. Detecting diagnostic accuracy of two biomarkers through a bivariate log-normal ROC curve. Journal of Applied Statistics, 42(12):2671–2685, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1046823.

Vitanov:2015:TTH

[3149] Nikolay K. Vitanov and Marcel Ausloos. Test of two hypotheses explaining the size of populations in a system of cities. Journal of Applied Statistics, 42 (12):2686–2693, December 2015. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1047744.

Garay:2015:BAC

[3150] Aldo M. Garay, Heleno Bolfarine, Victor H. Lachos, and Celso R. B. Cabral. Bayesian analysis of censored linear regression models with scale mixtures of normal distributions.

Journal of Applied Statistics, 42(12): 2694–2714, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1048671.

Lu:2015:RCI

[3151] H. Lu, P. Yin, R. X. Yue, and J. Q. Shi. Robust confidence intervals for trend estimation in meta-analysis with publication bias. Journal of Applied Statistics, 42(12):2715–2733, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1048672.

Hutson:2015:MBD

[3152] Alan D. Hutson, Gregory E. Wilding, Terry L. Mashtare, and Albert Vexler. Measures of biomarker dependence using a copula-based multivariate epsilon-skew-normal family of distributions. Journal of Applied Statistics, 42(12):2734–2753, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1049130.

Chakrabarty:2015:HBS

[3153] Manisha Chakrabarty, Amita Majumder, and Jeffrey Racine. Household budget-share distributions and welfare implications: an application of multivariate distributional statistics. Journal of Applied Statistics, 42(12): 2754–2768, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1049132.

Neuhauser:2015:CTW

[3154] Markus Neuhäuser. Combining the t test and Wilcoxon’s rank-sum test. Journal of Applied Statistics, 42(12): 2769–2775, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1070809.

Anonymous:2015:EB

[3155] Anonymous. Editorial board. Journal of Applied Statistics, 42(12): ebi, December 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1119986.

Attanasio:2016:SE

[3156] Massimo Attanasio and Vincenza Capursi. Statistics in education. Journal of Applied Statistics, 43(1):1– 2, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1104890.

Adelfio:2016:DCC

[3157] Giada Adelfio and Giovanni Boscaino.

Degree course change and student performance: a mixed-effect model approach. Journal of Applied Statistics, 43(1):3–15, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1018673.

Crippa:2016:DFA

[3158] F. Crippa, M. Mazzoleni, and M. Zenga. Departures from the formal of actual students’ university careers: an application of nonhomogeneous fuzzy Markov chains. Journal of Applied Statistics, 43 (1):16–30, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1091446.

Raponi:2016:BAU

[3159] Valentina Raponi, Francesca Martella, and Antonello Maruotti. A biclustering approach to university performances: an Italian case study. Journal of Applied Statistics, 43 (1):31–45, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1009005.

Enea:2016:AMB

[3160] Marco Enea and Massimo Attanasio. An association model for bivariate data with application to the analysis of university students’ success. Journal of Applied Statistics, 43(1):46–57, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2014.998407.

Dehbi:2016:AOQ

[3161] Hakim-Moulay Dehbi, Mario Cortina-Borja, and Marco Geraci. Aranda– Ordaz quantile regression for student performance assessment. Journal of Applied Statistics, 43(1):58– 71, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1025724.

Maia:2016:APS

[3162] Rafael Pimentel Maia, Hildete Prisco Pinheiro, and Aluísio Pinheiro. Academic performance of students from entrance to graduation via quasi Ustatistics: a study at a Brazilian research university. Journal of Applied Statistics, 43(1):72–86, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1077939.

Ratnaningsih:2016:CAC

[3163] Dewi Juliah Ratnaningsih and Imas Sukaesih Sitanggang. Comparative analysis of classification methods in determining non-active student characteristics in Indonesia Open University. Journal of Applied Statistics, 43(1):87–97, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1077940.

Contini:2016:PDI

[3164] Dalit Contini and Davide Azzolini. Performance and decisions: immigrantnative gaps in educational transitions in Italy. Journal of Applied Statistics, 43(1):98–114, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1036845.

Vitale:2016:EES

[3165] Maria Prosperina Vitale, Giovanni C.

Porzio, and Patrick Doreian. Examining the effect of social influence on student performance through network autocorrelation models. Journal of Applied Statistics, 43(1):115– 127, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1049517.

Monyai:2016:AML

[3166] Simon Monyai, ’Maseka Lesaoana, Timotheus Darikwa, and Philimon Nyamugure. Application of multinomial logistic regression to educational factors of the 2009 General Household Survey in South Africa. Journal of Applied Statistics, 43(1): 128–139, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1077941.

Castellano:2016:PED

[3167] Rosalia Castellano and Gennaro Punzo. Patterns of earnings differentials across three conservative European welfare regimes with alternative education systems. Journal of Applied Statistics, 43(1):140– 168, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1049518.

Capecchi:2016:IDJ

[3168] Stefania Capecchi and Domenico Piccolo. Investigating the determinants of job satisfaction of Italian graduates: a model-based approach. Journal of Applied Statistics, 43(1): 169–179, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1036844.

Fasola:2016:FLT

[3169] S. Fasola, O. Giambalvo, and C. Romano. Flexible latent trait aggregation to analyze employability after the Ph.D. in Italy. Journal of Applied Statistics, 43(1):180– 194, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1077797.

Fouskakis:2016:BHM

[3170] D. Fouskakis, G. Petrakos, and I. Vavouras. A Bayesian hierarchical model for comparative evaluation of teaching quality indicators in higher education. Journal of Applied Statistics, 43(1):195–211, January 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1054793.

Telcs:2016:UOD

[3171] András Telcs, Zsolt Tibor Kosztyán, and Ádám Török. Unbiased onedimensional university ranking — application-based preference ordering. Journal of Applied Statistics, 43(1): 212–228, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2014.998180.

Gross:2016:CDM

[3172] J. Groß, A. Robitzsch, and A. C.

George. Cognitive diagnosis models for baseline testing of educational standards in math. Journal of Applied Statistics, 43(1):229– 243, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2014.1000841.

Adegboye:2016:MMM

[3173] Oyelola A. Adegboye and Asadullah Jawid. Multivariate multilevel models for attitudes toward statistics: multidisciplinary settings in Afghanistan. Journal of Applied Statistics, 43(1): 244–261, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1091445.

DeStefano:2016:CAN

[3174] Domenico De Stefano and Susanna Zaccarin. Co-authorship networks and scientific performance: an empirical analysis using the generalized extreme value distribution. Journal of Applied Statistics, 43(1):262– 279, January 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1017719.

Meshkani:2016:BAC

[3175] Mohammadreza Meshkani, Afshin Fallah, and Amir Kavousi. Bayesian analysis of covariance under inverse Gaussian model. Journal of Applied Statistics, 43(2):280–298, February 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1049131.

Toffoli:2016:EOI

[3176] Sonia Ferreira Lopes Toffoli, Dalton Francisco de Andrade, and Antonio Cezar Bornia. Evaluation of open items using the many-facet Rasch model. Journal of Applied Statistics, 43 (2):299–316, February 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1049938.

Vinod:2016:NBI

[3177] H. D. Vinod. New bootstrap inference for spurious regression problems. Journal of Applied Statistics, 43(2): 317–335, February 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1049939.

Saha:2016:TTI

[3178] Chandan Saha and Michael P. Jones.

Type I and Type II error rates in the last observation carried forward method under informative dropout. Journal of Applied Statistics, 43(2): 336–350, February 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1063112.

Ambach:2016:STW

[3179] Daniel Ambach. Short-term wind speed forecasting in Germany. Journal of Applied Statistics, 43(2):351– 369, February 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1063113.

Karaoglan:2016:NPP

[3180] Aslan Deniz Karaoglan and Nihat Celik. A new painting process for vessel radiators of transformer: wet-on-wet. Journal of Applied Statistics, 43 (2):370–386, February 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1063114.

Baker:2016:EBP

[3181] Rose D. Baker and Ian G. McHale. An empirical Bayes’ procedure for ranking players in Ryder Cup golf. Journal of Applied Statistics, 43(3): 387–395, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1043869.

Choy:2016:RBA

[3182] S. T. Boris Choy, Jennifer S. K. Chan, and Udi E. Makov. Robust Bayesian analysis of loss reserving data using scale mixtures distributions. Journal of Applied Statistics, 43(3):396–411, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1063115.

Noughabi:2016:TGF

[3183] Hadi Alizadeh Noughabi and Narayanaswamy Balakrishnan. Tests of goodness of fit based on Phi-divergence. Journal of Applied Statistics, 43(3):412– 429, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1063116.

Armero:2016:BLM

[3184] C. Armero, A. Forte, and H. Perpiñán.

Bayesian longitudinal models for paediatric kidney transplant recipients. Journal of Applied Statistics, 43(3): 430–440, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1063117.

Li:2016:IMF

[3185] Qing Li. Indirect membership function assignment based on ordinal regression. Journal of Applied Statistics, 43(3):441–460, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1070802.

Li:2016:TLN

[3186] Chin-Shang Li. A test for the linearity of the nonparametric part of a semiparametric logistic regression model. Journal of Applied Statistics, 43(3):461–475, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1070803.

Emami:2016:NID

[3187] Hadi Emami and Mostafa Emami. New influence diagnostics in ridge regression. Journal of Applied Statistics, 43(3):476–489, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1070804.

Tyssedal:2016:FSN

[3188] John Tyssedal and Shahrukh Hussain. Factor screening in nonregular twolevel designs based on projection-based variable selection. Journal of Applied Statistics, 43(3):490–508, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1070805.

Nurunnabi:2016:ICM

[3189] A. A. M. Nurunnabi, M. Nasser, and A. H. M. R. Imon. Identification and classification of multiple outliers, high leverage points and influential observations in linear regression. Journal of Applied Statistics, 43(3):509–525, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1070806.

Ma:2016:VNN

[3190] Xuejun Ma, Xiaoqun He, and Xiaokang Shi. A variant of K nearest neighbor quantile regression. Journal of Applied Statistics, 43(3):526–537, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1070807.

Karagulle:2016:TDE

[3191] S. Karagulle and Z. Kalaylioglu. A test for detecting etiologic heterogeneity in epidemiological studies. Journal of Applied Statistics, 43 (3):538–549, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1070808.

Shih:2016:NIT

[3192] Shu-Man Shih, Wei-Hwa Wu, and Hsin-Neng Hsieh. A non-inferiority test for diagnostic accuracy in the absence of the golden standard test based on the paired partial areas under receiver operating characteristic curves. Journal of Applied Statistics, 43(3):550–562, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1070810.

Alavi:2016:MLE

[3193] Sayed Mohammad Reza Alavi and Mahboobeh Tajodini. Maximum likelihood estimation of sensitive proportion using repeated randomized response techniques. Journal of Applied Statistics, 43(3):563–571, March 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1070811.

Cancho:2016:UMS

[3194] Vicente G. Cancho, Dipak K. Dey, and Francisco Louzada. Unified multivariate survival model with a surviving fraction: an application to a Brazilian customer churn data. Journal of Applied Statistics, 43(3): 572–584, March 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1071341.

Wenren:2016:CCP

[3195] Cheng Wenren and Junfeng Shang. Conditional conceptual predictive statistic for mixed model selection. Journal of Applied Statistics, 43 (4):585–603, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1071342.

Otranto:2016:SED

[3196] Edoardo Otranto, Massimo Mucciardi, and Pietro Bertuccelli. Spatial effects in dynamic conditional correlations. Journal of Applied Statistics, 43(4):604–626, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1071343.

Leiva:2016:DER

[3197] Víctor Leiva, Shuangzhe Liu, Lei Shi, and Francisco José A. Cysneiros. Diagnostics in elliptical regression models with stochastic restrictions applied to econometrics. Journal of Applied Statistics, 43(4):627–642, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1072140.

Zhang:2016:QAV

[3198] Junying Zhang, Riquan Zhang, and Zhiping Lu. Quantile-adaptive variable screening in ultra-high dimensional varying coefficient models. Journal of Applied Statistics, 43(4):643– 654, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1072141.

Bogdanova:2016:WBA

[3199] Boryana Bogdanova and Ivan Ivanov.

A wavelet-based approach to the analysis and modelling of financial time series exhibiting strong long-range dependence: the case of Southeast Europe. Journal of Applied Statistics, 43(4):655–673, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1077370.

Schepers:2016:RMD

[3200] Jan Schepers. On regression modelling with dummy variables versus separate regressions per group: Comment on Holgersson et al. Journal of Applied Statistics, 43(4):674–681, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1077371.

Zinn:2016:SAA

[3201] S. Zinn and A. Würbach. A statistical approach to address the problem of heaping in self-reported income data. Journal of Applied Statistics, 43(4):682–703, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1077372.

Zhang:2016:DLT

[3202] Yue Zhang and Kiros Berhane. Dynamic latent trait models with mixed hidden Markov structure for mixed longitudinal outcomes. Journal of Applied Statistics, 43(4):704–720, April 2016.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1077373.

Christen:2016:BDO

[3203] J. Andrés Christen, Bruno Sansó, Mario Santana-Cibrian, and Jorge X. Velasco-Hernández. Bayesian deconvolution of oil well test data using Gaussian processes. Journal of Applied Statistics, 43(4):721–737, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1077374.

Kuvvetli:2016:DSS

[3204] Ümit Kuvvetli, Ali Riza Firuzan, Süleyman Alpaykut, and Atakan Gerger. Determining Six Sigma success factors in Turkey by using structural equation modeling. Journal of Applied Statistics, 43(4):738–753, April 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1077375.

Beccarini:2016:BCT

[3205] Andrea Beccarini. Bias correction through filtering omitted variables and instruments. Journal of Applied Statistics, 43(4):754–766, April 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1077376.

Ma:2016:MVI

[3206] Xiaofei Ma and Qiuyan Zhong. Missing value imputation method for disaster decision-making using K nearest neighbor. Journal of Applied Statistics, 43(4):767–781, April 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1077377.

Chen:2016:DDD

[3207] Hsiu-Wen Chen, Weng Kee Wong, and Hongquan Xu. Data-driven desirability function to measure patients’ disease progression in a longitudinal study. Journal of Applied Statistics, 43(5):783–795, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1077378.

Wang:2016:VSH

[3208] Xiuli Wang and Mingqiu Wang. Variable selection for high-dimensional generalized linear models with the weighted elastic-net procedure. Journal of Applied Statistics, 43(5):796– 809, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1078300.

Nores:2016:BHT

[3209] M. L. Nores and M. P. Díaz. Bootstrap hypothesis testing in generalized additive models for comparing curves of treatments in longitudinal studies. Journal of Applied Statistics, 43(5):810–826, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1078301.

Lyu:2016:TNP

[3210] Meng-Ning Lyu, Qing-Shan Yang, Na Yang, and Siu-Seong Law. Tourist number prediction of historic buildings by singular spectrum analysis. Journal of Applied Statistics, 43(5):827–846, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1078302.

Costa:2016:MPI

[3211] Leonardo Costa, Adrian Pizzinga, and Rodrigo Atherino. Modeling and predicting IBNR reserve: extended chain ladder and heteroscedastic regression analysis. Journal of Applied Statistics, 43(5):847–870, May 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1079305.

Degiannakis:2016:ERV

[3212] Stavros Degiannakis and Alexandra Livada. Evaluation of realized volatility predictions from models with leptokurtically and asymmetrically distributed forecast errors. Journal of Applied Statistics, 43(5):871–892, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1079306.

Aghdam:2016:IGR

[3213] Rosa Aghdam, Mojtaba Ganjali, Parisa Niloofar, and Changiz Eslahchi. Inferring gene regulatory networks by an order independent algorithm using incomplete data sets. Journal of Applied Statistics, 43 (5):893–913, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1079307.

Baycan:2016:EER

[3214] Ismail Onur Baycan. The effects of exchange rate regimes on economic growth: evidence from propensity score matching estimates. Journal of Applied Statistics, 43(5):914–924, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1080669.

Asar:2016:FOM

[3215] Özgür Asar and Ozlem Ilk. Firstorder marginalised transition random effects models with probit link function. Journal of Applied Statistics, 43(5):925–942, May 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1080670.

Zhou:2016:RIV

[3216] Diwei Zhou, Ian L. Dryden, Alexey A.

Koloydenko, Koenraad M. R. Audenaert, and Li Bai. Regularisation, interpolation and visualisation of diffusion tensor images using non-Euclidean statistics. Journal of Applied Statistics, 43(5):943–978, May 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1080671.

Chen:2016:PACa

[3217] T. Chen, K. Knox, J. Arora, W. Tang, J. Kowalski, and X. M. Tu. Power analysis for clustered non-continuous responses in multicenter trials. Journal of Applied Statistics, 43(6):979– 995, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1089218.

Shen:2016:EAP

[3218] Pao sheng Shen. Estimation of association parameters in copula models for bivariate left-truncated and rightcensored data. Journal of Applied Statistics, 43(6):996–1010, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1089219.

Shrestha:2016:IUQ

[3219] Rijan Shrestha and Tomasz Kozlowski. Inverse uncertainty quantification of input model parameters for thermal-hydraulics simulations using expectation-maximization under Bayesian framework. Journal of Applied Statistics, 43(6):1011– 1026, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1089220.

Yiqi:2016:IDW

[3220] Bao Yiqi, Cibele Maria Russo, Vicente G. Cancho, and Francisco Louzada. Influence diagnostics for the Weibull-negative-binomial regression model with cure rate under latent failure causes. Journal of Applied Statistics, 43(6):1027–1060, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1089221.

Lee:2016:CPB

[3221] Sang Eun Lee and Key-Il Shin.

The cut-off point based on underlying distribution and cost function. Journal of Applied Statistics, 43(6): 1061–1073, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1089222.

Yan:2016:MPM

[3222] Guohua Yan, M. Tariqul Hasan, and Renjun Ma. Modeling proportions and marginal counts simultaneously for clustered multinomial data with random cluster sizes. Journal of Applied Statistics, 43(6):1074–1087, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1089223.

Martin-Barragan:2016:FBB

[3223] B. Martin-Barragan, R. E. Lillo, and J. Romo. Functional boxplots based on epigraphs and hypographs. Journal of Applied Statistics, 43(6): 1088–1103, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1092108.

Chen:2016:PACb

[3224] T. Chen, N. Lu, J. Arora, I. Katz, R. Bossarte, H. He, Y. Xia, H. Zhang, and X. M. Tu. Power analysis for cluster randomized trials with binary outcomes modeled by generalized linear mixed-effects models. Journal of Applied Statistics, 43(6): 1104–1118, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1092109.

Maansson:2016:RLE

[3225] Kristofer Månsson, B. M. Golam Kibria, and Ghazi Shukur. A restricted Liu estimator for binary regression models and its application to an applied demand system. Journal of Applied Statistics, 43(6):1119–1127, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1092110.

Ghosh:2016:IFT

[3226] Himadri Ghosh, S. Chowdhury, and Prajneshu. An improved fuzzy timeseries method of forecasting based on L–R fuzzy sets and its application. Journal of Applied Statistics, 43 (6):1128–1139, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1092111.

Pehlivanli:2016:NFS

[3227] Ayça Çakmak Pehlivanli. A novel feature selection scheme for highdimensional data sets: Four-Staged Feature Selection. Journal of Applied Statistics, 43(6):1140–1154, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1092112.

Saraiva:2016:PGE

[3228] E. F. Saraiva, A. K. Suzuki, F. Louzada, and L. A. Milan. Partitioning gene expression data by data-driven Markov chain Monte Carlo. Journal of Applied Statistics, 43(6):1155–1173, June 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1092113.

Dagne:2016:GMT

[3229] Getachew A. Dagne. A growth mixture tobit model: application to AIDS studies. Journal of Applied Statistics, 43(7):1174–1185, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1092114.

Voigt:2016:DBV

[3230] Tobias Voigt, Roland Fried, Wolfgang Rhode, and Fabian Temme. Distancebased variable generation with applications to the FACT experiment. Journal of Applied Statistics, 43(7): 1186–1197, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1092115.

Melo:2016:GLS

[3231] Oscar O. Melo, Jorge Mateu, and Carlos E. Melo. A generalised linear space–time autoregressive model with space–time autoregressive disturbances. Journal of Applied Statistics, 43(7):1198–1225, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1092506.

Bocci:2016:MLD

[3232] Chiara Bocci and Emilia Rocco. Modelling the location decisions of manufacturing firms with a spatial point process approach. Journal of Applied Statistics, 43(7):1226–1239, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1093612.

Tian:2016:CFM

[3233] Yuzhu Tian, Manlai Tang, and Maozai Tian. A class of finite mixture of quantile regressions with its applications. Journal of Applied Statistics, 43(7):1240–1252, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1094035.

Caudill:2016:ECS

[3234] Steven B. Caudill and Franklin G. Mixon Jr. Estimating class-specific parametric models using finite mixtures: an application to a hedonic model of wine prices. Journal of Applied Statistics, 43(7):1253– 1261, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1094036.

Ajadi:2016:ISM

[3235] Jimoh Olawale Ajadi, Muhammad Riaz, and Khalid Al-Ghamdi. On increasing the sensitivity of mixed EWMA–CUSUM control charts for location parameter. Journal of Applied Statistics, 43(7):1262–1278, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1094453.

Shan:2016:ECI

[3236] Guogen Shan. Exact confidence intervals for randomized response strategies. Journal of Applied Statistics, 43 (7):1279–1290, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1094454.

Roth:2016:TSR

[3237] M. Roth, G. Jongbloed, and T. A.

Buishand. Threshold selection for regional peaks-over-threshold data. Journal of Applied Statistics, 43(7): 1291–1309, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
02664763.2015.1100589.

Vieira:2016:IIC

[3238] Marcel de Toledo Vieira, Maria de Fátima Salgueiro, and Peter W. F. Smith. Investigating impacts of complex sampling on latent growth curve modelling. Journal of Applied Statistics, 43(7):1310–1321, July 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/02664763.2015.1100590.

Bayoud:2016:TST

[3239] Husam Awni Bayoud. Testing the similarity of two normal populations with application to the bioequivalence problem. Journal of Applied Statistics, 43 (7):1322–1334, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1100591.

Rivaz:2016:OND

[3240] Firoozeh Rivaz. Optimal network design for Bayesian spatial prediction of multivariate non-Gaussian environmental data. Journal of Applied Statistics, 43(7):1335–1348, July 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http:/

                                        /www.tandfonline.com/doi/abs/10.
                                        1080/02664763.2015.1100592.

Lasek:2016:HIT

[3241] Jan Lasek, Zoltán Szlávik, Marek Gagolewski, and Sandjai Bhulai. How to improve a team’s position in the FIFA ranking? A simulation study. Journal of Applied Statistics, 43(7): 1349–1368, July 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). URL http://www.

                                        tandfonline.com/doi/abs/10.1080/
                                        02664763.2015.1100593.

Alhaji:2016:BAM

[3242] Baba B. Alhaji, Hongsheng Dai, Yoshiko Hayashi, Veronica Vinciotti, Andrew Harrison, and Berthold Lausen. Bayesian analysis for mixtures of discrete distributions with a nonparametric component. Journal of Applied Statistics, 43(8):1369–1385, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Telmoudi:2016:CRE

[3243] Fedya Telmoudi, Mohamed El Ghourabi, and Mohamed Limam. On conditional risk estimation considering model risk.

Journal of Applied Statistics, 43(8): 1386–1399, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ershadi:2016:ESD

[3244] M. J. Ershadi, R. Noorossana, and S. T. A. Niaki. Economic-statistical design of simple linear profiles with variable sampling interval. Journal of Applied Statistics, 43(8):1400–1418, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Machalova:2016:PCL

[3245] J. Machalová, K. Hron, and G. S.

Monti. Preprocessing of centred logratio transformed density functions using smoothing splines. Journal of Applied Statistics, 43(8):1419–1435, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ka:2016:RGE

[3246] Ndéné Ka and Stéphane Mussard. l_{1} regressions: Gini estimators for fixed effects panel data. Journal of Applied Statistics, 43(8):1436–1446, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jaffa:2016:MGL

[3247] Miran A. Jaffa, Mulugeta Gebregziabher, Deirdre K. Luttrell, Louis M. Luttrell, and Ayad A. Jaffa. Multivariate generalized linear mixed models with random intercepts to analyze cardiovascular risk markers in Type-1 diabetic patients. Journal of Applied Statistics, 43(8):1447–1464, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Robbiano:2016:IPP

[3248] Sylvain Robbiano, Matthieu Saumard, and Michel Curé. Improving prediction performance of stellar parameters using functional models. Journal of Applied Statistics, 43(8):1465–1476, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roy:2016:BOA

[3249] Soumya Roy and Chiranjit Mukhopadhyay. Bayesian D-optimal accelerated life test plans for series systems with competing exponential causes of failure. Journal of Applied Statistics, 43 (8):1477–1493, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cebula:2016:TID

[3250] Richard J. Cebula, Fiorentina Angjellari-Dajci, and Russell Kashian. Are there interregional differences in the response of cigarette smoking to state cigarette excise taxes in the USA? Exploratory analysis. Journal of Applied Statistics, 43(8):1494–1507, August 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rondon:2016:BAG

[3251] Luz Marina Rondon and Heleno Bolfarine. Bayesian analysis of generalized elliptical semi-parametric models. Journal of Applied Statistics, 43 (8):1508–1524, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2016:PEE

[3252] Junying Chen, Haoyu Zeng, and Fei Yang. Parameter estimation for employee stock ownerships preference experimental design. Journal of Applied Statistics, 43(8):1525–1540, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pallmann:2016:AMG

[3253] Philip Pallmann and Ludwig A.

Hothorn. Analysis of means: a generalized approach using R. Journal of Applied Statistics, 43(8):1541–1560, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bozza:2016:PLR

[3254] Silvia Bozza and Franco Taroni. Posterior likelihood ratios for evaluation of forensic trace evidence given a two-level model on the data by Alberink et al. (2013). Journal of Applied Statistics, 43(8):1561–1563, August 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Holgersson:2016:RMD

[3255] Thomas Holgersson, Louise Nordström, and Özge Öner. On regression modelling with dummy variables versus separate regressions per group: comment on Holgersson et al. Journal of Applied Statistics, 43(8):1564–1565, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alawadhi:2016:BCP

[3256] F. A. Alawadhi and D. Alhulail.

Bayesian change points analysis for earthquakes body wave magnitude. Journal of Applied Statistics, 43(9): 1567–1582, September 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

UlHassan:2016:FPD

[3257] Mahmood Ul Hassan and Pär Stockhammar. Fitting probability distributions to economic growth: a maximum likelihood approach. Journal of Applied Statistics, 43(9):1583–1603, September 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ludwig:2016:UFT

[3258] Alexander Ludwig. On the usability of the fluctuation test statistic to identify multiple cointegration break points. Journal of Applied Statistics, 43(9):1604–1624, September 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hand:2016:BAT

[3259] Austin L. Hand, John A. Scott, Phil D. Young, James D. Stamey, and Dean M. Young. Bayesian adaptive two-stage design for determining person-time in Phase II clinical trials with Poisson data. Journal of Applied Statistics, 43 (9):1625–1635, September 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Achcar:2016:BDB

[3260] J. A. Achcar, N. Davarzani, and R. M. Souza. Basu-dhar bivariate geometric distribution in the presence of covariates and censored data: a Bayesian approach. Journal of Applied Statistics, 43(9):1636–1648, September 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2016:ACS

[3261] Zheng Xu. An alternative circular smoothing method to nonparametric estimation of periodic functions.

Journal of Applied Statistics, 43(9): 1649–1672, September 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhu:2016:ABT

[3262] Hongjian Zhu, Dejian Lai, and Nils P.

Johnson. Agreement between two diagnostic tests when accounting for testretest variation: application to FFR versus iFR. Journal of Applied Statistics, 43(9):1673–1689, September 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2016:BMS

[3263] Jaehee Kim and Chulwoo Jeong. A Bayesian multiple structural change regression model with autocorrelated errors. Journal of Applied Statistics, 43 (9):1690–1705, September 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sampson:2016:MAM

[3264] Joshua N. Sampson, Charles E.

Matthews, Laurence S. Freedman, Raymond J. Carroll, and Victor Kipnis. Methods to assess measurement error in questionnaires of sedentary behavior. Journal of Applied Statistics, 43 (9):1706–1721, September 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rakhmawati:2016:LID

[3265] Trias Wahyuni Rakhmawati, Geert Molenberghs, Geert Verbeke, and Christel Faes. Local influence diagnostics for incomplete overdispersed longitudinal counts. Journal of Applied Statistics, 43(9):1722–1737, September 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hernandez:2016:OSA

[3266] M. Álvarez Hernández, A. Martín Andrés, and I. Herranz Tejedor. Onesided asymptotic inferences for a proportion. Journal of Applied Statistics, 43(9):1738–1752, September 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Afzal:2016:NWO

[3267] Saima Afzal and Muhammad Mutahir Iqbal. A new way to order independent components. Journal of Applied Statistics, 43(9):1753–1764, September 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sanchez-Rodriguez:2016:IEF

[3268] M. I. Sánchez-Rodríguez, E. M. Sánchez-López, A. Marinas, J. M. Caridad, F. J. Urbano, and J. M. Marinas. Improving the estimations of fatty acids in several Andalusian PDO olive oils from NMR spectral data. Journal of Applied Statistics, 43(10): 1765–1793, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rubio:2016:SLD

[3269] Francisco J. Rubio and Yili Hong. Survival and lifetime data analysis with a flexible class of distributions. Journal of Applied Statistics, 43(10): 1794–1813, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bayes:2016:BIM

[3270] Cristian L. Bayes and Luis Valdivieso.

A beta inflated mean regression model for fractional response variables. Journal of Applied Statistics, 43(10):1814– 1830, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

SalmeronGomez:2016:CDA

[3271] Roman Salmerón Gómez, José García Pérez, María Del Mar López Martín, and Catalina García García. Collinearity diagnostic applied in ridge estimation through the variance inflation factor. Journal of Applied Statistics, 43 (10):1831–1849, October 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Noorian:2016:BTH

[3272] S. Noorian, M. Ganjali, and E. Bahrami Samani. A Bayesian test of homogeneity of association parameter using transition modelling of longitudinal mixed responses. Journal of Applied Statistics, 43(10):1850–1863, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oyeyemi:2016:DPM

[3273] Gafar Matanmi Oyeyemi, George Chinanu Mbaeyi, Saheed Ishola Salawu, and Bernard Olagboyega Muse. On discrimination procedure with mixtures of continuous and categorical variables. Journal of Applied Statistics, 43(10): 1864–1873, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cha:2016:ARC

[3274] Ji Hwan Cha. Analysis of reliability characteristics in the acceptance sampling tests. Journal of Applied Statistics, 43(10):1874–1891, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Turkan:2016:NMJ

[3275] Semra Türkan and Gamze Özel. A new modified Jackknifed estimator for the Poisson regression model. Journal of Applied Statistics, 43(10):1892–1905, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nugroho:2016:BCR

[3276] Didit B. Nugroho and Takayuki Morimoto. Box–Cox realized asymmetric stochastic volatility models with generalized Student’s t-error distributions. Journal of Applied Statistics, 43(10): 1906–1927, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Denhere:2016:REF

[3277] Melody Denhere and Huybrechts F. Bindele. Rank estimation for the functional linear model. Journal of Applied Statistics, 43(10):1928–1944, October 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2016:ECE

[3278] Housila P. Singh and Surya K. Pal. An efficient class of estimators of finite population variance using quartiles. Journal of Applied Statistics, 43 (10):1945–1958, October 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brentari:2016:CRD

[3279] Eugenio Brentari, Livia Dancelli, and Marica Manisera. Clustering ranking data in market segmentation: a case study on the Italian McDonald’s customers’ preferences. Journal of Applied Statistics, 43(11):1959–1976, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kasap:2016:SAC

[3280] Pelin Kasap, Birdal Senoglu, and Olcay Arslan. Stochastic analysis of covariance when the error distribution is long-tailed symmetric. Journal of Applied Statistics, 43(11):1977–1997, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lombardo:2016:PIA

[3281] R. Lombardo and E. J. Beh. The prediction index of aggregate data. Journal of Applied Statistics, 43(11): 1998–2018, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gao:2016:SSP

[3282] Hsiao-Hsian Gao and Li-Shan Huang.

Sample size planning for testing significance of curves. Journal of Applied Statistics, 43(11):2019–2028, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2016:NMD

[3283] Chien-Chia L. Huang, Yow-Jen Jou, and Hsun-Jung Cho. A new multicollinearity diagnostic for generalized linear models. Journal of Applied Statistics, 43(11):2029–2043, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lechner:2016:DDE

[3284] Michael Lechner, Nuria Rodriguez-Planas, and Daniel Fernández Kranz. Difference-in-difference estimation by FE and OLS when there is panel non-response. Journal of Applied Statistics, 43(11):2044–2052, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2016:GFT

[3285] Kuo-Chin Lin and Yi-Ju Chen. Goodness-of-fit tests of generalized linear mixed models for repeated ordinal responses. Journal of Applied Statistics, 43(11):2053–2064, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Corbier:2016:HFA

[3286] Christophe Corbier. Huberian function applied to neurodegenerative disorder gait rhythm. Journal of Applied Statistics, 43(11):2065–2084, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rizzato:2016:BAA

[3287] Fernanda B. Rizzato, Roseli A. Leandro, Clarice G. B. Demétrio, and Geert Molenberghs. A Bayesian approach to analyse overdispersed longitudinal count data. Journal of Applied Statistics, 43(11):2085–2109, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiao:2016:EPG

[3288] Xun Xiao, Amitava Mukherjee, and Min Xie. Estimation procedures for grouped data — a comparative study. Journal of Applied Statistics, 43(11): 2110–2130, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shao:2016:MNB

[3289] Jiguang Shao and Sheng Fu. On the modes of the negative binomial distribution of order k. Journal of Applied Statistics, 43(11):2131–2149, November 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Manghi:2016:EMM

[3290] Roberto F. Manghi, Gilberto A. Paula, and Francisco José A. Cysneiros. On elliptical multilevel models. Journal of Applied Statistics, 43(12):2150–2171, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Seo:2016:OBA

[3291] Jung-In Seo and Suk-Bok Kang. An objective Bayesian analysis of the two-parameter half-logistic distribution based on progressively Type-II censored samples. Journal of Applied Statistics, 43(12):2172–2190, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Faisal:2016:CSS

[3292] Muhammad Faisal, Andreas Futschik, Ijaz Hussain, and Mitwali Abdel.Moemen. Choosing summary statistics by least angle regression for approximate Bayesian computation.

Journal of Applied Statistics, 43(12): 2191–2202, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Razie:2016:LVM

[3293] F. Razie, E. Bahrami Samani, and M. Ganjali. A latent variable model for analyzing mixed longitudinal (k, l)- inflated count and ordinal responses. Journal of Applied Statistics, 43(12): 2203–2224, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2016:EIC

[3294] Chen Chen and Feng Guo. Evaluating the influence of crashes on driving risk using recurrent event models and Naturalistic Driving Study data. Journal of Applied Statistics, 43(12): 2225–2238, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haile:2016:PGD

[3295] S. R. Haile, J.-H. Jeong, X. Chen, and Y. Cheng. A 3-parameter Gompertz distribution for survival data with competing risks, with an application to breast cancer data. Journal of Applied Statistics, 43(12):2239–2253, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vaitheeswaran:2016:COL

[3296] K. Vaitheeswaran, M. Subbiah, R. Ramakrishnan, and T. Kannan. A comparison of ordinal logistic regression models using Classical and Bayesian approaches in an analysis of factors associated with diabetic retinopathy. Journal of Applied Statistics, 43(12):

2254–2260, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coolen-Maturi:2016:NWR

[3297] Tahani Coolen-Maturi. New weighted rank correlation coefficients sensitive to agreement on top and bottom rankings. Journal of Applied Statistics, 43(12): 2261–2279, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elsawah:2016:AUD

[3298] A. M. Elsawah and Hong Qin. Asymmetric uniform designs based on mixture discrepancy. Journal of Applied Statistics, 43(12):2280–2294, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sharpe:2016:UFP

[3299] James Sharpe and Nick Fieller. Uncertainty in functional principal component analysis. Journal of Applied Statistics, 43(12):2295–2309, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jhwueng:2016:ATE

[3300] D.-C. Jhwueng and V. Maroulas.

Adaptive trait evolution in random environment. Journal of Applied Statistics, 43(12):2310–2324, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2016:CID

[3301] Yu-Mei Chang, Pao-Sheng Shen, and Guan-Wei Liu. Confidence intervals for the difference between two median survival times for clustered survival data. Journal of Applied Statistics, 43(12):

2325–2345, December 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aykroyd:2016:E

[3302] Robert G. Aykroyd, M. Rosario Gonzales-Rogriguez, Biagio Simonetti, and Massimo Squillante. Editorial. Journal of Applied Statistics, 43(13): 2347–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DiPalma:2016:CMA

[3303] M. A. Di Palma and M. Gallo. A co-median approach to detect compositional outliers. Journal of Applied Statistics, 43(13):2348–??, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Romano:2016:DDI

[3304] A. A. Romano and G. Scandurra. Divergences in the determinants of investments in renewable energy sources: hydroelectric vs. other renewable sources. Journal of Applied Statistics, 43(13): 2363–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khan:2016:SGG

[3305] Md Hasinur Rahaman Khan, A. M. Azharul Islam, and Faisal Ababneh. Substantial gender gap reduction in Bangladesh explained by the proximity measure of literacy and life expectancy. Journal of Applied Statistics, 43(13):2377–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gonzalez-Rodriguez:2016:CSR

[3306] M. Rosario González-Rodríguez, M. Carmen Díaz Fernández, and Biagio Simonetti. Corporate social responsibility perception versus human values: a structural equation modeling approach. Journal of Applied Statistics, 43(13):2396–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Najafabadi:2016:BEC

[3307] Amir T. Payandeh Najafabadi and Maryam Omidi Najafabadi. On the Bayesian estimation for Cronbach’s alpha. Journal of Applied Statistics, 43(13):2416–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scippacercola:2016:OPC

[3308] Sergio Scippacercola and Enrica Sepe.

Ordinal principal component analysis for a common ranking of stochastic frontiers. Journal of Applied Statistics, 43(13):2442–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Figliola:2016:EMC

[3309] Alejandra Figliola and Lucas Catalano. Evolution of multifractal crosscorrelations between the Argentina MERVAL Index and international commodities prices. Journal of Applied Statistics, 43(13):2452–??, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chandrasekhar:2016:PRR

[3310] C. K. Chandrasekhar, H. Bagyalakshmi, M. R. Srinivasan, and M. Gallo.

Partial ridge regression under multicollinearity. Journal of Applied Statistics, 43(13):2462–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demir:2016:DFA

[3311] Mehmet Ozer Demir, Murat Alper Basaran, and Biagio Simonetti. Determining factors affecting healthcare service satisfaction utilizing fuzzy rulebased systems. Journal of Applied Statistics, 43(13):2474–??, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sarnacchiaro:2016:COV

[3312] Pasquale Sarnacchiaro, Antonello D’Ambra, and Luigi D’Ambra. CATANOVA for ordinal variables using orthogonal polynomials with different scoring methods. Journal of Applied Statistics, 43(13):2490–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lai:2016:MFM

[3313] Dongbing Lai, Huiping Xu, Daniel Koller, Tatiana Foroud, and Sujuan Gao. A multivariate finite mixture latent trajectory model with application to dementia studies. Journal of Applied Statistics, 43(14):2503–??, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ding:2016:BRN

[3314] Jianhua Ding and Zhongzhan Zhang. Bayesian regression on non-parametric mixed-effect models with shaperestricted Bernstein polynomials. Journal of Applied Statistics, 43(14):2524– ??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hossain:2016:IGM

[3315] Ahmed Hossain and Hafiz T. A. Khan.

Identification of genomic markers correlated with sensitivity in solid tumors to Dasatinib using sparse principal components. Journal of Applied Statistics, 43(14):2538–??, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Magnussen:2016:AMS

[3316] S. Magnussen, G. Frazer, and M. Penner. Alternative mean-squared error estimators for synthetic estimators of domain means. Journal of Applied Statistics, 43(14):2550–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bagnato:2016:KLA

[3317] L. Bagnato, L. De Capitani, and A. Punzo. The Kullback–Leibler autodependogram. Journal of Applied Statistics, 43(14):2574–??, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fan:2016:ERM

[3318] Yan Fan and Yukun Liu. Evaluating the relative merits of competing models based on empirical likelihood ratio test. Journal of Applied Statistics, 43(14):2595–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Afify:2016:WFD

[3319] Ahmed Z. Afify, Haitham M. Yousof, Gauss M. Cordeiro, Edwin M. M. Ortega, and Zohdy M. Nofal. The Weibull Fréchet distribution and its applications. Journal of Applied Statistics, 43(14):2608–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See correction [4205].

Cabras:2016:BNP

[3320] Stefano Cabras and Juan de Dios Tena Horrillo. A Bayesian nonparametric modeling to estimate student response to ICT investment. Journal of Applied Statistics, 43(14):2627– ??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2016:EDA

[3321] Cheolwoo Park, Yongho Jeon, and Kee-Hoon Kang. An exploratory data analysis in scale-space for intervalvalued data. Journal of Applied Statistics, 43(14):2643–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coban:2016:TVM

[3322] Onur Çoban, Talha Kivanç, Mustafa Özgür Bora, Burcu Özcan, Tamer Sinmazçelik, and Sinan Fidan. Thermal, viscoelastic and mechanical properties’ optimization of polyphenylene sulfide via optimal processing parameters using the Taguchi method. Journal of Applied Statistics, 43(14):2661–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kumar:2016:OVA

[3323] C. Satheesh Kumar and A. Riyaz. An order k version of the alternative zeroinflated logarithmic series distribution and its applications. Journal of Applied Statistics, 43(14):2681–??, 2016.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2016:Ca

[3324] Anonymous. Corrigendum. Journal of Applied Statistics, 43(14):2696–??, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jana:2016:BEE

[3325] Nabakumar Jana, Somesh Kumar, and Kashinath Chatterjee. Bayes estimation for exponential distributions with common location parameter and applications to multi-state reliability models. Journal of Applied Statistics, 43(15):2697–2712, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2016:JAN

[3326] Xiaosun Lu, Yangxin Huang, and Rong Zhou. Joint analysis of nonlinear heterogeneous longitudinal data and binary outcome: an application to AIDS clinical studies. Journal of Applied Statistics, 43(15):2713–2728, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jeyaseelan:2016:LDA

[3327] Visalakshi Jeyaseelan, Tunny Sebastian, Jeyaseelan Lakshmanan, and Shrikant I. Bangdiwala. Longitudinal data analysis of mean passage time among malnutrition states: an application of Markov chains. Journal of Applied Statistics, 43(15):2729–2739, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cai:2016:CPA

[3328] Xia Cai, Khamis Khalid Said, and Wei Ning. Change-point analysis with bathtub shape for the exponential distribution. Journal of Applied Statistics, 43(15):2740–2750, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Choi:2016:BVS

[3329] Taeryon Choi, Hea-Jung Kim, and Seongil Jo. Bayesian variable selection approach to a Bernstein polynomial regression model with stochastic constraints. Journal of Applied Statistics, 43(15):2751–2771, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harkanen:2016:SHM

[3330] Tommi Härkänen, Juha Karvanen, Hanna Tolonen, Risto Lehtonen, Kari Djerf, Teppo Juntunen, and Seppo Koskinen. Systematic handling of missing data in complex study designs — experiences from the Health 2000 and 2011 Surveys. Journal of Applied Statistics, 43(15):2772–2790, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Acar:2016:IMB

[3331] T. Söküt Açar and M. R. Özkale. Influence measures based on confidence ellipsoids in general linear regression model with correlated regressors. Journal of Applied Statistics, 43(15):2791– 2812, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ayele:2016:CMS

[3332] Dawit G. Ayele and Temesgen T. Zewotir. Childhood mortality spatial distribution in Ethiopia. Journal of Applied Statistics, 43(15):2813–2828, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marchant:2016:DMG

[3333] Carolina Marchant, Víctor Leiva, Francisco José A. Cysneiros, and Juan F. Vivanco. Diagnostics in multivariate generalized Birnbaum–Saunders regression models. Journal of Applied Statistics, 43(15):2829–2849, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2016:SBA

[3334] Kiranmoy Das. A semiparametric Bayesian approach for joint modeling of longitudinal trait and event time. Journal of Applied Statistics, 43(15):2850– 2865, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tutz:2016:GAM

[3335] Gerhard Tutz and Sebastian Petry.

Generalized additive models with unknown link function including variable selection. Journal of Applied Statistics, 43(15):2866–2885, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Atem:2016:MIR

[3336] Folefac D. Atem, Jing Qian, Jacqueline E. Maye, Keith A. Johnson, and Rebecca A. Betensky. Multiple imputation of a randomly censored covariate improves logistic regression analysis. Journal of Applied Statistics, 43(15):

2886–2896, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2016:MTS

[3337] Hongyue Guo, Xiaodong Liu, and Zhubin Sun. Multivariate time series prediction using a hybridization of VARMA models and Bayesian networks. Journal of Applied Statistics, 43(16):2897–2909, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rad:2016:MEE

[3338] N. Nakhaei Rad, G. R. Mohtashami Borzadaran, and G. H. Yari. Maximum entropy estimation of income share function from generalized Gini index. Journal of Applied Statistics, 43(16):2910–2921, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2016:JAL

[3339] T.-F. C. Lu, C.-M. Hsu, K.-H. Shu, S.- C. Weng, and C.-M. Chen. Joint analysis of longitudinal data and competing terminal events in the presence of dependent observation times with application to chronic kidney disease. Journal of Applied Statistics, 43(16):2922– 2940, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dai:2016:DMH

[3340] Xianhua Dai, Wolfgang Karl Härdle, and Keming Yu. Do maternal health problems influence child’s worrying status? Evidence from the British Cohort Study. Journal of Applied Statistics, 43(16):2941–2955, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Louzada:2016:MIF

[3341] F. Louzada and P. H. Ferreira. Modified inference function for margins for the bivariate Clayton copula-based SUN Tobit Model. Journal of Applied Statistics, 43(16):2956–2976, 2016. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mota:2016:MSS

[3342] Pedro P. Mota and Manuel L. Esquível.

Model selection for stock prices data. Journal of Applied Statistics, 43(16): 2977–2987, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Noughabi:2016:ECD

[3343] Hadi Alizadeh Noughabi and Albert Vexler. An efficient correction to the density-based empirical likelihood ratio goodness-of-fit test for the inverse Gaussian distribution. Journal of Applied Statistics, 43(16):2988–3003, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cafri:2016:SCS

[3344] Guy Cafri. Sparsely clustered survival data: application to the evaluation of the safety and effectiveness of medical devices. Journal of Applied Statistics, 43(16):3004–3014, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Canale:2016:BMU

[3345] Antonio Canale, Euloge Clovis Kenne Pagui, and Bruno Scarpa. Bayesian modeling of university first-year students’ grades after placement test. Journal of Applied Statistics, 43(16): 3015–3029, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Perez-Carceles:2016:DCI

[3346] María Concepción Pérez-Cárceles, Juan Cándido Gómez-Gallego, and Juan Gómez-García. Distribution of cost inefficiency in stochastic frontier approach: evidence from Spanish banking. Journal of Applied Statistics, 43(16):3030–3041, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2016:Cb

[3347] Anonymous. Corrigendum. Journal of Applied Statistics, 43(16):3042, 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2017:E

[3348] Jie Chen. Editorial. Journal of Applied Statistics, 44(1):1–2, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cai:2017:ISS

[3349] Jing Cai, Yimin Shi, and Bin Liu. Inference for a series system with dependent masked data under progressive interval censoring. Journal of Applied Statistics, 44(1):3–15, January 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Best:2017:ARD

[3350] D. J. Best and J. C. W. Rayner. Analysis of ranked data in randomized blocks when there are missing values. Journal of Applied Statistics, 44(1):16–23, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McNeish:2017:MDM

[3351] Daniel McNeish. Missing data methods for arbitrary missingness with small samples. Journal of Applied Statistics, 44(1):24–39, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shieh:2017:ETA

[3352] Gwowen Shieh. The equivalence of two approaches to incorporating variance uncertainty in sample size calculations for linear statistical models. Journal of Applied Statistics, 44(1):40–56, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2017:BAR

[3353] Shengping Yang and Zhide Fang. Beta approximation of ratio distribution and its application to next generation sequencing read counts. Journal of Applied Statistics, 44(1):57–70, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2017:CTH

[3354] Shaochuan Lu. A continuous-time HMM approach to modeling the magnitude-frequency distribution of earthquakes. Journal of Applied Statistics, 44(1):71–88, January 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bersimis:2017:RTM

[3355] Sotirios Bersimis, Stavros Degiannakis, and Dimitrios Georgakellos. Real-time monitoring of carbon monoxide using value-at-risk measure and control charting. Journal of Applied Statistics, 44(1):89–108, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dlugosz:2017:CRQ

[3356] Stephan Dlugosz, Simon M. S. Lo, and Ralf A. Wilke. Competing risks quantile regression at work: in-depth exploration of the role of public child support for the duration of maternity leave. Journal of Applied Statistics, 44 (1):109–122, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aslam:2017:CCM

[3357] Muhammad Aslam, G. Srinivasa Rao, Liaquat Ahmad, and Chi-Hyuck Jun. A control chart for multivariate Poisson distribution using repetitive sampling. Journal of Applied Statistics, 44 (1):123–136, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hambuckers:2017:RSA

[3358] J. Hambuckers and C. Heuchenne. A robust statistical approach to select adequate error distributions for financial returns. Journal of Applied Statistics, 44(1):137–161, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hossain:2017:SEL

[3359] Shakhawat Hossain and Hatem A. Howlader. Shrinkage estimation in lognormal regression model for censored data. Journal of Applied Statistics, 44 (1):162–180, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2017:FDF

[3360] Jianmin Liu, Xiaomei Hu, and Jinguang Wu. Fiscal decentralization, financial efficiency and upgrading the industrial structure: an empirical analysis of a spatial heterogeneity model. Journal of Applied Statistics, 44(1): 181–196, January 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miljkovic:2017:RSB

[3361] Tatjana Miljkovic, Saleem Shaik, and Dragan Miljkovic. Redefining standards for body mass index of the US population based on BRFSS data using mixtures. Journal of Applied Statistics, 44(2):197–211, February 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hung:2017:WVK

[3362] Wen-Liang Hung, De-Hua Chen, and Jenn-Hwai Yang. Weighting variables in Kohonen competitive learning algorithms. Journal of Applied Statistics, 44(2):212–232, February 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pescim:2017:NLL

[3363] Rodrigo R. Pescim, Edwin M. M. Ortega, Gauss M. Cordeiro, and Morad Alizadeh. A new log-location regression model: estimation, influence diagnostics and residual analysis. Journal of Applied Statistics, 44(2):233–252, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tu:2017:BVS

[3364] Shiyi Tu, Min Wang, and Xiaoqian Sun. Bayesian variable selection and estimation in maximum entropy quantile regression. Journal of Applied Statistics, 44(2):253–269, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kombo:2017:MIO

[3365] A. Y. Kombo, H. Mwambi, and G. Molenberghs. Multiple imputation for ordinal longitudinal data with monotone missing data patterns. Journal of Applied Statistics, 44(2): 270–287, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hwang:2017:WGL

[3366] Y. T. Hwang, C. H. Huang, W. L. Yeh, and Y. D. Shen. The weighted general linear model for longitudinal medical cost data — an application in colorectal cancer. Journal of Applied Statistics, 44(2):288–307, February 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luo:2017:SRS

[3367] YangXia Luo. Statistics and recognition for software birthmark based on clustering analysis. Journal of Applied Statistics, 44(2):308–324, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Keefe:2017:MFR

[3368] Matthew J. Keefe, Christopher T. Franck, and William H. Woodall. Monitoring foreclosure rates with a spatially risk-adjusted Bernoulli CUSUM chart for concurrent observations. Journal of Applied Statistics, 44(2):325–341, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Celik:2017:NTD

[3369] Resit Çelik. A new test to detect monotonic and non-monotonic types of heteroscedasticity. Journal of Applied Statistics, 44(2):342–361, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sun:2017:THV

[3370] Huihui Sun. Testing of homogeneity of variance and autocorrelation coefficients of nonlinear mixed models with AR(1) errors based on M-estimation. Journal of Applied Statistics, 44(2): 362–375, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Malathi:2017:EDA

[3371] D. Malathi and S. Muthulakshmi. Economic design of acceptance sampling plans for truncated life test using Frechet distribution. Journal of Applied Statistics, 44(2):376–384, February 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vovan:2017:DCP

[3372] Tai Vovan. L^{1}-distance and classification problem by Bayesian method. Journal of Applied Statistics, 44(3): 385–401, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kvaalseth:2017:CVS

[3373] Tarald O. Kvålseth. Coefficient of variation: the second-order alternative. Journal of Applied Statistics, 44(3):402–415, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ko:2017:ATG

[3374] Feng shou Ko. Application of trajectories from growth curve in identification of longitudinal biomarker for the multivariate survival data. Journal of Applied Statistics, 44(3):416–426, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saengseedam:2017:STM

[3375] Panudet Saengseedam and Nantachai Kantanantha. Spatio-temporal model for crop yield forecasting. Journal of Applied Statistics, 44(3):427–440, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luwanda:2017:NME

[3376] A. G. Luwanda and H. G. Mwambi. A nonlinear mixed-effects model for multivariate longitudinal data with partially observed outcomes with application to HIV disease dynamics. Journal of Applied Statistics, 44(3):441–456, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2017:SBR

[3377] Lili Yu, Hani Samawi, Daniel Linder, Arpita Chatterjee, Yisong Huang, and Robert Vogel. On stratified bivariate ranked set sampling with optimal allocation for naïve and ratio estimators. Journal of Applied Statistics, 44(3):457–473, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhou:2017:ELI

[3378] Xiaoshuang Zhou, Peixin Zhao, and Xiuli Wang. Empirical likelihood inferences for varying coefficient partially nonlinear models. Journal of Applied Statistics, 44(3):474–492, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kulahci:2017:SPD

[3379] Murat Kulahci and John Tyssedal.

Split-plot designs for multistage experimentation. Journal of Applied Statistics, 44(3):493–510, March 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hamedani:2017:DBD

[3380] Hamideh D. Hamedani and Sara Sadat Moosavi. Dealing with big data: comparing dimension reduction and shrinkage regression methods. Journal of Applied Statistics, 44(3):511–532, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Drosou:2017:PSV

[3381] Krystallenia Drosou and Christos Koukouvinos. Proximal support vector machine techniques on medical prediction outcome. Journal of Applied Statistics, 44(3):533–553, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Owen:2017:EPP

[3382] A. Owen, A. C. Smith, P. Osei-Owusu, A. Harland, and J. R. Roberts. Elite players’ perceptions of football playing surfaces: a mixed effects ordinal logistic regression model of players’ perceptions. Journal of Applied Statistics, 44(3):554–570, March 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aboagye-Sarfo:2017:ITS

[3383] Patrick Aboagye-Sarfo, James Cross, and Ute Mueller. Intervention time series analysis of voluntary, counselling and testing on HIV infections in West African sub-region: the case of Ghana. Journal of Applied Statistics, 44(4):571–582, April 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nguyentrang:2017:FCP

[3384] Thao Nguyentrang and Tai Vovan. Fuzzy clustering of probability density functions. Journal of Applied Statistics, 44(4):583–601, April 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rienzner:2017:CVI

[3385] Michele Rienzner and Francesca Ieva. Critical values improvement for the standard normal homogeneity test by combining Monte Carlo and regression approaches. Journal of Applied Statistics, 44(4):602–619, April 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rakhmawati:2017:LID

[3386] Trias Wahyuni Rakhmawati, Geert Molenberghs, Geert Verbeke, and Christel Faes. Local influence diagnostics for generalized linear mixed models with overdispersion. Journal of Applied Statistics, 44(4):620–641, April 2017.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2017:SMI

[3387] Feng Xu. Statistical measurement of the inventory shortage cost. Journal of Applied Statistics, 44(4):642–648, April 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Afshari:2017:NBW

[3388] M. Afshari, F. Lak, and B. Gholizadeh.

A new Bayesian wavelet thresholding estimator of nonparametric regression. Journal of Applied Statistics, 44(4):649–666, April 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kakarantza:2017:SUS

[3389] Eudoxia Kakarantza and Spyridon D.

Symeonides. Seemingly unrelated systems of econometric equations. Journal of Applied Statistics, 44(4):667–684, April 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ham:2017:CDA

[3390] Seunghon Ham, Sunju Kim, Naroo Lee, Pilje Kim, Igchun Eom, Byoungcheun Lee, Perng-Jy Tsai, Kiyoung Lee, and Chungsik Yoon. Comparison of data analysis procedures for real-time nanoparticle sampling data using classical regression and ARIMA models. Journal of Applied Statistics, 44(4):685–699, April 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dhhan:2017:HBH

[3391] Waleed Dhhan, Sohel Rana, and Habshah Midi. A high breakdown, high efficiency and bounded influence modified GM estimator based on support vector regression. Journal of Applied Statistics, 44(4):700–714, April 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2017:DAT

[3392] Maosheng Li, Zhengqiu Liu, Yonghong Zhang, Weijun Liu, and Feng Shi. Distribution analysis of train interval journey time employing the censored model with shifting character. Journal of Applied Statistics, 44(4):715–733, April 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Templ:2017:ETO

[3393] M. Templ, K. Hron, and P. Filzmoser. Exploratory tools for outlier detection in compositional data with structural zeros. Journal of Applied Statistics, 44(4):734–752, April 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samawi:2017:MEL

[3394] Hani M. Samawi, Haresh Rochani, Daniel Linder, and Arpita Chatterjee. More efficient logistic analysis using moving extreme ranked set sampling. Journal of Applied Statistics, 44(4):753–766, April 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alizadeh:2017:NFP

[3395] Mojtaba Alizadeh, Seyyed Fazel Bagheri, Mohammad Alizadeh, and Saralees Nadarajah. A new fourparameter lifetime distribution. Journal of Applied Statistics, 44(5):767– 797, May 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rubio:2017:FOB

[3396] Francisco J. Rubio and Keming Yu.

Flexible objective Bayesian linear regression with applications in survival analysis. Journal of Applied Statistics, 44(5):798–810, May 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2017:BML

[3397] Kyeongjun Lee and Youngseuk Cho.

Bayesian and maximum likelihood estimations of the inverted exponentiated half logistic distribution under progressive Type II censoring. Journal of Applied Statistics, 44(5):811–832, May 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Johnson:2017:GLM

[3398] Nels Johnson and Inyoung Kim. Generalized linear models with covariate measurement error and unknown link function. Journal of Applied Statistics, 44(5):833–852, May 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alshaybawee:2017:BEN

[3399] Taha Alshaybawee, Habshah Midi, and Rahim Alhamzawi. Bayesian elastic net single index quantile regression. Journal of Applied Statistics, 44(5):853–871, May 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Adnan:2017:SDP

[3400] Mian Arif Shams Adnan and Shongkour Roy. A sequential discrimination procedure for two almost identically shaped wrapped distributions. Journal of Applied Statistics, 44(5):872–881, May 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miranda:2017:PMC

[3401] Douglas Moura Miranda and Samuel Vieira Conceição. A practical method to calculate probabilities: illustrative example from the electronic industry business. Journal of Applied Statistics, 44(5):882–896, May 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Das:2017:CRE

[3402] Rabindra Nath Das and Anis Chandra Mukhopadhyay. Correlated random effects regression analysis for a lognormally distributed variable. Journal of Applied Statistics, 44(5):897–915, May 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2017:BEP

[3403] Sukhdev Singh, Yogesh Mani Tripathi, and Shuo-Jye Wu. Bayesian estimation and prediction based on lognormal record values. Journal of Applied Statistics, 44(5):916–940, May 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2017:ICM

[3404] Xu Guo, Hecheng Wu, Gaorong Li, and Qiuyue Li. Inference for the common mean of several Birnbaum–Saunders populations. Journal of Applied Statistics, 44(5):941–954, May 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kurz:2017:FPE

[3405] D. Kurz, H. Lewitschnig, and J. Pilz.

Failure probability estimation under additional subsystem information with application to semiconductor burnin. Journal of Applied Statistics, 44 (6):955–967, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2017:EPL

[3406] Yunlu Jiang. S-estimator in partially linear regression models. Journal of Applied Statistics, 44(6):968–977, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hung:2017:LBE

[3407] Wen-Liang Hung and Shou-Jen Chang-Chien. Learning-based EM algorithm for normal-inverse Gaussian mixture model with application to extrasolar planets. Journal of Applied Statistics, 44(6):978–999, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gadre:2017:CSA

[3408] M. P. Gadre, S. B. Adnaik, and R. N.

Rattihalli. Continuous single attribute control chart for Markov-dependent processes. Journal of Applied Statistics, 44(6):1000–1012, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2017:IAR

[3409] S. Wang, N. G. Cadigan, and H. P.

Benoît. Inference about regression parameters using highly stratified survey count data with over-dispersion and repeated measurements. Journal of Applied Statistics, 44(6):1013–1030, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deLara:2017:POT

[3410] I. A. R. de Lara, J. P. Hinde, A. C. de Castro, and I. J. O. da Silva. A proportional odds transition model for ordinal responses with an application to pig behaviour. Journal of Applied Statistics, 44(6):1031–1046, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oliveira:2017:NVC

[3411] I. R. C. Oliveira, G. Molenberghs, G. Verbeke, C. G. B. Demétrio, and C. T. S. Dias. Negative variance components for non-negative hierarchical data with correlation, over-, and/or underdispersion. Journal of Applied Statistics, 44(6):1047–1063, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pertsinidou:2017:AHS

[3412] C. E. Pertsinidou, G. Tsaklidis, E. Papadimitriou, and N. Limnios. Application of hidden semi-Markov models for the seismic hazard assessment of the North and South Aegean Sea, Greece. Journal of Applied Statistics, 44(6):1064–1085, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jaeger:2017:SFE

[3413] Byron C. Jaeger, Lloyd J. Edwards, Kalyan Das, and Pranab K. Sen. An R^{2} statistic for fixed effects in the generalized linear mixed model. Journal of Applied Statistics, 44(6):1086–1105, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahdizadeh:2017:NGF

[3414] M. Mahdizadeh and Ehsan Zamanzade. New goodness of fit tests for the Cauchy distribution. Journal of Applied Statistics, 44(6):1106–1121, June 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fang:2017:NTS

[3415] Rui Fang and Xiaohu Li. Nonparametric tests for strictly increasing virtual valuations. Journal of Applied Statistics, 44(6):1122–1136, June 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stewart:2017:AMD

[3416] Connie Stewart. An approach to measure distance between compositional diet estimates containing essential zeros. Journal of Applied Statistics, 44 (7):1137–1152, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gallardo:2017:NCR

[3417] Diego I. Gallardo, Héctor W. Gómez, and Heleno Bolfarine. A new cure rate model based on the Yule–Simon distribution with application to a melanoma data set. Journal of Applied Statistics, 44(7):1153–1164, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kubota:2017:BPU

[3418] Tatsuya Kubota and Takeshi Kurosawa. Bayesian prediction of unobserved values for Type-II censored data. Journal of Applied Statistics, 44(7):1165–1180, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guneri:2017:PAD

[3419] Öznur Isçi Güneri, Atilla Göktas, and Ugur Kayali. Path analysis and determining the distribution of indirect effects via simulation. Journal of Applied Statistics, 44(7):1181–1210, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Carmona-Benitez:2017:CBE

[3420] Rafael Bernardo Carmona-Benítez and María Rosa Nieto. Comparison of bootstrap estimation intervals to forecast arithmetic mean and median air passenger demand. Journal of Applied Statistics, 44(7):1211–1224, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Santos:2017:UJP

[3421] Izabella A. R. A. Santos, Denise Duarte, and Marcelo Azevedo Costa. Use of jump process to model mobility in massive multiplayer on-line games. Journal of Applied Statistics, 44(7):1225–1247, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nakajima:2017:BMD

[3422] Jouchi Nakajima, Tsuyoshi Kunihama, and Yasuhiro Omori. Bayesian modeling of dynamic extreme values: extension of generalized extreme value distributions with latent stochastic processes. Journal of Applied Statistics, 44(7):1248–1268, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cammarota:2017:ETP

[3423] Camillo Cammarota. Estimating the turning point location in shifted exponential model of time series. Journal of Applied Statistics, 44(7):1269–1281, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Moghimbeygi:2017:LSA

[3424] M. Moghimbeygi and M. Golalizadeh.

Longitudinal shape analysis by using the spherical coordinates. Journal of Applied Statistics, 44(7):1282–1295, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Masci:2017:BMM

[3425] C. Masci, F. Ieva, T. Agasisti, and A. M. Paganoni. Bivariate multilevel models for the analysis of mathematics and reading pupils’ achievements. Journal of Applied Statistics, 44(7):1296–1317, July 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hawkins:2017:CCD

[3426] Douglas M. Hawkins and F. Lombard.

Cusum control for data following the von Mises distribution. Journal of Applied Statistics, 44(8):1319–1332, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baker:2017:NPI

[3427] Rebecca M. Baker, Tahani Coolen-Maturi, and Frank P. A. Coolen. Nonparametric predictive inference for stock returns. Journal of Applied Statistics, 44(8):1333–1349, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Men:2017:CAS

[3428] Zhongxian Men, Don McLeish, Adam W. Kolkiewicz, and Tony S. Wirjanto. Comparison of asymmetric stochastic volatility models under different correlation structures. Journal of Applied Statistics, 44(8):1350–1368, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Davarzani:2017:EDR

[3429] N. Davarzani, L. Golparvar, A. Parsian, and R. Peeters. Estimation on dependent right censoring scheme in an ordinary bivariate geometric distribution. Journal of Applied Statistics, 44 (8):1369–1384, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahabbati:2017:DST

[3430] A. Mahabbati, A. Izady, M. Mousavi Baygi, K. Davary, and S. M. Hasheminia. Daily soil temperature modeling using ‘panel-data’ concept. Journal of Applied Statistics, 44(8):1385–1401, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2017:GPQ

[3431] Philip L. H. Yu, Thomas Mathew, and Yuanyuan Zhu. A generalized pivotal quantity approach to portfolio selection. Journal of Applied Statistics, 44 (8):1402–1420, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhu:2017:ETF

[3432] Xiaojuan Zhu, William Seaver, Rapinder Sawhney, Shuguang Ji, Bruce Holt, Gurudatt Bhaskar Sanil, and Girish Upreti. Employee turnover forecasting for human resource management based on time series analysis. Journal of Applied Statistics, 44(8):1421–1440, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Altarazi:2017:DAM

[3433] Safwan A. Altarazi and Rula M. Allaf.

Designing and analyzing a mixture experiment to optimize the mixing proportions of polyvinyl chloride composites. Journal of Applied Statistics, 44 (8):1441–1465, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lei:2017:CDD

[3434] Yang Lei, Susan Carlson, Lisa N. Yelland, Maria Makrides, Robert Gibson, and Byron J. Gajewski. Comparison of dichotomized and distributional approaches in rare event clinical trial design: a fixed Bayesian design. Journal of Applied Statistics, 44(8):1466–1478, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mohammed:2017:EBP

[3435] Heba S. Mohammed, Saieed F. Ateya, and Essam K. Al-Hussaini. Estimation based on progressive first-failure censoring from exponentiated exponential distribution. Journal of Applied Statistics, 44(8):1479–1494, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shuweihdi:2017:CFU

[3436] Farag Shuweihdi, Charles C. Taylor, and Arief Gusnanto. Classification of form under heterogeneity and non-isotropic errors. Journal of Applied Statistics, 44(8):1495–1508, August 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghorbel:2017:DBO

[3437] Ahmed Ghorbel, Wajdi Hamma, and Anis Jarboui. Dependence between oil and commodities markets using timevarying Archimedean copulas and effectiveness of hedging strategies. Journal of Applied Statistics, 44(9):1509–1542, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2017:BIR

[3438] Fang Liu, Peng Zhang, Ibrahim Erkan, and Dylan S. Small. Bayesian inference for random coefficient dynamic panel data models. Journal of Applied Statistics, 44(9):1543–1559, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arbia:2017:SAH

[3439] Giuseppe Arbia, Giuseppe Espa, Diego Giuliani, and Rocco Micciolo. A spatial analysis of health and pharmaceutical firm survival. Journal of Applied Statistics, 44(9):1560–1575, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahmed:2017:EPG

[3440] Essam A. Ahmed. Estimation and prediction for the generalized inverted exponential distribution based on progressively first-failure-censored data with application. Journal of Applied Statistics, 44(9):1576–1608, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barnoy:2017:CAS

[3441] Eran A. Barnoy, Hyun J. Kim, and David W. Gjertson. Complexity in applying spatial analysis to describe heterogeneous air-trapping in thoracic imaging data. Journal of Applied Statistics, 44(9):1609–1629, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

daPaz:2017:BEM

[3442] Rosineide Fernando da Paz, Jorge Luis Bazán, and Luis Aparecido Milan. Bayesian estimation for a mixture of simplex distributions with an unknown number of components: HDI analysis in Brazil. Journal of Applied Statistics, 44(9):1630–1643, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2017:CRR

[3443] Ying Liu, Fang Luo, Danhui Zhang, and Hongyun Liu. Comparison and robustness of the REML, ML, MIVQUE estimators for multi-level random mediation model. Journal of Applied Statistics, 44(9):1644–1661, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silva:2017:PLS

[3444] Eliud Silva and Víctor M. Guerrero. Penalized least squares smoothing of two-dimensional mortality tables with imposed smoothness. Journal of Applied Statistics, 44(9):1662–1679, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bersimis:2017:ISF

[3445] Fragkiskos Bersimis, Demosthenes Panagiotakos, and Malvina Vamvakari. Investigating the sensitivity function’s monotony of a health-related index. Journal of Applied Statistics, 44(9): 1680–1706, September 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dhindsa:2017:PCA

[3446] I. S. Dhindsa, R. Agarwal, and H. S. Ryait. Principal component analysis-based muscle identification for myoelectric-controlled exoskeleton knee. Journal of Applied Statistics, 44(10):1707–1720, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Che:2017:SCC

[3447] JinXing Che and YouLong Yang. Stochastic correlation coefficient ensembles for variable selection. Journal of Applied Statistics, 44(10):1721–1742, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Longford:2017:IAD

[3448] Nicholas T. Longford. Inflated assessments of disability. Journal of Applied Statistics, 44(10):1743–1760, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peterson:2017:SPC

[3449] Geoffrey Colin L. Peterson, Dong Li, Brian J. Reich, and Donald Brenner. Spatial prediction of crystalline defects observed in molecular dynamic simulations of plastic damage. Journal of Applied Statistics, 44(10):1761–1784, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cugnata:2017:BNA

[3450] F. Cugnata, G. Perucca, and S. Salini.

Bayesian networks and the assessment of universities’ value added. Journal of Applied Statistics, 44(10):1785–1806, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martin:2017:MZI

[3451] Jacob Martin and Daniel B. Hall.

Marginal zero-inflated regression models for count data. Journal of Applied Statistics, 44(10):1807–1826, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Costa:2017:MMA

[3452] André G. F. C. Costa, Enrico A.

Colosimo, Aline B. M. Vaz, José Luiz P. Silva, and Leila D. Amorim. Marginal models for the association structure of hierarchical binary responses. Journal of Applied Statistics, 44(10):1827–1838, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2017:RVS

[3453] Tao Wang, Lin Zheng, Zhonghua Li, and Haiyang Liu. A robust variable screening method for high-dimensional data. Journal of Applied Statistics, 44 (10):1839–1855, October 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sankaran:2017:NEM

[3454] P. G. Sankaran and N. N. Midhu. Nonparametric estimation of mean residual quantile function under right censoring. Journal of Applied Statistics, 44(10):1856–1874, October 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baione:2017:CRT

[3455] Fabio Baione, Paolo De Angelis, Massimiliano Menzietti, and Agostino Tripodi. A comparison of risk transfer strategies for a portfolio of life annuities based on RORAC. Journal of Applied Statistics, 44(10):1875–1892, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2017:DDP

[3456] Ying-Hsiu Chen and Po-Lin Lai. Does diversification promote risk reduction and profitability raise? Estimation of dynamic impacts using the pooled mean group model. Journal of Applied Statistics, 44(10):1893–1901, October 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2017:FAH

[3457] Xiaoqi Jiang, Steven Wink, Bob van de Water, and Annette Kopp-Schneider. Functional analysis of high-content high-throughput imaging data. Journal of Applied Statistics, 44(11):1903–1919, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Inan:2017:JGM

[3458] G. Inan and R. Yucel. Joint GEEs for multivariate correlated data with incomplete binary outcomes. Journal of Applied Statistics, 44(11):1920–1937, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2017:CIC

[3459] Wei Wang. Checking identifiability of covariance parameters in linear mixed effects models. Journal of Applied Statistics, 44(11):1938–1946, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alhamzawi:2017:ITL

[3460] Rahim Alhamzawi. Inference with three-level prior distributions in quantile regression problems. Journal of Applied Statistics, 44(11):1947–1959, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2017:MIV

[3461] Hongmei Lin, Wenchao Xu, Riquan Zhang, Jianhong Shi, and Yuedong Wang. Multiple-index varyingcoefficient models for longitudinal data. Journal of Applied Statistics, 44(11): 1960–1978, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wan:2017:VCA

[3462] Alan T. K. Wan, Shangyu Xie, and Yong Zhou. A varying coefficient approach to estimating hedonic housing price functions and their quantiles. Journal of Applied Statistics, 44(11): 1979–1999, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhou:2017:LBA

[3463] Guohai Zhou, Lang Wu, Rollin Brant, and J. Mark Ansermino. A likelihoodbased approach for multivariate onesided tests with missing data. Journal of Applied Statistics, 44(11):2000–2016, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stojkovic:2017:EFF

[3464] Milan Stojković, Stevan Prohaska, and Nikola Zlatanović. Estimation of flood frequencies from data sets with outliers using mixed distribution functions. Journal of Applied Statistics, 44 (11):2017–2035, November 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kerns:2017:SCB

[3465] Lucy Kerns and John T. Chen. Simultaneous confidence bands for restricted logistic regression models. Journal of Applied Statistics, 44(11):2036–2051, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vilca:2017:BIS

[3466] Filidor Vilca, Caio L. N. Azevedo, and N. Balakrishnan. Bayesian inference for sinh-normal/independent nonlinear regression models. Journal of Applied Statistics, 44(11):2052–2074, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liang:2017:DFM

[3467] Wenjuan Liang, Xiaolong Pu, and Dongdong Xiang. A distribution-free multivariate CUSUM control chart using dynamic control limits. Journal of Applied Statistics, 44(11):2075–2093, November 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cannas:2017:VCD

[3468] M. Cannas, C. Conversano, F. Mola, and E. Sironi. Variation in Caesarean delivery rates across hospitals: a Bayesian semi-parametric approach. Journal of Applied Statistics, 44(12): 2095–2107, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Naji:2017:LIS

[3469] Z. Naji, A. Rasekh, and E. L. Boone.

Local influence in seemingly unrelated regression model with ridge estimate. Journal of Applied Statistics, 44(12): 2108–2124, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ferreira:2017:PLM

[3470] Clécio S. Ferreira, Camila B. Zeller, Aparecida M. S. Mimura, and Júlio C. J. Silva. Partially linear models and their applications to change point detection of chemical process data. Journal of Applied Statistics, 44(12): 2125–2141, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Calzolari:2017:SSD

[3471] Giorgio Calzolari and Antonino Di Pino. Self-selection and direct estimation of across-regime correlation parameter. Journal of Applied Statistics, 44(12):2142–2160, December 2017.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2017:DBM

[3472] Chien-Chia L. Huang, Yow-Jen Jou, and Hsun-Jung Cho. Differencebased matrix perturbation method for semi-parametric regression with multicollinearity. Journal of Applied Statistics, 44(12):2161–2171, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gorecki:2017:MAV

[3473] T. Górecki and L. Smaga. Multivariate analysis of variance for functional data. Journal of Applied Statistics, 44 (12):2172–2189, December 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Longford:2017:AMC

[3474] Nicholas T. Longford, José Rafael Tovar Cuevas, and Carlos Alvear. Analysis of a marker for cancer of the thyroid with a limit of detection. Journal of Applied Statistics, 44(12): 2190–2203, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yan:2017:NCI

[3475] Liang Yan, Rui Wang, and Xingzhong Xu. A new confidence interval in errorsin-variables model with known error variance. Journal of Applied Statistics, 44(12):2204–2221, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Seo:2017:OBA

[3476] Jung In Seo and Yongku Kim. Objective Bayesian analysis based on upper record values from two-parameter Rayleigh distribution with partial information. Journal of Applied Statistics, 44(12):2222–2237, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guney:2017:MOD

[3477] Yesim Güney, Yetkin Tuaç, and Olcay Arslan. Marshall–Olkin distribution: parameter estimation and application to cancer data. Journal of Applied Statistics, 44(12):2238–2250, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ozkale:2017:ERE

[3478] M. Revan Özkale and Funda Can. An evaluation of ridge estimator in linear mixed models: an example from kidney failure data. Journal of Applied Statistics, 44(12):2251–2269, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2017:MLR

[3479] Jiajia Chen, Xiaoqin Zhang, and Shengjia Li. Multiple linear regression with compositional response and covariates. Journal of Applied Statistics, 44(12):2270–2285, December 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ara:2017:SMW

[3480] Anderson Ara, Francisco Louzada, and Carlos A. R. Diniz. Statistical monitoring of a web server for error rates: a bivariate time-series copula-based modeling approach. Journal of Applied Statistics, 44(13):2287–2300, 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lo:2017:JMB

[3481] Yungtai Lo. Joint modeling of bottle use, daily milk intake from bottles, and daily energy intake in toddlers. Journal of Applied Statistics, 44(13):2301–2316, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pasanen:2017:EST

[3482] Leena Pasanen, Päivi Laukkanen-Nevala, Ilkka Launonen, Sergey Prusov, Lasse Holmström, Eero Niemelä, and Jaakko Erkinaro. Extraction of sea temperature in the Barents Sea by a scale space multiresolution method — prospects for Atlantic salmon. Journal of Applied Statistics, 44(13):2317–2336, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2017:CAS

[3483] Haiyong Zhang and Xinyu Wang. Combined asymmetric spatial weights matrix with application to housing prices. Journal of Applied Statistics, 44(13): 2337–2353, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2017:JMS

[3484] Tao Lu. Jointly modeling skew longitudinal survival data with missingness and mismeasured covariates. Journal of Applied Statistics, 44(13):2354–2367, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Asgharzadeh:2017:GBE

[3485] A. Asgharzadeh, Hassan S. Bakouch, and M. Habibi. A generalized binomial exponential 2 distribution: modeling and applications to hydrologic events. Journal of Applied Statistics, 44(13):2368–2387, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zeng:2017:LFS

[3486] Bilin Zeng, Xuerong Meggie Wen, and Lixing Zhu. A link-free sparse group variable selection method for singleindex model. Journal of Applied Statistics, 44(13):2388–2400, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jalilian:2017:MCS

[3487] Abdollah Jalilian. Modelling and classification of species abundance: a case study in the Barro Colorado Island plot. Journal of Applied Statistics, 44(13):2401–2409, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

FerrazdoNascimento:2017:BMM

[3488] Fernando Ferraz do Nascimento and Wyara Vanesa Moura e Silva. A Bayesian model for multiple change point to extremes, with application to environmental and financial data. Journal of Applied Statistics, 44(13): 2410–2426, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pellegrini:2017:MPZ

[3489] Tiago R. Pellegrini, M. Tariqul Hasan, and Renjun Ma. Modeling of paired zero-inflated continuous data without breaking down paired designs. Journal of Applied Statistics, 44(13):2427–2443, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ekiz:2017:ODM

[3490] Meltem Ekiz and O. Ufuk Ekiz. Outlier detection with Mahalanobis square distance: incorporating small sample correction factor. Journal of Applied Statistics, 44(13):2444–2457, 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cakmakyapan:2017:MCB

[3491] Selen Cakmakyapan and Haydar Demirhan. A Monte Carlo-based pseudo-coefficient of determination for generalized linear models with binary outcome. Journal of Applied Statistics, 44(14):2458–2482, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2017:MEI

[3492] Xiaoqin Wang, Weimin Ye, and Li Yin. Measuring and estimating the interaction between exposures on a dichotomous outcome for observational studies. Journal of Applied Statistics, 44(14):2483–2498, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nasr:2017:PEF

[3493] Arwa Nasr, Soufiane Gasmi, and Fayçal Ben Hmida. Parameter estimation of the flexible Weibull distribution for type i censored samples. Journal of Applied Statistics, 44(14):2499–2512, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Suleman:2017:FCA

[3494] Abdul Suleman. A fuzzy clustering approach to evaluate individual competencies from REFLEX data. Journal of Applied Statistics, 44(14):2513–2533, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peng:2017:MPE

[3495] Chen Peng, Maochao Xu, Shouhuai Xu, and Taizhong Hu. Modeling and predicting extreme cyber attack rates via marked point processes. Journal of Applied Statistics, 44(14):2534–2563, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2017:LRR

[3496] Rand R. Wilcox. Linear regression: robust heteroscedastic confidence bands that have some specified simultaneous probability coverage. Journal of Applied Statistics, 44(14):2564–2574, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Panahi:2017:EBT

[3497] Hanieh Panahi. Estimation of the Burr type III distribution with application in unified hybrid censored sample of fracture toughness. Journal of Applied Statistics, 44(14):2575–2592, 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2017:SRP

[3498] Xiaojie Xu. Short-run price forecast performance of individual and composite models for 496 corn cash markets. Journal of Applied Statistics, 44(14):2593–2620, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sawadogo:2017:TOB

[3499] Amadou Sawadogo, Simplice Dossou-Gbété, and Dominique Lafon. Ties in one block comparison experiments: a generalization of the Mallows–Bradley– Terry ranking model. Journal of Applied Statistics, 44(14):2621–2644, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hoegh:2017:MSM

[3500] Andrew Hoegh, Ian Crandell, Scott Klopfer, and Mike Fies. Model selection with missing covariates for policy considerations in fox enclosures. Journal of Applied Statistics, 44(15):2645– 2658, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wongrin:2017:GPL

[3501] Weerinrada Wongrin and Winai Bodhisuwan. Generalized Poisson-Lindley linear model for count data. Journal of Applied Statistics, 44(15):2659–2671, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Michis:2017:CSS

[3502] Antonis A. Michis and Guy P. Nason. Case study: shipping trend estimation and prediction via multiscale variance stabilisation. Journal of Applied Statistics, 44(15):2672–2684, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Moisheer:2017:UND

[3503] A. S. Al-Moisheer. Updating a nonlinear discriminant function estimated from a mixture of two Burr type III distributions. Journal of Applied Statistics, 44(15):2685–2696, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alawieh:2017:PMF

[3504] Hiba Alawieh, Nicolas Wicker, Baydaa Al Ayoubi, and Luc Moulinier. Penalized multidimensional fitting for protein movement detection. Journal of Applied Statistics, 44(15):2697–2715, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mirkamali:2017:GLM

[3505] Sayed Jamal Mirkamali and Mojtaba Ganjali. A general location model with zero-inflated counts and skew normal outcomes. Journal of Applied Statistics, 44(15):2716–2728, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mendes:2017:CTN

[3506] Susana Mendes, M. José Fernández-Gómez, Sónia Cotrim Marques, Miguel Ângelo Pardal, Ulisses Miranda Azeiteiro, and M. Purificación Galindo-Villardón. CO-tucker: a new method for the simultaneous analysis of a sequence of paired tables. Journal of Applied Statistics, 44(15):2729–2755, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hof:2017:SML

[3507] M. H. Hof, J. Z. Musoro, R. B. Geskus, G. H. Struijk, I. J. M. ten Berge, and A. H. Zwinderman. Simulated maximum likelihood estimation in joint models for multiple longitudinal markers and recurrent events of multiple types, in the presence of a terminal event. Journal of Applied Statistics, 44(15):2756–2777, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2017:AIC

[3508] Yang-Jin Kim. Analysis of interval censored failure time data with competing risk. Journal of Applied Statistics, 44(15):2778–2790, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mollica:2017:BBQ

[3509] Cristina Mollica and Lea Petrella. Bayesian binary quantile regression for the analysis of Bachelor-to-Master transition. Journal of Applied Statistics, 44(15):2791–2812, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baghfalaki:2017:SPM

[3510] Taban Baghfalaki, Mojtaba Ganjali, and Geert Verbeke. A shared parameter model of longitudinal measurements and survival time with heterogeneous random-effects distribution. Journal of Applied Statistics, 44(15):2813–2836, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2017:FIG

[3511] L. Feng and Y. Shi. Fractionally integrated GARCH model with tempered stable distribution: a simulation study. Journal of Applied Statistics, 44(16):2837–2857, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ayyildiz:2017:MAA

[3512] Ezgi Ayyildiz, Melih Agraz, and Vilda Purutçuoğlu. MARS as an alternative approach of Gaussian graphical model for biochemical networks. Journal of Applied Statistics, 44(16):2858– 2876, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lian:2017:EST

[3513] Tonghui Lian, Caihua Yu, Zhongqun Wang, and Zhiping Hou. The evaluation study on tourism websites: from the perspective of triangular intuitionistic fuzzy multiple attribute group decision making. Journal of Applied Statistics, 44(16):2877–2889, 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Islam:2017:ADI

[3514] M. Ataharul Islam, Rafiqul I. Chowdhury, and K. S. Sultan. Analyzing dependence in incidence of diabetes and heart problem using generalized bivariate geometric models with covariates. Journal of Applied Statistics, 44(16):2890–2907, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cho:2017:CIM

[3515] Wendy K. Tam Cho. Causal inferences from many experiments. Journal of Applied Statistics, 44(16):2908–2922, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiong:2017:ELB

[3516] Wenjun Xiong, You Su, and Juan Ding.

Empirical likelihood-based robust tests for genetic association analysis with quantitative traits. Journal of Applied Statistics, 44(16):2923–2935, 2017. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bulut:2017:ESE

[3517] Hasan Bulut and Yüksel Öner. The evaluation of socio-economic development of development agency regions in Turkey using classical and robust principal component analyses. Journal of Applied Statistics, 44(16):2936–2948, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gallardo:2017:CCR

[3518] Diego I. Gallardo, Heleno Bolfarine, and Atonio Carlos Pedroso de Lima. A clustering cure rate model with application to a sealant study. Journal of Applied Statistics, 44(16):2949–2962, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Famoye:2017:EEG

[3519] Felix Famoye and Carl Lee. Exponentiatedexponential geometric regression model. Journal of Applied Statistics, 44(16): 2963–2977, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Domma:2017:NGW

[3520] Filippo Domma, Francesca Condino, and Bozidar V. Popović. A new generalized weighted Weibull distribution with decreasing, increasing, upsidedown bathtub, N-shape and M-shape hazard rate. Journal of Applied Statistics, 44(16):2978–2993, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Groenitz:2017:VER

[3521] Heiko Groenitz. Valid estimates for repeated randomized response methods. Journal of Applied Statistics, 44(16):

2994–3010, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Montano:2017:MHB

[3522] Diego Montano. Multivariate hierarchical Bayesian models and choice of priors in the analysis of survey data. Journal of Applied Statistics, 44(16):3011– 3032, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ferreira:2017:EDS

[3523] Clécio S. Ferreira and Gilberto A.

Paula. Estimation and diagnostic for skew-normal partially linear models. Journal of Applied Statistics, 44(16): 3033–3053, 2017. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vijayabhama:2018:STV

[3524] M. Vijayabhama, R. Jaisankar, S. Varadha Raj, and K. Baranidharan. Spatialtemporal variation of casuarina spread in Cauvery delta and north eastern zone of Tamil Nadu, India: a spatial autoregressive model. Journal of Applied Statistics, 45(1):1–7, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dunder:2018:VSG

[3525] Emre Dunder, Serpil Gumustekin, and Mehmet Ali Cengiz. Variable selection in gamma regression models via artificial bee colony algorithm. Journal of Applied Statistics, 45(1):8–16, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cengiz:2018:EPE

[3526] Mehmet Ali Cengiz, Emre Dünder, and Talat Senel. Energy performance evaluation of OECD countries using Bayesian stochastic frontier analysis and Bayesian network classifiers. Journal of Applied Statistics, 45(1):17–25, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2018:LTE

[3527] Feng Li and Yiqiang Lu. Lasso-type estimation for covariate-adjusted linear model. Journal of Applied Statistics, 45(1):26–42, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kachiashvili:2018:VBS

[3528] Kartlos Joseph Kachiashvili and Archil Iveri Prangishvili. Verification in biometric systems: problems and modern methods of their solution. Journal of Applied Statistics, 45(1):43–62, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2018:SEI

[3529] Hang J. Kim, Jerome P. Reiter, and Alan F. Karr. Simultaneous editimputation and disclosure limitation for business establishment data. Journal of Applied Statistics, 45(1):63–82, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cox:2018:SC

[3530] T. F. Cox and D. S. Arnold. Simple components. Journal of Applied Statistics, 45(1):83–99, January 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2018:RRI

[3531] Rand R. Wilcox. Robust regression:

an inferential method for determining which independent variables are most important. Journal of Applied Statistics, 45(1):100–111, January 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2018:FDA

[3532] Tao Chen and Qingliang Fan. A functional data approach to model score difference process in professional basketball games. Journal of Applied Statistics, 45(1):112–127, January 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodrigues:2018:PED

[3533] Giovani Carrara Rodrigues, Francisco Louzada, and Pedro Luiz Ramos. Poisson-exponential distribution: different methods of estimation. Journal of Applied Statistics, 45(1):128– 144, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barros:2018:GTM

[3534] Michelli Barros, Manuel Galea, Víctor Leiva, and Manoel Santos-Neto. Generalized Tobit models: diagnostics and application in econometrics. Journal of Applied Statistics, 45(1):145–167, January 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DeBastiani:2018:GMR

[3535] Fernanda De Bastiani, Robert A.

Rigby, Dimitrios M. Stasinopoulous, Audrey H. M. A. Cysneiros, and Miguel A. Uribe-Opazo. Gaussian Markov random field spatial models in GAMLSS. Journal of Applied Statistics, 45(1):168–186, January 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lillo:2018:MBS

[3536] Camilo Lillo, Víctor Leiva, Orietta Nicolis, and Robert G. Aykroyd. Lmoments of the Birnbaum–Saunders distribution and its extreme value version: estimation, goodness of fit and application to earthquake data. Journal of Applied Statistics, 45(2): 187–209, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Islam:2018:CRI

[3537] S. Islam, S. Anand, M. McQueen, J. Hamid, L. Thabane, S. Yusuf, and J. Beyene. Classification rules for identifying individuals at high risk of developing myocardial infarction based on ApoB, ApoA1 and the ratio were determined using a Bayesian approach. Journal of Applied Statistics, 45(2): 210–224, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rocha:2018:STR

[3538] Francisco M. M. Rocha and Julio M. Singer. Selection of terms in random coefficient regression models. Journal of Applied Statistics, 45(2):225–242, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lalitha:2018:DPO

[3539] S. Lalitha and Pratyasha Tripathi. Detection of a pair of outliers in a sample from a Gumbel distribution with known scale parameter. Journal of Applied Statistics, 45(2):243–254, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tasdan:2018:CSM

[3540] Feridun Tasdan and Ozgur Yeniay.

A comparative simulation of multiple testing procedures in circular data problems. Journal of Applied Statistics, 45(2):255–269, February 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Danish:2018:BIR

[3541] Muhammad Yameen Danish, Irshad Ahmad Arshad, and Muhammad Aslam. Bayesian inference for the randomly censored Burr-type XII distribution. Journal of Applied Statistics, 45 (2):270–283, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oliveira:2018:UDP

[3542] Amílcar Oliveira, Teresa Oliveira, and Antonio Seijas-Macías. The uniform distribution product: an approach to the (Q, r) inventory model using R. Journal of Applied Statistics, 45(2): 284–297, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fry:2018:BHM

[3543] J. T. Fry, Andrew Hoegh, Scotland Leman, and Matthew Montesano. A Bayesian hierarchical model for estimating the cost of postponing the cyclo-cross national championships. Journal of Applied Statistics, 45(2):298–305, February 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nassiri:2018:EDS

[3544] Mohammadreza Nassiri, Mahdi Elahi Torshizi, Shahrokh Ghovvati, and Mohammad Doosti. Evaluation of different statistical methods using SAS software: an in silico approach for analysis of real-time PCR data. Journal of Applied Statistics, 45(2):306–319, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Paraiba:2018:BTB

[3545] Carolina Costa Mota Paraíba, Natalia Bochkina, and Carlos Alberto Ribeiro Diniz. Bayesian truncated beta nonlinear mixed-effects models. Journal of Applied Statistics, 45(2):320–346, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Matioli:2018:NAC

[3546] L. C. Matioli, S. R. Santos, M. Kleina, and E. A. Leite. A new algorithm for clustering based on kernel density estimation. Journal of Applied Statistics, 45(2):347–366, February 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2018:RMR

[3547] Tingting Yu and Lang Wu. Robust modelling of the relationship between CD4 and viral load for complex AIDS data. Journal of Applied Statistics, 45 (2):367–383, February 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ortega:2018:HLE

[3548] Edwin M. M. Ortega, Artur J.

Lemonte, Gauss M. Cordeiro, Vicente G. Cancho, and Fábio L. Mialhe. Heteroscedastic log-exponentiated Weibull regression model. Journal of Applied Statistics, 45(3):384–408, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2018:CGG

[3549] Bei Wang and Jeffrey R. Wilson.

Comparative GMM and GQL logistic regression models on hierarchical data. Journal of Applied Statistics, 45(3):409–425, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zamanzade:2018:EPP

[3550] Ehsan Zamanzade and M. Mahdizadeh.

Estimating the population proportion in pair ranked set sampling with application to air quality monitoring. Journal of Applied Statistics, 45(3):426– 437, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bowman:2018:BAC

[3551] Dale Bowman and E. Olusegun George.

A Bayesian analysis of clustered discrete and continuous outcomes. Journal of Applied Statistics, 45(3):438– 449, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dai:2018:TDN

[3552] Wenlin Dai, Yuejin Zhou, and Tiejun Tong. Testing discontinuities in nonparametric regression. Journal of Applied Statistics, 45(3):450–473, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2018:IMA

[3553] Shuo Wang, Sungsu Kim, and Thelge Buddika Peiris. An improved meta-analysis for analyzing cylindricaltype time series data with applications to forecasting problem in environmental study. Journal of Applied Statistics, 45(3):474–486, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Solea:2018:SLE

[3554] Eftychia Solea, Bing Li, and Aleksandra Slavković. Statistical learning on emerging economies. Journal of Applied Statistics, 45(3):487–507, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chadee:2018:DUI

[3555] Samantha Chadee and Valerie Stoute. Development of an urban intensity index to facilitate urban ecosystem studies in Trinidad and Tobago. Journal of Applied Statistics, 45(3):508–527, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bergtold:2018:ILR

[3556] Jason S. Bergtold, Elizabeth A. Yeager, and Allen M. Featherstone. Inferences from logistic regression models in the presence of small samples, rare events, nonlinearity, and multicollinearity with observational data. Journal of Applied Statistics, 45(3):528–546, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haji-Maghsoudi:2018:APM

[3557] Saiedeh Haji-Maghsoudi, Azam Rastegari, Behshid Garrusi, and Mohammad Reza Baneshi. Addressing the problem of missing data in decision tree modeling. Journal of Applied Statistics, 45(3):547–557, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deAlmeida:2018:LCP

[3558] M. F. de Almeida, Aldina Correia, and Eliana Costa e Silva. Layered clays in PP polymer dispersion: the effect of the processing conditions. Journal of Applied Statistics, 45(3):558–567, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Onicescu:2018:SES

[3559] Georgiana Onicescu, Andrew B. Lawson, Jiajia Zhang, Mulugeta Gebregziabher, Kristin Wallace, and Jan M. Eberth. Spatially explicit survival modeling for small area cancer data. Journal of Applied Statistics, 45(3): 568–585, March 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiao:2018:BCE

[3560] Yan-Ting Xiao and Zhan-Shou Chen.

Bias-corrected estimations in varyingcoefficient partially nonlinear models with measurement error in the nonparametric part. Journal of Applied Statistics, 45(4):586–603, April 2018.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2018:BFM

[3561] Qing Li, Feng Guo, Inyoung Kim, Sheila G. Klauer, and Bruce G. Simons-Morton. A Bayesian finite mixture change-point model for assessing the risk of novice teenage drivers. Journal of Applied Statistics, 45(4):604–625, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Irimata:2018:IIC

[3562] Kyle M. Irimata and Jeffrey R. Wilson. Identifying intraclass correlations necessitating hierarchical modeling. Journal of Applied Statistics, 45(4):626–641, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Palmeros:2018:CEC

[3563] Oscar Palmeros, Jose A. Villaseñor, and Elizabeth González. On computing estimates of a change-point in the Weibull regression hazard model. Journal of Applied Statistics, 45(4):642– 648, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Casadei:2018:OBA

[3564] Diego Casadei, Cornelius Grunwald, Kevin Kröninger, and Florian Mentzel. Objective Bayesian analysis of counting experiments with correlated sources of background. Journal of Applied Statistics, 45(4):649–667, April 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilms:2018:AMG

[3565] I. Wilms and C. Croux. An algorithm for the multivariate group lasso with covariance estimation. Journal of Applied Statistics, 45(4):668–681, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gurung:2018:ITD

[3566] Bishal Gurung, K. N. Singh, Ravindra Singh Shekhawat, and Md Yeasin. An insight into technology diffusion of tractor through Weibull growth model. Journal of Applied Statistics, 45(4):682–696, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hu:2018:DCA

[3567] Zicheng Hu, Jessica N. Lancaster, Lauren I. R. Ehrlich, and Peter Müller. Detecting T cell activation using a varying dimension Bayesian model. Journal of Applied Statistics, 45(4):697–713, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kaur:2018:ATM

[3568] Amanpreet Kaur, Ravinder Agarwal, and Amod Kumar. Adaptive threshold method for peak detection of surface electromyography signal from around shoulder muscles. Journal of Applied Statistics, 45(4):714–726, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alhusseini:2018:BCT

[3569] Fadel Hamid Hadi Alhusseini and Vasile Georgescu. Bayesian composite Tobit quantile regression. Journal of Applied Statistics, 45(4):727–739, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dolinsky:2018:RES

[3570] Ján Dolinský, Kei Hirose, and Sadanori Konishi. Readouts for echo-state networks built using locally regularized orthogonal forward regression. Journal of Applied Statistics, 45(4):740–762, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nichols:2018:IAE

[3571] K. Nichols, E. Trevino, N. Ikeda, D. Philo, A. Garcia, and D. Bowman. Interdependency amongst earthquake magnitudes in Southern California. Journal of Applied Statistics, 45(4):763–774, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Espejo:2018:BRP

[3572] Mariano Ruiz Espejo and Adalbert Marqués Vilallonga. Book review: Principles of Scientific Methods[, by Mark Chang, ISBN 1-4822-3810-1 (ebook)]. Journal of Applied Statistics, 45(4):775–776, April 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Atkinson:2018:CDC

[3573] Anthony C. Atkinson, Marco Riani, and Andrea Cerioli. Cluster detection and clustering with random start forward searches. Journal of Applied Statistics, 45(5):777–798, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Choo-Wosoba:2018:ACC

[3574] Hyoyoung Choo-Wosoba and Somnath Datta. Analyzing clustered count data with a cluster-specific random effect zero-inflated Conway–Maxwell– Poisson distribution. Journal of Applied Statistics, 45(5):799–814, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alhamzawi:2018:BQR

[3575] Rahim Alhamzawi and Haithem Taha Mohammad Ali. Bayesian quantile regression for ordinal longitudinal data. Journal of Applied Statistics, 45(5):815–828, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Waken:2018:FES

[3576] R. J. Waken, Joon Jin Song, Soohyun Kwon, Ki-Hong Min, and GyuWon Lee. A flexible and efficient spatial interpolator for radar rainfall estimation. Journal of Applied Statistics, 45(5):829–844, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Man:2018:CAJ

[3577] Georg Man. Critical appraisal of jointness concepts in Bayesian model averaging: evidence from life sciences, sociology, and other scientific fields. Journal of Applied Statistics, 45(5):845– 867, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Flores:2018:HTF

[3578] Ramón Flores, Rosa Lillo, and Juan Romo. Homogeneity test for functional data. Journal of Applied Statistics, 45(5):868–883, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pierdzioch:2018:DAI

[3579] Christian Pierdzioch, Monique B. Reid, and Rangan Gupta. On the directional accuracy of inflation forecasts: evidence from South African survey data. Journal of Applied Statistics, 45(5):884–900, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ye:2018:FPC

[3580] Jun Ye, Juan Xi, and Richard L. Einsporn. Functional principal component analysis in age–period–cohort analysis of body mass index data by gender and ethnicity. Journal of Applied Statistics, 45(5):901–917, May 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goldstein:2018:BMM

[3581] Harvey Goldstein, William J. Browne, and Christopher Charlton. A Bayesian model for measurement and misclassification errors alongside missing data, with an application to higher education participation in Australia. Journal of Applied Statistics, 45(5):918–931, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Desousa:2018:TBS

[3582] Mário F. Desousa, Helton Saulo, Víctor Leiva, and Paulo Scalco. On a tobit-Birnbaum–Saunders model with an application to medical data. Journal of Applied Statistics, 45(5):932–955, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2018:GIS

[3583] Wei-Ya Wu, Wei-Hwa Wu, Hsin-Neng Hsieh, and Meng-Chih Lee. The generalized inference on the sign testing problem about the normal variances. Journal of Applied Statistics, 45(5):956–970, May 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yoon:2018:TNM

[3584] Jisu Yoon and Tatyana Krivobokova.

Treatments of non-metric variables in partial least squares and principal component analysis. Journal of Applied Statistics, 45(6):971–987, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mallick:2018:BBR

[3585] Himel Mallick and Nengjun Yi.

Bayesian bridge regression. Journal of Applied Statistics, 45(6):988–1008, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mullensiefen:2018:UCR

[3586] Daniel Müllensiefen, Christian Hennig, and Hedie Howells. Using clustering of rankings to explain brand preferences with personality and sociodemographic variables. Journal of Applied Statistics, 45(6):1009–1029, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ogundimu:2018:PPP

[3587] Emmanuel O. Ogundimu and Gary S.

Collins. Predictive performance of penalized beta regression model for continuous bounded outcomes. Journal of Applied Statistics, 45(6):1030–1040, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mokhtar:2018:CAD

[3588] Nurkhairany Amyra Mokhtar, Yong Zulina Zubairi, and Abdul Ghapor Hussin. A clustering approach to detect multiple outliers in linear functional relationship model for circular data. Journal of Applied Statistics, 45(6):1041–1051, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Smith:2018:BAB

[3589] Adam L. Smith and Sofía S. Villar. Bayesian adaptive bandit-based designs using the Gittins index for multiarmed trials with normally distributed endpoints. Journal of Applied Statistics, 45(6):1052–1076, June 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2018:VSM

[3590] Yingzhen Chen, Xuejun Ma, and Jingke Zhou. Variable selection for mode regression. Journal of Applied Statistics, 45(6):1077–1084, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haselimashhadi:2018:NBR

[3591] H. Haselimashhadi, V. Vinciotti, and K. Yu. A novel Bayesian regression model for counts with an application to health data. Journal of Applied Statistics, 45(6):1085–1105, June 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Makinde:2018:MDC

[3592] Olusola Samuel Makinde and Olusoga Akin Fasoranbaku. On maximum depth classifiers: depth distribution approach. Journal of Applied Statistics, 45(6):1106–1117, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kvaalseth:2018:MAB

[3593] Tarald O. Kvålseth. Measuring association between nominal categorical variables: an alternative to the Goodman– Kruskal lambda. Journal of Applied Statistics, 45(6):1118–1132, June 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hammer:2018:SMS

[3594] Hugo Lewi Hammer. Statistical models for short- and long-term forecasts of snow depth. Journal of Applied Statistics, 45(6):1133–1156, June 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Betancourt:2018:ICF

[3595] Brenda Betancourt, Abel Rodríguez, and Naomi Boyd. Investigating competition in financial markets: a sparse autologistic model for dynamic network data. Journal of Applied Statistics, 45(7):1157–1172, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patino:2018:WIP

[3596] Jairo A. Fúquene Patiño, Brenda Betancourt, and João B. M. Pereira. A weakly informative prior for Bayesian dynamic model selection with applications in fMRI. Journal of Applied Statistics, 45(7):1173–1192, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sarnacchiaro:2018:SRM

[3597] Pasquale Sarnacchiaro and Flavio Boccia. Some remarks on measurement models in the structural equation model: an application for socially responsible food consumption. Journal of Applied Statistics, 45(7):1193–1208, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jamalinia:2018:DCE

[3598] Hamid Jamalinia, Saber Khalouei, Vahideh Rezaie, Samad Nejatian, Karamolah Bagheri-Fard, and Hamid Parvin. Diverse classifier ensemble creation based on heuristic dataset modification. Journal of Applied Statistics, 45(7):1209–1226, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prematilake:2018:EPP

[3599] Chalani Prematilake and Leif Ellingson. Evaluation and prediction of polygon approximations of planar contours for shape analysis. Journal of Applied Statistics, 45(7):1227–1246, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Famoye:2018:MZI

[3600] Felix Famoye and John S. Preisser. Marginalized zero-inflated generalized Poisson regression. Journal of Applied Statistics, 45(7):1247–1259, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2018:BRM

[3601] Li-An Lin, Sheng Luo, and Barry R.

Davis. Bayesian regression model for recurrent event data with eventvarying covariate effects and event effect. Journal of Applied Statistics, 45 (7):1260–1276, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zimmer:2018:HII

[3602] David M. Zimmer. The heterogeneous impact of insurance on health care demand among young adults: a panel data analysis. Journal of Applied Statistics, 45(7):1277–1291, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Madden:2018:ANR

[3603] Gary Madden, Nicholas Apergis, Paul Rappoport, and Aniruddha Banerjee. An application of nonparametric regression to missing data in large market surveys. Journal of Applied Statistics, 45(7):1292–1302, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramires:2018:FSR

[3604] Thiago G. Ramires, Edwin M. M. Ortega, Niel Hens, Gauss M. Cordeiro, and Gilberto A. Paula. A flexible semiparametric regression model for bimodal, asymmetric and censored data. Journal of Applied Statistics, 45(7):1303–1324, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yuan:2018:ACM

[3605] Ao Yuan, Yuan Guo, Nawar M. Shara, Barbara V. Howard, and Ming T. Tan.

An additive Cox model for coronary heart disease study. Journal of Applied Statistics, 45(7):1325–1346, July 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DiPalma:2018:RPM

[3606] M. A. Di Palma, P. Filzmoser, M. Gallo, and K. Hron. A robust Parafac model for compositional data. Journal of Applied Statistics, 45(8): 1347–1369, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shim:2018:SCL

[3607] Jooyong Shim, Changha Hwang, Sunjoo Jeong, and Insuk Sohn. Semivarying coefficient least-squares support vector regression for analyzing highdimensional gene-environmental data. Journal of Applied Statistics, 45(8): 1370–1381, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mignone:2018:DOP

[3608] Flavio Mignone and Fabio Rapallo. Detection of outlying proportions. Journal of Applied Statistics, 45(8):1382– 1395, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Casin:2018:CML

[3609] Philippe Casin. Categorical multiblock linear discriminant analysis. Journal of Applied Statistics, 45(8):1396–1409, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2018:THD

[3610] Xi Shen and Chang-Xing Ma. Testing homogeneity of difference of two proportions for stratified correlated paired binary data. Journal of Applied Statistics, 45(8):1410–1425, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jayalath:2018:AMA

[3611] Kalanka P. Jayalath and Hon Keung Tony Ng. Analysis of means approach for random factor analysis. Journal of Applied Statistics, 45(8): 1426–1446, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luca:2018:MCS

[3612] Stijn Luca. Modified chain sampling plans for lot inspection by variables and attributes. Journal of Applied Statistics, 45(8):1447–1464, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chipperfield:2018:SQD

[3613] James O. Chipperfield, Margo L. Barr, and David. G. Steel. Split questionnaire designs: collecting only the data that you need through MCAR and MAR designs. Journal of Applied Statistics, 45(8):1465–1475, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shojaei:2018:GAD

[3614] S. Reza H. Shojaei, Yadollah Waghei, and Mohsen Mohammadzadeh. Geostatistical analysis of disease data: a case study of tuberculosis incidence in Iran. Journal of Applied Statistics, 45 (8):1476–1483, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scollnik:2018:BAQ

[3615] David P. M. Scollnik. Bayesian analysis of a quarantine inspection model. Journal of Applied Statistics, 45(8): 1484–1496, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Allaj:2018:TSM

[3616] Erindi Allaj. Two simple measures of variability for categorical data. Journal of Applied Statistics, 45(8):1497–1516, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lui:2018:SSD

[3617] Kung-Jong Lui. Sample size determination for testing equality in frequency data under an incomplete block crossover design. Journal of Applied Statistics, 45(8):1517–1529, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pandit:2018:MCS

[3618] Mahesh Pandit, Krishna P. Paudel, and Roger Hinson. Market channel selections by US nursery plant producers: a multivariate nonparametric fractional regression analysis. Journal of Applied Statistics, 45(8):1530–1546, August 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhou:2018:BMP

[3619] Qi Zhou, Yoo-Mi Chin, James D. Stamey, and Joon Jin Song. Bayesian misclassification and propensity score methods for clustered observational studies. Journal of Applied Statistics, 45(9):1547–1560, September 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schauberger:2018:AIF

[3620] Gunther Schauberger, Andreas Groll, and Gerhard Tutz. Analysis of the importance of on-field covariates in the German Bundesliga. Journal of Applied Statistics, 45(9):1561–1578, September 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kharroubi:2018:VEH

[3621] Samer A. Kharroubi. Valuations of EQ-5D health states: could United Kingdom results be used as informative priors for the United States. Journal of Applied Statistics, 45(9): 1579–1594, September 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Spezia:2018:SHM

[3622] Luigi Spezia, Nial Friel, and Alessandro Gimona. Spatial hidden Markov models and species distributions. Journal of Applied Statistics, 45(9):1595–1615, September 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Malhotra:2018:RVB

[3623] Akash Malhotra and Shailesh Krishna.

Release velocities and bowler performance in cricket. Journal of Applied Statistics, 45(9):1616–1627, September 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kidwell:2018:DAC

[3624] Kelley M. Kidwell, Nicholas J. Seewald, Qui Tran, Connie Kasari, and Daniel Almirall. Design and analysis considerations for comparing dynamic treatment regimens with binary outcomes from sequential multiple assignment randomized trials. Journal of Applied Statistics, 45(9):1628–1651, September 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zou:2018:EBE

[3625] Yajie Zou, John E. Ash, Byung-Jung Park, Dominique Lord, and Lingtao Wu. Empirical Bayes estimates of finite mixture of negative binomial regression models and its application to highway safety. Journal of Applied Statistics, 45 (9):1652–1669, September 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Morais:2018:UCD

[3626] Joanna Morais, Christine Thomas-Agnan, and Michel Simioni. Using compositional and Dirichlet models for market share regression. Journal of Applied Statistics, 45(9):1670–1689, September 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kaminski:2018:PPR

[3627] Bogumil Kamiński and Przemyslaw Szufel. On parallel policies for ranking and selection problems. Journal of Applied Statistics, 45(9):1690–1713, September 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2018:USR

[3628] W. Tang, H. He, W. J. Wang, and D. G. Chen. Untangle the structural and random zeros in statistical modelings. Journal of Applied Statistics, 45 (9):1714–1733, September 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2018:NBA

[3629] Chun-Xia Zhang, Jiang-She Zhang, Guan-Wei Wang, and Nan-Nan Ji. A novel bagging approach for variable ranking and selection via a mixed importance measure. Journal of Applied Statistics, 45(10):1734–1755, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Falley:2018:BEL

[3630] Brandi N. Falley, James D. Stamey, and A. Alexander Beaujean. Bayesian estimation of logistic regression with misclassified covariates and response. Journal of Applied Statistics, 45(10): 1756–1769, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hsieh:2018:QRL

[3631] Jin-Jian Hsieh and Jian-Lin Wang.

Quantile residual life regression based on semi-competing risks data. Journal of Applied Statistics, 45(10):1770–1780, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Loquiha:2018:JMM

[3632] Osvaldo Loquiha, Niel Hens, Emilia Martins-Fonteyn, Herman Meulemans, Edwin Wouters, Marleen Temmerman, Nafissa Osman, and Marc Aerts. Joint models for mixed categorical outcomes: a study of HIV risk perception and disease status in Mozambique. Journal of Applied Statistics, 45(10):1781–1798, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khan:2018:GBC

[3633] Shahedul A. Khan and Setu C. Kar. Generalized bent-cable methodology for changepoint data: a Bayesian approach. Journal of Applied Statistics, 45(10):1799–1812, October 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alencar:2018:SHD

[3634] Airlane P. Alencar. Seasonality of hospitalizations due to respiratory diseases: modelling serial correlation all we need is Poisson. Journal of Applied Statistics, 45(10):1813–1822, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Paez:2018:MLC

[3635] M. S. Paez and S. G. Walker. Modeling with a large class of unimodal multivariate distributions. Journal of Applied Statistics, 45(10):1823–1845, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goncalves:2018:BIQ

[3636] D. S. Gonçalves, C. L. N. Azevedo, C. Lavor, and M. A. Gomes-Ruggiero. Bayesian inference for quantum state tomography. Journal of Applied Statistics, 45(10):1846–1871, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghodsi:2018:VRS

[3637] Mansi Ghodsi, Hossein Hassani, Donya Rahmani, and Emmanuel Sirimal Silva. Vector and recurrent singular spectrum analysis: which is better at forecasting? Journal of Applied Statistics, 45(10):1872–1899, October 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Du:2018:MAB

[3638] Jiang Du, Xiuping Chen, Eddy Kwessi, and Zhimeng Sun. Model averaging based on rank. Journal of Applied Statistics, 45(10):1900–1919, October 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Duarte:2018:MSY

[3639] Gislaine V. Duarte, Altemir Braga, Daniel L. Miquelluti, and Vitor A. Ozaki. Modeling of soybean yield using symmetric, asymmetric and bimodal distributions: implications for crop insurance. Journal of Applied Statistics, 45(11):1920–1937, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2018:MSD

[3640] Puyu Wang, Hai Zhang, and Yong Liang. Model selection with distributed SCAD penalty. Journal of Applied Statistics, 45(11):1938–1955, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Swanson:2018:HTN

[3641] D. M. Swanson, C. D. Anderson, and R. A. Betensky. Hypothesis tests for Neyman’s bias in case-control studies. Journal of Applied Statistics, 45(11):

1956–1977, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Su:2018:SAL

[3642] Wen Su and Hangjin Jiang. Semiparametric analysis of longitudinal data with informative observation times and censoring times. Journal of Applied Statistics, 45(11):1978–1993, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2018:MSL

[3643] Yunlu Jiang, Yu Conglian, and Ji Qinghua. Model selection for the localized mixture of experts models. Journal of Applied Statistics, 45(11): 1994–2006, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ko:2018:DIS

[3644] Feng shou Ko. Discussion on the issue of sample size determination for a targeted to an untargeted and to a mixed effect model-based clinical trial design. Journal of Applied Statistics, 45(11): 2007–2019, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coscia:2018:GMC

[3645] Cristina Coscia, Roberto Fontana, and Patrizia Semeraro. Graphical models for complex networks: an application to Italian museums. Journal of Applied Statistics, 45(11):2020–2038, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mattos:2018:LBI

[3646] Thalita do Bem Mattos, Aldo M.

Garay, and Victor H. Lachos. Likelihoodbased inference for censored linear regression models with scale mixtures of skew-normal distributions. Journal of Applied Statistics, 45(11):2039–2066, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2018:RIM

[3647] Jiajia Chen, Xiaoqin Zhang, Karel Hron, Matthias Templ, and Shengjia Li. Regression imputation with Qmode clustering for rounded zero replacement in high-dimensional compositional data. Journal of Applied Statistics, 45(11):2067–2080, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coelho-Barros:2018:IPL

[3648] Emílio A. Coelho-Barros, Josmar Mazucheli, Jorge A. Achcar, Kelly Vanessa Parede Barco, and José Rafael Tovar Cuevas. The inverse power Lindley distribution in the presence of left-censored data. Journal of Applied Statistics, 45(11):2081–2094, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shirazi:2018:PHC

[3649] Aliakbar Mastani Shirazi and Aluisio Pinheiro. A proportional hazard cure model for ordinal responses by selfmodeling regression. Journal of Applied Statistics, 45(11):2095–2106, November 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Manso:2018:MLE

[3650] Rubén Manso, Rafael Calama, Marta Pardos, and Mathieu Fortin. A maximum likelihood estimator for lefttruncated lifetimes based on probabilistic prior information about time of occurrence. Journal of Applied Statistics, 45(12):2107–2127, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2018:VCS

[3651] Xin Zhao and Zhensheng Huang. Varying-coefficient single-index measurement error model. Journal of Applied Statistics, 45(12):2128–2144, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kwak:2018:BIC

[3652] Sang Gyu Kwak, Balgobin Nandram, and Dal Ho Kim. Bayesian inference on contingency tables with uncertainty about independence for small areas. Journal of Applied Statistics, 45(12): 2145–2163, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2018:HLR

[3653] Jiajia Chen, Xiaoqin Zhang, and Shengjia Li. Heteroskedastic linear regression model with compositional response and covariates. Journal of Applied Statistics, 45(12):2164–2181, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahabadi:2018:SME

[3654] S. Eftekhari Mahabadi and E. Rahimi Jafari. Skew-mixed effects model for multivariate longitudinal data with categorical outcomes and missingness. Journal of Applied Statistics, 45(12): 2182–2201, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ausloos:2018:IYS

[3655] Marcel Ausloos and Roy Cerqueti. Intriguing yet simple skewness: kurtosis relation in economic and demographic data distributions, pointing to preferential attachment processes. Journal of Applied Statistics, 45(12): 2202–2218, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zullo:2018:ECF

[3656] Anthony Zullo, Mathieu Fauvel, and Frédéric Ferraty. Experimental comparison of functional and multivariate spectral-based supervised classification methods in hyperspectral image. Journal of Applied Statistics, 45(12): 2219–2237, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lorah:2018:EIL

[3657] Julie Ann Lorah. Estimating individual-level interaction effects in multilevel models: a Monte Carlo simulation study with application. Journal of Applied Statistics, 45(12):2238–2255, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beath:2018:MBA

[3658] Ken J. Beath. A mixture-based approach to robust analysis of generalised linear models. Journal of Applied Statistics, 45(12):2256–2268, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vassilopoulos:2018:DSN

[3659] Achilleas Vassilopoulos, Andreas C. Drichoutis, Rodolfo M. Nayga Jr., and Panagiotis Lazaridis. Does the supplemental nutrition assistance program really increase obesity? The importance of accounting for misclassification errors. Journal of Applied Statistics, 45 (12):2269–2278, December 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cafri:2018:PRA

[3660] Guy Cafri, Luo Li, Elizabeth W. Paxton, and Juanjuan Fan. Predicting risk for adverse health events using random forest. Journal of Applied Statistics, 45 (12):2279–2294, December 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kirschenmann:2018:BEC

[3661] Thomas Kirschenmann, Paul Damien, and Stephen Walker. Bayesian estimation of the Cox model under different hazard rate shape assumptions via slice sampling. Journal of Applied Statistics, 45(12):2295–2306, December 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Desgagne:2018:PIA

[3662] A. Desgagné and P. Lafaye de Micheaux. A powerful and interpretable alternative to the Jarque–Bera test of normality based on 2nd-power skewness and kurtosis, using the Rao’s score test on the APD family. Journal of Applied Statistics, 45(13):2307–2327, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bokusheva:2018:UCR

[3663] Raushan Bokusheva. Using copulas for rating weather index insurance contracts. Journal of Applied Statistics, 45(13):2328–2356, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghebremichael:2018:ETI

[3664] Musie Ghebremichael and Semere Habtemicael. Effect of tuberculosis on immune restoration among HIVinfected patients receiving antiretroviral therapy. Journal of Applied Statistics, 45(13):2357–2364, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhatta:2018:BTI

[3665] Dilli Bhatta, Balgobin Nandram, and Joseph Sedransk. Bayesian testing for independence of two categorical variables under two-stage cluster sampling with covariates. Journal of Applied Statistics, 45(13):2365–2393, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhati:2018:DPI

[3666] Deepesh Bhati and Sreenivasan Ravi.

Diagnostic plots for identifying max domains of attraction under power normalization. Journal of Applied Statistics, 45(13):2394–2410, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cencic:2018:DRN

[3667] Oliver Cencic and Rudolf Frühwirth.

Data reconciliation of nonnormal observations with nonlinear constraints. Journal of Applied Statistics, 45(13): 2411–2428, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2018:TSB

[3668] Jun Zhang, Jing Zhang, Xuehu Zhu, and Tao Lu. Testing symmetry based on empirical likelihood. Journal of Applied Statistics, 45(13):2429–2454, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2018:CSA

[3669] Xiaojie Xu. Causal structure among US corn futures and regional cash prices in the time and frequency domain. Journal of Applied Statistics, 45(13):2455–2480, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sulewski:2018:PAI

[3670] Piotr Sulewski. Power analysis of independence testing for three-way contingency tables of small sizes. Journal of Applied Statistics, 45(13):2481–2498, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balamurali:2018:ODR

[3671] S. Balamurali, P. Jeyadurga, and M. Usha. Optimal design of repetitive group sampling plans for Weibull and gamma distributions with applications and comparison to the Birnbaum– Saunders distribution. Journal of Applied Statistics, 45(14):2499–2520, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Du:2018:SCF

[3672] Ruofei Du and Zhide Fang. Statistical correction for functional metagenomic profiling of a microbial community with short NGS reads. Journal of Applied Statistics, 45(14):2521–2535, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wright:2018:ESE

[3673] Daniel B. Wright. Estimating school effectiveness with student growth percentile and gain score models. Journal of Applied Statistics, 45(14):2536–2547, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kowalski:2018:MVS

[3674] J. Kowalski, S. Hao, T. Chen, Y. Liang, J. Liu, L. Ge, C. Feng, and X. M. Tu. Modern variable selection for longitudinal semi-parametric models with missing data. Journal of Applied Statistics, 45(14):2548–2562, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Punzo:2018:FIE

[3675] Antonio Punzo, Angelo Mazza, and Antonello Maruotti. Fitting insurance and economic data with outliers: a flexible approach based on finite mixtures of contaminated gamma distributions. Journal of Applied Statistics, 45(14):2563–2584, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fonseca:2018:BBS

[3676] Rodney V. Fonseca and Francisco Cribari-Neto. Bimodal Birnbaum– Saunders generalized autoregressive score model. Journal of Applied Statistics, 45(14):2585–2606, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Moni:2018:PSD

[3677] Md. Moniruzzaman Moni and M. Iftakhar Alam. Patient-specific dose finding in phase i clinical trials. Journal of Applied Statistics, 45(14):2607–2618, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghasemzadeh:2018:BCM

[3678] S. Ghasemzadeh, M. Ganjali, and T. Baghfalaki. A Bayesian conditional model for bivariate mixed ordinal and skew continuous longitudinal responses using quantile regression. Journal of Applied Statistics, 45(14):2619–2642, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alshaybawee:2018:BVS

[3679] Taha Alshaybawee, Rahim Alhamzawi, Habshah Midi, and Intisar Ibrahim Allyas. Bayesian variable selection and coefficient estimation in heteroscedastic linear regression model. Journal of Applied Statistics, 45(14):2643–2657, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reid:2018:GFS

[3680] Stephen Reid, Aaron M. Newman, Maximilian Diehn, Ash A. Alizadeh, and Robert Tibshirani. Genomic feature selection by coverage design optimization. Journal of Applied Statistics, 45(14):2658–2676, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2018:KPC

[3681] Jihwan Oh, Faye Zheng, R. W. Doerge, and Hyonho Chun. Kernel partial correlation: a novel approach to capturing conditional independence in graphical models for noisy data. Journal of Applied Statistics, 45(14):2677–2696, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

SilvaeSilva:2018:CPM

[3682] Luís G. Silva e Silva and Renato M.

Assunção. COWORDS: a probabilistic model for multiple word clouds. Journal of Applied Statistics, 45(15):2697– 2717, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peng:2018:MMC

[3683] Chen Peng, Maochao Xu, Shouhuai Xu, and Taizhong Hu. Modeling multivariate cybersecurity risks. Journal of Applied Statistics, 45(15):2718–2740, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Perez-Ruiz:2018:JRM

[3684] Luis Carlos Pérez-Ruiz and Gabriel Escarela. Joint regression modeling for missing categorical covariates in generalized linear models. Journal of Applied Statistics, 45(15):2741–2759, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bose:2018:PPM

[3685] Mausumi Bose and Kajal Dihidar. Privacy protection measures for randomized response surveys on stigmatizing continuous variables. Journal of Applied Statistics, 45(15):2760–2772, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Matthews:2018:CML

[3686] Gregory J. Matthews, Juliet K. Brophy, Maxwell Luetkemeier, Hongie Gu, and George K. Thiruvathukal. A comparison of machine learning techniques for taxonomic classification of teeth from the family Bovidae. Journal of Applied Statistics, 45(15):2773–2787, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mirshahi:2018:AOP

[3687] Mina Mirshahi, Vahid Partovi Nia, and Luc Adjengue. Automatic odor prediction for electronic nose. Journal of Applied Statistics, 45(15):2788–2799, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ju:2018:RPE

[3688] Cheng Ju, Aurélien Bibaut, and Mark van der Laan. The relative performance of ensemble methods with deep convolutional neural networks for image classification. Journal of Applied Statistics, 45(15):2800–2818, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aboagye-Sarfo:2018:SAE

[3689] Patrick Aboagye-Sarfo and Qun Mai. Seasonal analysis of emergency department presentations in Western Australia, 2009/10–2014/15. Journal of Applied Statistics, 45(15):2819–2830, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fischer:2018:SMP

[3690] Svenja Fischer. A seasonal mixed-POT model to estimate high flood quantiles from different event types and seasons. Journal of Applied Statistics, 45(15):2831–2847, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rocha:2018:PIH

[3691] Conceição Rocha and Pedro Quelhas Brito. Profiles identification on hierarchical tree structure data sets. Journal of Applied Statistics, 45(15):2848–2863, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hapfelmeier:2018:SIR

[3692] Alexander Hapfelmeier, Kurt Ulm, and Bernhard Haller. Subgroup identification by recursive segmentation. Journal of Applied Statistics, 45(15):2864–2887, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhu:2018:VDM

[3693] Xuwen Zhu. Variable diagnostics in model-based clustering through variation partition. Journal of Applied Statistics, 45(16):2888–2905, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhattacharya:2018:BSA

[3694] Durba Bhattacharya and Sourabh Bhattacharya. A Bayesian semiparametric approach to learning about gene-gene interactions in case-control studies. Journal of Applied Statistics, 45(16):2906–2928, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2018:IAL

[3695] Dal Ho Kim, Jayoun Lee, and Yongku Kim. Information analysis of local suppression scheme based on a spatialtemporal model. Journal of Applied Statistics, 45(16):2929–2942, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsionas:2018:BCS

[3696] Mike G. Tsionas and Marwan Izzeldin. Bayesian CV@R/super-quantile regression. Journal of Applied Statistics, 45(16):2943–2957, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rosenkranz:2018:EBE

[3697] Gerd K. Rosenkranz. Empirical Bayes estimators in hierarchical models with mixture priors. Journal of Applied Statistics, 45(16):2958–2980, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Andrade:2018:EKE

[3698] J. A. A. Andrade and J. P. Gosling. Expert knowledge elicitation using item response theory. Journal of Applied Statistics, 45(16):2981–2998, 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2018:HBN

[3699] Jae-Cheol Park and Jea-Young Lee. How to build nomogram for type 2 diabetes using a naïve Bayesian classifier technique. Journal of Applied Statistics, 45(16):2999–3011, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2018:MAA

[3700] Longmei Chen, Alan T. K. Wan, Geoffrey Tso, and Xinyu Zhang. A model averaging approach for the ordered probit and nested logit models with applications. Journal of Applied Statistics, 45(16):3012–3052, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2018:LR

[3701] Anonymous. List of reviewers. Journal of Applied Statistics, 45(16):3053–3056, 2018. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Righetto:2019:PWI

[3702] Ana J. Righetto, Thiago G. Ramires, Luiz R. Nakamura, Pedro L. D. B. Castanho, Christel Faes, and Taciana V. Savian. Predicting weed invasion in a sugarcane cultivar using multispectral image. Journal of Applied Statistics, 46(1):1–12, January 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahn:2019:DBO

[3703] Jeongyoun Ahn, Myung Hee Lee, and Jung Ae Lee. Distance-based outlier detection for high dimension, low sample size data. Journal of Applied Statistics, 46(1):13–29, January 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mitra:2019:BHM

[3704] Riten Mitra, Ryan Gill, Sinjini Sikdar, and Susmita Datta. Bayesian hierarchical model for protein identifications. Journal of Applied Statistics, 46(1):30–46, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Godichon-Baggioni:2019:CTC

[3705] Antoine Godichon-Baggioni, Cathy Maugis-Rabusseau, and Andrea Rau.

Clustering transformed compositional data using K-means, with applications in gene expression and bicycle sharing system data. Journal of Applied Statistics, 46(1):47–65, January 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cekic:2019:MBS

[3706] Sezen Cekic, Didier Grandjean, and Olivier Renaud. Multiscale Bayesian state-space model for Granger causality analysis of brain signal. Journal of Applied Statistics, 46(1):66–84, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Makri:2019:CRO

[3707] Frosso S. Makri, Zaharias M. Psillakis, and Anastasios N. Arapis. On the concentration of runs of ones of length exceeding a threshold in a Markov chain. Journal of Applied Statistics, 46(1):85–100, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bourguignon:2019:EPP

[3708] Marcelo Bourguignon, Josemar Rodrigues, and Manoel Santos-Neto. Extended Poisson INAR(1) processes with equidispersion, underdispersion and overdispersion. Journal of Applied Statistics, 46(1):101–118, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bercu:2019:NSP

[3709] Bernard Bercu, Sami Capderou, and Gilles Durrieu. A nonparametric statistical procedure for the detection of marine pollution. Journal of Applied Statistics, 46(1):119–140, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gomez:2019:RLM

[3710] Yolanda M. Gómez, Diego I. Gallardo, Yuri Iriarte, and Heleno Bolfarine. The Rayleigh–Lindley model: properties and applications. Journal of Applied Statistics, 46(1):141–163, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jaeger:2019:SCM

[3711] Byron C. Jaeger, Lloyd J. Edwards, and Matthew J. Gurka. An R^{2} statistic for covariance model selection in the linear mixed model. Journal of Applied Statistics, 46(1):164–184, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2019:Ca

[3712] Anonymous. Corrigendum. Journal of Applied Statistics, 46(1):185–186, January 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2019:RAI

[3713] Da Xu, Shishun Zhao, Tao Hu, Mengzhu Yu, and Jianguo Sun. Regression analysis of informative current status data with the semiparametric linear transformation model. Journal of Applied Statistics, 46(2):187–202, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Woldu:2019:AFD

[3714] Henok Woldu, Timothy G. Heckman, Andreas Handel, and Ye Shen. Applying functional data analysis to assess tele-interpersonal psychotherapy’s efficacy to reduce depression. Journal of Applied Statistics, 46(2):203–216, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mashford:2019:BHS

[3715] John Mashford, Yong Song, Q. J. Wang, and David Robertson. A Bayesian hierarchical spatio-temporal rainfall model. Journal of Applied Statistics, 46(2):217–229, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2019:SBM

[3716] Xiao Li, Michele Guindani, Chaan S. Ng, and Brian P. Hobbs. Spatial Bayesian modeling of GLCM with application to malignant lesion characterization. Journal of Applied Statistics, 46(2):230–246, February 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chien:2019:MCV

[3717] Li-Chu Chien. A method for combining p-values in meta-analysis by gamma distributions. Journal of Applied Statistics, 46(2):247–261, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peng:2019:CWC

[3718] J. Peng, W. Liu, F. Bretz, and A. J. Hayter. Counting by weighing: construction of two-sided confidence intervals. Journal of Applied Statistics, 46 (2):262–271, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gold:2019:PSC

[3719] David Gold, Lixin Lang, and Kim Zerba. Practical statistical considerations for investigating anti-tumor treatments in mice. Journal of Applied Statistics, 46(2):272–285, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Russell:2019:IPE

[3720] Brook T. Russell. Investigating precipitation extremes in south Carolina with focus on the state’s October 2015 precipitation event. Journal of Applied Statistics, 46(2):286–303, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sunecher:2019:CSM

[3721] Yuvraj Sunecher, Naushad Mamode Khan, and Vandna Jowaheer. A case study of MCB and SBMH stock transaction using a novel BINMA(1) with non-stationary NB correlated innovations. Journal of Applied Statistics, 46 (2):304–323, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Drosou:2019:SIS

[3722] K. Drosou, C. Koukouvinos, and A. Lappa. Sure independence screening for real medical Poisson data. Journal of Applied Statistics, 46(2): 324–350, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Montgomery:2019:EPC

[3723] R. N. Montgomery, A. S. Watts, N. C. Burns, E. D. Vidoni, and J. D.

Mahnken. Evaluating paired categorical data when the pairing is lost. Journal of Applied Statistics, 46(2): 351–363, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Perry:2019:DTC

[3724] Marcus B. Perry. On the detection of transitive clusters in undirected networks. Journal of Applied Statistics, 46 (2):364–384, February 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Phillips:2019:DMA

[3725] Caleb Phillips, Ryan Elmore, Jenny Melius, Pieter Gagnon, and Robert Margolis. A data mining approach to estimating rooftop photovoltaic potential in the US. Journal of Applied Statistics, 46(3):385–394, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodrigues:2019:ANH

[3726] Eliane R. Rodrigues, Mario H. Tarumoto, and Guadalupe Tzintzun. Application of a non-homogeneous Markov chain with seasonal transition probabilities to ozone data. Journal of Applied Statistics, 46(3):395–415, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nakamura:2019:NCD

[3727] Luiz R. Nakamura, Pedro H. R. Cerqueira, Thiago G. Ramires, Rodrigo R. Pescim, R. A. Rigby, and Dimitrios M. Stasinopoulos. A new continuous distribution on the unit interval applied to modelling the points ratio of football teams. Journal of Applied Statistics, 46(3):416–431, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2019:FGP

[3728] Haiqing Chen, Weihu Cheng, Yaohua Rong, and Xu Zhao. Fitting the generalized Pareto distribution to data based on transformations of order statistics. Journal of Applied Statistics, 46(3):432–448, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deOliveira:2019:DCB

[3729] Ricardo Puziol de Oliveira, Jorge Alberto Achcar, Danielle Peralta, and Josmar Mazucheli. Discrete and continuous bivariate lifetime models in presence of cure rate: a comparative study under Bayesian approach. Journal of Applied Statistics, 46(3):449–467, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhan:2019:SSC

[3730] Yihong Zhan, Yanan Zhang, Jiajia Zhang, Bo Cai, and James W. Hardin. Sample size calculation for a proportional hazards mixture cure model with nonbinary covariates. Journal of Applied Statistics, 46(3):468–483, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scudilio:2019:DMI

[3731] Juliana Scudilio, Vinicius F. Calsavara, Ricardo Rocha, Francisco Louzada, Vera Tomazella, and Agatha S. Rodrigues. Defective models induced by gamma frailty term for survival data with cured fraction. Journal of Applied Statistics, 46(3):484–507, March 2019.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2019:LSS

[3732] Wei Wang and Zhuo Fang. Linear scalar-on-surface random effects regression models. Journal of Applied Statistics, 46(3):508–521, March 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2019:PSD

[3733] Weiwei Wang, Xianyi Wu, Xiaoqi Zhang, Xiaobing Zhao, and Xian Zhou. Partial sufficient dimension reduction on joint model of recurrent and terminal events. Journal of Applied Statistics, 46(3):522–541, March 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hosseini:2019:ACM

[3734] Fatemeh Hosseini and Omid Karimi. Approximate composite marginal likelihood inference in spatial generalized linear mixed models. Journal of Applied Statistics, 46(3):542–558, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Danisman:2019:CSM

[3735] Ozgur Danisman and Umay Uzunoglu Kocer. Construction of a semi-Markov model for the performance of a football team in the presence of missing data. Journal of Applied Statistics, 46(3):559–576, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2019:Cb

[3736] Anonymous. Corrigendum. Journal of Applied Statistics, 46(3):577–579, March 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

King:2019:BSQ

[3737] Clay King and Joon Jin Song. Bayesian spatial quantile regression for areal count data, with application on substitute care placements in Texas. Journal of Applied Statistics, 46(4):580–597, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gumedze:2019:ULR

[3738] Freedom N. Gumedze. Use of likelihood ratio tests to detect outliers under the variance shift outlier model. Journal of Applied Statistics, 46(4):598–620, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Inan:2019:PSW

[3739] Gul Inan, Mahbub A. H. M. Latif, and John Preisser. A PRESS statistic for working correlation structure selection in generalized estimating equations. Journal of Applied Statistics, 46(4):621–637, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McElroy:2019:CEE

[3740] Tucker McElroy, Osbert Pang, and George Sheldon. Custom epoch estimation for surveys. Journal of Applied Statistics, 46(4):638–663, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2019:DCP

[3741] M. Xu, Y. Wu, and B. Jin. Detection of a change-point in variance by a weighted sum of powers of variances test. Journal of Applied Statistics, 46(4):664–679, April 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2019:NDC

[3742] C. I. Li, J. N. Pan, and M. H. Huang. A new demerit control chart for monitoring the quality of multivariate Poisson processes. Journal of Applied Statistics, 46(4):680–699, April 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mazucheli:2019:OPU

[3743] Josmar Mazucheli, André Felipe Berdusco Menezes, and Subrata Chakraborty. On the one parameter unit-Lindley distribution and its associated regression model for proportion data. Journal of Applied Statistics, 46(4):700–714, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vidovic:2019:MPM

[3744] Zoran Vidović. On MLEs of the parameters of a modified Weibull distribution based on record values. Journal of Applied Statistics, 46(4):715–724, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2019:PKE

[3745] Jen-Hao Chen and Wen-Liang Hung. Parametrizing the Kepler exoplanet period-radius distribution with the bivariate normal inverse Gaussian distribution. Journal of Applied Statistics, 46(4):725–736, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lucas:2019:NMI

[3746] Pablo Lucas and Issam Malki. A note on the modelling and interpretation of a public goods game experiment. Journal of Applied Statistics, 46(4):737–753, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Iddrisu:2019:ASA

[3747] Abdul-Karim Iddrisu and Freedom Gumedze. Application of sensitivity analysis to incomplete longitudinal CD4 count data. Journal of Applied Statistics, 46(4):754–769, April 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ma:2019:PII

[3748] Yanbin Ma and Wenhao Gui. Pivotal inference for the inverse Rayleigh distribution based on general progressively Type-II censored samples. Journal of Applied Statistics, 46(5):771– 797, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marrero:2019:BAS

[3749] Osvaldo Marrero. Bayesian analysis of seasonal variation when the sample size and the amplitude are small. Journal of Applied Statistics, 46(5):798–813, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jana:2019:BGC

[3750] Sayantee Jana, Narayanaswamy Balakrishnan, and Jemila S. Hamid. Bayesian growth curve model useful for high-dimensional longitudinal data. Journal of Applied Statistics, 46(5):814–834, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakar:2019:RME

[3751] S. A. Abu Bakar and S. Nadarajah. Risk measure estimation under two component mixture models with trimmed data. Journal of Applied Statistics, 46(5):835–852, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nian:2019:PEP

[3752] Hui Nian, Chang Yu, Juan Ding, Huiyun Wu, William D. Dupont, Steve Brunwasser, Tebeb Gebretsadik, Tina V. Hartert, and Pingsheng Wu. Performance evaluation of propensity score methods for estimating average treatment effects with multi-level treatments. Journal of Applied Statistics, 46(5):853–873, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakouch:2019:PLD

[3753] Hassan S. Bakouch, Muhammad Nauman Khan, Tassaddaq Hussain, and Christophe Chesneau. A power log-Dagum distribution: estimation and applications. Journal of Applied Statistics, 46(5):874–892, May 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pecanka:2019:MAB

[3754] J. Pecanka, A. W. van der Vaart, and M. A. Jonker. Modeling association between multivariate correlated outcomes and high-dimensional sparse covariates: the adaptive SVS method. Journal of Applied Statistics, 46(5):893–913, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Du:2019:EGP

[3755] Jiang Du, Ruiyuan Cao, Eddy Kwessi, and Zhongzhan Zhang. Estimation for generalized partially functional linear additive regression model. Journal of Applied Statistics, 46(5):914–925, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qiu:2019:RLR

[3756] Yanping Qiu and Baosheng Liang.

Robust logistic regression of family data in the presence of missing genotypes. Journal of Applied Statistics, 46(5):926–945, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wilcox:2019:MRR

[3757] Rand R. Wilcox. Multicolinearity and ridge regression: results on type I errors, power and heteroscedasticity. Journal of Applied Statistics, 46(5):946–957, May 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2019:MTS

[3758] G. N. Singh and S. Suman. A modified two-stage randomized response model for estimating the proportion of stigmatized attribute. Journal of Applied Statistics, 46(6):958–978, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hao:2019:RFS

[3759] Meiling Hao, Yuanyuan Lin, Xianhui Liu, and Wenlu Tang. Robust feature screening for high-dimensional survival data. Journal of Applied Statistics, 46(6):979–994, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bountzis:2019:EEO

[3760] P. Bountzis, E. Papadimitriou, and G. Tsaklidis. Estimating the earthquake occurrence rates in Corinth Gulf (Greece) through Markovian arrival process modeling. Journal of Applied Statistics, 46(6):995–1020, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tapia:2019:SAL

[3761] Alejandra Tapia, Viviana Giampaoli, Maria del Pilar Diaz, and Victor Leiva. Sensitivity analysis of longitudinal count responses: a local influence approach and application to medical data. Journal of Applied Statistics, 46(6):1021–1042, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ozanne:2019:BCM

[3762] Marie V. Ozanne, Grant D. Brown, Jacob J. Oleson, Iraci D. Lima, Jose W. Queiroz, Selma M. B. Jeronimo, Christine A. Petersen, and Mary E. Wilson. Bayesian compartmental model for an infectious disease with dynamic states of infection. Journal of Applied Statistics, 46(6):1043–1065, June 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Malekzadeh:2019:ICM

[3763] Ahad Malekzadeh and Mahmood Kharrati-Kopaei. Inferences on the common mean of several heterogeneous log-normal distributions. Journal of Applied Statistics, 46(6):1066–1083, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lemonte:2019:MRM

[3764] Artur J. Lemonte and Germán Moreno-Arenas. On a multivariate regression model for rates and proportions. Journal of Applied Statistics, 46(6):1084– 1106, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Walden:2019:CBC

[3765] A. T. Walden and L. Zhuang. Constructing brain connectivity group graphs from EEG time series. Journal of Applied Statistics, 46(6):1107–1128, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonamente:2019:PMC

[3766] Massimiliano Bonamente. Probability models of chance fluctuations in spectra of astronomical sources with applications to X-ray absorption lines. Journal of Applied Statistics, 46(6):1129– 1154, June 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Campbell:2019:IER

[3767] Harlan Campbell. Is it even rainier in North Vancouver? A non-parametric rank-based test for semicontinuous longitudinal data. Journal of Applied Statistics, 46(7):1155–1176, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2019:SUR

[3768] Jaeoh Kim and HyungJun Cho. Seemingly unrelated regression tree. Journal of Applied Statistics, 46(7):1177–1195, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahmadi:2019:PAL

[3769] Mohammad Vali Ahmadi and Mahdi Doostparast. Pareto analysis for the lifetime performance index of products on the basis of progressively firstfailure-censored batches under balanced symmetric and asymmetric loss functions. Journal of Applied Statistics, 46(7):1196–1227, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2019:STI

[3770] Yilin Li, Ming Zheng, and Wen Yu. A semiparametric type II Tobit model for time-to-event data with application to a merger and acquisition study. Journal of Applied Statistics, 46(7):1228–1245, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kafle:2019:DEM

[3771] Ram C. Kafle, Keshav P. Pokhrel, Netra Khanal, and Chris P. Tsokos. Differential equation model of carbon dioxide emission using functional linear regression. Journal of Applied Statistics, 46(7):1246–1259, July 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Punzo:2019:NLI

[3772] Antonio Punzo. A new look at the inverse Gaussian distribution with applications to insurance and economic data. Journal of Applied Statistics, 46(7):1260–1287, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elsawah:2019:COF

[3773] A. M. Elsawah and Kai-Tai Fang.

A catalog of optimal foldover plans for constructing u-uniform minimum aberration four-level combined designs. Journal of Applied Statistics, 46(7): 1288–1322, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kozlowski:2019:PRU

[3774] Steven E. Kozlowski and Thaddeus Sim. Predicting recessions using trends in the yield spread. Journal of Applied Statistics, 46(7):1323–1335, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kirschstein:2019:AMV

[3775] T. Kirschstein and Steffen Liebscher.

Assessing the market values of soccer players — a robust analysis of data from German 1. and 2. Bundesliga. Journal of Applied Statistics, 46(7):1336–1349, July 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ospina-Galindez:2019:FRC

[3776] Johann Ospína-Galindez, Ramón Giraldo, and Mercedes Andrade-Bejarano. Functional regression concurrent model with spatially correlated errors: application to rainfall ground validation. Journal of Applied Statistics, 46(8):1350–1363, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fernandez:2019:SCC

[3777] Daniel Fernández, Daphna Harel, Panos Ipeirotis, and Tara McAllister.

Statistical considerations for crowdsourced perceptual ratings of human speech productions. Journal of Applied Statistics, 46(8):1364–1384, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kazemi:2019:SIS

[3778] M. Kazemi, D. Shahsavani, and M. Arashi. A sure independence screening procedure for ultra-high dimensional partially linear additive models. Journal of Applied Statistics, 46 (8):1385–1403, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xampeny:2019:SSE

[3779] Rafel Xampeny, Pere Grima, and Xavier Tort-Martorell. Selecting significant effects in factorial designs: Lenth’s method versus the Box–Meyer approach. Journal of Applied Statistics, 46(8):1404–1416, August 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bowater:2019:BCH

[3780] Russell J. Bowater and Ludmila E. Guzmán-Pantoja. Bayesian, classical and hybrid methods of inference when one parameter value is special. Journal of Applied Statistics, 46(8):1417–1437, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sulewski:2019:SCP

[3781] Piotr Sulewski. Some contributions to practice of 2 × 2 contingency tables. Journal of Applied Statistics, 46 (8):1438–1455, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jung:2019:NRM

[3782] Yoonsuh Jung. Nonlinear regression models for heterogeneous data with massive outliers. Journal of Applied Statistics, 46(8):1456–1477, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Niu:2019:THH

[3783] Xiaoqing Niu, Pengfei Li, and Yuejiao Fu. Testing homogeneity in a heteroscedastic contaminated normal mixture. Journal of Applied Statistics, 46 (8):1478–1491, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Budhiraja:2019:ORA

[3784] Sonal Budhiraja and Biswabrata Pradhan. Optimum reliability acceptance sampling plans under progressive type-I interval censoring with random removal using a cost model. Journal of Applied Statistics, 46(8):1492–1517, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Corduas:2019:SPR

[3785] Marcella Corduas and Giancarlo Ragozini. A statistical procedure for representing state fragility and transition paths. Journal of Applied Statistics, 46(8):1518–1528, August 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beretta:2019:VSP

[3786] Alessandro Beretta and Cédric Heuchenne.

Variable selection in proportional hazards cure model with time-varying covariates, application to US bank failures. Journal of Applied Statistics, 46 (9):1529–1549, September 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez-Camblor:2019:IBD

[3787] P. Martínez-Camblor, S. Pérez-Fernández, and S. Díaz-Coto. Improving the biomarker diagnostic capacity via functional transformations. Journal of Applied Statistics, 46(9): 1550–1566, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chowdhury:2019:GRZ

[3788] Shrabanti Chowdhury, Saptarshi Chatterjee, Himel Mallick, Prithish Banerjee, and Broti Garai. Group regularization for zero-inflated Poisson regression models with an application to insurance ratemaking. Journal of Applied Statistics, 46(9):1567–1581, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tutz:2019:FUM

[3789] Gerhard Tutz and Micha Schneider. Flexible uncertainty in mixture models for ordinal responses. Journal of Applied Statistics, 46(9):1582–1601, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pereira:2019:NME

[3790] Marcos Antonio Alves Pereira and Cibele Maria Russo. Nonlinear mixedeffects models with scale mixture of skew-normal distributions. Journal of Applied Statistics, 46(9):1602–1620, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sukumaran:2019:MAR

[3791] Anjana Sukumaran and Isha Dewan. Modelling and analysis of recallbased competing risks data. Journal of Applied Statistics, 46(9):1621–1635, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

OShea:2019:MTS

[3792] Amy M. J. O’Shea and Jeffrey D. Dawson. Modeling time series data with semi-reflective boundaries. Journal of Applied Statistics, 46(9):1636–1648, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harrar:2019:CRN

[3793] Solomon W. Harrar, Fabrizio Ronchi, and Luigi Salmaso. A comparison of recent nonparametric methods for testing effects in two-by-two factorial designs. Journal of Applied Statistics, 46 (9):1649–1670, September 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chakraborty:2019:TSC

[3794] Hrishikesh Chakraborty, Akhtar Hossain, and Mahbub A. H. M. Latif. A three-state continuous time Markov chain model for HIV disease burden. Journal of Applied Statistics, 46(9): 1671–1688, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haller:2019:IAT

[3795] Bernhard Haller, Hans-Henning Eckstein, Peter A. Ringleb, and Kurt Ulm.

Investigation of age-treatment interaction in the SPACE trial using different statistical approaches. Journal of Applied Statistics, 46(9):1689–1701, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Smith:2019:GRS

[3796] Abigail R. Smith, Nathan P. Goodrich, Charlotte A. Beil, Qian Liu, Robert M. Merion, Brenda W. Gillespie, and Jarcy Zee. Graphical representation of survival curves in the presence of time-dependent categorical covariates with application to liver transplantation. Journal of Applied Statistics, 46 (9):1702–1713, September 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Carbo:2019:ECE

[3797] Jose M. Carbo, Daniel J. Graham, Anupriya, Daniel Casas, and Patricia C. Melo. Evaluating the causal economic impacts of transport investments: evidence from the Madrid– Barcelona high speed rail corridor. Journal of Applied Statistics, 46(9): 1714–1723, September 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2019:SRC

[3798] Yanqin Feng, Shurong Lin, and Yang Li. Semiparametric regression of clustered current status data. Journal of Applied Statistics, 46(10):1724–1737, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2019:RSR

[3799] Shuang Xu and Chun-Xia Zhang.

Robust sparse regression by modeling noise as a mixture of Gaussians. Journal of Applied Statistics, 46(10): 1738–1755, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Votsi:2019:CIM

[3800] I. Votsi and A. Brouste. Confidence interval for the mean time to failure in semi-Markov models: an application to wind energy production. Journal of Applied Statistics, 46(10): 1756–1773, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2019:CMF

[3801] Ruhao Wu and Bo Wang. Coherent mortality forecasting by the weighted multilevel functional principal component approach. Journal of Applied Statistics, 46(10):1774–1791, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prataviera:2019:EPE

[3802] F. Prataviera, J. C. S. Vasconcelos, G. M. Cordeiro, E. M. Hashimoto, and E. M. M. Ortega. The exponentiated power exponential regression model with different regression structures: application in nursing data. Journal of Applied Statistics, 46(10): 1792–1821, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fischer:2019:STD

[3803] Katinka Fischer, Christoph Bothung, Felix Lieder, Stefan Wolfart, and Holger Schwender. A statistical test for detecting discordance in rankings between k groups. Journal of Applied Statistics, 46(10):1822–1842, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dey:2019:BCI

[3804] Sanku Dey and Mahendra Saha. Bootstrap confidence intervals of generalized process capability index C_{pyk} using different methods of estimation. Journal of Applied Statistics, 46(10): 1843–1869, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alam:2019:SSL

[3805] Mohammad Samsul Alam, Syed Shahadat Hossain, and Farha Ferdous Sheela. Spatial smoothing of low birth weight rate in Bangladesh using Bayesian hierarchical model. Journal of Applied Statistics, 46(10):1870–1885, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Safari:2019:IEO

[3806] Muhammad Aslam Mohd Safari, Nurulkamal Masseran, and Kamarulzaman Ibrahim. On the identification of extreme outliers and dragon–kings mechanisms in the upper tail of income distribution. Journal of Applied Statistics, 46(10):1886–1902, October 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McAndrew:2019:HCM

[3807] Thomas McAndrew, Bjorn Redfors, Aaron Crowley, Yiran Zhang, Shmuel Chen, Mordechai Golomb, Maria C.

Alu, Dominic P. Francese, Ori Ben-Yehuda, Akiko Maehara, Gary S. Mintz, Gregg W. Stone, and Paul L. Jenkins. How Cox models react to a study-specific confounder in a patientlevel pooled dataset: random effects better cope with an imbalanced covariate across trials unless baseline hazards differ. Journal of Applied Statistics, 46(10):1903–1916, October 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wei:2019:PMC

[3808] Zheng Wei and Erin M. Conlon. Parallel Markov chain Monte Carlo for Bayesian hierarchical models with big data, in two stages. Journal of Applied Statistics, 46(11):1917–1936, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

M:2019:POM

[3809] Dileep Kumar M., Sankaran P. G., and Unnikrishnan Nair N. Proportional odds model — a quantile approach. Journal of Applied Statistics, 46(11): 1937–1955, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2019:CFP

[3810] Fengqing Zhang and Jiangtao Gou.

Control of false positive rates in clusterwise fMRI inferences. Journal of Applied Statistics, 46(11):1956–1972, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wei:2019:BIA

[3811] Yunfei Wei and Shifeng Xiong.

Bayesian integrative analysis for multifidelity computer experiments. Journal of Applied Statistics, 46(11):1973–1987, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shi:2019:DAC

[3812] Yue Shi, Chi Tim Ng, Zhiguo Feng, and Ka-Fai Cedric Yiu. A descent algorithm for constrained LAD-lasso estimation with applications in portfolio selection. Journal of Applied Statistics, 46(11):1988–2009, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hajrajabi:2019:NSA

[3813] A. Hajrajabi and M. Maleki. Nonlinear semiparametric autoregressive model with finite mixtures of scale mixtures of skew normal innovations. Journal of Applied Statistics, 46(11): 2010–2029, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fong:2019:HMS

[3814] Tom Fong, Ceara Hui, and Alfred Y.- T. Wong. How might sovereign bond yields in Asia Pacific react to US monetary normalisation under turbulent market conditions? Journal of Applied Statistics, 46(11):2030–2055, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2019:CUS

[3815] Yong Wang and Xuxu Wang. Classification using semiparametric mixtures. Journal of Applied Statistics, 46(11): 2056–2074, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Amiri:2019:MGL

[3816] Leila Amiri, Mojtaba Khazaei, and Mojtaba Ganjali. Mixtures of general location model with factor analyzer covariance structure for clustering mixed type data. Journal of Applied Statistics, 46(11):2075–2100, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sabharwal:2019:TDB

[3817] Alka Sabharwal, Gurprit Grover, and Sakshi Kaushik. Testing the difference between bipolar disorder and schizophrenia on the basis of the severity of symptoms with C(α) test. Journal of Applied Statistics, 46(11): 2101–2110, November 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Farahani:2019:LDA

[3818] Zahra Sadat Meshkani Farahani, Esmaile Khorram, Mojtaba Ganjali, and Taban Baghfalaki. Longitudinal data analysis in the presence of informative sampling: weighted distribution or joint modelling. Journal of Applied Statistics, 46(12):2111–2127, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2019:PIQ

[3819] Minjung Lee. Parametric inference for quantile event times with adjustment for covariates on competing risks data. Journal of Applied Statistics, 46(12): 2128–2144, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rahman:2019:MBC

[3820] M. Shafiqur Rahman and Afrin Sadia Rumana. A model-based concordancetype index for evaluating the added predictive ability of novel risk factors and markers in the logistic regression models. Journal of Applied Statistics, 46(12):2145–2163, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chun:2019:GRT

[3821] Young H. Chun. Generalized run tests for statistical process control. Journal of Applied Statistics, 46(12): 2164–2179, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Theodosiadou:2019:SPM

[3822] O. Theodosiadou, V. Polimenis, and G. Tsaklidis. A semi-parametric method for estimating the beta coefficients of the hidden two-sided asset return jumps. Journal of Applied Statistics, 46(12):2180–2197, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Naranjo:2019:BAM

[3823] Lizbeth Naranjo, Carlos J. Pérez, Jacinto Martín, Timothy Mutsvari, and Emmanuel Lesaffre. A Bayesian approach for misclassified ordinal response data. Journal of Applied Statistics, 46(12):2198–2215, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ju:2019:PSP

[3824] Cheng Ju, Mary Combs, Samuel D. Lendle, Jessica M. Franklin, Richard Wyss, Sebastian Schneeweiss, and Mark J. van der Laan. Propensity score prediction for electronic healthcare databases using super learner and high-dimensional propensity score methods. Journal of Applied Statistics, 46(12):2216–2236, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nani:2019:VRE

[3825] Asma Nani, Imed Gamoudi, and Mohamed El Ghourabi. Value-at-risk estimation by LS-SVR and FS-LS-SVR based on GAS model. Journal of Applied Statistics, 46(12):2237–2253, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2019:SFB

[3826] Yonggang Lu and Peter Westfall. Simple and flexible Bayesian inferences for standardized regression coefficients. Journal of Applied Statistics, 46(12): 2254–2288, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Miller:2019:ITA

[3827] Nikolay Miller, Yiming Yang, Bruce Sun, and Guoyi Zhang. Identification of technical analysis patterns with smoothing splines for bitcoin prices. Journal of Applied Statistics, 46(12): 2289–2297, December 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cai:2019:BAL

[3828] Jingheng Cai, Zhibin Liang, Rongqian Sun, Chenyi Liang, and Junhao Pan. Bayesian analysis of latent Markov models with non-ignorable missing data. Journal of Applied Statistics, 46(13):2299–2313, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mozgunov:2019:LFR

[3829] P. Mozgunov, T. Jaki, and M. Gasparini. Loss functions in restricted parameter spaces and their Bayesian applications. Journal of Applied Statistics, 46(13):2314–2337, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Muller:2019:OAS

[3830] Cord A. Müller. Optimal acceptance sampling for modules F and F1 of the European Measuring Instruments Directive. Journal of Applied Statistics, 46(13):2338–2356, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hwang:2019:JML

[3831] Yi-Ting Hwang, Chia-Hui Huang, Chun-Chao Wang, Tzu-Yin Lin, and Yi-Kuan Tseng. Joint modelling of longitudinal binary data and survival data. Journal of Applied Statistics, 46 (13):2357–2371, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

daSilva:2019:BEF

[3832] Marcelo A. da Silva, Anne C. Huggins-Manley, José A. Mazzon, and Jorge L. Bazán. Bayesian estimation of a flexible bifactor generalized partial credit model to survey data. Journal of Applied Statistics, 46(13):2372–2387, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsai:2019:OLT

[3833] Chih-Chun Tsai. Optimal lamination test of ethylene vinyl acetate sheets for solar modules. Journal of Applied Statistics, 46(13):2388–2408, 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baghestani:2019:IFW

[3834] A. R. Baghestani and F. S. Hosseini-Baharanchi. An improper form of Weibull distribution for competing risks analysis with Bayesian approach. Journal of Applied Statistics, 46(13): 2409–2417, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Christou:2019:EVR

[3835] Eliana Christou and Michael Grabchak.

Estimation of value-at-risk using single index quantile regression. Journal of Applied Statistics, 46(13):2418–2433, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Calsavara:2019:ZAD

[3836] Vinicius F. Calsavara, Agatha S. Rodrigues, Ricardo Rocha, Francisco Louzada, Vera Tomazella, Ana C. R. L. A. Souza, Rafaela A. Costa, and Rossana P. V. Francisco. Zero-adjusted defective regression models for modeling lifetime data. Journal of Applied Statistics, 46(13):2434–2459, 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Badade:2019:PFR

[3837] Meena Badade and T. V. Ramanathan.

Probabilistic frontier regression models for binary type output data. Journal of Applied Statistics, 46(13):2460–2480, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kourtzidis:2019:CTD

[3838] Stavros Kourtzidis, Panayiotis Tzeremes, and Nickolaos G. Tzeremes. Conditional time-dependent nonparametric estimators with an application to healthcare production function. Journal of Applied Statistics, 46(13):2481– 2490, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qiu:2019:THR

[3839] Shi-Fang Qiu, Li-Xuan Guo, G. Y. Zou, and Dan Yu. Tests for homogeneity of risk differences in stratified design with correlated bilateral data. Journal of Applied Statistics, 46(14):2491–2513, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2019:SOF

[3840] Zhi Liu and Tie Zhang. A second-order fuzzy time series model for stock price analysis. Journal of Applied Statistics, 46(14):2514–2526, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cobre:2019:MTT

[3841] Juliana Cobre, Fabiana Arca Cruz Tortorelli, and Sandra Cristina de Oliveira. Modelling two types of heterogeneity in the analysis of student success. Journal of Applied Statistics, 46(14):2527–2539, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2019:DHS

[3842] Soeun Kim and Woojoo Lee. Discovering hidden statistical issues through individual-level models in ecological studies. Journal of Applied Statistics, 46(14):2540–2552, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ha:2019:PLA

[3843] Il Do Ha, Jong-Min Kim, and Takeshi Emura. Profile likelihood approaches for semiparametric copula and frailty models for clustered survival data. Journal of Applied Statistics, 46(14): 2553–2571, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ehwerhemuepha:2019:MPU

[3844] Louis Ehwerhemuepha, Heng Sok, and Cyril Rakovski. A more powerful unconditional exact test of homogeneity for 2 × c contingency table analysis. Journal of Applied Statistics, 46(14): 2572–2582, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zaidi:2019:PHC

[3845] F. S. Zaidi, P. Castagliola, K. P. Tran, and M. B. C. Khoo. Performance of the Hotelling T^{2} control chart for compositional data in the presence of measurement errors. Journal of Applied Statistics, 46(14):2583–2602, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2019:MPP

[3846] Chenlong Li, Zhanjie Song, Wenjun Wang, and Xu (Sunny) Wang. Mining periodic patterns and cascading bursts phenomenon in individual e-mail communication. Journal of Applied Statistics, 46(14):2603–2626, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bayisa:2019:MBC

[3847] Fekadu L. Bayisa and Jun Yu. Modelbased computed tomography image estimation: partitioning approach. Journal of Applied Statistics, 46(14):2627– 2648, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hasegawa:2019:TUM

[3848] Hikaru Hasegawa and Kazuhiro Ueda. Time use of married couples: Bayesian approach. Journal of Applied Statistics, 46(14):2649–2665, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2019:DRM

[3849] Minkyung Oh and Jungwon Mun. Diagnostics for repeated measurements in generalized linear mixed effects models. Journal of Applied Statistics, 46(14): 2666–2676, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qiu:2019:NAS

[3850] Mingfei Qiu, Robert Paige, and Vic Patrangenaru. A nonparametric approach to 3D shape analysis from digital camera images — II. Journal of Applied Statistics, 46(15):2677–2699, 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramyamol:2019:ODV

[3851] P. C. Ramyamol and M. Kumar. Optimal design of variable acceptance sampling plans for mixture distribution. Journal of Applied Statistics, 46(15): 2700–2721, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Votsi:2019:CFO

[3852] Irene Votsi. Conditional failure occurrence rates for semi-Markov chains. Journal of Applied Statistics, 46(15): 2722–2743, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramires:2019:NSW

[3853] Thiago G. Ramires, Luiz R. Nakamura, Ana J. Righetto, Rodrigo R. Pescim, Josmar Mazucheli, and Gauss M. Cordeiro. A new semiparametric Weibull cure rate model: fitting different behaviors within GAMLSS. Journal of Applied Statistics, 46(15):2744– 2760, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hidaka:2019:SDA

[3854] Akinori Hidaka, Kenji Watanabe, and Takio Kurita. Sparse discriminant analysis based on estimation of posterior probabilities. Journal of Applied Statistics, 46(15):2761–2785, 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marange:2019:MBE

[3855] C. S. Marange and Y. Qin. A momentbased empirical likelihood ratio test for exponentiality using the probability integral transformation. Journal of Applied Statistics, 46(15):2786–2803, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Viviani:2019:JMG

[3856] S. Viviani. Joint models for a GLMtype longitudinal response and a timeto-event with smooth random effects. Journal of Applied Statistics, 46(15): 2804–2818, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khachatryan:2019:MTB

[3857] Davit Khachatryan and Brigitte Muehlmann. Measuring technological breadth and depth of patent documents using Rao’s quadratic entropy. Journal of Applied Statistics, 46(15):2819–2844, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ubaidillah:2019:MFH

[3858] Azka Ubaidillah, Khairil Anwar Notodiputro, Anang Kurnia, and I. Wayan Mangku. Multivariate Fay– Herriot models for small area estimation with application to household consumption per capita expenditure in Indonesia. Journal of Applied Statistics, 46(15):2845–2861, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2019:SMZ

[3859] Tian Chen, Hui Zhang, and Bo Zhang. A semiparametric marginalized zeroinflated model for analyzing healthcare utilization panel data with missingness. Journal of Applied Statistics, 46(16):2862–2883, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ma:2019:QRM

[3860] Huijuan Ma, Limin Peng, and Haoda Fu. Quantile regression modeling of latent trajectory features with longitudinal data. Journal of Applied Statistics, 46(16):2884–2904, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weiss:2019:SMO

[3861] Christian H. Weiß. On some measures of ordinal variation. Journal of Applied Statistics, 46(16):2905–2926, 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jamal:2019:GIK

[3862] Farrukh Jamal, Muhammad Arslan Nasir, Gamze Ozel, M. Elgarhy, and Naushad Mamode Khan. Generalized inverted Kumaraswamy generated family of distributions: theory and applications. Journal of Applied Statistics, 46(16):2927–2944, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Giussani:2019:MOF

[3863] A. Giussani and M. Bonetti. Marshall– Olkin frailty survival models for bivariate right-censored failure time data. Journal of Applied Statistics, 46(16): 2945–2961, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yildiz:2019:EDF

[3864] Tugba Ozkal Yildiz and Yusuf Can Sevil. Empirical distribution function estimators based on sampling designs in a finite population using single auxiliary variable. Journal of Applied Statistics, 46(16):2962–2974, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhattacharya:2019:DVS

[3865] Ritwik Bhattacharya and Muhammad Aslam. Design of variables sampling plans based on lifetime-performance index in presence of hybrid censoring scheme. Journal of Applied Statistics, 46(16):2975–2986, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reynolds:2019:SCD

[3866] Evan Reynolds, Brian Callaghan, and Mousumi Banerjee. SVM–CART for disease classification. Journal of Applied Statistics, 46(16):2987–3007, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jana:2019:ISS

[3867] Nabakumar Jana, Somesh Kumar, and Kashinath Chatterjee. Inference on stress-strength reliability for exponential distributions with a common scale parameter. Journal of Applied Statistics, 46(16):3008–3031, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Young:2019:DSC

[3868] Timothy M. Young, Olga Khaliukova, Nicolas André, Alexander Petutschnigg, Timothy G. Rials, and Chung-Hao Chen. Detecting specialcause variation ‘events’ from process data signatures. Journal of Applied Statistics, 46(16):3032–3043, 2019. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akhundjanov:2019:PLD

[3869] Sherzod B. Akhundjanov and Lauren Chamberlain. The power-law distribution of agricultural land size. Journal of Applied Statistics, 46(16):3044–3056, 2019. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rivas:2020:IAG

[3870] Luisa Rivas and Manuel Galea. Influence analysis for the generalized Waring regression model. Journal of Applied Statistics, 47(1):1–27, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2020:SRA

[3871] Mengmeng Guo, Jingyong Su, Li Sun, and Guofeng Cao. Statistical regression analysis of functional and shape data. Journal of Applied Statistics, 47(1):28–44, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dai:2020:QRG

[3872] Xiaowen Dai, Zhen Yan, Maozai Tian, and ManLai Tang. Quantile regression for general spatial panel data models with fixed effects. Journal of Applied Statistics, 47(1):45–60, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fan:2020:KNQ

[3873] Caiyun Fan, Gang Ding, and Feipeng Zhang. A kernel nonparametric quantile estimator for right-censored competing risks data. Journal of Applied Statistics, 47(1):61–75, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2020:RPD

[3874] Xiaohui Liu and Yang He. RR-plot:

a descriptive tool for regression observations. Journal of Applied Statistics, 47(1):76–90, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2020:THT

[3875] Sanying Feng, Gaorong Li, Tiejun Tong, and Shuanghua Luo. Testing for heteroskedasticity in two-way fixed effects panel data models. Journal of Applied Statistics, 47(1):91–116, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tian:2020:LBQ

[3876] Yuzhu Tian, Liyong Wang, Manlai Tang, Yanchao Zang, and Maozai Tian. Likelihood-based quantile autoregressive distributed lag models and its applications. Journal of Applied Statistics, 47(1):117–131, January 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hamaz:2020:IMD

[3877] Abdelghani Hamaz, Ouerdia Arezki, and Farida Achemine. Impact of missing data on the prediction of random fields. Journal of Applied Statistics, 47 (1):132–149, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2020:REL

[3878] Joonsung Kang. Robust estimation for longitudinal data based upon minimum Hellinger distance. Journal of Applied Statistics, 47(1):150–159, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Krishnamoorthy:2020:CIP

[3879] K. Krishnamoorthy, Dustin Waguespack, and Ngan Hoang-Nguyen-Thuy. Confidence interval, prediction interval and tolerance limits for a twoparameter Rayleigh distribution. Journal of Applied Statistics, 47(1):160– 175, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goswami:2020:DEA

[3880] Suchismita Goswami and Edward J.

Wegman. Detection of excessive activities in time series of graphs. Journal of Applied Statistics, 47(1):176–200, January 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nguyen:2020:VSU

[3881] Van Cuong Nguyen and Chi Tim Ng. Variable selection under multicollinearity using modified log penalty. Journal of Applied Statistics, 47(2): 201–230, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2020:ESC

[3882] Bing Wang, Xiaoguang Wang, and Lixin Song. Estimation in the single change-point hazard function for interval-censored data with a cure fraction. Journal of Applied Statistics, 47 (2):231–247, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghilagaber:2020:BCP

[3883] Gebrenegus Ghilagaber and Parfait Munezero. Bayesian change-point modelling of the effects of 3-points-for-a-win rule in football. Journal of Applied Statistics, 47(2):248–264, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lemonte:2020:ZIB

[3884] Artur J. Lemonte, Germán Moreno-Arenas, and Fredy Castellares. Zeroinflated Bell regression models for count data. Journal of Applied Statistics, 47(2):265–286, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zandkarimi:2020:RIM

[3885] Eghbal Zandkarimi, Abbas Moghimbeigi, Hossein Mahjub, and Reza Majdzadeh. Robust inference in the multilevel zero-inflated negative binomial model. Journal of Applied Statistics, 47 (2):287–305, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ramos:2020:EPF

[3886] Pedro L. Ramos, Dipak K. Dey, Francisco Louzada, and Victor H. Lachos. An extended Poisson family of life distribution: a unified approach in competitive and complementary risks. Journal of Applied Statistics, 47(2): 306–322, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2020:DFM

[3887] Mengta Yang, Reza Modarres, and Lingzhe Guo. Depth functions and mutidimensional medians on minimal spanning trees. Journal of Applied Statistics, 47(2):323–336, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Burch:2020:ZIB

[3888] Brent Burch and Jesse Egbert. Zeroinflated beta distribution applied to word frequency and lexical dispersion in corpus linguistics. Journal of Applied Statistics, 47(2):337–353, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

El-Morshedy:2020:NTP

[3889] M. El-Morshedy, M. S. Eliwa, and H. Nagy. A new two-parameter exponentiated discrete Lindley distribution: properties, estimation and applications. Journal of Applied Statistics, 47(2):354–375, February 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Modak:2020:UCE

[3890] Soumita Modak, Tanuka Chattopadhyay, and Asis Kumar Chattopadhyay. Unsupervised classification of eclipsing binary light curves through kmedoids clustering. Journal of Applied Statistics, 47(2):376–392, February 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goual:2020:VBX

[3891] Hafida Goual and Haitham M. Yousof.

Validation of Burr XII inverse Rayleigh model via a modified chi-squared goodness-of-fit test. Journal of Applied Statistics, 47(3):393–423, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DallaValle:2020:BAI

[3892] Luciana Dalla Valle, Fabrizio Leisen, Luca Rossini, and Weixuan Zhu. Bayesian analysis of immigration in Europe with generalized logistic regression. Journal of Applied Statistics, 47(3):424–438, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mitra:2020:AIC

[3893] Debanjan Mitra, Ujjwal Das, and Kalyan Das. Analysis of intervalcensored competing risks data under missing causes. Journal of Applied Statistics, 47(3):439–459, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chew:2020:ERR

[3894] XinYing Chew, Khai Wah Khaw, and Wai Chung Yeong. The efficiency of run rules schemes for the multivariate coefficient of variation: a Markov chain approach. Journal of Applied Statistics, 47(3):460–480, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tucker:2020:GAC

[3895] J. Derek Tucker, John R. Lewis, Caleb King, and Sebastian Kurtek. A geometric approach for computing tolerance bounds for elastic functional data. Journal of Applied Statistics, 47(3):481–505, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kumar:2020:SAZ

[3896] C. Satheesh Kumar and Rakhi Ramachandran. On some aspects of a zero-inflated overdispersed model and its applications. Journal of Applied Statistics, 47(3):506–523, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2020:ABS

[3897] Xueying Wang, Chunxiao Zhou, Kepher Makambi, Ao Yuan, and Jaeil Ahn. Analysis of batched service time data using Gaussian and semiparametric kernel models. Journal of Applied Statistics, 47(3):524–540, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2020:CTP

[3898] Navdeep Singh and Parminder Singh.

A closed testing procedure for comparison between successive variances. Journal of Applied Statistics, 47(3):541– 553, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Capecchi:2020:MMA

[3899] Stefania Capecchi and Maurizio Curtarelli. A mixture model to assess perception of discrimination on grounds of sexual orientation for policy considerations. Journal of Applied Statistics, 47(3):554–567, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2020:DSV

[3900] Lili Zhang, Herman Ray, Jennifer Priestley, and Soon Tan. A descriptive study of variable discretization and cost-sensitive logistic regression on imbalanced credit data. Journal of Applied Statistics, 47(3):568–581, March 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nakajima:2020:SSF

[3901] Jouchi Nakajima. Skew selection for factor stochastic volatility models. Journal of Applied Statistics, 47(4):582–601, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2020:SAM

[3902] Zhihuang Yang and Jiahua Chen. Small area mean estimation after effect clustering. Journal of Applied Statistics, 47(4):602–623, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balakrishnan:2020:NOD

[3903] Narayanaswamy Balakrishnan and Chengwei Qin. Nonparametric optimal designs for degradation tests. Journal of Applied Statistics, 47(4):624–641, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2020:NPR

[3904] Mian Huang and Shangbing Yu. A new procedure for resampled portfolio with shrinkaged covariance matrix. Journal of Applied Statistics, 47(4):642–652, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marozzi:2020:IDT

[3905] Marco Marozzi, Amitava Mukherjee, and Jan Kalina. Interpoint distance tests for high-dimensional comparison studies. Journal of Applied Statistics, 47(4):653–665, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soleymani:2020:AIE

[3906] Mehdi Soleymani and Paul S. F. Yip. Assessing the impact of the economic crises in 1997 and 2008 on suicides in Hong Kong, Taiwan and South Korea using a strata-bootstrap algorithm. Journal of Applied Statistics, 47(4):666–684, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Nasser:2020:ASP

[3907] Amjad D. Al-Nasser and Mohammed Obeidat. Acceptance sampling plans from truncated life test based on Tsallis q-exponential distribution. Journal of Applied Statistics, 47(4):685–697, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Behrens:2020:DGE

[3908] Christoph Behrens, Christian Pierdzioch, and Marian Risse. Do German economic research institutes publish efficient growth and inflation forecasts? A Bayesian analysis. Journal of Applied Statistics, 47(4):698–723, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2020:ETH

[3909] Guanfu Liu, Yuejiao Fu, Jianjun Zhang, Xiaolong Pu, and Boying Wang. EM-test for homogeneity in a two-sample problem with a mixture structure. Journal of Applied Statistics, 47(4):724–738, April 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Naeeim:2020:STS

[3910] Nurul Syafiah Abd Naeeim, Nuzlinda Abdul Rahman, and Fatin Afiqah Muhammad Fahimi. A spatialtemporal study of dengue in Peninsular Malaysia for the year 2017 in two different space–time model. Journal of Applied Statistics, 47(4):739–756, April 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodrigo:2020:BMN

[3911] Hansapani Rodrigo and Chris Tsokos. Bayesian modelling of nonlinear Poisson regression with artificial neural networks. Journal of Applied Statistics, 47(5):757–774, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Farhat:2020:NMD

[3912] Naha J. Farhat, Edward L. Boone, and David J. Edwards. A new method for determining the benchmark dose tolerable region and endpoint probabilities for toxicology experiments. Journal of Applied Statistics, 47(5):775–803, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Farzammehr:2020:SNB

[3913] Mohadeseh Alsadat Farzammehr, Mohammad Reza Zadkarami, Geoffrey J. McLachlan, and Sharon X. Lee. Skewnormal Bayesian spatial heterogeneity panel data models. Journal of Applied Statistics, 47(5):804–826, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2020:SEV

[3914] Yang Yang, Jiwei Zhao, Gregory Wilding, Melissa Kluczynski, and Leslie Bisson. Stability enhanced variable selection for a semiparametric model with flexible missingness mechanism and its application to the ChAMP study. Journal of Applied Statistics, 47(5):827–843, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Espinosa:2020:SPT

[3915] Alexandra M. Espinosa and Luís Horna. The statistical properties of the threshold model and the feedback leadership condition. Journal of Applied Statistics, 47(5):844–864, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Piriaei:2020:BEC

[3916] Hassan Piriaei, Gholamhossein Yari, and Rahman Farnoosh. On E-Bayesian estimations for the cumulative hazard rate and mean residual life under generalized inverted exponential distribution and type-II censoring. Journal of Applied Statistics, 47(5):865–889, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ariyo:2020:BMS

[3917] Oludare Ariyo, Adrian Quintero, Johanna Muñoz, Geert Verbeke, and Emmanuel Lesaffre. Bayesian model selection in linear mixed models for longitudinal data. Journal of Applied Statistics, 47(5):890–913, May 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Seo:2020:NCP

[3918] Ju-Hyun Seo, Hyun-Ji Kim, and Jea-Young Lee. Nomogram construction to predict dyslipidemia based on a logistic regression analysis. Journal of Applied Statistics, 47(5):914–926, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Verdoy:2020:STH

[3919] Pablo Juan Verdoy. Spatio-temporal hierarchical Bayesian analysis of wildfires with stochastic partial differential equations. A case study from Valencian Community (Spain). Journal of Applied Statistics, 47(5):927–946, May 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2020:MUL

[3920] Saralees Nadarajah and Stephen Chan. On moments of the unit Lindley distribution. Journal of Applied Statistics, 47(5):947–949, May 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See discussion [3921].

Mazucheli:2020:DCM

[3921] J. Mazucheli, A. F. B. Menezes, and S. Chakraborty. Discussion comments on ‘On moments of the unit Lindley distribution’. Journal of Applied Statistics, 47(5):950–953, May 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [3920].

Mazucheli:2020:UWD

[3922] J. Mazucheli, A. F. B. Menezes, L. B. Fernandes, R. P. de Oliveira, and M. E. Ghitany. The unit-Weibull distribution as an alternative to the Kumaraswamy distribution for the modeling of quantiles conditional on covariates. Journal of Applied Statistics, 47(6):954–974, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dey:2020:CME

[3923] Sanku Dey and Mazen Nassar. Classical methods of estimation on constant stress accelerated life tests under exponentiated Lindley distribution. Journal of Applied Statistics, 47(6):975–996, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Son:2020:SGM

[3924] Sungtaek Son, Cheolwoo Park, and Yongho Jeon. Sparse graphical models via calibrated concave convex procedure with application to fMRI data. Journal of Applied Statistics, 47(6): 997–1016, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2020:CBE

[3925] Xiaoning Kang, Chaoping Xie, and Mingqiu Wang. A Cholesky-based estimation for large-dimensional covariance matrices. Journal of Applied Statistics, 47(6):1017–1030, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mantalos:2020:IVM

[3926] P. Mantalos, A. Karagrigoriou, L. Strelec, P. Jordanova, P. Hermann, J. Kiselák, J. Hudák, and M. Stehlík. On improved volatility modelling by fitting skewness in ARCH models. Journal of Applied Statistics, 47(6): 1031–1063, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2020:TTG

[3927] Jinyu Li, Yutong Lai, Chi Zhang, and Qi Zhang. TGCnA: temporal gene coexpression network analysis using a low-rank plus sparse framework. Journal of Applied Statistics, 47(6):1064–1083, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mondal:2020:BIW

[3928] Shuvashree Mondal and Debasis Kundu. A bivariate inverse Weibull distribution and its application in complementary risks model. Journal of Applied Statistics, 47(6):1084–1108, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gkillas:2020:JBR

[3929] Konstantinos Gkillas, Rangan Gupta, Chi Keung Marco Lau, and Muhammad Tahir Suleman. Jumps beyond the realms of cricket: India’s performance in One Day Internationals and stock market movements. Journal of Applied Statistics, 47(6):1109–1127, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hassani:2020:FIR

[3930] Hossein Hassani, Mohammad Reza Yeganegi, Juncal Cuñado, and Rangan Gupta. Forecasting interest rate volatility of the United Kingdom: evidence from over 150 years of data. Journal of Applied Statistics, 47(6): 1128–1143, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2020:Ca

[3931] Anonymous. Correction. Journal of Applied Statistics, 47(6):i–ii, June 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Templ:2020:ERO

[3932] M. Templ, J. Gussenbauer, and P. Filzmoser. Evaluation of robust outlier detection methods for zero-inflated complex data. Journal of Applied Statistics, 47(7):1144–1167, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2020:TPB

[3933] Sha Li, Daniel A. Griffith, and Hong Shu. Temperature prediction based on a space-time regression-kriging model. Journal of Applied Statistics, 47(7): 1168–1190, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rolland:2020:SIM

[3934] Antoine Rolland and Jairo Cugliari.

Sensitivity index to measure dependence on parameters for rankings and top-k rankings. Journal of Applied Statistics, 47(7):1191–1207, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhati:2020:JEL

[3935] Deepesh Bhati and Sudheesh K. Kattumannil. Jackknife empirical likelihood test for testing one-sided Lévy distribution. Journal of Applied Statistics, 47(7):1208–1219, July 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2020:PAN

[3936] Qiaohui Lin, Brenda Betancourt, Benjamin A. Goldstein, and Rebecca C. Steorts. Prediction of appointment noshows using electronic health records. Journal of Applied Statistics, 47(7): 1220–1234, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Abushilah:2020:GBD

[3937] Samira F. Abushilah, Charles C. Taylor, and Arief Gusnanto. Geometrybased distance for clustering amino acids. Journal of Applied Statistics, 47(7):1235–1250, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shinoda:2020:MAO

[3938] Satoru Shinoda, Kouji Yamamoto, Kouji Tahata, and Sadao Tomizawa. A measure of asymmetry for ordinal square contingency tables with an application to modified LANZA score data. Journal of Applied Statistics, 47(7):1251–1260, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fernandez:2020:ASO

[3939] Daniel Fernández, Ivy Liu, Roy Costilla, and Peter Yongqi Gu. Assigning scores for ordered categorical responses. Journal of Applied Statistics, 47(7):1261–1281, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yamashita:2020:PPA

[3940] Haruka Yamashita and Yoshinobu Kawahara. Principal points analysis via p-median problem for binary data. Journal of Applied Statistics, 47(7):1282–1297, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dobra:2020:IMV

[3941] Adrian Dobra, Katherine Buhikire, and Joachim G. Voss. Identifying mediating variables with graphical models: an application to the study of causal pathways in people living with HIV. Journal of Applied Statistics, 47(7):1298–1314, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Betoko:2020:AME

[3942] Aisha Betoko, Chike Nwabuo, Bharath Ambale Venkatesh, Erin P. Ricketts, Sejong Bae, Colin Wu, Samuel S. Gidding, Kiang Liu, João A. C. Lima, and Christopher Cox. Application of measurement error models to correct for systematic differences among readers and vendors in echocardiography measurements: the CARDIA study. Journal of Applied Statistics, 47(7):1315–1324, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2020:Cb

[3943] Anonymous. Correction. Journal of Applied Statistics, 47(7):i, July 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Merola:2020:SFR

[3944] Giovanni Maria Merola. SIMPCA: a framework for rotating and sparsifying principal components. Journal of Applied Statistics, 47(8):1325–1353, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hibbah:2020:MCA

[3945] El Houcine Hibbah, Hamid El Maroufy, Christiane Fuchs, and Taib Ziad. An MCMC computational approach for a continuous time state-dependent regime switching diffusion process. Journal of Applied Statistics, 47(8): 1354–1374, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qiu:2020:CDP

[3946] Shi-Fang Qiu, Jie He, Ji-Ran Tao, Man-Lai Tang, and Wai-Yin Poon.

Comparison of disease prevalence in two populations under double-sampling scheme with two fallible classifiers. Journal of Applied Statistics, 47(8): 1375–1401, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lodhi:2020:IPT

[3947] Chandrakant Lodhi and Yogesh Mani Tripathi. Inference on a progressive type I interval-censored truncated normal distribution. Journal of Applied Statistics, 47(8):1402–1422, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mustafa:2020:IBS

[3948] Abdelfattah Mustafa. Improving the bridge structure by using linear failure rate distribution. Journal of Applied Statistics, 47(8):1423–1438, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Collins:2020:PCS

[3949] Gavin Q. Collins, Matthew J. Heaton, and Leiqiu Hu. Physically constrained spatiotemporal modeling: generating clear-sky constructions of land surface temperature from sparse, remotely sensed satellite data. Journal of Applied Statistics, 47(8):1439–1459, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alam:2020:CAC

[3950] Mohammad Samsul Alam and Sangita Paul. A comparative analysis of clustering algorithms to identify the homogeneous rainfall gauge stations of Bangladesh. Journal of Applied Statistics, 47(8):1460–1481, August 2020.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Naveed:2020:DCC

[3951] Muhammad Naveed, Muhammad Azam, Nasrullah Khan, and Muhammad Aslam. Designing a control chart of extended EWMA statistic based on multiple dependent state sampling. Journal of Applied Statistics, 47(8): 1482–1492, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2020:RAG

[3952] Mei Li and Houbao Xu. Reliability analysis for gap null gate based on model comparison criterion. Journal of Applied Statistics, 47(8):1493–1509, August 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Thackham:2020:MLE

[3953] Mark Thackham and Jun Ma. On maximum likelihood estimation of the semi-parametric Cox model with timevarying covariates. Journal of Applied Statistics, 47(9):1511–1528, September 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez-Camblor:2020:RVM

[3954] P. Martínez-Camblor, S. Pérez-Fernández, and S. Díaz-Coto. The role of the p-value in the multitesting problem. Journal of Applied Statistics, 47 (9):1529–1542, September 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2020:SIB

[3955] Sukhdev Singh, Sanku Dey, and Devendra Kumar. Statistical inference based on generalized Lindley record values. Journal of Applied Statistics, 47(9):1543–1561, September 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guedes:2020:MLR

[3956] Ana C. Guedes, Francisco Cribari-Neto, and Patrícia L. Espinheira. Modified likelihood ratio tests for unit gamma regressions. Journal of Applied Statistics, 47(9):1562–1586, September 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2020:CDA

[3957] Bo Chang and Harry Joe. Copula diagnostics for asymmetries and conditional dependence. Journal of Applied Statistics, 47(9):1587–1615, September 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marcy:2020:MMS

[3958] Peter W. Marcy, Scott A. Vander Wiel, Curtis B. Storlie, Veronica Livescu, and Curt A. Bronkhorst. Modeling material stress using integrated Gaussian Markov random fields. Journal of Applied Statistics, 47(9):1616–1636, September 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patra:2020:IEF

[3959] Lakshmi Kanta Patra. On the improved estimation of a function of the scale parameter of an exponential distribution based on doubly censored sample. Journal of Applied Statistics, 47(9):1637–1651, September 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Haq:2020:MPD

[3960] Abdul Haq and Michael B. C. Khoo.

Multivariate process dispersion monitoring without subgrouping. Journal of Applied Statistics, 47(9):1652–1675, September 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2020:NPA

[3961] Tsai-Yu Lin, Chen-Tuo Liao, and Yen-Ya Liao. Nonconformance probability of an air quality index. Journal of Applied Statistics, 47(9):1676–1689, September 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ferreira:2020:IDH

[3962] Clécio da Silva Ferreira, Víctor H.

Lachos, and Aldo M. Garay. Inference and diagnostics for heteroscedastic nonlinear regression models under skew scale mixtures of normal distributions. Journal of Applied Statistics, 47 (9):1690–1719, September 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vidotto:2020:MIL

[3963] Davide Vidotto, Jeroen K. Vermunt, and Katrijn Van Deun. Multiple imputation of longitudinal categorical data through Bayesian mixture latent Markov models. Journal of Applied Statistics, 47(10):1720–1738, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2020:CLI

[3964] Seongoh Park, Johan Lim, Hyejeong Choi, and Minjung Kwak. Clustering of longitudinal interval-valued data via mixture distribution under covariance separability. Journal of Applied Statistics, 47(10):1739–1756, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hu:2020:ALP

[3965] Xuehua Hu and Wenhao Gui. Assessing the lifetime performance index with Lomax distribution based on progressive type I interval censored sample. Journal of Applied Statistics, 47(10): 1757–1775, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Moral:2020:LSM

[3966] R. A. Moral, J. Hinde, E. M. M. Ortega, C. G. B. Demétrio, and W. A. C. Godoy. Location-scale mixed models and goodness-of-fit assessment applied to insect ecology. Journal of Applied Statistics, 47(10):1776–1793, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nguelifack:2020:RSR

[3967] Brice M. Nguelifack and Isabelle Kemajou-Brown. Robust signed-rank estimation and variable selection for semi-parametric additive partial linear models. Journal of Applied Statistics, 47(10):1794–1819, October 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2020:BTC

[3968] Min Wang, Fang Chen, Tao Lu, and Jianping Dong. Bayesian t-tests for correlations and partial correlations. Journal of Applied Statistics, 47(10): 1820–1832, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pereira:2020:CRO

[3969] Gustavo H. A. Pereira, Juliana Scudilio, Manoel Santos-Neto, Denise A. Botter, and Mônica C. Sandoval. A class of residuals for outlier identification in zero adjusted regression models. Journal of Applied Statistics, 47 (10):1833–1847, October 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alkhairy:2020:QCP

[3970] Ibrahim Alkhairy, Samantha Low-Choy, Justine Murray, Junhu Wang, and Anthony Pettitt. Quantifying conditional probability tables in Bayesian networks: Bayesian regression for scenario-based encoding of elicited expert assessments on feral pig habitat. Journal of Applied Statistics, 47(10): 1848–1884, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Noughabi:2020:TFG

[3971] Hadi Alizadeh Noughabi and Jalil Jarrahiferiz. Tests of fit for the Gumbel distribution: EDF-based tests against entropy-based tests. Journal of Applied Statistics, 47(10):1885–1900, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saffari:2020:BNB

[3972] Seyed Ehsan Saffari and John Carson Allen. Bivariate negative binomial regression model with excess zeros and right censoring: an application to Indonesian data. Journal of Applied Statistics, 47(10):1901–1914, October 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schulz:2020:STM

[3973] Jörn Schulz, Jan Terje Kvaløy, Kjersti Engan, Trygve Eftestøl, Samwel Jatosh, Hussein Kidanto, and Hege Ersdal. State transition modeling of complex monitored health data. Journal of Applied Statistics, 47(11): 1915–1935, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2020:SAT

[3974] Yuantao Zhang and Yue Kuen Kwok. Saddlepoint approximations to tail expectations under non-Gaussian base distributions: option pricing applications. Journal of Applied Statistics, 47 (11):1936–1956, November 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kwon:2020:DPC

[3975] Junhyeon Kwon, Hee-Seok Oh, and Yaeji Lim. Dynamic principal component analysis with missing values. Journal of Applied Statistics, 47(11): 1957–1969, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Schluter:2020:TPA

[3976] Stephan Schlüter and Milena Kresoja. Two preprocessing algorithms for climate time series. Journal of Applied Statistics, 47(11):1970–1989, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Garcia:2020:RJP

[3977] Catalina B. García, Román Salmerón, Claudia García, and José García. Residualization: justification, properties and application. Journal of Applied Statistics, 47(11):1990–2010, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Graffelman:2020:GFF

[3978] Jan Graffelman. Goodness-of-fit filtering in classical metric multidimensional scaling with large datasets. Journal of Applied Statistics, 47(11):2011–2024, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yavuz:2020:VSE

[3979] Fulya Gokalp Yavuz and Olcay Arslan. Variable selection in elliptical linear mixed model. Journal of Applied Statistics, 47(11):2025–2043, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonamente:2020:DSA

[3980] Massimiliano Bonamente. Distribution of the C statistic with applications to the sample mean of Poisson data. Journal of Applied Statistics, 47(11): 2044–2065, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Klein:2020:CGA

[3981] Nadja Klein, Andrew Entwistle, Albert Rosenberger, Thomas Kneib, and Heike Bickeböller. Candidate-gene association analysis for a continuous phenotype with a spike at zero using parent-offspring trios. Journal of Applied Statistics, 47(11):2066–2080, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cappelli:2020:NWH

[3982] Christopher J. Cappelli, Audrey J. Leroux, and Congying Sun. A new way for handling mobility in longitudinal data. Journal of Applied Statistics, 47(11): 2081–2096, November 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Korkmaz:2020:NHT

[3983] Mustafa Ç. Korkmaz. A new heavytailed distribution defined on the bounded interval: the logit slash distribution and its application. Journal of Applied Statistics, 47(12):2097–2119, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Metulini:2020:MSP

[3984] Rodolfo Metulini and Mael Le Carre. Measuring sport performances under pressure by classification trees with application to basketball shooting. Journal of Applied Statistics, 47(12):2120– 2135, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ganji:2020:FPC

[3985] Zainab Abbasi Ganji and Bahram Sadeghpour Gildeh. Fuzzy process capability indices for simple linear profile. Journal of Applied Statistics, 47(12): 2136–2158, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hashimoto:2020:MLR

[3986] E. M. Hashimoto, E. M. M. Ortega, G. M. Cordeiro, A. K. Suzuki, and M. W. Kattan. The multinomial logistic regression model for predicting the discharge status after liver transplantation: estimation and diagnostics analysis. Journal of Applied Statistics, 47 (12):2159–2177, December 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fan:2020:NTS

[3987] Xiaodong Fan, Shi shun Zhao, Qingchun Zhang, and Jianguo Sun. Nonparametric tests for stratified additive hazards model based on current status data. Journal of Applied Statistics, 47(12):2178–2191, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akca:2020:MIM

[3988] Elif Akça and Ceylan Yozgatligil. Mutual information model selection algorithm for time series. Journal of Applied Statistics, 47(12):2192–2207, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gomez-Deniz:2020:QBZ

[3989] E. Gómez-Déniz, D. I. Gallardo, and H. W. Gómez. Quasi-binomial zeroinflated regression model suitable for variables with bounded support. Journal of Applied Statistics, 47(12):2208– 2229, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Triantafyllou:2020:DFM

[3990] Ioannis S. Triantafyllou and Nikolaos I.

Panayiotou. Distribution-free monitoring schemes based on order statistics: a general approach. Journal of Applied Statistics, 47(12):2230–2257, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qasim:2020:NPL

[3991] Muhammad Qasim, B. M. G. Kibria, Kristofer Månsson, and Pär Sjölander. A new Poisson Liu Regression Estimator: method and application. Journal of Applied Statistics, 47(12):2258–2271, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tapak:2020:REE

[3992] Leili Tapak, Omid Hamidi, Payam Amini, and Geert Verbeke. Random effect exponentiated-exponential geometric model for clustered/longitudinal zero-inflated count data. Journal of Applied Statistics, 47(12):2272–2288, December 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stehlik:2020:ESI

[3993] M. Stehlík, L. M. Grilo, and P. K. Jordanova. Editorial to special issue V WCDANM 2018. Journal of Applied Statistics, 47(13–15):2289–2298, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arnold:2020:BPE

[3994] Barry C. Arnold and Matthew A. Arvanitis. On bivariate pseudoexponential distributions. Journal of Applied Statistics, 47(13–15):2299– 2311, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peterson:2020:OQN

[3995] Ryan A. Peterson and Joseph E. Cavanaugh. Ordered quantile normalization: a semiparametric transformation built for the cross-validation era. Journal of Applied Statistics, 47(13–15):2312–2327, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tomarchio:2020:DUC

[3996] Salvatore D. Tomarchio and Antonio Punzo. Dichotomous unimodal compound models: application to the distribution of insurance losses. Journal of Applied Statistics, 47(13–15):2328– 2353, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barbu:2020:SIG

[3997] Vlad Stefan Barbu, Alex Karagrigoriou, and Andreas Makrides. Statistical inference for a general class of distributions with time-varying parameters. Journal of Applied Statistics, 47(13–15):2354–2373, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Antunes:2020:EAM

[3998] Patrícia Antunes, Sandra S. Ferreira, Dário Ferreira, Célia Nunes, and João Tiago Mexia. Estimation in additive models and ANOVA-like applications. Journal of Applied Statistics, 47(13–15):2374–2383, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Carter:2020:BSM

[3999] Knute D. Carter and Joseph E. Cavanaugh. Best-subset model selection based on multitudinal assessments of likelihood improvements. Journal of Applied Statistics, 47(13–15):2384– 2420, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Santos:2020:MCO

[4000] C. Santos, C. Nunes, C. Dias, and J. T. Mexia. Models with commutative orthogonal block structure: a general condition for commutativity. Journal of Applied Statistics, 47(13–15):2421– 2430, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Flournoy:2020:BIA

[4001] Nancy Flournoy and Assaf P. Oron. Bias induced by adaptive dose-finding designs. Journal of Applied Statistics, 47(13–15):2431–2442, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nooghabi:2020:PCI

[4002] M. Jabbari Nooghabi. Process capability indices in normal distribution with the presence of outliers. Journal of Applied Statistics, 47(13–15):2443–2478, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Toure:2020:ANT

[4003] Aboubacar Y. Touré, Simplice Dossou-Gbété, and Célestin C. Kokonendji. Asymptotic normality of the test statistics for the unified relative dispersion and relative variation indexes. Journal of Applied Statistics, 47(13–15):2479– 2491, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahmed:2020:IPT

[4004] Essam A. Ahmed, Ziyad Ali Alhussain, Mukhtar M. Salah, Hanan Haj Ahmed, and M. S. Eliwa. Inference of progressively type-II censored competing risks data from Chen distribution with an application. Journal of Applied Statistics, 47(13–15):2492–2524, 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fermin:2020:VSA

[4005] Lisandro J. Fermín and Jacques Lévy-Véhel. Variability and singularity arising from a piecewise-deterministic Markov process applied to model poor patient compliance in the multi-IV case. Journal of Applied Statistics, 47(13–15):2525–2545, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silva:2020:IBI

[4006] Isabel Silva, Maria Eduarda Silva, and Cristina Torres. Inference for bivariate integer-valued moving average models based on binomial thinning operation. Journal of Applied Statistics, 47(13–15):2546–2564, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cullan:2020:CEP

[4007] Michael Cullan, Scott Lidgard, and Beckett Sterner. Controlling the error probabilities of model selection information criteria using bootstrapping. Journal of Applied Statistics, 47(13–15):2565–2581, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Riedle:2020:RVL

[4008] Benjamin Riedle, Andrew A. Neath, and Joseph E. Cavanaugh. Reconceptualizing the p-value from a likelihood ratio test: a probabilistic pairwise comparison of models based on Kullback– Leibler discrepancy measures. Journal of Applied Statistics, 47(13–15):2582– 2609, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nikitin:2020:GFT

[4009] Ya. Yu. Nikitin and I. A. Ragozin. Goodness-of-fit tests for the logistic location family. Journal of Applied Statistics, 47(13–15):2610–2622, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2020:LAD

[4010] M. Li, Q. Guo, W. J. Zhai, and B. Z. Chen. The linearized alternating direction method of multipliers for lowrank and fused LASSO matrix regression model. Journal of Applied Statistics, 47(13–15):2623–2640, 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nunes:2020:CSS

[4011] Célia Nunes, Elsa Moreira, Sandra S. Ferreira, Dário Ferreira, and João T. Mexia. Considering the sample sizes as truncated Poisson random variables in mixed effects models. Journal of Applied Statistics, 47(13–15):2641–2657, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Giacomazzo:2020:BES

[4012] Mario Giacomazzo and Yiannis Kamarianakis. Bayesian estimation of subset threshold autoregressions: short-term forecasting of traffic occupancy. Journal of Applied Statistics, 47(13–15):2658–2689, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arrue:2020:NTB

[4013] Jaime Arrué, Reinaldo B. Arellano-Valle, Héctor W. Gómez, and Víctor Leiva. On a new type of Birnbaum– Saunders models and its inference and application to fatigue data. Journal of Applied Statistics, 47(13–15):2690– 2710, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Romero:2020:AUV

[4014] Rafael Romero, Jose M. Pavía, Jorge Martín, and Gerardo Romero. Assessing uncertainty of voter transitions estimated from aggregated data. Application to the 2017 French presidential election. Journal of Applied Statistics, 47(13–15):2711–2736, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oliveira:2020:BPB

[4015] Sandra Oliveira, Célia Nunes, Elsa Moreira, Miguel Fonseca, and João T. Mexia. Balanced prime basis factorial fixed effects model with random number of observations. Journal of Applied Statistics, 47(13–15):2737–2748, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Witkovsky:2020:CED

[4016] Viktor Witkovský. Computing the exact distribution of the Bartlett’s test statistic by numerical inversion of its characteristic function. Journal of Applied Statistics, 47(13–15):2749–2764, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Marques:2020:TSD

[4017] F. J. Marques and C. A. Coelho. Testing simultaneously different covariance block diagonal structures — the multisample case. Journal of Applied Statistics, 47(13–15):2765–2784, 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sharifipanah:2020:NCW

[4018] Najme Sharifipanah, Rahim Chinipardaz, and Gholam Ali Parham. A new class of weighted bimodal distribution with application to gamma-ray burst duration data. Journal of Applied Statistics, 47(13–15):2785–2807, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Economou:2020:RIU

[4019] P. Economou, G. Tzavelas, and A. Batsidis. Robust inference under r-size-biased sampling without replacement from finite population. Journal of Applied Statistics, 47(13–15):2808–2824, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Penalva:2020:LMO

[4020] Helena Penalva, M. Ivette Gomes, Frederico Caeiro, and M. Manuela Neves. Lehmer’s mean-of-order-p extreme value index estimation: a simulation study and applications. Journal of Applied Statistics, 47(13–15):2825– 2845, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gomes:2020:EIB

[4021] D. Prata Gomes and M. Manuela Neves. Extremal index blocks estimator: the threshold and the block size choice. Journal of Applied Statistics, 47(13–15):2846–2861, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tchorbadjieff:2020:MIC

[4022] A. Tchorbadjieff and P. Mayster. Models induced from critical birth-death process with random initial conditions. Journal of Applied Statistics, 47(13–15):2862–2878, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

CostaeSilva:2020:LRM

[4023] Eliana Costa e Silva, Isabel Cristina Lopes, Aldina Correia, and Susana Faria. A logistic regression model for consumer default risk. Journal of Applied Statistics, 47(13–15):2879–2894, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Makinde:2020:RDC

[4024] Olusola Samuel Makinde. On rank distribution classifiers for highdimensional data. Journal of Applied Statistics, 47(13–15):2895–2911, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oliveira:2020:MEM

[4025] A. N. Oliveira, R. Menezes, S. Faria, and P. Afonso. Mixed-effects modelling for crossed and nested data: an analysis of dengue fever in the state of Goiás, Brazil. Journal of Applied Statistics, 47(13–15):2912–2926, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silva:2020:NDF

[4026] Isabel Silva and Hugo Alonso. New developments in the forecasting of monthly overnight stays in the North Region of Portugal. Journal of Applied Statistics, 47(13–15):2927–2940, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yin:2020:VSF

[4027] Junhui Yin, Liucang Wu, and Lin Dai. Variable selection in finite mixture of regression models using the skewnormal distribution. Journal of Applied Statistics, 47(16):2941–2960, 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kosztyan:2020:ACS

[4028] Zs. T. Kosztyán, É. Orbán-Mihálykó, Cs. Mihálykó, V. V. Csányi, and A. Telcs. Analyzing and clustering students’ application preferences in higher education. Journal of Applied Statistics, 47(16):2961–2983, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

vanDoorn:2020:BRB

[4029] J. van Doorn, A. Ly, M. Marsman, and E.-J. Wagenmakers. Bayesian rankbased hypothesis testing for the rank sum test, the signed rank test, and Spearman’s ρ. Journal of Applied Statistics, 47(16):2984–3006, 2020. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hashemi:2020:SFA

[4030] Farzane Hashemi, Mehrdad Naderi, Ahad Jamalizadeh, and Tsung-I Lin. A skew factor analysis model based on the normal mean-variance mixture of Birnbaum–Saunders distribution. Journal of Applied Statistics, 47(16):3007–3029, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maghami:2020:BRE

[4031] Mohammad Mahdi Maghami, Mohammad Bahrami, and Farkhondeh Alsadat Sajadi. On bias reduction estimators of skew-normal and skew-t distributions. Journal of Applied Statistics, 47(16):3030–3052, 2020. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

PhD:2021:E

[4032] Jie Chen PhD. Editorial. Journal of Applied Statistics, 48(1):1–3, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nawaz:2021:EDR

[4033] Muhammad Shujaat Nawaz, Muhammad Azam, and Muhammad Aslam. EWMA and DEWMA repetitive control charts under non-normal processes. Journal of Applied Statistics, 48(1):4–40, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2021:CTM

[4034] Xu Zhang, Sean Barnes, Bruce Golden, and Paul Smith. A continuous-time Markov model for estimating readmission risk for hospital inpatients. Journal of Applied Statistics, 48(1):41–60, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Koshti:2021:PEC

[4035] Rohan D. Koshti and Kirtee K. Kamalja. Parameter estimation of Cambanis-type bivariate uniform distribution with Ranked Set Sampling. Journal of Applied Statistics, 48(1):61– 83, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gagnon:2021:ARB

[4036] Philippe Gagnon, Mylène Bédard, and Alain Desgagné. An automatic robust Bayesian approach to principal component regression. Journal of Applied Statistics, 48(1):84–104, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dong:2021:SFM

[4037] Zhihang Dong, Yen-Chi Chen, and Adrian Dobra. A statistical framework for measuring the temporal stability of human mobility patterns. Journal of Applied Statistics, 48(1):105–123, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chesneau:2021:CGD

[4038] Christophe Chesneau, Hassan S. Bakouch, Tassaddaq Hussain, and Bilal A. Para. The cosine geometric distribution with count data modeling. Journal of Applied Statistics, 48(1):124–137, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aslam:2021:MCC

[4039] Muhammad Aslam, Muhammad Ali Raza, Rehan Ahmad Khan Sherwani, Muhammad Farooq, Jun Yong Jeong, and Chi-Hyuck Jun. A mixed control chart for monitoring failure times under accelerated hybrid censoring. Journal of Applied Statistics, 48(1):138– 153, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2021:BIP

[4040] Chen Feng, Paul Griffin, Shravan Kethireddy, and Yajun Mei. A boosting inspired personalized threshold method for sepsis screening. Journal of Applied Statistics, 48(1):154–175, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

vandenBroek:2021:MRP

[4041] Jan van den Broek. Modelling the reproductive power function. Journal of Applied Statistics, 48(1):176–190, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hassani:2021:TCR

[4042] Hossein Hassani, Mansi Ghodsi, Xu Huang, and Emmanuel Sirimal Silva. Is there a causal relationship between oil prices and tourist arrivals? Journal of Applied Statistics, 48(1):191–202, January 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Borges:2021:ETP

[4043] Patrick Borges. Estimating the turning point of the log-logistic hazard function in the presence of long-term survivors with an application for uterine cervical cancer data. Journal of Applied Statistics, 48(2):203–213, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deSousa:2021:RPC

[4044] J. de Sousa, K. Hron, K. Facevicová, and P. Filzmoser. Robust principal component analysis for compositional tables. Journal of Applied Statistics, 48 (2):214–233, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jiang:2021:ODR

[4045] Yunlu Jiang, Yan Wang, Jiantao Zhang, Baojian Xie, Jibiao Liao, and Wenhui Liao. Outlier detection and robust variable selection via the penalized weighted LAD-LASSO method. Journal of Applied Statistics, 48(2): 234–246, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ali:2021:MTM

[4046] Sajid Ali. Monitoring time and magnitude based on the renewal reward process with a random failure threshold. Journal of Applied Statistics, 48 (2):247–284, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bahari:2021:RSP

[4047] Fayyaz Bahari, Safar Parsi, and Mojtaba Ganjali. Reliability of a soccer player based on the bivariate Rayleigh distribution with right censored and ignorable missing data. Journal of Applied Statistics, 48(2):285–300, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gabrio:2021:BHM

[4048] Andrea Gabrio. Bayesian hierarchical models for the prediction of volleyball results. Journal of Applied Statistics, 48(2):301–321, February 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DAlberto:2021:HMR

[4049] Riccardo D’Alberto and Meri Raggi.

How much reliable are the integrated ‘live’ data? A validation strategy proposal for the non-parametric micro statistical matching. Journal of Applied Statistics, 48(2):322–348, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vasconcelos:2021:NRM

[4050] Julio Cezar Souza Vasconcelos, Gauss Moutinho Cordeiro, Edwin Moises Marcos Ortega, and Édila Maria de Rezende. A new regression model for bimodal data and applications in agriculture. Journal of Applied Statistics, 48(2): 349–372, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martins:2021:ITA

[4051] Rui Martins, Jorge Caldeira, Inês Lopes, and José João Mendes. Improving teeth aesthetics using a spatially shared-parameters model for independent regular lattices. Journal of Applied Statistics, 48(2):373–392, February 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cha:2021:VAR

[4052] Ji Hwan Cha and F. G. Badía. Variables acceptance reliability sampling plan for items subject to inverse Gaussian degradation process. Journal of Applied Statistics, 48(3):393–409, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2021:IEN

[4053] Xia Wang, Joseph Pancras, and Dipak K. Dey. Investigating emergent nested geographic structure in consumer purchases: a Bayesian dynamic multi-scale spatiotemporal modeling approach. Journal of Applied Statistics, 48(3):410–433, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2021:DFE

[4054] Shu Wu, Philippe Castagliola, and Giovanni Celano. A distribution-free EWMA control chart for monitoring time-between-events-and-amplitude data. Journal of Applied Statistics, 48(3): 434–454, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2021:PRC

[4055] Cheng Zhang and Tapan K. Nayak. Post-randomization for controlling identification risk in releasing microdata from general surveys. Journal of Applied Statistics, 48(3):455–470, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodriguez:2021:NPS

[4056] Alejandro Rodriguez, Gabriel Pino, and Rodrigo Herrera. A nonparametric statistic for testing conditional heteroscedasticity for unobserved component models. Journal of Applied Statistics, 48(3):471–497, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nikiforova:2021:LHD

[4057] Nedka Dechkova Nikiforova, Rossella Berni, Gabriele Arcidiacono, Luciano Cantone, and Pierpaolo Placidoli. Latin hypercube designs based on strong orthogonal arrays and kriging modelling to improve the payload distribution of trains. Journal of Applied Statistics, 48(3):498–516, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Javed:2021:HOM

[4058] Farrukh Javed, Stepan Mazur, and Edward Ngailo. Higher order moments of the estimated tangency portfolio weights. Journal of Applied Statistics, 48(3):517–535, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Basak:2021:ODP

[4059] Sancharee Basak, Ayanendranath Basu, and M. C. Jones. On the ‘optimal’ density power divergence tuning parameter. Journal of Applied Statistics, 48(3):536–556, March 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vasconcelos:2021:GSP

[4060] Julio Cezar Souza Vasconcelos and Cristian Villegas. Generalized symmetrical partial linear model. Journal of Applied Statistics, 48(3):557–572, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ogwang:2021:ORB

[4061] Tomson Ogwang and Danny I. Cho.

Olympic rankings based on objective weighting schemes. Journal of Applied Statistics, 48(3):573–582, March 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shang:2021:BBE

[4062] Han Lin Shang. Bayesian bandwidth estimation and semi-metric selection for a functional partial linear model with unknown error density. Journal of Applied Statistics, 48(4):583–604, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kaymaz:2021:PTP

[4063] Özlem Kaymaz, Khaled Alqahtani, Henry M. Wood, and Arief Gusnanto. Prediction of tumour pathological subtype from genomic profile using sparse logistic regression with random effects. Journal of Applied Statistics, 48(4):605–622, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pal:2021:ORC

[4064] Ayan Pal, Sharmishtha Mitra, and Debasis Kundu. Order restricted classical inference of a Weibull multiple stepstress model. Journal of Applied Statistics, 48(4):623–645, April 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ownuk:2021:HSH

[4065] Jamil Ownuk, Hossein Baghishani, and Ahmad Nezakati. Heavy or semi-heavy tail, that is the question. Journal of Applied Statistics, 48(4):646–668, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2021:TQL

[4066] Guangbao Guo. Taylor quasi-likelihood for limited generalized linear models. Journal of Applied Statistics, 48(4):669–692, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nduka:2021:REU

[4067] Uchenna C. Nduka, Tobias E. Ugah, and Chinyeaka H. Izunobi. Robust estimation using multivariate t innovations for vector autoregressive models via ECM algorithm. Journal of Applied Statistics, 48(4):693–711, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chakraborty:2021:DAG

[4068] Subrata Chakraborty, Dhrubajyoti Chakravarty, Josmar Mazucheli, and Wesley Bertoli. A discrete analog of Gumbel distribution: properties, parameter estimation and applications. Journal of Applied Statistics, 48(4): 712–737, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2021:JMB

[4069] Yang-Jin Kim. Joint model for bivariate zero-inflated recurrent event data with terminal events. Journal of Applied Statistics, 48(4):738–749, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2021:MMT

[4070] Qingzhao Yu and Bin Li. A multivariate multiple third-variable effect analysis with an application to explore racial and ethnic disparities in obesity. Journal of Applied Statistics, 48(4):750–764, April 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tabrizi:2021:GLM

[4071] Elham Tabrizi, Ehsan Bahrami Samani, and Mojtaba Ganjali. General location multivariate latent variable models for mixed correlated bounded continuous, ordinal, and nominal responses with non-ignorable missing data. Journal of Applied Statistics, 48(5):765–785, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alqawba:2021:CBM

[4072] Mohammed Alqawba and Norou Diawara. Copula-based Markov zeroinflated count time series models with application. Journal of Applied Statistics, 48(5):786–803, May 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Puindi:2021:DSM

[4073] António Casimiro Puindi and Maria Eduarda Silva. Dynamic structural models with covariates for short-term forecasting of time series with complex seasonal patterns. Journal of Applied Statistics, 48(5):804–826, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maneejuk:2021:STL

[4074] Paravee Maneejuk and Woraphon Yamaka. Significance test for linear regression: how to test without P - values? Journal of Applied Statistics, 48(5):827–845, May 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Du:2021:RAC

[4075] Mingyue Du, Qingning Zhou, Shishun Zhao, and Jianguo Sun. Regression analysis of case-cohort studies in the presence of dependent interval censoring. Journal of Applied Statistics, 48(5):846–865, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jung:2021:MCL

[4076] Yoonsuh Jung, Steven N. MacEachern, and Hang Joon Kim. Modified check loss for efficient estimation via model selection in quantile regression. Journal of Applied Statistics, 48(5):866–886, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2021:TVF

[4077] Mingjing Chen, Xiangyong Tan, and Jian Wu. Time varying factor models with possibly strongly correlated noises. Journal of Applied Statistics, 48(5):887–906, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pires:2021:ICD

[4078] Magda Carvalho Pires, Enrico Antônio Colosimo, Guilherme Augusto Veloso, and Raquel de Souza Borges Ferreira. Interval-censored data with misclassification: a Bayesian approach. Journal of Applied Statistics, 48(5):907–923, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kuran:2021:IMP

[4079] Özge Kuran and M. Revan Özkale. Improvement of mixed predictors in linear mixed models. Journal of Applied Statistics, 48(5):924–942, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vecerova:2021:IIO

[4080] Sárka Vecerová, Arnost Komárek, Luk Bruyneel, and Emmanuel Lesaffre. Identifying influential observations in a Bayesian multi-level mediation model. Journal of Applied Statistics, 48(5): 943–960, May 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2021:OWT

[4081] Xu Guo, Yan Wang, Niwen Zhou, and Xuehu Zhu. Optimal weighted twosample t-test with partially paired data in a unified framework. Journal of Applied Statistics, 48(6):961–976, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Harun:2021:BSM

[4082] Nusrat Harun, Bo Cai, and Yu Shen. A Bayesian semiparametric method for analyzing length-biased data. Journal of Applied Statistics, 48(6):977–992, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2021:LBC

[4083] Yu Zhang and Xiangzhong Fang. On the lengths of t-based confidence intervals. Journal of Applied Statistics, 48(6):993–1008, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Phadnis:2021:SSC

[4084] Milind A. Phadnis and Matthew S.

Mayo. Sample size calculations for noninferiority trials for time-to-event data using the concept of proportional time. Journal of Applied Statistics, 48(6):1009–1032, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2021:SIS

[4085] Zhensheng Huang and Wen Lou. Statistical inferences for single-index models with measurement errors. Journal of Applied Statistics, 48(6):1033–1052, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Paynter:2021:TPS

[4086] Alex Paynter and Amy D. Willis. Tuning parameter selection for a penalized estimator of species richness. Journal of Applied Statistics, 48(6):1053–1070, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nekoukhou:2021:BDI

[4087] Vahid Nekoukhou, Ashkan Khalifeh, and Hamid Bidram. A bivariate discrete inverse resilience family of distributions with resilience marginals. Journal of Applied Statistics, 48(6):1071– 1090, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xia:2021:SED

[4088] Fang Xia, Stephen L. George, Jing Ning, Liang Li, and Xuelin Huang. A signature enrichment design with Bayesian adaptive randomization. Journal of Applied Statistics, 48(6):

1091–1110, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghasemi:2021:DIP

[4089] Erfan Ghasemi, Alireza Akbarzadeh Baghban, Farid Zayeri, Asma Pourhoseingholi, and Seyed Mohammadreza Safavi. A doubly-inflated Poisson regression for correlated count data. Journal of Applied Statistics, 48(6): 1111–1127, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Muse:2021:CDC

[4090] Rhoda Nandai Muse and Satheesh Aradhyula. Correlated discrete and continuous outcomes with endogeneity and lagged effects: past season yield impact on improved corn seed adoption. Journal of Applied Statistics, 48 (6):1128–1153, June 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Teimouri:2021:EAM

[4091] Mahdi Teimouri. EM algorithm for mixture of skew-normal distributions fitted to grouped data. Journal of Applied Statistics, 48(7):1154–1179, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Klein:2021:TAO

[4092] Ingo Klein and Monika Doll. Tests on asymmetry for ordered categorical variables. Journal of Applied Statistics, 48(7):1180–1198, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Simkova:2021:CIB

[4093] T. Simková. Confidence intervals based on L-moments for quantiles of the GP and GEV distributions with application to market-opening asset prices data. Journal of Applied Statistics, 48(7):1199–1226, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tripathi:2021:DGA

[4094] Harsh Tripathi, Sanku Dey, and Mahendra Saha. Double and group acceptance sampling plan for truncated life test based on inverse log-logistic distribution. Journal of Applied Statistics, 48(7):1227–1242, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shongwe:2021:CMS

[4095] Sandile Charles Shongwe, Jean-Claude Malela-Majika, and Philippe Castagliola. A combined mixed-s-skip sampling strategy to reduce the effect of autocorrelation on the X̄ scheme with and without measurement errors. Journal of Applied Statistics, 48(7):1243–1268, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See correction [4100].

Nasini:2021:PID

[4096] Stefano Nasini and Victor Martínez de Albéniz. Pairwise influences in dynamic choice: network-based model and application. Journal of Applied Statistics, 48(7):1269–1302, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bordbar:2021:TIM

[4097] Pejman Bordbar and Sodeif Ahadpour. Type-I intermittency from Markov binary block visibility graph perspective. Journal of Applied Statistics, 48 (7):1303–1318, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alevizakos:2021:MZI

[4098] Vasileios Alevizakos and Christos Koukouvinos. Monitoring of zeroinflated binomial processes with a DEWMA control chart. Journal of Applied Statistics, 48(7):1319–1338, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sewell:2021:POC

[4099] Daniel K. Sewell, Journey Penney, Melissa Jay, Ying Zhang, and Jane S. Paulsen. Predicting an optimal composite outcome variable for Huntington’s disease clinical trials. Journal of Applied Statistics, 48(7):1339–1348, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2021:CCM

[4100] Anonymous. Correction: A combined mixed-s-skip sampling strategy to reduce the effect of autocorrelation on the X̄ scheme with and without measurement errors. Journal of Applied Statistics, 48(7):i, July 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [4095].

Jin:2021:ESI

[4101] Zhezhen Jin and Jianguo Sun. Editorial to special issue Frontiers of Data Analysis. Journal of Applied Statistics, 48 (8):1349–1351, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

He:2021:NBS

[4102] Xinwei He, Xiaoqiang Sun, and Yongzhao Shao. Network-based survival analysis to discover target genes for developing cancer immunotherapies and predicting patient survival. Journal of Applied Statistics, 48(8): 1352–1373, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2021:EES

[4103] Myeonggyun Lee, Anne Zeleniuch-Jacquotte, and Mengling Liu. Empirical evaluation of sub-cohort sampling designs for risk prediction modeling. Journal of Applied Statistics, 48 (8):1374–1401, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2021:GLR

[4104] Yang Liu, JunJia Zhu, and Dennis K. J.

Lin. A generalized likelihood ratio test for monitoring profile data. Journal of Applied Statistics, 48(8):1402–1415, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2021:ECB

[4105] Antai Wang, Xieyang Jia, and Zhezhen Jin. Estimation of the cumulative baseline hazard function for dependently right-censored failure time data. Journal of Applied Statistics, 48(8): 1416–1428, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2021:BAB

[4106] Chunjie Wang, Jingjing Jiang, Linlin Luo, and Shuying Wang. Bayesian analysis of the Box–Cox transformation model based on left-truncated and right-censored data. Journal of Applied Statistics, 48(8):1429–1441, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2021:SCI

[4107] Zhengyu Yang, Guo-Liang Tian, Xiaobin Liu, and Chang-Xing Ma. Simultaneous confidence interval construction for many-to-one comparisons of proportion differences based on correlated paired data. Journal of Applied Statistics, 48(8):1442–1456, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2021:DHR

[4108] Bojuan Barbara Zhao and Ruijuan Su. Determinants of the heavily righttailed residential housing price in Tianjin. Journal of Applied Statistics, 48 (8):1457–1474, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2021:DOA

[4109] Yuna Zhao, Dennis K. J. Lin, and Min-Qian Liu. Designs for order-of-addition experiments. Journal of Applied Statistics, 48(8):1475–1495, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhu:2021:MLE

[4110] Liang Zhu, Xingwei Tong, Dingjiao Cai, Yimei Li, Ryan Sun, Deo K. Srivastava, and Melissa M. Hudson.

Maximum likelihood estimation for the proportional odds model with mixed interval-censored failure time data. Journal of Applied Statistics, 48(8): 1496–1512, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2021:TVM

[4111] Jaehong Yu and Hua Zhong. Time varying mixed effects model with fused lasso regularization. Journal of Applied Statistics, 48(8):1513–1526, August 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Singh:2021:OAS

[4112] Satya Prakash Singh and Pradeep Yadav. Optimal allocation of subjects in a matched pair cluster-randomized trial with fixed number of heterogeneous clusters. Journal of Applied Statistics, 48(9):1527–1540, September 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hoang-Nguyen-Thuy:2021:EPC

[4113] Ngan Hoang-Nguyen-Thuy and K. Krishnamoorthy. Estimation of the probability content in a specified interval using fiducial approach. Journal of Applied Statistics, 48(9):1541–1558, September 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alheety:2021:NKS

[4114] M. I. Alheety, Kristofer Månsson, and B. M. Golam Kibria. A new kind of stochastic restricted biased estimator for logistic regression model. Journal of Applied Statistics, 48(9): 1559–1578, September 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsai:2021:CLC

[4115] Shin-Fu Tsai and Tse-Le Huang. Confidence limits for conformance proportions in normal mixture models. Journal of Applied Statistics, 48(9): 1579–1602, September 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Song:2021:SSM

[4116] Kai-Sheng Song. Simultaneous statistical modelling of excess zeros, over/underdispersion, and multimodality with applications in hotel industry. Journal of Applied Statistics, 48 (9):1603–1627, September 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ye:2021:OCC

[4117] Yunfei Ye and Dong Han. An optimal control chart for finite matrix sequences at some unknown change point. Journal of Applied Statistics, 48 (9):1628–1643, September 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cammarota:2021:VSI

[4118] Camillo Cammarota and Alessandro Pinto. Variable selection and importance in presence of high collinearity: an application to the prediction of lean body mass from multi-frequency bioelectrical impedance. Journal of Applied Statistics, 48(9):1644–1658, September 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Burgard:2021:ENP

[4119] J. P. Burgard, J. Krause, and R. Münnich. An elastic net penalized small area model combining unit- and area-level data for regional hypertension prevalence estimation. Journal of Applied Statistics, 48(9):1659–1674, September 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Petersen:2021:EDS

[4120] Anne Helby Petersen, Bo Markussen, and Karl Bang Christensen. Exploratory data structure comparisons: three new visual tools based on principal component analysis. Journal of Applied Statistics, 48(9):1675–1695, September 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Carpita:2021:DAB

[4121] Maurizio Carpita and Silvia Golia. Discovering associations between players’ performance indicators and matches’ results in the European Soccer Leagues. Journal of Applied Statistics, 48(9): 1696–1711, September 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karamikabir:2021:WTB

[4122] H. Karamikabir, M. Afshari, and F. Lak. Wavelet threshold based on Stein’s unbiased risk estimators of restricted location parameter in multivariate normal. Journal of Applied Statistics, 48(10):1712–1729, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2021:BMF

[4123] Joseph Wu, Mayetri Gupta, Amira I. Hussein, and Louis Gerstenfeld. Bayesian modeling of factorial timecourse data with applications to a bone aging gene expression study. Journal of Applied Statistics, 48(10):1730–1754, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2021:MFS

[4124] Jing Zhang and Yanyan Liu. Modelfree slice screening for ultrahighdimensional survival data. Journal of Applied Statistics, 48(10):1755–1774, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wickramasinghe:2021:MBE

[4125] Lahiru Wickramasinghe, Alexandre Leblanc, and Saman Muthukumarana. Model-based estimation of baseball batting metrics. Journal of Applied Statistics, 48(10):1775–1797, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Azhad:2021:ECL

[4126] Qazi J. Azhad, Mohd. Arshad, and Amit Kumar Misra. Estimation of common location parameter of several heterogeneous exponential populations based on generalized order statistics. Journal of Applied Statistics, 48(10): 1798–1815, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oyedele:2021:EBM

[4127] Opeoluwa F. Oyedele. Extension of biplot methodology to multivariate regression analysis. Journal of Applied Statistics, 48(10):1816–1832, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2021:QSI

[4128] Soyoung Park and Alicia Carriquiry.

Quantifying the similarity of 2D images using edge pixels: an application to the forensic comparison of footwear impressions. Journal of Applied Statistics, 48(10):1833–1860, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhou:2021:MAC

[4129] Thomas J. Zhou, Sughra Raza, and Kerrie P. Nelson. Methods of assessing categorical agreement between correlated screening tests in clinical studies. Journal of Applied Statistics, 48 (10):1861–1881, October 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Puurbalanta:2021:CGG

[4130] Richard Puurbalanta. A Clipped Gaussian Geo-Classification model for poverty mapping. Journal of Applied Statistics, 48(10):1882–1895, October 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bastos:2021:BSS

[4131] Fernando de Souza Bastos and Wagner Barreto-Souza. Birnbaum-Saunders sample selection model. Journal of Applied Statistics, 48(11):1896–1916, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gardner-Lubbe:2021:LDA

[4132] Sugnet Gardner-Lubbe. Linear discriminant analysis for multiple functional data analysis. Journal of Applied Statistics, 48(11):1917–1933, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2021:VEB

[4133] Xingli Yang, Yu Wang, Wennan Yan, and Jihong Li. Variance estimation based on blocked 3×2 cross-validation in high-dimensional linear regression. Journal of Applied Statistics, 48(11): 1934–1947, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Eliwa:2021:EOC

[4134] M. S. Eliwa, M. El-Morshedy, and Sajid Ali. Exponentiated odd Chen-g family of distributions: statistical properties, Bayesian and non-Bayesian estimation with applications. Journal of Applied Statistics, 48(11):1948–1974, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shirozhan:2021:DCI

[4135] Masoumeh Shirozhan and Mehrnaz Mohammadpour. A dependent counting INAR model with serially dependent innovation. Journal of Applied Statistics, 48(11):1975–1997, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Silva:2021:ALR

[4136] Ana R. S. Silva, Caio L. N. Azevedo, Jorge L. Bazán, and Juvêncio S. Nobre. Augmented-limited regression models with an application to the study of the risk perceived using continuous scales. Journal of Applied Statistics, 48(11): 1998–2021, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elliott:2021:FMM

[4137] Corrine F. Elliott, Joshua W. Lambert, Arnold J. Stromberg, Pei Wang, Ting Zeng, and Katherine L. Thompson. Feasibility as a mechanism for model identification and validation. Journal of Applied Statistics, 48(11): 2022–2041, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sun:2021:TCB

[4138] Zequn Sun and Thomas J. Fisher. Testing for correlation between two time series using a parametric bootstrap. Journal of Applied Statistics, 48(11): 2042–2063, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2021:SRN

[4139] Anonymous. Statement of retraction:

A novel alpha power transformed exponential distribution with real-life applications. Journal of Applied Statistics, 48(11):2064, November 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ijaz:2021:RAN

[4140] Muhammad Ijaz, Wali Khan Mashwani, Atilla Göktas, and Yuksel Akay Unvan. RETRACTED ARTICLE: A novel alpha power transformed exponential distribution with real-life applications. Journal of Applied Statistics, 48(11):I–XVI, November 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Slaoui:2021:TNN

[4141] Yousri Slaoui. Two new nonparametric kernel distribution estimators based on a transformation of the data. Journal of Applied Statistics, 48(12): 2065–2091, December 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sulewski:2021:EWH

[4142] Piotr Sulewski. Equal-bin-width histogram versus equal-bin-count histogram. Journal of Applied Statistics, 48(12):2092–2111, December 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Davies:2021:SEA

[4143] Katherine Davies, Suvra Pal, and Joynob A. Siddiqua. Stochastic EM algorithm for generalized exponential cure rate model and an empirical study. Journal of Applied Statistics, 48(12): 2112–2135, December 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Acitas:2021:MML

[4144] Sukru Acitas, Ismail Yenilmez, Birdal Senoglu, and Yeliz Mert Kantar. Modified maximum likelihood estimator under the Jones and Faddy’s skew t - error distribution for censored regression model. Journal of Applied Statistics, 48(12):2136–2151, December 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sheng:2021:MEE

[4145] Chong Sheng, Matthias Hwai Yong Tan, and Lu Zou. Maximum expected entropy transformed Latin hypercube designs. Journal of Applied Statistics, 48(12):2152–2177, December 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tran:2021:MCV

[4146] Phuong Hanh Tran, Cédric Heuchenne, Huu Du Nguyen, and Hélène Marie. Monitoring coefficient of variation using one-sided run rules control charts in the presence of measurement errors. Journal of Applied Statistics, 48(12): 2178–2204, December 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2021:ENP

[4147] Q. F. Xu, X. H. Ding, C. X. Jiang, K. M. Yu, and L. Shi. An elastic-net penalized expectile regression with applications. Journal of Applied Statistics, 48(12):2205–2230, December 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goktas:2021:ESI

[4148] A. Göktas and Ö. Akkus. Editorial to the special issue: Recent statistical methods for data analysis, applied economics, business & finance. Journal of Applied Statistics, 48(13-15):2231– 2238, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gunduz:2021:OBE

[4149] Necla Gündüz and Celal Aydin. Optimal bandwidth estimators of kernel density functionals for contaminated data. Journal of Applied Statistics, 48(13-15):2239–2258, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yilmaz:2021:CDE

[4150] Asuman Yilmaz, Mahmut Kara, and Onur Özdemir. Comparison of different estimation methods for extreme value distribution. Journal of Applied Statistics, 48(13-15):2259–2284, 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Eini:2021:TCM

[4151] Esmat Jamshidi Eini and Hamid Khaloozadeh. Tail conditional moment for generalized skew-elliptical distributions. Journal of Applied Statistics, 48(13-15):2285–2305, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arnold:2021:SID

[4152] Barry C. Arnold and B. G. Manjunath. Statistical inference for distributions with one Poisson conditional. Journal of Applied Statistics, 48(13-15):2306– 2325, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Turksen:2021:NPP

[4153] Özlem Türksen. A novel perspective for parameter estimation of seemingly unrelated nonlinear regression. Journal of Applied Statistics, 48(13-15):2326– 2347, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yamaguchi:2021:MEV

[4154] Naoya Yamaguchi, Yuka Yamaguchi, and Ryuei Nishii. Minimizing the expected value of the asymmetric loss function and an inequality for the variance of the loss. Journal of Applied Statistics, 48(13-15):2348–2368, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cakmak:2021:OBR

[4155] Berivan Çakmak and Fatma Zehra Dogru. Optimal b-robust estimators for the parameters of the power Lindley distribution. Journal of Applied Statistics, 48(13-15):2369–2388, 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akdur:2021:ULM

[4156] Hatice Tul Kubra Akdur. Unit-Lindley mixed-effect model for proportion data. Journal of Applied Statistics, 48(13-15):2389–2405, 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Evkaya:2021:CVM

[4157] O. Ozan Evkaya, Ceylan Yozgatligil, and A. Sevtap Selcuk-Kestel. CD-vine model for capturing complex dependence. Journal of Applied Statistics, 48(13-15):2406–2420, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Taylan:2021:NOD

[4158] Pakize Taylan, Fatma Yerlikaya-Özkurt, Burcu Bilgiç Uçak, and Gerhard-Wilhelm Weber. A new outlier detection method based on convex optimization: application to diagnosis of Parkinson’s disease. Journal of Applied Statistics, 48(13-15):2421–2440, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deliorman:2021:BPS

[4159] G. Deliorman and D. Inan. Binary particle swarm optimization as a detection tool for influential subsets in linear regression. Journal of Applied Statistics, 48(13-15):2441–2456, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Goktas:2021:NRR

[4160] Atila Göktas, Özge Akkus, and Aykut Kuvat. A new robust ridge parameter estimator based on search method for linear regression model. Journal of Applied Statistics, 48(13-15):2457– 2472, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ebegil:2021:SSE

[4161] Meral Ebegil, Yaprak Arzu Özdemir, and Fikri Gökpinar. Some shrinkage estimators based on median ranked set sampling. Journal of Applied Statistics, 48(13-15):2473–2498, 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sancar:2021:NAE

[4162] Nuriye Sancar and Deniz Inan. A new alternative estimation method for Liu-type logistic estimator via particle swarm optimization: an application to data of collapse of Turkish commercial banks during the Asian financial crisis. Journal of Applied Statistics, 48(13-15):2499–2514, 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Altun:2021:LRM

[4163] Emrah Altun. The Lomax regression model with residual analysis: an application to insurance data. Journal of Applied Statistics, 48(13-15):2515– 2524, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Munoz-Pichardo:2021:MPR

[4164] J. M. Muñoz-Pichardo, R. Pino-Mejías, J. García-Heras, F. Ruiz-Muñoz, and M. Luz González-Regalado. A multivariate Poisson regression model for count data. Journal of Applied Statistics, 48(13-15):2525–2541, 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gil-Alana:2021:TFU

[4165] Luis A. Gil-Alana and OlaOluwa S.

Yaya. Testing fractional unit roots with non-linear smooth break approximations using Fourier functions. Journal of Applied Statistics, 48(13-15):2542– 2559, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Eroglu:2021:PVR

[4166] Burak A. Erolu and Selim Yildirim.

On the performance of the variance ratio unit root tests with flexible Fourier form. Journal of Applied Statistics, 48(13-15):2560–2579, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cetin:2021:CFA

[4167] Beyza Cetin and Idil Yavuz. Comparison of forecast accuracy of Ata and exponential smoothing. Journal of Applied Statistics, 48(13-15):2580– 2590, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Evkaya:2021:FDU

[4168] O. Ozan Evkaya and Fatma Sevinç Kurnaz. Forecasting drought using neural network approaches with transformed time series data. Journal of Applied Statistics, 48(13-15):2591–2606, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Balci:2021:HCU

[4169] Mehmet Ali Balci, Ömer Akgüller, and Serdar Can Güzel. Hierarchies in communities of UK stock market from the perspective of Brexit. Journal of Applied Statistics, 48(13-15):2607–2625, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yildirim:2021:RSC

[4170] Julide Yildirim, Baris Alpaslan, and Erdener Emin Eker. The role of social capital in environmental protection efforts: evidence from Turkey. Journal of Applied Statistics, 48(13-15):2626– 2642, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Atukalp:2021:DRB

[4171] M. Esra Atukalp. Determining the relationship between stock return and financial performance: an analysis on Turkish deposit banks. Journal of Applied Statistics, 48(13-15):2643–2657, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Metin:2021:RCW

[4172] Cenker Burak Metin, Sinem Tugba Sahin Tekin, and Yaprak Arzu Özdemir. Restricted calibration and weight trimming approaches for estimation of the population total in business statistics. Journal of Applied Statistics, 48(13-15):2658–2672, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Momtaz:2021:EMD

[4173] Paul P. Momtaz. Econometric models of duration data in entrepreneurship with an application to start-ups’ time-to-funding by venture capitalists (VCs). Journal of Applied Statistics, 48(13-15):2673–2694, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kulekci:2021:ALR

[4174] Bükre Yildirim Külekci and A. Sevtap Selcuk-Kestel. Assessment of longevity risk: credibility approach. Journal of Applied Statistics, 48(13-15):2695– 2713, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2021:SCC

[4175] Ding-Geng Chen, Haipeng Gao, Chuanshu Ji, and Xinguang Chen. Stochastic cusp catastrophe model and its Bayesian computations. Journal of Applied Statistics, 48(13-15):2714– 2733, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ehrhardt:2021:RIM

[4176] Adrien Ehrhardt, Christophe Biernacki, Vincent Vandewalle, Philippe Heinrich, and Sébastien Beben. Reject inference methods in credit scoring. Journal of Applied Statistics, 48(13-15):2734–2754, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Evren:2021:SDI

[4177] Atif Evren, Elif Tuna, Erhan Ustaoglu, and Busra Sahin. Some dominance indices to determine market concentration. Journal of Applied Statistics, 48(13-15):2755–2775, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kabran:2021:TSM

[4178] Fatma Basoglu Kabran and Kamil Demirberk Ünlü. A two-step machine learning approach to predict SP 500 bubbles. Journal of Applied Statistics, 48(13-15):2776–2794, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guler:2021:MLE

[4179] Huseyin Guler and Ebru Ozgur Guler. Mixed Lasso estimator for stochastic restricted regression models. Journal of Applied Statistics, 48(13-15):2795– 2808, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yildirim:2021:TSM

[4180] Asiye Nur Yildirim, Eren Bas, and Erol Egrioglu. Threshold single multiplicative neuron artificial neural networks for non-linear time series forecasting. Journal of Applied Statistics, 48(13-15):2809–2825, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yerlikaya-Ozkurt:2021:EPN

[4181] Fatma Yerlikaya-Özkurt, Pakize Taylan, and Müjgan Tez. Estimation in the partially nonlinear model by continuous optimization. Journal of Applied Statistics, 48(13-15):2826–2846, 2021.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Erdis:2021:MDM

[4182] Atacan Erdis, M. Akif Bakir, and Muhammed I. Jaiteh. A method for detection of Mode-Mixing problem. Journal of Applied Statistics, 48(13-15):2847–2863, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

OBrien:2021:EED

[4183] Timothy E. O’Brien and Jack Silcox.

Efficient experimental design for dose response modelling. Journal of Applied Statistics, 48(13-15):2864–2888, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gerger:2021:TBC

[4184] Atakan Gerger and Ali Riza Firuzan.

Taguchi based case study in the automotive industry: nonconformity decreasing with use of Six Sigma methodology. Journal of Applied Statistics, 48(13-15):2889–2905, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yazici:2021:CSC

[4185] Berna Yazici and Mustafa Cavus. A comparative study of computation approaches of the generalized f-test. Journal of Applied Statistics, 48(13-15): 2906–2919, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rusyda:2021:DMC

[4186] H. A. Rusyda, L. Noviyanti, A. Z.

Soleh, A. Chadidjah, and F. Indrayatna. The design of multiple crop insurance in Indonesia based on revenue risk using the copula model approach. Journal of Applied Statistics, 48(13-15):2920–2930, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Karaibrahimoglu:2021:CAD

[4187] Adnan Karaibrahimoglu, Seren Ayhan, Mustafa Karaagac, and Mehmet Artac. Circular analyses of dates on patients with gastric carcinoma. Journal of Applied Statistics, 48(13-15):2931– 2943, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2021:HMM

[4188] Hao Chen, Minguang Zhang, Lanshan Han, and Alvin Lim. Hierarchical marketing mix models with sign constraints. Journal of Applied Statistics, 48(13-15):2944–2960, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gunduz:2021:USE

[4189] Necla Gündüz and Ernest Fokoué. Understanding students’ evaluations of professors using non-negative matrix factorization. Journal of Applied Statistics, 48(13-15):2961–2981, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deOliveira:2021:AEE

[4190] Ricardo Puziol de Oliveira and Jorge Alberto Achcar. Accurate estimation for extra-Poisson variability assuming random effect models. Journal of Applied Statistics, 48(16):2982–3001, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bakouch:2021:NCS

[4191] Hassan S. Bakouch, Meitner Cadena, and Christophe Chesneau. A new class of skew distributions with climate data analysis. Journal of Applied Statistics, 48(16):3002–3024, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2021:NBP

[4192] Indranil Ghosh, Filipe Marques, and Subrata Chakraborty. A new bivariate Poisson distribution via conditional specification: properties and applications. Journal of Applied Statistics, 48(16):3025–3047, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

doNascimento:2021:RMD

[4193] Fernando Ferraz do Nascimento, Andreson Almeida Azevedo, and Valmaria Rocha da Silva Ferraz. Regression models to dependence for exceedance. Journal of Applied Statistics, 48(16):3048–3059, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guzman:2021:LCR

[4194] Daniel C. F. Guzmán, Clécio S. Ferreira, and Camila B. Zeller. Linear censored regression models with skew scale mixtures of normal distributions. Journal of Applied Statistics, 48(16):3060– 3085, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghosh:2021:BCP

[4195] Pallab Kumar Ghosh. Box–Cox power transformation unconditional quantile regressions with an application on wage inequality. Journal of Applied Statistics, 48(16):3086–3101, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sreedevi:2021:NIP

[4196] E. P. Sreedevi and P. G. Sankaran. Nonparametric inference for panel count data with competing risks. Journal of Applied Statistics, 48(16):3102– 3115, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2021:IKF

[4197] Tao Chen, Qingliang Fan, Kai Liu, and Lingshan Le. Identifying key factors in momentum in basketball games. Journal of Applied Statistics, 48(16):3116– 3129, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2021:PLS

[4198] Jiajia Chen, Xiaoqin Zhang, and Karel Hron. Partial least squares regression with compositional response variables and covariates. Journal of Applied Statistics, 48(16):3130–3149, 2021. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Capdeville:2021:BFM

[4199] Vitor Capdeville, Kelly C. M. Gonçalves, and João B. M. Pereira. Bayesian factor models for multivariate categorical data obtained from questionnaires. Journal of Applied Statistics, 48(16):3150–3173, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guerra:2021:UEW

[4200] Renata Rojas Guerra, Fernando A. Peña-Ramírez, and Marcelo Bourguignon. The unit extended Weibull families of distributions and its applications. Journal of Applied Statistics, 48(16):3174–3192, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2021:MEF

[4201] Sungsu Kim and Ashis SenGupta. Multimodal exponential families of circular distributions with application to daily peak hours of pm2.5 level in a large city. Journal of Applied Statistics, 48(16):3193–3207, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Smith:2021:CDU

[4202] Thomas J. Smith, David A. Walker, and Cornelius M. McKenna. A coefficient of discrimination for use with nominal and ordinal regression models. Journal of Applied Statistics, 48(16): 3208–3219, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kruppa:2021:CSM

[4203] Jochen Kruppa and Ludwig Hothorn.

A comparison study on modeling of clustered and overdispersed count data for multiple comparisons. Journal of Applied Statistics, 48(16):3220–3232, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lopez-Laborda:2021:EEC

[4204] Julio López-Laborda, Carmen Marín-González, and Jorge Onrubia-Fernández. Estimating Engel curves: a new way to improve the SILC-HBS matching process using GLM methods. Journal of Applied Statistics, 48(16):3233–3250, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Azimi:2021:CWF

[4205] Reza Azimi and Mahdy Esmailian. Correction to: The Weibull Fréchet distribution and its applications. Journal of Applied Statistics, 48(16):3251–3252, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [3319].

Anonymous:2021:C

[4206] Anonymous. Correction. Journal of Applied Statistics, 48(16):3253–3254, 2021. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2022:ZIM

[4207] Cindy Feng. Zero-inflated models for adjusting varying exposures: a cautionary note on the pitfalls of using offset. Journal of Applied Statistics, 49(1):1–23, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pardo:2022:EEL

[4208] M. Carmen Pardo, Ying Lu, and Alba M. Franco-Pereira. Extensions of empirical likelihood and chi-squared-based tests for ordered alternatives. Journal of Applied Statistics, 49(1):24– 43, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ma:2022:CMF

[4209] Haiqiang Ma, Jin Yang, Sheng Xu, Chao Liu, and Qinyi Zhang. Combination of multiple functional markers to improve diagnostic accuracy. Journal of Applied Statistics, 49(1):44–63, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Najarzadeh:2022:CCI

[4210] Dariush Najarzadeh. Conservative confidence intervals on multiple correlation coefficient for high-dimensional elliptical data using random projection methodology. Journal of Applied Statistics, 49(1):64–85, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dibaj:2022:EUI

[4211] Shiva S. Dibaj, Alan D. Hutson, Graham W. Warren, and Gregory E. Wilding. Exact unconditional inference for analyzing contingency tables in finite populations. Journal of Applied Statistics, 49(1):86–97, January 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scaldelai:2022:MNA

[4212] D. Scaldelai, L. C. Matioli, S. R. Santos, and M. Kleina. MulticlusterKDE: a new algorithm for clustering based on multivariate kernel density estimation. Journal of Applied Statistics, 49 (1):98–121, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2022:RAM

[4213] Liang Wang, Ke Wu, Yogesh Mani Tripathi, and Chandrakant Lodhi. Reliability analysis of multicomponent stress-strength reliability from a bathtub-shaped distribution. Journal of Applied Statistics, 49(1):122–142, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Esteban:2022:SAE

[4214] María Dolores Esteban, María José Lombardía, Esther López-Vizcaíno, Domingo Morales, and Agustín Pérez. Small area estimation of expenditure means and ratios under a unit-level bivariate linear mixed model. Journal of Applied Statistics, 49(1):143–168, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jana:2022:EPI

[4215] Nabakumar Jana and Samadrita Bera. Estimation of parameters of inverse Weibull distribution and application to multi-component stress-strength model. Journal of Applied Statistics, 49(1):169–194, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prataviera:2022:JRM

[4216] F. Prataviera, A. M. Batista, P. L. Libardi, G. M. Cordeiro, and E. M. M. Ortega. Joint regression modeling of location and scale parameters of the skew t distribution with application in soil chemistry data. Journal of Applied Statistics, 49(1):195–213, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sciandra:2022:MBA

[4217] Mariangela Sciandra and Irene Carola Spera. A model-based approach to Spotify data analysis: a Beta GLMM. Journal of Applied Statistics, 49(1):214–229, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2022:SMD

[4218] Han Yu and Rachael Hageman Blair.

Scalable module detection for attributed networks with applications to breast cancer. Journal of Applied Statistics, 49(1):230–247, January 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vasconcelos:2022:SRM

[4219] J. C. S. Vasconcelos, G. M. Cordeiro, and E. M. M. Ortega. The semiparametric regression model for bimodal data with different penalized smoothers applied to climatology, ethanol and air quality data. Journal of Applied Statistics, 49(1):248–267, January 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arachchige:2022:RAC

[4220] Chandima N. P. G. Arachchige, Luke A. Prendergast, and Robert G. Staudte. Robust analogs to the coefficient of variation. Journal of Applied Statistics, 49(2):268–290, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mante:2022:DGE

[4221] C. Manté, S. Cornu, D. Borschneck, C. Mocuta, and R. van den Bogaert. Detecting the Guttman effect with the help of ordinal correspondence analysis in synchrotron X-ray diffraction data analysis. Journal of Applied Statistics, 49(2):291–316, February 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Loumponias:2022:KFC

[4222] Kostas Loumponias and George Tsaklidis. Kalman filtering with censored measurements. Journal of Applied Statistics, 49(2):317–335, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

F:2022:BGI

[4223] Paulo C. Marques F., Helton Graziadei, and Hedibert F. Lopes. Bayesian generalizations of the integer-valued autoregressive model. Journal of Applied Statistics, 49(2):336–356, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Labadi:2022:BEE

[4224] Luai Al-Labadi and Sean Berry. Bayesian estimation of extropy and goodness of fit tests. Journal of Applied Statistics, 49(2):357–370, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sharma:2022:ETD

[4225] Vikas Kumar Sharma, Sudhanshu V. Singh, and Komal Shekhawat. Exponentiated Teissier distribution with increasing, decreasing and bathtub hazard functions. Journal of Applied Statistics, 49(2):371–393, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahmadi:2022:PAJ

[4226] Mohammad Vali Ahmadi and Mahdi Doostparast. Predictive analysis for joint progressive censoring plans: a Bayesian approach. Journal of Applied Statistics, 49(2):394–410, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tan:2022:FSB

[4227] Hongwei Tan, Guodong Wang, Wendong Wang, and Zili Zhang. Feature selection based on distance correlation: a filter algorithm. Journal of Applied Statistics, 49(2):411–426, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2022:TMO

[4228] Lingzhe Guo and Reza Modarres. Two multivariate online change detection models. Journal of Applied Statistics, 49(2):427–448, February 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ward:2022:MPS

[4229] Caitlin Ward, Jacob Oleson, J. Bruce Tomblin, and Elizabeth Walker. Modeling population and subject-specific growth in a latent trait measured by multiple instruments over time using a hierarchical Bayesian framework. Journal of Applied Statistics, 49(2): 449–465, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jimenez:2022:QTD

[4230] José L. Jiménez. Quantifying treatment differences in confirmatory trials under non-proportional hazards. Journal of Applied Statistics, 49(2): 466–484, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nielsen:2022:SME

[4231] Mikkel Slot Nielsen and Victor Rohde. A surrogate model for estimating extreme tower loads on wind turbines based on random forest proximities. Journal of Applied Statistics, 49 (2):485–497, February 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2022:TVC

[4232] Xiaoxue Li, Stewart J. Anderson, Saul Shiffman, and Bo Zhang. Time-varying coefficient cumulative gap time models for intensive longitudinal ecological momentary assessment data with missingness. Journal of Applied Statistics, 49(2):498–521, February 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bonamente:2022:SAS

[4233] Massimiliano Bonamente and David Spence. A semi-analytical solution to the maximum-likelihood fit of Poisson data to a linear model using the Cash statistic. Journal of Applied Statistics, 49(3):522–552, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anastasopoulou:2022:ECC

[4234] Maria Anastasopoulou and Athanasios C. Rakitzis. EWMA control charts for monitoring correlated counts with finite range. Journal of Applied Statistics, 49(3):553–573, March 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2022:RRE

[4235] Qingguo Tang, Rohana J. Karunamuni, and Boxiao Liu. Regularized robust estimation in binary regression models. Journal of Applied Statistics, 49 (3):574–598, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qin:2022:GFT

[4236] Xinyan Qin, Jiao Yu, and Wenhao Gui.

Goodness-of-fit test for exponentiality based on spacings for general progressive Type-II censored data. Journal of Applied Statistics, 49(3):599–620, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2022:NPH

[4237] Junhyung Park, Adam W. Chaffee, Ryan J. Harrigan, and Frederic Paik Schoenberg. A non-parametric Hawkes model of the spread of Ebola in west Africa. Journal of Applied Statistics, 49(3):621–637, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baghfalaki:2022:BSP

[4238] T. Baghfalaki, M. Ganjali, A. Kabir, and A. Pazouki. A Bayesian shared parameter model for joint modeling of longitudinal continuous and binary outcomes. Journal of Applied Statistics, 49(3):638–655, March 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jayalath:2022:SAI

[4239] Kalanka. P. Jayalath and Raj S.

Chhikara. Survival analysis for the inverse Gaussian distribution with the Gibbs sampler. Journal of Applied Statistics, 49(3):656–675, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kiesse:2022:BIB

[4240] T. Senga Kiessé, Etienne Rivot, Christophe Jaeger, and Joël Aubin. Bayesian inference in based-kernel regression: comparison of count data of condition factor of fish in pond systems. Journal of Applied Statistics, 49(3):676–693, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hamdeni:2022:PHM

[4241] Tasnime Hamdeni and Soufiane Gasmi. A proportional-hazards model for survival analysis and long-term survivors modeling: application to amyotrophic lateral sclerosis data. Journal of Applied Statistics, 49(3):694–708, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sheikhi:2022:GTE

[4242] Ayyub Sheikhi, Fatemeh Bahador, and Mohammad Arashi. On a generalization of the test of endogeneity in a two stage least squares estimation. Journal of Applied Statistics, 49(3):709–721, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baillo:2022:NSC

[4243] A. Baíllo and J. E. Chacón. A new selection criterion for statistical home range estimation. Journal of Applied Statistics, 49(3):722–737, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oh:2022:RSA

[4244] Man-Suk Oh and Chee Kyung Park. Regional source apportionment of PM2.5 in Seoul using Bayesian multivariate receptor model. Journal of Applied Statistics, 49(3):738–751, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Azman:2022:GFL

[4245] Shafiqah Azman and Dharini Pathmanathan. The GLM framework of the Lee–Carter model: a multi-country study. Journal of Applied Statistics, 49(3):752–763, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bastien:2022:SMS

[4246] B. Bastien, T. Boukhobza, H. Dumond, A. Gégout-Petit, A. Muller-Gueudin, and C. Thiébaut. A statistical methodology to select covariates in high-dimensional data under dependence. Application to the classification of genetic profiles in oncology. Journal of Applied Statistics, 49 (3):764–781, March 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiong:2022:TEB

[4247] Peihan Xiong, Weiwei Zhuang, and Guoxin Qiu. Testing exponentiality based on the extropy of record values. Journal of Applied Statistics, 49(4):782–802, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hatefi:2022:EEC

[4248] Armin Hatefi and Amirhossein Alvandi. Efficient estimators with categorical ranked set samples: estimation procedures for osteoporosis. Journal of Applied Statistics, 49(4):803–818, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Malec:2022:RDC

[4249] Lukás Malec. On the rank-deficient canonical correlation technique solved by analytic spectral decomposition. Journal of Applied Statistics, 49(4): 819–830, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zavez:2022:MEM

[4250] Alexis Zavez, Emeir M. McSorley, Alison J. Yeates, and Sally W. Thurston. Modeling the effects of multiple exposures with unknown group memberships: a Bayesian latent variable approach. Journal of Applied Statistics, 49(4):831–857, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fang:2022:SMC

[4251] Zijian Fang, Peng Zhao, Maochao Xu, Shouhuai Xu, Taizhong Hu, and Xing Fang. Statistical modeling of computer malware propagation dynamics in cyberspace. Journal of Applied Statistics, 49(4):858–883, April 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Agogo:2022:TPR

[4252] George O. Agogo and Alexander K. Muoka. A three-part regression calibration to handle excess zeroes, skewness and heteroscedasticity in adjusting for measurement error in dietary intake data. Journal of Applied Statistics, 49(4):884–901, April 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2022:REM

[4253] Yuexia Zhang, Guoyou Qin, Zhongyi Zhu, and Bo Fu. Robust estimation of models for longitudinal data with dropouts and outliers. Journal of Applied Statistics, 49(4):902–925, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pathak:2022:BIW

[4254] Anurag Pathak, Manoj Kumar, Sanjay Kumar Singh, and Umesh Singh. Bayesian inference: Weibull Poisson model for censored data using the expectation-maximization algorithm and its application to bladder cancer data. Journal of Applied Statistics, 49(4):926–948, April 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scollnik:2022:BAE

[4255] David P. M. Scollnik. Bayesian analyses of an exponential-Poisson and related zero augmented type models. Journal of Applied Statistics, 49(4):949–967, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Govindarajulu:2022:OPP

[4256] Usha Govindarajulu and Thaddeus Tarpey. Optimal partitioning for the proportional hazards model. Journal of Applied Statistics, 49(4):968–987, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baek:2022:RFD

[4257] Seungchul Baek, Yewon Kim, Junyong Park, and Jong Soo Lee. Revisit to functional data analysis of sleeping energy expenditure. Journal of Applied Statistics, 49(4):988–1002, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Unlu:2022:MMP

[4258] Hande Konsuk Ünlü, Derek S. Young, Ayten Yigiter, and L. Hilal Özcebe. A mixture model with Poisson and zerotruncated Poisson components to analyze road traffic accidents in Turkey. Journal of Applied Statistics, 49(4): 1003–1017, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boscaino:2022:MSP

[4259] Giovanni Boscaino, Gianluca Sottile, and Giada Adelfio. Migration and students’ performance: detecting geographical differences following a curves clustering approach. Journal of Applied Statistics, 49(4):1018–1032, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Presno:2022:NPS

[4260] María José Presno, Manuel Landajo, and Paula Fernandez-Gonzalez. Nonparametric panel stationarity testing with an application to crude oil production. Journal of Applied Statistics, 49(4):1033–1048, April 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xavier:2022:SSR

[4261] Thomas Xavier and Joby K. Jose. Stress-strength reliability estimation involving paired observation with ties using bivariate exponentiated halflogistic model. Journal of Applied Statistics, 49(5):1049–1064, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Seitshiro:2022:QMR

[4262] M. B. Seitshiro and H. P. Mashele.

Quantification of model risk that is caused by model misspecification. Journal of Applied Statistics, 49(5): 1065–1085, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Murugeswari:2022:ODT

[4263] N. Murugeswari, P. Jeyadurga, and S. Balamurali. Optimal designing of two-level skip-lot sampling reinspection plan. Journal of Applied Statistics, 49(5):1086–1104, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fang:2022:SSS

[4264] Kuangnan Fang, Peng Wang, Xiaochen Zhang, and Qingzhao Zhang. Structured sparse support vector machine with ordered features. Journal of Applied Statistics, 49(5):1105–1120, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stark:2022:TDS

[4265] Florian Stark and Sven Otto. Testing and dating structural changes in copula-based dependence measures. Journal of Applied Statistics, 49(5): 1121–1139, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hao:2022:VBI

[4266] Xiaojuan Hao, Kent M. Eskridge, and Dong Wang. Variational Bayesian inference for association over phylogenetic trees for microorganisms. Journal of Applied Statistics, 49(5):1140–1153, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2022:UDS

[4267] Li-Pang Chen. Ultrahigh-dimensional sufficient dimension reduction for censored data with measurement error in covariates. Journal of Applied Statistics, 49(5):1154–1178, May 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beyaztas:2022:RBP

[4268] Ufuk Beyaztas and Han Lin Shang. Robust bootstrap prediction intervals for univariate and multivariate autoregressive time series models. Journal of Applied Statistics, 49(5):1179–1202, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Koloydenko:2022:MSB

[4269] Alexey Koloydenko, Kristi Kuljus, and Jüri Lember. MAP segmentation in Bayesian hidden Markov models: a case study. Journal of Applied Statistics, 49(5):1203–1234, May 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nguyen:2022:AIS

[4270] T. H. A. Nguyen, T. Laurent, C. Thomas-Agnan, and A. Ruiz-Gazen. Analyzing the impacts of socio-economic factors on French departmental elections with CoDa methods. Journal of Applied Statistics, 49 (5):1235–1251, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Oliveira:2022:BBS

[4271] Maria Ioneris Oliveira, Michelli Barros, Joelson Campos, and Francisco José A. Cysneiros. Bivariate Birnbaum–Saunders accelerated lifetime model: estimation and diagnostic analysis. Journal of Applied Statistics, 49(5):1252–1276, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ossai:2022:RTE

[4272] Everestus O. Ossai, Mbanefo S.

Madukaife, and Abimibola V. Oladugba. A review of tests for exponentiality with Monte Carlo comparisons. Journal of Applied Statistics, 49(5):1277– 1304, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barkhordar:2022:BAT

[4273] Zahra Barkhordar, Mohsen Maleki, Zahra Khodadadi, Darren Wraith, and Farajollah Negahdari. A Bayesian approach on the two-piece scale mixtures of normal homoscedastic nonlinear regression models. Journal of Applied Statistics, 49(5):1305–1322, May 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2022:AAM

[4274] Yonghui Liu, Chaoxuan Mao, Víctor Leiva, Shuangzhe Liu, and Waldemiro A. Silva Neto. Asymmetric autoregressive models: statistical aspects and a financial application under COVID-19 pandemic. Journal of Applied Statistics, 49(5):1323–1347, May 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2022:HPI

[4275] Jiefei Wang and Jeffrey C. Miecznikowski. High precision implementation of Steck’s recursion method for use in goodness-of-fit tests. Journal of Applied Statistics, 49(6):1348–1363, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hemavathi:2022:SAT

[4276] M. Hemavathi, Eldho Varghese, Shashi Shekhar, and Seema Jaggi. Sequential asymmetric third order rotatable designs (SATORDs). Journal of Applied Statistics, 49(6):1364–1381, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Munezero:2022:DBA

[4277] Parfait Munezero and Gebrenegus Ghilagaber. Dynamic Bayesian adjustment of anticipatory covariates in retrospective data: application to the effect of education on divorce risk. Journal of Applied Statistics, 49(6):1382–1401, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qi:2022:APP

[4278] Qianya Qi, Li Yan, and Lili Tian. Analyzing partially paired data: when can the unpaired portion(s) be safely ignored? Journal of Applied Statistics, 49(6):1402–1420, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-kaabawi:2022:BHM

[4279] Zainab Al-kaabawi, Yinghui Wei, and Rana Moyeed. Bayesian hierarchical models for linear networks. Journal of Applied Statistics, 49(6):1421–1448, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jing:2022:DMD

[4280] Hongying Jing, Jian Li, and Kaizong Bai. Directional monitoring and diagnosis for covariance matrices. Journal of Applied Statistics, 49(6):1449–1464, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2022:AKS

[4281] Xin Liu and Wenqing He. Adaptive kernel scaling support vector machine with application to a prostate cancer image study. Journal of Applied Statistics, 49(6):1465–1484, June 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Curley:2022:NME

[4282] B. Curley. A nonlinear measurement error model and its application to describing the dependency of health outcomes on dietary intake. Journal of Applied Statistics, 49(6):1485–1518, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lalloue:2022:SCB

[4283] Benoît Lalloué, Jean-Marie Monnez, and Eliane Albuisson. Streaming constrained binary logistic regression with online standardized data. Journal of Applied Statistics, 49(6):1519–1539, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Engmann:2022:OCE

[4284] Gideon Mensah Engmann and Dong Han. The optimized CUSUM and EWMA multi-charts for jointly detecting a range of mean and variance change. Journal of Applied Statistics, 49(6):1540–1558, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Papukdee:2022:PLA

[4285] Nipada Papukdee, Jeong-Soo Park, and Piyapatr Busababodhin. Penalized likelihood approach for the fourparameter kappa distribution. Journal of Applied Statistics, 49(6):1559–1573, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Burzala:2022:PTN

[4286] Milda Maria Burzala. The process of transferring negative impulses in capital markets — a wavelet analysis. Journal of Applied Statistics, 49 (6):1574–1597, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Carvalho:2022:RAB

[4287] Marcela de Marillac Carvalho and Thelma Sáfadi. Risk analysis in the Brazilian stock market: copula-APARCH modeling for value-at-risk. Journal of Applied Statistics, 49(6): 1598–1610, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2022:CNa

[4288] Anonymous. Correction notice. Journal of Applied Statistics, 49(6):1611, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anonymous:2022:CNb

[4289] Anonymous. Correction notice. Journal of Applied Statistics, 49(6):1612– 1614, June 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tahir:2022:NFG

[4290] M. H. Tahir, M. Adnan Hussain, and Gauss M. Cordeiro. A new flexible generalized family for constructing many families of distributions. Journal of Applied Statistics, 49(7):1615–1635, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2022:RDH

[4291] Yujie Zhao, Hao Yan, Sarah Holte, and Yajun Mei. Rapid detection of hotspots via tensor decomposition with applications to crime rate data. Journal of Applied Statistics, 49(7):1636–1662, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Piradl:2022:REM

[4292] Sajjad Piradl, Ali Shadrokh, and Masoud Yarmohammadi. A robust estimation method for the linear regression model parameters with correlated error terms and outliers. Journal of Applied Statistics, 49(7):1663–1676, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Milad:2022:TDM

[4293] Mohamed Milad and Gayla R. Olbricht. Testing differentially methylated regions through functional principal component analysis. Journal of Applied Statistics, 49(7):1677–1691, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Some:2022:BSA

[4294] Sobom M. Somé and Célestin C. Kokonendji. Bayesian selector of adaptive bandwidth for multivariate gamma kernel estimator on [0, ∞)^{d}. Journal of Applied Statistics, 49(7):1692–1713, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ronchi:2022:MNM

[4295] Fabrizio Ronchi, Solomon W. Harrar, and Luigi Salmaso. Multivariate nonparametric methods in two-way balanced designs: performances and limitations in small samples. Journal of Applied Statistics, 49(7):1714–1741, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Azimi:2022:AMC

[4296] Seyede Sedighe Azimi, Ehsan Bahrami Samani, and Mojtaba Ganjali. Analysis of mixed correlated overdispersed binomial and ordinal longitudinal responses: LogLindley-binomial and ordinal random effects model. Journal of Applied Statistics, 49(7):1742–1768, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Koley:2022:CSD

[4297] Tamalika Koley and Anup Dewanji. Current status data with two competing risks and missing failure types: a parametric approach. Journal of Applied Statistics, 49(7):1769–1783, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rehman:2022:CSH

[4298] H. Rehman, N. Chandra, Fatemeh Sadat Hosseini-Baharanchi, Ahmad Reza Baghestani, and Mohamad Amin Pourhoseingholi. Cause-specific hazard regression estimation for modified Weibull distribution under a class of non-informative priors. Journal of Applied Statistics, 49(7):1784–1801, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Momenyan:2022:CRM

[4299] Somayeh Momenyan, Farzane Ahmadi, and Jalal Poorolajal. Competing risks model for clustered data based on the subdistribution hazards with spatial random effects. Journal of Applied Statistics, 49(7):1802–1820, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2022:MMI

[4300] Cong Li, Haixiang Zhang, and Dehui Wang. Modelling and monitoring of INAR(1) process with geometrically inflated Poisson innovations. Journal of Applied Statistics, 49(7):1821–1847, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2022:CRS

[4301] Xiaolin Chen, Chenguang Li, Tao Zhang, and Zhenlong Gao. On correlation rank screening for ultra-high dimensional competing risks data. Journal of Applied Statistics, 49(7):1848– 1864, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deregnaucourt:2022:SCG

[4302] Thomas Deregnaucourt, Chafik Samir, Sebastian Kurtek, and Anne-Francoise Yao. Shape-constrained Gaussian process regression for surface reconstruction and multimodal, non-rigid image registration. Journal of Applied Statistics, 49(7):1865–1889, July 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cho:2022:HMP

[4303] Seonghun Cho, Johan Lim, and Woncheol Jang. How many people participated in candlelight protests? Counting the size of a dynamic crowd. Journal of Applied Statistics, 49(7):1890– 1899, July 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Corona:2022:DFS

[4304] Francisco Corona, Nelson Muriel, and Graciela González-Farías. Dynamic factor structure of team performances in Liga MX. Journal of Applied Statistics, 49(7):1900–1912, July 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ghanbari:2022:ESS

[4305] Sara Ghanbari, Abdolhamid Rezaei Roknabadi, and Mahdi Salehi. Estimation of stress-strength reliability for Marshall–Olkin distributions based on progressively Type-II censored samples. Journal of Applied Statistics, 49 (8):1913–1934, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Altun:2022:OPC

[4306] Emrah Altun, Gauss M. Cordeiro, and Miroslav M. Ristić. An oneparameter compounding discrete distribution. Journal of Applied Statistics, 49(8):1935–1956, August 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weiss:2022:EAE

[4307] Christian H. Weiß, Annika Homburg, Layth C. Alwan, Gabriel Frahm, and Rainer Göb. Efficient accounting for estimation uncertainty in coherent forecasting of count processes. Journal of Applied Statistics, 49(8):1957–1978, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2022:STP

[4308] Han Yu, Shanhe Jiang, and Hong Huang. Spatio-temporal parse network-based trajectory modeling on the dynamics of criminal justice system. Journal of Applied Statistics, 49 (8):1979–2000, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jabbari:2022:NIT

[4309] Yasaman Jabbari and Robert Cribbie. Negligible interaction test for continuous predictors. Journal of Applied Statistics, 49(8):2001–2015, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jadhav:2022:NLR

[4310] Nileshkumar H. Jadhav. A new linearized ridge Poisson estimator in the presence of multicollinearity. Journal of Applied Statistics, 49(8):2016–2034, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vasconcelos:2022:NHR

[4311] J. C. S. Vasconcelos, E. M. M. Ortega, J. S. Vasconcelos, G. M. Cordeiro, A. L. Vivan, and M. A. M. Biaggioni. A new heteroscedastic regression to analyze mass loss of wood in civil construction in Brazil. Journal of Applied Statistics, 49(8):2035–2051, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2022:BAC

[4312] Hyunsung Kim and Yaeji Lim. Bootstrap aggregated classification for sparse functional data. Journal of Applied Statistics, 49(8):2052–2063, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahto:2022:IPO

[4313] Amulya Kumar Mahto, Chandrakant Lodhi, Yogesh Mani Tripathi, and Liang Wang. Inference for partially observed competing risks model for Kumaraswamy distribution under generalized progressive hybrid censoring. Journal of Applied Statistics, 49(8): 2064–2092, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nandy:2022:RIS

[4314] Amarnath Nandy, Ayanendranath Basu, and Abhik Ghosh. Robust inference for skewed data in health sciences. Journal of Applied Statistics, 49 (8):2093–2123, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lukman:2022:MRT

[4315] Adewale F. Lukman, Benedicta Aladeitan, Kayode Ayinde, and Mohamed R. Abonazel. Modified ridge-type for the Poisson regression model: simulation and application. Journal of Applied Statistics, 49(8):2124–2136, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Correia:2022:CSD

[4316] H. E. Correia and A. Abebe. Capturing spatiotemporal dynamics of Alaskan groundfish catch using signed-rank estimation for varying coefficient models. Journal of Applied Statistics, 49 (8):2137–2156, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nascimento:2022:EEB

[4317] Marcus Gerardus L. Nascimento, Ralph S. Silva, Mario Jorge Mendonça, and Amaro Olimpio Pereira Jr. Estimating the efficiency of Brazilian electricity distribution utilities. Journal of Applied Statistics, 49(8):2157–2166, August 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhatti:2022:MBI

[4318] Fiaz Ahmad Bhatti and Munir Ahmad.

On the modified Burr IV model for hydrological events: development, properties, characterizations and applications. Journal of Applied Statistics, 49 (9):2167–2188, September 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2022:SST

[4319] Ryan Wu, Mihye Ahn, and Hojin Yang. Spike-and-slab type variable selection in the Cox proportional hazards model for high-dimensional features. Journal of Applied Statistics, 49 (9):2189–2207, September 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vittorietti:2022:IRM

[4320] Martina Vittorietti, Javier Hidalgo, Jilt Sietsma, Wei Li, and Geurt Jongbloed. Isotonic regression for metallic microstructure data: estimation and testing under order restrictions. Journal of Applied Statistics, 49(9): 2208–2227, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dutta:2022:JML

[4321] Srimanti Dutta, Geert Molenberghs, and Arindom Chakraborty. Joint modelling of longitudinal response and time-to-event data using conditional distributions: a Bayesian perspective. Journal of Applied Statistics, 49(9): 2228–2245, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luo:2022:SEC

[4322] Jingqin Luo, Feng Gao, Jingxia Liu, Guoqiao Wang, Ling Chen, Anne M. Fagan, Gregory S. Day, Jonathan Vöglein, Jasmeer P. Chhatwal, Chengjie Xiong, and the Dominantly Inherited Alzheimer Network (Dian) Steering Committee. Statistical estimation and comparison of group-specific bivariate correlation coefficients in family-type clustered studies. Journal of Applied Statistics, 49 (9):2246–2270, September 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ozdemir:2022:ELE

[4323] Senay Özdemir, Yesim Güney, Yetkin Tuaç, and Olcay Arslan. Empirical likelihood estimation for linear regression models with AR(p) error terms with numerical examples. Journal of Applied Statistics, 49(9):2271–2286, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Al-Kzzaz:2022:IPM

[4324] H. S. Al-Kzzaz and M. M. E. Abd El-Monsef. Inverse power Maxwell distribution: statistical properties, estimation and application. Journal of Applied Statistics, 49(9):2287–2306, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ekin:2022:ISD

[4325] Tahir Ekin and R. M. Musal. Integrated statistical and decision models for multi-stage health care audit sampling. Journal of Applied Statistics, 49 (9):2307–2325, September 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pokojovy:2022:UFI

[4326] Michael Pokojovy and J. Marcus Jobe.

Univariate fast initial response statistical process control with taut strings. Journal of Applied Statistics, 49(9): 2326–2348, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2022:ASB

[4327] Jingyu Liu, Walter W. Piegorsch, A. Grant Schissler, Rachel R. Mc-Caster, and Susan L. Cutter. Adjusting statistical benchmark risk analysis to account for non-spatial autocorrelation, with application to natural hazard risk assessment. Journal of Applied Statistics, 49(9):2349–2369, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jreich:2022:RBS

[4328] Rana Jreich, Christine Hatte, and Eric Parent. Review of Bayesian selection methods for categorical predictors using JAGS. Journal of Applied Statistics, 49(9):2370–2388, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gachau:2022:HMD

[4329] Susan Gachau, Edmund Njeru Njagi, Nelson Owuor, Paul Mwaniki, Matteo Quartagno, Rachel Sarguta, Mike English, and Philip Ayieko. Handling missing data in a composite outcome with partially observed components: simulation study based on clustered paediatric routine data. Journal of Applied Statistics, 49(9):2389–2402, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sozen:2022:ITD

[4330] Çaglar Sözen and Yüksel Öner. The investigation of temperature data in Turkey’s Black Sea Region using functional data analysis. Journal of Applied Statistics, 49(9):2403–2415, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

CavenaguedeSouza:2022:LNZ

[4331] Hayala Cristina Cavenague de Souza, Francisco Louzada, Mauro Ribeiro de Oliveira, Bukola Fawole, Adesina Akintan, Lawal Oyeneyin, Wilfred Sanni, and Gleici da Silva Castro Perdoná. The log-normal zero-inflated cure regression model for labor time in an African obstetric population. Journal of Applied Statistics, 49(9):

2416–2429, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Suarez-Sierra:2022:ANH

[4332] Biviana M. Súarez-Sierra, Eliane R.

Rodrigues, and Guadalupe Tzintzun. An application of a non-homogeneous Poisson model to study PM_{2.5} exceedances in Mexico City and Bogota. Journal of Applied Statistics, 49(9): 2430–2445, September 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rahman:2022:SES

[4333] Rezoanoor Rahman and M. Iftakhar Alam. Stopping for efficacy in singlearm phase II clinical trials. Journal of Applied Statistics, 49(10):2447–2466, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Eliwa:2022:OPD

[4334] M. S. Eliwa and M. El-Morshedy. A one-parameter discrete distribution for over-dispersed data: statistical and reliability properties with applications. Journal of Applied Statistics, 49(10): 2467–2487, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jin:2022:MBG

[4335] Zhongnan Jin, Lu Lu, Khaled Bedair, and Yili Hong. Modeling bivariate geyser eruption system with covariate-adjusted recurrent event process. Journal of Applied Statistics, 49 (10):2488–2509, October 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sosa:2022:TVC

[4336] Juan Sosa and Lina Buitrago. Timevarying coefficient model estimation through radial basis functions. Journal of Applied Statistics, 49(10):2510–2534, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shan:2022:CCI

[4337] Guogen Shan. Conservative confidence intervals for the intraclass correlation coefficient for clustered binary data. Journal of Applied Statistics, 49(10): 2535–2549, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rashid:2022:DOH

[4338] Abdullah Mohammed Rashid, Habshah Midi, Waleed Dhhan, and Jayanthi Arasan. Detection of outliers in high-dimensional data using ν-support vector regression. Journal of Applied Statistics, 49(10):2550–2569, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Diop:2022:CAM

[4339] S. Arona Diop, Thierry Duchesne, Steven G. Cumming, Awa Diop, and Denis Talbot. Confounding adjustment methods for multi-level treatment comparisons under lack of positivity and unknown model specification. Journal of Applied Statistics, 49 (10):2570–2592, October 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ding:2022:CSC

[4340] Saisai Ding, Hongyan Fang, Xiang Dong, and Wenzhi Yang. The CUSUM statistics of change-point models based on dependent sequences. Journal of Applied Statistics, 49(10):2593–2611, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Salmeron:2022:BBN

[4341] Diego Salmerón. Bayesian beta nonlinear models with constrained parameters to describe ruminal degradation kinetics. Journal of Applied Statistics, 49(10):2612–2628, October 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2022:NMD

[4342] Chi Zhang, Guo-Liang Tian, Kam Chuen Yuen, Pengyi Liu, and Man-Lai Tang. A new multivariate t distribution with variant tail weights and its application in robust regression analysis. Journal of Applied Statistics, 49(10):2629–2656, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2022:BDL

[4343] Yingzi Li, Huinan Liu, and Nairanjana Dasgupta. Binary Dynamic Logit for Correlated Ordinal: estimation, application and simulation. Journal of Applied Statistics, 49(10):2657–2673, October 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Arboretti:2022:MLD

[4344] Rosa Arboretti, Riccardo Ceccato, Luca Pegoraro, Luigi Salmaso, Chris Housmekerides, Luca Spadoni, Elisabetta Pierangelo, Sara Quaggia, Catherine Tveit, and Sebastiano Vianello. Machine learning and design of experiments with an application to product innovation in the chemical industry. Journal of Applied Statistics, 49(10):2674–2699, October 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kazianka:2022:ADP

[4345] Hannes Kazianka, Anna Morgenbesser, and Thomas Nowak. Assessing the discriminatory power of loss given default models. Journal of Applied Statistics, 49(10):2700–2716, October 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2022:GPI

[4346] Chenhui Zhang, Dehui Wang, Kai Yang, Han Li, and Xiaohong Wang. Generalized Poisson integer-valued autoregressive processes with structural changes. Journal of Applied Statistics, 49(11):2717–2739, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deng:2022:MZI

[4347] Dianliang Deng, Yiguang Sun, and Guo-Liang Tian. A multivariate zeroinflated binomial model for the analysis of correlated proportional data. Journal of Applied Statistics, 49(11): 2740–2766, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2022:IPE

[4348] Dongyu Li and Lei Wang. Improved kth power expectile regression with nonignorable dropouts. Journal of Applied Statistics, 49(11):2767–2788, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2022:STS

[4349] Huiping Xu, Xiaochun Li, and Shaun Grannis. A simple two-step procedure using the Fellegi–Sunter model for frequency-based record linkage. Journal of Applied Statistics, 49(11):2789– 2804, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vigas:2022:NNT

[4350] Valdemiro P. Vigas, Edwin M. M. Ortega, Gauss M. Cordeiro, Adriano K. Suzuki, and Giovana O. Silva. The new Neyman type A generalized odd log-logistic-G-family with cure fraction. Journal of Applied Statistics, 49(11): 2805–2824, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maswadah:2022:IML

[4351] M. Maswadah. Improved maximum likelihood estimation of the shapescale family based on the generalized progressive hybrid censoring scheme. Journal of Applied Statistics, 49(11): 2825–2844, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2022:SZI

[4352] Chin-Shang Li and Minggen Lu. Semiparametric zero-inflated Bernoulli regression with applications. Journal of Applied Statistics, 49(11):2845–2869, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Soleimani:2022:ENN

[4353] Paria Soleimani and Shervin Asadzadeh.

Effect of non-normality on the monitoring of simple linear profiles in twostage processes: a remedial measure for gamma-distributed responses. Journal of Applied Statistics, 49(11):2870–2890, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dey:2022:INH

[4354] Sanku Dey, Liang Wang, and Mazen Nassar. Inference on Nadarajah– Haghighi distribution with constant stress partially accelerated life tests under progressive type-II censoring. Journal of Applied Statistics, 49(11): 2891–2912, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McVittie:2022:RSA

[4355] J. H. McVittie and V. Addona. A risk set adjustment for proportional hazards modeling of combined cohort data. Journal of Applied Statistics, 49(11): 2913–2927, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

El-Morshedy:2022:DAO

[4356] M. El-Morshedy, M. S. Eliwa, and Abhishek Tyagi. A discrete analogue of odd Weibull-G family of distributions: properties, classical and Bayesian estimation with applications to count data. Journal of Applied Statistics, 49(11): 2928–2952, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lin:2022:SUH

[4357] Ting Hsiang Lin and Min-Hsiao Tsai. Solving unobserved heterogeneity with latent class inflated Poisson regression model. Journal of Applied Statistics, 49 (11):2953–2963, November 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sathe:2022:EPS

[4358] Aastha M. Sathe and N. S. Upadhye. Estimation of the parameters of symmetric stable ARMA and ARMA-GARCH models. Journal of Applied Statistics, 49(11):2964–2980, November 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahmad:2022:BNB

[4359] Hanan Haj Ahmad, Mukhtar M. Salah, M. S. Eliwa, Ziyad Ali Alhussain, Ehab M. Almetwally, and Essam A. Ahmed. Bayesian and non-Bayesian inference under adaptive type-II progressive censored sample with exponentiated power Lindley distribution. Journal of Applied Statistics, 49(12): 2981–3001, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2022:DDP

[4360] Qiushi Yu. Dynamic Dirichlet process mixture model for identifying voting coalitions in the United Nations General Assembly human rights roll call votes. Journal of Applied Statistics, 49 (12):3002–3021, December 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Naderi:2022:SIS

[4361] Mehrdad Naderi, Elham Mirfarah, Matthew Bernhardt, and Ding-Geng Chen. Semiparametric inference for the scale-mixture of normal partial linear regression model with censored data.

Journal of Applied Statistics, 49(12): 3022–3043, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vila:2022:BWD

[4362] Roberto Vila and Mehmet Niyazi Çankaya. A bimodal Weibull distribution: properties and inference. Journal of Applied Statistics, 49(12): 3044–3062, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2022:MPJ

[4363] Yangxin Huang, Nian-Sheng Tang, and Jiaqing Chen. Multivariate piecewise joint models with random changepoints for skewed-longitudinal and survival data. Journal of Applied Statistics, 49(12):3063–3089, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Adeoti:2022:HWM

[4364] Olatunde Adebayo Adeoti, Jean-Claude Malela-Majika, Sandile Charles Shongwe, and Muhammad Aslam. A homogeneously weighted moving average control chart for Conway–Maxwell Poisson distribution. Journal of Applied Statistics, 49(12):3090–3119, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Du:2022:SIA

[4365] Yuge Du and Wenhao Gui. Statistical inference of adaptive type II progressive hybrid censored data with dependent competing risks under bivariate exponential distribution. Journal of Applied Statistics, 49(12):3120–3140, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yilmaz:2022:HRM

[4366] Ayfer Ezgi Yilmaz. How reliable are the multiple comparison methods for odds ratio? Journal of Applied Statistics, 49 (12):3141–3163, December 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ahmadi:2022:TCM

[4367] A. Ahmadi, T. Baghfalaki, M. Ganjali, A. Kabir, and A. Pazouki. A transition copula model for analyzing multivariate longitudinal data with missing responses. Journal of Applied Statistics, 49(12):3164–3177, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deSouza:2022:BAZ

[4368] Hayala Cristina Cavenague de Souza, Francisco Louzada, Pedro Luiz Ramos, Mauro Ribeiro de Oliveira Júnior, and Gleici da Silva Castro Perdoná. A Bayesian approach for the zero-inflated cure model: an application in a Brazilian invasive cervical cancer database. Journal of Applied Statistics, 49(12): 3178–3194, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Giese:2022:MNP

[4369] Gill Giese, Dayna P. Saldaña Zepeda, Jacquelin Beacham, and Ciro Velasco Cruz. Modeling nematode population dynamics using a multivariate Poisson model with spike and slab variable selection. Journal of Applied Statistics, 49(12):3195–3214, December 2022.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2022:BSA

[4370] Lijuan Feng and Murat Munkin. Bayesian semiparametric analysis on the relationship between BMI and income for rural and urban workers in China. Journal of Applied Statistics, 49 (12):3215–3235, December 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Katona:2022:VRB

[4371] A. I. Katona. Validation of riskbased quality control techniques: a case study from the automotive industry. Journal of Applied Statistics, 49(12): 3236–3255, December 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2022:ILR

[4372] Lili Zhang, Trent Geisler, Herman Ray, and Ying Xie. Improving logistic regression on the imbalanced data by a novel penalized log-likelihood function. Journal of Applied Statistics, 49(13):3257–3277, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kreutzmann:2022:FHM

[4373] Ann-Kristin Kreutzmann, Philipp Marek, Marina Runge, Nicola Salvati, and Timo Schmid. The Fay–Herriot model for multiply imputed data with an application to regional wealth estimation in Germany. Journal of Applied Statistics, 49(13):3278–3299, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:2022:EDG

[4374] Samuel I. Watson. Efficient design of geographically-defined clusters with spatial autocorrelation. Journal of Applied Statistics, 49(13):3300–3318, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bouaziz:2022:RMI

[4375] Olivier Bouaziz, Eva Lauridsen, and Grégory Nuel. Regression modelling of interval censored data based on the adaptive ridge procedure. Journal of Applied Statistics, 49(13):3319–3343, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bapat:2022:MPE

[4376] Sudeep R. Bapat and Yan Zhuang.

Maximum precision estimation for a step-stress model using two-stage methodologies. Journal of Applied Statistics, 49(13):3344–3360, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2022:JMA

[4377] Jung Yeon Park, Melanie M. Wall, Irini Moustaki, and Arnold H. Grossman. A joint modeling approach for longitudinal outcomes and non-ignorable dropout under population heterogeneity in mental health studies. Journal of Applied Statistics, 49(13):3361–3376, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rahmati:2022:CPF

[4378] Maryam Rahmati, Parisa Rezanejad Asl, Javad Mikaeli, Hojjat Zeraati, and Aliakbar Rasekhi. Compound Poisson frailty model with a gamma process prior for the baseline hazard: accounting for a cured fraction. Journal of Applied Statistics, 49(13):3377–3391, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rey:2022:ECB

[4379] Alexey Rey, Olga Rozanova, and Sergey Zhuk. Estimating correlations between vaccine clinical trial outcomes. Journal of Applied Statistics, 49(13):3392–3413, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qiu:2022:CIA

[4380] Shi-Fang Qiu and Ji-Ran Tao. Confidence intervals for assessing equivalence of two treatments with combined unilateral and bilateral data. Journal of Applied Statistics, 49(13):3414–3435, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nascimento:2022:BVS

[4381] Marcus G. L. Nascimento and Kelly C. M. Gonçalves. Bayesian variable selection in quantile regression with random effects: an application to Municipal Human Development Index. Journal of Applied Statistics, 49(13):3436– 3450, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shakhatreh:2022:ILP

[4382] Mohammed K. Shakhatreh, Sanku Dey, and Devendra Kumar. Inverse Lindley power series distributions: a new compounding family and regression model with censored data. Journal of Applied Statistics, 49(13):3451–3476, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2022:OHT

[4383] Seongoh Park and Johan Lim. An overview of heavy-tail extensions of multivariate Gaussian distribution and their relations. Journal of Applied Statistics, 49(13):3477–3494, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chorozoglou:2022:ICS

[4384] D. Chorozoglou, D. Kugiumtzis, and E. Papadimitriou. Investigation of the correlation of successive earthquakes preceding main shocks in the Greek territory. Journal of Applied Statistics, 49(13):3495–3512, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zubair:2022:SMI

[4385] Seema Zubair and Sanjoy K. Sinha.

Semiparametric methods for incomplete longitudinal count data with an application to health and retirement study. Journal of Applied Statistics, 49(14):3513–3535, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Williams:2022:WAV

[4386] Jeffrey Williams, Raymond R. Hill, Joseph J. Pignatiello Jr., and Eric Chicken. Wavelet analysis of variance box plot. Journal of Applied Statistics, 49(14):3536–3563, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Barry:2022:NGM

[4387] Amadou Barry, Karim Oualkacha, and Arthur Charpentier. A new GEE method to account for heteroscedasticity using asymmetric least-square regressions. Journal of Applied Statistics, 49(14):3564–3590, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Biswas:2022:BCE

[4388] Aniket Biswas, Gaurangadeb Chattopadhyay, and Aditya Chatterjee. Bias-corrected estimators for proportion of true null hypotheses: application of adaptive FDR-controlling in segmented failure data. Journal of Applied Statistics, 49(14):3591–3613, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martinez-Florez:2022:FBD

[4389] Guillermo Martínez-Flórez, Eliseo Martínez, Roger Tovar-Falón, and Héctor W. Gómez. A family of bimodal distributions generated by distributions with positive support. Journal of Applied Statistics, 49(14):3614– 3637, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zafari:2022:MDF

[4390] Babak Zafari, Tahir Ekin, and Fabrizio Ruggeri. Multicriteria decision frontiers for prescription anomaly detection over time. Journal of Applied Statistics, 49(14):3638–3658, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Opheim:2022:DML

[4391] Timothy Opheim and Anuradha Roy. Doubly multivariate linear models with block exchangeable distributed errors and site-dependent covariates. Journal of Applied Statistics, 49(14):3659–3676, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2022:REV

[4392] Zhuoran Yang, Liya Fu, You-Gan Wang, Zhixiong Dong, and Yunlu Jiang. A robust and efficient variable selection method for linear regression. Journal of Applied Statistics, 49(14):3677–3692, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ji:2022:SEF

[4393] Yonggang Ji and Haifang Shi. Shrinkage estimation of fixed and random effects in linear quantile mixed models. Journal of Applied Statistics, 49(14): 3693–3716, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sheng:2022:LRT

[4394] Zhen Sheng, Yukun Liu, Pengfei Li, and Jing Qin. Likelihood ratio test for genetic association study with casecontrol data under probit model. Journal of Applied Statistics, 49(14):3717– 3731, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Panikian:2022:TSM

[4395] Garo Panikian, Gabby Colmenares Reverol, Jayne Rhodes, Emma McLarnon, Sarah Keast, and Kokouvi Gamado. Time series modelling methods to forecast the volume of self-assessment tax returns in the UK. Journal of Applied Statistics, 49(14):3732–3749, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Huang:2022:CMA

[4396] Jiasheng Huang, Yehua Li, Angelique G. Brellenthin, Duck chul Lee, Xuemei Sui, and Steven N. Blair. Causal mediation analysis between resistance exercise and reduced risk of cardiovascular disease based on the Aerobics Center Longitudinal Study. Journal of Applied Statistics, 49(14): 3750–3767, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2022:EEB

[4397] Junyi Lu and Sebastian Meyer. An endemic-epidemic beta model for time series of infectious disease proportions. Journal of Applied Statistics, 49(15): 3769–3783, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deFreitas:2022:NAM

[4398] João Victor B. de Freitas, Juvêncio S. Nobre, Marcelo Bourguignon, and Manoel Santos-Neto. A new approach to modeling positive random variables with repeated measures. Journal of Applied Statistics, 49(15):3784–3803, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhuang:2022:SIR

[4399] Weiwei Zhuang, Yadong Li, and Guoxin Qiu. Statistical inference for a relaxation index of stochastic dominance under density ratio model. Journal of Applied Statistics, 49(15):3804– 3822, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

He:2022:ECM

[4400] Xiaohong He and Lei Wang. Ensemble and calibration multiply robust estimation for quantile treatment effect. Journal of Applied Statistics, 49(15):3823– 3845, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2022:GBA

[4401] Xin Yang, Alan D. Hutson, and Dongliang Wang. A generalized BLUE approach for combining location and scale information in a meta-analysis. Journal of Applied Statistics, 49(15): 3846–3867, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sliwka:2022:MSC

[4402] Piotr Sliwka. Markov (set) chains application to predict mortality rates using extended Milevsky–Promislov generalized mortality models. Journal of Applied Statistics, 49(15):3868–3888, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Poythress:2022:DWS

[4403] J. C. Poythress, Cheolwoo Park, and Jeongyoun Ahn. Dimension-wise sparse low-rank approximation of a matrix with application to variable selection in high-dimensional integrative analyzes of association. Journal of Applied Statistics, 49(15):3889–3907, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DAmico:2022:CSD

[4404] Guglielmo D’Amico and Riccardo De Blasis. Confidence sets for dynamic poverty indexes. Journal of Applied Statistics, 49(15):3908–3927, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xie:2022:OSE

[4405] FuPeng Xie, Philippe Castagliola, Yu-Long Qiao, XueLong Hu, and Jin-Sheng Sun. A one-sided exponentially weighted moving average control chart for time between events. Journal of Applied Statistics, 49(15):3928–3957, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2022:IME

[4406] Qingzhao Yu, Lu Zhang, Xiaocheng Wu, and Bin Li. Inference on moderation effect with third-variable effect analysis — application to explore the trend of racial disparity in oncotype DX test for breast cancer treatment. Journal of Applied Statistics, 49(15):3958–3975, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2022:REE

[4407] Qiang Zhao, Liang Chen, and Jingjing Wu. Robust and efficient estimation of GARCH models based on Hellinger distance. Journal of Applied Statistics, 49(15):3976–4002, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2022:ESL

[4408] Yi-Ching Lee and Junfeng Shang. Estimation and selection in linear mixed models with missing data under compound symmetric structure. Journal of Applied Statistics, 49(15):4003–4027, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shen:2022:SMN

[4409] Xun Shen and Pongsathorn Raksincharoensak. Statistical models of nearaccident event and pedestrian behavior at non-signalized intersections. Journal of Applied Statistics, 49(15):4028–4048, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sinclair:2022:EMA

[4410] David G. Sinclair and Giles Hooker.

An expectation maximization algorithm for high-dimensional model selection for the Ising model with misclassified states*. Journal of Applied Statistics, 49(16):4049–4068, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ersel:2022:AKA

[4411] Derya Ersel and Yasemin Kayhan Atilgan. An approach for knowledge acquisition from a survey data by conducting Bayesian network modeling, adopting the robust coplot method. Journal of Applied Statistics, 49(16):4069–4096, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saha:2022:PIP

[4412] Mahendra Saha, Sanku Dey, and Saralees Nadarajah. Parametric inference of the process capability index for exponentiated exponential distribution. Journal of Applied Statistics, 49(16):4097–4121, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fang:2022:OMH

[4413] Zhiwen Fang, Wendong Li, Xin Liu, Xiaolong Pu, and Dongdong Xiang. Online monitoring of high-dimensional binary data streams with application to extreme weather surveillance. Journal of Applied Statistics, 49(16):4122–4136, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prataviera:2022:NRM

[4414] F. Prataviera, G. M. Cordeiro, E. M. M. Ortega, E. M. Hashimoto, and V. G. Cancho. A new regression model for rates and proportions data with applications. Journal of Applied Statistics, 49(16):4137–4161, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beretta:2022:CRP

[4415] A. Beretta, C. Heuchenne, and M. Restaino. Competing risks proportional-hazards cure model and generalized extreme value regression: an application to bank failures and acquisitions in the United States. Journal of Applied Statistics, 49(16):4162–4180, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qasim:2022:NCE

[4416] Muhammad Qasim, Kristofer Månsson, Pär Sjölander, and B. M. Golam Kibria. A new class of efficient and debiased two-step shrinkage estimators: method and application. Journal of Applied Statistics, 49(16):4181–4205, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Reyes:2022:MSL

[4417] Jimmy Reyes, Jaime Arrué, Osvaldo Venegas, and Héctor W. Gómez. The modified slash Lindley–Weibull distribution with applications to nutrition data. Journal of Applied Statistics, 49(16):4206–4224, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cunha:2022:LSQ

[4418] Danúbia R. Cunha, Jose Angelo Divino, and Helton Saulo. On a logsymmetric quantile tobit model applied to female labor supply data. Journal of Applied Statistics, 49(16):4225–4253, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nusrat:2022:DSN

[4419] Nowrin Nusrat and M. S. Rahman.

Dealing with separation or near-to-separation in the model for multinomial response with application to childhood health seeking behavior data from a complex survey. Journal of Applied Statistics, 49(16):4254–4277, 2022. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2022:RRL

[4420] Jinjuan Wang, Yunpeng Zhao, Larry L.

Tang, Claudius Mueller, and Qizhai Li. A resample-replace lasso procedure for combining high-dimensional markers with limit of detection. Journal of Applied Statistics, 49(16):4278–4293, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jayasinghe:2022:RMU

[4421] J. M. Thilini Jayasinghe, Leif Ellingson, and Chalani Prematilake. Regression models using the LINEX loss to predict lower bounds for the number of points for approximating planar contour shapes. Journal of Applied Statistics, 49(16):4294–4313, 2022. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2023:TGM

[4422] Vera Liu and Stephen Walker. Testing for genetic mutation of seasonal influenza virus. Journal of Applied Statistics, 50(1):1–18, January 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chatterjee:2023:MPM

[4423] Kashinath Chatterjee, Christos Koukouvinos, and Angeliki Lappa. Monitoring process mean and dispersion with one double generally weighted moving average control chart. Journal of Applied Statistics, 50(1):19–42, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dong:2023:JMM

[4424] Jianghu (James) Dong, Haolun Shi, Liangliang Wang, Ying Zhang, and Jiguo Cao. Jointly modelling multiple transplant outcomes by a competing risk model via functional principal component analysis. Journal of Applied Statistics, 50(1):43–59, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2023:ESA

[4425] Honggang Zhang, Jingyong Su, Linlin Tang, and Anuj Srivastava. Elastic statistical analysis of interval-valued time series. Journal of Applied Statistics, 50(1):60–85, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Agnihotri:2023:IMI

[4426] Shubham Agnihotri, Sanjay Kumar Singh, and Umesh Singh. Inferences for multiple interval type-I censoring scheme. Journal of Applied Statistics, 50(1):86–105, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brocker:2023:TRF

[4427] J. Bröcker. Testing the reliability of forecasting systems. Journal of Applied Statistics, 50(1):106–130, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Korkmaz:2023:NAQ

[4428] Mustafa Ç. Korkmaz, Christophe Chesneau, and Zehra Sedef Korkmaz. A new alternative quantile regression model for the bounded response with educational measurements applications of OECD countries. Journal of Applied Statistics, 50(1):131–154, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2023:MBD

[4429] Chun-Chao Wang, Yi-Ting Hwang, Chung-Chuan Chou, and Hui-Ling Lee. Misspecification of a binary dependent variable in the logistic model controlling for the repeated longitudinal measures. Journal of Applied Statistics, 50 (1):155–169, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Perdikis:2023:ESC

[4430] Theodoros Perdikis, Stelios Psarakis, Philippe Castagliola, Athanasios C. Rakitzis, and Petros E. Maravelakis.

The EWMA sign chart revisited: performance and alternatives without and with ties. Journal of Applied Statistics, 50(1):170–194, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tanis:2023:CRT

[4431] Caner Tanis and Bugra Saraçoglu. Cubic rank transmuted generalized Gompertz distribution: properties and applications. Journal of Applied Statistics, 50(1):195–213, January 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bi:2023:SPB

[4432] Ran Bi and Peng Liu. A semiparametric Bayesian approach for detection of gene expression heterosis with RNA-seq data. Journal of Applied Statistics, 50(1):214–230, January 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Boone:2023:MSM

[4433] Edward L. Boone, Abdel-Salam G. Abdel-Salam, Indranil Sahoo, Ryad Ghanam, Xi Chen, and Aiman Hanif. Monitoring SEIRD model parameters using MEWMA for the COVID-19 pandemic with application to the state of Qatar. Journal of Applied Statistics, 50 (2):231–246, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ding:2023:NGV

[4434] Hao Ding, Yan Zhang, and Yuehua Wu. A novel group VIF regression for group variable selection with application to multiple change-point detection. Journal of Applied Statistics, 50 (2):247–263, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Emura:2023:STB

[4435] Takeshi Emura, Wei-Chern Hsu, and Wen-Chi Chou. A survival tree based on stabilized score tests for highdimensional covariates. Journal of Applied Statistics, 50(2):264–290, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2023:CSE

[4436] Xinyi Wang and Zhenghui Feng. Component selection for exponential power mixture models. Journal of Applied Statistics, 50(2):291–314, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mulder:2023:BFT

[4437] Joris Mulder and John P. T. M. Gelissen. Bayes factor testing of equality and order constraints on measures of association in social research. Journal of Applied Statistics, 50(2):315–351, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2023:FOB

[4438] Zezhun Chen, Angelos Dassios, and George Tzougas. A first-order binomial-mixed Poisson integer-valued autoregressive model with serially dependent innovations. Journal of Applied Statistics, 50(2):352–369, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2023:IOO

[4439] Tao Wang, Xiaona Yang, Yunfei Guo, and Zhonghua Li. Identification of outlying observations for largedimensional data. Journal of Applied Statistics, 50(2):370–386, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lu:2023:BAV

[4440] Zihang Lu and Wendy Lou. Bayesian approaches to variable selection in mixture models with application to disease clustering. Journal of Applied Statistics, 50(2):387–407, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Altinisik:2023:AOZ

[4441] Yasin Altinisik. Addressing overdispersion and zero-inflation for clustered count data via new multilevel heterogeneous hurdle models. Journal of Applied Statistics, 50(2):408–433, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bettinger:2023:BRL

[4442] Robert A. Bettinger. Bimodality in reentry latitude predictions for spacecraft in prograde orbits. Journal of Applied Statistics, 50(2):434–450, February 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2023:ESI

[4443] Yichuan Zhao, Chi-Hua Chen, Feng Feng, and Dragan Pamucar. Editorial to the special issue: Statistical Approaches for Big Data and Machine Learning. Journal of Applied Statistics, 50(3):451–455, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2023:TSB

[4444] Liang Zhang, Tianming Zhu, and Jin-Ting Zhang. Two-sample Behrens– Fisher problems for high-dimensional data: a normal reference scaleinvariant test. Journal of Applied Statistics, 50(3):456–476, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pustokhin:2023:EDR

[4445] Denis A. Pustokhin, Irina V. Pustokhina, Phuoc Nguyen Dinh, Son Van Phan, Gia Nhu Nguyen, Gyanendra Prasad Joshi, and Shankar K. An effective deep residual network based class attention layer with bidirectional LSTM for diagnosis and classification of COVID-19. Journal of Applied Statistics, 50(3):477–494, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yuan:2023:PTS

[4446] Mingao Yuan and Qian Wen. A practical two-sample test for weighted random graphs. Journal of Applied Statistics, 50(3):495–511, March 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2023:MES

[4447] Mingtao Zhao, Xiaoli Xu, Yanling Zhu, Kongsheng Zhang, and Yan Zhou. Model estimation and selection for partial linear varying coefficient EV models with longitudinal data. Journal of Applied Statistics, 50(3):512–534, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Weber:2023:RRW

[4448] Matthias Weber, Jonas Striaukas, Martin Schumacher, and Harald Binder. Regularized regression when covariates are linked on a network: the 3CoSE algorithm. Journal of Applied Statistics, 50(3):535–554, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiaoyue:2023:DMS

[4449] Xie Xiaoyue, Jian Shi, and Kai Song. A distributed multiple sample testing for massive data. Journal of Applied Statistics, 50(3):555–573, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Thielmann:2023:UDC

[4450] Anton Thielmann, Christoph Weisser, Astrid Krenz, and Benjamin Säfken. Unsupervised document classification integrating web scraping, one-class SVM and LDA topic modelling. Journal of Applied Statistics, 50(3):574– 591, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fan:2023:OML

[4451] Fang Fan, Shu-Chuan Chu, Jeng-Shyang Pan, Chuang Lin, and Huiqi Zhao. An optimized machine learning technology scheme and its application in fault detection in wireless sensor networks. Journal of Applied Statistics, 50(3):592–609, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2023:MMC

[4452] Mingyue Zhang Wu, Jinzhu Luo, Xing Fang, Maochao Xu, and Peng Zhao. Modeling multivariate cyber risks: deep learning dating extreme value theory. Journal of Applied Statistics, 50(3):610–630, March 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2023:AML

[4453] Wei Zhang, Colin O. Wu, Xiaoyang Ma, Xin Tian, and Qizhai Li. Analysis of multivariate longitudinal data using dynamic lasso-regularized copula models with application to large pediatric cardiovascular studies. Journal of Applied Statistics, 50(3):631–658, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhi:2023:NID

[4454] Xiaobin Zhi, Tongjun Yu, Longtao Bi, and Yalan Li. Noise-insensitive discriminative subspace fuzzy clustering. Journal of Applied Statistics, 50 (3):659–674, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kang:2023:CHV

[4455] Ilsuk Kang, Cheolwoo Park, Young Joo Yoon, Changyi Park, Soon-Sun Kwon, and Hosik Choi. Classification of histogram-valued data with support histogram machines. Journal of Applied Statistics, 50(3):675–690, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2023:ING

[4456] Xiaotong Liu, Guoliang Tian, and Zhenqiu Liu. Identification of novel genes for triple-negative breast cancer with semiparametric gene-based analysis. Journal of Applied Statistics, 50 (3):691–702, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhi:2023:GNR

[4457] X. Zhi, J. Liu, S. Wu, and C. Niu. A generalized l_{2,p} -norm regression based feature selection algorithm. Journal of Applied Statistics, 50(3):703–723, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cheng:2023:AEB

[4458] Yujia Cheng, Yang Li, Matthew Lee Smith, Changwei Li, and Ye Shen. Analyzing evidence-based falls prevention data with significant missing information using variable selection after multiple imputation. Journal of Applied Statistics, 50(3):724–743, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhou:2023:SAU

[4459] Mi Zhou and Weixin Yao. Sensitivity analysis of unmeasured confounding in causal inference based on exponential tilting and super learner. Journal of Applied Statistics, 50(3):744–760, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dagdoug:2023:MAE

[4460] Mehdi Dagdoug, Camelia Goga, and David Haziza. Model-assisted estimation in high-dimensional settings for survey data. Journal of Applied Statistics, 50(3):761–785, March 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2023:HHD

[4461] Sixia Chen and Chao Xu. Handling high-dimensional data with missing values by modern machine learning techniques. Journal of Applied Statistics, 50(3):786–804, March 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jin:2023:MRS

[4462] Jin Jin, Lin Zhang, Ethan Leng, Gregory J. Metzger, and Joseph S. Koopmeiners. Multi-resolution super learner for voxel-wise classification of prostate cancer using multi-parametric MRI. Journal of Applied Statistics, 50(3):805–826, March 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Suzuki:2023:DFP

[4463] Akira Suzuki, Hidetoshi Murakami, and Amitava Mukherjee. Distributionfree Phase-I scheme for location, scale and skewness shifts with an application in monitoring customers’ waiting time. Journal of Applied Statistics, 50(4):827–847, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Paul:2023:MPT

[4464] Erina Paul, Ram C. Tiwari, Shrabanti Chowdhury, and Samiran Ghosh. A more powerful test for three-arm noninferiority via risk difference: Frequentist and Bayesian approaches. Journal of Applied Statistics, 50(4):848–870, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

daPaz:2023:FMM

[4465] Rosineide da Paz, Jorge Luis Bazán, Victor Hugo Lachos, and Dipak Dey. A finite mixture mixed proportion regression model for classification problems in longitudinal voting data. Journal of Applied Statistics, 50(4):871–888, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Korkmaz:2023:ULL

[4466] Mustafa Ç. Korkmaz and Zehra Sedef Korkmaz. The unit log-log distribution: a new unit distribution with alternative quantile regression modeling and educational measurements applications. Journal of Applied Statistics, 50(4):889–908, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Venkatasubramaniam:2023:FDC

[4467] A. Venkatasubramaniam, L. Evers, P. Thakuriah, and K. Ampountolas. Functional distributional clustering using spatio-temporal data. Journal of Applied Statistics, 50(4):909–926, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2023:TQT

[4468] Guanfu Liu and Zongliang Hu. Testing quantitative trait locus effects in genetic backcross studies with double recombination occurring. Journal of Applied Statistics, 50(4):927–944, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alsarray:2023:MWS

[4469] Rusul Mohsin Moharib Alsarray, Jaber Kazempoor, and Adel Ahmadi Nadi.

Monitoring the Weibull shape parameter under progressive censoring in presence of independent competing risks. Journal of Applied Statistics, 50(4):945–962, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2023:CFJ

[4470] Qianyi Li, Jianbo Li, Yongran Cheng, and Riquan Zhang. Curve fitting and jump detection on nonparametric regression with missing data. Journal of Applied Statistics, 50(4):963–983, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mohtashami-Borzadaran:2023:EPD

[4471] V. Mohtashami-Borzadaran, M. Amini, and J. Ahmadi. Estimating the parameters of a dependent model and applying it to environmental data set. Journal of Applied Statistics, 50(4):984– 1016, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Atif:2023:AMT

[4472] Muhammad Atif, Muhammad Shafiq, and Friedrich Leisch. Applications of monitoring and tracing the evolution of clustering solutions in dynamic datasets. Journal of Applied Statistics, 50(4):1017–1035, April 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aoki:2023:UCM

[4473] Reiko Aoki, Dorival Leão, Juan P. Mamani Bustamante, and Filidor Vilca. Ultrastructural calibration model for proficiency testing. Journal of Applied Statistics, 50(5):1037–1059, May 2023.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mou:2023:HTR

[4474] Ke-Yi Mou, Chang-Xing Ma, and Zhi-Ming Li. Homogeneity test of relative risk ratios for stratified bilateral data under different algorithms. Journal of Applied Statistics, 50(5):1060–1077, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2023:PGF

[4475] Shuhua Chang, Deli Li, and Yongcheng Qi. Pearson’s goodness-of-fit tests for sparse distributions. Journal of Applied Statistics, 50(5):1078–1093, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ren:2023:GFM

[4476] Pengcheng Ren, Guanfu Liu, and Xiaolong Pu. Generalized fiducial methods for testing the homogeneity of a three-sample problem with a mixture structure. Journal of Applied Statistics, 50(5):1094–1114, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kulasekera:2023:PTS

[4477] K. B. Kulasekera, Sudaraka Tholkage, and Maiying Kong. Personalized treatment selection using observational data. Journal of Applied Statistics, 50(5):1115–1127, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Flagg:2023:INL

[4478] Kenneth Flagg and Andrew Hoegh. The integrated nested Laplace approximation applied to spatial log-Gaussian Cox process models. Journal of Applied Statistics, 50(5):1128–1151, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kharazmi:2023:LMD

[4479] O. Kharazmi, G. G. Hamedani, and G. M. Cordeiro. Log-mean distribution: applications to medical data, survival regression, Bayesian and non-Bayesian discussion with MCMC algorithm. Journal of Applied Statistics, 50(5):1152–1177, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tang:2023:MBJ

[4480] Lin Tang, Pengcheng Zeng, Jian Qing Shi, and Won-Seok Kim. Model-based joint curve registration and classification. Journal of Applied Statistics, 50 (5):1178–1198, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vasconcelos:2023:RER

[4481] J. C. S. Vasconcelos, G. M. Cordeiro, E. M. M. Ortega, and G. O. Silva. A random effect regression based on the odd log-logistic generalized inverse Gaussian distribution. Journal of Applied Statistics, 50(5):1199–1214, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ruth:2023:FNS

[4482] David M. Ruth, Nicholas L. Wood, and Douglas N. VanDerwerken. Fully nonparametric survival analysis in the presence of time-dependent covariates and dependent censoring. Journal of Applied Statistics, 50(5):1215–1229, May 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Delgado-Vences:2023:IDS

[4483] F. Delgado-Vences, F. Baltazar-Larios, A. Ornelas Vargas, E. Morales-Bojórquez, V. H. Cruz-Escalona, and C. Salomón Aguilar. Inference for a discretized stochastic logistic differential equation and its application to biological growth. Journal of Applied Statistics, 50(6):1231–1254, June 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ferreira:2023:HPL

[4484] Clécio S. Ferreira, Camila Borelli Zeller, and Rafael R. de Oliveira Garcia. Heteroscedastic partially linear model under skew-normal distribution with application in ragweed pollen concentration. Journal of Applied Statistics, 50(6):1255–1282, June 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Eguasa:2023:LLR

[4485] O. Eguasa, E. Edionwe, and J. I. Mbegbu. Local linear regression and the problem of dimensionality: a remedial strategy via a new locally adaptive bandwidths selector. Journal of Applied Statistics, 50(6):1283–1309, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Papoutsis:2023:BHM

[4486] Panayotis Papoutsis, Tarn Duong, Bertrand Michel, and Anne Philippe. Bayesian hierarchical models for the prediction of the driver flow and passenger waiting times in a stochastic carpooling service. Journal of Applied Statistics, 50(6):1310–1333, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2023:TCM

[4487] Diana Q. Chen, Si-Qi Mao, and Xu-Feng Niu. Tests and classification methods in adaptive designs with applications. Journal of Applied Statistics, 50(6):1334–1357, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nascimento:2023:CTP

[4488] A. D. C. Nascimento, Leandro C. Rêgo, and Jonas W. A. Silva. Compound truncated Poisson gamma distribution for understanding multimodal SAR intensities. Journal of Applied Statistics, 50(6):1358–1377, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liang:2023:SRL

[4489] Jinwen Liang and Maozai Tian. Sparse regression for low-dimensional timedynamic varying coefficient models with application to air quality data. Journal of Applied Statistics, 50(6): 1378–1399, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2023:PMG

[4490] Jie Liu, Haojie Chen, and Yang Yang.

Prediction models with graph kernel regularization for network data. Journal of Applied Statistics, 50(6):1400– 1417, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tzeng:2023:SRC

[4491] ShengLi Tzeng, Chun-Shu Chen, Yu-Fen Li, and Jin-Hua Chen. On summary ROC curve for dichotomous diagnostic studies: an application to metaanalysis of COVID-19. Journal of Applied Statistics, 50(6):1418–1434, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Potthoff:2023:SIB

[4492] Richard F. Potthoff. Spot It! and balanced block designs: keys to better debate architecture for a plethora of candidates in presidential primaries? Journal of Applied Statistics, 50(6): 1435–1454, June 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2023:RAP

[4493] Woojoo Lee, Jeonghwan Kim, and Donghwan Lee. Revisiting the analysis pipeline for overdispersed Poisson and binomial data. Journal of Applied Statistics, 50(7):1455–1476, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wu:2023:RLC

[4494] Chunjie Wu, Zhijun Wang, Steven MacEachern, and Jingjing Schneider. A robust latent CUSUM chart for monitoring customer attrition. Journal of Applied Statistics, 50(7):1477–1495, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Anzarmou:2023:KIF

[4495] Youssef Anzarmou, Abdallah Mkhadri, and Karim Oualkacha. The Kendall interaction filter for variable interaction screening in high dimensional classification problems. Journal of Applied Statistics, 50(7):1496–1514, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

McElroy:2023:VTR

[4496] T. S. McElroy and Thomas Trimbur. Variable targeting and reduction in large vector autoregressions with applications to workforce indicators. Journal of Applied Statistics, 50 (7):1515–1537, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saini:2023:IMS

[4497] Shubham Saini, Sachin Tomer, and Renu Garg. Inference of multicomponent stress-strength reliability following Topp–Leone distribution using progressively censored data. Journal of Applied Statistics, 50(7):1538–1567, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gomes:2023:FIR

[4498] José Clelto Barros Gomes, Reiko Aoki, Victor Hugo Lachos, Gilberto Alvarenga Paula, and Cibele Maria Russo. Fast inference for robust nonlinear mixed-effects models. Journal of Applied Statistics, 50(7):1568–1591, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Burch:2023:CIR

[4499] Brent Burch and Jesse Egbert. Confidence intervals for ratios of means applied to corpus-based word frequency classes. Journal of Applied Statistics, 50(7):1592–1610, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2023:SAE

[4500] Qihuang Zhang and Grace Y. Yi. Sensitivity analysis of error-contaminated time series data under autoregressive models with the application of COVID-19 data. Journal of Applied Statistics, 50(7):1611–1634, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Paleologou:2023:IDB

[4501] Suzanna-Maria Paleologou. Income and democracy: a bivariate copula approach. Journal of Applied Statistics, 50(7):1635–1649, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ueki:2023:BNB

[4502] Masao Ueki. Beta-negative binomial nonlinear spatio-temporal random effects modeling of COVID-19 case counts in Japan. Journal of Applied Statistics, 50(7):1650–1663, July 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hashimoto:2023:RPI

[4503] E. M. Hashimoto, E. M. M. Ortega, G. M. Cordeiro, V. G. Cancho, and I. Silva. The re-parameterized inverse Gaussian regression to model length of stay of COVID-19 patients in the public health care system of Piracicaba, Brazil. Journal of Applied Statistics, 50 (8):1665–1685, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2023:FLM

[4504] Mingming Liu, Jing Yang, Yushi Liu, Bochao Jia, Yun-Fei Chen, Luna Sun, and Shujie Ma. A fusion learning method to subgroup analysis of Alzheimer’s disease. Journal of Applied Statistics, 50(8):1686–1708, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Attwood:2023:ASE

[4505] Kristopher Attwood, Surui Hou, and Alan Hutson. Application of the skew exponential power distribution to ROC curves. Journal of Applied Statistics, 50 (8):1709–1724, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Key:2023:STP

[4506] Melissa C. Key, Susanne Ragg, and Benzion Boukai. A statistical testing procedure for validating class labels. Journal of Applied Statistics, 50 (8):1725–1749, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kokaew:2023:LCE

[4507] Angkana Kokaew, Winai Bodhisuwan, Su-Fen Yang, and Andrei Volodin. Logarithmic confidence estimation of a ratio of binomial proportions for dependent populations. Journal of Applied Statistics, 50(8):1750–1771, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kersey:2023:DAM

[4508] Jing Kersey, Hani Samawi, Jingjing Yin, Haresh Rochani, and Xinyan Zhang. On diagnostic accuracy measure with cut-points criterion for ordinal disease classification based on concordance and discordance. Journal of Applied Statistics, 50(8):1772–1789, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Du:2023:APL

[4509] Ruofei Du, Li Luo, Laurie G. Hudson, Sara Nozadi, and Johnnye Lewis. An adjusted partial least squares regression framework to utilize additional exposure information in environmental mixture data analysis. Journal of Applied Statistics, 50(8):1790–1811, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Neelon:2023:MSM

[4510] Brian Neelon, Chun-Che Wen, and Sara E. Benjamin-Neelon. A multivariate spatiotemporal model for tracking COVID-19 incidence and death rates in socially vulnerable populations. Journal of Applied Statistics, 50 (8):1812–1835, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kaombe:2023:NOS

[4511] Tsirizani M. Kaombe and Samuel O. M. Manda. A novel outlier statistic in multivariate survival models and its application to identify unusual underfive mortality sub-districts in Malawi. Journal of Applied Statistics, 50(8): 1836–1852, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Poonvoralak:2023:BMC

[4512] Wantanee Poonvoralak. Bayesian Markov Chain Monte Carlo for reparameterized stochastic volatility models using Asian FX rates during Covid-19. Journal of Applied Statistics, 50(8):1853–1875, August 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yadav:2023:CRS

[4513] C. P. Yadav, Sanjeev K. Tomer, and M. S. Panwar. A competing risk study of menarcheal age distribution based on non-recall current status data. Journal of Applied Statistics, 50(9): 1877–1899, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Guo:2023:LCO

[4514] Guangbao Guo, Yue Sun, Guoqi Qian, and Qian Wang. LIC criterion for optimal subset selection in distributed interval estimation. Journal of Applied Statistics, 50(9):1900–1920, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2023:GFI

[4515] Yanqin Feng, Jie Wang, and Yang Li.

Goodness-of-fit inference for the additive hazards regression model with clustered current status data. Journal of Applied Statistics, 50(9):1921–1941, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bak:2023:ISS

[4516] Kwan-Young Bak, Jae-Kyung Shin, and Ja-Yong Koo. Intrinsic spherical smoothing method based on generalized Bézier curves and sparsity inducing penalization. Journal of Applied Statistics, 50(9):1942–1961, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2023:ICA

[4517] Jung Wun Lee and Ofer Harel. Incomplete clustering analysis via multiple imputation. Journal of Applied Statistics, 50(9):1962–1979, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2023:DEE

[4518] Xinhua Liu and Zhezhen Jin. On detecting the effect of exposure mixture. Journal of Applied Statistics, 50(9): 1980–1991, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kim:2023:DDC

[4519] Hyune-Ju Kim, Huann-Sheng Chen, Douglas Midthune, Bill Wheeler, Dennis W. Buckman, Donald Green, Jeffrey Byrne, Jun Luo, and Eric J. Feuer. Data-driven choice of a model selection method in joinpoint regression. Journal of Applied Statistics, 50(9): 1992–2013, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2023:BNB

[4520] Xun Li, Joyee Ghosh, and Gabriele Villarini. Bayesian negative binomial regression model with unobserved covariates for predicting the frequency of North Atlantic tropical storms. Journal of Applied Statistics, 50(9):2014–2035, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2023:STF

[4521] Xin Zhao, Stuart Barber, Charles C.

Taylor, Xiaokai Nie, and Wenqian Shen. Spatio-temporal forecasting using wavelet transform-based decision trees with application to air quality and covid-19 forecasting. Journal of Applied Statistics, 50(9):2036–2054, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2023:EPC

[4522] Shu Wang, Elias Sayour, and Ji-Hyun Lee. Evaluation of phase I clinical trial designs for combinational agents along with guidance based on simulation studies. Journal of Applied Statistics, 50(9):2055–2078, September 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alevizakos:2023:DGW

[4523] Vasileios Alevizakos, Kashinath Chatterjee, Christos Koukouvinos, and Angeliki Lappa. A double generally weighted moving average control chart for monitoring the process variability. Journal of Applied Statistics, 50(10): 2079–2107, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Morales:2023:SNP

[4524] Fidel Ernesto Castro Morales and Daniele Torres Rodrigues. Spatiotemporal nonhomogeneous Poisson model with a seasonal component applied to the analysis of extreme rainfall. Journal of Applied Statistics, 50(10):

2108–2126, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vipin:2023:SPS

[4525] N. Vipin, Indranil Ghosh, and S. M. Sunoj. Some properties of stoploss moments under biased sampling. Journal of Applied Statistics, 50(10): 2127–2150, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Su:2023:EEC

[4526] Pei-Fang Su, Junjiang Zhong, Yi-Chia Liu, Tzu-Hsuan Lin, and Huang-Tz Ou. Efficient estimation of a Cox model when integrating the subgroup incidence rate information. Journal of Applied Statistics, 50(10):2151–2170, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

vanOppen:2023:MDP

[4527] Yulan B. van Oppen, Gabi Milder-Mulderij, Christophe Brochard, Rink Wiggers, Saskia de Vries, Wim P. Krijnen, and Marco A. Grzegorczyk. Modeling dragonfly population data with a Bayesian bivariate geometric mixedeffects model. Journal of Applied Statistics, 50(10):2171–2193, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Saraiva:2023:HBA

[4528] E. F. Saraiva, L. Sauer, and C. A. B. Pereira. A hierarchical Bayesian approach for modeling the evolution of the 7-day moving average of the number of deaths by COVID-19. Journal of Applied Statistics, 50(10):2194–2208, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Brant:2023:FLS

[4529] Simon Boge Brant and Ingrid Hobæk Haff. The fraud loss for selecting the model complexity in fraud detection. Journal of Applied Statistics, 50(10): 2209–2227, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Best:2023:EDN

[4530] Ana F. Best, Yaakov Malinovsky, and Paul S. Albert. The efficient design of nested group testing algorithms for disease identification in clustered data. Journal of Applied Statistics, 50(10): 2228–2245, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ries:2023:RBM

[4531] Daniel Ries and Alicia Carriquiry. The relationship between moderate to vigorous physical activity and metabolic syndrome: a Bayesian measurement error approach. Journal of Applied Statistics, 50(10):2246–2266, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elayouty:2023:FDA

[4532] A. Elayouty and H. Abou-Ali. Functional data analysis of the relationship between electricity consumption and climate change drivers. Journal of Applied Statistics, 50(10):2267–2285, October 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Stromberg:2023:ESI

[4533] Arnold Stromberg, Jie Chen, Teresa Paula Costa Azinheira Oliveira, Yichuan Zhao, Ramin Moghaddass, and Milan Stehlik. Editorial to the special issue: statistical perspectives on analytics for COVID-19 data. Journal of Applied Statistics, 50(11-12): 2287–2293, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Gajardo:2023:PPM

[4534] Álvaro Gajardo and Hans-Georg Müller. Point process models for COVID-19 cases and deaths. Journal of Applied Statistics, 50(11-12): 2294–2309, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rawat:2023:STS

[4535] Siddharth Rawat and Soudeep Deb. A spatio-temporal statistical model to analyze COVID-19 spread in the USA. Journal of Applied Statistics, 50(11-12):2310–2329, November 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Atanasov:2023:SMC

[4536] D. Atanasov, Vessela Stoimenova, and Nikolay M. Yanev. Statistical modelling of COVID-19 pandemic development applying branching processes. Journal of Applied Statistics, 50(11-12):2330–2342, November 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tchorbadjieff:2023:RCC

[4537] A. Tchorbadjieff, L. P. Tomov, V. Velev, G. Dezhov, V. Manev, and P. Mayster. On regime changes of COVID-19 outbreak. Journal of Applied Statistics, 50(11-12):2343–2359, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Akdi:2023:ICC

[4538] Yilmaz Akdi, Yunus Emre Karamanoglu, Kamil Demirberk Ünlü, and Cem Bas. Identifying the cycles in COVID-19 infection: the case of Turkey. Journal of Applied Statistics, 50(11-12):2360–2372, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yang:2023:TFC

[4539] Hou-Cheng Yang, Yishu Xue, Yuqing Pan, Qingyang Liu, and Guanyu Hu. Time fused coefficient SIR model with application to COVID-19 epidemic in the United States. Journal of Applied Statistics, 50(11-12):2373–2387, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Skamnia:2023:HSI

[4540] E. Skamnia, P. Economou, S. Bersimis, M. Frouda, A. Politis, and P. Alexopoulos. Hot spot identification method based on Andrews curves: an application on the COVID-19 crisis effects on caregiver distress in neurocognitive disorder. Journal of Applied Statistics, 50(11-12):2388–2407, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Wang:2023:CIU

[4541] Guannan Wang, Zhiling Gu, Xinyi Li, Shan Yu, Myungjin Kim, Yueying Wang, Lei Gao, and Li Wang. Comparing and integrating US COVID-19 data from multiple sources with anomaly detection and repairing. Journal of Applied Statistics, 50(11-12): 2408–2434, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Khmaladze:2023:EMS

[4542] Estate Khmaladze and Giorgi Kvizhinadze. On evolution model for SARS-Cov-2-infected population: the case of new Zealand. Journal of Applied Statistics, 50(11-12):2435–2449, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Elhambakhsh:2023:LSM

[4543] Fatemeh Elhambakhsh, Kamyar Sabri-Laghaie, and Rassoul Noorossana. A latent space model and Hotelling’s T^{2} control chart to monitor the networks of Covid-19 symptoms. Journal of Applied Statistics, 50(11-12): 2450–2472, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2023:CSP

[4544] Teng Chen, Paweł Polak, and Stanislav Uryasev. Classification and severity progression measure of COVID-19 patients using pairs of multi-omic factors. Journal of Applied Statistics, 50(11-12):2473–2503, November 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Du:2023:CAS

[4545] Peipei Du, Peihua Cao, Xiaodong Yan, Daihai He, Xiaotong Zhang, Weixiang Chen, Jiawei Luo, Ziqian Zeng, Yaolong Chen, Lin Yang, Shu Yang, and Xixi Feng. A continuous age-specific standardized mortality ratio for estimating the unascertained rates in the early epidemic of COVID-19 in different regions. Journal of Applied Statistics, 50(11-12):2504–2517, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Donnat:2023:MHC

[4546] Claire Donnat and Susan Holmes.

Modeling the heterogeneity in COVID-19’s reproductive number and its impact on predictive scenarios. Journal of Applied Statistics, 50(11-12): 2518–2546, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tian:2023:DDA

[4547] Ting Tian, Jingwen Zhang, Shiyun Lin, Yukang Jiang, Jianbin Tan, Zhongfei Li, and Xueqin Wang. Data-driven analysis of the simulations of the spread of COVID-19 under different interventions of China. Journal of Applied Statistics, 50(11-12):2547–2560, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2023:FWD

[4548] Xiaolei Zhang and Renjun Ma. Forecasting waved daily COVID-19 death count series with a novel combination of segmented Poisson model and ARIMA models. Journal of Applied Statistics, 50(11-12):2561–2574, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bacci:2023:LTM

[4549] S. Bacci, R. Fabbricatore, and Maria Iannario. Latent trait models for perceived risk assessment using a Covid-19 data survey. Journal of Applied Statistics, 50(11-12):2575–2598, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cai:2023:EID

[4550] Bryan Cai, John P. A. Ioannidis, Eran Bendavid, and Lu Tian. Exact inference for disease prevalence based on a test with unknown specificity and sensitivity. Journal of Applied Statistics, 50(11-12):2599–2623, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2023:IFI

[4551] Peiyi Zhang, Tianning Dong, Ninghui Li, and Faming Liang. Identification of factors impacting on the transmission and mortality of COVID-19. Journal of Applied Statistics, 50(11-12): 2624–2647, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Maleki:2023:RCC

[4552] Mohsen Maleki, Hamid Bidram, and Darren Wraith. Robust clustering of COVID-19 cases across U.S. counties using mixtures of asymmetric time series models with time varying and freely indexed covariates. Journal of Applied Statistics, 50(11-12): 2648–2662, November 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dayaratna:2023:HBS

[4553] Kevin D. Dayaratna, Drew Gonshorowski, and Mary Kolesar. Hierarchical Bayesian spatio-temporal modeling of COVID-19 in the United States. Journal of Applied Statistics, 50(11-12):2663–2680, November 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kundu:2023:SWP

[4554] Debasis Kundu. A stationary Weibull process and its applications. Journal of Applied Statistics, 50(13):2681–2700, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Qu:2023:BMS

[4555] Kai Qu and Jonathan R. Bradley.

Bayesian models for spatial count data with informative finite populations with application to the American community survey. Journal of Applied Statistics, 50(13):2701–2716, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Patil:2023:AMU

[4556] S. M. Patil and H. V. Kulkarni.

Analysis of medians under two-way model with and without interaction for Birnbaum–Saunders distributed response. Journal of Applied Statistics, 50(13):2717–2738, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhang:2023:JML

[4557] Zili Zhang, Christiana Charalambous, and Peter Foster. Joint modelling of longitudinal measurements and survival times via a multivariate copula approach. Journal of Applied Statistics, 50(13):2739–2759, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Martel:2023:BHM

[4558] M. Martel, M. A. Negrín, and F. J. Vázquez-Polo. Bayesian heterogeneity in a meta-analysis with two studies and binary data. Journal of Applied Statistics, 50(13):2760–2776, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ries:2023:AAP

[4559] Daniel Ries and Alicia Carriquiry. Assessing adult physical activity and compliance with 2008 CDC guidelines using a Bayesian two-part measurement error model. Journal of Applied Statistics, 50(13):2777–2795, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Periyasamypandian:2023:DNM

[4560] Jeyadurga Periyasamypandian and Saminathan Balamurali. Determination of new multiple deferred state sampling plan with economic perspective under Weibull distribution. Journal of Applied Statistics, 50(13):2796–2816, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zahariah:2023:MRC

[4561] Siti Zahariah and Habshah Midi. Minimum regularized covariance determinant and principal component analysisbased method for the identification of high leverage points in high dimensional sparse data. Journal of Applied Statistics, 50(13):2817–2835, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Velthoen:2023:FVS

[4562] Jasper Velthoen, Juan-Juan Cai, and Geurt Jongbloed. Forward variable selection for random forest models. Journal of Applied Statistics, 50(13):2836– 2856, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mei:2023:ESI

[4563] Yajun Mei (Principal Guest Editor), Jay Bartroff (Co-Editor), Jie Chen (Co-Editor), Georgios Fellouris (Co-Editor), and Ruizhi Zhang (Co-Editor). Editorial to the special issue: modern streaming data analytics. Journal of Applied Statistics, 50(14): 2857–2861, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baron:2023:DEM

[4564] Michael Baron and Sergey V. Malov.

Detection and estimation of multiple transient changes. Journal of Applied Statistics, 50(14):2862–2888, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hu:2023:ARA

[4565] Jiuyun Hu, Yajun Mei, Sarah Holte, and Hao Yan. Adaptive resources allocation CUSUM for binomial count data monitoring with application to COVID-19 hotspot detection. Journal of Applied Statistics, 50(14):2889–2913, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lue:2023:IPS

[4566] Heng-Hui Lue and ShengLi Tzeng. Interpretable, predictive spatio-temporal models via enhanced pairwise directions estimation. Journal of Applied Statistics, 50(14):2914–2933, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lun:2023:STR

[4567] David Lun, Svenja Fischer, Alberto Viglione, and Günter Blöschl. Significance testing of rank cross-correlations between autocorrelated time series with short-range dependence. Journal of Applied Statistics, 50(14):2934–2950, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tian:2023:ALB

[4568] Hongzhen Tian, Reuven Zev Cohen, Chuck Zhang, and Yajun Mei. Active learning-based multistage sequential decision-making model with application on common bile duct stone evaluation. Journal of Applied Statistics, 50(14):2951–2969, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsang:2023:KPF

[4569] Ka Wai Tsang, Fugee Tsung, and Zhihao Xu. Knockoff procedure for false discovery rate control in highdimensional data streams. Journal of Applied Statistics, 50(14):2970–2983, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhan:2023:ROP

[4570] Yinglun Zhan, Ruizhi Zhang, Yuzhen Zhou, Vincent Stoerger, Jeremy Hiller, Tala Awada, and Yufeng Ge. Rapid online plant leaf area change detection with high-throughput plant image data. Journal of Applied Statistics, 50(14):2984–2998, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhao:2023:HSD

[4571] Yujie Zhao, Xiaoming Huo, and Yajun Mei. Hot-spots detection in count data by Poisson assisted smooth sparse tensor decomposition. Journal of Applied Statistics, 50(14):2999–3029, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Krahn:2023:EEA

[4572] Jody Krahn, Shakhawat Hossain, and Shahedul Khan. An efficient estimation approach to joint modeling of longitudinal and survival data. Journal of Applied Statistics, 50(15):3031–3047, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Montgomery:2023:FTE

[4573] R. N. Montgomery, L. T. Ptomey, and J. D. Mahnken. A flexible test for early-stage studies with multiple endpoints. Journal of Applied Statistics, 50(15):3048–3061, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shafie:2023:GFT

[4574] Termeh Shafie. Goodness of fit tests for random multigraph models. Journal of Applied Statistics, 50(15):3062–3087, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cefis:2023:OHP

[4575] Mattia Cefis. Observed heterogeneity in players’ football performance analysis using PLS-PM. Journal of Applied Statistics, 50(15):3088–3107, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Birbicer:2023:PES

[4576] Ismet Birbiçer and Ali I. Genç. On parameter estimation of the standard omega distribution. Journal of Applied Statistics, 50(15):3108–3124, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sayyari:2023:CIS

[4577] Yamin Sayyari, Mohammad Reza Molaei, and Adel Mehrpooya. Complexities of information sources. Journal of Applied Statistics, 50(15):3125–3141, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kalina:2023:TEM

[4578] Jan Kalina and Patrik Janácek. Testing exchangeability of multivariate distributions. Journal of Applied Statistics, 50(15):3142–3156, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bui:2023:PLM

[4579] Minh Thu Bui, Cornelis J. Potgieter, and Akihito Kamata. Penalized likelihood methods for modeling count data. Journal of Applied Statistics, 50(15):3157–3176, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lam:2023:MMM

[4580] Ka Kin Lam and Bo Wang. Multipopulation mortality modelling and forecasting: the weighted multivariate functional principal component approaches. Journal of Applied Statistics, 50(15):3177–3198, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Darzi:2023:TIP

[4581] Fatemeh Hassantabar Darzi, Samaneh Eftekhari Mahabadi, and Firoozeh Haghighi. Type-II progressive censoring with GLM-based random removal mechanism dependent on the experimental conditions. Journal of Applied Statistics, 50(16):3199–3228, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chaudhuri:2023:STM

[4582] Somnath Chaudhuri, Pablo Juan, and Jorge Mateu. Spatio-temporal modeling of traffic accidents incidence on urban road networks based on an explicit network triangulation. Journal of Applied Statistics, 50(16):3229–3250, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2023:EVW

[4583] Zhan Liu, Junbo Zheng, Chaofeng Tu, and Yingli Pan. Estimation for volunteer web survey samples using a modelaveraging approach. Journal of Applied Statistics, 50(16):3251–3271, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Coenders:2023:TAS

[4584] Germà Coenders and Michael Greenacre.

Three approaches to supervised learning for compositional data with pairwise logratios. Journal of Applied Statistics, 50(16):3272–3293, 2023. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2023:NNA

[4585] Soyun Park, Jihnhee Yu, Hwa-Hyoung Woo, and Chun Gun Park. A novel network architecture combining centralperipheral deviation with image-based convolutional neural networks for diffusion tensor imaging studies. Journal of Applied Statistics, 50(16):3294–3311, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xiong:2023:RSL

[4586] Wei Xiong, Maozai Tian, Manlai Tang, and Han Pan. Robust and sparse learning of varying coefficient models with high-dimensional features. Journal of Applied Statistics, 50(16):3312– 3336, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Camli:2023:VSL

[4587] Onur Camli, Zeynep Kalaylioglu, and Ashis SenGupta. Variable selection in linear-circular regression models. Journal of Applied Statistics, 50(16):3337– 3361, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luque:2023:BSV

[4588] Carolina Luque and Juan Sosa. A Bayesian spatial voting model to characterize the legislative behavior of the Colombian Senate 2010-2014. Journal of Applied Statistics, 50(16):3362–3383, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rathnayake:2023:GCS

[4589] Nirosha Rathnayake, Hongying Daisy Dai, Richard Charnigo, Kendra Schmid, and Jane Meza. A general class of small area estimation using calibrated hierarchical likelihood approach with applications to COVID-19 data. Journal of Applied Statistics, 50(16):3384–3404, 2023. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Dutta:2024:EPB

[4590] Subhankar Dutta and Suchandan Kayal. Estimation and prediction for Burr type III distribution based on unified progressive hybrid censoring scheme. Journal of Applied Statistics, 51(1):1–33, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2024:NME

[4591] Wenchao Xu, Hongmei Lin, Tiejun Tong, and Riquan Zhang. A new method for estimating Sharpe ratio function via local maximum likelihood. Journal of Applied Statistics, 51(1):34–52, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Hao:2024:MAE

[4592] Hao Hao, Bai Huang, and Tae hwy Lee. Model averaging estimation of panel data models with many instruments and boosting. Journal of Applied Statistics, 51(1):53–69, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ip:2024:MRF

[4593] Ryan H. L. Ip and K. Y. K. Wu. A Markov random field model with cumulative logistic functions for spatially dependent ordinal data. Journal of Applied Statistics, 51(1):70–86, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fanjul-Hevia:2024:TCC

[4594] A. Fanjul-Hevia, J. C. Pardo-Fernández, I. Van Keilegom, and W. González-Manteiga. A test for comparing conditional ROC curves with multidimensional covariates. Journal of Applied Statistics, 51(1):87–113, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Molinari:2024:BDN

[4595] Marco Molinari, Andrea Cremaschi, Maria De Iorio, Nishi Chaturvedi, Alun Hughes, and Therese Tillin. Bayesian dynamic network modelling: an application to metabolic associations in cardiovascular diseases. Journal of Applied Statistics, 51(1):114–138, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chang:2024:CCS

[4596] Peter Chang, Rongzi Liu, Tingting Hou, Xinyu Yan, and Guogen Shan. Continuity corrected score confidence interval for the difference in proportions in paired data. Journal of Applied Statistics, 51(1):139–152, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodriguez:2024:DMD

[4597] Carlos E. Rodríguez, Luis E. Nieto-Barajas, and Carlos S. Pérez-Pérez.

Dealing with missing data under stratified sampling designs where strata are study domains. Journal of Applied Statistics, 51(1):153–167, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2024:TSL

[4598] Cindy Feng and Xi Chen. A twostage latent factor regression method to model the common and unique effects of multiple highly correlated exposure variables. Journal of Applied Statistics, 51(1):168–192, January 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sun:2024:BRN

[4599] Jianguo Sun. Book review: New advances in statistics and data science, edited by Ding-Geng Chen, Gang Li, Aiyi Liu, Zhezhen Jin, Yi Li and Yichuan Zhao, Springer, Cham, Switzerland, 2018, 348 pp., $24.36 (ebook), ISBN 978-3-319-69416-0, 3-319-69416-2. Journal of Applied Statistics, 51(1):193–195, January 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sosa:2024:BAD

[4600] Juan Sosa and Abel Rodríguez. A Bayesian approach for de-duplication in the presence of relational data. Journal of Applied Statistics, 51(2): 197–215, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mohsin:2024:SMN

[4601] Muhammad Mohsin and Albrecht Gebhardt. A stochastic model for NFL games and point spread assessment.

Journal of Applied Statistics, 51(2): 216–229, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mahdi:2024:BPM

[4602] Esam Mahdi and Thomas J. Fisher. Bootstrapping a powerful mixed portmanteau test for time series. Journal of Applied Statistics, 51(2):230–255, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ibrar:2024:CSD

[4603] Faryal Ibrar, Sajid Ali, and Ismail Shah. A comparison of single- and double-threshold ROC plots for mixture distributions. Journal of Applied Statistics, 51(2):256–278, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Salaki:2024:MAC

[4604] Deiby Tineke Salaki, Anang Kurnia, Bagus Sartono, I. Wayan Mangku, and Arief Gusnanto. Model averaging in calibration of near-infrared instruments with correlated high-dimensional data. Journal of Applied Statistics, 51(2): 279–297, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Xu:2024:SMA

[4605] Qian Xu, Demetra Antimisiaris, and Maiying Kong. Statistical methods for assessing drug interactions using observational data. Journal of Applied Statistics, 51(2):298–323, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nascimento:2024:KBR

[4606] A. D. C. Nascimento, P. M. Almeida-Junior, J. M. Vasconcelos, and A. P. M. Borges-Junior. K-Bessel regression model for speckled data. Journal of Applied Statistics, 51(2):324–347, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mohamed:2024:SLM

[4607] Heba Soltan Mohamed, Gauss M.

Cordeiro, R. Minkah, Haitham M. Yousof, and Mohamed Ibrahim. A size-of-loss model for the negatively skewed insurance claims data: applications, risk analysis using different methods and statistical forecasting. Journal of Applied Statistics, 51 (2):348–369, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ogbagaber:2024:HMM

[4608] Semhar B. Ogbagaber, Yifan Cui, Kaigang Li, Ronald J. Iannotti, and Paul S. Albert. A hidden Markov modeling approach combining objective measure of activity and subjective measure of self-reported sleep to estimate the sleep-wake cycle. Journal of Applied Statistics, 51(2):370–387, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Feng:2024:JMM

[4609] Shengxiao Vincent Feng, Willem van den Boom, Maria De Iorio, Gladi J. Thng, Jerry K. Y. Chan, Helen Y. Chen, Kok Hian Tan, and Michelle Z. L. Kee. Joint modelling of mental health markers through pregnancy: a Bayesian semi-parametric approach.

Journal of Applied Statistics, 51(2): 388–405, February 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kumar:2024:ACR

[4610] Pushkal Kumar, Manas Ranjan Tripathy, and Somesh Kumar. Alternative classification rules for two inverse Gaussian populations with a common mean and order restricted scale-like parameters. Journal of Applied Statistics, 51(3):407–429, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chen:2024:SPL

[4611] Le Chen, Ruochen Tian, Guanjie Chen, Ao Yuan, Chuan-Ming Li, Amy R. Bentley, Howard J. Hoffman, and Charles Rotimi. Semiparametric partial linear modeling of risk factors for ear infections: the Early Childhood Longitudinal Study. Journal of Applied Statistics, 51(3):430–450, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lv:2024:SIG

[4612] Qi Lv, Yajie Tian, and Wenhao Gui. Statistical inference for Gompertz distribution under adaptive type-II progressive hybrid censoring. Journal of Applied Statistics, 51(3):451–480, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2024:IPB

[4613] Han Yu and Alan D. Hutson. Inferential procedures based on the weighted Pearson correlation coefficient test statistic. Journal of Applied Statistics, 51(3):481–496, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Samawi:2024:PDT

[4614] Hani Samawi, Ding-Geng Chen, Jingjing Yin, and Marwan Alsharman. Performance of diagnostic tests based on continuous bivariate markers. Journal of Applied Statistics, 51 (3):497–514, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Diks:2024:NNM

[4615] Cees Diks and Marcin Wolski. New nonparametric measures for instantaneous and Granger-causality tail codependence. Journal of Applied Statistics, 51(3):515–533, March 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shat:2024:ODS

[4616] Helmi Shat. Optimal design of stress levels in accelerated degradation testing for multivariate linear degradation models. Journal of Applied Statistics, 51(3):534–554, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zeng:2024:RTM

[4617] Ting Zeng and Solomon W. Harrar.

Robust tests for multivariate repeated measures with small samples. Journal of Applied Statistics, 51(3):555–580, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Roy:2024:ICP

[4618] Archi Roy, Soudeep Deb, and Divya Chakarwarti. Impact of COVID-19 on public social life and mental health: a statistical study of Google Trends data from the USA. Journal of Applied Statistics, 51(3):581–605, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Skripnikov:2024:PCG

[4619] A. Skripnikov. Partially constrained group variable selection to adjust for complementary unit performance in American college football. Journal of Applied Statistics, 51(3):606–620, March 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ge:2024:BAQ

[4620] Wentao Ge and Junfeng Shang. Bootstrap-adjusted quasi-likelihood information criteria for mixed model selection. Journal of Applied Statistics, 51(4):621–645, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Zhong:2024:HES

[4621] Zihang Zhong, Min Yang, Senmiao Ni, Lixin Cai, Jingwei Wu, Jianling Bai, and Hao Yu. The heterogeneity effect of surveillance intervals on progression free survival. Journal of Applied Statistics, 51(4):646–663, April 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vila:2024:MBR

[4622] Roberto Vila, Lucas Alfaia, André F. B. Menezes, Mehmet N. Çankaya, and Marcelo Bourguignon. A model for bimodal rates and proportions. Journal of Applied Statistics, 51(4):664–681, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sharma:2024:SRC

[4623] Padma Sharma. Selection of random coefficients in ordered response models: a framework to detect heterogeneity in household surveys. Journal of Applied Statistics, 51(4):682–700, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

doNascimento:2024:RMF

[4624] Fernando Ferraz do Nascimento and Aline Raquel Assunção Nunes. Regression models for the full distribution to exceedance data. Journal of Applied Statistics, 51(4):701–720, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Geneyro:2024:BNM

[4625] Emiliano Geneyro and Gabriel Núñez-Antonio. A Bayesian nonparametric model for bounded directional data on the positive orthant of the unit sphere. Journal of Applied Statistics, 51(4):721–739, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Scaldelai:2024:TCM

[4626] D. Scaldelai, L. C. Matioli, and S. R.

Santos. TreeKDE: clustering multivariate data based on decision tree and using one-dimensional kernel density estimation. Journal of Applied Statistics, 51(4):740–758, April 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Luo:2024:SES

[4627] Lin Luo, Jinzhao Yu, and Hui Zhao.

The sparse estimation of the semiparametric linear transformation model with dependent current status data. Journal of Applied Statistics, 51(4): 759–779, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

deUna-Alvarez:2024:CRD

[4628] J. de Uña-Álvarez, A. I. Martínez-Senra, M. S. Otero-Giráldez, and M. A. Quintás. Cox regression with doubly truncated responses and timedependent covariates: the impact of innovation on firm survival. Journal of Applied Statistics, 51(4):780–792, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vorpe:2024:CRD

[4629] Katherine Vorpe, Sierra Hessinger, Rebekah Poth, and Tatjana Miljkovic. Clustering regions with dynamic time warping to model obesity prevalence disparities in the United States. Journal of Applied Statistics, 51(4):793– 807, April 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Han:2024:OCC

[4630] Dong Han, Fugee Tsung, and Lei Qiao. The optimal CUSUM control chart with a dynamic non-random control limit and a given sampling strategy for small samples sequence. Journal of Applied Statistics, 51(5):809–825, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Prataviera:2024:NFR

[4631] F. Prataviera, E. M. Hashimoto, E. M. M. Ortega, G. M. Cordeiro, V. G. Cancho, and R. Vila. A new flexible regression model with application to recovery probability Covid-19 patients. Journal of Applied Statistics, 51(5):826–844, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Salam:2024:LSM

[4632] Abdul Salam and Marco Grzegorczyk.

Learning the structure of the mTOR protein signaling pathway from protein phosphorylation data. Journal of Applied Statistics, 51(5):845–865, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Egbon:2024:MSP

[4633] Osafu Augustine Egbon and Ezra Gayawan. Modeling the spatial patterns of antenatal care utilization in Nigeria with inference based on Pólya-Gamma mixtures. Journal of Applied Statistics, 51(5):866–890, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Siriwardhana:2024:SOP

[4634] Chathura Siriwardhana, K. B. Kulasekera, and Somnath Datta. Selection of the optimal personalized treatment from multiple treatments with rightcensored multivariate outcome measures. Journal of Applied Statistics, 51(5):891–912, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Imran:2024:IZS

[4635] Muhammad Imran, Jinsheng Sun, Xuelong Hu, Fatima Sehar Zaidi, and Anan Tang. Investigating zero-state and steady-state performance of MEWMA-CoDa control chart using variable sampling interval. Journal of Applied Statistics, 51(5):913–934, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2024:APU

[4636] Yifan Li, Chunjie Wu, Zhijun Wang, and Zhiming Hu. Aggregated parameter update schemes for monitoring binary profiles. Journal of Applied Statistics, 51(5):935–957, May 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Sousa:2024:BAS

[4637] Pedro Henrique T. O. Sousa, Camila P. E. de Souza, and Ronaldo Dias. Bayesian adaptive selection of basis functions for functional data representation. Journal of Applied Statistics, 51(5):958–992, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Choiruddin:2024:KCK

[4638] Achmad Choiruddin, Tabita Yuni Susanto, Ahmad Husain, and Yuniar Mega Kartikasari. kppmenet: combining the kppm and elastic net regularization for inhomogeneous Cox point process with correlated covariates. Journal of Applied Statistics, 51(5):993–1006, May 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vasconcelos:2024:ASM

[4639] Julio C. S. Vasconcelos, Thiago da Costa Travassos, Edwin M. M. Ortega, Gauss M. Cordeiro, and Leonardo Oliveira Reis. Alternative statistical modeling for radical prostatectomy data. Journal of Applied Statistics, 51(5):1007–1022, May 2024.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Acharyya:2024:MSI

[4640] Satwik Acharyya, Debdeep Pati, Shumei Sun, and Dipankar Bandyopadhyay. A monotone single index model for missing-at-random longitudinal proportion data. Journal of Applied Statistics, 51(6):1023–1040, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yu:2024:SRI

[4641] Zhou Yu, Jie Yang, and Hsin-Hsiung Huang. Smoothing regression and impact measures for accidents of traffic flows. Journal of Applied Statistics, 51(6):1041–1056, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liang:2024:IMT

[4642] Jinwen Liang and Maozai Tian. Imputed mean tensor regression for nearsited spatial temporal data. Journal of Applied Statistics, 51(6):1057–1075, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cakar:2024:HSM

[4643] Serenay Cakar and Fulya Gokalp Yavuz. Hybrid statistical and machine learning modeling of cognitive neuroscience data. Journal of Applied Statistics, 51(6):1076–1097, June 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peng:2024:VSQ

[4644] Bo Peng, Kai Yang, and Xiaogang Dong. Variable selection for quantile autoregressive model: Bayesian methods versus classical methods. Journal of Applied Statistics, 51(6):1098–1130, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Diop:2024:ECP

[4645] Mamadou Lamine Diop and William Kengne. Epidemic change-point detection in general integer-valued time series. Journal of Applied Statistics, 51(6):1131–1150, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Li:2024:PRL

[4646] Canhui Li, Weirong Li, and Wensheng Zhu. Penalized robust learning for optimal treatment regimes with heterogeneous individualized treatment effects. Journal of Applied Statistics, 51(6):1151–1170, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alevizakos:2024:DFP

[4647] Vasileios Alevizakos, Kashinath Chatterjee, and Christos Koukouvinos. Distribution-free phase II triple EWMA control chart for joint monitoring the process location and scale parameters. Journal of Applied Statistics, 51 (6):1171–1190, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Angeles:2024:GFS

[4648] David Angeles, Sebastian Kurtek, Elizabeth Klein, Marielle Brinkman, and Amy Ferketich. Geometric framework for statistical analysis of eye tracking heat maps, with application to a tobacco waterpipe study. Journal of Applied Statistics, 51(6):1191–1209, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tupper:2024:CCC

[4649] Laura L. Tupper, Charles R. Keese, and David S. Matteson. Classifying contaminated cell cultures using time series features. Journal of Applied Statistics, 51(6):1210–1226, June 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Baladezaei:2024:BDB

[4650] Seyedeh Mahbubeh Hoseini Baladezaei, Einolah Deiri, and Ezzatallah Baloui Jamkhaneh. The balanced discrete Burr-hatke model and mixing INAR(1) process: properties, estimation, forecasting and COVID-19 applications. Journal of Applied Statistics, 51(7): 1227–1250, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ma:2024:EAH

[4651] Rui Ma, Shishun Zhao, Jianguo Sun, and Shuying Wang. Estimation of accelerated hazards models based on case K informatively interval-censored failure time data. Journal of Applied Statistics, 51(7):1251–1270, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ali:2024:BSS

[4652] Sajid Ali, Mariyam Waheed, Ismail Shah, and Syed Muhammad Muslim Raza. Bayesian sample size determination for coefficient of variation of normal distribution. Journal of Applied Statistics, 51(7):1271–1286, July 2024.

CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ye:2024:FPC

[4653] Jun Ye. Functional principal component models for sparse and irregularly spaced data by Bayesian inference. Journal of Applied Statistics, 51(7):1287–1317, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Liu:2024:RAM

[4654] Yonghui Liu, Jing Wang, Víctor Leiva, Alejandra Tapia, Wei Tan, and Shuangzhe Liu. Robust autoregressive modeling and its diagnostic analytics with a COVID-19 related application. Journal of Applied Statistics, 51(7):1318–1343, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Chahine:2024:ASO

[4655] Rouba A. Chahine and Inmaculada Aban. Analysis of survival outcomes using likelihood ratio test in trials incorporating patient’s treatment choice. Journal of Applied Statistics, 51(7):1344–1358, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Beyaztas:2024:RSF

[4656] Ufuk Beyaztas, Mujgan Tez, and Han Lin Shang. Robust scalar-on-function partial quantile regression. Journal of Applied Statistics, 51(7): 1359–1377, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cortes:2024:NFQ

[4657] Isaac E. Cortés, Mário de Castro, and Diego I. Gallardo. A new family of quantile regression models applied to nutritional data. Journal of Applied Statistics, 51(7):1378–1398, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Surjanovic:2024:IHL

[4658] Nikola Surjanovic and Thomas M.

Loughin. Improving the Hosmerlemeshow goodness-of-fit test in large models with replicated Bernoulli trials. Journal of Applied Statistics, 51 (7):1399–1411, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Poongadan:2024:FTS

[4659] Sameer Poongadan and M. C. Lineesh.

Forecasting of the true satellite carbon monoxide data with ensemble empirical mode decomposition, singular value decomposition and moving average. Journal of Applied Statistics, 51(7):1412–1426, July 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jin:2024:RES

[4660] Jun Jin, Shuangzhe Liu, and Tiefeng Ma. Robust and efficient subsampling algorithms for massive data logistic regression. Journal of Applied Statistics, 51(8):1427–1445, August 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vila:2024:GCT

[4661] Roberto Vila, Victor Serra, Mehmet Niyazi Çankaya, and Felipe Quintino. A general class of trimodal distributions: properties and inference. Journal of Applied Statistics, 51(8):1446–1469, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shi:2024:ETG

[4662] Weihua Shi and Wenhao Gui. Estimation for two Gompertz populations under a balanced joint progressive type-II censoring scheme. Journal of Applied Statistics, 51(8):1470–1496, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Jarquin:2024:CPG

[4663] D. Jarquin, A. Roy, B. Clarke, and S. Ghosal. Combining phenotypic and genomic data to improve prediction of binary traits. Journal of Applied Statistics, 51(8):1497–1523, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Amin:2024:FBA

[4664] Ayman A. Amin. Full Bayesian analysis of double seasonal autoregressive models with real applications. Journal of Applied Statistics, 51(8):1524–1544, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Aoki:2024:CNC

[4665] Reiko Aoki, Juan P. Mamani Bustamante, Cibele M. Russo, and Gilberto A. Paula. Conformal normal curvature and detection of masked observations in multivariate null intercept measurement error models. Journal of Applied Statistics, 51(8):1545–1569, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Petrovich:2024:IVE

[4666] Justin Petrovich, Bahaeddine Taoufik, and Zachary George Davis. Instrumental variable estimation for functional concurrent regression models. Journal of Applied Statistics, 51(8):1570–1589, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Park:2024:REE

[4667] Chanseok Park, Xuehong Gao, and Min Wang. Robust explicit estimators using the power-weighted repeated medians. Journal of Applied Statistics, 51 (8):1590–1608, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Langworthy:2024:EIC

[4668] Benjamin W. Langworthy, Zhaoxun Hou, Gary C. Curhan, Sharon G. Curhan, and Molin Wang. Estimating intracluster correlation for ordinal data. Journal of Applied Statistics, 51 (8):1609–1617, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Modak:2024:BRF

[4669] Soumita Modak. Book review: Finding groups in data: an introduction to cluster analysis, authored by Leonard Kaufman and Peter J. Rousseeuw, John Wiley and Sons, 2005, ISBN: 0-47-1-73578-7. Journal of Applied Statistics, 51(8):1618–1620, August 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Cai:2024:MMS

[4670] Yantao Cai, Liu Liu, and Zhonghua Li.

Modelling and monitoring social network change based on exponential random graph models. Journal of Applied Statistics, 51(9):1621–1641, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vigas:2024:GOL

[4671] Valdemiro P. Vigas, Edwin M. M. Ortega, Adriano K. Suzuki, Gauss M. Cordeiro, and Paulo C. dos Santos Junior. The generalized odd log-logistic-g regression with interval-censored survival data. Journal of Applied Statistics, 51(9):1642–1663, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nassar:2024:EPO

[4672] Mazen Nassar and Ahmed Elshahhat. Estimation procedures and optimal censoring schemes for an improved adaptive progressively type-II censored Weibull distribution. Journal of Applied Statistics, 51(9):1664–1688, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Koley:2024:CSD

[4673] Tamalika Koley and Anup Dewanji. Current status data with two competing risks and time-dependent missing failure types. Journal of Applied Statistics, 51(9):1689–1708, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Fan:2024:TSC

[4674] Yawen Fan, Xiaohui Liu, Ting Luo, Yao Rao, and Hanqing Li. Testing serial correlation in a general d-factor model with possible infinite variance. Journal of Applied Statistics, 51(9):

1709–1728, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Vila:2024:FBD

[4675] Roberto Vila, Narayanaswamy Balakrishnan, Helton Saulo, and Peter Zörnig. Family of bivariate distributions on the unit square: theoretical properties and applications. Journal of Applied Statistics, 51(9):1729–1755, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rauschenberger:2024:PDO

[4676] Armin Rauschenberger and Enrico Glaab. Predicting dichotomised outcomes from high-dimensional data in biomedicine. Journal of Applied Statistics, 51(9):1756–1771, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Peralta:2024:HBA

[4677] Danielle Peralta, Ricardo Puziol de Oliveira, and Jorge Alberto Achcar. A hierarchical Bayesian analysis for bivariate Weibull distribution under leftcensoring scheme. Journal of Applied Statistics, 51(9):1772–1791, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deng:2024:LBM

[4678] Dianliang Deng and Xiaoqing Zhang. A Lindley-binomial model for analyzing the proportions with sparseness and excessive zeros. Journal of Applied Statistics, 51(9):1792–1817, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Lima:2024:PAT

[4679] S. Lima, A. M. Gonçalves, and M. Costa. Predictive accuracy of time series models applied to economic data: the European countries retail trade. Journal of Applied Statistics, 51(9): 1818–1841, September 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yan:2024:CTL

[4680] Han Yan, Shuhan Lu, and Sanguo Zhang. The cluster D-trace loss for differential network analysis. Journal of Applied Statistics, 51(10):1843–1860, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Alver:2024:MCT

[4681] Sarah Alver and Guoyi Zhang. Multiple comparisons of treatment against control under unequal variances using parametric bootstrap. Journal of Applied Statistics, 51(10):1861–1877, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Atandoh:2024:SCD

[4682] Paul H. Atandoh and Kevin H. Lee. Statistical clustering of documents via stochastic blockmodels. Journal of Applied Statistics, 51(10):1878–1893, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kuran:2024:KMK

[4683] Özge Kuran and Seçil Yalaz. Kernel mixed and kernel stochastic restricted ridge predictions in the partially linear mixed measurement error models: an application to COVID-19.

Journal of Applied Statistics, 51(10): 1894–1918, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bhootna:2024:GPE

[4684] Niharika Bhootna and Arun Kumar.

GARTFIMA process and its empirical spectral density based estimation. Journal of Applied Statistics, 51(10): 1919–1945, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Workman:2024:SAI

[4685] Austin Workman and Joon Jin Song.

Spatial analysis for interval-valued data. Journal of Applied Statistics, 51 (10):1946–1960, October 2024. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Kumari:2024:GFT

[4686] Aditi Kumari and Deepesh Bhati.

Goodness-of-fit test for the one-sided Lévy distribution. Journal of Applied Statistics, 51(10):1961–1975, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bal:2024:PER

[4687] Debasmita Bal, Manas Ranjan Tripathy, and Somesh Kumar. Point estimation and related classification problems for several Lindley populations with application using COVID-19 data. Journal of Applied Statistics, 51(10): 1976–2006, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Watson:2024:PME

[4688] G. L. Watson, C. E. Reid, M. Jerrett, and D. Telesca. Prediction and model evaluation for space-time data. Journal of Applied Statistics, 51(10): 2007–2024, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yan:2024:NTW

[4689] Xingyu Yan, Jiaqian Yu, Weiyong Ding, Hao Wang, and Peng Zhao. A novel two-way functional linear model with applications in human mortality data analysis. Journal of Applied Statistics, 51(10):2025–2038, October 2024. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ryan:1997:MRM

[4690] Thomas P. Ryan. Modern Regression Methods. Wiley series in probability and statistics. Applied probability and statistics. Wiley, New York, NY, USA, 1997. ISBN 0-471-52912-5 (hardcover). xix + 515 pp. LCCN QA278.2 .R93 1997. URL

http://www.loc.gov/catdir/bios/
wiley047/96006368.htm;   http:/
/www.loc.gov/catdir/description/
wiley033/96006368.htm;   http:/
/www.loc.gov/catdir/toc/onix03/
96006368.htm.

Schaeffer:2000:CRM

[4691] H. A. (Hal A.) Schaeffer. Credit Risk Management: a Guide to Sound Business Decisions. Wiley, New York, NY, USA, 2000. ISBN 0-471-35020-6. viii + 280 pp. LCCN HG3751 .S33 2000. URL

http://www.loc.gov/catdir/bios/
wiley042/00267999.htm;   http:/
/www.loc.gov/catdir/description/

wiley033/00267999.htm;   http:/
/www.loc.gov/catdir/toc/onix04/
00267999.htm.

Dalgaard:2002:ISR

[4692] Peter Dalgaard. Introductory Statistics with R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95475-9. xv + 267 pp. LCCN QA276.4 .D33 2002.

Crowley:2006:HSC

[4693] John Crowley and Donna Pauler Ankerst, editors. Handbook of Statistics in Clinical Oncology. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2006. ISBN 0-8247-2339-2. 617 pp. LCCN RC267

.H36 2006. URL http://catdir.loc.
gov/catdir/enhancements/fy0648/
2006273842-d.html;   http://catdir.
loc.gov/catdir/toc/fy0606/2006273842.
html;   http://www.zentralblatt-
math.org/zmath/en/search/?an=1140.
62347.

Doncaster:2007:AVC

[4694] C. Patrick Doncaster and Andrew J. H.

Davey. Analysis of Variance and Covariance: How to Choose and Construct Models for the Life Sciences. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-86562-X (hardcover), 0-521-68447-1 (paperback). xiii + 287 pp. LCCN QH323.5

.D655 2007. URL http://www.loc.
gov/catdir/enhancements/fy0808/
2008270343-b.htm;   http://www.
loc.gov/catdir/enhancements/fy0808/
2008270343-d.htm;   http://www.
loc.gov/catdir/enhancements/fy0808/
2008270343-t.htm.

Bailey:2008:DCE

[4695] R. (Rosemary) Bailey. Design of Comparative Experiments. Cambridge series on statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-86506-9 (hardcover), 0-521-68357-2 (paperback). xiii + 330 pp. LCCN QA279 .B35 2008.

Casella:2008:SD

[4696] George Casella. Statistical Design. Springer texts in statistics. Springer Science+Business Media, New York, NY, USA, 2008. ISBN 0-387-75964-6 (hardcover), 0-387-75965-4 (e-book). xiii + 307 pp. LCCN QA279 .C375 2008.

Chambers:2008:SDA

[4697] John M. Chambers. Software for Data Analysis: Programming with R. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-75935-2 (paperback), 0-387-75936-0 (e-book). xiv + 498 pp. LCCN QA276.45.R3 C43 2008.

Claeskens:2008:MSM

[4698] Gerda Claeskens and Nils Lid Hjort. Model Selection and Model Averaging. Cambridge series in statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85225-0. xvii + 312 pp. LCCN QA276.18

                                        .C53 2008. URL http://www.loc.
                                        gov/catdir/enhancements/fy0901/
                                        2008006507-b.htm;   http://www.
                                        loc.gov/catdir/enhancements/fy0901/
                                        2008006507-d.htm;   http://www.

loc.gov/catdir/enhancements/fy0901/
2008006507-t.htm.

Cohen:2008:SDR

[4699] Yosef Cohen and Jeremiah Y. Cohen.

Statistics and Data with R: an Applied Approach Through Examples. Wiley, New York, NY, USA, 2008. ISBN 0-470-75805-8 (hardcover). xviii + 599 pp. LCCN QA276.45.R3 C64 2008.

URL http://www.loc.gov/catdir/
toc/ecip0824/2008032153.htm.

Dalgaard:2008:ISR

[4700] Peter Dalgaard. Introductory Statistics with R. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 0-387-79053-5 (paperback). ISSN 1431-8784. xvi + 363 pp. LCCN QA276.45.R3 D35 2008.

DasGupta:2008:ATS

[4701] Anirban DasGupta. Asymptotic Theory of Statistics and Probability. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-75970-0, 0-387-75971-9 (ebook). xxvii + 722 pp. LCCN QA276 .D3254 2008.

deJong:2008:GLM

[4702] Piet de Jong and Gillian Z. Heller.

Generalized Linear Models for Insurance Data. International series on actuarial science. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-87914-0 (hardcover). x + 196 pp. LCCN HG8782 .J66 2008. URL http:

//www.loc.gov/catdir/toc/fy0805/
2008274156.htm.

deLeeuw:2008:HMA

[4703] Jan de Leeuw and Erik Meijer, editors. Handbook of Multilevel Analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-73183-0, 0-387-73186-5 (e-book). xiii + 493 pp. LCCN H61.25 .H357 2008. URL http:

                                        //www.loc.gov/catdir/toc/fy0804/
                                        2007938348.htm.

Dodge:2008:CES

[4704] Yadolah Dodge. The Concise Encyclopedia of Statistics. Springer reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-31742-2, 0-387-32833-5 (e-book), 0-387-33828-4 (print and electronic bundle). viii + 616 pp. LCCN HA17 .D63 2008.

Hand:2008:SVS

[4705] David J. Hand. Statistics: a Very Short Introduction. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-923356-X. 124 pp. LCCN ????

Marasinghe:2008:SDA

[4706] Mervyn G. Marasinghe and William J. (William Jo) Kennedy. SAS for Data Analysis: Intermediate Statistical Methods. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-77371-1 (hardcover). xii + 557 pp. LCCN QA276.4 .M266 2008.

Spector:2008:DMR

[4707] Phil Spector. Data Manipulation with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008.

ISBN 0-387-74730-3 (paperback). ix + 152 pp. LCCN QA276.45.R3

S64 2008. URL http://www.loc.
gov/catdir/enhancements/fy0904/
2008921862-d.htm;   http://www.
loc.gov/catdir/enhancements/fy0904/
2008921862-t.htm.

Ugarte:2008:PSR

[4708] María Dolores Ugarte, Ana F. Militino, and Alan T. Arnholt. Probability and Statistics with R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2008. ISBN 1-58488-891-1. xxvi + 700 pp. LCCN QA273.19.E4 U53 2008.

Weinberg:2008:SUS

[4709] Sharon L. Weinberg and Sarah Knapp Abramowitz. Statistics Using SPSS: an Integrative Approach. Cambridge University Press, Cambridge, UK, second edition, 2008. ISBN 0-521-89922-2 (hardcover), 0-521-67637-1 (paperback). xv + 764 pp. LCCN QA276

.W4423 2008. URL http://www.loc.
gov/catdir/enhancements/fy0803/
2007031423-b.htm;   http://www.
loc.gov/catdir/enhancements/fy0803/
2007031423-d.htm;   http://www.
loc.gov/catdir/enhancements/fy0803/
2007031423-t.htm.

Agung:2009:TSD

[4710] Ign Agung. Time Series Data Analysis using EViews. Wiley, New York, NY, USA, 2009. ISBN 0-470-82367-4 (hardcover). ???? pp. LCCN QA280 .A265 2009. URL http://

www.loc.gov/catdir/toc/ecip0825/
2008035077.htm.

Albert:2009:BCR

[4711] Jim Albert. Bayesian computation with R. Use R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 0-387-92297-0. 300 (est.) pp. LCCN ????

Broughton:2009:DFA

[4712] S. Allen Broughton and Kurt Bryan. Discrete Fourier Analysis and Wavelets: Applications to Signal and Image Processing. Wiley, New York, NY, USA, 2009. ISBN 0-470-29466-3 (hardcover). xv + 337 pp. LCCN QA403.5 .B77

                                        2009. URL http://www.loc.gov/
                                        catdir/toc/ecip0813/2008011012.
                                        htm.

Cowpertwait:2009:ITS

[4713] Paul S. P. Cowpertwait and Andrew V. Metcalfe. Introductory Time Series with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-88697-4. xv + 254 pp. LCCN QA280

                                        .C69 2009. URL http://www.loc.
                                        gov/catdir/enhancements/fy1002/
                                        2009928496-d.htm;   http://www.
                                        loc.gov/catdir/enhancements/fy1002/
                                        2009928496-t.htm.

Hulsizer:2009:GTS

[4714] Michael R. Hulsizer and Linda M. Woolf. A Guide to Teaching Statistics: Innovations and Best Practices, volume 3 of Teaching psychological science. Wiley-Blackwell Publishers, Malden, MA, USA, 2009. ISBN 1-4051-5573-6 (hardcover), 1-4051-5574-4 (paperback). xix + 253 pp. LCCN QA276.18 .H86 2009. URL http://

www.loc.gov/catdir/toc/ecip0813/
2008010968.html.

Klemens:2009:MDT

[4715] Ben Klemens. Modeling with Data:

Tools and Techniques for Scientific Computing. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13314-X (hardcover). xiii + 454 pp. LCCN QA276 .K546 2009.

URL http://www.loc.gov/catdir/
toc/ecip0821/2008028341.htm.

Mounfield:2009:SCM

[4716] Craig Mounfield. Synthetic CDOs:

Modelling, Valuation and Risk Management. Mathematics, finance and risk. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-89788-2. xv + 369 pp. LCCN HG6024.A3 M69 2009.

Petris:2009:DLM

[4717] Giovanni Petris, Sonia Petrone, and Patrizia Campagnoli. Dynamic Linear Models with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-77237-5, 0-387-77238-3. xiii + 251 pp. LCCN QA276.45.R3 P48 2009. URL http://

www.loc.gov/catdir/enhancements/
fy1108/2009926480-d.htm.

Ryan:2009:MRM

[4718] Thomas P. Ryan. Modern Regression Methods. Wiley series in probability and statistics. Wiley, New York, NY, USA, second edition, 2009. ISBN 0-470-08186-4 (hardcover). xix + 642 pp. LCCN QA278.2 .R93 2009.

URL http://www.loc.gov/catdir/
toc/ecip0825/2008035085.htm.

Stewart:2009:PMC

[4719] William J. Stewart. Probability, Markov Chains, Queues, and Simulation: the Mathematical Basis of Performance Modeling. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14062-6 (hardcover). xviii + 758 pp. LCCN QA273 .S7532 2009.

                                        URL http://www.loc.gov/catdir/
                                        toc/ecip0828/2008041122.htm.

Wainer:2009:PUW

[4720] Howard Wainer. Picturing the Uncertain World: How to Understand, Communicate, and Control Uncertainty Through Graphical Display. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13759-5 (hardcover). xviii + 244 + 16 pp. LCCN Q375 .W35 2009.

Engle:2010:VTS

[4721] R. F. (Robert F.) Engle, Mark W. Watson, Tim Bollerslev, and Jeffrey R. Russell, editors. Volatility and Time Series Econometrics: Essays In Honor of Robert F. Engle. Advanced texts in econometrics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-954949-4 (hardcover). xi + 419 pp. LCCN HB139 .V65 2010.

Prado:2010:TSM

[4722] Raquel Prado and Mike West. Time Series: Modeling, Computation, and Inference. Chapman and Hall/CRC texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-9336-3. xx + 353 pp. LCCN QA280 .P723 2010.

Balan:2011:EUF

[4723] Raluca Balan and Gilles Lamothe. Expect the Unexpected: a First Course in Biostatistics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4291-32-3. ix + 242 pp. LCCN QH323.5 .B353 2011.

Kupper:2011:ESB

[4724] Lawrence L. Kupper, Sean M. O’Brien, and Brian H. Neelon. Exercises and Solutions in Biostatistical Theory. Chapman and Hall/CRC texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-58488-722-2 (paperback). xvii + 402 pp. LCCN QH323.5 .K87 2011.

Lipton:2011:OHC

[4725] Alexander Lipton and Andrew Rennie, editors. The Oxford Handbook of Credit Derivatives. Oxford handbooks in finance. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-954678-9 (hardcover). xxvi + 677 pp. LCCN HG6024.A3 O94 2011.

Aladag:2012:ATS

[4726] Cagdas Hakan Aladag and Erol Eğrioğlu, editors. Advances in Time Series Forecasting. Bentham eBooks, Oak Park, IL, USA, 2012. ISBN 1-60805-522-1, 1-60805-373-3 (e-book). iii + 135 pp. LCCN HF5500.A1 .H384 2012.

Bell:2012:ETS

[4727] William R. Bell, Scott H. Holan, and Tucker McElroy, editors. Economic Time Series: Modeling and Seasonality. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-4657-X. xvii + 535 pp. LCCN HB141 .E255 2012.

Chaumont:2012:EPG

[4728] Loïc Chaumont and Marc Yor. Exercises in Probability: a Guided Tour from Measure Theory to Random Processes, via Conditioning. Cambridge Series in Statistical and Probabilistic Mathematics. Cambridge University Press, Cambridge, UK, second edition, 2012. ISBN 1-107-60655-1 (Paperback). xx + 279 pp. LCCN ????

Crawley:2012:RB

[4729] Michael J. Crawley. The R Book. Wiley, New York, NY, USA, second edition, 2012. ISBN 0-470-97392-7 (hardback). xxiv + 1051 pp. LCCN QA276.45.R3 C73 2013.

Crowley:2012:HSC

[4730] John Crowley and Antje Hoering, editors. Handbook of Statistics in Clinical Oncology. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2012. ISBN 1-4398-6200-1 (hardcover). xix + 635 pp. LCCN RC267 .H36

2012. URL http://catalogue.bnf.
fr/ark:/12148/cb42648647v.

Fenton:2012:RAD

[4731] Norman E. Fenton and Martin (Martin D.) Neil. Risk Assessment and Decision Analysis with Bayesian Networks. Taylor and Francis, Boca Raton, FL, USA, 2012. ISBN 1-4398-0910-0 (hardcover). xix + 503 pp. LCCN QA279.5 .F46 2012.

Gray:2012:RMG

[4732] Roger J. Gray and Susan M. Pitts.

Risk Modelling in General Insurance: from Principles to Practice. International series on actuarial science. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-139-03375-1 (ebook), 0-521-86394-5 (hardcover). xiv + 393 pp. LCCN HG8054.5 .G735

2012eb. URL http://site.ebrary.
com/id/10578228.

Kumar:2012:SMC

[4733] V. Kumar and J. Andrew Petersen.

Statistical Methods in Customer Relationship Management. Wiley, New York, NY, USA, 2012. ISBN 1-119-99320-2 (hardcover), 1-118-34920-2 (ebook). xiii + 272 pp. LCCN HF5415.5 .K864 2012.

Lee:2012:BSI

[4734] Peter M. Lee. Bayesian Statistics: an Introduction. Wiley, New York, NY, USA, fourth edition, 2012. ISBN 1-118-33257-1 (paperback). xxiii + 462 pp. LCCN QA279.5 .L44 2012.

McCool:2012:UWD

[4735] John McCool. Using the Weibull Distribution: Reliability, Modeling, and Inference. Wiley series in probability and statistics. Wiley, New York, NY, USA, 2012. ISBN 1-118-21798-5. xix + 336 pp. LCCN QA273 .6 M129u 2012.

Pruessner:2012:SOC

[4736] Gunnar Pruessner. Self-organised Criticality: Theory, Models, And Characterisation. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-139-52799-1, 0-521-85335-4, 1-139-52560-3 (e-book), 0-511-97767-0 (ebook). LCCN QC174.85.S34 P78 2011.

Schoenberg:2012:IPT

[4737] Frederic Paik Schoenberg. Introduction to Probability with Texas Hold’em Examples. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-2768-0 (paperback). x + 189 pp. LCCN QA273 .S377 2012.

Tang:2012:ACC

[4738] Wan. Tang, Hua. He, and Xin M. Tu. Applied Categorial and Count Data Analysis. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-0624-1. xx + 363 pp. LCCN ????

Agresti:2013:CDA

[4739] Alan Agresti. Categorical Data Analysis. Wiley series in probability and statistics. Wiley, New York, NY, USA, third edition, 2013. ISBN 0-470-46363-5 (hardcover), 1-118-71094-0 (e-book). xvi + 714 pp. LCCN QA278 .A353 2013.

Givens:2013:CS

[4740] Geof H. Givens and Jennifer A. (Jennifer Ann) Hoeting. Computational Statistics. Wiley series in computational statistics. Wiley, New York, NY, USA, second edition, 2013. ISBN 0-470-53331-5 (hardcover). xviii + 469 pp. LCCN QA276.4 .G58 2013.

Indrayan:2013:MB

[4741] Abhaya Indrayan. Medical Biostatistics. Chapman and Hall/CRC biostatistics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2013. ISBN 1-4398-8414-5 (hardcover). lxi + 962 pp. LCCN ????

Jureckova:2013:MRN

[4742] Jana Jurečková, Pranab Kumar Sen, and Jan Picek. Methodology in Robust and Nonparametric Statistics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4398-4068-7 (hardcover). xv + 394 pp. LCCN QA276 .J8675 2013.

Mateu:2013:STD

[4743] Jorge Mateu and W. G. (Werner G.)

Müller, editors. Spatio-Temporal Design: Advances in Efficient Data Acquisition. Statistics in practice. Wiley, New York, NY, USA, 2013. ISBN 0-470-97429-X. xx + 349 pp. LCCN QA276.6 .S63 2013.

URL http://www.loc.gov/catdir/
enhancements/fy1307/2012027161-
d.html;   http://www.loc.gov/catdir/
enhancements/fy1307/2012027161-
t.html.

Pewsey:2013:CSR

[4744] Arthur Pewsey, Markus Neuhäuser, and Graeme D. Ruxton. Circular Statistics in R. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-967113-3. xiv + 183 pp. LCCN QA276

.P45 2013. URL http://www.loc.
gov/catdir/enhancements/fy1410/
2013940576-b.html;   http://www.
loc.gov/catdir/enhancements/fy1410/
2013940576-d.html;   http://www.
loc.gov/catdir/enhancements/fy1410/
2013940576-t.html.

Sclove:2013:CSF

[4745] Stanley L. Sclove. A Course on Statistics for Finance. CRC Press/Taylor and Francis Group, Boca Raton, FL, USA, 2013. ISBN 1-4398-9254-7 (hardcover). xxvii + 251 pp. LCCN ????

Tsay:2013:IAF

[4746] Ruey S. Tsay. An Introduction to Analysis of Financial Data with R. Wiley, New York, NY, USA, 2013. ISBN 0-470-89081-9 (hardcover). xiv + 390 pp. LCCN HG106 .T76 2013.

Westfall:2013:UAS

[4747] Peter H. Westfall and Kevin S. S. Henning. Understanding Advanced Statistical Methods. Texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4665-1210-5 (hardcover). xxv + 543 pp. LCCN QA276 .W4537 2013.

Chang:2014:PSM

[4748] Mark Chang. Principles of Scientific Methods. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4822-3810-1, 1-4822-3809-8. ix + 237 pp. LCCN Q175 .C456 2014.

Ho:2014:HUM

[4749] Robert Ho. Handbook of Univariate and Multivariate Data Analysis with IBM SPSS. Taylor and Francis, Boca Raton, FL, USA, second edition, 2014. ISBN 1-4398-9021-8 (hardcover). xxiii + 561 pp. LCCN QA279 .H6 2014.